A Complete Bibliography of Publications in Concurrency and Computation: Practice and

Experience: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

28 May 2024 Version 2.03

Title word cross-reference

(1, |N_{i}|) [KK21a]. (k, m, t) [PKS22]. (n, k) [GQ24]. (σ, τ) [KKM21]. 2 [QMC^{+}20, RJH^{+}23]. 3 [AJNS22, CSL21, FAM22, GBB22, HSL^{+}22, LHWT20, TLH^{+}22, YYM^{+}24, ZQW^{+}21, ZZW^{+}22, ZKZ^{+}23, ZX23b]. ^{2} [Faz23, LLR^{+}21, RJ22]. ^{3} [CdRNB23]. _{2} [xZlGCzJ20]. _{i} [RN22]. c [KSS21]. ℓ [ST23b]. γ [YMZD21]. GF(2^{8}) [SV22d]. K [KUK22, LS23b, NJ21, PS22b, RRJ23, RP21c, VR22, AK22c, Ben21, CT21, ESS23, FYH^{+}21, FM22, GZY^{+}22, HVB22, LX22, WQY^{+}22, WGLL23]. N [FHH^{+}20, PJKO24]. n/2 [XHST20]. P [LML^{+}23]. P (X > Y) [Çet23]. π [VPQ22]. Q [SKR^{+}24, KB22]. t [WSL^{+}20].

* [BWS^{+}21]. *-Predictable [BWS^{+}21].

-anonymity [PKS22, WGLL23]. -ary [ST23b]. -based [SMBM24, SS21, KSS21]. -calculus [VPQ22]. -critical [KK21a]. -cubes [FHH^{+}20]. -degree [LX22]. -directed [KKM21]. -distributed [WSL^{+}20]. -frame [KB22]. -learning [SKR^{+}24]. -map [YMZD21]. -means [CT21, ESS23, FM22, HVB22, KUK22, LS23b, NJ21, PS22b, RP21c, VR22]. -nearest [Ben21, FYH^{+}21, GZY^{+}22, WQY^{+}22, RRJ23]. -Queens [PJKO24]. -slice [RN22]. -star [GQ24]. -wave [LML^{+}23].

/manycore [HOS^{+}21].

1 [SPHP21]. 16th [LM20a]. 1788 [BFR24]. 1788-2015 [BFR24]. 18 [SKÇA23]. 19 [ATS22, ARS22b, Asl22, Boz22b, Cek22, CPYC21, CBK23, DÖD22, DR23, DSSS22, EUYY22, GP22b, JD22, Kab22, KKKS23, LLM^{+}24, MRGP22, MSA22, ÖUG22, PS22c, PKVS21, SJA^{+}22, SVS22, SAAAA22, SP22b, UYÖ^{+}22, UAS^{+}22, ZARR23, ZSS22]. 19th [Bad23].

2 [NMT^{+}23]. 2015 [BFR24]. 2019-2020 [SCH22a]. 2021 [SKE22]. 2022 [KSY23]. 20th [IS23]. 27th [LM20a].

3000 [VBP^{+}24]. 3D [PBD23]. 3D-AB [PBD23]. 3DCNN [WLZ^{+}21b]. 3DCNN-ConvLSTM [WLZ^{+}21b]. 3Dconvolutional [TFZC23]. 3DES [APTT22].

4G [HYK21].

5G [BJK^{+}24, BTDJ21, CAS24, CÇ22c, DL23, DJF21, KKS22, PCG^{+}21, SKPV22, SCS^{+}21, WLCW20, rSN21]. 5G-IoT [CAS24].

6 [TYL22]. 6-DOF [TYL22]. 60GHz [WXCY20].

8 [CLWX21]. 802.21 [XLXZ20].

9 [YKW20].

A-level [ZK24]. A-PGRD [ZLX^{+}24]. A2Cloud [SKB^{+}20]. A2Cloud-RF [SKB^{+}20]. A64FX [AML^{+}22]. AAEE [LZSC23]. AAEE-Net [LZSC23]. ability [HLC23, LGW^{+}22a]. Abnormal [SG22c]. abnormalities [OE22a, OE22b]. abnormality [KGM23, YRV^{+}23]. absence [Shi22b]. absolute [CC22b]. Abstract [MdAA^{+}21, HL23]. abstracted [PYC^{+}20]. abstraction [IGYS23, dCJAAdOD21, PMC^{+}21, RF21]. abstractions [WPK^{+}22, ZYPD24]. Abstractive [EA22, CKL20]. AC [TPT^{+}22]. AC-based [TPT^{+}22]. academic [RT21, RT22a]. accelerate [SZZ^{+}24a, STH^{+}20]. accelerated [APTT22, DE22, HIN23, Liu24b, LYI^{+}20, SKSP20, ZZL^{+}22, DS22b, MM21, PK22b]. Accelerating [BABS21, MRK^{+}23, SMCM^{+}22, TRM^{+}24, VCFZ20, WMC^{+}23, WSM^{+}24, YKW20, GCF^{+}20, PRS23, PVP^{+}20, QMC^{+}20, RGM22, ZDL^{+}22].

Acceleration [VMFL23, JPL22, SSN21]. Accelerator [RSV24, ALB^{+}20, BCP^{+}23, GS21, ZYH^{+}23]. accelerator-based [ALB^{+}20]. accelerators [GGA24, NMS^{+}22, ÖGS22, PBC^{+}23, PSF23, RZVC21, ZSC^{+}21]. acceptance [VBP^{+}24]. Access [LGW^{+}22b, AA23a, Agr21, AAD20, BA24, CRB23, CLC^{+}20, DZCL22, EI22, HAAF22, dCJBP20, KM23a, KW21, KS24, LWL^{+}23, LYI^{+}20, NA22a, NET21b, NKST24, PS22c, PP20a, RSMCP22, SS23e, SLJ23, XSGL20, XZW^{+}20, XZYW23, YSS^{+}21, YF20, ZAB22]. access-control [dCJBP20, PS22c]. accesses [KHPH20]. accessibility [PEGP23, SA22c]. accessing [KKM20]. accident [GKM22, WHH^{+}23]. accidents [DWZ^{+}20b, DWY^{+}21]. according [AAA22b, YY22, gZWfL^{+}20]. accountability [PI21]. accounting [NSKS21]. accounts [ZWL^{+}20]. accuracy [Aka22b, AAB23, DS23b, KGK22, KMD23, OCD22]. accurate [AMBAJ22, AKUA22, CXX^{+}22, DSJ22, DBD23, HVB22, HXST22, KK21b, LS23b, LZSC23, LLC^{+}22b, MLVGR24, NBS^{+}22, RK23a, SZL^{+}22, VK23b, XYLW21, ZZW^{+}22]. achieving [YYM^{+}24]. acid [ABB22, RK23b]. acid-Schiff-stained [RK23b]. Acknowledgment [SGS^{+}23]. ACO [KS20]. ACO-based [KS20]. acoustic [GA23, Kum22, SR20a]. Acquisition [YGS^{+}23, HKP21, LP21, VPB^{+}23, WZMJ23a, ZX21b]. acronym [RV21]. acronym-definitions [RV21]. across [ASJ21, BWTJ20, HVT23, PSF23, Sin23, ZWL^{+}20]. action [CLX^{+}21, KDC22, RR23a, ZGG^{+}23]. actions [SP21a]. activate [KCP^{+}22]. activation [LGT^{+}23, OKJ^{+}21]. activations [MNR^{+}22]. Active [WJL^{+}20, LW20b, LHX^{+}23, RKL21]. activities [ASJ24, CZTC22, GA22, KSK^{+}22, KR22b]. activity [AR22a, AFM22, BYTG22, GXL^{+}20, KSJK21, KSK23, QZB^{+}23, TSD23, ZLD22]. actor [ECIB20]. acute [BP23a, JSAA22]. acyclic [KKM21, SSV24]. Ad [AMRH21, AK21, BKS22, CD22, DCR23, FLB23, KGGM22, LCW23b, NAR^{+}22, Ngu21, PDRM24, PA21, PD23, RGKK21, RSS20, RAN22, SK22a, SPS23, SWK22, SSH22, TT22, TK22b, WJD22, VPSM22]. ad-hoc [AK21, CD22, NAR^{+}22, PDRM24, RGKK21, TK22b]. Ada [SS22c]. Ada-Boost [SS22c]. AdaBoost [DBD22, Hem22, SS22b]. Adam [DM22d, VRS22]. ADAPT [WBZ21]. adaptability [RK21b]. adaptable [FPXM21, LGM21]. adaptation [ALR22, FKK23, RS21, VGDF22, WFHC21]. Adapted [SB21, CON23]. adapting [JT23]. adaption [ZZZ^{+}20]. Adaptive [AI22, CLYG22, Cud20, FC21, JM23, JKP22, LYF^{+}23, LZQF24, NTK22, NSKS21, RRB22, RM23, RKC22, SA22a, SLG^{+}20, SCH^{+}24, SP22a, URK^{+}22, WBZ21, XL21, YRV^{+}23, ZLZ^{+}22a, ACJ21a, AFKS23, BSN22, BL22, BN21, CK24a, CSWZ22, DRR22, DJ20, ESS23, FCMM20, GWA^{+}24, HPS23, ITO21, IAQ20, KTM22, KS22a, KV22b, LST22, LXW^{+}23, MS21c, NLG^{+}20, PSMM22, SWCB20, SRGK24, SHA^{+}22, SKP22, SMBM24, SHBC20, SZS20, SP22b, SS23f, SSI22, SKRS21, SRG^{+}21, TA23a, TS22, WK20, XCY22, XLL^{+}23, YLZW23, YZ21, ZTMC22, ZTMC23, ZYY^{+}24, uZKH^{+}20]. adder [GM22a, MKRK23, SPSP23, SN22a]. adder-subtractor [GM22a]. adding [VAB22]. addition [GWA^{+}23]. address [NY22, ZSC^{+}21]. addressed [KDFK23]. addresses [JAM24]. addressing [DCR23]. ADER [DGB24]. ADER-DG [DGB24]. adhoc [ST22b]. adjacency [TP20]. Adjoin [TP20]. adjustable [LNI^{+}24]. adjusted [AAK21c]. adjusting [FKO22]. adjustment [WJJ23]. Adleman [DP22b]. Admission [EABZB21, JR24, SV22c, SCS^{+}21]. adoption [BC21, KLJ21, KVP21, OMA^{+}23]. adulteration [MCT22a]. Advance [AAN^{+}21]. Advanced [JT23, Ogi20b, SM24a, AE22b, AFBM^{+}23, CZTC22, SSS23a, SZ24, VP22a, Yil21b, ZABT^{+}20, LF23]. advancements [BS22, WS21]. Advances [TLS22, WS24, YIB22b, ABZS20, BP23b, FZT22, MLKD20, WC22]. Advancing [BFM^{+}23]. advantages [ATS22]. adverbs [HAA^{+}21]. adversarial [AWS^{+}22, AGM23, CWW^{+}23b, DSSS22, DCWM20, Eke22, JAM24, KH22, KSJ23a, LG23, LSD21, LLL^{+}22, LWHW22, NGD^{+}22, PKK23, PSK23, ST23a, SYT^{+}21, VKSS23, WZSZ20, XLY^{+}23, ZDH^{+}22, ZQW^{+}21, ZLT21, ZCZ^{+}22]. adversaries [HZQ^{+}24]. adversary [XHST20]. advertising [SWZW20, ZR24]. advising [SZY^{+}22]. advisor [LMBK23]. advisors [PMS^{+}21]. AE [RZ24]. AE-Integrated [RZ24]. aerial [BEKS22, ETKD23, FJCFB23, JYH^{+}24, KGK22, KS22b, LLX^{+}21, LGL^{+}22, LMM^{+}22, PLX^{+}24, XPLX23]. aeroacoustic [FZ23]. AES [MS21a, CL22a]. AES-RC4 [MS21a]. affected [DSJ22, RK23a]. affinity [ALBZ21, KRJS22]. AFMC [dCMA23]. Africa [Ano21-42]. African [KMR22]. AG [Pan24]. against [AA23c, HZQ^{+}24, YYZS22, ZTP^{+}23]. AGCNN [ZLZ^{+}22a]. AGE [PLZL24]. AGE-MOEA [PLZL24]. Agent [ZWCS20, ALR22, DCZ^{+}22, EI22, FGJ^{+}21, GR22b, HBB20, KBR24, QPS20, RS23, SMD22b, XAC^{+}20, ZTL^{+}21, ZHWY22]. Agent-based [ZWCS20, ALR22, EI22, HBB20, KBR24, QPS20, XAC^{+}20]. agents [USK23, HNOI24, LGX^{+}23, RF21]. AGFT [NTK22]. agglomerative [SZZ^{+}21]. aggregate [PM23a]. aggregation [ASRN23, BZT^{+}23a, CLC^{+}23, JS22a, LGZ^{+}22, LZSC23, NTK22, SS23d, SZG^{+}24, YLZT23, YLGY20, YLW^{+}21, ZYZC23]. aggregative [CT22]. aggregator [SSRA23]. aggregators [ZTP^{+}23]. agile [GM22c]. agility [DDB^{+}21]. AGINFRA [ABC^{+}21]. aging [NGD^{+}22]. agreement [BZK^{+}21, CBK23, FCX^{+}22, KK23a, NAR^{+}22, NR22, PDRM24, VB22a]. agreements [SB21]. agri [ABC^{+}21, BK22a]. agri-food [ABC^{+}21]. agri-product [BK22a]. agricultural [BP23b, Liu21, SHA^{+}22, VDL23]. agriculture [LP21, MTSU22, YLW^{+}21]. agroecological [SP22d]. AGSO [SMBM24]. AI [GKLS23, LMBK23, MTY21, OAS23, YFdP^{+}23]. AI-centric [YFdP^{+}23]. AI-driven [MTY21]. aid [TLX22, ZFH^{+}23]. aided [CYDW20, NKST24, SMKA22, XZL^{+}20a]. AIoT [WZMJ23a, LCCT22]. air [BTDD20, KA21b, LZL^{+}20b]. airborne [ML20]. alarm [UE22]. Alder [BSPM23]. AlexNet [ATS22, MMR^{+}22, PS23b, SSMT22b, GM22b].

AlexNet-DrpXLm [MMR^{+}22]. AlexNet-extreme [PS23b]. alfalfa [JA23]. algebra [GCF^{+}20, HV21]. Algorithm [HH23, PI21, WWG21, AB22b, AB20a, AZM20, ABESEh20, AFF22, AMAT22, AE22b, AA21, ATS22, AK22c, AFBM^{+}23, ASL20, AV23, AEMS24, AS23b, ABA22, ASC22, Ano21-38, AB21, AMA22, AFNH21, AO23, BM22a, BMcKGK22, BBF21, BDK22, BP24, Ben21, Ben22, BRS^{+}22, BVP22, BYL20, Bul22, CGW^{+}20, CEN22, CDC20, CYQ^{+}20, CWL^{+}21, CWL^{+}22, CBFS23, Cho20, CG22, CÇM22b, CZ21, CZL^{+}22, DCM21, DMS^{+}21, DÖD22, DBD22, DM22b, DS21, DBPC22, DZCL22, DS22a, DR23, DP22b, DG21, DR20, DWZ20a, DSSS22, DWZ^{+}20b, DCWM20, DQF^{+}23, DLC^{+}21, DD24, DJGF21, DPSJ22, FSA^{+}22, FLG^{+}22, FGZC23, FZZZ23, GMM22, GKÇ22, GFA21, GDFDF22, GR22b, GBS24, GSG^{+}23, GK23a, GVSS22, Gui22, Gün23b, GAM23, HYG20, HCH^{+}21, HVB22, HHC^{+}22, HYC^{+}24, HK22b, HAK22, HGMK21, HXY^{+}22b, HHP23, Hua20, HFH^{+}21, HWY^{+}23, IAA20]. algorithm [ITO21, IAAA22, JT23, JMV24, JNM^{+}22, JSA^{+}20, JPL22, JM22, JZC^{+}23, JYH^{+}24, KUK22, KRKM22, KK24, KK21a, KV22a, KKKS23, KÖ22, KA21a, KS22a, KCP^{+}22, Kha22a, KFML20, KEMZ22, KJHM21, KKM21, KAO22, KWL^{+}24, KK22e, KT22b, KK22f, KS22d, KDS^{+}23, KAAR23, KA22c, LHJ22, LK22a, LJBS23, LSS^{+}21, LSSQ22, LS20, LZL^{+}20a, LCZ^{+}20a, LFWJ22, LYC22, LLW^{+}23, LLTZ24, Lia24, LJP^{+}21, LZC21, LSX21, LZY^{+}21, LXC^{+}22a, Liu24a, LAC21, LZQ^{+}22, LYG^{+}21, LHL^{+}20, LZT^{+}23, MSJ22, MNDK22, MTK^{+}21, MKA23, MS21c, MSK22, MB22, MMS24, MT22, MC20, MHB^{+}22b, MAB22, MdARS^{+}23, MSL22, NM23, NKKM21, NLB22, NSBT21, NAK^{+}22, NJ21, NMS^{+}21, ODK^{+}23, PBK23, Pal22, Pan23, PBD23, PZD^{+}21, PL21, PD22b, PB22b, PKR22, PSK23, PS23b, AST22, QCC^{+}23, RJH^{+}23, RR23a, RP21a, RPPK23, RPMA22, RCK22, Rav23, RP22, RZCA21]. algorithm [RSJ22, SAPC21, SEMA^{+}22, ST22a, SY23, SA22a, SKA23, SSMT22a, ST22b, SAC22, SRG^{+}22, SRGK24, SS22a, SEC22, SV22b, SVD22a, SMR21, SEM^{+}20, SM22b, SCP20, SZZ^{+}24a, SKHL22, SS22c, SZqWZ20, SP22a, Shi22b, SP22b, Shu22, SS22d, SS23f, SMKA22, SP22c, SR22c, SXC^{+}23, wSYyCsD23, SD23, SKSP23, SAMS23, SCL^{+}20, SBS22, TA23a, TTM^{+}22, TJZ23, TWW^{+}21, TAB21, TÖK21, TB23, TLX22, TO22, TK24, USI21, VKSS23, VP22a, VG21, WRJ20, Wan20, WLJ20, WZHL21, WSL21, WQY^{+}22, WWA22, WZC^{+}22a, WFY22b, WJJ23, WQS^{+}23, Wan24, WLLZ20, WN21, WZX^{+}22, WN23, WCWG21, WYLL24, XCY22, yXlLyGX21, XLZ20, XGZ^{+}20, XL21, XWW^{+}21, XLZL22, YZYT21, YLLL20, YCL^{+}22a, YLZC23, Yu21, Yüc22, ZTM21, ZX20, ZSX21, ZGLS21, ZLTX21, ZYCS22, ZWCC23, ZGWZ23, yZyWD^{+}21, ZZLZ21, ZCL21, ZZLZ22b, ZHW^{+}20, Zhu21, ZZS^{+}21b, Zhu22b, ZGH^{+}22, ZCW^{+}23, LST22, LSL20, Yil21c, ZFH^{+}23]. algorithm-based [AFF22, MKA23, ODK^{+}23, SRGK24, VG21]. algorithm-driven [SS22d]. Algorithm-II-based [Yil21c]. algorithm-least [XCY22]. algorithm-random [MHB^{+}22b]. algorithm/hardware [SZZ^{+}24a].

Algorithmic [WC22, KTU^{+}21, SP23a, WS21]. algorithmics [DJF21]. Algorithms [AKLK24, HK21, KLL^{+}21, LM20a, AAMAA22, ABZS20, AK21, AIA22, AB22a, AFG^{+}22, AR22b, Bad23, BLI20, BEKS22, BS23b, BF22, BKGC23, CZUPA23, CCCR21, CHI^{+}23, CBR22, CK23, ÇGB23, DS23b, EUYY22, FSdP^{+}23, FAM22, GDSS22, GKM22, GMN21, GKZ23, HIEH22, HZD^{+}22, HC22, HYT^{+}21, IS23, JCGT24, JXL^{+}23, KB22, KNM21, KK21b, KC22, Kay22b, KML23, KGGM22, KY22, KA21b, KA22b, LLW^{+}20, LYBZ23, LHPG21, MQEK21, MS22b, MLZ21b, NGB23, OO21, PR24, PVP^{+}20, PKM21, RTSK23, RAK22, SSS^{+}20, SV21, SYG22, SKSN22, TSR22, TCW^{+}23, TT21, WXLD21, WSL^{+}20, YGZ^{+}21, xZlGCzJ20, ZX21a]. aligning [ZWL^{+}20]. alignment [EASR22, GCQ^{+}24, LLM^{+}24, PQKDT21, SY23, STH^{+}20, dCMA23]. all-path [XZL^{+}20b]. alleviate [GK23b]. alleviating [YBJ^{+}23]. alliance [WMQ^{+}22]. allocating [YHL^{+}23]. Allocation [CJC22, ALBZ21, AKS^{+}22, BP24, CWW23a, DR21, DSC^{+}21a, FPXM21, FMNF22, GSZ^{+}20, GSG^{+}23, GMSM21, HZW^{+}23, JW21, JZC^{+}23, JR23a, KRKM22, KS22a, KKS22, KS20, KHR22, KDS^{+}23, LTLX22, LB21, Mir22, MTRA23, MV24, NKST24, PK23b, RR22, RZCA21, Sao22, SKPV22, SZJ21, SSPG20, TS21, VSS23, VB22b, WLCW20, WGY20, WJD22, YS23, Zha22, dSNdL^{+}23]. allocations [CYK^{+}21, CÇM22b]. AllReduce [NWT21]. Almost [LLO21, LFW20]. ALO [TK24]. alphabet [LAE^{+}22]. AltaRica [WXH^{+}22, WSL21]. alternating [BZT^{+}23a]. alternative [BC21]. alternatives [AÜ23, RS24]. Alzheimer [ABCP23]. amalgamated [WWF^{+}22]. amalgamation [SD23]. Amazon [TMB24]. ambient [WHDS22]. AMD [AT22, HMR24, SCKU23, XWD^{+}22]. AMD-based [HMR24]. AMD-CNN [AT22]. American [SLL^{+}23]. among [AA23a, KSJ23b, RDB22, YLW^{+}22]. amorphous [TK24]. amount [LYI^{+}20]. Amplified [CYW^{+}22]. amplifier [SA22b]. amplitude [YYZS22]. amplitude-damping [YYZS22]. analogy [SBB21]. analogy-based [SBB21]. analyse [ÖTT23]. analyses [DAdBM^{+}24]. Analysis [Ano21a, BS22, DM22c, Jia22, LWZY23, SHZY20, YYLL22, YRSO23, YF20, ZMLW23, AC22, AKEC20, AMRH21, AAK^{+}21a, AIA22, AOACAQ21, AS22c, APM^{+}21, AJS23, Ano21e, AS24, AHV21, Aru22, AS23c, Bai23, BMK^{+}20, BK21, BC23b, BABLH21, BDG^{+}23, BBP^{+}24, BMJ^{+}20, CD24a, CZG^{+}20, CK21b, CPYC21, CMLL21, DCM21, DDZ^{+}20, DBD22, DR23, DSS^{+}23, DE20, EHST21, EI22, GDA22, wGTC22, GXH^{+}21, GN21b, GK23c, HAA^{+}21, Haj20, HZD^{+}22, HNG22, HBB20, HGMK21, HKP21, JAR^{+}24, JB21, JZB20, Kab22, KS23a, KS20, KSA22, KPF^{+}20, Kim21a, KK22d, KJ20, Kum21, KPP^{+}22, KDS^{+}23, KS23c, LYL20, LMM23, LZC^{+}20, LJZ21, LLSC22, LDS^{+}23, LHPG21, LWCM21, LLL^{+}22, LDH^{+}22, LSW21, MM24, MS22b, MAW^{+}22, MRKY22, MBT^{+}20, NMS^{+}21, ORP21, OLF21, OS21b, ÖUG22, Ona21, PMC^{+}21, RR23b, RBDS22, RZ21, ST22a, SRG^{+}22, SG21]. analysis [SKK21, SAAAA22, SSK^{+}23a, SZGR21, SSP22, SM21, TTB^{+}22, TI22, TTZX22, TC24, UKS22, UDS21, UYÖ^{+}22, VSS23, VPB^{+}23, VRB21, WY20, WGZ^{+}20, WLZ^{+}21b, WMC21, WSL21, WLG^{+}22, WZX^{+}22, XHZHXBQX22, YB23a, YKW20, YJY^{+}21, YÇC22, YFdP^{+}23, YSC^{+}20, YSL^{+}21b, ZSS22, ZR24, ZLD22, ZZLZ22a, ZWJG21, ZJSJ20, ZJSJ21, ZHT^{+}23, dOdMC^{+}20]. analysis-Challenges [Kum21]. Analytic [AHW22, KUK22, PMS^{+}21, RAN21, RG23, VR22]. analytical [KSB23, XSZ^{+}20]. Analytics [DPS21, FPÁ^{+}20, AJAA21, CJ21b, DA22, DYF20, JAC^{+}21, KKR23, LLJ^{+}20, MG21a, MMKA23, PLP22, RC22, RMS24, RGPC23, YÇC22]. analyze [BEJD22]. Analyzing [ATC23, EUYY22, GCP22, HLCH20, AS22a, KUK^{+}24, YLT^{+}21]. anchor [SA22a]. Andhra [DM22d]. Android [AT22, BD21, KRJS22, LZC^{+}20, SKK21, XYLW21, YLS^{+}23]. ANFIS [JM23, AA23b, GKAO20, KEK^{+}20]. ANFIS-based [AA23b]. angle [LDS^{+}23, LLJ^{+}22b, YZ21, Zhu22b]. angle-based [Zhu22b]. animal [ETKD23, LHJ22]. animals [LLLX20]. animation [WZ24]. animats [QPS20]. Anisotropic [LCZY20, KSVP22]. ANN [JHG23, RK23a]. annealing [KT20, SBB21, SCP20, SB23b]. annealing-dominant [SB23b]. annotation [RCYuRH21, ZJ21]. anomalies [AYKE21, GMSM21]. anomalous [ARS22a, JDG22, PMR^{+}21]. Anomaly [ZHX^{+}23, CR23, GSS23b, ND23, PD20, SKPD23, SZZ^{+}21, TBNPQ23, VRS22, ZX20]. anonymity [LLM^{+}22, PKS22, WGLL23]. anonymization [AM22, CSV22, JB24, LX22]. Anonymous [DLG24, FCX^{+}22, HSR23, PDRM24, VDL23, VB22a]. anopheles [ATS22]. ANOVA [ZR24]. answer [YSK22, ZLM22]. answering [Che24, FDH22, KC20, YSK22]. ant [AD22b, JCL^{+}20, PKK23, RP22, MT21]. antenna [GAM23]. antennae [PB22b]. anti [AB20b, CAAHC23, SS22b]. anti-counterfeiting [AB20b]. anti-jamming [CAAHC23]. anti-phishing [SS22b]. antispoofing [LCW^{+}23a]. AOA [MKA23]. AODV [NM20]. AOMDV [SR22a]. Apache [ASL20, BVM22, DHSG23, HXY20, HTZT23, RZ24]. aperture [GDSS22, MRS^{+}21, SPHP21, XCX^{+}20]. APEX2S [CSWC20]. API [CPQ^{+}22, CJP^{+}21, CM21]. API-to-network [CM21]. app [RM22]. appearance [KCP^{+}22]. applicable [GM22a]. Application [ANP^{+}20, AAA22b, Ano21c, CI20, CMK22, Cud20, GKM22, KAO22, Li21, LPZ^{+}22, ST21, VRB21, ZSX21, AE22b, AS22c, AQP^{+}22, Ano21f, AÜ23, AYJ^{+}22, BA20, BRNR23, Boz22a, Cek22, CRB23, CYC21, CON23, Cob22, CMY21, DBPC22, GKLS23, HMR24, HC22, JQ22, JCL^{+}20, JQGL20, KM21b, KGE^{+}20, LYG^{+}21, MKRK23, MAAA22, NdSSSN20, OBTC20, OWB^{+}20, ÖSTY22, PEGP23, RJH^{+}23, RSM21, RKMS24, SPSP23, SRRM23, ST22c, Shu22, SK22d, SCdLV20, TS22, UK23, WZL^{+}24, WZZ^{+}22, Wan24, XYY24, XY21, XJW^{+}22, XHM22, YJY^{+}21, YLKC24, YK23, Yüc22, ZBZ^{+}20, dOdMC^{+}20]. application-centric [SCdLV20]. application-level [AÜ23]. Applications [AA22b, Ano21-41, HFFA20, JB20, PJ21, SCH22b, SV22d, SKE22, Vin21, XZ20b, AB20a, AZM20, ABZS20, AAN^{+}21, ASJ21, AA23b, ALR22, Ano21b, AYB21, BZEM20, BS23b, BFM^{+}21, BAR21, BKGC23, BPM^{+}22, BWS^{+}21, BPB21, CLLB20, CLE^{+}20, CM20, CL22a, CWFK23, CDP^{+}21, CDN^{+}21, DK21, DAK22, EBLM22, EAS23, FLB23, GBBS21, GDGN23, GMS^{+}21, GA23, GMSM21, GRC^{+}23, GMK^{+}21, GQ21, HA22, HZY^{+}21, HT21, IAT^{+}23, JAC^{+}21, KGK22, KML23, KAP20, KM23b, KM23c, KK22e, LBG^{+}20, LBFT22, LYF22, LXW^{+}23, LZZ^{+}20, Liu24b, LRC^{+}22, LAP23, MTSU22, MVR23, MMR21, MDSG^{+}23, MC24, Ogi20a, OO20, OLF21, PPM^{+}20, RSM23, RK23c, SV22a, SAB22, SPQM20, SMD22b, SCD^{+}23, SMD^{+}21, SLD24, SCH22a, SK20b, SK21b, SKRS21, SRS^{+}21, SZ24, ST22d, TTA20, TLS22, VP22a, ASB23, WCSG20, XYLW21, XZ20a, YSL^{+}21a, YY20a, YIB22b]. applications [Yüc22, ZWT22, dSNdL^{+}23]. Applied [ST21, ASA^{+}21, JC21a, KKE^{+}22, PBD23, SAF^{+}23]. Applying [MPB^{+}22, Kha22a]. approach [AM22, AS22a, AS22b, APM22, AZI20, AD22b, AJ21, ACC^{+}20, ASMS21, ARS22a, AGSN23, AR22b, AMM^{+}20, AYKE21, AHT^{+}20, ASANR22, BD22, BF22, BLT22, BABLH21, BBB^{+}20a, BD21, BJWY20, BSBF22, BMSD23, BN21, CD24a, CF21, CCPP21, CÇY22, CSC22, CLT^{+}21, CW22, CZ22, CWW^{+}23b, CMLL21, CCD^{+}20b, CMY21, DPB22a, DBD22, DL23, DBK21, DPBP24, DP22b, DBD23, DS22c, DPSJ22, DS24b, EGB21, FCMM20, GJPS23, GAS23, GSZ^{+}20, GB20, GGS^{+}22, GK22b, GK23b, HML21, HV21, HM20, HNG22, HBB20, HE23, HXY^{+}22a, JAM24, JD22, JS23, JNMG21, KPJ^{+}21, Kan22, KMR22, VK24, KGM23, KKP24, KV22b, KA22a, KHK^{+}23, KKE^{+}22, KKC22, KKP20, KBJ21, KP21, KAR^{+}23, KSS24a, LMGG20, LMR22, LCSR21, Le23, LYL21, LSZ24, LLKS21, LF23, MNR^{+}22, MS21b, MDSG^{+}23, MYCH22, Mir22, MC20, MRKY22, Mon21]. approach [MCNR20, NA22b, NNJC23, NJK22, OHRS21, PS22b, PP21, PPM22, PK22d, QMC^{+}20, QHL^{+}24, RS23, RSM22, RK23b, RSKA23, RCYuRH21, RVAE21, RVJ^{+}22, RZ21, RSJ22, SYRS^{+}22, SJ22, SAN24, SC22a, ST22b, SO22, SK21a, SMBM24, SMD^{+}21, SAHAN22, SV22c, SKHL22, SR24, Shi22a, ST23c, SPKK22, SA22d, SR22c, SRIB23, SSPG20, TSCM22, TT21, TSV^{+}22, TÖ21, UYÖ^{+}22, UWF^{+}21, WBZ21, WHH^{+}20, WGDP23, WWF^{+}22, YPY21, YYM^{+}24, YA22a, Yil21c, YHL^{+}23, ZDH^{+}22, ZWCS20, ZZY^{+}22b, ZYPD24, ZWL^{+}20, uZKH^{+}20, vdSTC21, KKM20]. approaches [APP^{+}21, AJAA21, CL22b, ENB^{+}20, HK24, IAT^{+}23, Kal22, KIAA^{+}22, Ogi20b, Ogi21, YT21, Yil21b]. Approaching [TL21, JYL^{+}23]. Approximate [AAB23, AJP22, CSD23, HYG20, MKRK23, MW21, Pan23, SPSP23, SK20b, WQY^{+}22, XLWX20]. approximately [LFW20]. approximation [HTZ^{+}22, KC22, MTK^{+}21]. apps [DKA^{+}21, HLG24, MAB22]. Apriori [SKS22]. APT [SKPD23]. Aquila [Den24, AC22, LM22b]. AR-DenseNet [GA22]. Arabic [TH22]. Arabica [RTBC23]. arbitrable [TY22]. ARCHER2 [LCR^{+}24]. archetype [MMR^{+}22]. Archimedes [KKKS23]. Architectural [GQ20, JQ22, LJ22, ZZC24]. Architecture [BR21, HXZS23, SK20a, AMR^{+}21, APV23, Ano21b, Ano21i, BPAE20, BMG22, CRB23, CC22b, CD22, DWDG20, DR22c, DMD^{+}20, DGSB20, DS24b, ENB^{+}20, FSFM22, GDA^{+}21, HNSS22, HPS23, JSAA22, KKR23, KAP20, KSK23, LAH^{+}22, LLC^{+}22b, MS21a, MTY21, MCT22b, NFF21, PD20, PAN22, PK23a, PKB^{+}23, PK22c, PD22b, RS23, RTBC23, SRGK24, Shi22a, SG22e, SGHL20, TTB^{+}22, TAB21, TTZX22, TA22b, WHH^{+}23, XW24, YV23, YLM21, YPO21, ZYH^{+}23, dCMM21]. Architecture-oriented [HXZS23]. Architectures [KLL^{+}21, Ano21-38, ATC23, BKGC23, BVP22, DJF21, DJ20, FFG^{+}23, GDGN23, GA23, HK24, HOS^{+}21, HVB22, IG24, KKE^{+}22, LGA^{+}20, NMS^{+}21, OCR^{+}23, SSN21, SCKU23, SPC^{+}21, SSP24, SKÇA23, VMFL23, XZXV21]. Area [MS21a, BZWH21, CCD^{+}20b, DK21, GDA22, GQ24, GRMP23, NVV24, TKG^{+}23, VB21]. Area-efficient [MS21a]. ARFC [AAN^{+}21]. argument [Aru22]. argumentation [ALNJ21]. ARIMA [TT22]. arithmetic [AE22b, BFR24, CSD23, FFLM21, VKSS23, ZZY22a, ZSQ22]. arithmetical [SP22a]. arm [NBHN22, TA22b, HNSS22, SGHL20, ZDL^{+}22, GHL^{+}23, JOW23, MSPPD20]. Arm-based [JOW23, MSPPD20]. ARmed [FDH22]. ArmurMimus [BK22c]. ARO [CK24a]. arrangements [FGJ^{+}21]. array [HZW^{+}23, NWW^{+}22, RRIL22, TAB21]. array-unit [HZW^{+}23]. arrays [SM24b]. Arrhythmia [AS20, Hem22, KNM21]. arrival [DWZ^{+}20b, FKGO22, LML^{+}23]. art [CI20, ELIG23, MTK^{+}21]. article [GU23]. artifact [KYZ20]. artifact-centric [KYZ20]. artifacts [PCC^{+}21]. Artificial [BJK^{+}24, GSS23b, KCM^{+}22, PS23c, SAD^{+}21, SRBH22, WZMJ23a, AAK^{+}21a, AGSN23, AMM^{+}20, BTP^{+}21, BPB21, CK23, Che24, CR23, DS22a, FSdP^{+}23, JMV24, JR22, KPJ^{+}21, KSJK21, KAO22, KRSR23, LZQ^{+}22, LYG^{+}21, ND23, NSK^{+}24, PKR22, QPS20, RP21a, RRIL22, WJJ23, XWW^{+}21]. ary [ST23b]. Asian [KSJ23b]. aspect [KS22c, KS23c, SSP22, XGX^{+}21]. aspect-based [KS22c]. aspect-category [XGX^{+}21]. aspects [GQ20, HNS^{+}21, HKMS21, NSKS21, YFdP^{+}23]. aspiration [NS23]. ASRNN [KKKS23]. Assamese [DM21]. assembly [MW21]. assess [MYCH22]. Assessing [Kim21a, MFA^{+}21, WHH^{+}20, YA22b, DJJR22]. Assessment [LM20b, DPdS^{+}23, JMY21, LC20, LGW^{+}22a, LLN^{+}23, NBHN22, SGS21a, SP22b, SKK22, VP22b, WKY22, XLZZ22, ZWZ^{+}21]. assets [PSHJ20]. assign [JHG23]. assigner [GD22]. assignment [APP^{+}21, BZWH21, IAQ20, MSS^{+}20, SGS21b]. assimilation [CD24a]. assistance [SD23]. assistant [CZZ^{+}22, CCdCC21, RRJ23]. assisted [Agr21, AP23, CDR^{+}23, DZW23, DKK24, GS22, HT21, JKKL21, LA22, LGL^{+}21, LF23, MC20, PS21, QHL^{+}24, SS23d, SP23b, SPK^{+}22a, SSS23b, SKR^{+}24, TZ22, VB22a, WGYZ22, XW24, Zhu22b]. assisted-Internet [VB22a]. associated [FEH22, GLW22]. Association [KRSR23, LK22a, AA22b, LJC23, MS22b, SM21, WWG^{+}20]. associations [TC24]. assurance [AADS21, BPT^{+}23, ZARR23]. assured [TT23a]. asthma [IDA22]. Astra [GHL^{+}23]. asymmetric [Ano21-38, BSPM23]. asynchronous [BPAE20, JLZ^{+}24, KAQ23, LYL21]. ATMathCoreLib [KHK^{+}24]. atmosphere [NAK^{+}22]. atom [KA22b]. atomic [BOI23]. atrial [AST22]. atrous [RMMK23]. Attack [Gün23a, AP22, AS22b, AA23b, AI21, CCBA23, CSR24, CZTC22, Dah23, DP22a, HK22a, HK24, JDG22, JD23, K22, VK24, LK23, LPW^{+}21, LMM^{+}22, MP23, NM20, PA23, PD23, RGKK21, SNBAI24, SV22b, SD23, TK22a, VJ22, VRV23, WKL^{+}22, XXSL23, ZLM22]. Attack-wise [Gün23a]. attacker [DRV22]. attacking [SAC22]. attacking-based [SAC22]. attacks [AA23c, BMZ^{+}22, CSC22, CYZX23, DCZ^{+}22, GQ21, HAA23, KKP24, KSTV21, LLM^{+}22, MES23, NJK22, PKK21, SKPD23, SSS^{+}20, SR22c, TPT^{+}22, VD21, ZWW^{+}21, ZCZ^{+}22]. Attention [AWS^{+}22, LZSC23, ZLX^{+}24, AAA22b, BKJ22, CZG^{+}20, CPQ^{+}22, CL23, CKL20, HR22a, HLZ23, HZZC22, KH22, KXL^{+}21, KKKS23, LLL^{+}24, LZY^{+}21, LLX^{+}24, SMR23, SSP22, SSP23, SZG^{+}24, SBS22, DM22a, THW21, TWXL21, VKSS23, WMS^{+}23, Yil22b, ZFF^{+}21, ZYX^{+}21, ZLCS21, ZRY^{+}22, ZCZC23, ZKZ^{+}23, ZCX^{+}21, ZDJ^{+}21]. Attention-based [ZLX^{+}24, KXL^{+}21]. Attention-guided [LZSC23]. attentive [QWL24, YLWL20]. attestation [VCBB20]. attraction [Fio20, ZZS^{+}21b]. attraction-repulsion [Fio20]. attractiveness [LWZY23]. Attribute [LJZ21, CXX^{+}22, CYDW20, HSR23, HLC^{+}21, KGM23, LFG^{+}22, MG20, NA22a, RT22a, SAN24, VK23a, XZW^{+}20, ZHDW24]. attribute-based [CYDW20, HSR23, HLC^{+}21, MG20, XZW^{+}20, ZHDW24]. attributes [HC22, LHL^{+}22, MSA22, RN22, ZCL21]. attribution [GADM20]. Auction [BC23a, HJZ^{+}22, KRKM22, KHR22, SZJ21, XHZ^{+}21, XWC^{+}22]. auction-based [KRKM22]. audio [JYW^{+}20, JJZ^{+}21, KDL20, YSP23]. audio-visual [JJZ^{+}21]. auditing [AR23, AS24, BV22b, GFQ20, LLJR21, TY22, YR21, YZXL22, ZWO^{+}20, ZLO^{+}21]. auditory [JYC^{+}21]. augmentation [APV23, LAE^{+}22, LWHW22, SET^{+}22]. Augmented [HK22b, KJS^{+}20, RR22, JHD^{+}21]. augmenting [BLT22]. Aurora [VMFL23]. authentic [GPR^{+}22]. authenticated [AMAT22, TJC24, WGYZ22, ZWC^{+}22]. Authentication [LF23, BT21, BK22c, CYC21, CCM22a, CBK23, DAT23, DK21, EI22, FCX^{+}22, GAH^{+}22, GJPS23, GMP^{+}20, HSR23, DR22a, JPA^{+}23, JPH23, KGRG24, LYL20, LAH^{+}22, DPD^{+}22, NVV24, NT23, PSP24, PS22c, PCK23, PDRM24, PA21, RAN22, RÖ22, SS23c, SS23e, SG24, SS22e, SHV24, SJ23b, VDL23, VB22a, ZZK^{+}22]. author [HBB20]. author-topic [HBB20]. authority [SAN24]. authorization [AAD20, WSZ^{+}23]. authors [MKBB22]. Autism [YY22, BGG24, RM23]. Auto [VB24, AV23, AFKS23, Anb22, CLLB20, GVSS22, KSVP22, KDA^{+}22, PPA22, SPK22b]. auto-encoder [AV23]. auto-metric [GVSS22]. Auto-scaling [VB24, AFKS23, CLLB20]. autodidactic [RSJ22]. autoencoder [CD24b, MWH^{+}23, RRB22, RZ24, XY21, ZHX^{+}23]. autoencoders [XLL^{+}21]. autoencoders-based [XLL^{+}21]. automata [AG22, HLO^{+}21, KA22a, KYP21, WZ20b]. Automated [BWTJ20, GM22c, HD23, KTM22, Kay22a, LCR^{+}24, NSBT21, BCK22, CMK22, ÇG21, DM22c, GSS23b, HML21, HXY20, JM23, JK22a, Kan22, KKC22, KY22, KST23, MNDK22, RBWB21, RD23, SGS21a, SLL22, YSP23, YT21]. Automatic [KHK^{+}24, LSLY20, SBA22, VRR^{+}22, Yil22a, AAMAA22, AMM^{+}20, CCPP21, CZZ^{+}22, DR22c, FKK23, HYT^{+}21, LPHK20, TNP21, XAC^{+}20, ZLX^{+}24]. Automatically [FKO23]. automating [ZYPD24]. automation [MCD^{+}23, MLC^{+}21, SP22c, SKSP23]. automaton [WZ21]. automaton-based [WZ21]. automobiles [UE22]. automotive [Ano21-43, BPB21]. autonomic [KLK23, MMR21, PR24, SAM^{+}23, SM22d, TS22]. autonomous [ATG23, BF24, ÖGS22, PPM22, PB22b, WZL^{+}22, WHH^{+}23, YZYT21]. autonomy [MTY21]. autoregressive [SMD^{+}21]. Autoscaling [SP21a, CMT20, KSSK22, Mon21]. Autotuning [BP20, WKB^{+}22, OHFF20]. autoupdate [KTU^{+}21]. auxiliary [BK22b, HUC^{+}22, YAR22]. available [LZZ^{+}23, SYG22]. average [CA22, DTK24]. avoidance [LSL20, SWK22, WLLX21, ZWC^{+}23]. Avoiding [ESB20]. AVS [LLAV22]. AVX [SKSP20]. Aware [Vin21, AAK^{+}21b, ACJ21a, ACC^{+}20, ARS22b, AGSN23, BC21, BC23a, BR21, BABLH21, BJGF20, BVM22, BZL^{+}22, BPK24, CSV22, CCCR21, CH21, CPLX21, CLC^{+}23, CD22, Cza23, DAT23, DS24b, FM20, FDY21, GMA20, GJBM22, HAS^{+}22, HMK23, UZAA21, USK23, HGM^{+}23, IA22, IAQ20, KB21, KK21b, KVV20, KA23, KHHK21, KSKS22, KLK23, KSSK22, KSB23, LMGG20, LBA23, LLX^{+}21, LXW^{+}23, LWL^{+}24, LLH^{+}20, LWCZ22, LCL^{+}22, LZSC23, LXKW23, MSS22, MLZ^{+}20, NGOS22, NT23, NBPR22, NSKS21, NLG^{+}20, PYC^{+}20, PLZL24, PEGP23, PB22b, PS23c, RGKK21, RSM22, RSM23, RZCA21, RT22b, RYG^{+}21, SAQJ23, ST22b, SKH^{+}21, SPS23, SS23d, SCD^{+}23, SZZ^{+}22, SR20a, SCM22, SZG^{+}24, SAF^{+}23, SP23b, SK21b, SKRS21, SRG^{+}21, SPK^{+}22a, SSH22, SM22d, SKR^{+}24, TWQ^{+}21, TAH22, TRM^{+}24, UK23, WZHL21, WZZ^{+}22, WFY22b, XYFZ23, YHOY22, YLW^{+}22]. aware [YSS^{+}21, ZQX^{+}23, AASPR22, DJGF21, Mir22, SG22d]. aware-cost [RSM22]. awareness [KGE^{+}20, LHX^{+}23, XXD^{+}22, XLL^{+}23]. AWC [Gün23a]. AWC-NIDS [Gün23a].

B [RB22]. B-SCORE [RB22]. B2B [KY23]. B5G [SAGS24]. BAC [WMQ^{+}22]. back [FSFM22, Shi22b, TK22a]. background [LLSC22, XCY22]. backhaul [IAQ20]. backpropagation [Cho20]. backup [WFY^{+}22a, YLW^{+}22]. backward [ZLV22]. bacterial [CÇ22a, Rav23, SK22a]. bad [TPT^{+}22]. badger [DBD23, KMR22]. bag [SK20b, VPSM22, WLLZ20]. bag-of-tasks [SK20b]. Bagging [SP21b]. Bagging-support [SP21b]. bags [CKKK20]. bait [PA23]. balance [DÖD22, LCL^{+}22]. balance-aware [LCL^{+}22]. balanced [Liu24a]. balancer [AAK^{+}21b, EASN22, FKO22]. Balancing [MV22, AB22b, AET^{+}22, Agr21, AAG^{+}22, AFNH21, BM22a, CDV23, CWFK23, DDUK23, FKO23, GP22a, HAS^{+}22, HYK21, HLL^{+}21, JAM24, KK22a, LLAV22, LCZ^{+}20a, MSSM24, MG21a, MV24, NA22c, SHBC20, SMD^{+}21, SKSN22, SSV24, ZBZ^{+}20]. ball [EASN22]. balsam [LS20]. band [wSYyCsD23]. bandit [WWW^{+}20]. bandwidth [BZT^{+}23a, CON23, JYW^{+}20]. banking [AAE23, DR22a, RVF22]. bare [YKL23]. barnacles [DBK21]. barrier [EGGG23, YCY20, ZGH^{+}22]. BArt [PSHJ20]. bartering [KHR22]. BASARIM [TC22, KSY23]. BASARIM-2020 [TC22]. base [STH^{+}20]. Based [LWZC21, LCZY20, MV22, ZXL^{+}21b, AR22a, AC22, AM22, AFF22, AET^{+}22, AAMM23, AFK^{+}22, AV22a, AK22a, AMJK21, AKA^{+}22a, AMAT22, AR23, AMRH21, AK21, AP23, AAN^{+}21, AS22b, APM22, AZI20, Aka22b, ATS22, AK22c, AFBM^{+}23, ANAMSAR21, AOACAQ21, AY21, AV23, AFKS23, AD22b, ABCP23, AA23b, AAARR20, ABA22, AA22b, ALR22, AA23d, MS22a, AZA20, AYG^{+}21, AWS^{+}22, AJS23, AADS21, ASC22, Ano21c, Ano21g, Ano21h, Ano21j, Ano21-43, AS24, ASMS21, AKZA22, Aru22, AA23e, ASRN23, AMM^{+}20, AD22c, AYKE21, ALB^{+}20, AGM23, ACJ21b, AYJ^{+}22, AY23, AA24, BCM22, BSN22, BST^{+}22, Bai23, BEKS22, BK22a, BDK22, BKK22, BDR23, BTDD20, BXH^{+}23, BKJ22, BBD24, BP24, BEJD22, BC23b, Bei24, BLT22, BABLH21, BBB^{+}20a, Ben21, Ben22, BFM^{+}21, BKGC23, BJGF20, BKS22, BRNR23, BJS23]. based [BTT21, BZGM22, BJWY20, BKLY20, Boz22b, Bul22, CWLL20, CLLB20, CNG^{+}20, CLDY21, CdRNB23, CF21, CGS^{+}21, CHL^{+}24, CCD^{+}20a, CMS21, CÇY22, CÇ22a, CK23, CDV23, CK24a, CAAHC23, CJ21b, CH21, CSWC20, CDC20, CYQ^{+}20, CJY^{+}20, CLC^{+}20, CCGN20, CPLX21, CK21b, CYC21, CCZ^{+}21, CWL^{+}21, CCL^{+}22, CW22, CZ22, CWL^{+}22, CYZX23, CCZM23, CWW^{+}23b, CWW23a, CDR^{+}23, CBFS23, CHLD23, Che24, CLWX21, CLY^{+}21, CYZ^{+}21, CZTC22, CMJM22, CK24b, CYW^{+}22, CA22, CPH20, CD22, CPCK23, CWFK23, CÇM22b, Cob22, CMA^{+}21, CDN^{+}21, CPPP21, CMY21, CYDW20, CZL^{+}22, CWH^{+}24, Cza23, DR21, DDZ^{+}20, DBD22, DRV22, DS23a, DDH^{+}20, DM22b, DM23, DS21, DZCL22, DJST24, DPB22b, DDUK23, DK21, DKL21, DS22a, DCR23, DP22b, DG21, DR20, DWZ20a, DBD23, DXXL20, DZYY22, DCZ^{+}22, DSSS22, DK22b, DWY^{+}21, DQF^{+}23, DLC^{+}21, DSC^{+}21b, DSS^{+}23, DRMA22, DPSJ22, EVVR21, EGB21]. based [EASN22, EI22, EASR22, FJCFB23, FSA^{+}22, FPHZ20, FC21, FYH^{+}21, FCX^{+}22, FLG^{+}22, FGZC23, FQD^{+}23, FEH22, FM22, FMJ^{+}22, FM20, FSWW21, FHS^{+}22, FDD^{+}24, FRS^{+}23, FKO22, Fio20, FKGO22, FKY^{+}23, FAM22, GBBS21, GZ20, GPDB20, GDA22, GJPS23, GSS^{+}23a, GMM22, GCS23, GXL^{+}20, GWGR20, GZC^{+}22, GLN23, GAS23, GKÇ22, GPR^{+}22, GHRM21, Gha20, GBS24, GB20, Gho21, GGA24, GSG^{+}23, GSS23b, GRMP23, GCF^{+}20, GCQ^{+}24, GXH^{+}21, GMP^{+}20, Gui22, Gul22, GNS22, GATK22, GC20, GYL^{+}21, GZY^{+}22, GQ21, GSVS21, GK22b, GD22, GKAO20, HK24, HML21, HSR23, HVT23, HMR24, HH23, HLL^{+}21, HWG23, HYC^{+}24, HM20, HGNN22, HUJ22, HD23, HBB20, HNSS22, HLT23, HL20, HXZH21, HXY^{+}22b, HXY^{+}22a, HDS^{+}23, HML19, HHXH20, Hua20, HFH^{+}21, HLC^{+}21, HYT^{+}21, HSL^{+}22, HHYL22, HZZ^{+}23, HWY^{+}23, HUC^{+}22, HLO^{+}21, HAAF22, HAA23, IMNC22, IK22, IVP^{+}23]. based [JR22, JPA^{+}23, JB24, JPK22, JKB22, JNM^{+}22, JK22a, JAR^{+}24, JB21, JA23, JKP22, Jeo20, JOW23, JM22, JY20, JS23, JWL20, JZB20, JMY21, JJZ^{+}21, JYC^{+}21, JZC^{+}23, JYH^{+}24, Jia22, JW22, JKKL21, JDG22, JBCI20, JLE22, KDL20, KEK^{+}20, KSS21, KTM22, KUK22, KRKM22, KE21, Kab22, KT20, KMS^{+}21, KHEF22, KXL^{+}21, KT22a, KMR22, KKKS23, KT23, KB22, VK24, KDC22, KRK21, KKS22, KK22b, KGM23, KBR24, KBK^{+}22, KC22, KS20, KQK^{+}20, KBL^{+}21, KCP^{+}22, KV24, KAP20, KTU^{+}21, KHK^{+}23, KAQ23, KM23b, KK23a, KGGM22, KKC22, KEMZ22, KKP20, Kim21b, KJHM21, KMD23, KYP21, KY22, KAO22, KHY^{+}20, KK22d, KWL^{+}24, KT22b, KRSR23, KSSP22, KSA^{+}21, KSK^{+}20, KSTV21, KSVP22, KR22c, KS22c, KSKS22, KR23, KS23b, KSJ23b, KSY24, KSS24a, Kum22, KCM^{+}22, KBS^{+}22, KSS24b, KS23c, KSK23, LMGG20, LS23a]. based [LHJ22, LBA23, LMR22, LGM21, LK22a, LJ22, LK22b, LS22, LLR^{+}21, LKR^{+}22, LLAV22, LCSR21, LYL20, LMBK23, LSS^{+}21, LG23, LGDW22, LHWT20, LW20a, LN20, LCZ^{+}20a, LW20b, LZF20, LJZ21, LLZ^{+}21a, LYSC21, LLMX21, LH21, LPW^{+}21, LZL^{+}22, LM22a, LDZ^{+}22, LYC22, LM22b, LLC^{+}22a, LGL^{+}22, LLW^{+}22a, LGZ^{+}22, LHX^{+}23, LWZY23, LLW^{+}23, LSZ^{+}23, LDS^{+}23, LCW^{+}23a, LML^{+}23, LSZ24, LLTZ24, Lia24, LZZ^{+}23, LLW^{+}21, LWW23, LJB22, LC20, LJP^{+}21, LLKS21, LZC21, LSX21, LC21, LGW^{+}22a, Liq22, LS23b, LZY^{+}20, LGLZ20, LLLX20, LSL20, Liu21, LWCM21, LLT21, LZY^{+}21, LCW21, LZW^{+}21, LPC^{+}21, LPZ^{+}22, LWCZ22, LXC^{+}22a, LAH^{+}22, LCL^{+}22, LLL^{+}22, LXT^{+}22, LWHW22, LLJ^{+}22b, LYF^{+}23, LZLZ23, LZL^{+}23, LL23, LZQ^{+}22, LLL23, LWZ^{+}20, LSW21, LYG^{+}21, LLN^{+}23, LL24, LGX^{+}23, LCZ^{+}20b, LCM22, LLYZ23, LLC^{+}21b, MBM20a, MNR^{+}22, MKRK23, MM24, MGSB23, MSSM24, MBB22, MG20, MQEK21, MS21b, MK22b, MSJ22, MB21, MMR21, MSPPD20, MKA23]. based [MS22b, MS21c, MY24, MLKD20, MPV22, MSK22, Mis22, MMS24, MT22, MRS^{+}21, MHB^{+}22b, MSA21, MGN^{+}22, MG23, MSL22, MLC^{+}21, MSBR23, MM21, NA22a, NBK22, NSR22, NNVD22, NBS^{+}22, NT23, NM23, NET20, NET21b, NLB22, NA22b, NS23, ND23, NAR^{+}22, NR22, NdSSSN20, NGD^{+}22, NY22, NWW^{+}22, NR23, NBHN22, NMS^{+}21, ÖUG22, OO22, OKJ^{+}21, ODK^{+}23, Ona21, OMA^{+}23, OO23a, OBER22, ÖGS22, Pal22, PT22, PSP22, PAN22, Pan24, PLP22, PMS^{+}21, PMC^{+}21, PPM^{+}20, PR23, PBD23, PR22, PKB22, PK23b, PSB^{+}20, PST24, PJK23, PPM22, PP20a, PGL^{+}23, PL21, PA23, PQKDT21, PP20b, PKB^{+}23, PCC^{+}21, PEGP23, PK22c, PKK23, PJP21, PVRM22, PCR21, PM23b, QWW^{+}22, QPS20, RV21, RB22, RRB22, RJH^{+}23, RM21, RJ22, RGKK21, RSS20, RS23, RSM21, RSM23, RK23a, RT21, RT22a]. based [RSKA23, RRJ23, RAN22, RP21c, RK21b, RRIL22, RG22, RVJ^{+}22, RV23, RNRK22, RSJ21, RÖ22, RAG21, RG23, RMR^{+}22, RD23, RT22b, RCBTGE23, RKC22, SB23a, ST22a, SS23a, SR22a, SYRS^{+}22, SJ22, SA22a, SAN24, SMM22, SKPD23, SNET21, SSMT22a, SSMT22b, SP21a, SDR23, SKB^{+}20, SM22a, SK22a, SNBAI24, Sao22, SKPV22, SPS23, SO22, SAC22, SRG^{+}22, SCKU23, SRL23, SRGK24, SS23e, SS22a, SL20, SEC22, SMBM24, SHBC20, SPSP23, SN22a, SEM^{+}20, SMD^{+}21, SBB21, SSN22, SLJ23, SLG^{+}20, SH22, SVD^{+}22b, SN22b, SRRM23, SKHL22, ST22c, SM23a, SG24, SR22b, SR24, SHZY20, SZqWZ20, SDSW21, SZGR21, SMR23, SM22c, SM23b, SM23c, Shi22a, Shu22, SGH23, SV22d, ST23c, SKS^{+}23, SR23, SSK23b, SSV24, SA22d, SMKA22, SP22c, SM24b, SR22c, SXC^{+}23, SLL^{+}23, SSP24, SK22d]. based [SRIB23, SBB^{+}20, SAL22a, SSCN23, SM21, SD23, SP23b, SM24c, SR20b, SKSP23, SB21, SHVA23, SKRS21, SP22d, SJ23b, SPK22b, SSH22, SV22e, SM22e, SLHW20, SP21b, SFJ^{+}21, SZZ^{+}24b, SK21c, SKÇA23, SBS22, SSRA23, SDR20, SM23d, TA23a, TWW^{+}22, TWL^{+}20, TSLL21, TFZC23, TCRP23, TJC24, TLQ21, THW21, TWW^{+}21, TDZ^{+}24, TLH^{+}22, TS24a, TAB21, TCW^{+}23, TSG21, TK22a, TP20, TWXL21, TY22, TTZX22, TZ23, TT23a, TC24, TAH22, TT21, TPT^{+}22, TS22, TSV^{+}22, TT23b, TSB23, TG23, UDS21, USP^{+}23, UAS^{+}22, USI21, UPGCA22, VKSS23, VDL23, VRV23, VM23, VLVS22, VRS22, VB22a, VK23a, VD21, VKK^{+}24, VP22b, VK23b, VCBB20, VG21, VPQ22, Wan22a, WCL^{+}24, WZL^{+}24, WJL^{+}20, WWG^{+}20, WSM^{+}20, WGY20, WZ20a, WGZ^{+}20, WWG21, WMC21, WAY^{+}21, WLZ21a, WJLC21, WFHC21, WLL21a, WLLX21, WEH^{+}22, WWZ^{+}22, WQY^{+}22, WWA22, WZC^{+}22a]. based [WZZ^{+}22, WZMJ23a, WHH^{+}23, WMS^{+}23, WSZ^{+}23, WL23, WGLL23, WZMJ23b, WER^{+}23, WRE^{+}23, WQS^{+}23, WHH^{+}20, WLLZ20, WZ21, WN21, WZX^{+}22, WN23, WNZ24, WWL^{+}20, WYZAD20, WZG^{+}21, WLX21, WLY^{+}22, XCY22, XGX^{+}21, XAC^{+}20, XLL^{+}21, XW24, XLZ20, XJW^{+}22, XLZZ22, XGZ^{+}20, XZW^{+}20, XLXZ20, XLWX20, XL21, XWW^{+}21, XA22, XHM22, XDL22, XLZL22, XCD^{+}20, YAR22, YS22, YV23, YKL23, YTW^{+}24, YL20a, YLLL20, YMZ^{+}20, YLM21, YYLL22, YZX^{+}22, YSLX22, YYPR22, YSK22, YCL^{+}22a, YLS^{+}23, YFF22, YLKC24, YBJ^{+}23, Yil21b, Yil21c, YJJ23, YLGY20, YOWY22, YLJ22, YGS^{+}23, YYY^{+}23, YPO21, YF20, YLWL20, Yu21, YZ21, YWC^{+}23, ZARR23, ZABT^{+}20, ZBZ^{+}20, ZWCS20, ZWO^{+}20, ZZL^{+}20, xZlGCzJ20, ZHJ20, ZLW^{+}21, ZFF^{+}21, ZZWZ21, ZLCL21, ZQL^{+}21, ZGLS21, ZLCS21, ZLD22, ZLL^{+}22, ZYCS22, ZZY^{+}22b, ZZL^{+}22, ZZW^{+}22, ZZY22a, ZCZW22, ZMLW23, ZCZC23, ZYH^{+}23, ZWL^{+}23, ZKZ^{+}23, ZX23a, ZZC24, ZTF^{+}20, ZXL^{+}21a, ZX21b, ZWJG21, ZTS^{+}22, ZTMC22, ZGG^{+}23, ZTMC23]. based [ZCH^{+}23, ZHDW24, ZWY^{+}24, yZyWD^{+}21, ZPL21, ZNLL22, ZLC^{+}22, ZJSJ20, ZJSJ21, ZCL21, ZOS^{+}21, ZWZ^{+}21, ZFW23, ZWL^{+}20, ZZS^{+}21b, ZXLD21, Zhu22b, Zhu22a, ZHWY22, ZWC^{+}23, ZQX^{+}23, ZLX^{+}24, ZS22, ZGH^{+}22, dOPBdO21, rSN21, vdSTC21, GAH^{+}22, HE23, JM23, LLX^{+}21, NMS^{+}22, ÖTT23, SS21, SBSK22, Tur23, VSP22]. based-on [WJL^{+}20]. based-rectified [OKJ^{+}21]. basis [DT22a, Pan23, ZCR23]. basket [WWG^{+}20]. bat [AP23, CGW^{+}20, Kha22a, SS22a]. bat-based [AP23]. batch [DMPR24, GGCGS20, GRL^{+}22, HXST22, YBJ^{+}23, YJJ23, YZXL22]. batching [SRG^{+}21]. Battle [HHP23]. bay [FZ23]. Bayes [CDC20]. Bayesian [BM22b, CSL20, CdRNB23, LHL^{+}20, SR22a, SMR23, SP22c, WKB^{+}22, WLDW22, YZX^{+}22]. be [AYD21, KUK^{+}24, MLVGR24]. beaconing [RP22]. beam [GDFDF22, NKST24]. bearing [HLC23, wSYyCsD23, XLL^{+}21]. Bee [PS23c, AD22b, APM^{+}21, JR22, LZQ^{+}22, LYG^{+}21, RP21a, SK22a, WJJ23, XWW^{+}21]. beetle [BRS^{+}22, CEN22, PB22b, SB23a]. before [HSL^{+}22]. behaved [YLJH22, LSL20]. behavior [AADS21, BPB21, DPFC20, DM23, EVVR21, EI22, FRS^{+}23, FGL^{+}20, GO22, LBA23, LYW^{+}21, LFG^{+}22, LDZ^{+}22, LWZY23, OA22, RM23, SKPD23, SEM^{+}20, SG22c, SLL^{+}23, UMR23, WZZ^{+}22, XZY^{+}22, ZWCS20, ZYXX23]. behavior-aware [WZZ^{+}22]. behavior-based [SEM^{+}20]. behavioral [JAC^{+}21, ZWL^{+}20]. behaviors [SYRS^{+}22, SGS^{+}23, YYLL22]. Beijing [LWCM21, xZlGCzJ20]. belief [AK22a, BKK22, DP22a, USK23, MES23, MRKY22, SM23d, VS22b, WKY22, WWL^{+}20]. belief-desire-intention [USK23]. Belt [CLL^{+}21]. benchmark [CMS24, EA22, GSG20]. Benchmarking [BAR21, MSPPD20, Wri22]. benchmarks [WKB^{+}22, WT23]. benefit [Zha22]. BERT [Gho21]. best [MTD^{+}20, YJJ23]. best-effort [YJJ23]. beta [AA22a, MAAA22]. better [RS21, SC22b]. between [CJ21a, CMLL21, DPB22a, JZB20, LLSC22, LL21, SGHL20, TI22, WHH^{+}20, WK20, ZTL^{+}21]. betweenness [RTSK23]. beyond [Ano21b]. Bézier [AK22c, Bul22]. Bi [Asl22, KML23, PJK23, VK23b]. bi-directional [PJK23, VK23b]. bi-objective [KML23]. bias [VJ22, YB23a]. biased [CdRNB23, LAP23]. Bibliometric [RBDS22]. bidiagonalization [KSA^{+}21]. Bidirectional [XFH^{+}21, BJWY20, CL23, DSS^{+}23, JR23b, KH22, Kab22, SM23b, SGS^{+}23, SSP23, TYA22, ZLCS21]. bifolded [JB21]. Big [AJAA21, Haj20, KUK22, PKKL21, RC22, AAN^{+}21, AA21, AYB21, BZEM20, BJGF20, CZUPA23, CJ21b, COZ21, DA22, DYF20, DJGF21, EAS23, Faz23, HAR24, JSS22, JAC^{+}21, JQ22, KT20, KGW^{+}20, KKR23, KKAM21, KVV20, KLJ21, KMD23, LCKJ21, LMBK23, LSD21, LZM^{+}23, MG21a, MBO^{+}21, Ogi20a, OLF21, PS22a, PS22b, PLP22, PD22b, RCYuRH21, SM24a, SVD22a, SMD^{+}21, SSN22, UKS22, UGK^{+}22, VRS22, VR22, Xia20, XGZ^{+}20, ZLW^{+}23, ZS22, AFK^{+}22, MBM^{+}20b, TD21]. big.LITTLE [SGHL20]. bike [SVB23, XDH^{+}20]. bikes [SVB23]. Bilateral [KSVP22, Wan24]. bilinear [HHYL22]. BiLSTM [KXL^{+}21, KS22c, SMR23, SD23]. BiLSTM Attention [XLZL22]. bin [AYH^{+}22]. Binary [Kot23, PVVS22, Sao22, AS23a, MS22a, AI22, Ano21-37, BCM22, CPCK23, KIN^{+}23, KS22d, LNI^{+}24, RM23, SV22a, SS22a, SDSW21, SBS22]. Bing [SG22b]. binomial [AAA^{+}22a, Çet23]. Bio [Kim21b, VNP^{+}23, CO21, MMR^{+}22, SPQM20]. bio-engineering [SPQM20]. Bio-inspired [VNP^{+}23, CO21, MMR^{+}22]. Bio-signal-processing-based [Kim21b]. biochip [RP21a]. Bioinformatics [MCL^{+}20]. biological [HIEH22, JNMG21]. biomarkers [PP20b]. biomedical [ASAAAA22, BWW^{+}20]. Biometric [ZLZ^{+}22a, BT21, BK22c]. biometrics [FCX^{+}22, MG21b]. biometrics-based [FCX^{+}22]. bionic [CCZM23]. biophysical [CS21]. bipartite [VVK23]. bipolar [GS22]. bird [ODK^{+}23]. biserial [CT22]. Bit [KK21b, LNI^{+}24, MKL21]. Bitcoin [HR22a, BDG^{+}23, PS23a, Ano21a]. biterms [YHOY22]. BitFlow [HGDD20]. bitmap [Yil21a]. BitTorrent [MTT20]. bivariate [CT22]. Black [PD23, JR23b, RZVC21, XXSL23]. black-box [RZVC21]. Black-hole [PD23]. bleeding [WSM^{+}20]. Blind [DZCL22, LLN^{+}23, Gul22, KDL20, KM23b, RRB22, SA22d, SAL22a]. Block [BV22b, BDG^{+}23, CMS21, CA22, KKS23, MK22b, RKR22, SS23e, VVK23, WMY^{+}21, WMQ^{+}22, gZWfL^{+}20, AS24]. block-based [CMS21]. Block-Chain [AS24]. Blockchain [AR23, Ano21b, Ano21-43, DKK24, FN23, GDA^{+}21, JCG^{+}22, KDS^{+}20, KRK21, KAP20, KVP21, LCSR21, MKA23, PK22c, SSCN23, TG23, VP22b, ZARR23, AB20b, Ano21f, Ano21i, BK22a, BS23a, BZK^{+}21, CR23, CD22, DPB22b, DCK21, DGP20, EAvM20, FD20b, FZA22, GJPS23, HGDD20, JB21, KHEF22, KSA22, KM23b, KM23c, KSS24b, LS23a, LFX^{+}20, LZL^{+}22, LGL^{+}22, LLW^{+}22a, LZZ^{+}23, LWW23, LLM^{+}22, DPD^{+}22, MGN^{+}22, NBPR22, NAR^{+}22, PS23a, PGL^{+}23, QWW^{+}22, RB22, RS24, RSMCP22, SSS23a, SR23, SK22d, TWW^{+}22, TS24b, TXT24, VH22, VG20, WLL21a, WMQ^{+}22, WYL^{+}22, XWC^{+}22, XHST20, YLM21, ZML^{+}23, ZHDW24, AKA^{+}22a]. Blockchain-based [AR23, KRK21, TG23, VP22b, DPB22b, GJPS23, LZL^{+}22, LGL^{+}22, MGN^{+}22, PGL^{+}23, SK22d, TWW^{+}22, WLL21a, YLM21]. Blockchain-enabled [MKA23]. blockchain-powered [VH22]. blockchains [BAPS22, BANT20, DRMA22, FD20a]. blocking [EAvM20, NQ21]. blocking/faulty [NQ21]. blocks [BAPS22, LC21, WN23, WNZ24, BBBC22]. Blood [RPPK23, BP23a, BABS21, BDR23, CÇY22, LS23b, RK23a]. Bloom [FN23, KS24, WMY^{+}21]. Blue [BKM^{+}24]. blur [PGD^{+}22]. BMM [GRC^{+}23]. boar [JPO^{+}21]. bodies [CA22]. body [DK21, NVV24, TKG^{+}23]. Bolt [SXC^{+}23]. Boltzmann [CZL^{+}22, KDFK23, LLS22, MES23, WY20]. bone [KPP^{+}22, VB21]. Book [LHWT20, QMC^{+}20]. Boost [SS22c, CB22, VR22]. boosted [AYJ^{+}22, ZHW^{+}20]. Boosting [NTB23, CD24b, SKP22, WWW^{+}20, xZlGCzJ20]. boot [RK21a]. bootstrap [CT22]. Border [KBR24, BTT21]. Border-hunting [KBR24]. boring [GRC^{+}23]. Botnet [WKL^{+}22, BBD24, KH22]. botnets [WEH^{+}22]. bottleneck [LSLY20, MLS^{+}24]. bottlenecks [ZXL^{+}21a]. botulinum [KJS^{+}20]. bound [AHW22, KS22d]. boundary [JYL^{+}23, Mis22]. bounded [KK24, RPM22]. BOWTIE [WZX^{+}22]. box [RZVC21]. BP [PLX20]. BPaaS [HM20]. BPEL [YLZ20]. BPMN [YLZ20]. Brain [AJNS22, DBN^{+}22, AMM^{+}20, BCK22, GGJ^{+}24, IK22, KEK^{+}20, KSJ23a, Kot23, KK22f, LJBS23, MMR^{+}22, ODK^{+}23, PS23b, SET^{+}22, SAMS23, TAT22, TNP21, ZLD22]. BrainForge [VPB^{+}23]. brake [AZA20]. brand [YLKC24]. brands [KK22d]. breaches [MVR23]. Breast [SSMT22a, SKS22, AMV22, DSJ22, KUK22, KMD23, MH23, OE22a, OE22b, PBK23, PZD^{+}21, PP20b, PVVS22, TB23]. BRGraph [GRL^{+}22]. bricks [MCL^{+}20]. bridge [LHL^{+}20]. Bridging [OAS23]. brief [BEKS22, GKZ23, LL21]. bringing [NMQ22]. broadcast [BOI23]. broadcasting [Ngu21]. broker [MA22]. brokering [LGA^{+}20]. brown [HLO^{+}21]. brows [PB21]. browser [PCC^{+}21]. Brute [NMS^{+}21]. BSI [JCG^{+}22]. BSSM [AS24]. Budget [AAE23, CCCR21, GSVS21]. Budget-aware [CCCR21]. Buffalo [PJK23]. Buffalo-based [PJK23]. buffer [LLC^{+}21b, Shi22b]. build [KLJ21]. Building [MCD^{+}23, SPJI^{+}21, VPQ22, AFBM^{+}23, CRB23, PPM^{+}20, SMR21, SP22a, SCdLV20, TAT^{+}23, YLKC24]. buildings [LSZ^{+}23, YWQ^{+}21]. bulk [CLE^{+}20, Kal22, LZY^{+}20]. bulk-loading [Kal22]. bulk-synchronous [CLE^{+}20]. bundles [CCZM23]. buoyancy [WWW24]. burden [AV21]. burglary [ZLC^{+}22]. Burr [Çet23]. Burr-XII [Çet23]. Burrows [AD22c]. burstable [DAdBM^{+}24]. bursting [HXY20, SK21b]. business [Ano21e, BBB^{+}20a, ELIG23, ESB20, KMS^{+}21, KYZ20, LFG^{+}22, LPC^{+}21, SAM^{+}23, YYY^{+}23, ZLX^{+}24, YYY^{+}23]. Business-to-Business [YYY^{+}23]. businesses [KY23, LLSC22]. butterfly [BP23a, SMKA22, VM23]. BWA [PQKDT21]. BWA-MEM [PQKDT21]. byte [JBCI20]. Byzantine [ADK^{+}23, HNOI24, SKV22, VG20].

C [KV22a, DB23, GK23a, IGYS23, LS22, PEGP23, RBC20, RLdO20]. C-means [KV22a, GK23a]. C-ParGRES [RLdO20]. C/OpenCL [RBC20]. CaaS [PST24]. cabin [RKL21]. Cache [AML^{+}22, AZI20, CPLX21, KSD22, RPM22, SR22b, YMZ^{+}20]. Caches [SKS20]. caching [DZW23, HGW^{+}23]. calculus [VPQ22]. calibration [ASANR22, GMK^{+}21, SAHAN22]. call [SCS^{+}21]. Can [KUK^{+}24]. cancer [AMV22, BD22, BS23c, CGS^{+}21, DSJ22, ESS23, GM22b, JM23, JM22, KUK22, KMD23, LZY^{+}21, LM24, MGSB23, MQEK21, OE22a, PBK23, PAN22, MR23a, PZD^{+}21, PP20b, PVRM22, PVVS22, RR23b, RK23b, RPMA22, SMKA22, SKS22, TB23, WMS^{+}23, WT23, ZZQ^{+}22, ZWY^{+}24]. CANDAR [BI23]. candidate [AR22a]. CANDLE [WT23]. Canny [NS23, SXC^{+}23]. canonical [DDZ^{+}20, YSL^{+}21b]. capabilities [KTU^{+}21, Liu24a]. capability [NET21b, SYG22]. capability-based [NET21b]. capacity [LYF^{+}23, LGT^{+}23]. Capitalizing [OO23a]. capping [CDP^{+}21]. capricious [MKRK23]. capsule [KDA^{+}22, SR24]. CAPTCHA [JHD^{+}21]. captioning [CZ22, HGNN22]. capture [GBB22]. captured [SAGS24]. capturing [LBZ^{+}22]. capuchin [GSG^{+}23, GVSS22]. car [RKL21, YL20c]. card [GCS23, GQ21]. card-based [GQ21]. Cardiac [VLVS22, LJB22, VCFZ20]. cardinality [HXY^{+}22a, ZGWZ23]. cardiovascular [AR22b, Kan22, TD21]. cards [RVF22]. care [DM22c, KD22, KS24, LWZ22a, SAC22]. Carlo [YK23]. carrier [DZCL22]. carry [MKRK23, SPSP23]. carrying [LLKS21]. CAS [PD22a]. CAS-extended [PD22a]. cascade [SKP22]. Cascaded [AAARR20, AA23b, WSJ^{+}21, XLL^{+}20]. Case [EGB21, ZCD^{+}22, BSPM23, CPYC21, CCD^{+}20b, DM22c, DJST24, FSdP^{+}23, GADM20, LBFT22, LST22, LWCM21, LLA^{+}22, MPB^{+}22, RS23, SG22b, SSCN23, TTB^{+}22, TO22, WHDS22, ZDH^{+}22, ZCZW22]. case-based [RS23]. cases [UYÖ^{+}22, ZLC^{+}22]. cash [HGDD20]. cash-flow [HGDD20]. casino [Ano21b]. Cassandra [SLC20]. cassava [RKR22]. Cat [BMV22]. categorization [CMK22, JSAA22, SSMT22a]. category [XGX^{+}21]. catering [PLX20]. Catheter [KD22]. causal [GLW22, TP20, aWLY^{+}22]. Causality [LLZ^{+}21a, GLW22]. causality-associated [GLW22]. Causality-based [LLZ^{+}21a]. causally [KIW^{+}22]. Cautionary [SKP23, SKP22]. CCHybrid [Yu20]. CCIAI [PI21]. CCIAI-FKGA [PI21]. CCITL [GLN23]. CCSFLF [ZYY^{+}24]. CCTV [DDUK23]. CDMA [rSN21]. cell [AKRR20, CÇ22c, SAPC21]. cells [BABS21, BDR23, CÇY22, MAG^{+}20, ZCR23]. cellular [AG22, HLO^{+}21, KA22a, KYP21, Mir22, RSS20, SAPC21, SCS^{+}21, THT20, WZ20b, WZ21]. censored [Çet23, YK23]. censoring [SZS20]. Center [ZARR23, AAMM23, BM22a, FSFM22, HLL^{+}21, MS23, NFF21, SZZ^{+}22, WWW24, YLJ22, YTN^{+}20]. centered [AS20, Rav23]. centers [ASJ21, BZT23b, CK23, CWW23a, DCT^{+}23, FDY21, IPRS21, JW21, MT21, PKM21, WZC^{+}22b]. central [MT22]. centrality [CS23, CW22, HWG23, JS22c, MY24, RA21, RTSK23, SS23a, ST22d]. Centralized [PI21, BYTG22, NM24]. centric [dCJBP20, KYZ20, MAB22, MSB^{+}20, NAR^{+}22, SBGC21, SCdLV20, TAT22, XSGL20, YFdP^{+}23, YR21]. centroid [MG23, WWW24]. centroid-to-buoyancy [WWW24]. Cerebral [WSM^{+}20]. Certain [GDSS22]. Certificate [SK23, TG23]. Certificateless [DRMA22, AR23, DZYY22, GFQ20, JHS^{+}21, SM24a, TJC24]. cervical [JM23, PAN22]. CFD [OWB^{+}20, WZL^{+}24]. CH [RJ22]. chain [AD22b, BV22b, CA22, FMNF22, JAR^{+}24, JY20, KHEF22, KEMZ22, LS23a, MK22b, SS23e, SR23, WYL^{+}22, YYY^{+}23, ZTM21, AS24]. chain-based [JAR^{+}24]. Challenges [CR23, KMZ^{+}20, Kum21, MBM20a, AYH20, CY22, EAS23, GQ20, KK22a, KKK21, KK23b, PKK21, PUL20, RS24, YSP23]. chamber [JB24]. chameleon [ASC22]. Chang [ADA22]. change [BEÖ23, MFA^{+}21, WKY22]. changes [LM20b, SM22c]. Channel [LWZC21, BP24, IAQ20, RK23b, RKMS24, Sao22, SO22, SGS21b, SH22, WSZ^{+}23, WZSZ20, WJD22, ZDJ^{+}21]. Chaos [AA22b, Shu22, TDZ^{+}24, TSV^{+}22, WLJ20, GAH^{+}22]. chaos-based [TSV^{+}22]. chaos-RSA-based [GAH^{+}22]. Chaotic [rSN21, ABB22, DBD22, DWZ20a, JT23, KSS24b, LYSC21, LYC22, Lia24, MRKY22, RB22, SV22a, SEC22, SAL22a, TB23]. Chapel [HGM^{+}23]. character [LWZ^{+}20, SS23b, YKL23]. character-level [LWZ^{+}20]. characteristic [KCP^{+}22, LCZ^{+}20b]. characteristics [CPYC21, FCZ20, IA23, SZZ^{+}22, YSS^{+}21, ZWJG21]. Characterization [OE22a, PSF23, CSD23, RGPC23]. characterize [DGM21]. Characterizing [LGA^{+}20]. charge [NM23, SSRA23]. charging [CSWZ22, NM23, YWQ^{+}21]. chart [RMR^{+}22]. chat [ZPL21]. chat-response [ZPL21]. chatbot [VCS^{+}21]. Chatbots [HNS^{+}21, XZ22, KY23]. Chebyshev [NT23]. check [EVVR21]. check-in [EVVR21]. checker [SS23b, SA22c]. checking [IIK^{+}23, RF21, XCJ22, YSH^{+}22]. checkpoint [CZZ^{+}22, EBLM22, NKKM21]. checkpoint-rollback [NKKM21]. checkpoint/restart [CZZ^{+}22, EBLM22]. checkpointing [KPA23]. checkpoints [LFW20]. chemical [KA21a]. Chen [RÖ22]. Chennai [SM21]. chest [Boz22b, DSSS22, GP22b, LLM^{+}24, MRGP22, UAS^{+}22]. chi [RM21]. Chicken [PS23c, LZC21]. children [SX21, YY22]. China [ZWJG21]. Chinese [DWDG20, GLW22, WHJ^{+}20, WGZ^{+}20, ZLW^{+}21, ZFF^{+}21, ZBC^{+}21, ZLCL21, ZWL^{+}20]. Chip [KK22e, AY21, CC22b, Liq22, MPG20, SCP20, USP^{+}23, GBBS21, ZX21b]. chips [FHH^{+}20, ITO21]. chirplet [PSMM22]. choice [YJY^{+}21]. choices [KVP21]. CHOpinionMiner [WHJ^{+}20]. choreographies [SSCN23]. chromodynamics [AML^{+}22]. chronic [AYJ^{+}22, RV23, SCM22, VR22]. chunk [DZW23]. chunking [GGJ^{+}24, TLX22]. chunking-grouping [GGJ^{+}24]. churn [BF22]. cipher [BA20]. ciphertext [CDR^{+}23, HHYL22, SYL23, ZHDW24]. ciphertext-policy [ZHDW24]. circuit [CL22a]. circuits [CSD23, HLC^{+}21, RK21b, SCP20]. circular [ZTF^{+}20]. circumstance [TWL^{+}20]. cirrhosis [GKZ23]. Citation [GZC^{+}22, MBT^{+}20]. cities [AYH^{+}22, Ano21-40, BTDD20, DKL21, UZAA21, KKAM21, Kat23, KKE^{+}22, SJA^{+}22, ZUTK23]. citizens [DPS21]. City [ML20, YLKC24, AA23b, BS23b, CLDY21, KHEF22, KAP20, ZXL^{+}21a, LLA^{+}22]. Clang [WKB^{+}22]. Clang/Polly [WKB^{+}22]. Class [Par22, IK22, KFKD22, KSY24, LLO21, LHL^{+}20, NSSS22, SP23a, SKS22]. class-bridge-decomposable [LHL^{+}20]. Class-index [Par22]. classic [AKA22c]. classical [AAA22b, GSB21, UAS^{+}22]. Classification [AMV22, Alm22, BDR23, DLC^{+}21, GK23c, KA21a, KSK^{+}22, TA22b, AR22a, AAMAA22, Abb23, AV22a, AMJK21, Aka22b, ATS22, ASL20, AFKS23, ABCP23, ARS22b, AS20, AA23e, AMM^{+}20, AMA22, CO21, BP23a, BGG24, BMZ^{+}22, Ben21, CÇY22, CÇ22a, CKKK20, CMK22, CMJM22, CB22, CMLL21, DDZ^{+}20, DS23a, DPYS22, DS20b, DRR22, DCK21, FEH22, FMJ^{+}22, GDSS22, GDA^{+}21, GSTS22, GSB21, GKZ23, GU23, HAA^{+}21, HPCK22, HHXH20, IK22, JM23, JMV24, JR22, JSYAA20, JA23, JNS22, JYL^{+}23, KEK^{+}20, KH22, KE22, Kab23, KXL^{+}21, VK24, KDC22, KBBH21, KR22a, KSJ23a, KIAA^{+}22, KKC22, KY22, KMS^{+}22, KP21, KSSP22, KS22c, KAAR23, Kum22, KK23b, LJBS23, LLJ^{+}22a, LM22a, LDS^{+}23, LJB22, LSW^{+}20, LDH^{+}22, MQEK21, MMR^{+}22, MRK^{+}23, NSR22, ODK^{+}23, PBK23, PSP22, PKB22, PKKL21, PS23b, PVVS22, RR23b]. classification [RTBC23, RG22, RW24, SS21, SAQJ23, SS22a, SVD22a, SMBM24, SSN22, SRRM23, SG22c, SS22c, SR24, SSP22, SSP23, SSDY20, SSP24, SKS22, SI22, SD23, SLL22, SAMS23, SP21b, TWL^{+}20, TLL20, TSA21, TYA22, TB23, TSD23, UAS^{+}22, VY23, VLVS22, WZ20a, WAY^{+}21, WZY^{+}22, WZMJ23b, WWM^{+}23, WRE^{+}23, XGX^{+}21, XFH^{+}21, XLL^{+}23, YRV^{+}23, YMWA21, YSLX22, Yil22a, Yil21b, YOWY22, YMKH22, YRSO23, ZLCL21, ZLCS21, ZLL^{+}22, ZZQ^{+}22, ZLW^{+}20]. classification-based [Yil21b]. classification-deep [SR24]. classifier [ATS22, APM^{+}21, ARS22b, AY23, BJS23, BS23c, CSR24, CT22, Dah23, GMM22, Gün23b, JSAA22, KEK^{+}20, KSS21, KUK22, Kot23, KSTV21, LJBS23, MVR23, MSBR23, NBS^{+}22, PLP22, SHA^{+}22, SKP22, SBS22, Tur23, VPSM22, YY22]. classifier-based [BJS23]. classifiers [CMLL21, LHL^{+}20]. classify [KUK22, SSK23b, VR22]. Classifying [AMR^{+}21]. classroom [DB23, TTZX22, ZHT^{+}23]. classy [ASRN23]. clean [RK21a]. cleaning [LSZ24]. Clearance [TG23]. clickstream [OA22]. climate [LM20b, MCD^{+}23, SYRS^{+}22, WWG21, WKY22]. clinic [ZZW^{+}22]. clinical [Ano21h, Kan22, SAD^{+}21, YW22, ZLW^{+}21]. clique [DD24]. clock [BM22b, WHH^{+}23]. cloned [XYLW21]. Closed [MV22, IA23]. closed-form [IA23]. closeness [RTSK23]. closeness- [RTSK23]. Cloud [Agr21, AIA22, BZEM20, CCCR21, DMD^{+}20, DKA^{+}21, HT21, JC21b, LMGG20, LL21, MJAM23, NM24, Oh21, RSM21, SK21c, TC24, VB24, WRJ20, WGYZ22, XGZ^{+}20, ZYY^{+}24, AB22b, AM22, AR23, AMB23, AASPR22, AB22a, ABSS22, AFKS23, ASJ21, AA23b, AI21, AHS20, AS24, ARFA21, AB21, AYJ^{+}22, AFNH21, AYH20, AYB21, ASMK21, BC21, BV22b, BR21, BABLH21, BRS^{+}22, BVS20, BJS23, BCP^{+}23, BZL^{+}22, BN21, BPK24, CLLB20, CMT20, CdRNB23, CK23, CD24b, CH21, CSWC20, CWW23a, CDR^{+}23, CHLD23, CA22, CYDW20, DS24a, DP22a, DKL21, DE22, DYF20, DXXL20, DK22b, DWY^{+}21, DS22c, EI22, FZA22, FXX22, FKK23, GFQ20, GWGR20, GLN23, GN21a, GR22b, GP22a, GSG^{+}23, GWA^{+}24, GLRB21, GYL^{+}21, GSVS21, HAS^{+}22, HYK21, HMK23, HAR20, HXY20, HPCK22, HM20, HK22b, Hem22, HAK22, HTZ^{+}22, HLC^{+}21]. cloud [HAA23, JB24, JR24, JMY21, JW21, dCJAAdOD21, KMS^{+}21, KV22a, KS22a, KB21, KA23, KS20, KV22b, KQK^{+}20, KV24, KK23a, KABE^{+}20, KKP20, KJ21, KHR22, KP21, KSSP22, KLK23, KDS^{+}23, KS24, KSSK22, LA22, LMR22, LLAV22, LW20a, LZZ21, LYSC21, LLZ^{+}21b, LLW^{+}23, LCCT22, LLJR21, LSLY20, LLH^{+}20, LZW^{+}21, LPZ^{+}22, LXC^{+}22a, LXC^{+}22b, LWL^{+}23, LCM22, LLYZ23, MR23b, MSN22, MT21, MBM^{+}20b, MSSM24, MBB22, MHB22a, MG20, MG22, MK23, MMR21, MDSG^{+}23, MYCH22, MSK22, MHB^{+}22b, MRAM^{+}21, MV24, MA22, NA22a, NA22b, NSPdO21, NFF21, NN23, NAK^{+}22, NSKS21, NSK^{+}24, NGB23, NMT^{+}23, Ogi20c, Ogi20a, OO20, OT20, OS21a, OO21, OLF21, OBTC20, PPM^{+}20, PSP24, PK23b, PSB^{+}20, PST24, PBC^{+}23, PK22c, PKM21, PSK23, PI21, PS23c, dQtZWS22, RSM23, RK21a, RMS24, RBWB21, SB23a, SAM^{+}23, SM24a]. cloud [SSN21, SKB^{+}20, SKH^{+}21, SAC22, SHBC20, SEM^{+}20, SV22c, SKHL22, SS24, SM23a, SZY^{+}22, SJ23a, SSI22, SZJ21, SHVA23, SK21b, SR22d, SM22d, SCL^{+}20, TWL^{+}20, TDZ^{+}24, TTA20, TCW^{+}23, TKS22, TT23a, TT23b, VJ22, VK23a, VD21, VB22b, VKK^{+}24, WCWG20, WGY20, WBZ21, WWZ^{+}22, WZZ^{+}22, Wan22b, WKL^{+}22, WLLZ20, WNZ24, WWL^{+}20, WDW^{+}22, WCZ^{+}23, XZL^{+}20a, XCZ^{+}21, XDJY21, XA22, XZYW23, XYFZ23, YV23, YLW^{+}22, YJJ23, YZPL21, YR21, YSS^{+}21, YJY^{+}20, YZXL22, ZWO^{+}20, ZWW^{+}21, ZBC^{+}21, ZHJW21, ZLO^{+}21, ZWJG21, Zha22, ZOS^{+}21, ZFW23, ZS22, MC24]. Cloud-assisted [HT21, WGYZ22, CDR^{+}23]. Cloud-based [SK21c, XGZ^{+}20, AR23, AYJ^{+}22, CSWC20, DWY^{+}21, EI22, GLN23, HAA23, KQK^{+}20, LCM22, ZS22]. cloud-edge [Wan22b, XCZ^{+}21]. Cloud-edge-terminal [ZYY^{+}24]. cloud-fog [TT23b]. cloud-Internet [KB21]. cloudlet [TTA20]. clouds [BWTJ20, BF24, GSZ^{+}20, GFPGT21, KMZ^{+}20, LGA^{+}20, LXZ20, MSS22, MC24, NGOS22, NRMB23, PKKL21, RLdO20, RYG^{+}21, SK20b, WCSG20, YSLX22]. Cluster [PM23b, TS21, AAK^{+}21b, AD22b, CLC^{+}23, FPXM21, LZZ^{+}20, MS21c, PS22b, RJ22, RSS20, RZ21, RMS24, ST22a, SPS23, SS23d, SP23b, SM24c, UMR23, WWW^{+}23, ZZL^{+}20, ZHX^{+}21]. cluster-based [RJ22]. cluster-optimized [AD22b]. clustered [DDB^{+}21]. Clustering [KGGM22, AM22, AS22a, AP23, AN23, ANAMSAR21, AV23, ASRN23, ATG23, BYL20, CT21, CS22, CH21, DR23, DKH20, FYH^{+}21, FM20, GSG^{+}23, GK23a, Gui22, GBB22, HVB22, HYC^{+}24, HC22, JB22b, JSYAA20, JD23, JLE22, KUK22, KRJS22, KV24, KHY^{+}20, KDS22, LLAV22, LLW^{+}20, LZL^{+}20a, LWZY23, LWC^{+}21, LS23b, LYF^{+}23, LAC21, MLZ^{+}21a, MAR23b, MKA23, MS21c, MB22, NJ21, OHRS21, OBER22, PPA22, PS22b, PLP22, MR23a, RGKK21, RPPK23, RR22, RP21c, RSJ21, RSJ22, RF23, RZ23, SAPC21, SMM22, ST22b, SZZ^{+}21, SPK^{+}22a, SSH22, SCL^{+}20, TWW^{+}21, TNP21, VS22b, VRS22, VR22, WJL^{+}20, WYLL24, XLZ20, YYLL22, YSS^{+}21, YJY^{+}20, YZZA23, ZTF^{+}20, ZCL21, ZS22]. Clustering-based [KGGM22, AM22]. clusters [ASJ21, Ano21h, CdRNB23, CLYG22, FKO22, GK22a, OXBL23, RR22, SP21a, TBT^{+}21, ZFW23, NWT21]. CM3 [WWG21]. CNN [AT22, DR22b, DSS^{+}23, HK24, JM23, JR22, JAM24, NWZ^{+}21, ÖTT23, SKP23, SMBM24, SMR23, SD23, TA22b, VK23b, WHZL21, YLT^{+}21, ZYH^{+}23]. CNN-ANFIS-based [JM23]. CNN-based [ÖTT23]. CNN-BiLSTM [SMR23]. CNN-LSTM [NWZ^{+}21, SMBM24, JR22]. CNN-RNN [DSS^{+}23, VK23b]. CNN-SVM [TA22b]. CNN-wavelet [WHZL21]. CNNs [SZZ^{+}24a]. Co [MBT^{+}20, RS20, ACC^{+}23, CZG^{+}20, DJGF21, JDG22, NWX^{+}23, SZZ^{+}24a, SRIB23, Yu20, xZlGCzJ20]. co-attention [CZG^{+}20]. Co-citation [MBT^{+}20]. co-creation [ACC^{+}23]. co-evolutionary [DJGF21]. co-occurrence [SRIB23]. co-operative [JDG22]. co-optimized [SZZ^{+}24a]. Co-scheduling [RS20, Yu20]. co-training [NWX^{+}23]. coal [CZ21]. Coalition [MMS24, ACC^{+}20, QHL^{+}24, SMD22b]. coastal [GDA22, PMR^{+}21]. coaxiality [PHZ^{+}22a]. Code [LK22c, AKRR20, CL23, DK21, DB23, DKH20, GDCW22, HLZ23, HZW^{+}23, JKS20, KT23, KHPH20, KGE^{+}20, LPHK20, LLL^{+}21, MT22, OWB^{+}20, XAC^{+}20, XR21, ZWL^{+}23]. coded [Ano21-37]. codelet [PWJ^{+}21]. codes [Akb22, ALB^{+}20, FZ23, KGE^{+}20, NVV24, QNZP22]. codesign [MAW^{+}22]. coding [Sin23, MKL21]. coefficient [SAHAN22]. coffee [RTBC23]. cofferdam [WJLC21]. Cognitive [Ogi20c, Ogi20a, OO20, SDSW21, AMRH21, Ben21, Ben22, CYC21, CÇM22b, DRMA22, HV21, HH23, HL23, JPN21, LLW^{+}20, LG21, MSS^{+}20, OT20, OO21, Ogi21, Sao22, SX21, SS22f, TSB23, ZX23a, IVP^{+}23]. CogramDroid [BD21]. Cohen [TSLL21]. coherence [MPG20]. coherent [DSC^{+}21a, RPM22]. cohesion [LDZ^{+}22]. cold [ZWX21]. collaboration [IAO21, LGW^{+}22a, XFPY24]. Collaborative [WWG^{+}20, APL^{+}21, GCSZ24, LTLX22, MLZ^{+}20, NLB22, NMM21, NNJC23, NJ21, Wan24, XGCZ23, ZJ21, ZYY^{+}24]. collaboratively [DCZ^{+}22]. collection [HML21, LKR^{+}22, NGXZ21, ZWX21]. collections [FKO23, YCL^{+}22b, YÖT^{+}22]. collective [LFW20, MB21, SEM^{+}20]. collectives [DJJR22, HGX^{+}22, TRM^{+}24]. colliding [CA22]. collision [LSL20, SWK22, ZWC^{+}23]. collusion [AEMS24]. collusion-free [AEMS24]. colon [RR23b]. Colony [MT21, AD22b, JR22, JCL^{+}20, LZQ^{+}22, LYG^{+}21, RP21a, RP22, WJJ23, XWW^{+}21]. colony-based [JR22]. color [BCM22, CCGN20, Gul22, LDS^{+}23, OSK23, SO22, SBA22, SBB^{+}20, WN23, XWQ^{+}23]. coloring [XHZHXBQX22]. Column [AYG^{+}21, STISM21]. column-stores [STISM21]. columnar [IP20]. combat [KÖÖG22, MAAK23, MVR23, SJA^{+}22]. Combination [CL22b, GR22a, LLW^{+}20, DLC^{+}21, HHXH20, Kab22, YLS^{+}23]. combinational [CL22a, SCP20]. Combined [BYTG22, MT21, SP22b, BMZ^{+}22, KKP20, LM22b, PHZ^{+}22b, SSMT22a, SS23b, SBB21, WZ20a, ZWW^{+}21]. Combining [ATS22, AU22, CL23, EGGG23, MTD^{+}20, SYG22, CPQ^{+}22, EI22, LSSQ22, LFWJ22]. ComChain [VG20]. come [KLJ21]. comfort [HHP23]. Command [ZARR23]. commands [WEH^{+}22]. comments [XGX^{+}21]. commerce [WQS^{+}23, Zha22, Zhu22a]. commercial [NMT^{+}23, WNZ24]. commit [EAvM20]. commodity [WWW^{+}20]. common [Gha20]. Communication [GMA20, LHC^{+}23, ASC22, APV23, ACVK23, BDG^{+}23, CAS24, CJ21a, CBK23, DJJR22, DSS21, FK24, GJPS23, GRC^{+}23, IQS^{+}22, JP21, JGW20, LZZ21, LPW^{+}21, Liq22, MWS^{+}23, ORP21, PCK23, RS21, SA22b, SWK22, SS22e, TK22b, TAH22, TRM^{+}24, WK20, YLZ^{+}21, YBJ^{+}23, YL20c]. Communication-aware [GMA20]. Communication-efficient [LHC^{+}23]. communications [AZM20, ALB^{+}20, BXH^{+}23, JYW^{+}20, TTM^{+}22, XLXZ20]. communities [AJAA21, DGM21, KABE^{+}20]. Community [HFFA20, LHL^{+}22, Ano21-35, ABC^{+}21, BKD22, BVP22, GlSL^{+}23, KMR22, KLJ21, LSS^{+}21, LSSQ22, MCNR20, RD23, SPJI^{+}21, WZG^{+}21, WCWG21, ZZY22a]. community-discovery [LSSQ22]. Compact [LM21]. compaction [CO21]. companies [Ano21c]. Comparative [LMM23, SKCS23, SKÇA23, ZWT22, AK21, CMLL21, JK22a, KPP^{+}22, OE22b, RZVC21]. comparator [Jia22]. Comparing [CYK^{+}21, EK20]. Comparison [ÇGB23, KÖÖG22, KK22a, Kay22b, MT22, PKM21, DHSG23, ECIB20, GDA^{+}21, GKZ23, JNMG21, KE22, KB22, Mar23a, PS21]. comparisons [NMT^{+}23]. compass [CL22b]. compensation [LWZC21]. competent [RJ22]. competition [YJY^{+}21]. competitive [ABCP23, LGL^{+}21]. Compilation [RBC20, GDGN23]. Compiler [ZYPD24, CZZ^{+}22, dRdSZ^{+}23]. Compiler-driven [ZYPD24]. compilers [dRdSC^{+}21]. complementary [MLZ^{+}21a]. Complete [ATG23]. completion [LLX^{+}24, OBER22, ZXLD21]. Complex [WHZL21, ZLD22, BKD22, Boz22a, CS23, FD20b, HL23, JT23, KSS24a, LLW^{+}22a, MCNR20, SS23a, VKK^{+}24, WMC21, uZKH^{+}20]. complexity [Ano21-37, FSFM22, KK21b]. compliance [AADS21, BFR24]. Component [ECIB20, AAK^{+}21a, AU22, BKM^{+}24, EHST21, wGTC22, HNG22, HML20, dCJAAdOD21, KÖÖG22, LLH19, RR23b, WY20, YYLL22, Yil21c]. component-oriented [dCJAAdOD21]. composable [MPG20]. Composite [PZD^{+}21, YSL^{+}21b, SS23a]. composition [BJGF20, DJGF21, HPCK22, JQGL20, KQK^{+}20, KEMZ22, LXKW23, WRJ20, XHZ^{+}21, YFF22]. Compound [YHL^{+}21, CCZM23]. comprehension [YSK22]. Comprehensive [SK22b, ASA^{+}21, GKG^{+}20, Haj20, HJT^{+}20, KK23b, LK22b, LXC^{+}22b, PKK21, THT20, WZ24, WGDP23, ZX21a]. compressed [FKGO22, SZZ^{+}24a, XWD^{+}22, ZX23b, ZWZ^{+}22]. Compression [AAG^{+}22, CCD^{+}20a, AD22c, CWH^{+}24, EBDB22, Gui22, IAASK23, KMR22, KSVP22, OHRS21, PRS23, PMP23, PSF23, SRG^{+}21, WFY^{+}22a, WZX^{+}22, YCL^{+}22b]. Compression-based [CCD^{+}20a]. Compressive [Kha22a, LGM21, XFH^{+}21, YLGY20]. Computation [Ano21-41, LAK22, XZ20b, AMBAJ22, AA22b, CO21, BMK^{+}20, BPAE20, CC22b, DBS^{+}22, DJJR22, EVVR21, FZ23, KML21, LJZ21, MGB^{+}23, Ogi20a, OO22, QLL^{+}22, SG22d, TSLL21, WDW^{+}22, XWC^{+}22, XGCZ23, YZ21]. computation/communication [DJJR22]. Computational [AE22a, JB20, BLB^{+}24, EVVR21, FZ23, FSFM22, GN21a, GCS20, Jeo20, JC21a, JJZ^{+}21, KAQ23, MKBB22, NA22c, NdSSSN20, NMT^{+}23, RD23, SLLMTS^{+}24, SMAG22, VSK22, WLX21, XR21, YSH^{+}22]. computationally [SMD^{+}21]. computations [CMY21, QMC^{+}20, RBC20, SK20b]. Compute [GTM24, IGYS23, MDSG^{+}23, ÖGS22, TNI23]. computed [BS23c, DSSS22, WPL20]. Computer [GS22, JS23, MSA21, dCMM21, BMZ^{+}22, Gui22, JK22b, JKKL21, LCW21, MCT22a, MCT22b, YKL23]. computer-assisted [JKKL21]. computers [DS22b]. Computing [BI23, GS21, HMQO23, JSLL20, KSY23, LM20a, RSM21, TC22, Vin21, VS22c, AB22b, AA21, AIA22, AB22a, AFKS23, AEMS24, ACC^{+}20, AHS20, AB21, AFNH21, AYH20, Bad23, BZEM20, BV22b, BPW^{+}20, BP24, Ben21, BBB^{+}20b, BVS20, BTDJ21, BWS^{+}21, CdRNB23, CMS21, CMS24, CHMC21, CLT^{+}21, CJC22, CPH20, Cud20, CYDW20, DDB^{+}21, DDUK23, DA22, DKB20, DE22, DYF20, FSA^{+}22, FEK20, FGZC23, FZT22, FKY^{+}23, GWGR20, GAS23, GR22b, GFPGT21, GCS20, GSG^{+}23, GMSM21, GWA^{+}23, GN21b, HSR23, HAS^{+}22, HMR24, HYK21, HGW^{+}23, HAK22, HNSS22, HLC^{+}21, HZQ^{+}24, IS23, JB22b, JK22b, JSA^{+}20, JR24, JZL22, dCJAAdOD21, KS22a, KM23a, KK22a, KKK21, KKS22, KFML20, KML23, KAP20, KHPH20, KJ21, KCL^{+}20, KT22b, KP21, KSSP22, KJ20, LA22, LXZ20, LZZ21, LYSC21, LGL^{+}22, LXW^{+}23, LSZ^{+}23, LYBZ23, LWW23, LXC^{+}22b]. computing [LBZ^{+}22, LDCD22, LWL^{+}23, LDW^{+}23, Liu24b, LG21, MSN22, MBM^{+}20b, MSSM24, MHB22a, MK23, MG21a, MHB^{+}22b, MdARS^{+}23, MRAM^{+}21, MGN^{+}22, NA22b, NAK^{+}22, NSKS21, NMS^{+}21, Ogi20c, Ogi20a, OO20, OT20, OS21a, OLF21, PYC^{+}20, PJP21, PSK23, QLL^{+}22, RK21a, RBWB21, RZVC21, SB23a, SZI^{+}23, SKB^{+}20, SPQM20, SHBC20, SN22a, SEM^{+}20, ST22c, SM23a, SSW^{+}22, SKSN22, SPS22, SPKK22, SYRP22, SG22d, SD22, SHV24, SCL^{+}20, SWS^{+}24, TWL^{+}20, TWG^{+}21, TS24a, TSR22, TCW^{+}23, TS24b, TLS22, TS22, UK23, VY23, VS21, VCBB20, VG21, WCWG20, WGY20, WLLZ20, WNZ24, WCCC20, Wri22, WS24, XWC^{+}22, yXlLyGX21, XCZ^{+}21, XDJY21, XCG^{+}22, XA22, YLZ^{+}21, ZWW^{+}21, ZBC^{+}21, ZHJW21, ZLTX21, ZCN22, ZYH^{+}23, Zha22, ZGH^{+}22, dCMM21, dCMA23]. computing-based [KAP20]. concentration [WSL^{+}20]. Concept [KTK20, AYG^{+}21, CH21, NMS^{+}21, RSR^{+}22, YW22, ZNLL22]. concepts [DJF21, WHH^{+}20]. conceptual [MJAM23, ÖK22]. concrete [Kha22a]. Concurrency [Ano21-41, CY22, FSWW21, Gün23a, JJZ^{+}21, WHZL21, XZ20b]. concurrent [BJWY20, CY24, GB20, LALMGLG20, SNGK21, SZL^{+}22, VNP^{+}23, WCCC20, ZYPD24, dSNdL^{+}23]. concurrently [SLC20]. condition [BTT21, WJLC21]. Conditional [NGD^{+}22, ZFF^{+}21, BKK23, GLW21, LLL^{+}24, ÖK22, PA21, ZCL21]. conditions [YLZ^{+}21]. Conference [HFFA20, KSY23, SKE22, TC22]. confidence [LLMX21, WBL22]. confidence-based [LLMX21]. configurable [BBB^{+}20a, KYZ20]. Configuration [BMSD23, BP20, LYF22, WLDW22]. configurations [BC23b]. configuring [WZC^{+}22b]. conflict [JZB20, ZJ21]. conflict-free [ZJ21]. conformity [ZWCS20]. Congestion [NLG^{+}20, BS23b, CMJM22, LLW^{+}22b, ZXL^{+}21a]. Congestion-aware [NLG^{+}20]. congruence [CB22]. Connected [RT21, KMS^{+}22, MIN^{+}23, NNVD22, SC22b, VRR^{+}22, Yil22b]. connection [YMWA21]. conscious [LDCD22, ZLTX21]. consecutive [BB23]. Consensus [CF21, AB20b, HLCH20, JB21, SKV22, WMQ^{+}22]. Consensus-based [CF21]. conservation [HMK24, JKP22]. conserving [LGM21]. Considering [MSL22, KFKD22, LWZ^{+}22b, QCC^{+}23, SSSP21, YWQ^{+}21, ZWCS20, ZYXX23]. consistency [TP20, WCWG20, YMZ^{+}20]. Consistent [SLL22, BOI23, PSK23, YLZ20]. consolidated [HTZ^{+}22]. consolidation [BRS^{+}22, LWCZ22, PST24, PKM21, WCZ^{+}23]. consortium [JB21, LZZ^{+}23, WMQ^{+}22, ZHDW24]. constellation [Man21]. constrained [Ano21e, ASMK21, BZL^{+}22, CDV23, DCM21, GFA21, GSVS21, HXY^{+}22a, JXL^{+}23, RYG^{+}21, USI21, WFY22b]. constraint [AZM20, CZL^{+}22, LLH^{+}20, LLMZ21, MMR21, MSS22, NR22, WZ21, ZGWZ23, ZXLD21]. constraint-dividing [CZL^{+}22]. constraints [DPdS^{+}23, FJCFB23, JR23a, PLZL24, RS20, SSDY20, SWS^{+}24, TWQ^{+}21, USP^{+}23, XYFZ23]. constructing [DS24b, KKM21, ZHX^{+}21]. Construction [JCL^{+}20, Liu21, WZ20b, DPBP24, FSWW21, GKLS23, HNSS22, LPHK20, SSCN23, WJLC21, Yib22a, YLZ20]. consumed [RP22]. consumers [LLSC22]. consumption [ADGT22, ASC22, DÖD22, DTK24, FPÁ^{+}20, GSG20, KSA22, KK23a, LWZY23, MTT20, PKK23, PB22b, SZZ^{+}22, SZJ21, SM24c, Wan24, YYLL22, YFdP^{+}23]. contact [CZCM23]. container [ASJ21, CI20, IAA20, MV24, OXBL23, PST24, SAM^{+}23, SP21a, VB22b, YPY21, YJJ23, ZHWY22]. container-based [SP21a, YJJ23, ZHWY22]. containerized [HAA23, JGJ^{+}21]. containers [MGS^{+}20, MK23, SLC20]. containing [HZW^{+}23]. containment [SM24a]. Contemporary [MTSU22, RDB22]. Content [BCM22, PR22, TLQ21, Aru22, BKLY20, GK22a, IAO21, RRB22, SM22a, SKS20, YOWY22, YPO21, ZQX^{+}23]. Content-based [BCM22, PR22, TLQ21, SM22a, YPO21, ZQX^{+}23]. contention [SCD^{+}23]. contents [HKMS21, PSHJ20]. Context [BC21, BABLH21, DRR22, GJBM22, KHHK21, SAQJ23, UK23, Vin21, BC23a, BTT21, BJWY20, UZAA21, HD23, LMGG20, NSKS21, PYC^{+}20, PMC^{+}21, XXD^{+}22, YHOY22]. Context-Aware [Vin21, BC21, BABLH21, GJBM22, KHHK21, SAQJ23, UK23, BC23a, UZAA21, LMGG20, NSKS21, PYC^{+}20, YHOY22]. Contextual [dCJAAdOD21, uHAU23, SGH23]. continuity [ZXLD21]. Continuous [FKK23, KML23, LHK22, LZL^{+}23, RAN22, SMAG22, ZHXY23]. continuum [MDSG^{+}23]. contour [BTT21, LW20b]. contours [WJL^{+}20]. contract [BS23a, GKLS23, HLCH20, NAR^{+}22, SK22d, TG23, XFPY24]. contracts [Ano21i, BZK^{+}21, ÇKÇ22, dCJAAdOD21, SB21, TY22, ZOS^{+}21]. contrast [SBA22, WLJ20]. contributing [RS24]. contribution [SAC22]. contributions [BTP^{+}21]. Control [AZA20, GLM^{+}22, KBL^{+}21, LLW^{+}22b, LGW^{+}22b, MV22, Agr21, AAD20, ALR22, BZT^{+}23a, BWS^{+}21, CLLB20, CRB23, CY22, CCZ^{+}21, CYZ^{+}23, CG22, ÇGB23, DL23, EABZB21, EASN22, EI22, FKK23, GHL^{+}23, HAAF22, JR24, dCJBP20, KM23a, KKP20, KHHK21, KCP23, LZL^{+}20b, LWL^{+}23, NET20, NET21a, NET21b, PS22c, RMR^{+}22, RSMCP22, RKL21, SS23e, SV22c, SLJ23, SCS^{+}21, WEH^{+}22, XSGL20, XZW^{+}20, XL21, YYZS22, ZAB22, ZTMC22, ZTMC23, ZWC^{+}23]. ControlChain [AAD20]. controlled [BZGM22, MSS^{+}20]. controller [BZWH21, FJCFB23, HTZT23, IQS^{+}22, KT22a, SKK22, XZXV21, XCJ22]. controllers [HJT^{+}20]. Controlling [SPK^{+}22a]. conundrum [RS24]. conventional [IG24]. Convergence [AYB21, SPQM20, DWZ20a, LLT21]. conversion [LHWT20, ZPL21]. convex [CF21, QMC^{+}20, XL21]. ConvLSTM [SGS^{+}23, WLZ^{+}21b]. Convnet [VS21]. Convolution [RT21, AMV22, CZL^{+}22, GDSS22, HXST22, IK22, Kay22b, PKB22, RKR22, VRS22, ZYX^{+}21, ZDL^{+}22]. Convolutional [AJNS22, BRNR23, CBFS23, JNS22, ZLZ^{+}22a, AMB23, ATS22, AT22, AFM22, BKJ22, Bei24, CÇ22a, DRM22a, DBD23, DCWM20, DM21, GPDB20, GSTS22, GM22b, HLC23, JAM24, JPK22, JA23, JPO^{+}21, KTM22, KE21, Kab22, Kab23, Kar22, KR22a, KK22b, KBR24, Kim21b, KMD23, KMS^{+}22, Kot23, KSK23, LKR^{+}22, LLJ^{+}22a, LN20, LZF20, LZC^{+}20, LGW^{+}22a, LCL^{+}20, LL23, LWZ^{+}20, MSJ22, MWH^{+}23, MIN^{+}23, Mis22, MCT22b, NMQ22, OKJ^{+}21, OE22a, PBK23, PT22, PK22a, PD22b, AST22, RRB22, RMMK23, RK23b, RSKA23, SJ22, SSMT22a, SSMT22b, SC22a, SMD22a, SKP22, SBA22, SMBM24, SZZ^{+}24a, SZqWZ20, SSN23, SMKA22, SPHP21, SLHW20, SS22g, TZ22, TÖ21, VRR^{+}22, WSM^{+}20, WWJ^{+}20, WSJ^{+}21, XGX^{+}21, XCX^{+}20, XY21, YS22, YTW^{+}24, YW22]. Conway [SAB22]. Conway-Maxwell [SAB22]. cooling [LM22b]. cooperation [CCL^{+}22]. Cooperative [GPR^{+}22, ANP^{+}20, GR22b, JP21, NKKM21, SSW^{+}22, WCL^{+}24, WBZ21, ZX23a]. coordinate [BYTG22, LFW20]. Coordinated [WXLD21]. coordination [DWZ^{+}22, FGJ^{+}21, KMY24, RAN21]. coordinators [AEMS24]. copyright [SA22d]. core [AV21, AÜ23, CL22a, FKO22, HMQO23, HNSS22, Hua20, IGYS23, LCKJ21, LSSQ22, LYI^{+}20, LB21, EGGG23, YCY20, ZZL^{+}22, ZNLL22]. corner [WSL^{+}20]. corona [JS22b]. coronavirus [ZCR23]. corpus [Par22, RSJ21]. corpus-index [Par22]. correction [WHC^{+}20, ZWCC23]. Correlated [ZZLZ22a, OSK23, OAA22]. correlation [CH21, CWL^{+}21, CWL^{+}22, CYZX23, DDZ^{+}20, HMK23, LC20, SS22a, SSP24, SRIB23, TI22, WLY^{+}22, XYY24, XLZZ22, YSL^{+}21b]. correlation-aware [HMK23]. correlation-based [LC20]. correlations [LLY^{+}23]. correlative [MS22a]. correntropy [LLMZ21]. corridor [MK22a]. cosine [FLG^{+}22, Gul22, ZCW^{+}23]. Cost [BPK24, JHZ20, LCZY20, SK21b, WCSG20, ZTY^{+}22, BP24, BZL^{+}22, BN21, CPLX21, GWGR20, HYG^{+}23, KRKM22, KÖ22, KSB23, LYF22, LXW^{+}23, LSZ24, MSS22, OCD22, OO23a, RSM21, RSM22, RSM23, SKA23, SP22c, SR20b, WFY^{+}22a, WL23, WDW^{+}22, XHZ^{+}21, Zha22]. Cost- [BPK24]. Cost-aware [SK21b]. cost-based [BP24, GWGR20]. cost-effective [SKA23]. Cost-effectively [ZTY^{+}22]. Cost-efficiency [JHZ20]. cost-efficient [BN21, LXW^{+}23, WFY^{+}22a]. cost-energy [BZL^{+}22]. cost-oriented [WDW^{+}22]. cost-sensitive [LSZ24]. costs [NMT^{+}23]. count [JOW23]. counter [AG22, KA22a, NJK22, PD23]. counterfeiting [AB20b]. countermeasures [GQ21, SKK22]. counting [SZZ^{+}24b]. countrie [SR20b]. countries [HNG22]. country [TG23]. counts [KML21]. coupled [CLLB20, FM22, TB23, VPSM22, WXLD21]. CoV [NMT^{+}23]. Coverage [SP23b, AZI20, ATG23, DPFC20, GDCW22, URK^{+}22, ZGH^{+}22]. coverage-based [AZI20]. coverage-guided [GDCW22]. COVID [JS22b, Ele22, ATS22, Asl22, Boz22b, Cek22, CPYC21, CBK23, DÖD22, DR23, DSSS22, EUYY22, GP22b, JD22, Kab22, KKKS23, LLM^{+}24, MRGP22, MSA22, ÖUG22, PKVS21, SVS22, SAAAA22, SP22b, UAS^{+}22, ZARR23, ZSS22, ARS22b, PS22c, SJA^{+}22, UYÖ^{+}22]. COVID-19 [JS22b, Ele22, ATS22, Asl22, Boz22b, Cek22, CPYC21, CBK23, DÖD22, DR23, DSSS22, EUYY22, GP22b, JD22, Kab22, KKKS23, LLM^{+}24, MRGP22, MSA22, ÖUG22, PKVS21, SVS22, SAAAA22, SP22b, UAS^{+}22, ZARR23, ZSS22, ARS22b, PS22c, SJA^{+}22, UYÖ^{+}22]. Covid19 [Alm22]. Covid19MPD [Alm22]. CovidBChain [PS22c]. coyote [CYZ^{+}23, DBD22, MRKY22]. CPP11sort [LS22]. CPU [CS21, FD22, GDGN23, HSO^{+}21, HMB22, HVB22, HNSS22, JNMG21, KSKS22, LMM23, NdSSSN20, PRPD21, PSF23, RBC20, RT22c, SD22, SSPG20, WWL^{+}20, XAC^{+}20, Yu20]. CPU-FPGA [JNMG21]. CPU-GPU [FD22, HSO^{+}21, SD22]. CPU/GPU [LMM23, NdSSSN20, RBC20]. CPUs [CTA^{+}23, CHI^{+}23, IGYS23, KS21b, Yil23, ZDL^{+}22]. CRAC [CZZ^{+}22]. crack [WLL^{+}21b]. crafted [PR22]. crawler [JQ22]. crawling [GD22, KTU^{+}21]. Cray [BBB^{+}20c, HPH^{+}20, KMRR20, MGS^{+}20, RMS24, RT20, SSSR20, VBP^{+}24]. Crays [BRL^{+}20]. Creating [KGK22]. creation [ACC^{+}23]. credentials [BPT^{+}23]. credibility [LF23]. credit [GCS23, GDS23, MNR^{+}22]. Cricket [EBLM22]. Crime [KR22b, RNRK22]. crimes [JSS22]. criteria [DRR22, GK22b, NAK^{+}22, PKC^{+}20, SM24c, ZWCC23]. critical [AAB23, KK21a, Ker22, KMZ^{+}20, LLTZ24, MYCH22, OSMM23, RK23c, ST23c, ZLW^{+}20]. Crop [DM22d, JR23b, SHA^{+}22, SM22c, SM23b, SM23c]. crops [SP22d]. CROSA [BABLH21]. Cross [SSP22, SSP23, ZCZC23, GSS^{+}23a, LLZ^{+}20, LLMZ21, RK21a, SBGC21, SL20, TI22, UGK^{+}22, WLZ21a, WYL^{+}22, YWC^{+}23, ZZC24]. cross-architectural [ZZC24]. cross-blockchain [WYL^{+}22]. cross-correlation [TI22]. cross-correntropy [LLMZ21]. cross-device [SBGC21]. Cross-domain [SSP22, SSP23, LLZ^{+}20]. cross-layer [SL20, YWC^{+}23]. cross-lingual [GSS^{+}23a]. Cross-modal [ZCZC23]. cross-platform [UGK^{+}22]. cross-similarity-based [WLZ21a]. cross-view [RK21a]. crossbar [SM24b]. crossover [CZL^{+}22, GK23a, SS22a, YBJ^{+}23]. Crossover-SGD [YBJ^{+}23]. Crow [LST22]. crowd [SZZ^{+}24b, ZYXX23]. crowdsensing [JZC^{+}23, PGL^{+}23, ZYZC23]. crowdsourced [ARHT21, AYD21]. crowdsourcing [AS23c, RAG21, SG22a, XCZ^{+}21]. cry [JYC^{+}21]. cryptanalysis [FK24]. cryptocurrencies [FD20a, GDA^{+}21, LK23]. cryptocurrency [Ano21b]. Cryptographic [SV22d, AIA22, BKGC23, SV22a, TCW^{+}23]. cryptography [AMAT22, DK21, HK22b, HZZ^{+}23, HAAF22, LMR22, Ogi20c, PSP24, RKMS24, SDR23, SSS23a, VB22a, ZWT22]. cryptography-based [HZZ^{+}23, VB22a]. cryptosystem [DK21, GAH^{+}22, MS21a, PM23a, PA21]. CSFPA [SKHL22]. CSP [XZXV21]. CSS [CS22]. CT [DAK22, KKKS23, VRR^{+}22, Yl20b, ZX21a, ZZW^{+}22]. cubes [FHH^{+}20]. CUBIC [AAN^{+}21]. cuckoo [AI21, BMcKGK22, CNG^{+}20, DS20b, HR22a, RW24]. Cuculidae [RG22]. CUDA [Cza23, EBLM22, JPL22, LSS24, MNYN21, TO22]. CudaCHPre2D [QMC^{+}20]. cultivars [AA24]. Cultural [HKMS21, CCdCC21, WN21, WN23]. culture [JQ22]. cum [GLA^{+}22]. cumulative [JS22c]. cup [WLJ20]. Curation [BSML21]. current [CY24, CPH20, KK23b]. curvature [FMJ^{+}22, MGB^{+}23]. curve [AMAT22, HK22b, LCZ^{+}20b, NR22, PSP24, PSB^{+}20, PA21, SSS23a, VB22a, WGYZ22]. curve-based [PSB^{+}20]. CURVELET [EKS20]. Curves [SV22d, AE22a, Bul22, ZWT22]. Customer [COZ21, BF22, KY23, PLX20]. Customer-oriented [COZ21]. customizable [SY23]. customized [CPCK23, GCF^{+}20, Gün23a]. Customizing [PMS^{+}21]. Customs [TG23].

Cut [SKP23]. Cutting [LW20a]. Cyber [CZTC22, OSMM23, XZ20b, AR23, DAT23, DCT^{+}23, GKG^{+}20, HXZS23, Kab23, KGRG24, Ker22, LL23, PCG^{+}21, SS22c, SD23, SZ24, TS24b, Tur23, VRS22, XZ20a, ZX23a, ZALM23]. cyber-attack [SD23]. cyber-infrastructure [SZ24]. cyber-physical [AR23, DAT23, DCT^{+}23, GKG^{+}20, HXZS23, KGRG24, Ker22, LL23, SS22c, TS24b, ZX23a, ZALM23]. Cyberbullying [LWZ^{+}20]. cyberinfrastructure [KLJ21, ZABT^{+}20]. cybernetic [Boz22a]. cybersecurity [QZB^{+}23, vdSTC21]. Cycle [PSK23, BM22b, Boc21, DP22b, GK23a, KA22b, KS22c, LG23]. Cycle-consistent [PSK23]. cycled [GJBM22]. Cyclic [SV22d]. Czechia [PKVS21].

D [LHWT20, QMC^{+}20, AJNS22, CSL21, FAM22, GBB22, HSL^{+}22, HWY^{+}23, RJH^{+}23, TLH^{+}22, YYM^{+}24, ZQW^{+}21, ZZW^{+}22, ZKZ^{+}23, ZX23b]. D2D [CAS24, XLXZ20]. D4Science [ACC^{+}23]. DAG [PB22a]. Daily [GA22, NWZ^{+}21, XY21]. damage [XGZ^{+}20]. damaged [Li21]. damages [GNMELC21]. damping [YYZS22]. dark [SAF^{+}23, WZSZ20]. dark-silicon-aware [SAF^{+}23]. Darknet [SI22]. Dask [DHSG23]. Data [AMD20, DDB^{+}21, KBJ21, LFG^{+}22, LWCM21, NGOS22, PB22a, PK22d, SAAAA22, SZZ^{+}22, TAT22, TCW^{+}23, TAH22, VZR^{+}21, ABESEh20, AAMM23, AV22a, AKA^{+}22a, ARHT21, AAN^{+}21, AA21, AN23, ABSS22, ASAAAA22, AlS21, ASJ21, Ano21c, Ano21j, ASJ24, ARFA21, AJAA21, ARS22b, Aru22, ASRN23, ACJ21b, AYB21, BM22a, BYPO21, BBF21, BBD23, BZEM20, BPT^{+}23, BKM^{+}21, BVS20, BBP^{+}24, BZT23b, CD24a, CSL20, ÇKÇ22, CZUPA23, CSV22, CS22, Çet23, CK23, CJ21b, CH21, CSWC20, COZ21, CHMC21, CWW23a, CXXW24, CL22a, CB22, CPH20, CJP^{+}21, DBD22, DM22b, DSC^{+}21a, DCK21, DA22, DKC^{+}21, DPS21, DCT^{+}23, DKK24, DYF20, DLY^{+}23, EGB21, EUYY22, EAS23, FSFM22, Faz23, FDY21, GDA22, GJPS23, GRL^{+}22, GlSL^{+}23, Gha20, GBB22, GHT20, GK23b, GKAO20, GK23c, Haj20, HMK23, HGW^{+}23, HNS^{+}21, HLL^{+}21, HK22b, HC22, HAR24]. data [HD23, HR22b, HTVL22, HKP21, IPRS21, JSS22, JAM24, JS22a, JAC^{+}21, JS23, JQ22, JW21, JWT^{+}20, JSZS22, JR23b, JLE22, KUK22, KH22, KT20, KGW^{+}20, KKR23, KKAM21, KVV20, KGRG24, KLJ21, KFML20, KM23b, KM23c, KMD23, KK22d, KMZ^{+}20, KG22, KS24, KTK20, LA22, LMR22, LP21, LCKJ21, LMBK23, LLJ^{+}20, LFX^{+}20, LXZ20, LJZ21, LSD21, LH21, LXZ22, LLJR21, LX22, LZY^{+}20, LWHW22, LZM^{+}23, LWL^{+}23, LZLZ23, LAC21, LYI^{+}20, MT21, ML20, MCD^{+}23, MAW^{+}22, MDSG^{+}23, MG21a, MRAS^{+}23, MS23, MBO^{+}21, MFA^{+}21, MLC^{+}21, NVV24, NSPdO21, NFF21, NGXZ21, Ogi20c, Ogi20a, OT20, OLF21, OS21b, OHFF20, OA22, OGA^{+}22, OMA^{+}23, OHRS21, PD20, PS22a, PS22b, PM23a, PLP22, PYC^{+}20, PMC^{+}21, Par22, PSB^{+}20, MR23a, PS22c, PRS23, PZZ^{+}23, PL21, PKKL21, PKVS21, PMP23, PD22b, PKM21, PHDS22, PVRM22, PB22b]. data [PSF23, PKS22, RC22, RSR^{+}22, RSS20, RN22, RCYuRH21, RS20, RP21c, RW24, RNRK22, RSJ22, RYG^{+}21, SR22a, SNET21, SM24a, SSSR20, SK22a, SCKU23, SS23d, SK21a, SS23e, SV21, SVD22a, SMD^{+}21, SM22b, SSN22, SK22b, SZW^{+}22, SYL23, SLD24, SCH^{+}24, SZS20, SZZ^{+}21, SET^{+}22, SSM^{+}21, SKS^{+}23, SSK23b, SSDY20, SPHP21, SK22d, SRG^{+}21, SP21b, THW21, TWQ^{+}21, TTB^{+}22, TBT^{+}21, TLXX21, TY22, TZ23, TL21, TPT^{+}22, TD21, UKS22, UYÖ^{+}22, UGK^{+}22, VCFZ20, VRS22, VSS23, VK23a, VPB^{+}23, VR22, VNP^{+}23, VG21, WBZ21, WLZ^{+}21b, WMC21, WLL21a, WQY^{+}22, WZMJ23a, WCCC20, WZC^{+}22b, WZB21, WLDW22, Xia20, XLZZ22, XGZ^{+}20, XFH^{+}21, YLZT23, YA22a, YÇC22, YLGY20, YLJ22, Yl20b, YPO21, YR21, YSS^{+}21, YTN^{+}20, YLW^{+}21, YZXL22, YZZA23, ZABT^{+}20, ZJ21, ZHX^{+}23, ZHJ20, ZWW^{+}21, ZZLZ22a, ZZQ^{+}22, ZML^{+}23, ZYZC23, ZX23b, ZYPD24, ZX21b, ZWJG21]. data [ZWT22, ZLW^{+}23, ZLW^{+}20, ZCL21, ZWX21, ZLV22, Zhu22a, ZS22, AFK^{+}22, MBM^{+}20b, TD21]. data-aggregation [JS22a]. Data-aware [NGOS22]. Data-driven [KBJ21, SAAAA22, CJP^{+}21]. data-intensive [LXZ20, NSPdO21]. data-Mapping [MCD^{+}23]. Data-parallel [PB22a, TBT^{+}21]. data-placement [BYPO21]. data-prediction [JS22a]. databank [AA24]. database [AET^{+}22, APM22, AQP^{+}22, AJS23, DK22b, OCD22, OO23a, OAS23, SMD22a, VM23, YF20]. databases [BANT20, HYG^{+}23, LM21, RAaB21, XZD^{+}21]. datacenter [LLZ^{+}21b]. datacenters [ETH^{+}23, JC21b, KK23a, LLH^{+}20, YLW^{+}22]. dataflow [BCP^{+}23, LZZ21, MdAA^{+}21, WPK^{+}22, WGDP23]. dataflows [OCR^{+}23]. Datalog [JSZS22]. DATAM [LSW21]. DataPlane [CCBA23]. DataPlane-ML [CCBA23]. Dataset [Alm22, MLC^{+}21, AAMAA22, EA22, ÖSTY22, TPT^{+}22]. datasets [ARS22a, BD22, CU22, Kal22, KSY24, KKM20, LLL^{+}24, MFA^{+}21, SKSP23]. date [GGS^{+}22]. day [BMZ^{+}22, CZTC22]. DBA [OAS23]. DBSDS [ZML^{+}23]. DC [CDV23]. DCCA [xZlGCzJ20]. DCNMIX [CHL^{+}24]. DCNN [SMM22]. DCNN-SFO [SMM22]. DCT [KKS23]. DDL [YSL^{+}21a]. DDL-QoS [YSL^{+}21a]. DDoS [KKP24, FK20, GYZ^{+}20, HK24, HE23, JW22, KKP24, KHY^{+}20, PA23, SNBAI24, SSS^{+}20, VD21]. de-identification [XLY^{+}23]. DEA [ZST^{+}23]. Deadline [BZL^{+}22, CDV23, GSVS21, HAS^{+}22, JR23a, MMR21, MSS22, RYG^{+}21, TWQ^{+}21, XYFZ23]. Deadline-constrained [BZL^{+}22]. deadline-constraint [MMR21]. deadlock [GB20]. Dealing [ARHT21]. deaths [Cek22]. DEC [LLYZ23]. decay [LWZ^{+}22b]. deceiving [HAA23]. Decentralized [Ano21-41, DGP20, ALR22, AB20b, FEK20, HPS23, LFX^{+}20, LWL^{+}23, DPD^{+}22, ST22c, ZCN22]. Decision [Ano21c, TS24a, Ano21h, BM22c, BBB22, CYZ^{+}21, CMA^{+}21, DRR22, DCK21, FSdP^{+}23, KPJ^{+}21, Kan22, KGM23, KC22, KS23b, KA22c, LLX^{+}21, LCW23b, MMR21, MKBB22, MTY21, Ogi20b, SAD^{+}21, SDSW21, THW21, WWW^{+}20, ZARR23]. decision-making [LLX^{+}21, Ogi20b]. decisions [IAA20, MSS^{+}20]. decoder [YOWY22]. decoding [Ano21-37, SSP23, TNIK23]. decoding-enhanced [SSP23]. decomposable [LHL^{+}20]. decomposition [CD24a, FQD^{+}23, HGNN22, KDL20, KSP21, PP21, SG22e, WZZ^{+}22]. decomposition-based [HGNN22, WZZ^{+}22]. decoupled [ZSC^{+}21]. decreasing [GC20, SDSW21]. dedicated [SP23a, SWK22]. deduction [PL21]. Dedup [TLXX21]. deduplicated [ZWZ^{+}22]. Deduplication [PL21, SR22d, TLXX21, WFY^{+}22a, XZL^{+}20a]. dedusting [HLW^{+}21]. Deep [AC22, AV23, Bai23, CDV23, CSR24, DP22a, GSS^{+}23a, GK22a, GKAO20, JB22a, KK23b, LPHK20, LLJ^{+}20, LLM^{+}24, LLT21, OCD22, ÖGS22, PR22, PKSM24, RT21, RPMA22, RBWB21, SJA^{+}22, SR22b, SP22c, ZBC^{+}21, AYH^{+}22, AAMAA22, AK22a, AS22b, Akb22, ATS22, ABCP23, AA23b, AAA22b, AJS23, Anb22, AA23e, Asl22, AY23, BMcKGK22, BWW^{+}20, BST^{+}22, BP23a, BBD23, BKK22, BJS23, BYTG22, BS23c, Boz22b, CWLL20, CPQ^{+}22, CLT^{+}21, CWW^{+}23b, ÇG21, CA23, DDZ^{+}20, DWDG20, DL23, DRV22, DS23a, DSYF22, DRM22a, DSJ22, DR23, DK22a, DXXL20, DCWM20, ESS23, EA22, FK24, FC21, FSWW21, FRS^{+}23, FAM22, GCS23, GZC^{+}22, GDS23, GCP22, GGA24, GATK22, HXZH21, HFH^{+}21, HLG24, JMV24, JPK22, JDLP22, JKB22, JNM^{+}22, JA23, JLE22, KE22, KKR23, Kar22, KA21a, KT23, KS22a, KBR24, Kay22a, KIW^{+}22]. deep [KSK^{+}22, KIAA^{+}22, KK22d, KG22, KSVP22, KR22b, KPP^{+}22, KK22f, KSJ23b, KS23c, LK22b, LSZL23, LG23, LMM23, LWLZ20, LN20, LZF20, Li21, LWL^{+}24, LLKS21, LS23b, LLLX20, Liu21, LZW^{+}21, LXC^{+}22b, LCL^{+}22, LZM^{+}23, Liu24b, LL24, LLYZ23, MRGP22, MBM^{+}20b, MGSB23, MCT22a, MES23, MKBB22, MPV22, Mis22, MRS^{+}21, MRKY22, MMKA23, MRK^{+}23, NBK22, NSR22, NSBT21, NWT21, NZD^{+}24, OS21b, ÖUG22, ODK^{+}23, Ona21, ÖTT23, PBK23, PT22, PAN22, PSMM22, PK22a, PMR^{+}21, PA23, PGD^{+}22, PVRM22, PB22b, PVVS22, RPPK23, RTBC23, RK23b, RAaB21, RV23, SPA^{+}21, SJ22, SSMT22b, SM24a, SPQM20, SM22b, SSN22, SRRM23, SKCS23, SS24, SR24, SZGR21, SM22c, SET^{+}22, SSN23, SCM22, SSP24, SG22e, SP23b, SHVA23, SPK22b, SV22e, SSS23b, SRS23, SKR^{+}24, SM23d, TAT^{+}23, TMB24, UDS21, VPSM22, VS22b, VRS22]. deep [VS21, WKY22, WSZ^{+}23, WWJ^{+}20, WN23, WWL^{+}20, WSJ^{+}21, WHDS22, WT23, XZG^{+}23, XLL^{+}23, YKL23, Yal22, YWQ^{+}21, YLKC24, YBJ^{+}23, Yil22a, YZPL21, ZDH^{+}22, ZZL^{+}20, ZCZW22, ZCZC23, LM20b, JMV24]. deep-convolutional [LN20]. deep-learning [LLYZ23, TMB24]. deep-learning-based [HXZH21]. deep-Q [SM22c]. DeepCM [OCD22]. DeepFM [CBFS23]. DeepJoint [RPPK23]. DeepLab [SKÇA23]. DeepMUI [SVM24]. DeepNet [GP22b]. DeePOF [SJ22]. deer [SAC22]. defeasible [uHAU23]. defect [Anb22, CNG^{+}20, HWBZ21, HLG24, LSZ24, LSW^{+}20, RZ21, YLZC23]. defects [BMcKGK22]. defense [FK20, HE23, HAA23, K22, KHK^{+}23, LHX^{+}23, NY22]. deficit [AAA22b]. Defined [MV22, AAMM23, AA23d, CCBA23, CTA^{+}23, CK24b, CD22, CÇ22c, GYZ^{+}20, GSVS21, HLL^{+}21, IQS^{+}22, KT22a, LCZ^{+}20a, LZZ21, LLW^{+}22b, LGLZ20, LZXG22, MK22b, MS23, PA23, RG23, SKK22, SDR20, VRV23, XHZHXBQX22, YY20a, YLW^{+}22, VK24]. defined-data [MS23]. Defining [CWDM^{+}21]. Definition [ZZLZ22a]. definitions [RV21]. Deflate [NMQ22, TNIK23]. Deflate-inflate [NMQ22]. deformable [LTW^{+}24]. deformation [KWZ^{+}21]. deformation-driven [KWZ^{+}21]. degraded [BOI23]. degree [GXL^{+}20, LXYY21, LX22, MSL22]. dehazing [WZSZ20]. dehydration [WZX^{+}22]. dehydrocarbon [WZX^{+}22]. delay [AK22c, CPCK23, DZLH20, FJCFB23, FSFM22, JSA^{+}20, MNDK22, PB22b, QLL^{+}22, ST22b]. delay-tolerant [CPCK23]. Delayed [MW21]. delays [AZM20, BDG^{+}23, IQS^{+}22, TSLL21]. delete [SV22b]. deletion [TT23a]. delivery [CPCK23, DT22a, LLX^{+}21, LGL^{+}22, WLL21a, Zhu21]. delta [WFY^{+}22a, ZQL^{+}21, ZWZ^{+}22, ZK24]. delta-compressed [ZWZ^{+}22]. demand [CSWZ22, GZ20, JGW20, SPA^{+}21]. demand-side [GZ20]. demands [KA21b]. demographic [SLL^{+}23]. demosaicing [AS23b]. dengue [RK23a]. denial [CSC22, Dah23, DP22a, DCZ^{+}22, HK22a, PKK21, SR22c, TK22a, VJ22]. denial-of-service [TK22a]. denoising [Bei24, WHZL21, AS22d]. Dense [YMWA21, AM20b, KHPH20, KHK^{+}24, LXJ^{+}22, MM21, RKR22, SA22b, SKCS23, WWJ^{+}20]. Densely [RT21]. DenseNet [GA22, JSAA22, LZY^{+}21, WXS^{+}23]. DenseNet-ResNet [JSAA22]. Density [FYH^{+}21, ZTF^{+}20, LZL^{+}20a, LWZY23, LZQF24, SCL^{+}20, XCD^{+}20, YYLL22, Yüc22, ZHX^{+}23, ZZWZ21]. dental [VB21]. deoxyribonucleic [ABB22]. dependability [LSW21]. dependencies [HYG^{+}23]. dependent [BEÖ23, HD23, PP20b, RSKA23, Tur23]. deployment [ASJ21, Ano21-40, CG22, CPCK23, GA23, HM20, LCR^{+}24, LPZ^{+}22, OXBL23, RMS24, ST23b, YPY21]. depression [KG22, YSP23]. depth [KW21, LH21, RS24, Shi22b]. depth-buffer [Shi22b]. Derivation [IA23]. derivative [JS23]. dermoscopy [GM22b]. descent [AFK^{+}22, Anb22, AV22b, PS23b, RKK23, VPSM22]. descent-based [AFK^{+}22]. description [AS22a, GC20, MLC^{+}21, ZLX^{+}24]. descriptions [HLH^{+}20]. descriptor [AI22, BCM22]. descriptors [PKSM24, SMD22a]. desertification [ZDH^{+}22]. Design [Ano21d, BK21, DAdBM^{+}24, KDL20, KGW^{+}20, LTW^{+}24, MKRK23, MWS^{+}23, SPSP23, TD21, TT23b, ZX21b, AG22, AO23, BCC^{+}21, CRB23, CC22b, CD22, EASN22, FGL^{+}20, GM22a, GO22, HZD^{+}22, HL20, Jia22, KA22a, KWL^{+}24, LM22b, LGT^{+}23, MG20, ÖGS22, RT22b, SC22a, SAF^{+}23, TS22, WZL^{+}24, WN21, WNZ24, YL20c]. Designing [SN22a, SSI22, SCdLV20, TPT^{+}22, YNK^{+}23, GAM23, TL21]. desire [USK23]. desktop [Gui22]. desmoking [LJP^{+}21]. despite [BOI23, HNOI24]. destination [ZYXX23]. destructive [LK23]. detached [FZ23]. detached-eddy [FZ23]. detailed [SAQJ23]. detect [Gho21, SSS^{+}20]. Detecting [KM21a, RVF22, ZCZ^{+}22, ARS22a, AYKE21, DCZ^{+}22, KBK^{+}22, KIW^{+}22, KSTV21, PK22a, PKK21, SKPD23, XYLW21]. Detection [AK22b, AQP^{+}22, BMZ^{+}22, GM22b, HFH^{+}21, JAM24, JSAA22, JKS20, KKKS23, KK22f, LK22b, LMM^{+}22, LGW^{+}22b, AST22, SV22b, ST22d, TK22a, VB21, ZWW^{+}21, AM20a, AK22a, AMJK21, AS22b, AMB23, AA23b, AAARR20, AJNS22, AA23d, MS22a, AI21, Ano21-39, AKZA22, AKUA22, ARS22b, AT22, AGSN23, ATC23, BBA^{+}22, BLI20, BST^{+}22, BP23a, BBD24, BK21, BS23b, BKD22, BD21, BRNR23, BVP22, BJS23, Boc21, BS23c, CCBA23, CSR24, CK21a, CCGN20, CMJC23, CMJM22, CK24b, CR23, Cho20, ÇG21, CPA22, CZL^{+}22, Dah23, DRV22, DDH^{+}20, DSYF22, DPYS22, DP22a, DS20b, DD21, DRR22, DBD23, Ele22, ENB^{+}20, FM22, FHS^{+}22, FKY^{+}23, GPDB20, GMM22, GCS23, GDFDF22, GYZ^{+}20, GP22b, GSS23b, GRMP23, GVSS22, Gün23a, Gün23b, HK22a, HK24, HWBZ21, HA21, IDA22, JPK22, JD22, JNS22, JW22, JDG22]. detection [JD23, KV22a, KMR22, KK21b, KBR24, KKP24, Kay22a, KP21, KSSP22, KSTV21, KPP^{+}22, KDA^{+}22, KAAR23, KG23, LAE^{+}22, LPW^{+}21, LTLX22, LLM^{+}24, LJB22, LC21, LSLY20, LZW^{+}21, LHL^{+}22, LXC^{+}22b, LLL^{+}22, LXJ^{+}22, LLC^{+}22b, LAC21, LWZ^{+}20, LCZ^{+}20b, LCM22, LLYZ23, MRGP22, MSN22, MGSB23, ML20, MAR23b, MES23, MMR^{+}22, MPV22, Mis22, MT22, MHB^{+}22b, MCNR20, MP23, NBS^{+}22, ND23, NSBT21, NRP^{+}20, NJ21, ÖUG22, OE22b, PD20, PS22a, PAN22, PL21, PA23, PD22b, PSK23, PKSM24, RKR22, RSR^{+}22, RPPK23, RK23a, RMMK23, RK21a, RPMA22, RBDS22, RDB22, RZ24, SKPD23, SNBAI24, SBA22, SG21, SVD^{+}22b, SKK21, SG22c, SS24, SR22b, SR24, SZZ^{+}21, SET^{+}22, ST23c, SBSK22, SPKK22, SSN23, SMKA22, SCM22, SR22c, SRBH22, SXC^{+}23, SI22, SG22e, SD23, SS22f, SPK22b, SM22e, SFJ^{+}21, SS22g]. detection [SSS23b, SDR20, TZ22, TFZC23, THX22, TAT22, TBNPQ23, TNP21, TPT^{+}22, TÖ21, TSB23, UWF^{+}21, UE22, VKSS23, VY23, VJ22, VRV23, VRS22, VKK^{+}24, VS21, VK23b, WJL^{+}20, WFY^{+}22a, WMC^{+}23, WXS^{+}23, WKL^{+}22, WLLZ20, WSJ^{+}21, WLL^{+}21b, WZG^{+}21, WCWG21, XCX^{+}20, XZG^{+}23, XCJ22, XJW^{+}22, XGZ^{+}20, YS22, YKL23, YSP23, YZX^{+}22, YLZW23, YCL^{+}22a, YLS^{+}23, YMKH22, YY22, ZHX^{+}23, ZDH^{+}22, ZX20, ZWL^{+}23, ZKZ^{+}23, ZZC24, ZCX^{+}21, ZWY^{+}24, ZHWY22]. detector [BST^{+}22, DG21, Gho21]. Determination [OSK23, AYG^{+}21, HXZS23, Yüc22]. Determining [Pan23, ST22a]. deterministic [Boz22a]. Developing [AD22a, AA22a, AS24, LBFT22, SR20b, ZCR23]. Development [Aru22, ADA22, BMG22, CCZM23, KPJ^{+}21, DB23, GM22c, HNG22, PEGP23, SBB21, SX21, WC22]. device [CYK^{+}21, MMKA23, NY22, PK22c, RAN21, SBGC21]. device-specific [SBGC21]. devices [AFM22, HHP23, JPH23, JPAA21, KPM20, LAH^{+}22, MBM20a, SS23c, TT23b, WKL^{+}22]. DFA [HM20]. DFA-based [HM20]. DG [DGB24]. DHGAN [WZSZ20]. DHLink [LBFT22]. diabetes [DRM22b, GRG^{+}22, SM22e, VSP22]. diabetic [AK22a, DRM22a, LK22b, SK21c, VK23b]. diagnosability [JQGL20]. diagnose [GKZ23]. diagnosing [ZBC^{+}21]. diagnosis [AAA22b, AI22, BKK22, Boz22b, CDC20, DWDG20, ESS23, EUYY22, FDD^{+}24, GS22, HLC23, HT21, KNM21, KA22c, LAC21, SG22e, SK21c, SBS22, VPSM22, VLVS22, WY20, Yil22a, ZZW^{+}22]. diagnostic [KCL^{+}20, RK23a, YÇC22]. diagnostics [DBPC22]. diagonal [CWH^{+}24, HCG21]. diagram [DPBP24]. Dialect [TH22, WWM^{+}23]. diameter [TNI23, YNK^{+}23]. diamond [SRGK24]. diarization [SMM22]. dichotomous [DT22a]. DICOM [SN22b]. dictionaries [EGB21]. Dictionary [ZZS21a, ARS22b, LYSC21, RR23a]. diet [TC24]. difference [CdOO^{+}20, CC22b, KKS23, SYJL20, XGZ^{+}20]. different [Aka22b, AS22d, BKGC23, CMT20, Cob22, Kay22b, KMB23, LLJ^{+}22b, OXBL23, RDB22, SA22b, SZJ21]. Differential [AK22c, DLY^{+}23, HZZ^{+}23, AMJK21, CHLD23, FC21, LXYY21, LFWJ22, Pal22, Pan23, PBD23, PVVS22, SCH^{+}24, SYJL20, TZ23, WZC^{+}22a, WZMJ23b, XZD^{+}21, XDH^{+}20, YLZT23, YZZA23, ZZLZ22a]. Differentially [HDS^{+}23, SZY^{+}22]. Diffie [LF23, NR22]. Diffusion [KLJ21, CYQ^{+}20, ZXL^{+}21a]. DigiNet [DM21]. digital [AK22b, Aka22b, BBA^{+}22, DPBP24, DGP20, Jia22, LBFT22, MKRK23, MTY21, OE22b, PSHJ20, PKSM24, RP21a, RVAE21, SEC22, SR21, SLG^{+}20, SKS^{+}23, WNZ24, Yil21b, YRSO23, YÖT^{+}22, uZKH^{+}20]. digitalization [Ano21-42]. digits [DM21]. dihybrid [CO21]. Dijkstra [PCR21]. dilated [RKR22]. dimension [CEN22, KBBH21, THW21]. dimensional [ASAAAA22, BD22, CD24b, CS21, GBBS21, GAMT22, HD23, ITO21, KYP21, LHJ22, LZL^{+}20b, LCW23b, NKY23, SCH^{+}24, SP21b, WQY^{+}22, XDJY21, YZYT21, ZHX^{+}23, ZHJ20, ZHJW21]. dimensionality [AOACAQ21, RSJ22]. dimensions [SYRS^{+}22]. DiPETrans [BAPS22]. DiPOSH [CM21]. Direct [Dev21, FMNF22, HPH^{+}20, JKS20, KS21b]. Direct-Vertical [Dev21]. directed [GDGN23, KKM21, LX22, SSV24]. direction [Ano21-36, FKGO22, YSP23]. directional [PJK23, VK23b, WLX21]. directions [CR23, KKK21, KK23b, MKS^{+}20, PKK21, WZ24]. directive [WCL^{+}24]. directive-based [WCL^{+}24]. directly [KDFK23]. directories [SMR21]. directory [WZZ^{+}22]. Dirichlet [RR22, VSS23]. disabilities [SX21]. disaggregated [ETH^{+}23, SCK^{+}23]. disambiguation [KC20]. disaster [LMBK23, LLZ^{+}21b, SG22a, SAGS24, YLW^{+}22, Zhu21]. disastrous [KS22b]. discover [RV21]. Discovering [JS22c, LZL^{+}20b, Ano21-35]. Discovery [VZR^{+}21, BM22b, CSWC20, ELIG23, GlSL^{+}23, GK22b, LSSQ22, SPWX23, VCS^{+}21, aWLY^{+}22, ZZY22a]. discrete [Ano21e, AS22d, BA20, DPBP24, EKS20, FLG^{+}22, GFPGT21, Gul22, KSP21, LLO21, SO22, SD22, ZSQ22]. discrete-time [GFPGT21, LLO21]. discriminant [CB22]. Discrimination [YYZS22, SPHP21, WER^{+}23]. discriminative [KDC22, RR23a]. disease [ABCP23, AI22, AR22b, AYJ^{+}22, BP23b, DR22c, DS23a, GKZ23, HT21, IDA22, JMV24, JA23, Kan22, Kay22a, KAAR23, KA22c, LJ22, MPV22, MCT22b, PLP22, RKR22, RTBC23, RMMK23, RG22, RV23, SRRM23, SLL22, SSS23b, SBS22, TC24, VLVS22, YRV^{+}23, YÇC22, ZCX^{+}21]. disease-diet [TC24]. diseases [BM22c, CÇ22a, KSS21, PK22a, VPSM22, VR22, Yil22a, YMKH22, ZBC^{+}21]. disentangled [MLS^{+}24, SSP23]. Disentangling [ZLL^{+}22]. disjoint [LZXG22]. disjunctive [HKA20]. disk [HTZ^{+}22, JHZ20]. disorder [AAA22b, BGG24, GS22, SG22e, YY22]. disorders [LK22b, VY23]. dispatch [ZZLZ22b]. Dispersed [MS21b, Dev21]. displacement [BEÖ23, LTW^{+}24]. disrupt [SSK23b]. dissemination [FLB23, HLDT^{+}23, MCD^{+}23, PK22d, RSS20, SK22b, YPO21]. distance [DÖD22, Eke22, KK24, KGM23, KK23a, KW21, THW21]. distance-2 [KK24]. distances [KFKD22]. Distillation [LLX^{+}24, CHL^{+}24, LDW^{+}23, LGX^{+}23]. distortion [KK21b, LHWT20, RGKK21]. distraction [MSA21]. Distributed [BZK^{+}21, JR23a, LZM^{+}23, PKK21, RZ23, ST21, YWQ^{+}21, ZAB22, AS22a, ACJ21a, AEMS24, ASJ21, ACJ21b, APL^{+}21, BAPS22, BOI23, BWTJ20, BANT20, BZT23b, CF21, CK23, CSC22, CYC21, CLC^{+}23, CWFK23, Cud20, Dah23, DP22a, DCZ^{+}22, DJGF21, EBLM22, FQD^{+}23, FD20a, FCMM20, FKO23, wGTC22, Gün23a, HK22a, HGM^{+}23, HXST22, IMNC22, JZL22, KK24, KK21a, KKP24, KMY24, LHC^{+}23, LDCD22, MM24, MLKD20, MHPA21, NKKM21, NR23, PK22c, RN22, SA22b, SKH^{+}21, SZW^{+}22, SR22c, SS23g, SWS^{+}24, TTB^{+}22, TBT^{+}21, TK22a, TP20, TLS22, VJ22, WBZ21, WFHC21, WSL^{+}20, WWW^{+}23, XXD^{+}22, YBJ^{+}23, YF20, ZZL^{+}20, ZTL^{+}21, ZWC^{+}23]. distributing [JOW23]. distribution [AET^{+}22, AE22b, AI21, AC23, BK22a, DKP^{+}20, HOS^{+}21, HUC^{+}22, KABE^{+}20, LLY^{+}23, LSX21, RMR^{+}22, RF23, SVB23, VK23a, WBL22, XHZHXBQX22, ZWCC23, ZFW23]. district [SM21]. Diverse [HGNN22, AKEC20, HVT23, SK20a]. diversities [ZLM22]. diversity [LGL^{+}21, WL23, JB21]. dividing [CZL^{+}22]. Divisible [AB20a]. division [Ben22, JCL^{+}20, LXT^{+}22, SA22b]. DL [CDV23]. DL-DC [CDV23]. DLB [SSV24]. DLFC [MMS24]. DLFC-NN [MMS24]. DLIFT [HXZH21]. DLPM [ZHXY23]. DMAP [ST22c]. DML [LLM^{+}24]. DML-OFA [LLM^{+}24]. DNA [BPB21, KB22, PQKDT21]. DNFNet [NSR22]. do [MLVGR24]. Docker [YPY21]. document [CKL20, GSS^{+}23a, YÖT^{+}22]. does [WWW^{+}23]. DOF [TYL22]. dolphin [DR23]. DOM [NJK22]. domain [ASANR22, CD24a, CDN^{+}21, FZC20, GC20, KKS23, KSK^{+}20, KSP21, LGDW22, LLZ^{+}20, QWL24, RSJ22, SSP22, SSP23, SKS^{+}23, TSG21, VCS^{+}21]. domain-based [TSG21]. domain-science [VCS^{+}21]. domain-specific [CDN^{+}21]. domains [GLJ20, LLN^{+}23]. dominance [HXY^{+}22b, YZ21]. dominant [SB23b]. dominated [HXY^{+}22b, Yil21c]. dominating [Sin23]. domination [KK24]. Donelan [WLX21]. Dongguan [YLKC24]. Don’t [NdMP22]. DONUTS [KPA23]. dose [KKKS23]. dot [AG22, KA22a]. dots [GM22a]. Double [LWC^{+}21, FZA22, HXY^{+}22a, KHR22, PHZ^{+}22a, SKR^{+}24, WYLL24]. double-layer [FZA22, WYLL24]. Double-weighted [LWC^{+}21]. Doughnutie [NFF21]. Douglas [FZZZ23]. Douglas-Rachford [FZZZ23]. down [TSCM22]. downstairs [KBL^{+}21]. DPA [WMS^{+}23]. DPA-UNet [WMS^{+}23]. DR [KK22b]. DRA [LLW^{+}23]. DRA-MQoS [LLW^{+}23]. dragonfly [KK22f, ZZLZ21]. DRAM [KSD22]. drifting [RSR^{+}22]. drilling [YTW^{+}24]. DRIP [KMZ^{+}20]. drive [DZW23, Kum21, ZTY^{+}22]. drive-assisted [DZW23]. driven [AQP^{+}22, Anb22, BP23a, BJK^{+}24, CJP^{+}21, CA23, KGRG24, KBJ21, KWZ^{+}21, MCT22a, MCT22b, NGOS22, NRMB23, PKSM24, RF21, SAAAA22, SS22d, WHDS22, Xia20, ZDH^{+}22, ZYPD24, ZZLZ21, MTY21]. driver [MSA21, YL20c]. drivers [GDA^{+}21, KVP21]. drives [WWW^{+}23]. driving [Cho20, DKL21, WHH^{+}23, ZK24, ZWC^{+}23]. DRL [JDG22, LGX^{+}23]. DRNN [NBPR22]. drone [YRSO23]. drones [NT23, PCK23]. Dropout [PS23b, Anb22, CT22, CB22]. Drowsy [Cho20]. DrpXLm [MMR^{+}22]. Drug [DSS^{+}23]. drugs [GGS^{+}22, KA21a]. DSEL [CTA^{+}23]. DSP [WMY^{+}21]. DT [Faz23]. Dual [WXCY20, GAS23, HNSS22, HLT23, HZW^{+}23, HUC^{+}22, JT23, KA22a, LLO21, LLN^{+}23, MKRK23, TWXL21, ZML^{+}23, ZWCC23]. dual-blockchain [ZML^{+}23]. dual-domains [LLN^{+}23]. Dual-Hop [WXCY20]. dual-mode [KA22a]. dual-system [HNSS22]. dual-usage [HZW^{+}23]. duf [ADGT22]. dummy [CYZX23, MS21b, PR23, SSSP21]. dummy-based [CYZX23]. dummy-generation [PR23]. duplicate [GRMP23]. duplication [AA21, LNI^{+}24, PL21]. during [BK22a, KCM^{+}22, RDB22, SP22b]. duty [BM22b, GJBM22]. duty-cycled [GJBM22]. DVO [LmJdL^{+}22]. dwarf [AK22a]. DWDM [SLLMTS^{+}24]. DWT [SN22b]. Dynamic [Agr21, ADGT22, BZT23b, CWW^{+}23b, CLWX21, CG22, HOS^{+}21, HBB20, HL23, JB22a, KABE^{+}20, KMZ^{+}20, LLP^{+}22, LKR^{+}22, LZZ^{+}20, MR23b, NA22c, PK23a, RN22, SJ23b, WZG^{+}21, WJD22, XCG^{+}22, ZBZ^{+}20, ZXL^{+}21a, AN23, AS22c, AB20b, Ano21-39, Boc21, BN21, CKKK20, CON23, Den24, GlSL^{+}23, GC20, GD22, GK23b, HJZ^{+}22, JKB22, JPH23, JC21b, KBL^{+}21, KFML20, KK23a, LLX^{+}21, LLMX21, LFWJ22, LLL^{+}24, LK22c, MG22, MdAA^{+}21, MTRA23, MW21, NKKM21, NdSSSN20, PM23a, QKSK23, RTSK23, RYG^{+}21, SM22a, SAC22, SKS22, WLG^{+}22, YSL^{+}21a, YZPL21, YZ21, YSC^{+}20, ZHX^{+}23, ZAB22, ZZWZ21, ZLV22, ZHXY23, dRdSC^{+}21]. Dynamics [TSLL21, BEÖ23, FZ23, HLDT^{+}23, IA23, JPA^{+}23, KPF^{+}20, LYL21, VCFZ20, XR21, YZXL22]. dynamism [CJY^{+}20].

E-Bayesian [CSL20]. e-business [YYY^{+}23]. e-commerce [WQS^{+}23, Zha22]. e-healthcare [BVS20]. e-Infrastructure-based [CPPP21]. e-infrastructures [MBT^{+}20]. E-learning [PM22, PKB^{+}23, SP22b, VS22b]. E-maintenance [HL20]. e-science [MBT^{+}20]. E-shopping [MRKY22]. e-STLB [MSSM24]. e-TOALB [LS24]. e-waste [LKR^{+}22]. eagle [RG22]. ear [GPDB20, KCP^{+}22, YSS^{+}21]. earliest [DWZ^{+}20b]. Early [BBA^{+}22, VY23, YMKH22, AK22a, MSBR23, NSBT21, RK23a, VPSM22, VSP22]. earth [WXLD21, BBB^{+}20c]. EarthCube [VZR^{+}21]. earthworm [CPA22]. earthworm-rider [CPA22]. EASE [YLZ^{+}21]. East [HNG22]. EasyScienceGateway [MPLW24]. eavesdropper [CAS24]. Ebola [RV23]. eBPF [KLH24]. EC [RJ22]. ECG [AS20, IAASK23, KPM20]. Echo [SBB^{+}20, JB24]. economical [DP22a]. economically [KDS^{+}20]. ecosystem [DS22a, GKLS23, KV24, WFHC21]. ecosystems [BR21, DPB22b]. ECR [CPLX21]. eddy [FZ23]. Edge [Agr21, GlSL^{+}23, Liu24b, MMKA23, PKC^{+}20, SZI^{+}23, YYY^{+}23, ZALM23, BTDJ21, CHMC21, CLT^{+}21, CJC22, DDB^{+}21, DAT23, DPB22b, DDUK23, DG21, DKC^{+}21, DYF20, FGZC23, GAS23, GWA^{+}23, HGW^{+}23, JK22a, KM23a, KAP20, KKP20, KS24, LGL^{+}22, LLW^{+}23, LSZ^{+}23, LCCT22, LWW23, LZW^{+}21, LAH^{+}22, LLYZ23, MGN^{+}22, MP23, NMQ22, PRS23, PLZL24, PB22b, QLL^{+}22, RGPC23, SSN21, SSW^{+}22, SG22d, SXC^{+}23, TWL^{+}20, TWG^{+}21, TS24a, TS24b, VG21, Wan22b, WDW^{+}22, XWC^{+}22, XCZ^{+}21, XCG^{+}22, XLL^{+}23, YV23, YLZ^{+}21, ZCN22, ZYY^{+}24]. edge-based [JK22a, TWL^{+}20]. edge-cloud [LLYZ23, WDW^{+}22]. Edge-enabled [PKC^{+}20, PLZL24]. Edge-Fog-Cloud [Agr21]. edge/cloud [SSN21]. edges [ST22a]. editing [CXX^{+}22, Kim21b, NMM21]. Editorial [HFFA20, KS21a, Liu24b, WYZAD20, Wu22]. Edman [KT20]. EDTL [DSYF22]. education [Che24, HKMS21, LZZ^{+}23]. educational [DB23, RRJ23, SA22c, ZHDW24]. EduChain [LZZ^{+}23]. EEG [KKC22, LLJ^{+}20]. Effect [ECIB20, RKL21, EUYY22, LJZ21, MHL^{+}20, MSL22, UYÖ^{+}22, VS22c]. Effective [BM22b, CHMC21, JKKL21, KH22, KDA^{+}22, RSR^{+}22, AS23c, AD22c, BGG24, CÇM22b, DRM22a, Dev21, FK24, GWA^{+}24, HKMS21, LJ22, LWLZ20, LFG^{+}22, LCCT22, MMR^{+}22, MTY21, NTB23, RSS20, RSM22, RK23a, RK23b, RMS22, SAPC21, SKA23, SHA^{+}22, SSN22, SLD24, SCM22, SK22d, SHV24, TK22b, ZYH^{+}23]. effectively [ZTY^{+}22]. effectiveness [ÖTT23, SWZW20]. effects [VAB22]. efficiency [Cud20, FSdP^{+}23, IPRS21, JHZ20, JW21, KR23, Le23, LBFT22, LCW23b, MdARS^{+}23, NMS^{+}22, PMP23, SPC^{+}21, SP22a, SS22f, TTA20, VRB21, YLT^{+}21, ZYH^{+}23]. Efficient [AZM20, AM20a, AÜ23, BM22a, FKGO22, HKA20, JPN21, JLZ^{+}24, KML23, KA22a, KML21, LCSR21, LWL^{+}23, LLL^{+}21, NSK^{+}24, NWT21, PM23a, PL21, QNZP22, RTSK23, SET^{+}22, SKS20, SCS^{+}21, SSSP21, TNI23, VB24, WMY^{+}21, XXD^{+}22, ZDH^{+}22, ZLV22, AK22a, AA21, AMB23, AD22b, AAG^{+}22, APV23, ARFA21, ARS22a, ASRN23, AYJ^{+}22, BP24, BK21, BKD22, BFM^{+}23, BRS^{+}22, BKS22, BK22b, BN21, CÇY22, CKKK20, CLE^{+}20, CC22b, CWW23a, CDR^{+}23, CL22b, DS22a, DP22b, DGB24, ETKD23, FKY^{+}23, GBBS21, GFQ20, GPR^{+}22, GRL^{+}22, HYG20, HCG21, HUJ22, HTZT23, HLS^{+}20, JB22b, JS22a, JPH23, JZL21, KPA23, KLK23, LA22, LMR22, LM21, LHC^{+}23, LGDW22, LCZ^{+}20a, LZZ21, LYF22, LGZ^{+}22, LXW^{+}23, LJB22, LC20, LK22c, LZY^{+}20, LSW^{+}20, LZW^{+}21, LZSC23, LAC21, LS24, MR23b, MT21, MSSM24, MG20, MS21a, MNYN21, MS21c, MSL22]. efficient [MCNR20, NFF21, NA22c, NQ21, NSSS22, NZD^{+}24, OBTC20, OHRS21, PPA22, PT22, PP21, PK23b, MR23a, PHDS22, PB22b, PM22, PM23b, PDRM24, RC22, RJH^{+}23, RGKK21, RSJ21, RAG21, RYG^{+}21, ST22b, SPS23, SRGK24, SEC22, SK22b, SS22c, SB23b, SM23a, Shi22b, SNGK21, SKSN22, SS23f, SR20a, SBB^{+}20, SM24c, SGHL20, SR22d, SRS23, TAB21, TLXX21, TS22, URK^{+}22, VVK23, WCWG20, WFY^{+}22a, WXCY20, XYLW21, XWD^{+}22, XZW^{+}20, YV23, YLZ^{+}21, YMWA21, YZX^{+}22, YB23b, YM22, YPO21, YLW^{+}21, ZHX^{+}21, ZSQ22]. effort [GM22c, SBB21, YJJ23]. eigensolvers [KHK^{+}24]. eigenvalue [MM21]. eigenvector [HWG23]. einit [CLE^{+}20]. Elastic [ASJ21, YZPL21, LDW^{+}23, MSK22, dQtZWS22, SAM^{+}23, SWS^{+}24, WJJ23]. elasticity [KRB^{+}20]. Elderly [AR22a, LWZ22a]. election [DQF^{+}23, GAMT22, JP21]. electric [GZ20, KAR^{+}23, LST22, NM23, SSRA23, Xia20]. electrical [VCFZ20]. electricity [AJAA21, DÖD22, LWZY23, PKK23, YYLL22]. Electro [SVD^{+}22b]. electrocardiogram [ZWT22]. electroencephalogram [TÖ21]. electromyography [AY23, BZGM22, CLY^{+}21, LLJ^{+}22b]. electromyography-based [AY23]. electromyography-feature [CLY^{+}21]. electron [GS21]. electronic [KRK21, QWW^{+}22, SC22b, SHZY20, SK22d, YLM21]. element [KS22c, SD22]. elephant [AAMM23, BKD22, RKC22, SSK23b]. elevator [LCZ^{+}20a]. ElGamal [KV22b]. eligibility [CA23]. eliminating [GB23]. elimination [MP23]. elite [AO23, LYG^{+}21]. Elliptic [PSB^{+}20, SV22d, AE22a, AMAT22, HK22b, PSP24, PA21, SSS23a, VB22a, WGYZ22, ZWT22]. ELM [LLJ^{+}22b]. Elman [AFBM^{+}23, SM22c, VRV23, WWG21]. email [LDH^{+}22]. embed [CT22]. embedded [JK22a, JPAA21, LSZ^{+}23, LGX^{+}23, PT22, PKB^{+}23, RR22, ZCX^{+}21]. Embedding [WSL^{+}20, CYQ^{+}20, DPB22a, GK22a, IG24, KBBH21, NNJC23, RP21b, WZZ^{+}20, WZHL21, YT21, ZLCL21]. embeddings [HLH^{+}20, Ona21]. Emergency [ZWC^{+}23, DZLH20, DWY^{+}21]. Emerging [AYH20, DYF20, QZB^{+}23]. EMG [TA22b]. eminent [MKRK23]. Emoji [VM23]. emotion [AC22, BMV22, JWL20, JJZ^{+}21, JTY^{+}21, PS22a, SHL^{+}22, SKCS23, UDS21]. Emotional [SM22b, TÖ21]. emotions [PB21, RMS22]. emphasizing [HYG^{+}23]. Empirical [HUC^{+}22, AOACAQ21, GDA22, SG22e, TI22, WWW^{+}23]. empiricism [SR22b]. employee [VP22b]. employees [AV22b, JZB20, SM23d]. Employing [TSB23, HLS^{+}20, RG22, SBSK22, SD23]. empowered [KS24, SRBH22]. Enable [YZXL22, HPS23]. enabled [AGSN23, BV22b, BM22c, BTDJ21, GSS^{+}23a, GlSL^{+}23, JDLP22, JR24, JA23, KTM22, KKR23, KS20, KCP23, KK22e, LZZ21, MKA23, NLG^{+}20, PPA22, PKC^{+}20, PLZL24, SG22a, SRL23, Sin23, SSRA23, VDL23, VB22a, WZMJ23b, YYY^{+}23, ZCN22, ZZY^{+}22b, ZALM23]. Enabling [BS23a, CS21, GHL^{+}23, HGDD20, MC24, SSN21, SV21, ZBY^{+}23, ASC22, CCL^{+}22, LGA^{+}20, SS22e]. enactment [APL^{+}21]. encoder [AV23, Anb22, KDA^{+}22, SSP23, SPK22b, YOWY22, ZLZ^{+}22b]. encoder-decoder [YOWY22]. encoding [CC22b, KSVP22, LC21, NMM21, OHRS21, TNIK23, Yil21a]. encrypted [DCK21, ZLV22]. encryption [AMAT22, APTT22, BV22b, BVS20, CDR^{+}23, CYDW20, DP22b, DKK24, FZA22, GAH^{+}22, HSR23, KSS24b, LYSC21, LYC22, LWW23, LF23, MG20, NA22a, PZZ^{+}23, RB22, SAN24, SLG^{+}20, TDZ^{+}24, TSG21, TT23a, TSV^{+}22, VK23a, XZL^{+}20b, XZYW23, ZHDW24, ZWT22, LF23]. end [CJ21b, KUK^{+}24, LLW^{+}22b, KLL^{+}21]. end-to-end [LLW^{+}22b]. end-user [CJ21b, KUK^{+}24]. endoscopic [LJP^{+}21]. endpoints [WK20]. Energy [CC22b, DP22b, ETKD23, FKY^{+}23, HTZT23, IA22, KB21, KSSK22, LGM21, LLX^{+}21, LLH^{+}20, LWCZ22, LCW23b, MSSM24, MSS22, RJ22, ST22b, SPC^{+}21, SKR^{+}24, WXCY20, YLZ^{+}21, YLW^{+}22, AZM20, AAT21, AFBM^{+}23, AD22b, ASC22, AKLK24, AGSN23, ASRN23, BP24, BRS^{+}22, BKS22, BZL^{+}22, CJ21a, CWW23a, Cud20, DÖD22, DS22a, DA22, Faz23, FSdP^{+}23, GSG20, GPR^{+}22, HMK24, HLL^{+}21, HHP23, JS22a, JKP22, KSJK21, KFML20, KML23, KK23a, KLK23, KR23, KAR^{+}23, LK22c, LDCD22, LAC21, MR23b, MT21, MTT20, MNYN21, MS21c, NA22c, NMS^{+}22, PPA22, PHDS22, PB22b, PM23b, RGKK21, RP22, RZCA21, RT22b, RKC22, SPS23, SP22a, SSM^{+}21, SKSN22, SS23f, SR20a, SM24c, SR20b, SPK^{+}22a, SS22f, TCRP23, TAH22, TT23b, URK^{+}22, VRB21, WJL^{+}20, WFY22b, XLZ20, YWQ^{+}21, YLT^{+}21, YLJ22, YFdP^{+}23]. energy [YSS^{+}21, ZLTX21, ZCW^{+}23]. Energy-aware [IA22, KB21, KSSK22, LLX^{+}21, LLH^{+}20, LWCZ22, YLW^{+}22, AGSN23, RGKK21, RZCA21, RT22b, YSS^{+}21]. energy-conscious [LDCD22]. Energy-efficient [CC22b, DP22b, ETKD23, FKY^{+}23, MSSM24, WXCY20, YLZ^{+}21, BP24, BKS22, CWW23a, DS22a, JS22a, LK22c, LAC21, MR23b, MNYN21, MS21c, PPA22, PHDS22, SPS23, SKSN22, SS23f, SR20a, URK^{+}22]. energy-low [XLZ20]. energy-saving [YLJ22]. Enforcing [ZLZ^{+}22b, ZOS^{+}21]. engagement [Kim21a, LDH^{+}22]. engine [AJ21, IP20, SG22b, SDSW21, RMS24]. engineering [AMA22, AO23, GFA21, KLJ21, SPQM20, SMAG22]. engines [HMK24]. English [BKK23]. Enhance [SGS21b]. Enhanced [CO21, BKK23, DS20b, DSJ22, MS23, NA22a, PHZ^{+}22a, PKS22, RMS22, SS21, SR22a, SR24, VRV23, VK23b, DK21, GMM22, GLA^{+}22, JCGT24, KK22b, Liu24a, MK23, MS21c, MGGA20, MKS^{+}20, MMKA23, NNVD22, PR23, PGL^{+}23, PLX^{+}24, RPPK23, RCK22, SB23a, SS22c, SM23a, SSP23, SAMS23, SR22d, VSS23, WLZ^{+}21b, XZW^{+}20, YAR22, ZFZ^{+}20]. Enhancement [AS23b, SYG22, WSJ^{+}21, APA22, BMG22, HYC^{+}24, Hua20, HYT^{+}21, JKP22, JPAA21, LZSC23, LL23, TYLY23, TJZ23, VP22b, WLJ20, WCCC20, WJD22, YSK22, ZLT21]. enhancement-based [HYC^{+}24]. Enhancing [HTZ^{+}22, IAO21, MJAM23, NJK22, SNBAI24, VBM^{+}21, Ano21-40, BPAE20, Bei24, BKS22, JAR^{+}24, MRS^{+}21, SM24c, SS22f, YBJ^{+}23]. enlarging [PPA22]. enriching [TSCM22]. enrichment [EGB21, PMC^{+}21]. Ensemble [KIAA^{+}22, KP21, AS22b, BMV22, CSR24, CSC22, CD24b, CMLL21, DSYF22, DRR22, Ele22, GPDB20, GMM22, HYC^{+}24, KUK22, Kan22, KSSP22, LSZ24, NR23, NBHN22, RSJ21, RF23, SKPD23, SMD22a, SRL23, TBNPQ23, Tur23, TSD23, WLLZ20, WT23, XLZ20, YLZW23, ZWZ^{+}21]. ensemble-based [GPDB20]. ensembled [BS23c]. ensembles [DPdS^{+}23]. ensure [ÇKÇ22]. Ensuring [Ano21f, ZWZ^{+}22]. enterprise [BC21, CJY^{+}20, ZS22]. enterprises [ZK24, Zha22]. entire [XHZHXBQX22]. entities [GDFDF22]. entity [BKK23, GZC^{+}22, KC20, ZLW^{+}21]. entries [NTK22]. Entropy [KT22a, KSY24, Ano21g, CS23, DS24b, JMY21, KT20, LGL^{+}21, LWC^{+}21, SS23a, SNBAI24, WER^{+}23, ZCL21, LWZ22a]. entropy-assisted [LGL^{+}21]. entropy-based [SS23a, SNBAI24]. ENU [KSY24]. Environment [ZXL^{+}21b, Agr21, AI21, BV22b, BC23a, BP24, BKD22, BA24, BRS^{+}22, BJS23, BBB^{+}20c, COZ21, CHLD23, DR21, DDUK23, Dev21, DXXL20, DK22b, DS22c, EI22, FSA^{+}22, FEK20, FPXM21, FXX22, wGTC22, GR22b, GSG^{+}23, HAR20, HK22b, Hem22, HAA23, JPH23, JSP20a, JSP20b, JGJ^{+}21, JQ22, KK22a, KKP24, Kim21a, KLK23, KSSK22, LYW^{+}21, LDZ^{+}22, LSZ^{+}23, LPZ^{+}22, LM24, MSN22, MSSM24, MSK22, MV24, MA22, NTK22, NA22b, PSP24, PSK23, PI21, QCC^{+}23, RB22, RAN21, SB23a, SPQM20, ST22c, SZY^{+}22, SK21b, SJ23b, SM22d, TAT^{+}23, UK23, VDL23, VS22a, WCWG20, Wan22b, WKL^{+}22, WDW^{+}22, XCG^{+}22, XLL^{+}23, YR21, Yu20, YSC^{+}20, ZTS^{+}22, Zha22]. Environmental [LLC21a, CJY^{+}20, GDA22, MDSG^{+}23, SM22c]. environments [AKEC20, AA21, AB22a, ABC^{+}21, ACC^{+}23, BYPO21, BN21, CWDM^{+}21, CMS21, DWZ^{+}22, Gho21, GYL^{+}21, KMY24, KRB^{+}20, KHR22, LBZ^{+}22, LCM22, MG22, MGS^{+}20, MPLW24, MHPA21, MRAM^{+}21, PYC^{+}20, PJP21, SKH^{+}21, SBGC21, SHBC20, SWS^{+}24, TSR22, VKK^{+}24, WFY22b]. EPA [ZYH^{+}23]. EPF [RCK22]. epidemic [PPM^{+}20]. EPIDL [NZD^{+}24]. Epilepsy [SG22e, DLC^{+}21]. epileptic [AHT^{+}20]. epistasis [MLZ21b]. epistasis-tunable [MLZ21b]. EPK [SS21]. EPK-SVM [SS21]. EQK [LML^{+}23]. EQK-IncResNet [LML^{+}23]. equally [BA20]. equations [DS22b, IA23, Pan23]. Equilibrium [FCZ20, BD22]. equivalence [MdAA^{+}21]. era [AAT21, DAT23, SG24]. Erlang [JBBH21]. Erratum [Ano23a]. error [KK21b, LZSC23, RRIL22, RKL21, SPSP23, VSK22, WHC^{+}20]. error-aware [LZSC23]. errors [SK22c, VS22c, Wan22a]. ERSChain [ZHDW24]. ESLO [VK23b]. establishment [AASPR22]. estate [LLA^{+}22]. estimating [MA22, OHFF20]. Estimation [KSJK21, LSX21, SAHAN22, SK22c, AK22c, CC22b, DKK24, FKGO22, GM22c, HYG^{+}23, HWY^{+}23, HUC^{+}22, IAT^{+}23, Kar22, KG23, LP21, LWZC21, LZQF24, MSA22, ÖSTY22, PS21, RP21b, SAB22, SRGK24, SBB21, SS22g, TAB21, TYL22, TO22, VSK22, XCD^{+}20, YAR22, YJY^{+}20, ZZWZ21]. estimations [CSL20]. estimator [AAA^{+}22a, AA22a, AAK21c, HML19, HML20, KÖÖG22, LLH19, LAKA21, LAK22, MAAK23, OU22, OAA22, SU23, TÖK21, ZHX^{+}23, FPHZ20]. estimators [AD22a, AU22, ADA22, BC22, BK22b, BK23, Cek22, GLA^{+}22, LAP23, NSSS22, ÖK22, VS22c, YAR22, ZSS22, ZBY22]. Ethereum [FD20b, KSA22]. ETL [HD23, JWT^{+}20]. Euclidean [KGM23]. Evacuation [NKY23, DWZ^{+}20b, DWY^{+}21, LLZ^{+}21b]. evaluate [WKY22]. Evaluating [LTLX22, LM20b, MSA22, YLT^{+}21, Eke22, Le23, Mit20]. Evaluation [AA23c, AKRR20, CMT20, CSD23, GFPGT21, HML20, LLH19, LDH^{+}22, OGA^{+}22, PGD^{+}22, Ano21i, Boc21, CdOO^{+}20, CLC^{+}20, CYC21, CW22, FFLM21, GGCGS20, GSG20, GB23, HK22a, JGJ^{+}21, KLH24, KK22c, KPP^{+}22, LWZ22a, LGZ^{+}22, NET21a, NMS^{+}22, NGB23, PKC^{+}20, PLX20, PP20a, SLLMTS^{+}24, SA22c, SKCS23, SLC20, SWZW20, VMFL23, WZC^{+}22b, XGCZ23, ZWW^{+}21, ZST^{+}23]. evaluations [HGDD20, NKKM21]. evaporation [DP22b]. event [Ano21e, CJP^{+}21, GLW21, GLW22, KM21a, LPC^{+}21, NRMB23, QWL24, SCKU23]. event-driven [NRMB23]. events [ARS22a, DSS21, GLW21, JZB20, KIW^{+}22, PMR^{+}21, VV23, YW22]. Eventually [BOI23]. Eviction [CPLX21]. Eviction-cost-aware [CPLX21]. evidence [KHEF22, OMA^{+}23]. evidence-based [OMA^{+}23]. evidences [AYG^{+}21]. Evolution [WZ24, AMJK21, AKS^{+}22, CY22, CDR^{+}23, CHLD23, GlSL^{+}23, KUK^{+}24, LYL21, LFWJ22, Pal22, PBD23, PVVS22, SCP20, SYJL20, WFHC21, WZC^{+}22a, Xia20, ZZS21a, ZK24, AK22c]. evolution-vehicle [AKS^{+}22]. Evolutionary [LLSC22, NBHN22, YJY^{+}21, AA22b, AMA22, DZCL22, DJGF21, GLW21, GBS24, GBB22, HXY^{+}22b, JM22, K22, LWCZ22, MAR23b, MS22b, RGM22, WZG^{+}21, WYLL24, XL21, YZ21, ZX20, Zhu22b]. evolving [MRAS^{+}23]. EX [VBP^{+}24]. exact [KB22, KML23]. examination [ÖSTY22, RS24]. Examining [ZR24, KGK22]. ExaMPI [SBG20]. example [WZ21, WN23, WNZ24]. Exascale [SBG20, BBB^{+}20b, BBP^{+}24, SRS^{+}21, WSM^{+}24, WS24, ZBY^{+}23]. exchange [CLL^{+}21, SKS^{+}23, TJC24, WGYZ22, TI22]. excitation [LSZL23, ZKZ^{+}23]. Excogitating [BDK22]. executable [SR24]. executables [RF23]. executing [PPM^{+}20]. Execution [AML^{+}22, BEJD22, RSM21, AAB23, BAPS22, BP20, BLB^{+}24, BWS^{+}21, EBLM22, KSB23, LK22c, MC24, PYC^{+}20, QKSK23, SCD^{+}23, ZZL^{+}20, dSNdL^{+}23, dOPBdO21]. Execution-Cache-Memory [AML^{+}22]. executions [KVV20, MdARS^{+}23]. exhaustive [PR24]. exogenous [SMD^{+}21]. expansion [KC20, KS22d, RV21, SKSP23]. Expectation [ZXL^{+}21b]. expenditure [KSJK21]. expenses [CGS^{+}21]. Experience [DBS^{+}22, PP20a, Ano21-41, ABC^{+}21, ACC^{+}23, KY23, OAS23, PLX20, VG21, XZ20b]. Experience-based [PP20a]. Experiences [LBG^{+}20, RT22c]. Experimental [KDS^{+}23, BBP^{+}24, FPHZ19, KE22, KHK^{+}24, KPP^{+}22, NKKM21, Pan20]. expert [BBB22, TSA21]. experts [HGNN22]. expiry [GGS^{+}22]. Explainable [OAS23]. explanations [VAB22]. explanatory [ZYXX23]. explicit [CdOO^{+}20, KW21]. Explicitly [PHZ^{+}22b]. exploitation [DM22b, Liu24a]. Exploiting [ALBZ21, CY24, DGM21, GLM^{+}22, LJC23, LFW20, NMQ22, OHFF20, SWCB20]. Exploration [HSO^{+}21, BRL^{+}20, DWZ^{+}22, DPS21, Liu24a, SAF^{+}23, VZR^{+}21]. exploratory [KBJ21]. Explorer [DB23]. Exploring [BBBC22, HKMS21, KUK^{+}24, PRS23, SLL^{+}23, TLQ21, YPO21]. explosion [XSZ^{+}23]. exponential [GLA^{+}22, LLT21, THT20, ZSS22]. exponential-cum-ratio [GLA^{+}22]. exponentiated [WBL22, YK23]. Exports [TG23]. expose [YB23a]. exposure [ASMS21]. Expression [SV22e, JB22a, Kim21a, PVRM22, XFH^{+}21]. Extended [ASAAAA22, AMB23, JCL^{+}20, NGB23, PJP21, WWZ^{+}22, PD22a]. Extending [BF24, NNJC23]. extension [JYW^{+}20, PMC^{+}21, VMFL23]. Extensions [GTM24]. extensive [SKK22]. extent [AHV21]. external [AY23, CPYC21]. Extracting [YSS^{+}21, BWW^{+}20, VM23]. Extraction [LSZL23, PYC^{+}20, ATS22, BKM^{+}21, BBB22, CLX^{+}21, DLC^{+}21, GLW22, GRMP23, HWBZ21, Kay22b, KBS^{+}22, LN20, Li21, LWL^{+}24, Mis22, RA21, RMS22, SSMT22b, SGH23, SBB^{+}20, SBS22, TYLY23, TA22a, TA23b, VLVS22, WHJ^{+}20, XY21, ZNLL22]. Extraction-and-excitation [LSZL23]. Extreme [AJNS22, ANP^{+}20, APM^{+}21, BBA^{+}22, LLMZ21, PK22b, RR23b, RV23, SM23b, TS22, TUD21, VR22, PS23b]. extreme-scale [ANP^{+}20]. exudates [SBA22]. eye [PB21]. eyes [CCZM23, KKC22].

F [LWZ^{+}22b]. F-SWIR [LWZ^{+}22b]. Fabric [LZZ^{+}23]. facade [SP22a]. Face [XLY^{+}23, JB22a, JD22, Kim21a, LN20, LCW^{+}23a, NGD^{+}22, NNX^{+}23, RMS22, SMBM24, SPKK22, ZYCS22, ZLZ^{+}22b]. Facebook [HLDT^{+}23]. faces [AAK^{+}21a]. Facial [SV22e, CXX^{+}22, Kim21a, RMS22]. facilities [ZABT^{+}20]. factor [DWZ20a, KGRG24, NBK22, TWW^{+}21, TT21, WAY^{+}21]. factor-based [WAY^{+}21]. factored [AS23a]. factorization [CPQ^{+}22, GHRM21, JW22, WQS^{+}23, ZWY^{+}24]. factorizations [CHI^{+}23, HMQO23]. factors [CPYC21, LLTZ24, LXKW23, SLL^{+}23, UYÖ^{+}22, WZ21, ZK24]. fail [BPB21]. fail-operational [BPB21]. failed [DMPR24]. failure [ACJ21a, CYZ^{+}21, JHZ20, XZY^{+}22, XYY24, ZMLW23]. failure-aware [ACJ21a]. failures [FDY21]. fair [LFX^{+}20]. fairness [JW21, MBO^{+}21, PLZL24, SB23b]. fairness-aware [PLZL24]. fake [DBD23, JR22, KBK^{+}22, KUK^{+}24]. fall [KTM22, SHVA23]. False [TAT22, TPT^{+}22]. family [IAT^{+}23, YY22]. FANET [AD22b]. FANETs [SSV24]. Fang [RKK23]. far [WWM^{+}23]. far-field [WWM^{+}23]. farming [SP22c]. farmland [GFA21]. farms [KWL^{+}24]. Fast [AMBAJ22, BBF21, BCP^{+}23, DPFC20, HNOI24, HXST22, HXY^{+}22b, ZZWZ21, dRdSZ^{+}23, BEKS22, Ben21, CWL^{+}21, CLX^{+}21, CWL^{+}22, IAASK23, JZL21, KJHM21, LLC^{+}22b, MRAS^{+}23, PSP^{+}20, PS23b, RSJ22, SS23a, SRGK24, SR21, SKV22, WWF^{+}22, XCY22, ZTM21, ACVK23]. fast-evolving [MRAS^{+}23]. faster [SMKA22]. fat [WL23]. fat-trees [WL23]. FATM [ACJ21a]. Fault [BLT22, DR21, FHH^{+}20, FDD^{+}24, KMY24, ANP^{+}20, AM20b, ACJ21a, AFNH21, BSN22, CLE^{+}20, CDC20, HLC23, ITO21, JAR^{+}24, Jia22, LAC21, MB21, QNZP22, SM23a, SKV22, VG20, ZLV22]. Fault-tolerance [BLT22, FHH^{+}20, CLE^{+}20]. Fault-tolerant [KMY24, AM20b, ITO21, Jia22, MB21, QNZP22, VG20, ZLV22]. faulty [NQ21]. FBO [VM23]. FBO-RNN [VM23]. FCM [KSS21]. FDI [TPT^{+}22]. feasibility [KCL^{+}20]. Feature [Alm22, BD22, DPYS22, Gün23b, LN20, Li21, LLC^{+}21b, MS22b, RT21, TYLY23, ATS22, ASAAAA22, APM^{+}21, MS22a, Anb22, AI22, AV22b, AT22, AS23c, AMA22, CO21, BST^{+}22, BEKS22, BDK22, BBD24, BM22c, BYL20, CU22, CKKK20, CEN22, CSC22, CWL^{+}21, CLX^{+}21, CWL^{+}22, CLY^{+}21, CT22, CB22, Cob22, DDZ^{+}20, DLC^{+}21, DSC^{+}21b, Ele22, FEH22, FMJ^{+}22, FHS^{+}22, FDD^{+}24, GZC^{+}22, GRMP23, GK22a, GRG^{+}22, HK22a, HWBZ21, HLW^{+}21, HZZC22, IAAA22, JR22, JB22a, JM22, JYW^{+}20, JYC^{+}21, JTY^{+}21, KEK^{+}20, KTM22, Kay22b, KG22, KSSP22, KBS^{+}22, LLZ^{+}21a, LLW^{+}23, LWL^{+}24, LLM^{+}24, LM24, MQEK21, MPV22, Mis22, NR23, NBHN22, Pan24, Par22, PR22, PKB22, PKSM24, RKR22, RMS22, SSMT22b, SHL^{+}22, SM22a, SRL23, SS22a, SV21, SS24, SGH23, ST23c, SBSK22, SSK23b, SSDY20, SSP24, SBB^{+}20, SLHW20, SBS22, SM23d, VY23]. feature [VLVS22, VS21, WZY^{+}22, WER^{+}23, WRE^{+}23, WHH^{+}20, WLY^{+}22, XY21, XLY^{+}23, YLZW23, Yl20b, YLWL20, YSL^{+}21b, ZX20, ZYCS22, ZZQ^{+}22, ZRY^{+}22, ZKZ^{+}23, vdSTC21]. feature-based [FMJ^{+}22, vdSTC21]. feature-reinforced [Ele22]. feature-rich [KG22]. features [AK22b, AAA22b, AEM22, BEKS22, BT21, Boz22b, CMLL21, CLL^{+}21, DBN^{+}22, DRM22a, FLP20, GS22, GP22b, GSB21, HLC23, JW22, KDC22, Kum22, LW20b, LWC^{+}21, LWCM21, LPZ^{+}22, LLN^{+}23, NNVD22, ÖUG22, SVD22a, TUD21, TSD23, VPSM22, XJW^{+}22, YLS^{+}23, ZLC^{+}22]. features-based [Boz22b, LPZ^{+}22]. Fed [WZMJ23b]. Federated [ZZC24, BJK^{+}24, BF24, DLG24, HDS^{+}23, JLZ^{+}24, LLW^{+}23, LZM^{+}23, LZLZ23, MA22, PCG^{+}21, SZW^{+}22, THX22, WZMJ23b, ZCZ^{+}22, ZTP^{+}23, ZYY^{+}24]. federation [AASPR22]. feed [SS24, YYZS22]. feed-forward [YYZS22]. feedback [GPR^{+}22, JMV24, KUK^{+}24, SBGC21, YGZ^{+}21]. Feedbacks [DJST24]. Fengyun [ZWJG21]. FERNet [JB22a]. fertile [SM23c]. fertility [GFA21, RM21]. few [Abb23, ZLM22]. few-shot [Abb23]. FFANN [DÖD22]. FFNet [HLW^{+}21]. FFT [VMFL23]. FH [LWZ22a]. FH-entropy [LWZ22a]. Fi [MBM20a, MES23, NLG^{+}20, QZB^{+}23]. fiber [CCZM23]. fibrillation [AST22]. Fick [LWZ^{+}22b]. Fick-spreading [LWZ^{+}22b]. field [BKK23, TAB21, WWM^{+}23]. field-long [BKK23]. Fields [ZLZ^{+}22a, AE22a, NKY23]. fight [JD22]. Figure [YÖT^{+}22]. File [LGW^{+}22b, PUL20, CWL^{+}20, FLP20, GLJ20, IP20, SS23e, SS23g, TSA21, WWW^{+}23, ZCD^{+}22]. files [SR24, TA23b]. filter [DCM21, FN23, JYW^{+}20, KAQ23, KS24, LYW^{+}21, LSQW21, MKRK23, SPSP23, WMY^{+}21]. filtering [AZI20, ACJ21b, CWL^{+}21, CWL^{+}22, JPK22, KM21a, LXYY21, LGDW22, MH23, NLB22, NNJC23, WWG^{+}20, Yil21b]. filters [HC22, SR21]. final [VBP^{+}24]. financial [FSWW21, HNG22, HXZH21, YS22, YYY^{+}23]. finch [SSN23]. find [CS23, KSCL21]. finding [CZUPA23, UPGCA22]. Fine [CHI^{+}23, TT23a, LCL^{+}22, LZLZ23, MCNR20, ÖTT23]. Fine-grain [CHI^{+}23]. Fine-grained [TT23a, LCL^{+}22, LZLZ23, MCNR20]. fine-tuned [ÖTT23]. finger [DR22a]. fingerprint [Li21, MRK^{+}23]. finite [CdOO^{+}20, HUC^{+}22, NKY23, SPSP23, ZBY22]. finite-difference [CdOO^{+}20]. FIR [SR21]. fire [CK21a, UE22, PCG^{+}21]. Firefly [Liu24a, CDC20, KT20, KS22a, LZT^{+}23, MHB^{+}22b, PZD^{+}21, SM23a, Yu21, PI21]. firewalls [LHX^{+}23]. firmware [GDCW22]. first [GHL^{+}23, Ker22, LML^{+}23, MSPPD20, VBP^{+}24]. first-arrival [LML^{+}23]. fish [KA22c]. fitness [CA22, DÖD22, KT23, SP23a, Zhu22b]. fitness-based [CA22]. fitting [ZCW^{+}23]. five [KS22c]. five-element [KS22c]. FKGA [PI21]. flame [CCGN20, KS22d, NBPR22, SYG22, VB22b, WDW^{+}22]. flamingo [SBS22]. flash [CPLX21, ZWX21]. flash-based [CPLX21]. Flex [RMS24]. Flexible [BSPM23, PB22a, TWXL21, EI22, HGX^{+}22, LZQ^{+}22, MNDK22, PM23a, XHST20, ZSX21, ZHDW24]. flight [DBD23, LL24, SSK23b]. Flink [ZQL^{+}21]. flipping [IAAA22]. Floating [DKC^{+}21]. Flow [LZXG22, BCP^{+}23, BMJ^{+}20, CZ24, DBS^{+}22, GDGN23, HGDD20, JCL^{+}20, KÖ22, VK24, LGLZ20, NET20, NET21a, NET21b, NTK22, NWZ^{+}21, RP21c, Sin23, XLZL22, ZYZC23, CLX^{+}21]. flow-oriented [Sin23]. flower [GMM22, MSK22, SR20a, KKM20]. flows [AAMM23]. flowtime [RSM23]. flowtime-aware-cost [RSM23]. fluctuations [CBR22]. fluid [FZ23, NdSSSN20, XR21]. fly [BCK22, KR23, PMP23, ST22b, YPO21]. fly-based [KR23]. Flying [JZR22, AK21]. FMap [AAMM23]. Focus [HMK23, AY23, VRB21]. focusing [AHS20]. Fog [Agr21, JPH23, KKK21, AEMS24, ARC22, ACC^{+}20, DR21, FSA^{+}22, GLJ20, HSR23, JB22b, JSA^{+}20, JR24, KK22a, KGRG24, KA23, KV24, KT22b, KS24, LXW^{+}23, LS24, LF23, PJP21, SG22a, SS23c, SV22c, ST22c, SKS20, SKSN22, SHV24, SK21b, TSR22, TT23b, UK23, VB22a, WWA22, YV23, YR21, ZWW^{+}21, BP24]. fog-assisted [LF23]. fog-based [WWA22]. fog-centric [YR21]. fog-cloud [KA23, SK21b]. fog-driven [KGRG24]. fog-enabled [SG22a]. fog-integrated [SV22c]. foggy [KCM^{+}22]. FogiRecruiter [SG22a]. FogSec [SHV24]. food [Ano21f, ABC^{+}21, SR23]. foot [DRM22a]. footage [DDUK23]. football [KBJ21]. forage [Tur23]. foraging [LK22a, YS22]. Force [NMS^{+}21, TZZ^{+}23]. forces [ZZS^{+}21b]. Forecasting [BPAE20, The21, Tur23, WWG21, AHT^{+}20, CBR22, CCD^{+}20b, CLL^{+}21, DL23, DM22d, GCP22, JS22b, KKR23, KA21b, KNK22, KSJ23b, LWLZ20, LST22, LLL23, LLA^{+}22, NBK22, NWZ^{+}21, PKK23, SMR23, SSI22, THT20, UYÖ^{+}22, Yal22, YWQ^{+}21, ZCH^{+}23]. foreground [NBS^{+}22]. forensic [HAAF22]. Forensics [KHEF22, AKZA22, CPA22, Kum21, NM24, PCC^{+}21, RRB22, VRS22]. Forest [CK21a, SJ23b, AMJK21, Aka22b, AKA22c, ATS22, JSAA22, MHB^{+}22b, SKB^{+}20, WQY^{+}22]. forest-based [Aka22b]. forests [MCT22a].

Foreword [CSW20, FD20a, LXZ20, RT20, SBG20, WJS21]. forgery [PKSM24, SSN23]. forget [ADK^{+}23, NdMP22]. forgetting [WEH^{+}22]. Forgive [ADK^{+}23]. form [IA23]. Formal [KQK^{+}20, MM22, NMS^{+}21, YPY21]. formalism [AYKE21]. formalisms [USK23]. format [CWH^{+}24, GCF^{+}20, XWD^{+}22]. formation [ACC^{+}20, CPYC21, KTM22, KABE^{+}20, LJC23, LL24, QHL^{+}24, SAPC21, SMD22b, TS21, xZlGCzJ20]. formats [IP20, SA22b]. forming [NKST24]. formula [DPB22a]. formulas [DPB22a]. formulation [GMN21]. forward [HMK24, SS24, YYZS22, ZLV22]. forwarding [Li20, SSV24, TAH22]. fostered [PBK23, PSK23]. fostering [BTP^{+}21]. four [CCZ^{+}21]. four-wheel [CCZ^{+}21]. Fourier [ACVK23, ZSQ22]. FP [XYY24]. FP-JSC [XYY24]. FPDAPP [Ano21-41]. FPGA [AY21, BCP^{+}23, Cho20, JNMG21, KIN^{+}23, NMS^{+}22, PBC^{+}23, PQKDT21, SR21, SZZ^{+}24a, WMY^{+}21, WYE23, XAC^{+}20, ZYH^{+}23]. FPGA-based [NMS^{+}22, AY21, PQKDT21]. FPGAs [RCBTGE23]. FPHO [PST24]. fractal [DÖD22]. fractals [IAASK23]. Fractional [PS23c, RW24, KUK22, Kha22a, PKR22, VR22, PST24]. fracture [KPP^{+}22]. fragile [RÖ22]. fragmentation [KMS^{+}21]. fragments [DKH20, HM20, TSA21, TSA21]. Fragments-Expert [TSA21]. frame [DSJ22, HLS^{+}20, KB22, MNYN21, SSN23]. frames [JT23, ZGG^{+}23]. Framework [ABSS22, PS22c, AS22a, AFG^{+}22, Anb22, Ano21-40, AMA22, BAPS22, BV22a, BTP^{+}21, BFR24, BCK22, Boz22b, BCP^{+}23, BMJ^{+}20, CLDY21, Cza23, DLG24, EI22, FPXM21, GDCW22, GJBM22, GVSS22, GM21, GRG^{+}22, UZAA21, HAAF22, JPN21, JNM^{+}22, KTM22, KHEF22, Kar22, KRK21, KKS22, KBK^{+}22, KAQ23, KGE^{+}20, KR22b, KDS^{+}23, KST23, LJ22, LLZ^{+}20, LBFT22, LWZ22a, LZY^{+}20, LZW^{+}21, LBZ^{+}22, MBB22, MG20, MCT22a, MAW^{+}22, MGGA20, MPLW24, MJAM23, MBO^{+}21, MCT22b, NS23, NSBT21, NRP^{+}20, NR23, OBTC20, PPA22, PPM^{+}20, PS23a, PKB^{+}23, PKKL21, PEGP23, PM22, PSK23, RVAE21, RYG^{+}21, SSMT22a, SSMT22b, SHL^{+}22, SKB^{+}20, SG22b, SZY^{+}22, ST23c, SR23, SSI22, SPK22b, SDR20, TK22b, UAS^{+}22, VH22, VD21, VSK22, WCL^{+}24, WHZL21, WSZ^{+}23, WZMJ23b, XAC^{+}20, XSZ^{+}20, YZX^{+}22, YLZT23, YY20a, YCY20, YLZ20, YLW^{+}21, ZARR23]. framework [ZABT^{+}20, ZTP^{+}23, ZYY^{+}24, ZZZ^{+}20, ZQX^{+}23, ZHT^{+}23, dCMA23, uHAU23]. frameworks [AFK^{+}22, AKRR20, LMM23, PKK21]. fraud [GCS23, SG21, YS22]. fraudulent [RVF22]. Fréchet [Eke22]. free [AEMS24, DE20, JWT^{+}20, LHK22, RVF22, ZJ21]. Frei [RÖ22]. frequency [ADGT22, Ano21-37, Ben22, EK20, JYW^{+}20, KK23a, wSYyCsD23]. Frequent [TSR22, CZUPA23, Dev21, FEH22, LM21, LK22a, NWW^{+}22]. friendship [BV22a, SJ22]. fringes [TLH^{+}22]. frog [Ben22]. front [Shi22b]. front-to-back [Shi22b]. frontal [NNX^{+}23]. frontier [JH21, VBP^{+}24]. frontiers [AYH20]. frost [GNMELC21]. fruit [AA23e, BCK22, GK23b, KR23, ST22b]. fruits [GNMELC21]. FS [NR23]. FSOA [NSR22]. FSOA-DNFNet [NSR22]. FTPS [ZLV22]. FUCBR [AP23]. fuel [ZCR23]. full [SN22a]. full-adder [SN22a]. Fully [AMM^{+}20, ÇG21, KMS^{+}22, LWW23, MIN^{+}23, PZZ^{+}23, VRR^{+}22, Yil22b]. function [AAN^{+}21, ARC22, AJP22, CWLL20, DM22d, HUC^{+}22, KPJ^{+}21, KT23, KEMZ22, LZC21, LGT^{+}23, MG22, NT23, OKJ^{+}21, Pan23, PL21, RCK22, SP23a, VJ22, VP22a, WLX21, ZHX^{+}21, ZCR23]. function-based [PL21]. functioning [HKMS21]. functions [JXL^{+}23, LFWJ22, MLZ21b, MAAA22, PB21, RP21b, RKK23]. fund [HXZH21]. Fundamental [DJF21, IA23]. fundus [KK22b, LK22b, PGD^{+}22, SBA22]. fungal [CÇ22a]. Fused [DGB24, SV22e, AK22b, Aka22b, KH22, XWQ^{+}23]. Fusion [DRM22a, WZHL21, APM^{+}21, MS22a, AJK^{+}23, AEM22, BST^{+}22, CWL^{+}21, CWL^{+}22, CMLL21, DP22a, DAK22, DSS^{+}23, FHS^{+}22, GZC^{+}22, HWBZ21, HLC23, HLW^{+}21, HZZC22, JM23, KEK^{+}20, KDC22, KCL^{+}20, LWZ^{+}22b, LCW^{+}23a, LCL^{+}20, LLN^{+}23, MPV22, PR22, PHDS22, SHL^{+}22, SS23c, SZG^{+}24, SS22f, WWZ^{+}22, WZ20b, WN23, YCL^{+}22a, YLWL20, ZYCS22, ZZQ^{+}22, ZX23a, ZGG^{+}23, ZDJ^{+}21, Zhu22a]. fusion-based [KDC22, ZX23a]. FuSTM [Ano21g]. Future [Ano21-41, ASB23, BV22a, CWDM^{+}21, CFR^{+}21, CR23, GLW21, HPCK22, KKK21, KK22d, KK23b, MKS^{+}20, PKK21, PJ21, WZ24, YSP23]. futuristic [BKGC23]. fuzz [GDCW22]. fuzzing [GDCW22, KSCL21]. Fuzzy [AP23, BYL20, DRV22, JR22, NSR22, SKS^{+}23, SM24c, VM23, AP22, AAMM23, APM22, AD22b, ABCP23, Ano21g, BM22c, DS21, DCR23, DR23, DP22b, DG21, DBB24, GK23a, HH23, HR22b, JM23, JSS22, JKP22, JM22, KSS21, KUK22, KRKM22, KV22a, KV22b, KSTV21, KS23b, LK22a, LWC^{+}21, LXC^{+}22a, MGSB23, MRAM^{+}21, MCNR20, PPA22, PMS^{+}21, PSP24, PCR21, PB21, RPPK23, RG23, RKC22, SYRS^{+}22, SP22b, SMKA22, SP22c, SKSP23, SHVA23, TWW^{+}21, VR22, WJL^{+}20, WAY^{+}21, WER^{+}23, YV23, YB23b, ZTMC23]. fuzzy-analytic [PMS^{+}21]. Fuzzy-based [AP23, SKS^{+}23, DCR23, LXC^{+}22a, PCR21, YV23]. fuzzy-region [JM23]. Fw [GDCW22]. Fw-fuzz [GDCW22]. FWM [SLLMTS^{+}24].

GA [KSTV21, LLJ^{+}22b]. GA-ELM [LLJ^{+}22b]. GA-fuzzy [KSTV21]. Gabor [FLG^{+}22, KSVP22, ZLZ^{+}22a]. Gabriel [SSH22]. gain [ZLCS21]. gaining [KT22b]. game [ACC^{+}20, BPW^{+}20, CWW^{+}23b, DR21, FCZ20, GWGR20, HML21, K22, LLSC22, LWCZ22, QHL^{+}24, RNRK22, ST22c, Shi22b, YJY^{+}21, YWC^{+}23, ZOS^{+}21]. game-theoretic [ACC^{+}20]. game-theoretical [CWW^{+}23b]. games [DMS^{+}21]. GAMESS [SSPG20]. gamma [SU23, MdAA^{+}21]. GAN [CXXW24, JAM24, SKP23]. garbage [LLYZ23, ZWX21]. Garden [LS20, WNZ24]. Garson [HUJ22]. Gas [KSA22, PKSM24, AA23e, BRL^{+}20, ODK^{+}23, SPC^{+}21, WSL^{+}20]. gastric [CGS^{+}21, LZY^{+}21]. gate [AM20b, SN22a, TAB21]. gated [FRS^{+}23]. gates [QNZP22]. gateway [BSML21, CYC21, CJP^{+}21, EABZB21, NRMB23, OMA^{+}23, PKC^{+}20, PKB^{+}23, SPJI^{+}21, VZR^{+}21, VCS^{+}21]. Gateways [KS21a, CWDM^{+}21, KLJ21, MPLW24]. gateways/virtual [CWDM^{+}21]. gathering [HNOI24, PB22b]. gauging [PS22a]. Gaussian [JD23, LAKA21, MNR^{+}22, PP21, WHC^{+}20, ZLYS21]. GCR [YKL23]. GCR-Net [YKL23]. GEMMs [DGB24]. gender [Kay22a]. Gene [AV22a, PBC^{+}23, SS22a, EHST21, PS22b, MR23a, PVRM22, XFH^{+}21]. gene-expression [PVRM22]. General [MdAA^{+}21, AYKE21, KJMB22, WZL^{+}22]. generalization [HLC23, LDS^{+}23]. generalize [XPLX23]. Generalized [SB23b, Çet23, IMNC22, KAQ23, LWC^{+}21, PP20b, WBL22, YK23]. generate [GK23b, HD23, LNI^{+}24, PMC^{+}21, SS23a]. generated [CCPP21, LTL^{+}20, SSS23a]. generating [KST23]. Generation [BA20, PI21, SV22a, AE22a, AWS^{+}22, DBPC22, DSS^{+}23, ETH^{+}23, HJT^{+}20, HWG23, HIN23, LS23a, LK22a, LHK22, LK22c, MTK^{+}21, MSPPD20, NNX^{+}23, PR23, RBC20, SSSP21, SM21, TPT^{+}22, VNP^{+}23, WWG21, WS21, WC22, XAC^{+}20, XW24, XFPY24, Yil21b, ZFF^{+}21, ZQW^{+}21, ZPL21, ZLX^{+}24]. Generative [AGM23, WZSZ20, DSSS22, DCWM20, Eke22, JAM24, KH22, KSJ23a, LG23, LSD21, LLL^{+}22, LWHW22, NGD^{+}22, PKK23, PSK23, ST23a, SYT^{+}21, VKSS23, ZDH^{+}22, ZLT21, ZCZ^{+}22]. generator [BNB22, CZ22, WPL20]. generic [APV23, OBTC20, SNGK21, WMC21]. Genetic [PWJ^{+}21, Yil21c, AB20a, AFF22, CHMC21, DCWM20, GSG^{+}23, GK22b, IAA20, JPL22, KA21a, KEMZ22, KJHM21, KWL^{+}24, KK22e, LZL^{+}20a, MSJ22, MLZ21b, NLB22, AST22, SEMA^{+}22, SP23a, SCP20, SP22b, TAB21, TÖK21, USI21, WRJ20, Wan20, WN21, XCY22]. genetic-based [NLB22]. genetic-simulated [SCP20]. genetics [TCW^{+}23]. genome [ZZS21a]. genomes [YCL^{+}22b]. genres [Kum22]. geo [CON23, SWS^{+}24, YYM^{+}24]. geo-distributed [SWS^{+}24]. geo-indistinguishability [YYM^{+}24]. geo-replicated [CON23]. geodetic [KGK22]. geographic [AK21, CL22b, LH21, SSV24]. geographic-based [AK21]. geographical [KHY^{+}20]. geographically [ASJ21, BZT23b]. geohash [MKL21]. geoid [AKA22c, KAO22]. geomagnetic [BBB^{+}20c]. Geometric [YSLX22, AAEA20, DPBP24, PB21, RKK23]. geometrical [PHZ^{+}22b]. georeferencing [MRAS^{+}23]. geoscience [VZR^{+}21]. geospatial [WZB21]. gesticulated [YKL23]. Gesture [BKJ22, CLY^{+}21, DSC^{+}21b, JPA^{+}23, JKB22, KSK^{+}22, LLW^{+}21, PT22, YLWL20]. gestures [GMP^{+}20]. GFDL [WWG21]. GFDL-CM3 [WWG21]. GHub [SPJI^{+}21]. Giant [ECIB20]. Gibbs [KT20]. Gill [DS23b]. GIS [SM21]. GitHub [DB23]. given [WPL20]. GKM [RC22]. glacial [TA22a]. glaciology [SPJI^{+}21]. gland [SKÇA23]. glaucoma [NSBT21, RPPK23]. GLeSI [TA22a]. glioblastoma [DBN^{+}22]. Global [ST21, Den24, FKO22, HNG22, HZW^{+}23, HZZC22, KAO22, LW20b, LMM^{+}22, Rav23, RMS22, SY23, SKHL22, Shu22, SSK23b, WJL^{+}20, XGCZ23, ZZY^{+}22b, ZFH^{+}23, ZZLZ21, ZCW^{+}23]. globally [MGB^{+}23]. glowworm [RP22]. GMM [JYC^{+}21]. GMM-UBM [JYC^{+}21]. GNN [KS23b]. GNN-based [KS23b]. go [DKA^{+}21]. goal [KD22]. gold [BEÖ23]. Goldbach [JOW23]. golf [HKP21]. Golub [KSA^{+}21]. Google [RM22, UMR23]. gossip [BM22b, YBJ^{+}23]. gossip-based [YBJ^{+}23]. governor [KSKS22]. GPGPUs [PJKO24]. GPR [XGZ^{+}20]. GPS [ARHT21, LWCM21, RT22c, WMC21, YYM^{+}24]. GPU [APTT22, BAR21, BVP22, CLYG22, CS21, CWH^{+}24, Cza23, DPBP24, DS22b, DGB24, DE20, FD22, GRL^{+}22, HVT23, HSO^{+}21, HMR24, HMB22, HYG20, HVB22, IPRS21, KHPH20, LMM23, LYI^{+}20, MTK^{+}21, MM21, NdSSSN20, NWT21, PK22b, RBC20, RGM22, RT22c, Shi22b, SD22, STH^{+}20, SSPG20, TNIK23, TWH^{+}22, VCFZ20, XAC^{+}20, XWD^{+}22, XR21, YLT^{+}21, ZZL^{+}20, QMC^{+}20]. GPU-accelerated [DS22b, MM21, PK22b, LYI^{+}20]. GPU-based [HVT23]. GPU-clusters [NWT21]. GPU-FPGA [XAC^{+}20]. GPU-powered [CLYG22]. GPUAnimats [QPS20]. GPUs [BABS21, BP20, BFM^{+}23, DAW22, GGCGS20, INY^{+}23, IGYS23, JK22a, JSYAA20, KW21, KCL^{+}20, LALMGLG20, SMCM^{+}22, TCA23, YKW20]. Grab [SKP23]. graceful [ZHWY22]. grade [OSK23]. Gradient [RKK23, AFK^{+}22, Anb22, AV22b, CD24b, Jeo20, KAO22, PS23b, SSMT22b, SV22b, VPSM22, VR22, WWW^{+}20]. gradient-based [Jeo20, KAO22, SSMT22b]. grain [CHI^{+}23, LM22a]. grained [LCL^{+}22, LZLZ23, MCNR20, TT23a]. Grammar [ZLCL21, KEMZ22]. grammar-based [KEMZ22]. granularity [CMS21, FKO22]. Graph [Ano21h, CBFS23, JSLL20, LLJ^{+}22a, LGW^{+}22a, RMS24, ST21, XZL^{+}20b, APM22, AOACAQ21, AV23, AT22, CPQ^{+}22, CKKK20, CYQ^{+}20, CCL^{+}22, CCD^{+}20b, DS20a, GLW21, GLW22, GRL^{+}22, GQ24, GVSS22, HCH^{+}21, HHC^{+}22, HYG^{+}23, JZL22, JZL21, JW22, KGM23, KKM21, LM21, LZY^{+}20, LPC^{+}21, MLS^{+}24, MdARS^{+}23, RA21, RAK22, RZ23, Shi22a, SSV24, SSP24, SSH22, ST22d, TNI23, TLQ21, TL21, VVK23, WZC^{+}22a, WZMJ23a, XGX^{+}21, XLWX20, Yil23, yZyWD^{+}21, dRdSZ^{+}23]. Graph-based [Ano21h, AOACAQ21, AV23, CCL^{+}22, LPC^{+}21, Shi22a, SSV24, TLQ21]. Graphic [LZC21, yXlLyGX21]. graphical [DBS^{+}22, QPS20, TSA21, USI21]. Graphics [HIN23, JPL22, AS22c, AAG^{+}22, AAQOT23, FJCFB23, GR22a, HCG21, MT22, NdMP22, PRS23, SPC^{+}21, SAF^{+}23, SRG^{+}21, TNP21]. graphs [AS22a, AMD20, LM21, LYBZ23, LK22c, RVF22, TLQ21, YNK^{+}23]. GraphWare [CCL^{+}22]. grasp [WSJ^{+}21]. grass [APM^{+}21]. grasshopper [KA22b, SKSP23]. Gravitational [SS22d, SMBM24]. gray [AS23a, ASC22, AB21, GP22a, KV24, KGGM22, MCT22b, TS22, NNVD22, PVRM22, UDS21]. grayscale [Gul22]. Greedy [ARC22, CL22b, JXL^{+}23, KS22a, MTK^{+}21, MT22, SS22a]. Green [FSWW21, WLCW20, CZ21, MK22a, XA22, YFdP^{+}23, MKS^{+}20]. Grey [BJGF20, CZ24, DS20b, DPSJ22, SFJ^{+}21, TB23]. grid [AAT21, BXH^{+}23, BTDJ21, KC22, LLL23, NNVD22, NKY23, NA22c, OE22b, SG24, ZTF^{+}20, gZWfL^{+}20]. grid-based [KC22]. GRIDHPC [FEK20]. grids [AFG^{+}22, KDFK23, SAN24, SR20b]. grooming [RSJ22]. Grossberg [TSLL21]. ground [Yil21b]. Groundwater [MCR^{+}23]. Group [WJJ23, CMA^{+}21, Den24, GZC^{+}21, LZL^{+}22, LDZ^{+}22, MG22, NAR^{+}22, PA21, SG21, ZWCS20, RT20]. Group-guided [WJJ23]. group-key [PA21]. grouping [AAB23, GGJ^{+}24, LZL^{+}23]. growing [BTT21, JM23]. growth [YL20a]. GRU [LLL23, YA22b]. GRU1 [YA22b]. GRU2 [YA22b]. guarantee [HWG23, HLT23]. guaranteed [KB21]. Guaranteeing [XDH^{+}20]. guide [SKB^{+}20]. guided [DM22c, GDCW22, LYW^{+}21, LZSC23, MNDK22, VCS^{+}21, WJJ23, ZLCL21]. Guidelines [NdMP22]. guiding [BEKS22, vdSTC21]. Gulf [GDA22]. Gustafson [RM21]. GV [SM21]. GWO [DPSJ22, TK24]. gyrokinetic [ALB^{+}20].

H.265 [CC22b]. H.265/HEVC [CC22b]. H.323 [EABZB21]. habituation [ZHX^{+}23]. Hadamard [QNZP22]. Hadoop [BYPO21, IP20, SB23b, SS23g, VRS22, VR22, YTN^{+}20, ZGLS21]. Hadoop-big [VR22]. HAFTA [ITO21]. Hajj [AR22a]. half [CH21, IG24]. half-life [CH21]. halting [XHST20]. Hand [YLWL20, BT21, PT22, PR22, YKL23]. handcrafted [Boz22b, DRM22a]. handle [KSY24]. Handling [CS22, GHT20, VPGK23]. Handoff [MM24, IVP^{+}23]. Handover [CÇ22c, AA23a, IVP^{+}23, KS23b, PDRM24]. handovers [KGM23]. handshake [TTM^{+}22]. handwritten [DM21, KE21]. hard [LNI^{+}24, SS22f]. Hardware [CL22a, FGL^{+}20, DE22, GGA24, GMK^{+}21, MS21a, RS21, SZZ^{+}24a, SGHL20, TAB21, TBNPQ23, TWH^{+}22, WL23, YFdP^{+}23]. hardware-accelerated [DE22]. harmonic [KG23]. harness [KLJ21]. Harnessing [CXX^{+}22, Ano21-42]. Harris [ND23, VG21]. harvesting [AAT21, SSM^{+}21]. Hash [NVV24, Akb22, BS23a, EASR22, KB22, LMR22, PL21, SSS23a, SJ23b, SR22d, ZHX^{+}21, ZCZC23]. hash-based [EASR22]. hashing [AS22a, CY24, CYW^{+}22, HDS^{+}23, JKS20, LLT21, NMQ22, WQY^{+}22, ZHX^{+}21]. hashing-based [CYW^{+}22, WQY^{+}22]. hasten [RG22]. hate [CMJC23]. haul [SLLMTS^{+}24]. Hawai‘i [MCR^{+}23]. Hawk [VG21, ND23]. Hawks [PST24]. haze [LZF20, PP21, xZlGCzJ20]. hazelnut [AA24]. HC [AAK^{+}21b]. HCA [TRM^{+}24]. HCMonitor [SZL^{+}22]. HDF5 [MSB^{+}20]. head [LHWT20, LDS^{+}23, MS21c, SPS23, SM24c, TNP21]. healing [RS23]. Health [dCJBP20, ANP^{+}20, ALR22, BCC^{+}21, DM22c, GK22a, GB23, HK22b, HTVL22, KS24, LBFT22, Liu21, NA22a, ÖSTY22, SAC22, SC22b, SHZY20, wSYyCsD23, YLM21]. health-care [KS24]. Healthcare [SG21, AM22, AYJ^{+}22, BM22c, BVS20, BJK^{+}24, CM20, DPB22b, DK21, EI22, HT21, HK22b, JDLP22, KRK21, KS20, LA22, LJB22, NS23, OMA^{+}23, PLP22, Shi22a, SK22d, WGYZ22, WWA22, YR21]. Heap [KR22c]. Heap-based [KR22c]. hearing [RD23, SS21]. hearing-impaired [SS21]. Heart [MPV22, AI22, BJK^{+}24, LJ22, Yil22a]. heat [FDY21, SEC22]. heating [LM22b]. hedge [HV21]. hedonic [ZNDA22]. height [YGS^{+}23]. Hellman [LF23, NR22]. HELLP [MRAM^{+}21]. helmet [HFH^{+}21]. hemorrhage [JSAA22, KMS^{+}22]. hemorrhages [SBA22]. Henry [PKSM24, AA23e, ODK^{+}23]. hepatitis [KA22c]. herd [PBK23, ZCR23]. herding [BKD22, RKC22, SSK23b]. heritage [CCdCC21]. Hermitian [WPL20]. Heterogeneity [LM20a]. Heterogeneous [DKB20, HSO^{+}21, KLL^{+}21, LM20a, OCR^{+}23, PCR21, XZD^{+}21, YZPL21, AB20a, AET^{+}22, AA21, AAK^{+}21b, ABB22, Ano21j, Ano21-35, Bad23, BYPO21, BV22a, BA24, BVP22, BPK24, CLLB20, CMS21, CMS24, Cud20, DSC^{+}21a, DR20, FFG^{+}23, FPXM21, FKGO22, GFPGT21, GS21, GTM24, HMB22, IS23, JBBH21, JSP20a, JSP20b, JNMG21, JR23a, KGM23, KFML20, KML23, LBG^{+}20, LLTZ24, LK22c, LLH^{+}20, LBZ^{+}22, LLS22, LDCD22, LDW^{+}23, LLL^{+}21, MBM20a, MAG^{+}20, MTRA23, NdSSSN20, NMS^{+}21, PR24, PB22a, PMP23, PM23b, QCC^{+}23, RGPC23, SKA23, SSN21, VS22a, VSS23, WCL^{+}24, WZL^{+}24, WLG^{+}22, WFY22b, WGDP23, WC22, XAC^{+}20, XLXZ20, ZGLS21, ZLTX21, ZZQ^{+}22, ZTY^{+}22, uHAU23]. heterogeneously [ZZL^{+}20]. HeteroPar’21 [Bad23]. HeteroPar’22 [IS23]. HeteroPP [WCL^{+}24]. HetNets [XLXZ20, Mir22]. Heuristic [ARC22, AJNS22, APP^{+}21, AJ21, CGW^{+}20, CEN22, CK24a, EK20, GMN21, GRG^{+}22, HAK22, HD23, KA21b, MG21b, NN23, RSM22, SV21, SVD22a, VPGK23, YS23]. heuristically [GM22c]. Heuristics [LALMGLG20, KAR^{+}23, RZVC21]. HEVC [CC22b, SRGK24]. HHAMA20 [KLL^{+}21]. HHFDS [PCR21]. hidden [AAK^{+}21a, BB23, RRB22, ZWW^{+}21]. hiding [Wan22a]. Hierarchical [LSSQ22, YKW20, AD22c, CSL20, CK21b, DBB24, FQD^{+}23, FFG^{+}23, GSG^{+}23, HPS23, JLE22, LZZ^{+}20, LB21, MSB^{+}20, NAR^{+}22, SZZ^{+}21, TS21, WZY^{+}22, WRE^{+}23, XJW^{+}22, ZTY^{+}22, ZCR23]. hierarchy [AMA22, KS24, MS21c, PMS^{+}21, RAN21, RG23, XLZ20]. High [DG21, FJCFB23, FFLM21, GCS20, GLRB21, KIN^{+}23, KSY23, KFKD22, MRAS^{+}23, Oh21, SP21b, TBNPQ23, TC22, MKL21, WPK^{+}22, AMBAJ22, ASAAAA22, AY21, BD22, BTDD20, BVM22, CTA^{+}23, CD22, CS21, FEK20, FPHZ19, GFPGT21, GMSM21, GNMELC21, HGX^{+}22, HLC23, HIN23, JYW^{+}20, JNMG21, dCJAAdOD21, KFML20, KML23, LAE^{+}22, LHJ22, LCKJ21, LLY^{+}23, LHPG21, MBM^{+}20b, MS21a, MNYN21, MG21a, MdARS^{+}23, Pan20, PQKDT21, RK21b, RZVC21, SMCM^{+}22, SSSR20, SKB^{+}20, SCH^{+}24, SPS22, SPKK22, SZL^{+}22, TK22a, VS21, WQY^{+}22, WMC^{+}23, Wri22, WCWG21, YLT^{+}21, YCL^{+}22a, Yil21a, ZHX^{+}23, ZQW^{+}21, ZYH^{+}23, ZZZ^{+}20, ZWZ^{+}22, dCMM21, KLL^{+}21, KR22c]. High-dimensional [SP21b, ASAAAA22, BD22, CS21, LHJ22, SCH^{+}24, WQY^{+}22, ZHX^{+}23]. High-end [KLL^{+}21]. high-frame-rate [MNYN21]. high-level [JNMG21, RK21b]. high-mixed [WCWG21]. high-order [AMBAJ22, LLY^{+}23, SMCM^{+}22]. High-Performance [KSY23, Oh21, FJCFB23, FFLM21, GCS20, GLRB21, TBNPQ23, MKL21, WPK^{+}22, AY21, GMSM21, GNMELC21, KFML20, LCKJ21, MBM^{+}20b, MG21a, MdARS^{+}23, PQKDT21, SKB^{+}20, SPS22, SPKK22, YLT^{+}21]. high-precision [YCL^{+}22a]. High-quality [DG21, HIN23, ZQW^{+}21]. high-resolution [BTDD20, LAE^{+}22, LHPG21, WMC^{+}23]. high-security [CD22]. high-speed [MS21a, ZZZ^{+}20]. High-throughput [KIN^{+}23, MRAS^{+}23]. higher [HKMS21, LLJ^{+}22a]. higher-order [LLJ^{+}22a]. Highly [APL^{+}21, ITO21, LS23b, LZZ^{+}23, PBD^{+}21, SNGK21]. HighPerMeshes [AFG^{+}22]. HIPLZ [ZBY^{+}23]. Histogram [Pal22, LXYY21, TUD21]. historic [WZ20b, WZ21, WN23, WNZ24]. Historical [CY22, OHFF20]. history [YY22]. hitting [KSS24a]. HMDE [NR23]. HMDE-FS [NR23]. HML [KCP23]. HMM [AAEA20]. Hoc [AMRH21, AK21, BKS22, CD22, DCR23, FLB23, KGGM22, LCW23b, NAR^{+}22, Ngu21, PDRM24, PA21, PD23, RGKK21, RSS20, RAN22, SK22a, SPS23, SWK22, SSH22, TT22, TK22b, WJD22]. HOL [CMJC23]. hole [PD23, SP23b, XXSL23]. holistic [WZC^{+}22b]. Home [GLM^{+}22, JP21, JHS^{+}21, LHWT20, SB21, SJ23b]. homo [WZL^{+}24]. homo- [WZL^{+}24]. homogeneity [LLM^{+}22]. homogeneous [NR23]. homologycontinuity [NGD^{+}22]. homomorphic [BV22b, LWW23, PZZ^{+}23, QWW^{+}22, ZWO^{+}20]. homophily [DGM21]. honey [DBD23, KMR22]. honeypot [NRP^{+}20]. Hop [WXCY20]. hoppers [HLO^{+}21]. horde [Rav23]. horizon [DZCL22]. horizontal [YTW^{+}24]. hospital [DM22c]. hospitalized [The21]. Host [VB24, CSWC20, CYK^{+}21, YNK^{+}23]. host-pathogen [CSWC20]. host-switch [YNK^{+}23]. host/device [CYK^{+}21]. hotel [Wan24]. Hough [CCGN20]. House [ZNDA22]. HP* [DR20]. HPC [AFK^{+}22, DHSG23, JGJ^{+}21, LMMS24, LZZ^{+}20, MTRA23, MC24, NMS^{+}22, PRPD21, WS21, YSL^{+}21a]. HPE [VBP^{+}24]. HPE/Cray [VBP^{+}24]. HSCT [SS23g]. HTLC [BS23a]. HTM [SNGK21]. HTSA [MKA23]. HTTP [NLG^{+}20]. hub [CPPP21]. Hubbard [AKLK24]. Huber [DM22d]. Hugo [TBT^{+}21]. Hugo* [TBT^{+}21]. hull [QMC^{+}20]. Human [KNM21, KDC22, KSK23, ORP21, ARHT21, AADS21, AFM22, BYTG22, CCPP21, GA22, Kay22a, KSK^{+}22, KPP^{+}22, LYW^{+}21, MTY21, OS21b, QZB^{+}23, RR23a, SLL^{+}23, TNP21, TL21, TSD23, WMC21, YKL23, ZLZ^{+}22b]. human-computer [YKL23]. human-oriented [CCPP21, OS21b]. humanoid [KBL^{+}21]. Hungarian [NJ21]. hunter [SR22b]. hunting [KBR24, SAC22]. HXPLS [SMM22]. HXPLS-TMFCC [SMM22]. Hybrid [APM^{+}21, BF22, CÇ22a, CD24b, CZ24, Dah23, DS21, DK21, DK22b, GAM23, HAK22, IAAA22, JR23b, KKR23, KS22d, LSD21, MGSB23, MQEK21, MES23, Mis22, NA22b, PLP22, RCYuRH21, RRJ23, SS23b, SM22c, SM23b, SSRA23, VG21, WL23, XGX^{+}21, AZM20, ABESEh20, AMJK21, AS23a, AAARR20, Ano21i, Ano21-36, ASMS21, ASA^{+}21, ASMK21, BKK22, BYTG22, CNG^{+}20, CÇY22, Çet23, CYZ^{+}23, CS21, CZ21, CWH^{+}24, DÖD22, DR22b, DL23, DPYS22, DS22a, DP22b, DG21, FK24, GAH^{+}22, GKÇ22, GSG^{+}23, GYZ^{+}20, GCSZ24, GMN21, GVSS22, GM22c, HK24, HXY20, HXY^{+}22a, IA22, IDA22, JA23, KMR22, KA21a, KV22b, Kay22a, KV24, KCP23, KG22, KR23, KST23, KS23c, LJ22, LW20b, Lia24, LM20b, MRGP22, MSN22, MS21a, MLZ^{+}20, MG23, MSBR23, NSR22, NN23, NAK^{+}22, NNJC23, PS22a, PT22, RC22]. hybrid [RB22, RP21a, RSS20, RR23b, RSKA23, RZCA21, SAPC21, SJ22, SMM22, SKPD23, ST22b, SK22a, SAC22, SS23d, SEC22, SV22c, SKHL22, SZS20, SS23f, SCM22, SR22c, SP23b, SFJ^{+}21, SS22g, SSPG20, TA22b, Tur23, USI21, VY23, VP22a, VS21, WWG21, WWA22, WHH^{+}23, Yal22, YZX^{+}22, YMKH22, YY22, Yüc22, ZSX21, Zhu21, PD22b]. Hybrid-based [NA22b]. hybrid-compression-format [CWH^{+}24]. hybridization [ZHW^{+}20]. hybridized [PCR21, RK23b, SS22c, SYG22]. hybridizing [SRGK24]. Hydra [DKB20]. hydrological [LSD21, WWG21]. hydrological-Elman [WWG21]. hydropower [WWG21, ZZLZ22b]. hyena [NAK^{+}22]. Hyper [AJNS22, KNK22, YS23, HK22a, SHA^{+}22, SBS22]. Hyper-Heuristic [AJNS22, YS23]. Hyper-Layer [AJNS22]. hyper-parameter [HK22a]. Hyper-parametric [KNK22]. hyper-spectral [SHA^{+}22]. hyperactivity [AAA22b]. hypercube [IG24, NQ21]. hypercubes [WL23]. hyperelliptic [NR22]. hypergraph [KDS22, QWL24]. Hyperledger [LZZ^{+}23]. hyperparameters [LLC^{+}22a, OE22b, YT21]. hypersonic [XL21]. hyperspectral [KR22a, PKB22]. hypersphere [YZ21]. hypervisor [DE22, VD21]. hyponatremia [The21].

I/O [BPM^{+}22, LCKJ21, PMP23, RZVC21, YSL^{+}21a]. I/O-intensive [BPM^{+}22]. IaaS [NGOS22, NAK^{+}22]. iceberg [SPHP21]. ICP [HWY^{+}23]. ICS [AI21]. ICS-LD [AI21]. ICTS [SKP22]. ideal [DKP^{+}20, KGGM22]. Identification [KSS24a, MCT22b, AS20, CLT^{+}21, CSL21, DRM22a, DRM22b, HT21, JBCI20, KS22b, KPM20, LG23, MCT22a, MG21b, PLX^{+}24, PMR^{+}21, RKR22, SSMT22a, VRR^{+}22, WSM^{+}20, XLX^{+}21, XPLX23, XLY^{+}23, YLLL20, ZCZC23, ZXL^{+}21a, ZJSJ20, ZJSJ21]. identifier [JB24]. identifiers [GAMT22]. identify [SVM24]. Identifying [ALNJ21, AJAA21]. Identity [LAH^{+}22, CCM22a, CDR^{+}23, GJPS23, MGN^{+}22, RVAE21, SDR23]. identity-based [CDR^{+}23]. IDF [SS22b]. IDS [CD24b]. IDSDN [AA23d]. IEEE [BFR24, XLXZ20]. If [GLM^{+}22, KLJ21]. IFATA [JMV24]. IFATA-Deep [JMV24]. IFC [HL20]. II [GRG^{+}22, DCWM20, HXY^{+}22b, WHC^{+}20, Wan20, Yil21c]. IICA [SS23d]. IIoT [PLZL24, SLD24, TJC24]. IIR [SR21]. illness [Kar22]. illumination [SZG^{+}24]. illumination-aware [SZG^{+}24]. IM [SKSB20]. IM-SSO [SKSB20]. Image [BEKS22, CPA22, JPAA21, LYSC21, LC20, SLHW20, ZYX^{+}21, AB20a, Abb23, AMAT22, AK22b, ABB22, APA22, AJK^{+}23, AJS23, AS22d, AD22c, BCM22, BABS21, BMK^{+}20, BMG22, BYL20, BP23b, CZ22, CLY^{+}21, CMLL21, DBPC22, DAK22, GDSS22, GAH^{+}22, GSTS22, Gui22, Gul22, HGNN22, Hua20, HLW^{+}21, HYT^{+}21, HSL^{+}22, HIN23, JT23, JNS22, KEK^{+}20, KE22, KR22a, KS23a, KWZ^{+}21, KKS23, KSK^{+}20, KSVP22, KK23b, KSS24b, LK22b, LYL20, LGDW22, LW20b, LZF20, LYC22, LHPG21, LCL^{+}20, LLLX20, LLL^{+}22, LL23, LLN^{+}23, LLX^{+}24, MKRK23, MTK^{+}21, NSR22, NNX^{+}23, NBHN22, OSK23, OHRS21, PSMM22, PP21, PR22, PKB22, PKSM24, RB22, RRB22, RÖ22, SM22a, SO22, SEC22, SPSP23, SLG^{+}20, SZqWZ20, SA22d, SZG^{+}24, SBB^{+}20, SAL22a, SR22d, SS22g, TSV^{+}22, VB21, WHZL21, WMS^{+}23, WLL^{+}21b, WZSZ20, XLY^{+}23, YMWA21, Yil21b, Yil22c, Yl20b, ZX21a, ZLT21]. image [ZZY^{+}22b, ZZW^{+}22, ZRY^{+}22, ZFZ^{+}20, ZDJ^{+}21, ZCW^{+}23]. Image-based [BEKS22, LYL20, NBHN22]. image-enhancement [Hua20]. imagery [TA22a]. images [AK22b, Aka22b, AMV22, BP23a, BKJ22, Bei24, BS23c, Boz22b, CÇ22a, CK21a, DSYF22, DSSS22, FMJ^{+}22, GDSS22, GP22b, GRMP23, GM22b, IK22, KKKS23, KK22b, Kot23, KPP^{+}22, LAE^{+}22, LK22b, LYW^{+}21, LLM^{+}24, LJP^{+}21, MRGP22, MRS^{+}21, MH23, MSBR23, OE22a, OE22b, PBK23, PR22, RK23b, SSMT22b, SHA^{+}22, SBA22, SGH23, SHVA23, SKÇA23, SAGS24, TSG21, TB23, VRR^{+}22, VS21, WLJ20, WWJ^{+}20, XCX^{+}20, Yil21c, YMKH22, ZLZ^{+}22b]. imaging [Kot23, Liu21, ODK^{+}23]. IMAR [LLR^{+}21]. imbalance [JYL^{+}23]. imbalanced [CS22, KSY24, Par22, SSDY20, WAY^{+}21]. immune [CK23, YL20a, YLZC23, ZLM22]. immunity [ZCR23]. Impact [HAA^{+}21, CWDM^{+}21, IP20, LYW^{+}21, TTA20, WKY22, YA22b, ZR24, uZKH^{+}20]. impaired [RD23, SS21]. Impersonation [CCM22a]. Implementation [Ano21i, BZGM22, NET21a, SKA23, SLLMTS^{+}24, TÖK21, VMFL23, APA22, Ano21d, AKRR20, BS23a, CM21, DGB24, EASN22, HCG21, Hua20, KIN^{+}23, KY23, MWS^{+}23, NGB23, QNZP22, RAK22, RT22c, SPA^{+}21, SR21, SSPG20, MKL21, WYE23, YM22]. implementations [MT22, TNI23, TNIK23, WMY^{+}21]. implemented [ASL20, Cho20, TAB21]. implementing [TD21]. Implicit [KDFK23, JWL20]. Importance [Alm22, LJZ21]. Impression [SWZW20]. improve [BM22a, BC21, FLP20, OKJ^{+}21, OCD22, SAQJ23, VS22b, ZWX21]. Improved [AI21, Anb22, BST^{+}22, BRS^{+}22, DDUK23, GAH^{+}22, GN21a, JMV24, MK22b, PMP23, SKP23, SVD22a, SKV22, SS22f, WXS^{+}23, XCX^{+}20, ZLO^{+}21, ZZQ^{+}22, BMcKGK22, BD22, BM22b, CS22, CSR24, CYZ^{+}23, DZCL22, DS22a, GBS24, GP22a, GGJ^{+}24, GM22c, HK22a, HNG22, HHP23, HWY^{+}23, HUC^{+}22, IMNC22, JM23, JSAA22, JZC^{+}23, JYH^{+}24, JWT^{+}20, KUK22, KS20, KT22b, KNK22, KSY24, KG23, LHJ22, LZL^{+}20a, LM22b, LGZ^{+}22, Lia24, LLW^{+}21, LJP^{+}21, LL23, LYG^{+}21, NSSS22, PS22b, PK22a, RM23, SA22a, SM22a, SGS21b, SCP20, SS22b, SZZ^{+}21, ST23c, TFZC23, TLL20, TYA22, VB22b, VR22, WZ20a, WRE^{+}23, WZX^{+}22, WZG^{+}21, WDW^{+}22, XLZ20, YRV^{+}23, YLLL20, ZLT21, ZTS^{+}22, WWG21]. Improvement [TK24, CK23, HHP23, SH22, SP22a, SMAG22, WT23]. Improving [Aka22b, BLB^{+}24, BTDJ21, BPB21, CPCK23, DS23b, DMPR24, GDGN23, HLH^{+}20, IPRS21, JSYAA20, JSP20a, JSP20b, KY22, LM24, MY24, MdARS^{+}23, PEGP23, RP21b, SMKA22, TTA20, WWJ^{+}20, WWW^{+}20, YT21, AC23, BYPO21, BTT21, DRM22b, DBB24, HIEH22, HZZ^{+}23, KGM23, KW21, VG21, WL23, ZTY^{+}22]. impulse [SPSP23]. imputation [SK21a]. in-depth [KW21, RS24]. in-DRAM [KSD22]. In-memory [ZLW^{+}23, LM21, WCCC20]. in-painting [RCK22]. In-the-loop [FGJ^{+}21]. incentive [XHZ^{+}21, XCZ^{+}21]. InceptionV3 [KK22b]. including [DÖD22]. incoherence [RPM22]. incoming [CMT20]. incomplete [KABE^{+}20, LSD21, SZS20]. inconsistent [ST22a]. incorporated [ZKZ^{+}23]. incorporating [HLH^{+}20]. increasing [BTP^{+}21, CK23, KML23]. increment [KMD23]. Incremental [CKKK20, NSR22, BVS20, DKH20, GD22, NSPdO21, YMZ^{+}20]. IncResNet [LML^{+}23]. independent [CCZ^{+}21, CHLD23, INY^{+}23, WY20]. Index [KKM20, CWL^{+}20, CLL^{+}21, GWA^{+}23, GWA^{+}24, Par22, wSYyCsD23, WER^{+}23, XDJY21, Yil21a]. indexing [CY24, NSR22]. India [DM22d]. Indian [GKM22, GSB21, PJK23, SMD22a, SKP22, SKP23, TG23]. indicating [LJC23]. indication [LLKS21]. indicator [Zhu21]. indices [Li20, TI22]. indigenous [AAMAA22]. indistinguishability [YYM^{+}24]. individual [CPYC21, ZYXX23]. individuals [RVJ^{+}22, YB23a]. Indoor [ZXL^{+}21b, LSZ^{+}23, LLKS21, MBM20a, PS21]. indoor-positioning [LLKS21]. induction [QWL24, Yib22a]. Industrial [AlS21, JB20, AA23c, Jeo20, JCGT24, KCP23, LXKW23, SMD^{+}21, Wan20, WYLL24, XLL^{+}23, YWC^{+}23, OO22, ZTP^{+}23]. industry [Ano21-43, BF22, GKM22, MVR23, SAPC21, SSCN23]. inertia [GC20, SDSW21]. inertial [DWZ20a, FZZZ23]. Inexpensively [ZWX21]. infant [JYC^{+}21]. Inference [Çet23, EHST21, JKP22, MRK^{+}23, NRMB23, NMQ22, NZD^{+}24, RKC22, SZS20, SP22b, YLT^{+}21, YB23b]. inflate [NMQ22]. inflated [AAA^{+}22a]. Influence [AE22b, GXL^{+}20, RRIL22, THW21, SKSB20, WWW24, xZlGCzJ20, ZXL^{+}21a]. influential [CS23, JS22c, KSS24a, PS22a]. influenza [Kar22]. influenza-like [Kar22]. inform [FGL^{+}20]. informatics [Liu21, WHZL21]. Information [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano22a, Ano22b, Ano22-27, Ano22-28, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22z, Ano23b, Ano23c, Ano23-28, Ano23-29, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j]. Information [Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano23z, Ano23-27, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, HFFA20, dCJBP20, NET20, BK22b, CPYC21, CPH20, DE20, GCQ^{+}24, HWBZ21, HUC^{+}22, HLDT^{+}23, KABE^{+}20, KS22d, KSK23, LS23a, LH21, LWZ^{+}22b, LLKS21, LSX21, LSQW21, LRC^{+}22, MLZ^{+}21a, MLS^{+}24, MAR23b, MSL22, NET21a, NET21b, Ogi20b, Ogi21, PMC^{+}21, PI21, SM23c, SKSP23, VSS23, WWZ^{+}22, YGZ^{+}21, ZARR23, ZLCS21, ZZW^{+}22, ZLYS21]. information-based [SM23c]. infrared [SZG^{+}24, ZDJ^{+}21]. Infrastructure [CPPP21, KMZ^{+}20, CCCR21, CR23, CJP^{+}21, DGP20, HL20, MYCH22, OGA^{+}22, OSMM23, SSW^{+}22, SZ24]. infrastructures [MBT^{+}20, PCG^{+}21, RGPC23]. infrequency [LFG^{+}22]. ingestion [KS24, MRAS^{+}23]. inheritance [LXZ22]. inhibitive [TWW^{+}21]. INISTA [SKE22]. initialization [AE22b, OKJ^{+}21, PS22b]. initialized [AS23a, TNP21]. initiated [MSSM24]. injection [AQP^{+}22, KJS^{+}20, TPT^{+}22]. Innovations [HFFA20, SKE22, WZ24]. Innovative [HFFA20, Ogi20a, SD23, YIB22b, YT21]. input [AM20b, Ano21a, HTZ^{+}22, RRIL22, SN22a, SYJL20]. inrush [DWZ^{+}20b, DWY^{+}21]. insertion [TT23a]. insightful [DAW22]. Insights [FGL^{+}20]. inspired [AAK^{+}21a, CO21, BC23a, DZCL22, JC21b, KHA22b, LH21, MMR^{+}22, MG21b, NAK^{+}22, OE22a, VNP^{+}23, WWA22]. Instagram [KBK^{+}22]. instance [CB22, PLX^{+}24, ZTS^{+}22]. instantaneous [DTK24]. Instillation [IIK^{+}23]. Instruction [DAW22, SCD^{+}23]. instruction-aware [SCD^{+}23]. instructions [AY23, GO22, KML21]. instrument [GSB21]. instrumentation [FHS^{+}22, VRB21]. insulation [STJ^{+}20]. integrals [GS21]. integrated [BAR21, CCBA23, CSC22, CM20, CYZ^{+}21, CA22, CLL^{+}21, GAS23, Gün23b, HNG22, Hua20, JNM^{+}22, KAO22, KA22c, ST23a, SV22c, SRRM23, SZqWZ20, UAS^{+}22, VG21, YW22, ZX20, ZZW^{+}22, RZ24]. Integrating [YJY^{+}20, ZABT^{+}20, ZLW^{+}20]. Integration [RKR22, AD22b, Ano21j, KSK23, LLMX21, LM22a, LL21, LZT^{+}23, WZ21]. integrative [VPB^{+}23]. integrity [AS24, LLJR21, PS22c, PUL20, PI21, YR21, ZARR23]. Intel [BSPM23, LHPG21, MPB^{+}22, Mit20, TCA23]. intellectual [DKK24]. Intelligence [JB20, XZ20b, ASMS21, AGSN23, BTP^{+}21, Che24, CR23, DSYF22, DAT23, EVVR21, FM22, FSdP^{+}23, GlSL^{+}23, GSS23b, HKMS21, JAM24, Jeo20, JC21a, OSMM23, SRBH22, WZMJ23a, XZ20a, YYY^{+}23, ZALM23]. intelligence-based [EVVR21, GSS23b]. intelligence-enabled [GlSL^{+}23, YYY^{+}23, ZALM23]. Intelligent [ESS23, GDS23, HDXH20, JLE22, LCW21, OS21a, OBER22, SPA^{+}21, SV22c, SKE22, WNZ24, ZHJW21, Abb23, ABZS20, AFF22, AKUA22, ALNJ21, BTDD20, BC23b, BJS23, CLDY21, CCZ^{+}21, Che24, DYF20, DLC^{+}21, FGZC23, FSWW21, GLN23, GYZ^{+}20, HXZH21, IA22, JB22b, JHG23, KCM^{+}22, LHJ22, LLAV22, LLC21a, MK22a, MES23, NBK22, NKST24, ORP21, PMS^{+}21, PK23a, PA21, RSKA23, RK23c, SMD^{+}21, SKHL22, SR22c, SZ24, TK22b, VDL23, WLCW20, WJS21, WZ20b, WN21, YIB22b, ZWC^{+}23, CSW20]. intelligently [AS23a]. Intel’s [HYT^{+}21]. intensified [DBK21]. intensity [VV23]. intensive [BPM^{+}22, CYK^{+}21, KD22, LXZ20, NSPdO21, RS20]. intention [USK23]. inter [ALBZ21, KFKD22, KKS23, SS23d, XFPY24]. inter-block [KKS23]. inter-class [KFKD22]. inter-organizational [XFPY24]. inter-resource [ALBZ21]. interaction [Che24, LLC^{+}21b, SGHL20, YKL23, YL20c]. Interactional [FGJ^{+}21]. interactions [CSWC20, HLCH20]. Interactive [DB23, KWZ^{+}21, LHJ22, LPHK20, SGS21a, SP22c, WN21]. interchange [ZTL^{+}21]. interclass [CWLL20]. interconnect [FSFM22]. interconnected [HMK23]. interconnection [LGA^{+}20, WL23, YNK^{+}23]. Interdisciplinary [JWL20]. Interest [BKM^{+}21, SJ22]. interface [CYC21, SRS^{+}21, TSA21, TO22]. interfaces [WZZ^{+}22]. Interfacing [MSB^{+}20, KNM21]. interference [FPHZ19, IAQ20, Pan20, ZHWY22]. interference-aware [IAQ20]. interior [YA22b]. interleaved [ZX21b]. internal [AY23, CPYC21]. International [LM20a, SKE22, Bad23, IS23, BI23]. Internet [CSW20, OO22, ZTP^{+}23, AGSN23, CBK23, CÇM22b, DCZ^{+}22, DBB24, DLG24, JPAA21, KHEF22, KS22b, KS23b, LXKW23, NT23, RAN21, RV23, SSRA23, UE22, ASB23, YWC^{+}23, AYH^{+}22, Agr21, AS23a, AlS21, AADS21, AYB21, CM20, CYC21, CYZX23, CWW^{+}23b, CZTC22, DDH^{+}20, DPB22b, DKL21, DP22b, DKC^{+}21, FCZ20, GMM22, GZC^{+}21, GXH^{+}21, GQ20, HAR20, Hem22, HXZH21, IAASK23, JPH23, JZR22, JB21, JA23, JCG^{+}22, KH22, KB21, KVV20, KGRG24, KK22c, KK22e, KJMB22, LJ22, LLSC22, LGZ^{+}22, LYF^{+}23, MFE^{+}23, MK22b, Man21, MG23, NET21a, NET21b, NGXZ21, PPA22, PK23a, PZZ^{+}23, PCK23, PJ21, PHDS22, RPMA22, RZCA21, RSMCP22, SSN22, SRRM23, SK22b, ST22c, ST23c, SS22e, SRBH22, SWZW20, TWW^{+}22, TAH22, TT23b, VH22, VDL23, VRV23, VLVS22, VB22a, VBM^{+}21, VPQ22, WJS21, WGLL23, WKL^{+}22, XCY22]. Internet [XHM22, YF20, ZCN22, ZZY^{+}22b, ZZK^{+}22, ZWT22, vdSTC21]. Internet-of-Things-Fog [JPH23]. interoperability [BS23a, NS23, Shi22a]. Interoperable [KVV20]. interpolation [Aka22b, AKA22c, Jeo20, Yil22c]. interpretability [ZYXX23]. Intersection [SWK22, JCL^{+}20]. intersections [AK22c]. interval [BFR24, CLLB20, CDC20, GC20, HR22b, LYBZ23, TSLL21, WBL22]. Intra [CLC^{+}23, KFKD22, SS23d, SSN23]. intra-class [KFKD22]. Intra-cluster [CLC^{+}23]. intra-frame [SSN23]. intra-inter [SS23d]. intraclass [CWLL20]. Intracluster [JP21]. Intracranial [KMS^{+}22, JSAA22]. intraoperative [Bai23]. introduced [MBM20a]. Introducing [AFKS23, JBBH21]. introduction [KW21]. Intrusion [GVSS22, KSTV21, LXC^{+}22b, MHB^{+}22b, SPK22b, AM20a, AMB23, AAARR20, AA23d, AGSN23, BK21, BJS23, CSR24, CK24b, DS20b, DRR22, GMM22, Gün23a, Gün23b, HAA23, KV22a, KBR24, KSSP22, KDA^{+}22, LTLX22, LCZ^{+}20b, LCM22, MSN22, MES23, ND23, PD20, PD22b, PSK23, RSR^{+}22, RZ24, SS24, ST23c, SR22c, SDR20, THX22, TBNPQ23, VJ22, VRV23, VKK^{+}24, WLLZ20, XJW^{+}22, YLZW23, YWC^{+}23]. intrusive [FPÁ^{+}20, MAB22, QZB^{+}23]. invariant [BEKS22, GS22]. invasive [BWS^{+}21, JMV24, RW24, TWW^{+}22]. inventory [CA22]. inverse [HYG20, LAKA21]. inversion [CHI^{+}23, TZZ^{+}23, TO22]. Investigating [GMS^{+}21, SBGC21, YFdP^{+}23, SG22b]. Investigation [Yil21b, HK24, NM24, SAQJ23, ZHWY22]. Investigations [SA22b, GDSS22]. investment [AE22a, LLL23, ZST^{+}23]. IO [GLRB21]. IoMT [KS24, YÇC22]. IoT [GQ20, VLVS22, KCP23, SPK22b, AAT21, AKA^{+}22a, AMRH21, AP23, AAD20, AAN^{+}21, AS22b, AA23b, AA24, BZEM20, BC23a, BFM^{+}21, BM22c, BTDJ21, CAS24, CLC^{+}20, CL22a, DCK21, GPR^{+}22, GR22b, GK23a, GVSS22, GNMELC21, GKAO20, HA22, HA21, JDLP22, JAC^{+}21, JYW^{+}20, KS20, KKP20, KK22c, KDA^{+}22, LA22, LCW21, LS24, LCM22, LF23, MMKA23, PKC^{+}20, RB22, RT22b, SS23d, SH22, SYL23, SR23, SCS^{+}21, SP23b, SR20b, SPK^{+}22a, Tur23, VRS22, VSP22, YR21, ZZC24]. IoT-assisted [SS23d]. IoT-based [Tur23, VSP22, AAN^{+}21, AA23b, BFM^{+}21]. IoT-enabled [KS20]. IOT-HML [KCP23]. IoT-TDMA [KK22c]. IOTA [AEMS24]. IP [CL22a, HMK24, PSP^{+}20]. IPDS [Ano21j]. IPFS [SS23e]. IR [CM20]. IR-UWB [CM20]. Iris [RAN22, CSL21]. Iris-based [RAN22]. irrational [KVP21]. ISIT2017 [CSW20]. ISIT2018 [WJS21]. island [JPL22, WZ21]. islands [WZ20b]. isolated [LYBZ23]. Isolation [ZWX21, MK23, TKS22, ZHWY22]. isolation-enhanced [MK23]. isomorphism [HCH^{+}21, HHC^{+}22]. isotopic [WZB21]. ISSA [WWG21]. issuance [TG23]. Issue [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Ano20x, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-41, Ano22a, Ano22b, Ano22-27, Ano22-28, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22z, Ano23b, Ano23c, Ano23-28, Ano23-29, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i]. Issue [Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano23z, Ano23-27, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Bad23, FZT22, JB20, JSLL20, KS21a, KLL^{+}21, LMMS24, RT20, SCH22b, SBG20, SKE22, Vin21, XZ20b, ABZS20, BI23, CSW20, FD20a, IS23, JC21a, JAC^{+}21, KSY23, LXZ20, Liu24b, LG21, PJ21, SCH22a, TC22, WY20, WJS21, Wri22, WYZAD20, Wu22, XZ20a]. Issues [PUL20, AS23a, DJF21, FLB23, KK22a, YIB22b]. Istanbul [TI22]. Italy [UYÖ^{+}22]. item [Cob22]. items [UPGCA22]. Itemset [KR22c, LK22a, NWW^{+}22]. itemsets [BVM22, FEH22]. iterative [INY^{+}23, KSA^{+}21, LLZ^{+}20, LZY^{+}20, PVP^{+}20, RV21, WCCC20, WPK^{+}22].

jackknife [OU22]. Jacobian [BSN22]. jamming [CAAHC23, K22, MP23]. jaundice [YY22]. Java [FKO22, NGB23]. Java/PCJ [NGB23]. Jaya [DSSS22, KV24, Sao22]. Jaya-based [Sao22]. Jaya-tunicate [DSSS22]. JCA [IAT^{+}23]. Job [XYY24, ANAMSAR21, BLB^{+}24, DS20a, JSA^{+}20, JR23a, KRB^{+}20, LZQ^{+}22, XZY^{+}22, ZSX21]. job-shop [LZQ^{+}22]. jobs [ANAMSAR21, DMPR24, RS20, TBT^{+}21, VS22a, YJJ23]. JOCP [WYLL24]. join [JWT^{+}20, RCS20, STISM21, WWF^{+}22]. join-free [JWT^{+}20]. Joint [DDZ^{+}20, HLL^{+}21, KS21b, MV24, NKST24, JY20, MKA23, PKB22]. jointly [WYLL24]. JONSWAP [WLX21]. JPEG [SKSP20]. JSC [XYY24]. Judicious [SSM^{+}21]. jump [NY22, ZST^{+}23]. jumping [AAMM23]. Jupyter [VZR^{+}21]. just [GMP^{+}20, HLG24]. just-in-time [HLG24].

K-D [HWY^{+}23]. K-DDoS-SDN [KKP24]. K-ear [YSS^{+}21]. K-L [LAK22]. k-nearest [KSY24]. Kaciranlar [ADA22]. Kafka [RZ24]. Kahan [DS23b, KSA^{+}21]. Kai [RKK23]. Kalman [DCM21, JPK22, KAQ23, LWZC21]. Kalman-Based [LWZC21]. KC [RN22]. Kernel [WAY^{+}21, BBA^{+}22, BFM^{+}23, Hem22, LZQF24, RK21a, SS21, VMFL23, ZHX^{+}23, ZZWZ21, ZSZ^{+}22]. kernelized [CT22]. kernelled [PSMM22]. kernels [AHW22]. Kessel [RM21]. Key [KJHM21, PI21, SV22d, AMB23, BV22b, CBK23, FCX^{+}22, FCMM20, FXX22, GJPS23, GWA^{+}24, HLS^{+}20, LS23a, LLC^{+}22a, MS21a, MG22, NAR^{+}22, NR22, PM23a, PDRM24, PA21, SPS23, SLG^{+}20, SYL23, TJC24, VB22a, VK23a, WGYZ22, ZGG^{+}23]. key-aggregate [PM23a]. key-value [FCMM20]. keying [Ano21-37]. Keypose [GBB22]. keys [AMAT22]. keyword [CKL20, HKA20]. Khosher [BK22c]. Kibria [MAAK23, OU22, SU23]. kick [HSL^{+}22]. kidney [AYJ^{+}22, RV23, VKSS23, VR22, XLL^{+}20]. kind [MLZ21b]. KL [LAKA21]. KMFA2 [SH22]. knapsack [Yil22b]. knee [BKK22]. kNN [TWL^{+}20]. Knowledge [CHL^{+}24, ZTL^{+}21, AS22a, AMR^{+}21, AHS20, AMD20, BBB^{+}20a, EGB21, JWL20, KT22b, LDW^{+}23, LLX^{+}24, SZ24, TLQ21, VCS^{+}21, WFHC21, WLLX21, WZC^{+}22a, YSK22, uHAU23]. knowledge-based [BBB^{+}20a, JWL20, WFHC21, WLLX21]. Kokkos [WSM^{+}24]. KP [MG20]. Krill [PBK23]. Krylov [CMS21]. Kubernetes [HMB22, LCL^{+}22]. KubeSC [HMB22]. KubeSC-RTP [HMB22]. KubFBS [LCL^{+}22]. Kullback [RM21]. kurtosis [PS22b].

L [LAK22]. Label [LLY^{+}23, SP22d, MSJ22, TJZ23, WER^{+}23, WCWG21, WLY^{+}22]. labeled [KUK22, LLL^{+}24]. labeling [RV21]. labels [LLT21]. labor [JCL^{+}20]. ladle [STJ^{+}20]. Lagrange [JGW20, LZL^{+}20a]. lake [CBR22, BSPM23]. lake-water [CBR22]. lakes [TA22a]. LAMMPS [HVT23]. lamp [WN21]. LAN [AA23a]. land [Aka22b, KRSR23, LM20b, SM21, WKY22, ZDH^{+}22]. land-use [Aka22b]. LANDSAT [CK21a]. landscape [WZ20b, WZ21, WN21, WNZ24, WZB21]. Landslide [LG23]. Lane [YCL^{+}22a]. Language [RP21b, RVJ^{+}22, AQP^{+}22, CMJC23, CDN^{+}21, KG22, NMS^{+}21, RD23, ST23a, WSL21]. language-based [NMS^{+}21]. languages [UWF^{+}21]. Lap [ZQX^{+}23]. Laplace [YYM^{+}24]. Laplacian [LCL^{+}20]. Large [LDW^{+}23, WZL^{+}24, ZTS^{+}22, AE22a, BBF21, BMK^{+}20, CK24a, FPXM21, HXST22, Kal22, KKM20, LM21, LHC^{+}23, LGL^{+}21, LSZ^{+}23, LTW^{+}24, LB21, MDSG^{+}23, NKKM21, NAR^{+}22, NWT21, RC22, RBWB21, RZ23, SK22a, Shi22a, Shu22, TAT^{+}23, TA23b, WZZ^{+}22, Wu22, XLL^{+}21, YCL^{+}22b, YBJ^{+}23, YTN^{+}20, YÖT^{+}22, ZABT^{+}20]. Large-scale [LDW^{+}23, WZL^{+}24, BMK^{+}20, LHC^{+}23, LGL^{+}21, MDSG^{+}23, NKKM21, NWT21, RC22, RBWB21, SK22a, Shu22, TAT^{+}23, WZZ^{+}22, Wu22, XLL^{+}21]. large-scaled [LB21]. laser [Yüc22]. last [LGL^{+}22]. late [CMLL21, LJB22]. Latency [SRG^{+}21, CTA^{+}23, NFF21, NN23, SSSR20, SWS^{+}24, ZQX^{+}23]. Latency-aware [SRG^{+}21, ZQX^{+}23]. latent [RR22, RF23, VSS23, ZLZ^{+}22b, ZLYS21]. Latest [MLKD20]. Lattice [AML^{+}22, BXH^{+}23, DK21, DZYY22, KK24, KM23b, KDFK23, LLS22, SG24, TJC24, ZGWZ23]. lattice-based [BXH^{+}23, DK21, KM23b, SG24, TJC24]. Layer [AJNS22, CAAHC23, CSWC20, EBLM22, FZA22, LYF^{+}23, MIN^{+}23, Mir22, MRK^{+}23, Ngu21, RRB22, RS21, SL20, SN22a, SS23g, WSM^{+}20, WYLL24, XLL^{+}21, YYPR22, YWC^{+}23]. Layered [DPSJ22]. Layering [MP23]. layout [IG24]. LBBCLCT [MV22]. LBBESA [LCZ^{+}20a]. LBMA [LLR^{+}21]. LBP [TUD21]. LBR [DPSJ22]. LBR-GWO [DPSJ22]. LBSs [LNC^{+}20]. Lc [SWS^{+}24]. Lc-Stream [SWS^{+}24]. LCLMF [LmJdL^{+}22]. LD [AI21]. LDP [WZMJ23b]. LDP-Fed [WZMJ23b]. LEACH [CGW^{+}20]. Leader [GAMT22, Ano21-39, DQF^{+}23]. leadership [SYRS^{+}22]. leaf [CÇ22a, DS23a, KSS21, Kay22b, KAAR23, OSK23, PSP22, PK22a, RKR22, SRRM23, YRV^{+}23]. lean [CJY^{+}20]. leaping [Ben22]. learned [KMS^{+}22, SM24a, VBP^{+}24]. Learning [AJNS22, GKAO20, KY23, LLL^{+}24, TBT^{+}21, XPLX23, YZPL21, ZQW^{+}21, dOdMC^{+}20, AYH^{+}22, AAMAA22, Abb23, AET^{+}22, AS22b, AMR^{+}21, AAE23, AA23b, APM^{+}21, AAA22b, AA23d, Anb22, ASJ24, AJAA21, AS23c, AR22b, AHT^{+}20, Asl22, BBA^{+}22, BMcKGK22, BLI20, BWW^{+}20, Bai23, BP23a, BDK22, BV22a, BGG24, BLB^{+}24, BPT^{+}23, BBD24, BF22, BC23b, BA24, BFM^{+}21, BJK^{+}24, BYTG22, BS23c, BRL^{+}20, CWLL20, CF21, CÇY22, CBR22, CDV23, CAAHC23, CSWC20, CLT^{+}21, CWW^{+}23b, CB22, CPH20, CD22, ÇG21, Cob22, DR21, DL23, DS23a, DSYF22, DRM22b, DSJ22, DK22a, DXXL20, DLG24, Ele22, ENB^{+}20, EASN22, ESS23, ETKD23, EUYY22, EA22, FSWW21, FKY^{+}23, FAM22, GKM22, GSS^{+}23a, GCS23, GZC^{+}22, GR22b, GB20, GDS23, GCP22, GKZ23, Gün23a, GM22c, HK22a, HXZH21, HHXH20, HLH^{+}20, HFH^{+}21, IDA22, JB22a, JDLP22]. learning [JKB22, JS22b, JNM^{+}22, JLZ^{+}24, JLE22, KSS21, KE22, KKR23, Kan22, Kar22, KS22a, KNM21, VK24, KBBH21, KBK^{+}22, KIW^{+}22, Kay22b, KSK^{+}22, KIAA^{+}22, KAQ23, KKC22, Kim21a, KHHK21, KY22, KMRR20, KCP23, KK22d, KG22, KR22b, KK22f, KNK22, KSKS22, KS23b, KSJ23b, KK23b, KS23c, LHJ22, LK22b, LMM23, LLJ^{+}20, LWLZ20, LLZ^{+}20, Li21, LPW^{+}21, LTLX22, LLY^{+}23, LSZ24, Lia24, LLM^{+}24, LJB22, LLKS21, LS23b, LLLX20, Liu21, LLMZ21, LZW^{+}21, LCL^{+}22, LDH^{+}22, LZM^{+}23, LZLZ23, Liu24b, LZQ^{+}22, LLA^{+}22, LL24, LLYZ23, LHL^{+}20, MRGP22, MTSU22, MBM^{+}20b, MBC23, MPB^{+}22, MTK^{+}21, MKBB22, MRS^{+}21, MMKA23, MRK^{+}23, MSBR23, NBK22, NRMB23, NSPdO21, NSBT21, NWT21, NTB23, NZD^{+}24, NBHN22, OCR^{+}23, OA22, ODK^{+}23, OAS23, ÖGS22, ÖTT23, PSP22, PAN22, PLP22, PR22, PK23b, PRS23, PLX^{+}24, PMR^{+}21, PGD^{+}22, PKB^{+}23, PKKL21, PCC^{+}21]. learning [PK22b, PK22c, PB22b, PM22, PS23b, PVVS22, PK22d, RR23a, RR23b, RSM21, RSM23, RTBC23, RAaB21, RV23, RBWB21, SB23a, SJA^{+}22, SKPD23, SPQM20, SVS22, SNBAI24, SSS^{+}20, SRG^{+}22, SRL23, SM22b, SV22c, SKCS23, SR24, SX21, SZW^{+}22, SPWX23, SZGR21, SM23b, SM23c, SET^{+}22, SP22b, Shu22, SBSK22, SPS22, SPKK22, SCM22, SMAG22, SSP24, SRIB23, SAL^{+}22b, SG22e, SV22e, SM22e, SKR^{+}24, TJZ23, THX22, TMB24, TCW^{+}23, TSCM22, TS22, TUD21, TSB23, UDS21, UWF^{+}21, UE22, UAS^{+}22, URK^{+}22, VLVS22, VS22b, VSP22, Wan22a, WZZ^{+}20, WZY^{+}22, WZC^{+}22a, WZL^{+}22, WZMJ23b, WWM^{+}23, WRE^{+}23, WKL^{+}22, WN23, WHDS22, WTL23, WT23, XLX^{+}21, XZG^{+}23, XLZZ22, XDL22, YS22, YKL23, YWQ^{+}21, YZX^{+}22, YLKC24, YBJ^{+}23, YLW^{+}22, Yil22b, YHL^{+}23, YMKH22, YFdP^{+}23, YSL^{+}21b, ZNDA22, ZDH^{+}22, ZZL^{+}20, ZBC^{+}21, ZZW^{+}22, ZZY22a, ZCZW22, ZCZC23, ZCZ^{+}22, ZTP^{+}23]. learning [ZYY^{+}24, dOPBdO21]. learning-based [AET^{+}22, Bai23, BBD24, CDV23, CAAHC23, CWW^{+}23b, CD22, GB20, JLE22, VK24, KBK^{+}22, KY22, LLKS21, LL24, MRS^{+}21, MSBR23, ÖGS22, PCC^{+}21, SRL23, YKL23, dOPBdO21]. learning-driven [WHDS22, ZDH^{+}22]. learning-Internet [AYH^{+}22]. Least [LWZC21, DCM21, XCY22]. ledger [BOI23]. legality [NET21b]. length [DKP^{+}20, KA22b, OHRS21, AST22]. lesion [HZZC22, KIAA^{+}22, MMR^{+}22, SSMT22a]. Lessons [VBP^{+}24]. letter [KE21]. leukemia [BP23a, SCM22]. level [AÜ23, BZK^{+}21, CD24a, CBR22, CPLX21, CZCM23, GDA22, HLCH20, JNMG21, KEK^{+}20, KBBH21, KK23a, LWZ^{+}20, LLX^{+}24, MG20, MCNR20, ÖUG22, RKR22, RK21b, SS23c, SN22b, SB21, WJLC21, ZZW^{+}22, ZK24, ZCD^{+}22]. leveling [KAO22, YGS^{+}23]. levels [Kim21a, WHDS22]. Leverage [ZLYS21]. Leveraging [KKAM21, SKSN22, FDY21, RCBTGE23]. Levy [DBD23, SSK23b]. Lexical [EGB21, SSK^{+}23a]. lexicographic [Li20]. lexicon [DSS^{+}23]. Libfabric [SSSR20]. libraries [BFR24]. Library [BSML21, FKO22, IGYS23, MM21, PB22a, USP^{+}23]. LibreSocial [GM21]. license [QCC^{+}23]. Lidar [ML20]. life [CH21, GA22, QPS20, uZKH^{+}20]. lifecycle [SAL^{+}22b]. lifeline [FKO22]. lifeline-based [FKO22]. lifetime [MS21a, SK22a, SM24c, ZWX21, ZTY^{+}22]. light [GLN23, VH22, VPQ22]. light-weighted [VH22]. lighting [WNZ24]. Lightning [TCRP23]. Lightweight [BKGC23, DAT23, JPH23, LSW^{+}20, SWCB20, SL20, TWL^{+}20, TCW^{+}23, YR21, DBB24, Kha22a, LGT^{+}23, DPD^{+}22, NR22, SC22a, STJ^{+}20, TJC24, TZZ^{+}23, TMB24, TSG21, USP^{+}23, VB22a]. like [CMS24, Kar22]. limb [BZGM22, YHL^{+}21]. limit [JHG23]. limitation [HSR23, QCC^{+}23]. limited [KW21, SLLMTS^{+}24]. line [YLLL20, ZGH^{+}22]. Linear [ZGH^{+}22, AV22a, AOACAQ21, Ano21-36, AU22, BC22, CMY21, DS22b, GCF^{+}20, HNOI24, IMNC22, KML23, LGDW22, LFWJ22, LLO21, LAK22, MAAK23, Mon21, OKJ^{+}21, OAA22, ÖK22, SAHAN22, SDSW21, The21]. lines [FPHZ19, Pan20]. lingual [GSS^{+}23a]. linguistic [HV21, HH23, SRG^{+}22]. Link [WXCY20, KBS^{+}22, LZXG22, MSL22, MAAA22, PMP23, SEMA^{+}22, SSP24, YYPR22]. link-disjoint [LZXG22]. linkable [DRMA22]. linkage [AN23, SAPC21]. linked [BJWY20, DPS21]. linking [AMD20]. links [DKP^{+}20, SS22d]. linpack [CMS24]. linpack-like [CMS24]. lion [AFBM^{+}23, PKK23, PKR22, SAC22, SYG22, SSS23b, VK23b]. list [AA21, BJWY20, TP20]. listeners [SS21]. Listwise [LLMZ21]. Literature [JKB22, Haj20, KM23c, KJ20, LXC^{+}22b, LCM22, MBT^{+}20]. Liu [FPHZ20, AA22a, AAK21c, BC22, HML19, KÖÖG22]. Liu-type [AA22a]. live [GJK^{+}20, KSK^{+}22, NK22, SM21]. liver [DR22c, GKZ23, VRR^{+}22, VY23, YMKH22]. LLF [Ker22]. LLVM [GDGN23, WKB^{+}22]. LLVM-directed [GDGN23]. LMF [EGB21]. LMI [TSLL21]. Load [AB22b, GP22a, KK22a, MS21c, SHBC20, VB24, AB20a, AET^{+}22, Agr21, AAK^{+}21b, AAG^{+}22, AI21, AFNH21, BM22a, CDV23, CJ21a, CCZ^{+}21, CWFK23, DDUK23, FKO22, HAS^{+}22, HYK21, HMK24, HLL^{+}21, LLAV22, LWLZ20, LCZ^{+}20a, LST22, MV22, MSSM24, MG21a, MV24, NA22c, RSM22, SMR23, SKSN22, SSV24, YWQ^{+}21, ZBZ^{+}20, ZCH^{+}23, MV22]. load-balancing [LCZ^{+}20a]. Load-based [MS21c]. loading [Kal22]. Loads [WCZ^{+}23, LLJ^{+}22b]. loan [CA23, LZL^{+}22]. Local [JQGL20, KS22c, Pan24, AI22, Ano21f, BCM22, EASR22, GAMT22, KK21a, KAO22, LW20b, LXJ^{+}22, MSL22, PPM22, RA21, RMS22, SKHL22, SCH^{+}24, SBS22, WJL^{+}20, WAY^{+}21, WZMJ23b, YLZT23, YZZA23, ZWY^{+}24]. locality [AS22a, BYPO21, CYW^{+}22, HDS^{+}23, KW21, RYG^{+}21, SCKU23, WQY^{+}22, ZHX^{+}21]. locality-sensitive [CYW^{+}22, ZHX^{+}21]. localization [AK22b, AC23, FMJ^{+}22, GA23, LXJ^{+}22, MBM20a, RCYuRH21, SA22a, SSW^{+}22, SM24b, TK24]. localize [RP22]. Locating [KKM20, BEKS22]. Location [EVVR21, ZSZ^{+}22, ABSS22, CY24, FMNF22, GLA^{+}22, GZY^{+}22, HDXH20, LLM^{+}22, LZL^{+}23, MS21b, NT23, PR23, PS21, RSKA23, RKL21, SJ22, SSSP21, WGZ^{+}20, WGLL23, XCY22, YYM^{+}24, ZHXY23]. location-allocation-routing [FMNF22]. location-aware [NT23]. Location-based [EVVR21, LZL^{+}23, MS21b, PR23, SJ22, WGLL23]. location-dependent [RSKA23]. locator [Kab23]. lock [BS23a, YY20b]. locks [MTD^{+}20, NdMP22]. log [TA23b]. logarithmic [BK22b]. logging [ARHT21]. logic [APM22, CLWX21, DWZ^{+}22, DBB24, PPA22, RNRK22, SYRS^{+}22, SCP20, XCJ22, YSH^{+}22]. logic-based [APM22]. logical [QNZP22]. Logistic [SV22d, AV22b, KV22a, LHC^{+}23, PP20b, SHA^{+}22]. logistics [LGL^{+}22, ZK24]. logs [LPC^{+}21]. LoHDP [SCH^{+}24]. London [CCD^{+}20b]. long [AKS^{+}22, ATC23, BKK23, CL23, CÇ22c, DR22c, DM22d, HR22a, IIK^{+}23, JR23b, KTM22, KH22, Kab22, Kar22, KDC22, PT22, PVRM22, SLLMTS^{+}24, SV22b, SMBM24, SVD^{+}22b, SHZY20, SM23b, VMFL23, YRV^{+}23, YW22, ZLCL21, ZLCS21]. long-haul [SLLMTS^{+}24]. long-short [Kab22]. long-term [AKS^{+}22]. longevity [STJ^{+}20]. lookup [ZZS21a]. Loop [MV22, AV21, BL22, FGJ^{+}21, MCNR20, WKB^{+}22]. loop-level [MCNR20]. loops [DSS21]. loosely [CLLB20]. looseness [SXC^{+}23]. LoRaWAN [LSZ^{+}23]. Lorenz [LYC22]. loss [BYTG22, CWLL20, DM22d, LM22a, VB21]. lossesless [KSVP22]. lottery [LLR^{+}21]. Loudspeaker [LHWT20]. Louvain [BVP22]. Low [RAN21, AM20b, Ano21-37, AKLK24, BM22b, CTA^{+}23, FSFM22, JK22a, KKKS23, KK21b, KBBH21, LHWT20, MS21c, MCL^{+}20, NFF21, QLL^{+}22, SSSR20, SBA22, SP22c, SR20b, TK22a, TXT24, XLZ20, YNK^{+}23, ZXLD21, ZWZ^{+}22]. low-contrast [SBA22]. low-cost [SP22c, SR20b]. low-diameter [YNK^{+}23]. low-energy [MS21c]. low-latency [NFF21]. low-power [AM20b, MCL^{+}20]. low-rank [ZXLD21]. low-storage [TXT24]. lower [YHL^{+}21]. lowest [ANAMSAR21]. LPLB [SC22a]. LPWAN [DGP20, KKP20]. LPWAN-based [KKP20]. LR [DTK24]. LR-TCN [DTK24]. LRP [LGX^{+}23]. LRP-based [LGX^{+}23]. LS [FM22]. LS-SVM [FM22]. LSB [SAGS24]. LSB-XOR [SAGS24]. LSR [WQY^{+}22]. LSR-forest [WQY^{+}22]. LSSVM [YLJH22]. LSTM [Asl22, DR22b, FDD^{+}24, HK24, JR22, NWZ^{+}21, RT21, SMBM24, SD23, VM23, VKK^{+}24]. LSTM-based [SD23, VKK^{+}24]. LSTM-CNN [HK24]. LTIS [DT22b]. Lukman [MAAK23, OU22, SU23]. Lung [RR23b, MSBR23, RPMA22, SZqWZ20, SMKA22, TFZC23, ZZQ^{+}22]. Lustre [FLP20]. LWMPI [USP^{+}23]. lymphedema [NBHN22]. lymphoblastic [BP23a]. lymphocytic [SCM22].

M [PSP^{+}20]. M-REP [PSP^{+}20]. MABC [RCK22]. MABC-EPF [RCK22]. MAC [DK21, GXH^{+}21, Ngu21, RK23c]. MAC-MELBC [DK21]. Machine [AET^{+}22, AJNS22, AA23d, BRL^{+}20, CAAHC23, KAQ23, PK23b, SVS22, SSS^{+}20, SMAG22, SM22e, VSP22, WHDS22, dOdMC^{+}20, AE22b, AMR^{+}21, AAE23, AAA22b, AJAA21, AR22b, AHT^{+}20, AYJ^{+}22, BBA^{+}22, BLI20, BGG24, BLB^{+}24, BPT^{+}23, BBD24, BF22, BC23b, BA24, BFM^{+}21, BRS^{+}22, BS22, CÇY22, CBR22, CK21a, CSWC20, CPH20, CD22, Cob22, CZL^{+}22, DBD22, DRM22b, DP22a, DS20b, DCK21, DS22c, ENB^{+}20, EASN22, EUYY22, FGZC23, FKY^{+}23, GKM22, GKZ23, Gün23a, HK22a, Hem22, HAK22, HTVL22, HPS23, IDA22, JS22b, JMY21, JKS20, KSS21, KNM21, KB21, VK24, KR22a, KBK^{+}22, KS20, KIAA^{+}22, KKC22, KHA22b, KCP23, KNK22, LJBS23, LJB22, LPZ^{+}22, LWCZ22, LLA^{+}22, LZQF24, MR23b, MTSU22, MBC23, MNDK22, MPB^{+}22, MES23, MTK^{+}21, MKBB22, MSBR23, NK22, NRMB23, NTB23, OCR^{+}23, OAS23, PSP22, Pan24]. machine [PRS23, PKKL21, PCC^{+}21, PK22b, PK22c, PK22d, PS23c, RSR^{+}22, RR23b, RRJ23, RG22, RV23, SS21, SKPD23, SPQM20, SNBAI24, SEM^{+}20, SM22b, SV22c, SS22c, SM23a, SZW^{+}22, SM23b, SM23c, SBSK22, SPS22, SPKK22, SR22c, SAL^{+}22b, SP21b, TCW^{+}23, TNP21, TS22, TUD21, TSB23, UWF^{+}21, UE22, UMR23, VPSM22, VLVS22, WY20, WAY^{+}21, WJLC21, WKL^{+}22, WLLZ20, WTL23, XCY22, YZX^{+}22, YSK22, YGS^{+}23, YMKH22, YFdP^{+}23, YY22, ZNDA22, ZJSJ20, ZJSJ21, dOPBdO21, IVP^{+}23]. machines [AZM20, AFK^{+}22, BPM^{+}22, BBB^{+}20c, CPQ^{+}22, CON23, CWFK23, DAdBM^{+}24, EK20, GJK^{+}20, JBBH21, MNDK22, MK23, SLC20, WCZ^{+}23, XA22, YLJ22]. macro [CS21, LYL21]. Maghrebian [Ano21-40]. magnetic [IK22, Kot23, ODK^{+}23, SKÇA23, TNP21]. Maharashtra [SM23c]. main [STISM21, XLL^{+}21]. maintaining [SK22a]. maintenance [HL20]. major [TI22]. majority [AM20b, SN22a]. make [MKRK23]. makeup [CXXW24]. making [DRR22, KGM23, LLX^{+}21, MTY21, Ogi20b, ZARR23]. MaLeFICE [SMAG22]. Malicious [DD21, ZWL^{+}23, DPYS22, HZQ^{+}24, JKS20, KSTV21, LZW^{+}21, SVM24, TSB23, YZX^{+}22, ZTP^{+}23]. Malware [CZL^{+}22, LLL^{+}22, AKUA22, AT22, ATC23, BD21, DDH^{+}20, KRJS22, LZC^{+}20, MVR23, RF23, SVD^{+}22b, SR22b, SR24, SBSK22, SD23, ST22d, XCJ22, YLS^{+}23, ZZC24]. mammogram [AMV22, BBA^{+}22, MH23, PBK23, SSMT22b, TB23]. mammograms [SKS22]. mammography [SSMT22a]. manage [BBB^{+}20a]. management [AET^{+}22, ANP^{+}20, AMB23, AlS21, AS23c, BC21, BZK^{+}21, CPLX21, CA22, CÇ22c, DCT^{+}23, FKK23, GZ20, GLN23, GLM^{+}22, GWA^{+}24, GYL^{+}21, HTVL22, IAA20, IPRS21, JC21b, KMZ^{+}20, KLK23, KTK20, LS23a, LKR^{+}22, LGL^{+}21, MR23b, MM24, MG22, MV24, MSS^{+}20, OT20, OO21, OMA^{+}23, PB22a, PA21, RVAE21, RKC22, SAM^{+}23, SG22a, SPS23, SSM^{+}21, SR23, SM22d, XLXZ20, YY20a, ZARR23, ZSC^{+}21, ZFW23]. manager [RT22b, SV22c, AAE23]. managing [GFPGT21]. mandibular [VB21]. MANET [HUJ22, JKP22, NM20, SR22a, SL20]. MANETs [DD21, SK23]. manifold [JS23, TJZ23, XDL22]. mAnipulation [MdAA^{+}21]. Manipulator [LXT^{+}22, ZTMC22, ZTMC23]. Mannar [GDA22]. MANSOR [GCQ^{+}24]. manta [LK22a, YS22]. mantissa [GCF^{+}20]. Manual [SGS21a]. manufacturing [DXXL20, HPCK22, LW20a, MNDK22, SAPC21]. Many [YZ21, Zhu22b, AV21, CHI^{+}23, CZ21, FGZC23, FKO22, GSZ^{+}20, HK24, Hua20, IAA20, LCKJ21, LB21, XGCZ23, ZX20, ZWL^{+}23, ZWCC23, ZZC24, ZWY^{+}24, Zhu21]. many-core [AV21, FKO22, LCKJ21, LB21]. Many-objective [YZ21, Zhu22b, CZ21, FGZC23, IAA20, XGCZ23, ZX20, ZWL^{+}23, ZWCC23, ZZC24, ZWY^{+}24, Zhu21]. many-task [GSZ^{+}20]. many-threaded [CHI^{+}23]. many-to-many [HK24]. many-to-one [HK24]. manycore [CdOO^{+}20, HOS^{+}21, USP^{+}23]. Manycores [SCH22b, SCH22a]. Map [AV22a, Dev21, RT21, SV22d, VS22a, AAMM23, ABB22, HV21, HH23, JS23, NT23, YCL^{+}22a, YMZD21]. Map-based [RT21, NT23]. Map-Reduce [VS22a]. Mapping [KMB23, MCD^{+}23, AÜ23, DSS^{+}23, FHH^{+}20, GBBS21, HK21, HHYL22, LLW^{+}22a, LXJ^{+}22, RJH^{+}23, RP21b, SZI^{+}23, SCD^{+}23, WZHL21, WGDP23]. MapReduce [BJGF20, JSP20a, JSP20b, RKuH^{+}20]. MapReduce-based [BJGF20]. maps [BTDD20, HL23, KSS24b]. Marathi [SSK^{+}23a]. marble [TUD21]. Marine [AO23, BDK22, SSN22]. maritime [JH21]. mark [WLL^{+}21b]. market [KKR23, KBJ21, KRSR23, PJK23, WWG^{+}20, Xia20, XY21]. markets [KSJ23b, TI22, XCZ^{+}21]. marking [YCL^{+}22a]. Markov [AAK^{+}21a, AD22b, BB23, JAR^{+}24, JY20, KC22]. Mary [TBT^{+}21]. mashup [BMSD23, RR22]. Mask [SKP23, JD22]. masked [CSL20]. masks [CXX^{+}22, LDS^{+}23]. masses [SKS22]. massive [CYC21, PP20a]. Massively [MGB^{+}23, PJKO24, PBD23]. mat [KAR^{+}23]. mat-heuristics [KAR^{+}23]. Match [LCZY20]. Matching [LCZY20, BCC^{+}21, CJ21a, DGSB20, FGL^{+}20, GMA20, GRC^{+}23, HIEH22, JQGL20, KB22, Li21, LLW^{+}23, LWL^{+}24, LZSC23, Pan24, RGM22, ST22c]. matchmaking [SKRS21]. material [PHZ^{+}22a]. materials [Zhu21]. Mathematical [GMN21, PS22a]. mathematics [KLJ21]. mating [DBK21, GVSS22, SAMS23]. MATLAB [RBC20, TSA21]. matrices [AAEA20, HMQO23, WPL20]. Matrix [TI22, AML^{+}22, AAG^{+}22, AAQOT23, BFM^{+}23, CHI^{+}23, CWH^{+}24, DP22b, Eke22, GHRM21, HCG21, HXY^{+}22b, HLS^{+}20, JW22, KGM23, KS21b, MW21, OBER22, WQS^{+}23, WZG^{+}21, XWD^{+}22, ZTM21]. matrix-based [DP22b]. matrix-vector [AML^{+}22, AAG^{+}22, HCG21, KS21b]. matter [GAMT22]. matters [WHDS22]. max [PS22b, ZGH^{+}22]. maximal [KKM21, LM21]. Maximization [ZXL^{+}21b, GXL^{+}20, LGLZ20, SK22a, XCZ^{+}21, ZGWZ23, ZXL^{+}21a]. maximize [Zha22]. Maximizing [SSS23a, SKSB20, JXL^{+}23, MS21a]. maximum [BZWH21, DPFC20, DD24, INY^{+}23, WJL^{+}20]. Maxout [PKSM24, RPPK23, RPMA22, SRRM23, SSS23b]. Maxwell [SAB22]. mayfly [Anb22, DS22c, SD23]. mazes [HL23]. MBN [RC22]. MC [rSN21]. MC-CDMA [rSN21]. MCFCM [NA22b]. Mean [LWZC21, RMR^{+}22, BEÖ23, HUC^{+}22, LXYY21, MSA22, PS22b, RKK23, SK22c, VS22c, VSK22, ZBY22]. means [CT21, ESS23, FM22, GK23a, HVB22, KSS21, KUK22, KV22a, LS23b, NJ21, PS22b, RP21c, VR22]. measure [Ano21g, Par22, PD23, YYPR22]. measurement [BGPQ21, CW22, FPHZ20, GHL^{+}23, GN21b, HML19, LJZ21, LGL^{+}21, SK22c, SZL^{+}22, VS22c, YHOY22, ZCN22]. measurements [KAO22]. measures [GLA^{+}22, MY24, SS23a, ST22d, TK22a]. Measuring [CWDM^{+}21]. MEC [LLX^{+}21]. MEC-based [LLX^{+}21]. MECC [NM24]. mechanism [Agr21, BKS22, CL23, DRR22, DCZ^{+}22, DQF^{+}23, EABZB21, FPHZ19, FKO22, GYL^{+}21, HR22a, JH21, KV22b, LHJ22, LmJdL^{+}22, LSQW21, LLC^{+}21b, NT23, Pan20, PSP^{+}20, RSM23, SG22a, SPS23, SLJ23, SMR23, SSP22, SZJ21, TWG^{+}21, TWXL21, VP22b, WMQ^{+}22, WN21, XHZ^{+}21, Xia20, XZW^{+}20, XCZ^{+}21, YYM^{+}24, YMZD21, YLJ22, YZZA23, ZLCS21, ZRY^{+}22, ZKZ^{+}23, ZCX^{+}21, ZHXY23]. mechanisms [FK20, KHA22b, MJAM23, PCG^{+}21, SG22b, SSS23a]. media [APM^{+}21, ARS22b, AS23c, BKM^{+}21, DBD22, GlSL^{+}23, GK22a, HAR24, KM21a, KUK^{+}24, KM21b, LK22a, LWZ^{+}20, MS22b, RSJ21, RSJ22, TWG^{+}21, WGZ^{+}20, YLKC24, YRSO23, ZR24]. Median [KR22c]. mediator [Ano21j]. mediator-based [Ano21j]. Medical [AS22d, Che24, JB20, LCL^{+}20, YW22, ZWT22, Abb23, ABB22, AAEA20, AJK^{+}23, Bei24, BYL20, ÇKÇ22, CGS^{+}21, CBK23, DKK24, GCSZ24, IAASK23, KS23a, LJ22, LGDW22, LW20b, LCW21, LZLZ23, MSBR23, NS23, RB22, RP21c, RV23, SK22a, SSK23b, SJ23a, SK22d, SAL22a, SKSP23, TWW^{+}22, TSG21, WLCW20, WLJ20, WWA22, WWJ^{+}20, ZZW^{+}22]. medicine [DWDG20, ZBC^{+}21]. medium [DCT^{+}23]. medium-scale [DCT^{+}23]. meets [LDH^{+}22]. megastores [ASJ24]. melanoma [GM22b]. MELBC [DK21]. mellitus [SM22e]. MEM [PQKDT21]. member [HYC^{+}24]. membership [PB21]. Memetic [VS22b, SR20a]. memoization [AJP22]. memories [IIK^{+}23, KPA23, SCdLV20, ZTY^{+}22]. Memory [AML^{+}22, GMSM21, AHW22, AKZA22, ATC23, BKK23, CdOO^{+}20, CL23, CMJC23, CKL20, CYK^{+}21, CWFK23, CÇ22c, Cud20, DR22c, DM22d, FQD^{+}23, GMS^{+}21, GADM20, HR22a, HLT23, IGYS23, JR23b, KTM22, KH22, Kab22, Kar22, KDC22, KHPH20, KMB23, LM21, LYF22, PT22, PCC^{+}21, PVRM22, RPM22, STISM21, SV22b, SMBM24, SVD^{+}22b, SHZY20, SM23b, SCK^{+}23, SM24b, USP^{+}23, WCCC20, WK20, YRV^{+}23, YLKC24, ZSC^{+}21, ZLCS21, ZLW^{+}23]. memory-bound [AHW22]. memory-intensive [CYK^{+}21]. memristor [GGA24, SM24b]. mental [GK22a]. menu [TC24]. merged [Wan20]. Merkle [SSS23a]. mesh [KDS^{+}20, RJH^{+}23, SWCB20, SPS22, ZBZ^{+}20]. mesh-based [ZBZ^{+}20]. Mesos [HXY20]. mess [SGS21b]. message [AC23, Cud20, DK21, DSS21, FLB23, FGL^{+}20, GMA20, GRC^{+}23, LYL20, NVV24, SRS^{+}21, TO22]. message-passing [DSS21]. messages [BZWH21, Gho21]. messaging [ELIG23]. Meta [CK24a, GRG^{+}22, APP^{+}21, CEN22, GMN21, HAK22, KA21b, MG21b, NN23, RSM22, SV21, SVD22a, XPLX23]. Meta-heuristic [CK24a, GRG^{+}22, APP^{+}21, CEN22, GMN21, HAK22, KA21b, MG21b, RSM22, SVD22a]. meta-heuristic-oriented [NN23]. meta-transfer [XPLX23]. metadata [AJ21, FLP20]. Metaheuristic [AB22a, SM23d, ÇGB23, GSS^{+}23a, KAO22, OE22a, SD23, XA22]. metaheuristic-based [XA22]. metaheuristics [LM20b, PPM22, SEM^{+}20]. metamorphic [NKY23]. metanil [MCT22a]. meteorological [DM22b, TO22]. meter [SR20b]. metering [FPÁ^{+}20]. method [AAMAA22, AMJK21, Akb22, Ano21-37, Ano21-36, AHV21, ATG23, BWW^{+}20, BST^{+}22, BKK22, BGG24, BTT21, CLLB20, CNG^{+}20, ÇKÇ22, CEN22, CSC22, CDC20, CYZX23, CWW23a, CYZ^{+}21, CWH^{+}24, DT22b, DBB24, DGB24, DCWM20, DLY^{+}23, EASR22, FEH22, GZ20, GCS23, GZC^{+}22, GSTS22, GWA^{+}24, GCQ^{+}24, Gul22, GZY^{+}22, GGJ^{+}24, GU23, HNSS22, HLT23, HKP21, JT23, JB22b, JS23, JYH^{+}24, JBCI20, KT22a, KA21a, KKC22, KHHK21, KLH24, KD22, KSSP22, KPA23, KSY24, KG23, LZL^{+}20a, LW20b, LFG^{+}22, LM22a, LWZ22a, LCW^{+}23a, LCL^{+}20, LTL^{+}20, LXC^{+}22a, LLL^{+}22, LLS22, LWHW22, LLL23, LCZ^{+}20b, LZQF24, MLZ^{+}20, MLVGR24, NJ21, NY22, ÖTT23, PMC^{+}21, Par22, PLZL24, QKSK23, RVAE21, RZCA21, RÖ22, RKK23, SS23a, SEM^{+}20, SVM24, SZZ^{+}24a, SX21, SZZ^{+}21, SGH23, SCM22, SD22, TFZC23, TWQ^{+}21, THX22, TNP21, WLCW20, WZ20a, WWZ^{+}22, WWA22, WGLL23, WLL^{+}21b, WZG^{+}21, XLX^{+}21]. method [XZG^{+}23, XCJ22, XJW^{+}22, XLZZ22, XA22, XDL22, XPLX23, YWQ^{+}21, YS23, YOWY22, YGS^{+}23, YLJH22, YTN^{+}20, ZWW^{+}21, ZQL^{+}21, ZZQ^{+}22, ZX23b, Zha22, ZNLL22, gZWfL^{+}20]. method-based [XDL22]. methodological [SSI22]. Methodologies [KLL^{+}21]. methodology [AS20, BCC^{+}21, CA22, DJJR22, GM22b, PCC^{+}21, SD23, WZC^{+}22b]. Methods [SAB22, AE22b, Aka22b, AAA22b, Bei24, CdOO^{+}20, CMS21, CÇY22, DLC^{+}21, GDS23, GA23, GCSZ24, HPCK22, IMNC22, KW21, KDFK23, MM22, MH23, NK22, PGD^{+}22, PHDS22, RD23, SYRP22, SS22f, TÖK21, THT20, WHDS22, WTL23, ZLD22]. metric [CU22, GVSS22, LC20]. metrics [Mar23a, TS22]. metro [SZZ^{+}24b, ZYZC23, ZXL^{+}21a]. metro-city [ZXL^{+}21a]. Mexican [Alm22]. Mexican-Patient [Alm22]. MF [SSV24, xZlGCzJ20]. MF-DCCA [xZlGCzJ20]. MF-DLB [SSV24]. MFCC [Yil22a]. MFDC [LZF20]. MFDC-net [LZF20]. MH [CK24a]. MH-ARO [CK24a]. MiCADO [DKA^{+}21]. Micro [LYL21, CS21, SR20b, SRG^{+}21, WSM^{+}20]. micro-batching [SRG^{+}21]. micro-bleeding [WSM^{+}20]. Micro-macro [LYL21, CS21]. microarray [BD22, PS22b]. Microblog [MAR23b, XGX^{+}21]. microblogs [VV23]. microcomputer [Liq22]. microfluidic [RP21a]. microgrid [ASA^{+}21]. microgrids [AJAA21]. microscopy [BMK^{+}20]. microservice [DKC^{+}21, SKRS21]. microservice-based [SKRS21]. Microservices [AA24, AHS20, JGJ^{+}21, JC21b]. Microservices-based [AA24]. microstrip [GAM23]. Middle [HNG22]. middleware [CCL^{+}22, CWL^{+}20, OMA^{+}23]. migrate [KB21]. migration [BN21, FGZC23, GJK^{+}20, HAK22, HPS23, LHJ22, LLR^{+}21, LLX^{+}21, LXC^{+}22a, NK22, QLL^{+}22, SEM^{+}20, VB24, XA22]. mile [LGL^{+}22]. mileage [Ano21-43]. millionaire [AYD21]. MIMO [Ano21-38, BC23b, KR23]. MIMO-OFDM [KR23]. Min [PS22b, ZGH^{+}22]. Min-max [PS22b, ZGH^{+}22]. mine [DWZ^{+}20b, DWY^{+}21, WMQ^{+}22]. minHash [AS22a]. mini [LBG^{+}20, OWB^{+}20, TTZX22, YBJ^{+}23]. mini-application [OWB^{+}20]. mini-applications [LBG^{+}20]. mini-batch [YBJ^{+}23]. mini-Xception [TTZX22]. Minimal [BP24, TWW^{+}22, WRJ20, YY22]. minimization [WFY22b, WCSG20]. Minimize [IQS^{+}22]. Minimized [SBSK22]. minimizing [ASC22, RSM21, RSM23]. Minimum [ST22a, ASMS21, CCZ^{+}21, KÖ22, MR23a]. Mining [AMR^{+}21, BVM22, JB21, RV21, AA22b, Ano21c, Ano21g, Aru22, ASMK21, CGS^{+}21, CMK22, CPH20, CPPP21, Dev21, ER22, EGB21, GU23, JR23b, KIW^{+}22, KCM^{+}22, LM21, LK22a, LJC23, LFG^{+}22, LXZ22, LWCM21, MS22b, NWW^{+}22, SRL23, SM21, THW21, TSR22, VSS23, WWG^{+}20, WMC21, Yl20b, ZHJ20, ZLW^{+}20, vdSTC21, KR22c]. Mining-based [JB21, EGB21, WWG^{+}20]. mining-constrained [ASMK21]. Minkowski [RT21]. minor [YA22b]. Minority [Fio20, SP23a]. mirror [LTW^{+}24]. MIS [XLXZ20]. MIS-based [XLXZ20]. misallocations [ESB20]. misinformation [RDB22]. missing [GKAO20, HTVL22, SK21a, SS22d]. mission [RK23c]. mitigate [MAB22, RGKK21]. Mitigating [FLB23, KHY^{+}20, SCD^{+}23, ZCZ^{+}22]. Mitigation [AP22, PCG^{+}21, AS23a, CCBA23, CSR24, CSC22, DBK21, HJT^{+}20, LWZC21, PA23]. mixed [JCL^{+}20, KKM21, LAK22, ÖK22, PK22b, RVAE21, WCWG21, YZZA23, ZWCC23, ZCL21]. mixed-method [RVAE21]. mixture [HGNN22, JD23]. ML [AA23d, CCBA23, Faz23, KV24]. ML-based [KV24]. ML-IDSDN [AA23d]. MLCT [CZCM23]. MLP [XLL^{+}21]. MLPNN [BSN22]. MLPNN-RF [BSN22]. MLR [DÖD22]. MLR-FFANN [DÖD22]. MMF [PLZL24]. MMWD [LZW^{+}21]. MobEdge [DPB22b]. Mobile [DPB22b, NY22, ALR22, APV23, Ano21c, ATG23, BWTJ20, BM22b, BVS20, BKS22, Bul22, CY22, CJ21a, CHMC21, CJC22, CG22, DCR23, FGZC23, GXL^{+}20, GAS23, GKÇ22, Gho21, HLG24, HAAF22, KBR24, KJ21, LGL^{+}22, LZW^{+}21, LLL^{+}21, MRAM^{+}21, NSKS21, PPM22, PGL^{+}23, PB22b, QLL^{+}22, RJ22, RGKK21, RAN22, ST22b, SPS23, SVD^{+}22b, ST23b, SG22d, TCRP23, TWG^{+}21, TWQ^{+}21, TTA20, TT22, VCBB20, WLCW20, WGZ^{+}20, WBZ21, XWC^{+}22, XXD^{+}22, XHM22, ZYZC23, Zhu22a, ZGH^{+}22]. MobileNetV2 [MCT22b, SKÇA23]. Mobility [SVB23, TT22, MM24, PKVS21, WMC21, XLXZ20, YMZD21]. modal [DSC^{+}21b, ZCZC23]. modality [DAK22]. mode [CYZ^{+}21, DT22a, DZLH20, KA22a, LLKS21, SG22e, SV22e, THW21, TAH22, YLJ22, ZTMC22, ZTMC23]. mode-recycling [CYZ^{+}21]. Model [AS24, MdAA^{+}21, ÖK22, RF21, AD22a, AHW22, AAK^{+}21a, ACJ21a, AAA^{+}22a, AFKS23, AA22a, AAARR20, AAK21c, Ano21e, AKLK24, AHS20, AU22, Aru22, AMM^{+}20, Asl22, AFM22, BC22, BP23a, BC21, BPAE20, BEJD22, BKGC23, BMG22, BNB22, CSL20, CSV22, Çet23, CSWC20, CK21b, CWL^{+}21, CWL^{+}22, CMJC23, CKL20, CSL21, CZL^{+}22, DCM21, Dev21, DT22b, DSJ22, DPS21, DAW22, DBD23, DK22a, FM22, GPDB20, GDA22, GSS^{+}23a, GWGR20, GBS24, Gho21, GSG^{+}23, GYZ^{+}20, GK23a, GWA^{+}23, GWA^{+}24, GNS22, GYL^{+}21, GMK^{+}21, GK22a, GK22b, HK24, HGHD22, HUJ22, HD23, HBB20, HR22b, HHXH20, HYT^{+}21, IMNC22, JSS22, JS22a, JPL22, JY20, JCL^{+}20, JQGL20, JYW^{+}20, JMY21, JJZ^{+}21, JYC^{+}21, JD23, KUK22, KH22, K22, KE21, KK24, KM23a, KBBH21, KK22b, Kay22a, KIW^{+}22, KCP^{+}22, KKE^{+}22, Kim21b, KMY24, KAO22, KHR22, KG22]. model [KDA^{+}22, KDS22, KSB23, KS23c, LJ22, LLAV22, LPHK20, LSSQ22, LWLZ20, LW20b, LSD21, LYL21, LWZ^{+}22b, LGZ^{+}22, LHX^{+}23, LML^{+}23, LCCT22, LJP^{+}21, LS23b, LZY^{+}20, LPC^{+}21, LSQW21, LLS22, LSW21, LAKA21, LAP23, LSS24, MRGP22, MAAK23, MQEK21, MY24, MMS24, Mon21, MRAM^{+}21, MLC^{+}21, MAAA22, NET21b, OKJ^{+}21, OU22, OAA22, OE22a, OE22b, ÖSTY22, PT22, Pan23, PMC^{+}21, PWJ^{+}21, PD22a, PP20b, PVP^{+}20, PB22b, QPS20, RV21, RAN21, RN22, RK23a, RR22, RRIL22, RG22, RKuH^{+}20, RAG21, SR22a, SKPD23, SAB22, SHA^{+}22, SS23e, SVD22a, SSN22, SS22b, SS24, SZZ^{+}22, SU23, SZS20, SYG22, SBSK22, SSN23, SYRP22, SP22c, SWZW20, SGHL20, SR22d, SP21b, SZZ^{+}24b, SS22g, SBS22, SKR^{+}24, DM22a, THW21, TTA20, TZZ^{+}23, TMB24, TT22, TP20, TO22, TAH22, TS22, TSD23, VPGK23, VS22b, VSP22, VR22, WHC^{+}20]. model [WZ20a, WWG21, WYL^{+}22, WHH^{+}23, WYE23, WZ20b, WZX^{+}22, WN23, XSGL20, XCJ22, XHZHXBQX22, XWW^{+}21, YWQ^{+}21, YSK22, YLZW23, YB23b, YSH^{+}22, Yil22a, Yil21b, YYY^{+}23, YWC^{+}23, ZNDA22, ZZL^{+}22, ZML^{+}23, ZWL^{+}23, ZXL^{+}21a, ZYXX23, ZWY^{+}24, ZFZ^{+}20, ZOS^{+}21, Zhu22a, ZLX^{+}24, ZCW^{+}23]. model-based [AMM^{+}20, GBS24, GK22b, IMNC22, JYC^{+}21, KCP^{+}22, LSW21, YWC^{+}23]. model-checking [YSH^{+}22]. Modeling [CPYC21, FDY21, HSO^{+}21, USK23, HLDT^{+}23, K22, LDZ^{+}22, SYRS^{+}22, VPGK23, WZZ^{+}20, XZXV21, YSC^{+}20, AML^{+}22, AS22c, BPM^{+}22, BDG^{+}23, CLC^{+}20, COZ21, DZLH20, GKG^{+}20, HOS^{+}21, HMR24, KYZ20, KQK^{+}20, KKE^{+}22, KAO22, KK22e, LGT^{+}23, MNR^{+}22, PS22a, PBD^{+}21, RF23, ST23a, SAAAA22, SSI22, THT20, VV23, WGDP23, WZB21, Wri22, WLX21, WLDW22, WTL23, WT23, YLKC24, ZWCS20, ZWL^{+}20]. Modelization [WLG^{+}22]. Modelling [BBB^{+}20c, UZAA21]. Models [LM20a, Oh21, SCH22b, Aka22b, AK22c, ASRN23, ATC23, ADA22, Bad23, BTP^{+}21, BPT^{+}23, BC23b, BB23, CD24a, CS21, ESS23, FSWW21, HK22a, IS23, JNS22, KA22b, KNK22, LM22a, LAK22, MY24, NBK22, OCD22, OO23a, ÖK22, ÖTT23, SLLMTS^{+}24, SKCS23, SET^{+}22, SCH22a, WXLD21, YK23, Yil22b, ZBC^{+}21, ZST^{+}23]. modern [ASA^{+}21, BFM^{+}23, BAR21, FAM22, SJA^{+}22, TH22]. modes [BPW^{+}20]. modification [Cob22, LCZ^{+}20b, Pal22]. modifications [TLXX21, YA22b]. Modified [ARFA21, ARS22b, Cek22, NM20, OU22, SBS22, WBL22, ABESEh20, ASC22, BDR23, BJGF20, BMG22, CHLD23, DR22c, DK21, DR23, EKS20, GFA21, GM22b, HR22a, KK22f, LS23b, LF23, MSJ22, MSK22, MCT22b, NM23, Ngu21, RZ21, SKR^{+}24, XWW^{+}21, Yu21, KR22c]. modified-flower [MSK22]. Modular [ZZY^{+}22b, DSS^{+}23, FFLM21, ZSQ22]. Modular-based [ZZY^{+}22b]. modularity [CYZ^{+}21]. Modularized [KDS22]. modulation [Ano21-37, SA22b, SK20a]. module [GCQ^{+}24, KSJK21]. modulo [XCJ22]. MOEA [PLZL24]. mole [SS23f]. mole-rat [SS23f]. molecular [AKLK24, BEÖ23, KPF^{+}20, ZHW^{+}20]. molecule [GS21]. Moller [WZ21, DS23b]. moments [SAHAN22]. momentum [NBK22]. monarch [BP23a]. monetizing [AKA^{+}22a]. mongoose [AK22a]. monitor [ASJ24]. Monitoring [ST21, AFF22, AM20a, ANP^{+}20, wGTC22, GSS23b, HT21, Hem22, IMNC22, JAC^{+}21, KCP23, LBA23, LA22, LCR^{+}24, LLC21a, LYW^{+}21, LSZ^{+}23, MBO^{+}21, NS23, PUL20, TKG^{+}23, VSP22, WGYZ22, YAR22]. monkey [ASRN23, KK22d, KDS^{+}23, MK22b, RK21b, ZGLS21]. monotone [ZGWZ23]. monotonically [KML23]. monsoon [BBD23]. Monte [YK23]. morphism [OO22]. morphism-based [OO22]. morphological [GM22b]. moth [GNS22, KS22d, NBPR22, SYG22, VB22b, WDW^{+}22]. moth-flame [WDW^{+}22]. motifs [LZL^{+}20b]. Motion [HKP21, CC22b, GBB22, HSL^{+}22, SRGK24, TAB21, WWW24]. motor [MHL^{+}20]. mould [SS23f]. mouth [PB21]. movement [DM23, DKC^{+}21, SZGR21, ZGH^{+}22]. moving [CY24, DTK24, HAA23, JHZ20, KHK^{+}23, LHK22, NBS^{+}22]. MP [BS23a]. MP-HTLC [BS23a]. MPI [SBG20, BBB^{+}20b, CdOO^{+}20, CLE^{+}20, DJJR22, FGL^{+}20, GMA20, GADM20, HGX^{+}22, LFW20, LZZ^{+}20, MB21, NWT21, SWCB20, TRM^{+}24, USP^{+}23, WK20]. MPI-AllReduce [NWT21]. MPSoC [BWS^{+}21, GBBS21]. MQoS [LLW^{+}23]. MRI [AJNS22, AMM^{+}20, DAK22, KEK^{+}20, TAT22]. MRMR [SBSK22]. MRMR-based [SBSK22]. ms [XZL^{+}20a]. ms-PoSW [XZL^{+}20a]. MSO [HUJ22]. MSSA [HZZC22]. MSSA-Net [HZZC22]. MSSBiLS [DM22a]. MSU [JCGT24]. MSU-TSCH [JCGT24]. muliclass [dOdMC^{+}20]. Muller [QNZP22]. Multi [DQF^{+}23, ELIG23, FMNF22, GKLS23, GSG^{+}23, GP22b, GNS22, HLG24, JW21, KT22b, KSS24b, LGL^{+}21, LDS^{+}23, LCZY20, Liu24a, LZT^{+}23, MHB22a, MSK22, NBK22, NNX^{+}23, RS23, SAM^{+}23, TJZ23, WER^{+}23, XZXV21, XR21, Yl20b, ZZL^{+}22, ZZW^{+}22, AAEA20, ABCP23, Ano21a, AS23c, AMA22, BV22b, BKD22, BZT^{+}23a, BVP22, BK22b, BYL20, CdOO^{+}20, CNG^{+}20, CÇY22, CCL^{+}22, CZCM23, CZL^{+}22, Cza23, DWZ^{+}22, DRR22, DAK22, DCZ^{+}22, DGSB20, DSC^{+}21b, DS22c, ETKD23, FSA^{+}22, FC21, FM20, FHS^{+}22, FCZ20, GBBS21, GJPS23, GSS^{+}23a, GZC^{+}22, GR22b, GBS24, GCSZ24, GK22b, HOS^{+}21, HMR24, HKA20, HD23, HNSS22, Hua20, HZZC22, IGYS23, JPK22, JNM^{+}22, KRKM22, KGM23, KBL^{+}21, KW21, KSP21, LLL^{+}24, LZL^{+}20b, LYF^{+}23, LLX^{+}24, EGGG23, MK22b, MSJ22, MNDK22, MS22b, MPG20, Mir22, NR22, OO22, PLP22, PK23b, PPM22, PLZL24, PKR22, PKSM24]. multi [RKR22, RAaB21, SB23a, SAN24, SSMT22a, SMD22b, SBA22, SN22a, SLG^{+}20, SH22, SYJL20, Shi22a, SM24c, SS23g, SFJ^{+}21, SSRA23, TFZC23, TY22, TRM^{+}24, USI21, VS22a, VCFZ20, WHC^{+}20, WWZ^{+}22, WLY^{+}22, XLL^{+}21, XZL^{+}20a, XDJY21, XZYW23, XYFZ23, YYPR22, YNK^{+}23, YLZC23, YLKC24, YCY20, ZZQ^{+}22, ZCH^{+}23, Zhu22a, TWH^{+}22, BAR21, CS21, GN21a]. multi- [CdOO^{+}20, HOS^{+}21]. multi-/manycore [HOS^{+}21]. Multi-agent [RS23, DCZ^{+}22, GR22b, SMD22b]. multi-aggregator-based [SSRA23]. Multi-angle [LDS^{+}23]. multi-attribute [KGM23]. multi-authority [SAN24]. multi-auxiliary [BK22b]. multi-channel [SH22]. Multi-chaotic [KSS24b]. multi-chip [MPG20]. multi-classification [CÇY22]. multi-cloud [PK23b, XZYW23, XYFZ23]. multi-constraints [PLZL24]. Multi-controller [XZXV21]. Multi-core [ZZL^{+}22, HNSS22, IGYS23, EGGG23, YCY20]. multi-criteria [DRR22, GK22b, SM24c]. multi-dimensional [HD23, LZL^{+}20b, XDJY21]. multi-disease [PLP22]. multi-document [GSS^{+}23a]. Multi-feature [Yl20b, GZC^{+}22]. Multi-GPU [XR21, BVP22, HMR24, VCFZ20, TWH^{+}22, BAR21, CS21]. multi-HCA [TRM^{+}24]. multi-input [Ano21a, SYJL20]. multi-Internet [FCZ20]. multi-key [BV22b]. multi-keyword [HKA20]. Multi-label [TJZ23, WER^{+}23, MSJ22, WLY^{+}22]. multi-layer [LYF^{+}23, Mir22, SN22a, SS23g, XLL^{+}21, YYPR22]. Multi-level [ZZW^{+}22, CZCM23, LLX^{+}24, RKR22]. multi-machine [MNDK22]. multi-modal [DSC^{+}21b]. multi-modality [DAK22]. multi-node [BVP22, Cza23]. multi-object [JPK22]. Multi-objective [FMNF22, GSG^{+}23, GNS22, KT22b, Liu24a, LZT^{+}23, MHB22a, MSK22, SAM^{+}23, AMA22, CNG^{+}20, CZL^{+}22, DS22c, FC21, GBS24, KBL^{+}21, KSP21, MK22b, MS22b, PPM22, PKR22, SB23a, SSMT22a, Shi22a, WHC^{+}20, YLZC23]. multi-organ [LLL^{+}24]. multi-orientation [AAEA20]. multi-output [Ano21a, SYJL20]. Multi-party [GKLS23, NR22, OO22, GN21a]. multi-path [BZT^{+}23a, SAM^{+}23]. Multi-perspective [ELIG23]. multi-ported [YNK^{+}23]. multi-preference [JNM^{+}22]. multi-product [FMNF22]. multi-project [USI21]. multi-queue [FSA^{+}22]. multi-replica [TY22]. multi-resolution [BYL20]. Multi-resource [JW21]. multi-robot [CCL^{+}22, DWZ^{+}22, KRKM22]. Multi-scale [LCZY20, FHS^{+}22, Hua20, HZZC22, TFZC23, ZZQ^{+}22, ZCH^{+}23]. multi-server [XZL^{+}20a]. multi-set [AS23c]. multi-source [WWZ^{+}22, Zhu22a]. Multi-step [NBK22]. multi-strategy [LZT^{+}23]. Multi-strategy-based [DQF^{+}23]. Multi-swarm [LGL^{+}21, BKD22]. Multi-task [HLG24, GBBS21]. multi-tenant [RAaB21, SAM^{+}23]. Multi-texture [GP22b, PKSM24]. multi-threaded [DGSB20]. multi-unmanned [ETKD23]. multi-user [GCSZ24, SFJ^{+}21, VS22a]. multi-valued [FM20]. multi-verse [ABCP23]. Multi-view [NNX^{+}23]. multi-workgroup [KW21]. multi-WSN [GJPS23]. multiagent [SPA^{+}21, ZLD22]. multiarmed [WWW^{+}20]. Multiattribute [XLWX20]. multibox [BST^{+}22]. Multicast [PSP^{+}20]. multicasting [SDR23, SL20]. multicategory [KST23]. Multicellular [HYK21]. multichannel [LHWT20]. Multiclass [VPSM22, BJS23, FAM22, VKK^{+}24]. multicloud [CDN^{+}21]. multicollinearity [AU22, ADA22, KÖÖG22, MAAK23, SAB22]. multicore [AZM20, AAG^{+}22, BBF21, BEJD22, BSPM23, CTA^{+}23, JBBH21, LLTZ24, SKA23, SCKU23, SPC^{+}21, YY20b]. Multicores [SCH22b, SCH22a]. multicriteria [CMA^{+}21]. multicrossbar [AY21]. Multidimensional [Ano21-35, HWBZ21, LSD21, LHL^{+}20, PS22b, ZLC^{+}22]. Multidomain [ZZK^{+}22]. Multienvironment [PS21]. multifactor [CBK23]. Multifeature [LZF20, ZGG^{+}23, LCW^{+}23a]. Multifeature-based [LZF20]. multiforme [DBN^{+}22]. multigranularity [WZY^{+}22]. Multigraph [SPWX23]. multigrid [MW21]. Multihop [Ngu21, SPK^{+}22a, TK22b]. multiion [ZZS^{+}21b]. multikernal [APM^{+}21]. multikey [LWW23]. Multilabel [aWLY^{+}22, Akb22, DDZ^{+}20, SLL22]. multilabeled [AV22b]. multilateration [SM24b]. Multilayer [HLZ23, KM23a, GM22a, JSAA22, KSK23, LC21, OBTC20, The21]. Multilevel [ABB22, SS22e, BMK^{+}20, CCD^{+}20b, DMS^{+}21, GWA^{+}23, GWA^{+}24, KRSR23, KK22f, PCK23, THW21]. multilingual [KS22c]. Multimedia [MBM^{+}20b, ABESEh20, CJ21a, CL22b, DS24a, Ogi20b, ORP21, SKR^{+}24, WLZ^{+}21b]. Multimetric [SSV24]. Multimodal [SZqWZ20, APM^{+}21, AJK^{+}23, BK22c, CZG^{+}20, CZ22, DWDG20, LCW^{+}23a, MG21b, SS23c, SCM22]. multimodel [UE22]. Multimodular [APM^{+}21]. multinomial [LHC^{+}23]. Multiobjective [GWGR20, ZZY22a, BPW^{+}20, MKA23, TS22, VPGK23, YLW^{+}22]. multiomics [CSWC20]. multiparty [BS23a]. multipath [CL22b, HUJ22, LWZC21, RT22b]. Multiplatform [NJ21, SI22]. Multiple [DS20a, LLJR21, AMAT22, Akb22, AEMS24, AYG^{+}21, CWW23a, CB22, CMLL21, DSYF22, DZCL22, FZA22, FXX22, KSK^{+}20, KAR^{+}23, LM22a, LZXG22, MES23, NTK22, NRMB23, NKST24, NRP^{+}20, PMS^{+}21, RN22, SHL^{+}22, SGH23, SA22d, SM21, TSG21, ZWL^{+}20, dCMA23]. Multiple-replica [LLJR21]. multiple-session-keys [AMAT22]. multiple-watermarking [TSG21]. multiplexed [SA22b]. multiplexing [Ben22, SO22]. multiplexing-based [Ben22]. multiplication [AML^{+}22, BFM^{+}23, HCG21, KS21b, XWD^{+}22]. multiplicative [HR22a]. multiplier [MKRK23, SPSP23]. Multipopulation [XWW^{+}21]. multiprocessing [LLR^{+}21]. multiprocessor [DSC^{+}21a, TS21, WFY22b]. multiprocessors [ALBZ21]. multiproduct [GMN21]. multiprogramming [UWF^{+}21]. multirank [LCW^{+}23a]. multiresolution [XCX^{+}20]. Multiscale [CWL^{+}21, CWL^{+}22, SZG^{+}24, SYT^{+}21, ZDJ^{+}21, CLL^{+}21, KKE^{+}22, LLN^{+}23]. multisensor [YCL^{+}22a]. Multiset [MdAA^{+}21, YSL^{+}21b]. multisource [aWLY^{+}22]. Multispectral [BT21, LYW^{+}21]. multistage [LZY^{+}21]. multisurface [SBGC21]. multitask [CBFS23, PCR21, WWM^{+}23]. multitenancy [RYG^{+}21]. multitenant [MK23, ZFW23]. multithreaded [CDP^{+}21, Cza23, GB20, KS21b]. multithreading [AÜ23, Boz22a]. multitiered [LJ22]. multitopology [LHL^{+}22]. multiuser [CJ21b, WSZ^{+}23]. multivariate [KG22, The21]. multivector [AAQOT23]. Multiview [MLZ^{+}21a, YSL^{+}21b, ZWZ^{+}21, ZWL^{+}20]. MuMMI [WTL23]. murmuration [DCR23]. music [Kim21b, Kum22]. musical [GSB21]. Mutable [MTD^{+}20, AMAT22]. mutual [GYL^{+}21, JPH23, LLM^{+}24, LSX21, LRC^{+}22, NT23, PSP24, SHV24, WER^{+}23, ZFH^{+}23].

N [ZLYS21]. NAGNet [ZHT^{+}23]. naive [LHL^{+}20]. naked [SS23f]. name [BKK23]. Named [GDFDF22, TAH22, KC20, ZLW^{+}21]. names [HZW^{+}23]. Namib [CEN22]. naming [ZWL^{+}20]. NAND [ZWX21]. nano [Jia22, KA22a, SKS20]. Nano-Caches [SKS20]. nano-scale [Jia22]. nanocrystals [BEÖ23]. nanomagnets [AKLK24]. nanoscale [GM22a]. nanotechnology [AG22]. nascent [BBP^{+}24]. national [LXZ22, ÖSTY22, ZARR23]. native [CDC20, SKH^{+}21]. natural [VS21, RVJ^{+}22]. Nature [JC21b, KHA22b, LMMS24, MG21b, NAK^{+}22, OE22a, TT21, WWA22]. nature-based [TT21]. Nature-inspired [JC21b, KHA22b, MG21b, OE22a, WWA22]. naval [JH21]. navigable [XLWX20]. navigation [BSBF22, FPHZ19, KAO22, Pan20]. navigational [OA22]. NB [CLC^{+}20, SCS^{+}21]. NB-IoT [CLC^{+}20, SCS^{+}21]. Near [GRMP23, VKK^{+}24]. near-real-time [VKK^{+}24]. Nearest [NLB22, Ben21, FYH^{+}21, GZY^{+}22, KSY24, RRJ23, MKL21, WQY^{+}22, XLWX20]. NEC [VMFL23]. NEEBS [TAT^{+}23]. negative [AAA^{+}22a, JW22]. negotiated [BWTJ20]. negotiation [BF24, LMGG20, MBB22, OBTC20, ZTL^{+}21].

Neighbor [LSZ24, WSL^{+}20, Ben21, BM22b, GCQ^{+}24, GZY^{+}22, KSY24, RRJ23, MKL21, WQY^{+}22, XLWX20]. neighbor-based [RRJ23]. Neighborhood [CS23, WLY^{+}22, OO23b, MR23a, WER^{+}23, Yu21, ZLTX21]. neighbors [FYH^{+}21, LWZY23, NLB22]. neonatal [KD22]. NERSC [BBP^{+}24, YKW20]. NERSC-9 [YKW20]. nested [KSA^{+}21]. nesting [LW20a]. net [XWQ^{+}23, WZX^{+}22, BDR23, BMG22, HZZC22, KK22b, KMD23, LJP^{+}21, LZSC23, LLX^{+}24, XLL^{+}20, YKL23, ZCX^{+}21, JMV24, LZF20]. NET-BOWTIE [WZX^{+}22]. nets [YB23b]. Network [CBFS23, CON23, GMM22, HWG23, HFFA20, KSSP22, RT21, AC22, AKEC20, AK22a, AMRH21, AAK^{+}21a, Agr21, AS22b, AMB23, ATS22, AFBM^{+}23, AY21, AMV22, ABCP23, AA23c, AA23d, AZA20, AJS23, ASC22, Ano21a, AA23e, ASRN23, AGM23, AFM22, AY23, BSN22, BZWH21, BBD23, BKK22, BKJ22, BA24, BM22b, BM22c, BJK^{+}24, BKS22, BRNR23, BZGM22, BMJ^{+}20, CLDY21, CZG^{+}20, CÇ22a, CS23, CLC^{+}20, CPYC21, CYC21, CW22, CZ22, CCM22a, CCZM23, CYZ^{+}21, Cho20, CD22, CPCK23, CM21, CA23, DL23, DRV22, DRM22a, DJF21, DP22a, DT22a, DCR23, DRR22, DG21, DZLH20, DBD23, DSSS22, DCWM20, DM21, DM22d, EVVR21, EHST21, FK24, FRS^{+}23, GDSS22, GPDB20, GLW22, GDCW22, GRL^{+}22, GGA24, GVSS22, GM22b, GNS22, GATK22, Gün23a, Gün23b, HZD^{+}22, HLC23, HYG^{+}23, HUJ22, HBB20, HXST22, HJZ^{+}22, HLZ23, HLW^{+}21]. network [HZZC22, HSL^{+}22, DR22a, ITO21, IQS^{+}22, IK22, JMV24, JAM24, JPK22, JS22c, JS22a, JA23, JNS22, JQ22, JZB20, KTM22, KH22, KE21, Kab22, Kab23, KPJ^{+}21, KSJK21, KKKS23, KÖ22, KA21a, KT23, KVV20, KR22a, KBR24, Kay22b, Kha22a, KKP20, KMD23, KAO22, KMS^{+}22, Kot23, KP21, KBS^{+}22, KSK23, LKR^{+}22, LPHK20, LSZL23, LG23, LWLZ20, LN20, LZF20, LZC^{+}20, Li21, LSD21, LYL21, LLZ^{+}21b, LDZ^{+}22, LLW^{+}22a, LLL^{+}24, LLW^{+}22b, LX22, LSW^{+}20, LGLZ20, LWHW22, LLJ^{+}22b, LZSC23, LCW23b, LGT^{+}23, LL23, LLL23, LWZ^{+}20, LGX^{+}23, LHL^{+}20, MGSB23, MCT22a, MSJ22, MAR23b, MKA23, MY24, MLKD20, MPV22, Mis22, MS23, MRS^{+}21, MRKY22, MG23, MCT22b, MP23, MSBR23, NBK22, NSR22, NBPR22, ND23, NFF21, NSK^{+}24, NWZ^{+}21, NLG^{+}20, OKJ^{+}21, ÖGS22, PBK23, PT22, PSP22, PSMM22, PK22a, PKB22, PS23a, PLX20, PA23]. network [PD22b, PB22b, PSK23, PVVS22, AST22, QWL24, RKR22, RM21, RJ22, RGKK21, RPPK23, RSS20, RR23b, RMMK23, RK23b, RSKA23, RPMA22, RAN22, RRIL22, RW24, RZ24, SJ22, SKPD23, SSMT22a, SSMT22b, SC22a, ST22b, SMD22a, SRG^{+}22, SGS21b, SKP22, SK21a, SBA22, ST23a, SV22b, SMBM24, SWK22, SMD^{+}21, SVM24, SSN22, SVD^{+}22b, SRRM23, SS24, SR24, SZqWZ20, SM22c, SKK22, SPS22, SS23f, SSN23, SMKA22, SR20a, SP22c, SCS^{+}21, SPHP21, SZL^{+}22, SZG^{+}24, SSP24, SRIB23, SBB^{+}20, SP23b, SM24c, SHVA23, SAMS23, SPK22b, SSH22, SLHW20, SYT^{+}21, SS22g, SSS23b, SBS22, SRS23, SM23d, TZ22, TAT^{+}23, TFZC23, THX22, TB23, TC24, TÖ21, VKSS23, VRR^{+}22, VH22, VJ22, VRV23, VS22b, VKK^{+}24, WSM^{+}20, WLCW20, WZZ^{+}20, Wan20, WWG21, WMC21, WZHL21, WLZ21a, WSZ^{+}23, WWJ^{+}20, WWL^{+}20]. network [WZSZ20, WYZAD20, XGX^{+}21, XCX^{+}20, XHZHXBQX22, XSZ^{+}20, YRV^{+}23, YS22, YLLL20, YMWA21, YLZW23, YY20a, YLW^{+}22, Yib22a, YF20, YSC^{+}20, ZDH^{+}22, ZWCS20, ZLW^{+}21, ZYX^{+}21, ZSC^{+}21, ZLT21, ZLL^{+}22, ZMLW23, ZKZ^{+}23, ZCX^{+}21, ZZZ^{+}20, ZDJ^{+}21, Zhu22a, GBBS21, PKSM24, TCRP23, ZX21b]. network-attention [BKJ22]. network-aware [KVV20]. network-based [AC22, KTM22, KT23, MLKD20, PSP22, RSKA23, SMD^{+}21, SR24, SBB^{+}20]. network-driven [BJK^{+}24]. network-extreme [RR23b]. network-fractional [Kha22a]. network-long [PT22, SMBM24]. network-morphological [GM22b]. network-on-chip [AY21, GBBS21]. network-on-chips [ITO21]. network-taking [PLX20]. networked [ZZL^{+}20]. Networking [BI23, ETH^{+}23, MV22, TAH22, AAMM23, CK24b, KT22a, KJMB22, LCZ^{+}20a, LZZ21, LLW^{+}22b, LGLZ20, LZXG22, MK22b, SKK22, VK24]. Networking-based [TAH22, AAMM23]. Networks [AJNS22, WXCY20, ZLZ^{+}22a, AP22, AB20a, AM20a, AAMM23, AK21, ABSS22, AAARR20, Ano21-35, Ano21-39, ASMS21, AT22, AGSN23, AMM^{+}20, BV22a, BMZ^{+}22, BK21, Bei24, BLT22, Ben21, BKLY20, BGPQ21, CCBA23, CAS24, CK24a, CSWZ22, CLC^{+}23, CL22b, CG22, CD22, CÇ22c, DRV22, DGM21, DK21, DRMA22, DPSJ22, Eke22, EBDB22, ETKD23, FC21, FCX^{+}22, FLB23, FSFM22, FZT22, FD20b, GXL^{+}20, GJBM22, GSTS22, GA23, GYZ^{+}20, GSS23b, GM21, GMK^{+}21, HLL^{+}21, HLG24, IA22, IAQ20, JSS22, JPN21, JH21, JB22b, JCGT24, JHS^{+}21, JQGL20, JPO^{+}21, K22, KE22, KDS^{+}20, KMR22, KGM23, KSJ23a, KGGM22, Kim21b, KJHM21, KK22c, KRSR23, KSS24a, LCSR21, LLJ^{+}22a, LLW^{+}20, Li20, LPW^{+}21, LWZ^{+}22b, LL21, LGW^{+}22a, LTL^{+}20, LSQW21, LHL^{+}22, LLL^{+}22, LAC21, LS24, LLL^{+}21, LF23, MWH^{+}23, MTT20, MES23, MIN^{+}23, MS21c, MS23, MCNR20]. networks [MSS^{+}20, NVV24, NAR^{+}22, NR22, NMQ22, Ngu21, NGD^{+}22, Ona21, OCD22, OE22a, PPA22, PAN22, PKK23, PM23b, PDRM24, PA21, PD23, QLL^{+}22, QHL^{+}24, RC22, RMMK23, RK23c, RP22, RG23, RT22b, SS23a, SPA^{+}21, SJ22, SDR23, Sao22, SPS23, SAD^{+}21, SLJ23, SH22, SG24, SZGR21, ST23c, SKSB20, SS22d, SYRP22, SM24b, SCS^{+}21, SD23, SPK^{+}22a, SS22f, SM22d, SKR^{+}24, SDR20, SAGS24, TA23a, TTM^{+}22, TSLL21, TT22, TK22b, TBNPQ23, TSB23, TKG^{+}23, URK^{+}22, VB22a, VVK23, WKY22, WZMJ23a, WL23, WSJ^{+}21, WCWG21, WLDW22, WJD22, WYLL24, XLXZ20, XXD^{+}22, Yal22, YW22, YYPR22, YB23b, YNK^{+}23, YLGY20, ZX23a, ZCZ^{+}22, ZCH^{+}23, ZWZ^{+}21, ZDL^{+}22, ZWL^{+}20, ZWC^{+}22, ZCR23, dRdSZ^{+}23, uZKH^{+}20, IVP^{+}23]. networks-based [LGW^{+}22a]. Neural [AJNS22, CKL20, PSP22, ZLZ^{+}22a, AC22, AAK^{+}21a, AMB23, ATS22, AFBM^{+}23, AMV22, AAARR20, AZA20, AT22, AMM^{+}20, AFM22, AY23, BSN22, BBD23, BKJ22, BK21, Bei24, BM22c, BJK^{+}24, BRNR23, BZGM22, CÇ22a, CK21b, CCZM23, Cho20, CA23, DRV22, DRM22a, DS21, DG21, DBD23, DM21, EBDB22, FK24, FC21, FZT22, FRS^{+}23, GDSS22, GPDB20, GLW22, GRL^{+}22, GSTS22, GGA24, GVSS22, GM22b, GATK22, HLC23, HYG^{+}23, HUJ22, HXST22, HLG24, IK22, JSS22, JMV24, JAM24, JPK22, JA23, JNS22, JPO^{+}21, KTM22, KE21, KE22, Kab22, Kab23, KPJ^{+}21, KSJK21, KKKS23, KA21a, KT23, KR22a, KBR24, Kay22b, Kha22a, Kim21b, KMD23, KAO22, Kot23, KRSR23, KSK23, LKR^{+}22, LPHK20, LSZL23, LWLZ20, LN20, LZF20, LZC^{+}20, LGW^{+}22a, LLJ^{+}22b, LGT^{+}23, LL23, LLL23, LWZ^{+}20, MCT22a, MSJ22, MWH^{+}23, MIN^{+}23, Mis22, MS23]. neural [MRS^{+}21, MCT22b, MSBR23, NBK22, ND23, NSK^{+}24, NMQ22, NWZ^{+}21, OKJ^{+}21, Ona21, OCD22, OE22a, ÖGS22, PBK23, PT22, PAN22, PSMM22, PK22a, PKB22, PLX20, PA23, PD22b, PVVS22, AST22, QWL24, RM21, RK23b, RSKA23, RRIL22, RW24, SPA^{+}21, SJ22, SSMT22a, SSMT22b, SC22a, SMD22a, SRG^{+}22, SKP22, SK21a, SBA22, SMBM24, SAD^{+}21, SMD^{+}21, SSN22, SS24, SZqWZ20, SM22c, SSN23, SMKA22, SP22c, SPHP21, SSP24, SP23b, SHVA23, SAMS23, SPK22b, SM22d, SLHW20, SS22g, SBS22, SRS23, TZ22, TAT^{+}23, TSLL21, TFZC23, TB23, TBNPQ23, TÖ21, VRR^{+}22, VJ22, VRV23, VRS22, VVK23, WSM^{+}20, WWG21, WZMJ23a, WSZ^{+}23, WWJ^{+}20, WSJ^{+}21, XCX^{+}20, XW24, YRV^{+}23, YS22, Yal22, YLLL20, YF20, ZLW^{+}21, ZSC^{+}21, ZCH^{+}23, ZDL^{+}22, ZCR23, dRdSZ^{+}23]. neural-fuzzy [JSS22]. neural-network-based [CCZM23]. Neuro [MRAM^{+}21, ABCP23, JKP22, KS23b, MGSB23, NSR22, RKC22, SP22b]. Neuro-fuzzy [MRAM^{+}21, ABCP23, JKP22, KS23b, SP22b]. neuroimaging [DHSG23, VPB^{+}23]. neurological [VPSM22]. neuromorphic [PK23a]. neuron [WYE23]. neutrosophic [RNRK22]. news [CZ22, DBD23, KM21a, MSJ22]. Newton [MLVGR24]. next [ETH^{+}23, HJT^{+}20, WS21, WC22]. next-generation [HJT^{+}20, WS21, WC22]. NFV [PCG^{+}21]. ngrams [BD21]. NIDS [Gün23a]. Nimble [STISM21]. nine [WSM^{+}20]. nine-layer [WSM^{+}20]. Niyama [TKS22]. NLOS [AC23, ZXL^{+}21b]. NLP [AAE23]. NLPHub [CPPP21]. NLSP [BXH^{+}23]. NLU [AAE23]. NMIEDA [LSX21]. NN [MMS24]. NNPEC [CS23]. NO [xZlGCzJ20]. NoC [USP^{+}23]. NoC-based [USP^{+}23]. NoCs [RJH^{+}23]. Node [TKS22, WZHL21, AC23, BM22b, BVP22, CS23, CW22, CG22, Cza23, DD21, DT22b, DMPR24, ECIB20, HUJ22, HNSS22, KJHM21, LXYY21, LLJ^{+}22a, LMM23, MAG^{+}20, MS21c, MSL22, MP23, SA22a, SSM^{+}21, XGX^{+}21, XHST20]. node-based [XGX^{+}21]. Node-Fusion [WZHL21]. nodes [AS23a, DRV22, DP22b, FZC20, LSSQ22, NQ21, RP22, Sin23, VBM^{+}21, WZHL21]. nodule [TFZC23, ZKZ^{+}23]. nodules [MSBR23, ZZW^{+}22]. Noise [WWM^{+}23, ARHT21, MRS^{+}21, MH23, YYZS22]. noises [AS22d]. noisy [RKMS24]. NOMA [Mir22]. Nominator [ARC22]. Non [EAvM20, FPÁ^{+}20, HXY^{+}22b, JW22, Kim21a, KPA23, LGX^{+}23, RPM22, SDSW21, WLL21a, WPL20, Yil21c]. Non-blocking [EAvM20]. non-cache-coherent [RPM22]. non-dominance [HXY^{+}22b]. non-dominated [HXY^{+}22b, Yil21c]. non-face-to-face [Kim21a]. non-Hermitian [WPL20]. non-intrusive [FPÁ^{+}20]. non-linear [SDSW21]. non-negative [JW22]. non-repudiation [WLL21a]. non-robust [LGX^{+}23]. non-volatile [KPA23]. nonblocking [DJJR22, ZYPD24]. noncommunicative [DM23]. nonconvex [FZZZ23]. noncooperative [DZCL22, GWGR20]. Nondominated [Wan20, DCWM20]. Nonexpert [TAT^{+}23]. noninterference [HLT23]. nonlinear [DWZ20a, LFWJ22, SMD^{+}21, VP22a, YSL^{+}21b]. nonorthogonal [YYZS22]. nonsensitive [SAHAN22]. nonsmooth [FZZZ23]. nonsubsampled [AJK^{+}23]. nonsymmetric [MM21]. Nonvolatile [RK21a]. norm [PK22b]. normalized [LSX21, ZQW^{+}21]. Northwest [ZWJG21]. NoSQL [APM22]. notebooks [VZR^{+}21]. Novel [Ano21-37, BYPO21, BS23b, BFM^{+}21, BK23, CPH20, Kal22, KAAR23, LGT^{+}23, VK23a, AM20b, Akb22, AD22b, AA23b, AI22, AV22b, Ano21-38, AR22b, Asl22, BCM22, BWW^{+}20, BXH^{+}23, BCK22, ÇKÇ22, CYQ^{+}20, DL23, DR22c, DBK21, DP22b, DS24b, EKS20, FLG^{+}22, GA22, GK23a, GD22, GK23b, HYC^{+}24, HBB20, HZZC22, JYC^{+}21, KRKM22, KE21, Kab22, Kab23, KMR22, VK24, KBK^{+}22, KFML20, KKC22, KA21b, KR22c, KS23c, LTW^{+}24, LCZ^{+}20b, MK22a, MIN^{+}23, MDSG^{+}23, MYCH22, Mir22, MJAM23, NM23, NA22b, NN23, OHRS21, PS22a, Par22, PSP24, PS23a, PP20b, PB21, RTBC23, RMMK23, RZ21, RÖ22, SVM24, SSN22, SZY^{+}22, ST23c, SS23f, SJ23a, SYRP22, SAL22a, SR20b, SHVA23, TWW^{+}21, TT21, TSV^{+}22, TUD21, VPGK23, VKK^{+}24, YYPR22, YA22a, YLZC23, Yil23, ZHX^{+}21, ZFH^{+}23, Zhu22b, ZHT^{+}23]. novel-cascaded [AA23b]. novelty [WWW^{+}20]. NSCT [KSK^{+}20, SKS^{+}23, TSG21]. NSCT-RDWT-SVD [SKS^{+}23]. NSGA [GRG^{+}22, WHC^{+}20]. NSGA-II [WHC^{+}20]. NSGA-II-XGB [GRG^{+}22]. NUMA [LLR^{+}21, PEGP23, YCY20, YY20b]. NUMA-aware [PEGP23]. number [BA20, HNOI24, JOW23, OHFF20, RSKA23]. number-theoretic [JOW23]. numbers [AA22b]. Numerical [STJ^{+}20, WWW24, WLX21, XSZ^{+}23, AFG^{+}22, ÇGB23, GMSM21, HPH^{+}20, HOS^{+}21, SS23f]. numerics [TCA23]. nurse [SAC22]. nutrition [ÖSTY22]. NVIDIA [HMR24]. NVIDIA- [HMR24]. NX [ZZL^{+}22].

O [LCKJ21, PMP23, RZVC21, YSL^{+}21a]. O-intensive [BPM^{+}22]. O2O [PLX20]. obfuscation [ABSS22, CL22a, HK22b, HE23]. Object [BTT21, HWY^{+}23, JPK22, LWL^{+}24, SS22g, MS22a, BRNR23, FAM22, JNS22, LAE^{+}22, LXJ^{+}22, ML20, MSB^{+}20, NBS^{+}22, NET20, NN23, SI22, TAT22, WMC^{+}23, XZG^{+}23, ZQW^{+}21, ZWZ^{+}21]. Object-aware [LWL^{+}24]. object-based [NET20]. object-centric [MSB^{+}20]. objective [AMA22, CNG^{+}20, CZ21, CZL^{+}22, DS22c, FC21, FGZC23, FMNF22, GBS24, GSG^{+}23, GNS22, HPCK22, IAA20, KBL^{+}21, KML23, KT22b, KSP21, Liu24a, LZT^{+}23, MHB22a, MK22b, MS22b, MSK22, PPM22, PKR22, RZ21, SB23a, SAM^{+}23, SSMT22a, Shi22a, WHC^{+}20, XGCZ23, YLZC23, YZ21, ZX20, ZWL^{+}23, ZWCC23, ZZC24, ZWY^{+}24, Zhu21, Zhu22b]. objects [ABESEh20, BA20, CY24, LHK22, SG22c, Shi22b]. obligations [GLM^{+}22]. observable [CJC22, KC22]. observational [VSK22]. observations [DM23]. observed [SZS20]. observer [DD21]. obstacle [WLLX21]. obstacles [SSSP21]. OCCI [BF24]. occluded [ST23a, Shi22b]. Occlusion [LLW^{+}21]. occupancy [ENB^{+}20]. occurrence [SYRS^{+}22, SRIB23]. OCR [DK22a]. OCTRA [KKS22]. OCTRA-5G [KKS22]. ocular [PGD^{+}22]. odd [IAT^{+}23]. ODE [KW21]. OFA [LLM^{+}24]. OFDM [Ano21-38, KR23]. off [Ano21i, JW21]. off-blockchain [Ano21i]. offensive [CMJC23]. offline [GR22b]. offloading [BP24, LS24, LLL^{+}21, MNYN21, OO22, PLZL24, QLL^{+}22, QHL^{+}24, SWCB20, SG22d, TWQ^{+}21, TS24a, Wan22b, WDW^{+}22, XWC^{+}22, XGCZ23, YV23]. offshore [KWL^{+}24]. oil [BRL^{+}20, GMN21, SYRS^{+}22, SPC^{+}21]. OLAP [JWT^{+}20, RLdO20]. OMFTSA [RAG21]. on- [Ano21i]. On-chip [SCP20, USP^{+}23]. on-demand [CSWZ22, JGW20]. on-premises [MDSG^{+}23]. on-route [Sin23]. on-the-fly [PMP23, YPO21]. on-the-loop [FGJ^{+}21]. onboard [SK21c]. OnDemand [CFR^{+}21]. One [JGW20, ZGWZ23, CD24b, HK24, IAAA22, KW21, XLX^{+}21, ZZS21a]. one-dimensional [CD24b]. One-pass [ZGWZ23]. one-shot [XLX^{+}21]. One-sided [JGW20]. one-step [KW21]. oneAPI [MPB^{+}22]. onion [TT23a]. Online [HGW^{+}23, MB22, WLZ21a, WZY^{+}22, WRE^{+}23, AC22, ALNJ21, BABLH21, BKLY20, CYC21, DGM21, GM21, IA23, KMR22, LLZ^{+}21a, LYL21, LLM^{+}24, LGLZ20, LLC^{+}21b, MAW^{+}22, NJK22, PK22b, QCC^{+}23, RZVC21, SVM24, SLJ23, SYRP22, SV22e, VPB^{+}23, YB23b]. OnOff [uZKH^{+}20]. onto [FHH^{+}20]. ontological [HR22b]. ontologies [ZTL^{+}21]. Ontology [YL20a, ZTL^{+}21, FRS^{+}23, HL20, LMGG20, RF21, TSCM22]. ontology-based [LMGG20]. ontology-driven [RF21]. Opcode [LZC^{+}20, BD21, DDH^{+}20]. opcode-based [DDH^{+}20]. Open [CFR^{+}21, QWL24, DJF21, DPS21, FZ23, FMNF22, KK22a, LFX^{+}20, NMS^{+}21, SKK22, YSP23, OO23b]. Open-domain [QWL24]. OpenABLext [XAC^{+}20]. OpenACC [LBG^{+}20, XR21]. OpenCL [CZZ^{+}22, HSO^{+}21, RBC20, SI22]. OpenFAM [SCK^{+}23]. OpenFlow [AYKE21]. OpenMP [CdOO^{+}20, Cza23, DKB20, GTM24, LBG^{+}20]. OpenSHMEM [CM21]. OpenStreetMap [ML20]. OpenStreetMap-tagged [ML20]. operating [LCZ^{+}20b, SKK22, YLJ22]. Operation [ZARR23, DZLH20, FZC20, GZ20, ZZS21a]. operational [BPB21, MHPA21]. operations [ABB22, JH21, LFW20, MB21, RP21a]. operative [JDG22, Man21]. operator [DS20a, SSK23b]. operator-based [SSK23b]. ophthalmology [IIK^{+}23]. Opinion [Ano21-39, Aru22, CPYC21, CMK22, LYL21, MS22b, SSP22, WHJ^{+}20, YSC^{+}20]. opinions [ALNJ21, RM22]. Opportunistic [GJK^{+}20, Sin23, CK24a, XXD^{+}22]. opportunities [BB23]. opposition [BDK22, Lia24]. opposition-based [BDK22, Lia24]. Oppositional [KV22a]. Opt [CdRNB23]. OptDRN [MPV22]. optical [FSFM22]. Optimal [BZWH21, BKK22, CAS24, CJ21a, DT22b, Hem22, KEK^{+}20, KS22a, Kum22, LM22b, LST22, SS23d, ZZLZ22b, APM^{+}21, BV22b, ÇGB23, DKL21, EK20, FRS^{+}23, GNS22, GAM23, HK21, HAK22, IDA22, JB22a, KRKM22, KB21, Ker22, KKS23, LS23a, LL24, MGB^{+}23, MG23, MV24, NNVD22, NBS^{+}22, NM23, NKST24, PT22, PSMM22, PLP22, RCK22, SPS23, SVD22a, SLG^{+}20, SP23b, SRS23, SKR^{+}24, VY23, VB22b, VG21, YZYT21]. optimisation [BWTJ20, Gün23b]. Optimization [GWA^{+}23, GLJ20, IG24, KSS21, NAK^{+}22, PKSM24, PS23c, ST23a, SPS22, TZ22, AC22, ABESEh20, AFK^{+}22, AK22a, AE22b, AFBM^{+}23, ASL20, ASAAAA22, AS23b, APM^{+}21, ABA22, AZA20, AJ21, Anb22, AS22d, ASMS21, ASA^{+}21, AB21, AA23e, ASRN23, BSN22, BP23a, BD22, BKK22, BKD22, Ben22, BRS^{+}22, BM22c, CdRNB23, CF21, CEN22, CK23, CHMC21, CWW^{+}23b, CA22, CPCK23, ÇGB23, CDP^{+}21, CPA22, CZ21, CA23, DZW23, DBD22, DRV22, DS23a, DM22b, DS20a, DZCL22, DPYS22, Den24, DDUK23, DBK21, DS20b, DS22a, DCR23, DR23, DWZ20a, DK22b, DLC^{+}21, DD24, DPSJ22, FC21, FLG^{+}22, FGZC23, FDD^{+}24, GZ20, GSS^{+}23a, GCS23, GAS23, GKÇ22, GBS24, GK23a, GVSS22, GC20, GGJ^{+}24, GAM23, HK22a, HR22a, HAR20, HLL^{+}21, HGMK21, HXY^{+}22b, HXY^{+}22a, HHP23, HYT^{+}21, IAAA22, JT23, JB24, JSA^{+}20]. optimization [JA23, JZC^{+}23, JR23b, KIN^{+}23, KT20, KKR23, KMR22, KKKS23, KS23a, KBR24, KBL^{+}21, KV24, Kha22a, KFML20, KML23, KGGM22, KA22b, KK22d, Kot23, KWL^{+}24, KSP21, KK22f, KS22d, KR23, KDS^{+}23, KAR^{+}23, LHJ22, LK22a, LJBS23, LLAV22, LG23, LS20, LGL^{+}21, LNI^{+}24, LZC21, LXC^{+}22a, LCW23b, MT21, MGSB23, MK22b, Mir22, MG23, MCT22b, MV24, NSR22, NNVD22, NBPR22, ND23, NSKS21, NSK^{+}24, ODK^{+}23, PBK23, Pan24, PLP22, PKK23, PJP21, PVRM22, PSK23, PS23b, PD23, RSM21, RSM23, RM23, RPMA22, RP21c, RK21b, RG22, Rav23, RP22, RNRK22, RZCA21, RAG21, RZVC21, RKC22, SB23a, SA22a, SSMT22a, ST22b, SAC22, SS23d, SEC22, SMBM24, SBB21, SVD^{+}22b, SN22b, SRRM23, SKHL22, SB23b, SM23a, SDSW21, SP22a, SMR23, Shi22a, Shu22, SKSB20, SSK23b, SS23f, SMKA22, SCM22]. optimization [SKSP20, SR22c, wSYyCsD23, SD23, SKSP23, SFJ^{+}21, SSS23b, SBS22, SSRA23, SM23d, TÖK21, TT21, TS22, UDS21, VKSS23, VPGK23, VM23, VRS22, VP22a, VNP^{+}23, VK23b, VG21, Wan20, WGY20, WLJ20, WBZ21, WWL^{+}20, WKB^{+}22, WDW^{+}22, XL21, XDL22, XLZL22, XR21, YZYT21, YKW20, YLZC23, YCY20, YY22, ZGLS21, ZWCC23, ZFH^{+}23, ZX23a, ZZC24, ZWY^{+}24, ZZLZ21, ZZLZ22b, Zhu21, ZZS^{+}21b, ZCW^{+}23, KR22c, PST24]. optimization-aware [NBPR22]. optimization-based [AC22, AA23e, CPCK23, DCR23, DK22b, GCS23, JB24, LG23, MK22b, NSR22, NNVD22, PKK23, PVRM22, RK21b, VM23, ZGLS21, ZWY^{+}24]. optimization-enabled [JA23]. optimization-multikernal [APM^{+}21]. optimization-support [YY22]. optimizations [ZCD^{+}22, dRdSC^{+}21, dRdSZ^{+}23]. optimize [BRL^{+}20, CGW^{+}20, RP22, Yil21c]. Optimized [KMD23, KWL^{+}24, MSJ22, MG21a, PJKO24, PA23, SB23a, SRL23, VS22a, ZCR23, Agr21, ATS22, AD22b, AJK^{+}23, AJS23, Ano21c, AFM22, BS23c, BMV22, CPCK23, DÖD22, DBD23, GP22b, GK23a, GVSS22, HR22a, INY^{+}23, JAM24, JPK22, JB22b, JD23, KTM22, KR22b, KDA^{+}22, KS22d, KS22c, LKR^{+}22, LJBS23, LYC22, LM20b, LWHW22, MK22a, MPV22, Mis22, MRKY22, MCT22b, OE22a, PD22b, PSK23, PS23b, RW24, RT22b, SSMT22a, SSMT22b, SKP22, SZZ^{+}24a, ST22c, SJ23a, SSN23, SAMS23, TB23, VKSS23, VJ22, VCFZ20, VPSM22, WKY22, WYLL24, XLZL22, YZX^{+}22, YLZW23]. optimizer [AS23a, ABCP23, BJGF20, CYZ^{+}23, CZ24, Den24, DS20b, DM22d, GP22a, KAO22, LGL^{+}21, LM22b, SSMT22b, VRS22, VR22, ZCR23]. optimizer-based [ABCP23]. Optimizing [BMK^{+}20, IAA20, MHPA21, Mon21, NMT^{+}23, MR23a, PJP21, RMS24, WXH^{+}22, WK20, Yil21a, ZZL^{+}20, ZLW^{+}20, Mit20, SR22d, Yil23, dOPBdO21]. Optimum [KA22b, MNDK22, KK22e]. options [KAR^{+}23]. OQ [SS23d]. OQ-IICA [SS23d]. ORC [IP20]. Orchestrated [MK23, LLYZ23]. orchestration [CI20, PB22a]. order [AMBAJ22, Kha22a, KGGM22, LLJ^{+}22a, LSSQ22, LLY^{+}23, LLMZ21, SMCM^{+}22]. ordering [ZTM21]. Organ [ÇKÇ22, LLL^{+}24]. organisms [MNDK22, Pan23]. organization [JZB20]. organizational [SYRS^{+}22, XFPY24]. organizing [GFPGT21]. orientation [AAEA20]. orientational [BCM22]. orientational-based [BCM22]. oriented [BEKS22, CCPP21, COZ21, DLY^{+}23, HXZS23, dCJAAdOD21, LFG^{+}22, LCCT22, NN23, ORP21, OS21b, Sin23, WLCW20, WDW^{+}22]. origin [TG23]. original [KK21b]. ORNL [VBP^{+}24]. orthogonal [Ben22, RRIL22]. orthophotos [KGK22]. oscillating [BA20]. oscillations [DBK21]. Osmotic [KKS22]. OSPF [AKEC20]. osteoarthritis [BKK22]. Ottoman [DK22a]. out-of-core [HMQO23, LYI^{+}20]. outbreak [JS22b]. outcomes [GKM22]. outdoor [ZTS^{+}22]. outlier [AU22, ALNJ21, WAY^{+}21]. outliers [ADA22, MAAK23]. output [Ano21a, HTZ^{+}22, SYJL20]. Outsourced [HLC^{+}21, CT21, ZLV22]. outsourcing [Gha20]. Ovarian [BS23c, RK23b]. over-pressure [XSZ^{+}23]. overall [DBN^{+}22]. overflow [VK24, NTK22, AMR^{+}21]. overhead [GN21a, SBSK22, ZWZ^{+}22]. overheads [HK21]. overheating [PRPD21]. overlap [DJJR22, KSY24]. Overlapping [ALB^{+}20, AHW22, GlSL^{+}23]. overlays [BCP^{+}23]. overload [YLLL20]. Oversample [DRM22b]. Oversample-select-tune [DRM22b]. Oversampling [SP23a, Fio20, KFKD22]. overview [AAT21, GGA24, GQ20, NK22]. ownership [XZL^{+}20a]. oxide [ZCR23]. Özkale [ADA22].

P2P [ZMLW23]. Pack [LSL20, JYH^{+}24]. Packer [JBCI20]. packet [BKS22, SGS21b]. page [AV23, CPLX21, DPYS22]. page-level [CPLX21]. PageRank [ZLC^{+}22]. pages [JAM24]. painting [HIN23, RCK22]. pair [STH^{+}20]. paired [DZCL22]. pairwise [STH^{+}20]. Pakistan [ZARR23]. Pakistani [uZKH^{+}20]. palmprint [AEM22]. PAMPAR [GSG20]. pandemic [PKVS21, SP22b]. pandemics [BK22a, SJA^{+}22]. panoramic [VB21]. pansharpened [Yil21c]. Pansharpening [LYW^{+}21, LSZL23, Pal22, Yil22c]. Pap [PAN22]. paper [TLQ21, YOWY22]. paradigm [AYJ^{+}22, BSEN20, DAK22, FSWW21, HA21, MG21b, MKS^{+}20, OO20, PJK23, WZL^{+}24, WMC21, XSGL20]. paradigms [CdOO^{+}20, CPH20, Ogi20a]. Parallel [ASL20, BPW^{+}20, BKM^{+}21, Boc21, DBPC22, Dev21, EHST21, FZ23, FQD^{+}23, HVB22, HGM^{+}23, KÖ22, KPF^{+}20, KCL^{+}20, KSA^{+}21, LM20a, LTL^{+}20, MAG^{+}20, Oh21, RCS20, SD22, ST21, YCL^{+}22b, YSH^{+}22, AZM20, AHW22, AE22b, ABA22, AV21, AYKE21, ACVK23, Bad23, BAPS22, Ben21, BFM^{+}21, BPK24, CSL20, CHI^{+}23, CWL^{+}20, CWFK23, CMY21, CWH^{+}24, DMS^{+}21, DSS21, DR20, DS23b, DE22, DWZ^{+}20b, DD24, EK20, EASR22, FKGO22, GSG20, GMS^{+}21, GTM24, HMK24, IS23, JK22b, JSZS22, LS22, LHPG21, LK22c, LZY^{+}20, LLS22, MS21a, MTK^{+}21, MLKD20, MGB^{+}23, MT22, MCNR20, NdSSSN20, NWW^{+}22, NMS^{+}21, NGB23, Pan23, PJKO24, PBD23, PR24, PB22a, QKSK23, RTSK23, RLdO20, SEMA^{+}22, SY23, SWCB20, SLLMTS^{+}24, SPQM20, STISM21, SRGK24, SV21, SCD^{+}23, SMR21, SPHP21, SCdLV20, TNI23, TAB21, TBT^{+}21, TLS22, USI21, VGDF22]. parallel [WCL^{+}24, WZL^{+}24, WPL20, WT23, yXlLyGX21, YWQ^{+}21, YS23, YWC^{+}23, ZYX^{+}21, ZQL^{+}21, ZCD^{+}22, dSNdL^{+}23]. parallelism [CS21, HYT^{+}21, JBBH21, SPWX23, ZQX^{+}23]. Parallelization [Hua20, TO22]. parallelized [BNB22, PD20, RAK22]. Parameter [PLX^{+}24, AK22c, BMK^{+}20, HK22a, HML20, KÖÖG22, LLH19, ÖSTY22, QKSK23, TÖK21, XDL22, dOPBdO21]. parameterization [SMM22]. parameters [MHPA21, RMR^{+}22, SSMT22a, SSMT22b, SBS22, YAR22, Yal22]. parametric [CMY21, KNK22]. ParaMo [Oh21]. paraphernalia [NS23]. ParGRES [RLdO20]. parity [DMS^{+}21]. parking [ÇG21, UZAA21, ÖGS22, ZUTK23]. Parkinson [HT21, Kay22a, SSS23b]. parotid [SKÇA23]. Parquet [IP20]. part [DSJ22]. partial [NKKM21, Pan23, RCBTGE23, XGCZ23]. partially [KC22, LLL^{+}24, Shi22b]. participatory [KAP20]. Particle [KSS21, AS23a, ASL20, ASAAAA22, AB21, AKRR20, CK23, CHMC21, CPCK23, GC20, HAR20, IAAA22, KWL^{+}24, LG23, MT21, NSK^{+}24, PJP21, SA22a, SBB21, SB23b, SKSP20, WWL^{+}20, ZZY22a, ZX23a, ZZS^{+}21b]. particle-in-cell [AKRR20]. particles [SDSW21]. particulate [WHDS22]. parties [SJ23a]. partition [ZTF^{+}20]. partitioned [ALBZ21, MWS^{+}23, XZD^{+}21]. Partitioning [ST21, AD22c, ACJ21b, KFML20, LCKJ21, MDSG^{+}23, RJH^{+}23, WWF^{+}22, yZyWD^{+}21]. partitions [JOW23]. parts [YLJH22]. party [CT21, GKLS23, HZQ^{+}24, NR22, OO22, TLX22, GN21a]. pass [MKRK23, ZGWZ23]. passenger [KA21b, ZYZC23]. passing [Cud20, DSS21, SRS^{+}21, TO22]. password [WGYZ22]. patch [GAM23, KSJK21, RCK22]. patch-type [KSJK21]. patching [KJHM21]. path [BZT^{+}23a, Bul22, CM21, DWZ^{+}20b, ETKD23, GKÇ22, JYH^{+}24, KRKM22, KWL^{+}24, MY24, NQ21, PPM22, SA22a, SAM^{+}23, WL23, XZY^{+}22, XZL^{+}20b, YZYT21, YF20]. pathogen [CSWC20]. pathological [LZY^{+}21]. pathology [TNP21]. paths [LZXG22, MGB^{+}23, YYPR22]. Patient [Alm22, DR22c, DR23, HT21, dCJBP20, KKKS23, SGS^{+}23, TKG^{+}23, VSP22]. patient-centric [dCJBP20]. patients [Ano21h, CBK23, Hem22, IIK^{+}23, The21]. patrol [KWL^{+}24]. Pattern [SYJL20, BCM22, Dev21, HIEH22, MWH^{+}23, PHZ^{+}22b, SM22e, TSR22, XZYW23, ZK24]. Pattern2Vec [OA22]. patterns [CMT20, GRC^{+}23, SLL^{+}23, SBS22, WMC21]. payloads [CCPP21]. payment [XHM22]. PB [CdRNB23]. PCA [CLX^{+}21]. PCA-Flow [CLX^{+}21]. PCJ [NGB23]. PCOS [FM22]. PDEs [CdOO^{+}20]. peak [LWZY23, SCL^{+}20, YYLL22]. peaks [FYH^{+}21, ZTF^{+}20]. PECCO [WDW^{+}22]. pecker [SAMS23]. pedagogical [HKMS21]. peer [GM21, NET20, SG21]. peer-to-peer [GM21, NET20]. PEGASUS [ST21]. Pelican [PST24]. PEMFC [LM22b]. PEMFC-based [LM22b]. penalization [MGB^{+}23]. penalties [GMN21]. penetrating [TPT^{+}22]. Perceiving [RM22]. percentage [DS23a]. perception [CZTC22, JYW^{+}20, JTY^{+}21, TL21, XZG^{+}23]. perceptive [KR23]. perceptron [The21, XLL^{+}21]. perfect [HLC^{+}21, ST23b]. perform [WWW^{+}23]. Performance [AAMAA22, AKEC20, AMRH21, AIA22, AS22c, BC23b, BPM^{+}22, DPdS^{+}23, DHSG23, GGCGS20, GXH^{+}21, HK22a, HPH^{+}20, HC22, JB21, JZL22, KS20, KSY23, KLH24, LHPG21, MBC23, NGB23, Oh21, RGPC23, SA22c, SLC20, SAF^{+}23, TC22, WCCC20, Wri22, WTL23, YÇC22, ZST^{+}23, APP^{+}21, AHW22, AE22b, AY21, AML^{+}22, AAB23, AKRR20, ATC23, AY23, BR21, BPAE20, BZT^{+}23a, BANT20, BPK24, CLC^{+}20, CW22, CDP^{+}21, DPFC20, DS20a, DAW22, FJCFB23, FEK20, FFLM21, FLP20, GSG20, GDGN23, GFPGT21, GCS20, GMS^{+}21, GMSM21, GNMELC21, GLRB21, HVT23, HGX^{+}22, HZY^{+}21, HTZ^{+}22, HHP23, HML20, IP20, JSA^{+}20, JKP22, dCJAAdOD21, Kay22b, KFKD22, KFML20, KML23, KPF^{+}20, KHK^{+}24, KK22c, KDS^{+}23, KSB23, LCKJ21, LMM23, LLH19, LJZ21, LBZ^{+}22, MM24, MBM^{+}20b, Mar23a, MY24, MG21a, Mit20, MdARS^{+}23, MAAA22, NMS^{+}22]. performance [NTB23, OKJ^{+}21, OWB^{+}20, OE22b, PS21, PQKDT21, RT21, RT22a, RAK22, RZVC21, SA22b, SKB^{+}20, SRL23, SPS22, SPKK22, SSI22, SMAG22, STJ^{+}20, TMB24, TBNPQ23, TCA23, MKL21, VS22b, VS21, VP22b, VBM^{+}21, VRB21, WY20, WPK^{+}22, WT23, XSZ^{+}23, XR21, YKW20, YLT^{+}21, Yil21a, ZYH^{+}23, ZBY^{+}23, ZWX21, ZHWY22, ZCD^{+}22, ZWZ^{+}22, dCMM21, dRdSZ^{+}23]. performance-accuracy [AAB23]. performance-aware [BPK24]. performance-energy [KML23]. performances [Yil21b]. performant [RMS24]. period [DBN^{+}22, DLC^{+}21, PKVS21]. periodic [BMJ^{+}20, Ker22, RK23b, YSS^{+}21]. periodontal [VB21]. perishable [FMNF22]. Perlmutter [YKW20]. Permissioned [BANT20, RSMCP22]. permutation [CZ24]. permuted [LMR22]. persistent [CZTC22]. Person [CLT^{+}21, KWZ^{+}21, MG21b, PLX^{+}24, XLX^{+}21, XPLX23, ZCZC23]. Personal [KPM20, HDXH20, NA22a, SS23c, ZYXX23, uZKH^{+}20]. personality [YB23a]. Personalized [BKLY20, FRS^{+}23, BMSD23, JJZ^{+}21, LNC^{+}20, MSJ22, WZC^{+}22a]. perspective [CJC22, ELIG23, FD20b, KY23, YB23a, YLKC24]. Perspectives [Ano21-41]. persuasive [MSK22]. pertinent [AKLK24]. perturbation [XLY^{+}23]. perturbations [MC20]. pervasive [LMMS24]. petascale [GHL^{+}23, HPH^{+}20]. Petri [YB23b]. petroleum [GKM22]. petrophysical [BRL^{+}20]. PETSc [BBB^{+}20c]. PG [JDG22]. PG-DRL [JDG22]. PGASUS [PEGP23]. PGRD [ZLX^{+}24]. PGWO [LLAV22]. PGWO-AVS-RDA [LLAV22]. Pharmaceutical [GGS^{+}22]. phase [EAvM20, XYLW21]. PHAST [PB22a]. PHengLEI [WZL^{+}24]. phenotyping [DMD^{+}20, MRAS^{+}23]. Phi [HYT^{+}21, Mit20]. phishing [GB23, JAM24, NJK22, SS22b]. phone [JKKL21]. phone-sensitive [JKKL21]. photoelastic [TLH^{+}22]. photovoltaic [NNVD22]. phrase [ZLV22]. physical [AR23, CAAHC23, DAT23, DCT^{+}23, GKG^{+}20, HXZS23, KSJK21, KGRG24, Ker22, LL23, MBC23, NT23, SS22c, TS24b, YLJ22, ZX23a, ZALM23]. picking [LML^{+}23]. piecewise [LGDW22, LFWJ22]. pile [WJLC21]. pilgrims [AR22a]. PIN [GMP^{+}20]. PIN-based [GMP^{+}20]. pinned [CYK^{+}21]. pipeline [AFF22, DRM22b, GMN21, OO23a, TAT22, ZYH^{+}23]. pipelined [ZZL^{+}20]. pipelined-based [ZZL^{+}20]. pipelining [ZDL^{+}22]. pitch [KG23]. pixel [GRMP23]. pixel-based [GRMP23]. pixelwise [LGT^{+}23]. PixJS [TSV^{+}22]. PKD [LLX^{+}24]. PKD-Net [LLX^{+}24]. placement [ARC22, BYPO21, BZT23b, CHMC21, FCMM20, GLJ20, HAK22, KV24, KEMZ22, KHA22b, LXW^{+}23, PS23c, ST22c, SM23a, SLD24, TWG^{+}21, VG21]. plagiarism [MT22, UWF^{+}21]. Planner [KMZ^{+}20]. planning [ATG23, Bul22, DWZ^{+}20b, ETKD23, FGJ^{+}21, GKÇ22, GSZ^{+}20, HML21, HL23, JYH^{+}24, KD22, LKR^{+}22, LXT^{+}22, NGOS22, PPM22, YZYT21]. plant [DMD^{+}20, HLO^{+}21, PK22a, RKR22, YRV^{+}23, ZCX^{+}21]. plants [JA23, MCT22b, RTBC23]. plasma [DBPC22, KCL^{+}20]. plate [HWBZ21, RSKA23]. platform [AIA22, CFR^{+}21, DS24a, DCK21, DKC^{+}21, ENB^{+}20, HMB22, HA21, HL20, Hua20, JPL22, LMBK23, LFX^{+}20, LZZ^{+}23, Liq22, MBO^{+}21, RV23, SKA23, UKS22, UGK^{+}22, VCFZ20, VPB^{+}23, WZL^{+}22, WNZ24, ZHJW21]. Platforms [HSO^{+}21, ALB^{+}20, AYB21, Bad23, CCCR21, DR20, HVT23, HMR24, HTZ^{+}22, IS23, KFML20, KML23, LK22c, LCW21, LLS22, PSF23, SYRS^{+}22, SVM24, WZZ^{+}22, WC22, XAC^{+}20, YYY^{+}23, LM20a]. player [AYD21, KBJ21]. plots [KKC22]. plug [OMA^{+}23]. plug-in [OMA^{+}23]. plugin [Ano21g]. plus [AFBM^{+}23, ABC^{+}21]. PMC [JQGL20]. PMCSched [BSPM23]. PMIPv6 [MM24]. PMIPv6-based [MM24]. PN [LMMS24]. PN-HPC [LMMS24]. PoA [BK22a]. POI [SJ22]. point [DPFC20, GK23b, IAAA22, KSA^{+}21, MWS^{+}23, WK20, YSLX22, SJ22]. Point-of-Interest [SJ22]. point-of-sale [GK23b]. point-to-point [MWS^{+}23, WK20]. pointer [CZ22]. pointer-generator [CZ22]. points [AA23a, KCP^{+}22]. poisoning [ZCZ^{+}22]. Poisson [AD22a, AAK21c, IMNC22, LAP23, OU22, SAB22]. polarization [YSC^{+}20, ZWCS20]. polarizers [LSW^{+}20]. pole [MHL^{+}20]. policies [FKK23, RSMCP22]. policy [ACVK23, CPLX21, GLJ20, LGX^{+}23, PWJ^{+}21, ZHDW24]. political [BP23a, GCS23]. pollination [GMM22, MSK22, SR20a]. polling [Cud20, GXH^{+}21]. polling-based [GXH^{+}21]. pollution [BTDD20, GDA22, LZL^{+}20b]. Polly [WKB^{+}22]. PolyBench [WKB^{+}22]. polyhedral [CMY21]. polymorphic [DDH^{+}20]. Polynomial [LYBZ23, SBB21, FFLM21, HCH^{+}21, HHC^{+}22, MG22, SS21]. polynomial-time [HCH^{+}21, HHC^{+}22]. polynomials [AMBAJ22]. pooling [LLJ^{+}22a, WSM^{+}20]. Poor [DR23]. Popularity [CH21, YB23a, MSJ22]. Population [PKVS21, Cek22, HUC^{+}22, HLDT^{+}23, KML21, LFWJ22, MSA22, NSSS22, SK22c, UYÖ^{+}22, VS22c, XCD^{+}20, YAR22, ZBY22]. portability [AKRR20, CMS21, DKA^{+}21, HVT23, Mar23a, SPC^{+}21, ZBY^{+}23]. portable [CM21, IGYS23, KLH24, MWS^{+}23, RF23, TCA23]. portal [MCD^{+}23]. ported [YNK^{+}23]. portfolio [HXY^{+}22b, HXY^{+}22a]. Porting [HSO^{+}21, LBG^{+}20]. pose [HWY^{+}23, LDS^{+}23, TYL22]. positional [KML21]. Positioning [ZXL^{+}21b, KCP^{+}22, LMM^{+}22, LLKS21, LL24, WHC^{+}20, XZG^{+}23]. positive [TAT22]. possession [FXX22]. possibilities [FMNF22]. Post [HZQ^{+}24, SG24, WFY^{+}22a]. post-deduplication [WFY^{+}22a]. Post-quantum [HZQ^{+}24, SG24]. postdisaster [QHL^{+}24]. posterior [WJL^{+}20]. posts [AMR^{+}21, YSP23]. postural [ASJ24]. PoSW [XZL^{+}20a]. potential [Ano21-42, LJB22]. PoW [VH22]. powder [MCT22a]. Power [JB22b, LLL23, AM20b, ADGT22, CAS24, CYZ^{+}23, CDP^{+}21, Cza23, DPFC20, DBS^{+}22, DBK21, DTK24, FPÁ^{+}20, FDY21, GHL^{+}23, GO22, JK22a, LM22b, MCL^{+}20, MSS^{+}20, NBK22, NNVD22, NKST24, RAN21, SZZ^{+}22, SSPG20, TTA20, WTL23, WT23, YLLL20, Yüc22]. power-aware [Cza23, FDY21]. power-controlled [MSS^{+}20]. Power-efficient [JB22b]. powered [CD22, CLYG22, DCT^{+}23, MV22, MG22, VH22]. Practical [ZYXX23, Ano21d]. Practice [Ano21-41, XZ20b, LJZ21]. practices [CY22, DJF21, KPP^{+}22]. Pradesh [DM22d]. pragmas [WKB^{+}22]. Pre [ACJ21b, Yüc22, KE22, PLP22, ZFF^{+}21]. Pre-determination [Yüc22]. Pre-filtering [ACJ21b]. pre-processing [PLP22]. pre-trained [KE22, ZFF^{+}21]. precautions [DÖD22]. Precise [SCKU23]. precision [GCF^{+}20, LP21, SM22a, VDL23, YCL^{+}22a]. precisions [MW21]. preconditioning [HYG20]. predator [AO23, SSN22]. predators [BDK22]. Predatory [LK23]. Predicate [DWZ^{+}22]. predication [YTN^{+}20]. predict [BBD23, KUK22, KPJ^{+}21, KSB23, PP20b, VR22]. predictability [ZYXX23]. Predictable [BWS^{+}21]. Predicting [AFBM^{+}23, BB23, GLW21, HR22b, KG22, UMR23, XZY^{+}22, ZHW^{+}20, Asl22, BM22c, GKM22, JSS22, Kan22, SHA^{+}22, WHDS22, YW22]. Prediction [AKA22c, Alm22, BMcKGK22, CGS^{+}21, DÖD22, DM23, DM21, JS22b, LLMX21, VB24, AJS23, Anb22, AV22b, AR22b, AGM23, AHT^{+}20, AYJ^{+}22, BSN22, BLB^{+}24, BF22, BJK^{+}24, CNG^{+}20, CPQ^{+}22, CHL^{+}24, CWW23a, CT22, CB22, CA23, DR22c, DS23a, DBN^{+}22, DM22b, DT22a, DSSS22, DTK24, FM20, FRS^{+}23, GlSL^{+}23, GHRM21, GCP22, GRG^{+}22, GKAO20, HOS^{+}21, HR22a, HGHD22, HLG24, JS22a, JM22, JZL22, JHZ20, JR23b, KSS21, KTM22, KK21b, Kha22a, KMD23, KK22d, KRSR23, KR22b, KBS^{+}22, LJ22, LSD21, LM22a, LSZ24, LLJ^{+}22b, LAC21, LM24, MBB22, MCT22a, MBC23, MRAM^{+}21, MSL22, MSBR23, NTB23, OCD22, OWB^{+}20, PBD23, PJK23, PRPD21, PVRM22, PM22, RM21, RM23, RK23a, RT21, RT22a, RK23b, RRIL22, RV23, RZ21, SRL23, SPSP23, SKCS23, SZZ^{+}22, SZGR21, SM22c, SM23b, SVB23, SS22d]. prediction [SSP24, SHVA23, SM22e, SK21c, SRS23, DM22a, SM23d, TT22, VSP22, WJLC21, WHH^{+}23, WWL^{+}20, WSL^{+}20, WCZ^{+}23, WTL23, XCY22, XYY24, XLL^{+}21, XY21, XLZL22, YTW^{+}24, YYPR22, YLZC23, YLJH22, YF20, ZNDA22, ZYXX23, dRdSZ^{+}23]. prediction-based [CHL^{+}24]. predictions [XDH^{+}20]. Predictive [GMK^{+}21, BK23, HML20, JD23, LLH19, ÖK22, RAaB21, SAAAA22]. predictor [RAK22]. predictors [SAF^{+}23]. PreDiKT [uZKH^{+}20]. PreDiKT-OnOff [uZKH^{+}20]. preempt [MS21b]. preemption [BMJ^{+}20]. preemptive [RCBTGE23]. PreF [XZY^{+}22]. preference [GK22b, JNM^{+}22, KGGM22, MBB22, SLJ23, Wan24]. preferences [TLQ21]. prefetch [HK21]. prefetcher [GMS^{+}21]. prefiltering [WWF^{+}22]. prefix [NWW^{+}22, SR22d]. prefix-tree [NWW^{+}22]. Pregnancy [DT22a]. Preliminary [HYT^{+}21, CCD^{+}20b]. premises [MDSG^{+}23]. premixed [XSZ^{+}23]. preparation [MLC^{+}21]. preprocessing [EUYY22, QMC^{+}20, ZCZW22]. presence [ADA22, BDG^{+}23, CSL20, Faz23, HTVL22, PP20b, VSK22]. presentation [YLKC24]. preservation [ABSS22, DK22b, HJZ^{+}22, JB24, KHEF22, LmJdL^{+}22, MLZ^{+}20, SYL23, Wan22b, ZSZ^{+}22]. preserving [AM22, AR23, BXH^{+}23, CSV22, DP22b, FC21, HKA20, HAAF22, KM23a, LNC^{+}20, LZL^{+}23, MLS^{+}24, MJAM23, NZD^{+}24, PR23, PDRM24, PA21, RAG21, SSW^{+}22, SCL^{+}20, TWL^{+}20, WZMJ23b, XDJY21, YZZA23, ZYZC23, ZS22]. pressure [FSFM22, TXT24, XSZ^{+}23, YTW^{+}24]. pretopology [HBB20]. pretrained [JNS22]. prevent [VK24, LLM^{+}22, NTK22, NM20]. preventing [VD21]. prevention [CCM22a, GB20, KHK^{+}23, LLC^{+}22b]. Price [LLA^{+}22, HR22a, KRSR23, KSJ23b, NTB23, ZNDA22]. prices [KAR^{+}23]. pricing [BC21, KRSR23, NM23, ZNDA22]. prime [AE22a]. primitive [BXH^{+}23, SG21]. primitives [MWS^{+}23]. Principal [AAK^{+}21a, wGTC22, AU22, HNG22, HML20, KÖÖG22, LLH19, YYLL22]. principle [NGD^{+}22, RR23b]. principles [LBFT22]. prior [LLX^{+}24, WLLX21, WZSZ20]. prioritized [AY21, SKR^{+}24]. prioritizing [AZI20]. Priority [LGLZ20, BZWH21, BKS22, FSA^{+}22, GD22, NM23, RCK22]. Priority-based [LGLZ20, BKS22, FSA^{+}22, NM23]. PriTxt [XLZZ22]. Privacy [BR21, CT21, HJZ^{+}22, HAAF22, KS22b, MLS^{+}24, PR23, SSW^{+}22, YZZA23, ZS22, AM22, AR23, ABSS22, BXH^{+}23, CSV22, CYZX23, CZCM23, CYW^{+}22, DAT23, DPB22b, DP22b, DK22b, DLY^{+}23, DS24b, FC21, GZY^{+}22, HKA20, HNS^{+}21, HDXH20, HLC^{+}21, HZZ^{+}23, JB24, KM23a, LXYY21, LNC^{+}20, LmJdL^{+}22, LWW23, LZL^{+}23, MLZ^{+}20, MJAM23, MAB22, MGN^{+}22, NZD^{+}24, PGL^{+}23, PZZ^{+}23, PDRM24, PA21, PKS22, QZB^{+}23, RAG21, SZW^{+}22, SYL23, SCH^{+}24, SCL^{+}20, TWL^{+}20, TZ23, VRS22, Wan22b, WGLL23, WZMJ23b, XZD^{+}21, XDH^{+}20, XLZZ22, XZW^{+}20, XDJY21, YLZT23, YMZD21, YZXL22, ZZLZ22a, ZSZ^{+}22, ZML^{+}23, ZYZC23, ZLZ^{+}22b, ZLV22]. Privacy-aware [BR21, DAT23, DS24b]. privacy-edge [DPB22b]. privacy-enhanced [XZW^{+}20]. privacy-intrusive [QZB^{+}23]. Privacy-preserving [MLS^{+}24, PR23, SSW^{+}22, YZZA23, ZS22, AM22, AR23, BXH^{+}23, DP22b, HKA20, KM23a, LNC^{+}20, LZL^{+}23, MJAM23, NZD^{+}24, RAG21, SCL^{+}20, TWL^{+}20, WZMJ23b, XDJY21]. privacy-protection [ZML^{+}23]. Private [ZLZ^{+}22b, HDS^{+}23, HAA23, SZY^{+}22]. Private-encoder [ZLZ^{+}22b]. Proactive [AFNH21, CWFK23]. probabilistic [AP22, BSBF22, CCGN20, RM21, ZFZ^{+}20]. probabilities [STH^{+}20]. Probability [SNET21, ZFW23, AE22b, AAEA20, FSFM22, RMR^{+}22, SK22c, WJL^{+}20, XWW^{+}21]. problem [ABESEh20, AAMM23, APP^{+}21, AE22b, ASMS21, BKD22, CZ24, CZ21, DCR23, DD24, Fio20, FMNF22, GMN21, GADM20, HNG22, INY^{+}23, JYL^{+}23, KK21a, KÖ22, KWL^{+}24, LHJ22, Le23, LGL^{+}22, LZQ^{+}22, Shu22, USI21, WGDP23, XA22, XGCZ23, YLZC23, YBJ^{+}23, Yil22b, ZTM21, Zhu21, OO23b]. problems [AO23, ÇGB23, CMA^{+}21, FZZZ23, GFA21, KK24, LNI^{+}24, LZC21, MM21, SAPC21, SSK23b, ZSX21, dOdMC^{+}20]. Procedures [FZC20, Ogi21]. Process [AKZA22, ASMK21, CGS^{+}21, CSR24, CC22b, DCK21, ER22, ELIG23, wGTC22, GK23a, HD23, LS23a, LFG^{+}22, LZC21, LPC^{+}21, MNR^{+}22, OO23a, PMS^{+}21, PA23, dQtZWS22, RAN21, RG23, SAM^{+}23, Shi22a, WXLD21, XFPY24, XDL22, YLZ20, ZLYS21, ZLX^{+}24]. process-based [MNR^{+}22]. processes [Ano21e, BBB^{+}20a, ESB20, GK23c, IMNC22, KMS^{+}21, KYZ20, KC22, YLJH22, ZST^{+}23]. Processing [SM24b, AB20a, ACJ21a, AS22c, AAG^{+}22, ABA22, ACJ21b, BM22a, BFM^{+}21, CZUPA23, CY24, DBS^{+}22, DDB^{+}21, DAK22, ENB^{+}20, FJCFB23, GLRB21, HCG21, HIN23, JPL22, JZL21, JS23, Kim21b, LK22b, LZY^{+}20, LLC^{+}22b, MKRK23, MBM^{+}20b, MLKD20, MT22, NdMP22, OSK23, Ogi20b, PLP22, PRS23, QPS20, RLdO20, RT22c, SKH^{+}21, SPSP23, SPC^{+}21, SMD^{+}21, SAF^{+}23, SRG^{+}21, TBT^{+}21, TNP21, USI21, VB21, VGDF22, WWM^{+}23, WCCC20, WGDP23, WZB21, yXlLyGX21, ZLW^{+}23, ZWC^{+}22, RVJ^{+}22]. Processing-in-memory [SM24b]. processor [KMB23, RS20, YS23]. processors [AAG^{+}22, AAQOT23, AV21, LLTZ24, EGGG23, USP^{+}23]. produced [Aka22b]. product [AAG^{+}22, AB20b, BBF21, BK22a, CYZ^{+}21, FMNF22, HAA^{+}21, JR22, Li20, MRKY22, Ona21]. production [CJY^{+}20, CZ21, LXKW23, MSPPD20, OWB^{+}20, WZC^{+}22b, XDL22, YLJH22]. productivity [HGM^{+}23]. productivity-aware [HGM^{+}23]. products [AAQOT23, Ano21f, VVK23]. proficient [SPKK22]. Profile [CdRNB23]. Profile-based [CdRNB23]. profiles [KBK^{+}22, KML23, RVJ^{+}22]. Profiling [GRC^{+}23, BWTJ20, FD22, MS21b, MKBB22, RVJ^{+}22]. Profit [ACC^{+}20, WDW^{+}22]. Profit-aware [ACC^{+}20]. prognostic [HTVL22]. program [BP20, CZZ^{+}22, Kim21b]. programmable [GAMT22, TAB21, XCJ22]. Programming [Dev21, FFG^{+}23, HSO^{+}21, Oh21, SCH22b, SCK^{+}23, Ano21-36, CYC21, CMY21, Cza23, FKGO22, GTM24, GK22b, HXY^{+}22a, JPL22, KD22, LSS24, MGS^{+}20, Mon21, SCH22a, WCL^{+}24, WZZ^{+}22]. programs [BJWY20, DSS21, FD22, GB20, SPWX23]. progression [SP23a]. Progressive [EBDB22, LLZ^{+}21b, Çet23, CCGN20, SZS20, VKSS23, ZZY^{+}22b]. progressive-enabled [ZZY^{+}22b]. project [BBB^{+}20b, CFR^{+}21, USI21]. projected [CB22]. projection [RT22a, SP21b, ZZL^{+}22]. projects [DJF21]. ProM [Ano21g]. prominent [AIA22]. promise [KE22]. promote [KLJ21]. promotion [THW21, YM22]. pronunciation [JKKL21]. Proof [JB21, XHZ^{+}21, XZL^{+}20a, LK23]. Proof-of-Stake [LK23]. propagation [CS23, DZLH20, KRJS22, KDFK23, LSQW21, PBD^{+}21, WCWG21]. Properties [IAT^{+}23, SGHL20]. Property [VCBB20]. Property-based [VCBB20]. proportional [SK22c]. Proposed [HA22, TSD23, GRMP23, Yil22a]. Proposing [RVAE21]. prosthetics [BZGM22]. protected [KS22b]. protecting [HDXH20, KKP24, SSSP21]. Protection [PCG^{+}21, CYZX23, CWW^{+}23b, CYW^{+}22, DLY^{+}23, GZY^{+}22, HNS^{+}21, JQ22, KM23b, KM23c, LXZ22, LWW23, SA22d, WGLL23, WZ21, XZYW23, ZML^{+}23, ZHXY23]. protein [CSWC20, KT20, PBD23]. protein-protein [CSWC20]. proteomics [Ano21j]. Protocol [KSCL21, AMAT22, AP23, AB20b, Ano21d, BZT^{+}23a, BM22b, BDG^{+}23, CGW^{+}20, GLN23, GPR^{+}22, GN21a, GXH^{+}21, HA22, IA22, JP21, JCG^{+}22, KGRG24, KGGM22, LCSR21, LWW23, DPD^{+}22, MM24, MKA23, MPG20, MG23, NM20, PSP24, PB22b, PM23b, QWW^{+}22, RJ22, RKMS24, RK23c, RT22b, SNET21, SK22a, SPS23, SL20, SS22e, SR20a, SPK^{+}22a, SSH22, TKG^{+}23, VDL23, WLL21a, WGYZ22, WYLL24, XLZ20, XHST20, ZWO^{+}20, ZLO^{+}21]. protocol-urban [MM24]. protocols [GDCW22, GQ20, HZQ^{+}24, KJMB22, LGA^{+}20, LZZ^{+}20, Ngu21, OT20, SK22b]. prototype [Ano21-43]. Provably [TJC24, VDL23, NAR^{+}22, PSP24]. ProvBench [LBZ^{+}22]. provenance [LBZ^{+}22, NM24, SAL^{+}22b, TA23b, YA22a]. provider [FCZ20]. providers [ACC^{+}20, MA22, YJY^{+}21, dCMA23]. Providing [BPT^{+}23, TCA23, MPLW24, SK23]. provisioning [MMR21, OBTC20, WWL^{+}20, XCG^{+}22]. proxy [BVS20, PSB^{+}20, SRS^{+}21, TMB24]. pruned [HUJ22]. Pruning [LLC^{+}22a, OHFF20, JKS20, LGX^{+}23, WSZ^{+}23]. pseudo [BA20, HDXH20]. pseudo-location [HDXH20]. PSO [KSS21, SA22a, YLJH22, LLL23, NA22b]. PSO-FCM [KSS21]. PSO-GRU [LLL23]. PSONN [Yüc22]. psychological [RVJ^{+}22, THW21]. psychology [HGMK21, PVRM22]. Psychometric [RVJ^{+}22]. psychotherapy [XZ22]. PTNet [HZD^{+}22]. pub [KHHK21]. pub/sub [KHHK21]. public [AR23, BV22b, CPYC21, FXX22, GFQ20, HL20, MC24, SJ23a, UKS22, YSC^{+}20, YZXL22, ZWO^{+}20]. publication [LXYY21]. Publicly [KYP21, ZLO^{+}21]. publish [NET20, XSGL20, YPO21, ZQX^{+}23]. publish/subscribe [NET20, XSGL20, YPO21, ZQX^{+}23]. publishing [CSV22, DK22b, LX22, MLS^{+}24, RN22, SCH^{+}24, TZ23]. Pulmonary [ZKZ^{+}23, KPJ^{+}21, ZZW^{+}22]. Pulse [TB23]. punishment [GYL^{+}21, Ano21b]. purchase [THW21]. pure [ZST^{+}23]. purposes [DB23]. pursuit [RT22a, URK^{+}22]. pushdown [YSH^{+}22]. pyramid [LCL^{+}20, ZKZ^{+}23]. pyramidal [PP21].

Q [FSFM22, SB23a, SM22c]. Q-learning [SB23a]. QCA [AM20b]. QoE [LBA23, SG22d]. QoE-aware [SG22d, LBA23]. QoS [AMRH21, AASPR22, BJGF20, DJGF21, FM20, GCP22, HGHD22, LmJdL^{+}22, LXW^{+}23, Mir22, PS23c, RG23, SK23, SS23d, SH22, SR20a, SSH22, XHZ^{+}21, YSL^{+}21a, YJJ23]. QoS-aware [AASPR22, DJGF21, Mir22, BJGF20, LXW^{+}23, SS23d, SR20a]. QoS-based [RG23]. QRDF [JZL21]. quadraplet [LMR22]. quadratic [APP^{+}21, HXY^{+}22a, KIN^{+}23, LNI^{+}24, RKK23]. Qualitative [WZX^{+}22, SGS21a]. Quality [Ben22, SKRS21, BKS22, CPQ^{+}22, CHL^{+}24, CD22, DG21, EABZB21, GHRM21, HR22b, HIN23, KLK23, LC20, LLN^{+}23, MSPPD20, MRS^{+}21, NNVD22, OSK23, PPA22, RP21b, RKL21, SAQJ23, SK22a, WJD22, YFF22, YT21, Yil21c, YLJH22, ZQW^{+}21, ZWZ^{+}21]. quantifying [SGHL20]. quantitative [LLT21]. Quantity [JZC^{+}23]. quantization [JLZ^{+}24, KDL20, TA23a]. quantization-based [KDL20]. Quantum [GN21b, RKMS24, YLJH22, AML^{+}22, AG22, DZCL22, DAK22, FZT22, GM22a, HZQ^{+}24, JMV24, Jia22, KA22a, OBER22, QNZP22, RC22, SN22a, SG24, UAS^{+}22, VK23a, Yib22a, ZZLZ22b, LSL20]. Quantum-behaved [YLJH22, LSL20]. quantum-dot [AG22]. quantum-dots [GM22a]. quantum-inspired [DZCL22]. Quasi [JB24]. Quasi-identifier [JB24]. Queens [PJKO24]. queries [APM22, Akb22, GGCGS20, LHK22, LZL^{+}23, XZL^{+}20b, XDJY21, Yil21a, ZHXY23]. Query [HYG^{+}23, KC20, AQP^{+}22, CY24, CWL^{+}20, GSS^{+}23a, GZY^{+}22, JZL21, KS22d, PM22, RV21, RAG21, RLdO20, SKSP23, WQY^{+}22, ZYZC23, ZLW^{+}23, ZWC^{+}22]. querying [KRK21]. Question [Che24, YSK22, ZPL21, AWS^{+}22, FDH22, KC20, ZFF^{+}21, ZLM22]. question-answering [KC20]. questions [Che24]. queue [DKP^{+}20, FSA^{+}22]. queueing [FSFM22]. queues [DGSB20]. quicksort [LS22]. quintic [Bul22]. Quo [MTSU22]. Quorum [Rav23].

R [Faz23, SKP22, SKP23]. R-CNN [SKP23]. R-ICTS [SKP22]. RabbitMQ [KSCL21]. Rabin [SM24a]. Rachford [FZZZ23]. radar [GDSS22, MRS^{+}21, SYJL20, SPHP21, XCX^{+}20]. radial [DT22a, Pan23, VJ22, ZCR23]. radiation [KNK22, SEC22]. Radio [IVP^{+}23, AMRH21, AKS^{+}22, Ben21, Ben22, CTA^{+}23, CÇM22b, DRMA22, LLW^{+}20, MSS^{+}20, Sao22, SKPV22, SS22f, TSB23, ZX23a]. radiograph [UAS^{+}22, VB21]. radiographic [FMJ^{+}22]. radiographs [Ele22]. radiomic [DBN^{+}22]. radix [TZ23, yXlLyGX21, MRAS^{+}23]. Raft [DQF^{+}23]. rail [DZLH20, TZZ^{+}23]. railway [TZZ^{+}23]. rain [KTM22, SHVA23]. Rainbow [MPG20]. Rainfall [RP21c, BBD23, CCD^{+}20b, DM22b]. RALB [AAK^{+}21b]. RALB-HC [AAK^{+}21b]. Raman [SA22b]. Ramanujan [VVK23]. RAMs [WMY^{+}21]. Random [AKA22c, SJ23b, AMJK21, Aka22b, ATS22, BA20, BDK22, BKK23, CAS24, JSAA22, MCT22a, MHB^{+}22b, OE22b, SV22a, SKB^{+}20, SCM22, SP21b, TT22, WLX21, YLGY20, ZBY22, TI22]. random-grid [OE22b]. randomized [HTZ^{+}22, LYI^{+}20]. randomly [AE22a]. range [LNC^{+}20, SWK22]. range-sensitive [LNC^{+}20]. Rank [CT22, BV22a, HMQO23, LLMZ21, SS23c, SP22d, ZXLD21]. rank-level [SS23c]. rank-revealing [HMQO23]. ranked [BK22b, BK23, HHYL22, MS21c, ÖUG22, RMR^{+}22, VS22c]. Ranking [XW24, BV22a, BABLH21, HKA20, HNG22, JY20, KGM23, LPC^{+}21, MAB22, PVP^{+}20, SBSK22, SP22d, ZNLL22]. Ranking-based [XW24, SP22d]. Ransomware [BLI20, KHK^{+}23, NBPR22, AKZA22, NRP^{+}20, SKK21]. RAPSAMS [KRJS22]. rat [SS23f, VR22]. rate [CLL^{+}21, KK21b, Man21, MNYN21, NKST24, WEH^{+}22]. ratio [CPCK23, GLA^{+}22, KG23, WWW24, ZBY22]. ratio-type [ZBY22]. rationality [KVP21]. Raven [SSN23]. ray [Boz22b, GP22b, KPP^{+}22, LK22a, LLM^{+}24, SSN21, WSM^{+}24, YS22, ZZL^{+}22]. ray-tracing [SSN21]. rays [KCL^{+}20, MRGP22]. raytracing [TWH^{+}22]. RBF [AAARR20, AZA20]. RBPSO [WGY20]. RC4 [MS21a]. RCC [CLWX21]. RCC-8 [CLWX21]. RCU [SNGK21]. RCU-HTM [SNGK21]. rCUDA [RS21]. RDA [LLAV22]. RDF [AMD20, JZL21, yZyWD^{+}21]. RDWT [SKS^{+}23]. re [BVS20, CLT^{+}21, LMGG20, PSB^{+}20, PLX^{+}24, PVP^{+}20, XLX^{+}21, XPLX23, ZCZC23]. re-encryption [BVS20]. re-identification [CLT^{+}21, PLX^{+}24, XLX^{+}21, XPLX23, ZCZC23]. re-negotiation [LMGG20]. re-ranking [PVP^{+}20]. re-signcryption [PSB^{+}20]. reaction [BSEN20]. Reactive [CYZ^{+}23, CWFK23]. reactors [KCL^{+}20]. reader [CSL21]. reading [YSK22]. Real [ALBZ21, BBP^{+}24, CJ21b, CLX^{+}21, EASN22, KA23, KMZ^{+}20, PD20, PMR^{+}21, RZ24, TTZX22, YTW^{+}24, ARS22a, BEJD22, BWS^{+}21, CY22, ÇG21, GMS^{+}21, GMK^{+}21, HWG23, HGDD20, IAO21, JK22a, JAC^{+}21, LZL^{+}23, LLA^{+}22, RKuH^{+}20, SKA23, SKP22, SMD^{+}21, Shi22b, SK20b, SK21b, SYT^{+}21, SZZ^{+}24b, TS21, TYLY23, VP22a, VKK^{+}24, WMC^{+}23, WSJ^{+}21, XLL^{+}23, XYFZ23, YÇC22, ZHT^{+}23]. Real-Time [KMZ^{+}20, ALBZ21, BBP^{+}24, CJ21b, CLX^{+}21, EASN22, KA23, PD20, RZ24, TTZX22, YTW^{+}24, ARS22a, BEJD22, BWS^{+}21, CY22, GMK^{+}21, HWG23, HGDD20, IAO21, JK22a, JAC^{+}21, LZL^{+}23, RKuH^{+}20, SKA23, SKP22, SMD^{+}21, Shi22b, SK20b, SK21b, SZZ^{+}24b, TYLY23, WMC^{+}23, WSJ^{+}21, XLL^{+}23, XYFZ23, YÇC22, ZHT^{+}23]. real-world [SYT^{+}21]. reality [JHD^{+}21, KJS^{+}20, YLKC24]. Realizing [ABC^{+}21]. Reasoning [HV21, ARS22b, DWZ^{+}22, HH23, USK23, HYG^{+}23, RS23, RF21, YL20a, uHAU23]. Recalling [SAMS23]. Recalling-enhanced [SAMS23]. receding [DZCL22]. received [LLKS21]. receiver [LCZ^{+}20b]. Receptive [ZLZ^{+}22a]. Recharge [MCR^{+}23]. rechargeable [CSWZ22]. recharging [KAR^{+}23]. recirculation [FDY21]. Recognition [SV22e, ZLZ^{+}22a, AR22a, AEM22, AFM22, BKJ22, BKK23, BYL20, BP23b, BYTG22, BMV22, CCGN20, CLY^{+}21, CZTC22, DSC^{+}21b, FAM22, GZC^{+}22, GA22, JB24, JB22a, JKB22, JWL20, JYC^{+}21, JTY^{+}21, JPO^{+}21, KE21, KDC22, Kay22b, KSK^{+}22, KSK23, LN20, Li21, LLW^{+}21, LZY^{+}21, MSA21, NBPR22, PT22, PZD^{+}21, RR23a, RSKA23, RMS22, RD23, SS21, SHL^{+}22, SMD22a, ST23a, SMBM24, SM22b, SZqWZ20, SPKK22, TWXL21, TUD21, VPSM22, YKL23, Yl20b, YLWL20, YSL^{+}21b, ZLW^{+}21, ZLD22, ZYCS22, ZGG^{+}23]. Recognize [SKP22, AAK^{+}21a, PB21]. Recognizing [JHD^{+}21, SKP23]. Recommendation [LSS^{+}21, AV23, AJ21, BKLY20, CPQ^{+}22, CHL^{+}24, CBFS23, DR22b, ER22, FM20, GCP22, HDXH20, JY20, JKKL21, LmJdL^{+}22, LSLY20, MSJ22, MLZ^{+}20, NNJC23, PMS^{+}21, PM22, dQtZWS22, SJ22, SX21, SM22c, SM23c, SYRP22, SRIB23, SP22d, TLQ21, UPGCA22, VS22b, WWG^{+}20, WWZ^{+}22, WZZ^{+}22, WQS^{+}23, Wan24, YB23a, YGZ^{+}21, ZLYS21, Zhu22a]. recommendations [Ano21h, EVVR21, GK23b, Kum21, MWH^{+}23, RKuH^{+}20, TC24, WWW^{+}20]. Recommender [VCS^{+}21, CYQ^{+}20, CYW^{+}22, DS21, GK22b, HDS^{+}23, JNM^{+}22, Kat23, OBER22, dQtZWS22, RSJ21, VAB22]. Recommender-as-a-service [VCS^{+}21]. Reconfigurable [FHH^{+}20, HGX^{+}22, CC22b, OCR^{+}23, TBNPQ23]. reconfiguration [HK21, RCBTGE23, VG20]. Reconstruction [ZFZ^{+}20, TLH^{+}22, ZX21a, ZX23b]. record [PK22c, SC22b, SJ23a, YLM21]. records [Ano21-43, DR22c, KRK21, NA22a, SHZY20]. recoverable [XHST20]. recovering [KR22c]. recovery [LZZ^{+}20, NKKM21, RS23, SP21a]. rectal [WMS^{+}23]. rectangular [HLS^{+}20]. rectified [OKJ^{+}21]. recurrence [GQ24, KKC22]. recurrent [AC22, ASJ24, BZGM22, CA23, FRS^{+}23, HGNN22, HUJ22, KKKS23, Kha22a, MS23, PJK23, SRG^{+}22, SP23b, SAMS23, SBS22, YRV^{+}23, YF20, ZCH^{+}23]. recursive [SSMT22a]. recustomize [SPSP23]. recyclable [LLYZ23]. recycling [CYZ^{+}21]. Redefining [RSM23]. Redesign [CJY^{+}20]. Redesigning [RS21]. redistribution [RS20]. Reduce [AV22a, Dev21, VS22a, ADGT22, GNMELC21, GHT20, KK23a, PYC^{+}20]. reduced [GN21a, SPSP23, VPSM22]. ReducePartition [GHT20]. Reducing [AV21, KGM23, LYI^{+}20, FSFM22, HK21, MHL^{+}20, MTT20, WL23]. reduction [AOACAQ21, CEN22, HC22, JM22, LXYY21, LFWJ22, EGGG23, MAW^{+}22, PL21, RSJ22]. redundancy [WBZ21, ZTY^{+}22]. redundant [AÜ23, KSP21, WLG^{+}22]. Reed [QNZP22]. Refactoring [JBBH21, KT23]. reference [PKB^{+}23]. refinement [SWCB20]. reflecting [NKST24]. region [BTT21, JM23, SMKA22, XCX^{+}20, Ano21f]. region-based [SMKA22]. regional [GHRM21, WN21]. Regions [BKM^{+}21, AZI20, PMR^{+}21, SM23c]. Regions-of-Interest [BKM^{+}21]. register [HZW^{+}23]. registration [PSMM22]. regression [AD22a, AAA^{+}22a, AA22a, AAK21c, AV22b, AU22, ADA22, BC22, BC23b, KV22a, KRSR23, LHC^{+}23, LGDW22, LAKA21, LAP23, MAAK23, MAAA22, OU22, OAA22, SAB22, SHA^{+}22, SU23, SM21, The21, YK23, YJY^{+}20, ZLL^{+}22, WSL^{+}20]. regression-based [BC23b]. regressive [CT22, CB22]. regressors [OAA22]. regularization [LLT21]. Regularized [LLS22, PK22b]. regulation [JJZ^{+}21]. regulatory [EHST21, LTL^{+}20, PBC^{+}23]. regulizer [MQEK21]. rehabilitation [YHL^{+}21]. reinforced [Ele22]. Reinforcement [KS23b, Yil22b, AA23b, DR21, DXXL20, EBDB22, ETKD23, JDLP22, KS22a, KHHK21, KMRR20, LZQ^{+}22, NSPdO21, ÖGS22, RBWB21, SPS22, WZZ^{+}20, YLW^{+}22, YHL^{+}23]. related [ARS22b, KIW^{+}22, KHK^{+}24, LK22b]. relatedness [DPB22a]. relation [BWW^{+}20, HLH^{+}20, YSLX22]. relationship [KSJ23b, ZWCS20]. relative [HXY^{+}22b, WER^{+}23]. relay [GAS23]. relay-based [GAS23]. relevance [MS22a]. reliability [BKM^{+}24, CK23, FPHZ20, HML19, MSK22, WBL22, WFY22b, ZWZ^{+}22]. reliability-based [MSK22]. Reliable [HTVL22, PP20a, AEMS24, BZT^{+}23a, SL20, YYPR22, ZCN22]. relief [MQEK21, Zhu21]. relief-based [MQEK21]. ReliefF [SSDY20, WRE^{+}23]. relocatable [FKO23]. relocation [MR23b]. reLU [KS22c]. reLU-BiLSTM [KS22c]. reluctance [MHL^{+}20]. remanufacturing [CYZ^{+}21]. remote [GDA22, GSS23b, HT21, KK23b, LYW^{+}21, PRS23, SHA^{+}22, SGH23, YM22]. remote-sensing [LYW^{+}21]. Removal [MH23, LZF20, MRS^{+}21, ZXLD21]. removals [Çet23]. removing [PP21]. rendering [Shi22b]. renewable [Faz23]. REP [PSP^{+}20]. Repair [PSP^{+}20]. repairability [ZMLW23]. repeatable [WZHL21]. repellent [ZZS^{+}21b]. replacement [CON23, NR23]. replica [CON23, FCMM20, FXX22, GLJ20, LLJR21, TWG^{+}21, TY22]. replicated [CON23, ZJ21]. replication [BC23a, BZT23b, CH21, HMK23, RAaB21]. repository [BSML21]. Representation [OA22, AA22b, CPQ^{+}22, CHL^{+}24, LM21, LCL^{+}20, PHZ^{+}22b, PHZ^{+}22a, RF23, WLLZ20, ZZY22a]. representations [SSP23]. Representative [RR23a]. Representative-discriminative [RR23a]. represented [HL23]. repression [TAT22]. reproducible [MPLW24]. reproduction [ZHJ20, ZHJW21]. reprogramming [SK22b]. repudiation [WLL21a]. repulsion [Fio20]. reputation [BK22a, GHRM21, MA22, OO22]. reputation-based [GHRM21]. request [CWW23a]. requirement [CJ21b]. requirements [BCC^{+}21, MM22, PHZ^{+}22a]. reroute [PSP^{+}20]. resampling [NR23]. rescheduling [JC21b]. rescue [DDUK23, QHL^{+}24]. Research [BPAE20, Bei24, CCZ^{+}21, CLL^{+}21, GCSZ24, GU23, HZW^{+}23, HSL^{+}22, JY20, LXZ22, Liq22, LLLX20, LL24, SX21, wSYyCsD23, TDZ^{+}24, TYL22, WZ20a, WQS^{+}23, Wan24, WZ21, WN21, WN23, Xia20, XJW^{+}22, XHM22, YLLL20, Yib22a, ZCZW22, yZyWD^{+}21, gZWfL^{+}20, ZXL^{+}21b, Zhu22a, ABC^{+}21, ACC^{+}23, CWDM^{+}21, DMD^{+}20, DKK24, LBZ^{+}22, NWX^{+}23, THT20, YL20c]. resemblance [WFY^{+}22a]. Reservation [SKPV22]. reserving [YLJ22]. residential [SPA^{+}21]. Residual [ZCX^{+}21, AA23e, CD24b, CW22, HLZ23, MPV22, PAN22, RKR22, RMMK23, YW22, ZCH^{+}23]. resilience [SP22b]. resiliency [ANP^{+}20]. resist [CYZX23]. resistant [XHST20]. ResNet [JSAA22, SKÇA23]. ResNet-18 [SKÇA23]. ResNeXT [XLL^{+}20]. resolution [BTDD20, BYL20, LAE^{+}22, LGDW22, LHPG21, LL23, SYT^{+}21, WMC^{+}23, WWJ^{+}20, ZYX^{+}21]. resolutioned [PR22]. Resolving [Gha20]. resonance [IK22, Kot23, ODK^{+}23, SKÇA23, TNP21]. Resource [SKH^{+}21, AS22a, AAK^{+}21b, ALBZ21, AKS^{+}22, APV23, BC21, BP24, BWTJ20, BPK24, CWW23a, DR21, DPdS^{+}23, ESB20, GAS23, GSG^{+}23, HTZT23, JW21, JC21b, dCJAAdOD21, Kab23, KS22a, KKS22, KHR22, KLK23, KDS^{+}23, LLW^{+}23, LCCT22, MR23b, MMR21, Mir22, MV24, NA22b, NKST24, NJ21, OXBL23, PJP21, SKB^{+}20, SKPV22, SB23b, SSK^{+}23a, SM22d, TKS22, UGK^{+}22, USI21, VB22b, WLCW20, WGY20, WWL^{+}20, XCG^{+}22, ZHDW24, ZFW23, dSNdL^{+}23]. resource-aware [AAK^{+}21b]. resource-constrained [USI21]. resources [ARC22, CLLB20, CMS24, Cud20, FPXM21, LM20b, LDW^{+}23, MBM^{+}20b, MTRA23, Mon21, NdSSSN20, RSMCP22, SZJ21, WKY22, YHL^{+}23, YTN^{+}20]. response [AAN^{+}21, BM22a, Cob22, JR23a, LMBK23, SPA^{+}21, SPSP23, SLL22, TWH^{+}22, YWC^{+}23, ZPL21]. REST [MGGA20]. REST4Mobile [MGGA20]. restart [CZZ^{+}22, EBLM22]. restoration [LHPG21]. restore [DZW23]. restricted [CZL^{+}22, MES23, ÖK22, PP20a, WY20, ZLW^{+}20]. Restriction [KMS^{+}21]. Restriction-based [KMS^{+}21]. results [PMC^{+}21, ZHW^{+}20]. retail [WWW^{+}20]. retinal [KK22b, LK22b, WJL^{+}20]. Retinex [Hua20]. retinopathy [AK22a, LK22b, SK21c, VK23b]. Retracted [FPHZ19, FPHZ20, HML20, Pan20, ZJSJ21, CWL^{+}21, CWL^{+}22]. Retraction [Ano23-30, CLDY23, Che23, FSWW23, LH23, LCW23c, Liu23, THW23, WHZL23, XL23]. retrial [GXH^{+}21]. retrieval [Akb22, BCM22, GSTS22, HHYL22, KS22d, KTK20, MLC^{+}21, Ogi21, PR22, PKVS21, SM22a, SKS20, SKSP23]. retrospective [SR22b]. reuse [BMK^{+}20, HK21, LPHK20, LBFT22]. reusing [GRL^{+}22]. revealing [HMQO23, KE22]. reverse [KRKM22, RT22c]. reverse-GPS [RT22c]. reversible [JT23]. Review [HMK23, HZZ^{+}23, LHWT20, MKS^{+}20, QMC^{+}20, YSP23, AKA^{+}22a, ASA^{+}21, BKGC23, BP23b, CBR22, CR23, DA22, DSS^{+}23, GA23, GKZ23, GN21b, Haj20, JR22, JKB22, JZR22, KKAM21, KR22a, KKK21, KS23a, KM23c, KHA22b, KJ20, KK23b, LK22b, LL21, LXC^{+}22b, LCM22, MM22, MSA21, NWX^{+}23, PKK21, RDB22, RD23, SRBH22, TSR22, VGDF22, WZMJ23a, WZ24]. Reviewing [BBD24]. reviews [BABLH21, HAA^{+}21, KUK^{+}24, MRKY22, Ona21, RM22]. Revised [RP22, AHV21]. revisited [MLVGR24]. revocable [CDR^{+}23, CYDW20, ZHDW24]. revocation [MG20, SK23]. revolution [CPH20]. reward [GYL^{+}21, WLLX21, Ano21b]. reward-with-punishment [GYL^{+}21]. rewriting [ZLW^{+}20]. RF [BSN22, MKA23, SKB^{+}20]. RF-AOA-HTSA [MKA23]. RGB [BKJ22]. RGBD [LXJ^{+}22]. rice [DS23a, SRRM23]. rich [DR23, IAO21, KG22]. RideNN [MPV22]. RideNN-OptDRN [MPV22]. Rider [DM22b, CPA22]. ridge [AD22a, AU22, BC22, MAAA22, OAA22, ÖK22, SAB22, TÖK21]. ridge-type [OAA22]. ring [DZYY22, DRMA22, KM23b]. ripeness [AA23e]. ripple [MHL^{+}20]. RISC [VMFL23]. RISC-V [VMFL23]. Risk [CK24b, JMY21, MNR^{+}22, MSA22, WZX^{+}22, XLZZ22, YL20a, ZSS22]. River [ZK24]. Rivest [DP22b]. Riyadh [LLA^{+}22]. RNA [STH^{+}20]. RNN [DSS^{+}23, TT22, VM23, VK23b]. RNN-LSTM [VM23]. road [CMJM22, ZWC^{+}22, CLL^{+}21]. roadbed [XGZ^{+}20]. roadmap [PHDS22]. roadside [Ano21-40, ÇG21]. robot [CCL^{+}22, DWZ^{+}22, DM23, HL23, JHD^{+}21, JR23a, KRKM22, KBL^{+}21, LYG^{+}21, PPM22, PS21, TDZ^{+}24, YHL^{+}21]. robot-assisted [PS21]. robotic [LG21, NKY23]. robots [ADK^{+}23, ATG23, Bul22, GKÇ22, ST23b]. Robust [AEM22, BMG22, KRJS22, KKS23, LAP23, RF23, SFJ^{+}21, ZBY22, AD22a, AU22, AFM22, ADA22, BSN22, BEKS22, CWLL20, CF21, CXXW24, FAM22, Gul22, GLA^{+}22, HAR20, JYC^{+}21, Kab23, LGX^{+}23, MAAK23, MVR23, MM21, PA21, SAN24, SO22, SA22d, SSH22, WZMJ23b, WLDW22, YAR22]. Robustness [PGD^{+}22, AA23c, BZWH21]. role [AlS21, HAAF22, Yil22c]. role-based [HAAF22]. rollback [LZZ^{+}20, NKKM21]. rollback-recovery [LZZ^{+}20]. rolling [HLC23, wSYyCsD23]. Roofline [DAW22, YKW20]. room [TWH^{+}22]. root [JMV24, JA23]. rootkit [RK21a]. rotated [BEKS22]. rotation [SYL23]. Rough [Faz23, AP22, GKM22, KUK22, KA22c, MBB22, SRL23, TWW^{+}21, VR22, WLX21, WLY^{+}22]. roulette [HXY^{+}22a]. round [LYSC21]. rounding [MKRK23, SPSP23]. route [HML21, HE23, HL23, KAR^{+}23, Sin23]. routers [HMK24]. Routing [KJMB22, OO23b, AK21, AP23, AD22b, AGSN23, BM22b, CK24a, CLC^{+}23, CL22b, CD22, CÇM22b, DS22a, DCR23, DPSJ22, FMNF22, GPR^{+}22, GJBM22, GA23, GNS22, HA22, HZD^{+}22, HUJ22, IA22, ITO21, JAR^{+}24, JCG^{+}22, JDG22, KGGM22, LCSR21, LCW23b, MK22b, MS23, MG23, Ngu21, NQ21, PB22b, PKR22, PM23b, PD23, RJ22, RG23, RT22b, SAM^{+}23, SK22a, SPS23, SL20, SPS22, SSV24, SR20a, SP23b, SPK^{+}22a, SSH22, SKR^{+}24, VPGK23, Wan22a, XXD^{+}22, ZZWZ21]. Row [YMZ^{+}20, XWD^{+}22]. Row-based [YMZ^{+}20]. rows [KSD22]. Royal [HHP23]. RPA [ZLX^{+}24]. RPL [HA22]. RS [YLJH22]. RS-PSO-LSSVM [YLJH22]. RSA [GAH^{+}22]. RSCVC [YMZ^{+}20]. RSSI [SR22a]. RTP [HMB22]. Rule [BBB22, AA22b, AYKE21, JM22, KRSR23, LK22a, LJC23, MS22b, SM21, WWG^{+}20]. rule-based [KRSR23]. rules [BM22c, BBB22, Lia24]. Rumor [LWZ^{+}22b, RDB22, KM21a, LC21, RBDS22]. rumors [ARS22b, LLSC22].

Run [RS23, WS21, BBP^{+}24, OHRS21]. Run-time [RS23, WS21]. running [SLC20]. Runtime [GB20, SSPG20, BCP^{+}23, BNB22, DJST24, RK21b, YLZ^{+}21]. Russia [Cek22].

SA [ZZZ^{+}20]. SaaS [SK20b, WY20]. SACRO [DZW23]. saddle [KSA^{+}21]. Safe [PKR22, KCM^{+}22, SVS22]. safety [FLB23, HFH^{+}21, Liu21, Man21, SYRS^{+}22, WSL21, WHH^{+}23]. SageMaker [TMB24]. sailfish [GNS22]. sale [GK23b]. sales [COZ21, LDH^{+}22]. salesman [AAMM23, KWL^{+}24]. saliency [XWQ^{+}23]. Salp [BKK22, KR23, VG21, ZFH^{+}23]. same [Kay22a]. sample [ÇKÇ22, LWHW22, LZQF24]. sampled [CNG^{+}20]. samples [BA20, LWC^{+}21, RK23a]. Sampling [LPC^{+}21, BK22b, BK23, CS22, CYQ^{+}20, KSY24, SCKU23, SK22c, VS22c, YLZC23, ZBY22]. sampling-based [SCKU23]. Sanda [HSL^{+}22]. sandboxed [MK23]. sandpiper [GK23a]. sarcasm [PS22a]. SARS [NMT^{+}23]. SARS-CoV-2 [NMT^{+}23]. satellite [AJS23, KAO22, LAE^{+}22, TA22a, ZWJG21]. satisfiability [XCJ22]. save [ACVK23]. saving [CJ21a, HLL^{+}21, YLJ22]. SBHA [XXSL23]. SBM [AAE23]. Scalability [CD24a, PS23a, dQtZWS22, YBJ^{+}23]. Scalable [AS22a, BVP22, CLE^{+}20, KC22, KMRR20, UPGCA22, BBF21, DCM21, DMS^{+}21, FPXM21, HPS23, MSB^{+}20, NMQ22, NMM21, PBD^{+}21, RKuH^{+}20, SY23, SPSP23, SMR21, UKS22, YY20b, YM22]. scale [ANP^{+}20, BEKS22, BMK^{+}20, CK24a, DCT^{+}23, FHS^{+}22, HXST22, Hua20, HZZC22, Jia22, KH22, Kal22, KVV20, LHC^{+}23, LGL^{+}21, LCZY20, LDW^{+}23, MDSG^{+}23, MY24, NKKM21, NWT21, PMP23, RC22, RVF22, RBWB21, SK22a, Shi22a, Shu22, TAT^{+}23, TFZC23, TA23b, WZL^{+}24, WZZ^{+}22, Wu22, XLL^{+}21, YÖT^{+}22, ZZQ^{+}22, ZTS^{+}22, ZCH^{+}23, ZHWY22]. scale-free [RVF22]. scale-out [PMP23, ZHWY22]. scaled [LB21]. ScaleReactor [ZHWY22]. scales [ZXL^{+}21a]. scaling [AFKS23, ADGT22, CLLB20, EK20, KK23a, LSLY20, MBC23, VB24]. scans [TNP21]. scattering [GS21]. scenario [CF21, JP21, JD22, MM24]. scenarios [CCdCC21, WHZL21]. scene [Kim21b, KK23b, TWXL21]. scenes [TYL22]. schedule [TBT^{+}21]. scheduler [BL22, GD22, HMB22, JSP20a, JSP20b, MSK22, SKH^{+}21, SB23b, UGK^{+}22, YZPL21]. schedulers [AV21, DMPR24]. Scheduling [ANAMSAR21, CHLD23, LDCD22, SSSR20, SK20b, WXCY20, AB20a, AZM20, ABESEh20, AAMM23, AE22b, AA21, ALBZ21, AB22a, AB21, ACVK23, ASMK21, BZT^{+}23a, BKS22, BSPM23, BL22, BVM22, BZL^{+}22, CCCR21, CJ21b, CZ24, CLYG22, DSC^{+}21a, DR20, DXXL20, DS22c, EK20, FSA^{+}22, GWGR20, GR22b, GMN21, GSVS21, HAR20, IAA20, JSA^{+}20, JCGT24, KKK21, KKS22, KA23, Ker22, KRB^{+}20, KT22b, KLK23, LCZ^{+}20a, LLW^{+}23, LLTZ24, LCCT22, LLH^{+}20, LGLZ20, LCL^{+}22, LZQ^{+}22, LALMGLG20, MHB22a, MNDK22, MSS22, MHPA21, MMS24, MC20, NM23, NA22b, NSPdO21, NAK^{+}22, NSK^{+}24, NJ21, PWJ^{+}21, PJP21, PCR21, QCC^{+}23, RP21a, RSM22, RS20, RBWB21, RCBTGE23, RYG^{+}21, SB23a, SEMA^{+}22, SAM^{+}23, SAC22, SKHL22, SWS^{+}24, SSRA23, TTM^{+}22, TKS22, UK23, USI21, VS22a, VG21, WWA22, WFY22b, WXLD21, WCSG20, XCG^{+}22, XYFZ23, YLZ^{+}21, YSL^{+}21a, YS23, Yu20, ZSX21]. scheduling [ZGLS21, ZLTX21, gZWfL^{+}20]. schema [JWT^{+}20, QWL24, SR22b]. scheme [AA22b, AS22d, BYPO21, BKLY20, CSWZ22, CZCM23, DAT23, DPB22b, DDUK23, DG21, DKK24, DZYY22, FCX^{+}22, FZA22, GFQ20, Gha20, GSG^{+}23, GZC^{+}21, HKA20, HTZ^{+}22, HLS^{+}20, HHYL22, DR22a, IVP^{+}23, JPH23, JHS^{+}21, JGW20, KDL20, KM23b, KK22c, KDS22, LMR22, LGM21, LCZ^{+}20a, LNC^{+}20, LZXG22, LCW23b, LZLZ23, LZL^{+}23, LL23, MG22, MGN^{+}22, NTK22, Ngu21, OO22, PMC^{+}21, PK23a, PSB^{+}20, PGL^{+}23, PB22b, PDRM24, PA21, RG23, RKC22, SY23, SM23a, SAL22a, SHV24, SJ23b, SSH22, TWL^{+}20, TCRP23, TJC24, TLL20, TY22, TZ23, TT23a, UDS21, VB22a, WCWG20, WFY^{+}22a, WDW^{+}22, XWC^{+}22, XYLW21, XZL^{+}20a, YR21, YZXL22, ZLM22]. scheme-based [Gha20]. schemes [Gha20, HSR23, LLJR21, OS21a, RMR^{+}22, SK23, SG24]. Schiff [RK23b]. school [CT22, CB22]. Schur [SAL22a]. Science [BSML21, KS21a, OMA^{+}23, CWDM^{+}21, CJP^{+}21, GCS20, KLJ21, LA22, MPLW24, MBT^{+}20, PKB^{+}23, RBDS22, VCS^{+}21, Wu22]. scientific [AA21, CCCR21, CHMC21, CDN^{+}21, GCQ^{+}24, KSSK22, MFA^{+}21, NGOS22, NSPdO21, OGA^{+}22, SB23a, SKB^{+}20, SZZ^{+}21, SAL^{+}22b, TLQ21, WCSG20, ZBZ^{+}20, dOPBdO21]. scoliosis [FMJ^{+}22]. scope [DZLH20, YM22]. scorched [Anb22]. SCORE [RB22]. scores [MAB22]. scoring [GDS23, SY23]. scraping [DAdBM^{+}24, KTU^{+}21]. scratchpad [ZSC^{+}21]. screening [BGG24, LZQF24, SSMT22a]. scrutability [BPT^{+}23]. SDCCP [LLW^{+}22b]. SDN [MV22, AA23a, AKEC20, BMJ^{+}20, CLC^{+}20, HJT^{+}20, HE23, KKP24, MKS^{+}20, Mon21, NLG^{+}20, PCG^{+}21, SNBAI24, SSS^{+}20, Sin23, WCWG20, WYZAD20, XZXV21]. SDN-based [HE23, CLC^{+}20, WYZAD20]. SDN-enabled [NLG^{+}20]. SDN/NFV [PCG^{+}21]. SDR [SK20a]. SE [XLL^{+}20]. SE-ResNeXT [XLL^{+}20]. Sea [SAC22, AFBM^{+}23, SSS23b, VK23b, WLX21]. seagull [LJBS23, SSMT22a]. seamless [AA23a, IVP^{+}23]. Search [AK22c, LST22, WWG21, ATS22, AJ21, AI21, BMcKGK22, Bul22, CNG^{+}20, CL23, DÖD22, DDUK23, DS20b, DR20, DD24, GSS^{+}23a, GDFDF22, GSG^{+}23, GVSS22, HR22a, HKA20, HGM^{+}23, HLZ23, INY^{+}23, KA22b, KS22c, KSY24, LLC^{+}22a, Lia24, LYG^{+}21, MNDK22, OGA^{+}22, OO23b, Pan23, PR24, MR23a, PB22b, RCK22, RG22, RW24, SRGK24, SMBM24, SG22b, SVD^{+}22b, SNGK21, SS22d, SP22c, SBS22, MKL21, XW24, XLWX20, XZYW23, Yu21, YÖT^{+}22, ZLTX21, ZLV22, ZS22]. search-based [DR20]. Searchable [FZA22, MG20, XZYW23]. Searching [DKH20]. secret [Gha20, KYP21, RKMS24, ZZY^{+}22b]. section [KK21a]. sections [AAB23, LZY^{+}21]. sector [SAC22]. Secure [AS24, DCK21, GJPS23, KJ21, MG23, NR22, OO22, PCK23, SDR23, SS23c, VH22, XHZHXBQX22, AMAT22, AMB23, ASC22, AGSN23, BXH^{+}23, BT21, BVS20, CAS24, CL22a, CBK23, DAT23, EI22, FK24, HLT23, HLS^{+}20, HZQ^{+}24, DR22a, JPH23, JCG^{+}22, KGRG24, LMR22, LCSR21, LFX^{+}20, LWL^{+}23, MG22, NS23, NM20, NAR^{+}22, PM23a, PSP24, PK23b, PSB^{+}20, RGKK21, SK23, SAN24, SS23e, SL20, SS22e, SKS^{+}23, SJ23a, SHV24, SJ23b, TJC24, TSG21, TLXX21, TLX22, TXT24, TKG^{+}23, VDL23, VB22a, XZL^{+}20a, ZWO^{+}20, ZHDW24]. Secured [LS23a, CA22, FZC20, KRK21, NA22a, PKR22, RJ22, SM24a]. Securing [GMP^{+}20, RSMCP22, DRMA22, NVV24, PSK23, SEC22, SK22d, SDR20, SAGS24, YLM21]. Security [CK23, HNS^{+}21, JH21, MAB22, SN22b, AS23a, ABB22, BTDJ21, ÇKÇ22, CI20, CYC21, CWW^{+}23b, CD22, DBB24, EABZB21, Gho21, GLM^{+}22, HJT^{+}20, HXZS23, HK22b, HLT23, JZR22, JMY21, Kab23, KHEF22, KM23a, KV22b, KK22e, KSCL21, MVR23, MK22b, MM22, MMKA23, NSKS21, NJK22, NY22, Ogi20b, Ogi20c, OO20, OS21a, Ogi21, OLF21, OSMM23, PZZ^{+}23, PK22c, SR22a, SV21, SLG^{+}20, SSS23a, SG24, SZW^{+}22, SS23g, TDZ^{+}24, TCW^{+}23, VK23a, VP22b, VBM^{+}21, XHM22, YWC^{+}23, ZZK^{+}22, ZML^{+}23]. Security-based [SN22b]. Security-centric [MAB22]. Sed [TLXX21]. Sed-Dedup [TLXX21]. SeDaSc [ARFA21]. seed [DT22b]. segment [FXX22]. segmental [KKKS23, SBS22]. segmentation [AAEA20, AMM^{+}20, BABS21, BCK22, CWLL20, CXX^{+}22, DBN^{+}22, DSJ22, ESS23, FMJ^{+}22, GDSS22, GM22b, GCF^{+}20, HZZC22, JM23, KSS21, KK22b, KSJ23a, KIAA^{+}22, KAAR23, LLZ^{+}20, LW20b, LLL^{+}24, LS23b, LLLX20, LYG^{+}21, NBS^{+}22, NSBT21, PGD^{+}22, SET^{+}22, SGH23, SBB^{+}20, SS22g, SKÇA23, TYLY23, TNP21, VRR^{+}22, WMS^{+}23, XLL^{+}20, YRV^{+}23, YSLX22, ZRY^{+}22, ZTS^{+}22, ZFZ^{+}20, ZCW^{+}23]. segregation [AYH^{+}22]. Seismic [LML^{+}23, GLRB21]. SeisSol [DGB24]. seizures [AHT^{+}20]. select [CdRNB23, DRM22b, MKRK23]. selecting [AEMS24, SVD22a]. selection [ASAAAA22, APM^{+}21, Anb22, AV22b, BDK22, BD22, BBD24, BC23b, BA24, BPK24, CU22, CKKK20, CEN22, CSC22, Cob22, DDZ^{+}20, DT22b, ECIB20, FEH22, Faz23, GWA^{+}24, Gün23b, GRG^{+}22, HK22a, HUJ22, HC22, HXY^{+}22a, IAAA22, JB22a, JM22, KT22a, KJHM21, KK22e, KSSP22, LHWT20, LLZ^{+}21a, LLMX21, LM24, LLC^{+}21b, MQEK21, MS21b, MS21c, MC20, NM23, NR23, NBHN22, ÖUG22, OE22b, ÖK22, PKC^{+}20, PK23a, Par22, RT22a, RK23b, RAK22, SG22a, SKB^{+}20, SM22a, SPS23, SS22a, SV21, SS24, ST23c, SSK23b, SSDY20, SKS22, SM24c, SLHW20, SM23d, VY23, WZY^{+}22, WER^{+}23, WRE^{+}23, WLY^{+}22, WYLL24, XWW^{+}21, YLZW23, YFF22, ZX20, ZZQ^{+}22, ZLC^{+}22, Zhu21, dRdSC^{+}21, dRdSZ^{+}23, vdSTC21]. selection-integrated [Gün23b]. Selective [AG22, HMK24, NJK22]. Self [ADK^{+}23, VKSS23, VB22b, VGDF22, WFHC21, XLX^{+}21, ALR22, BL22, CL23, EGB21, FKO22, FKK23, GFPGT21, HLZ23, HZZC22, KK24, KK21a, KKM21, LAC21, LZQ^{+}22, PSMM22, PVVS22, RS23, SS23f, SSI22, DM22a, TNP21, TS22, WZC^{+}22a, ZYY^{+}24, ZZZ^{+}20]. Self-adaptation [VGDF22, WFHC21, ALR22, FKK23]. self-adaption [ZZZ^{+}20]. self-adaptive [SS23f, SSI22, TS22, ZYY^{+}24]. self-adjusting [FKO22].

Self-attention [VKSS23, CL23, HLZ23, HZZC22, DM22a]. self-diagnosis [LAC21]. self-enrichment [EGB21]. self-healing [RS23]. Self-improved [VB22b]. self-initialized [TNP21]. self-learning [LZQ^{+}22, WZC^{+}22a]. self-managing [GFPGT21]. self-organizing [GFPGT21]. self-scheduling [BL22]. Self-stabilizing [ADK^{+}23, KK24, KK21a, KKM21]. Self-training [XLX^{+}21]. selling [KRSR23, LK23]. Semantic [MFE^{+}23, AZI20, ALR22, Ano21j, AA24, CXX^{+}22, DPB22a, EGB21, GZC^{+}22, HZY^{+}21, JR22, KK22b, KS22d, LLLX20, LLX^{+}24, OS21a, OS21b, PHZ^{+}22b, PHZ^{+}22a, RA21, RCYuRH21, SBA22, SGH23, TYLY23, VSS23, WHH^{+}20, XLZZ22, YMZ^{+}20, ZRY^{+}22, ZNLL22, HL20]. semantic-weight [ZNLL22]. Semantics [SZ24, GR22a]. Semi [ST23b, AWS^{+}22, JGW20]. semi-Lagrange [JGW20]. semi-supervised [AWS^{+}22]. Semi-uniform [ST23b]. semisupervised [LLJ^{+}22a, ZFZ^{+}20]. SemRec [RSJ21]. sender [MSSM24]. sender-initiated [MSSM24]. sense [SGS^{+}23]. sensing [AFM22, FKGO22, GDA22, JPN21, KK23b, LGM21, LYW^{+}21, Rav23, SHA^{+}22, SGH23, XFH^{+}21, YLGY20, ZX23b, ZX23a]. Sensitive [SK22a, AS22a, CJC22, CYW^{+}22, HDS^{+}23, JZC^{+}23, JKKL21, KM23b, KM23c, LNC^{+}20, LSZ24, LLMZ21, NN23, RN22, SAHAN22, WQY^{+}22, YLS^{+}23, ZHX^{+}21]. Sensitivity [LXYY21, BMK^{+}20, LTLX22, PGD^{+}22]. Sensor [SM23c, AP22, AM20a, AS23a, ASC22, ASMS21, AGSN23, ASRN23, BK21, BLT22, BM22b, BGPQ21, CLDY21, CSWZ22, CLC^{+}23, CL22b, CG22, DPSJ22, ETKD23, FCX^{+}22, FZC20, GJBM22, GA23, GSS23b, GNS22, GKAO20, HSL^{+}22, IA22, JH21, JB22b, JS22a, JCGT24, K22, KSJK21, KK22c, LCSR21, LL21, LAC21, MKA23, MS21c, MRAS^{+}23, MG23, MP23, Ngu21, PPA22, PB22b, PM23b, RJ22, RK23c, RP22, RT22b, RKL21, SDR23, SV22b, SSM^{+}21, SS23f, SR20a, SM24b, SP23b, SPK^{+}22a, SKR^{+}24, TTM^{+}22, URK^{+}22, WYLL24, YLGY20, ZWW^{+}21]. sensor-cloud [ZWW^{+}21]. sensorial [dOdMC^{+}20]. sensors [BTDD20, RR23a, ZGH^{+}22]. sensory [HR22b, PYC^{+}20]. sentence [KBBH21]. Sentiment [DBD22, KK22d, Ona21, APM^{+}21, ARS22b, AS23c, AGM23, BABLH21, CZG^{+}20, CK21b, HAA^{+}21, Haj20, Kab22, KBBH21, KJ20, KS23c, MRKY22, RSJ21, SSK^{+}23a, SZGR21, SSP23, TTZX22, XGX^{+}21, ZLCL21, ZHT^{+}23]. sentiment-based [AGM23]. sentiment-embedding [KBBH21]. sentimental [Aru22, DSS^{+}23]. sentiments [Asl22, VM23]. Sentinel [SPHP21]. Sentinel-1 [SPHP21]. SentiWordNet [SSK^{+}23a]. separable [IK22, KMS^{+}22]. separating [ZWX21]. separation [DZCL22, KGE^{+}20]. sequence [ABB22, BA20, EASR22, JNMG21, LZC^{+}20, NMM21, PQKDT21, RV21, SY23, SAL22a, XYY24, XDH^{+}20, rSN21]. sequence-based [rSN21]. Sequences [SV22d, CZUPA23, HIEH22, JBCI20, KB22, LYSC21, NMT^{+}23, OA22, SV22a, UPGCA22]. sequencing [KT20]. sequential [IAAA22, SKCS23, SPWX23, YY22, ZTM21]. series [BBD23, CJP^{+}21, GCP22, JS22b, LZL^{+}20b, LWHW22, RCS20, SK21a, Yal22]. SERU [XLL^{+}20]. Server [CYDW20, HMR24, MV22, Mon21, OXBL23, SZY^{+}22, TLX22, XZL^{+}20a]. server-aid [TLX22]. Server-aided [CYDW20]. serverless [DS24a, KSB23, LYF22, MK23, NRMB23]. servers [LLR^{+}21, MPG20, SZZ^{+}22, YLJ22]. Service [KEMZ22, LWZ22a, SKRS21, ALR22, BC23a, BABLH21, Ben22, BJGF20, BKS22, BZK^{+}21, CCCR21, CHL^{+}24, CSC22, CD22, Dah23, DP22a, DCZ^{+}22, DWY^{+}21, DJGF21, EABZB21, FM20, FCZ20, GHRM21, GCP22, GWA^{+}23, GWA^{+}24, HK22a, HGHD22, HPCK22, JMY21, KXL^{+}21, KVV20, KQK^{+}20, KK23a, KABE^{+}20, KLK23, LCR^{+}24, LmJdL^{+}22, LGZ^{+}22, LSLY20, LXC^{+}22a, LXKW23, NSKS21, OO21, OO23a, PPA22, PPM^{+}20, PKK21, PK22c, dQtZWS22, RR22, SR22c, SB21, TK22a, VJ22, VCS^{+}21, WRJ20, WLCW20, WFHC21, WWZ^{+}22, WZZ^{+}22, WJD22, XHZ^{+}21, XCY22, YFF22, ZLCS21, ZMLW23, Zhu22a]. Service-aware [SKRS21, CD22]. service-based [OO23a, PPM^{+}20]. service-level [BZK^{+}21, KK23a]. Services [HFFA20, Ano21-40, AYJ^{+}22, CCM22a, CYDW20, DJ20, EKS20, GKLS23, HL20, dCJAAdOD21, KKP20, LBA23, LZL^{+}23, MS21b, MGGA20, PYC^{+}20, PK23a, PR23, SZL^{+}22, SB21, ASB23, WGLL23, dCMA23]. session [AMAT22]. set [AP22, AI22, AS23c, AD22c, BK22b, BK23, GKM22, INY^{+}23, JAM24, KA22c, LHK22, MBB22, RMR^{+}22, SM24a, SRL23, SYJL20, TWW^{+}21, VS22c, WLY^{+}22]. sets [PMC^{+}21]. settings [KABE^{+}20]. setups [HXY20]. SEV [ZCX^{+}21]. SEV-Net [ZCX^{+}21]. severity [BP23a, DS23a, ÖUG22, ZCX^{+}21]. SFO [SMM22]. SGD [YBJ^{+}23]. shading [HHP23]. Shadja [KG23]. shaming [AC22]. Shamir [DP22b]. Shanghai [WZ20b, WZ21, WN23, WNZ24]. Shannon [DS24b]. shaped [KMB23]. shaping [WLLX21]. sharding [NAR^{+}22, TXT24, XHST20]. share [DKK24]. shared [BTP^{+}21, BPT^{+}23, CdOO^{+}20, ÇKÇ22, Cud20, DLY^{+}23, LWZY23, PR24, RRB22, TBT^{+}21, WK20, XZL^{+}20a]. shared-hidden-layer [RRB22]. shared-memory [CdOO^{+}20]. Sharing [RPM22, ARFA21, BVS20, FYH^{+}21, Gha20, KGRG24, KRK21, KYP21, KT22b, LS23a, LMR22, LWL^{+}23, LZLZ23, OT20, OO21, PM23a, RKMS24, SVB23, SSP22, VPB^{+}23, XDH^{+}20, ZZY^{+}22b, ZML^{+}23, ZHDW24]. shark [LXC^{+}22a]. Shasta [RMS24]. shearlet [AJK^{+}23]. sheet [WJLC21]. shepherd [DS23a, GCS23, SSS23b]. shift [Ano21-37, PS22a]. shifting [SZJ21]. Ship [SPHP21, LSL20, XCX^{+}20]. shipment [FMNF22]. shoe [MHL^{+}20]. shop [CZ24, LZQ^{+}22, ZSX21]. shoplifting [ASJ24]. shopping [MRKY22]. shore [DKC^{+}21]. Short [SMR23, XLZL22, YHOY22, ZCH^{+}23, ATC23, BKK23, CL23, CU22, CÇ22c, CM21, DR22c, DM22d, HR22a, HZY^{+}21, IIK^{+}23, JR23b, KTM22, KH22, Kab22, Kar22, KDC22, LK23, LWLZ20, PT22, PVRM22, SV22b, SMBM24, SWK22, SVD^{+}22b, SHZY20, SM23b, YRV^{+}23, YMWA21, YW22, ZLCS21]. short-long [YW22]. short-range [SWK22]. Short-term [SMR23, XLZL22, ZCH^{+}23, ATC23, BKK23, CL23, CÇ22c, DM22d, IIK^{+}23, KTM22, Kar22, LWLZ20, SV22b, SMBM24, SVD^{+}22b, SHZY20, SM23b, YRV^{+}23, ZLCS21]. shortcuts [LWZ^{+}20]. shortest [MY24, MGB^{+}23, NQ21, SA22a]. shortest-path-based [MY24]. Shorting [LK23]. shot [Abb23, BST^{+}22, Mis22, WXS^{+}23, XLX^{+}21]. Shuffled [DS23a, GCS23, Ben22, SSS23b]. SI [DSYF22]. SI-EDTL [DSYF22]. Siamese [WLZ21a, ZLL^{+}22]. Siamese-based [ZLL^{+}22]. side [GZ20, ZLYS21]. sided [JGW20, MHL^{+}20]. sign [RD23, SMD22a]. Signal [RT22c, IDA22, JCL^{+}20, Kim21b, LLKS21, MK22a, PBD^{+}21, TSB23]. signalized [AK22c]. signals [AS20, BMV22, DZCL22, KK21b, QZB^{+}23, TA22b, TÖ21]. signature [CCM22a, DZYY22, DRMA22, GZC^{+}21, KM23b, LGZ^{+}22, SJ23b, VDL23]. signature-based [DRMA22, VDL23]. signatures [DGP20, HLC^{+}21, RB22]. signboard [JHG23]. signcryption [JHS^{+}21, PSB^{+}20, QWW^{+}22, SM24a]. significance [Kat23]. Signs [SKP23, SKP22]. silhouette [KWZ^{+}21]. silicon [SAF^{+}23]. Sim [WZL^{+}22]. SimAnMo [BNB22]. SimAnMo-A [BNB22]. SimAS [MC20]. SIMCard [SC22b]. SIMD [FFLM21, KML21]. similar [Ano21g, SJA^{+}22, UPGCA22]. similarity [CLT^{+}21, DPB22a, EVVR21, GB23, HZY^{+}21, KGGM22, PM22, RA21, RT22a, WLZ21a, WHH^{+}20, YYPR22, YHOY22, ZTF^{+}20]. Simple [INY^{+}23, DTK24, HCH^{+}21, HHC^{+}22, HLW^{+}21, ZBY22]. SimpleSync [ZQL^{+}21]. simplex [KÖ22]. simplification [BJWY20]. simulated [GMS^{+}21, KT20, SBB21, SCP20, SB23b, XDL22]. Simulating [BSEN20, HLO^{+}21, QPS20]. Simulation [ASANR22, JZB20, LAP23, MAAA22, AFBM^{+}23, BEÖ23, CPYC21, CS21, DM22c, FZ23, FPÁ^{+}20, GFPGT21, GCSZ24, HPH^{+}20, K22, LAK22, MBC23, MAG^{+}20, MC20, NGOS22, PPM^{+}20, SLLMTS^{+}24, STJ^{+}20, TWH^{+}22, WSL21, WZL^{+}22, WZ20b, Wri22, YK23, YSC^{+}20, ZBY22, ZZL^{+}22]. simulation-assisted [MC20]. simulation-driven [NGOS22]. simulations [GMSM21, MNYN21, PBD^{+}21, SPC^{+}21, VCFZ20, XAC^{+}20]. simulator [TWW^{+}22, WXH^{+}22]. simultaneous [MSS^{+}20]. Single [AN23, HMR24, HPCK22, LMM23, BST^{+}22, GGCGS20, HMK23, JCL^{+}20, LZF20, LYSC21, Liq22, MHL^{+}20, PP21, SAPC21, WXS^{+}23, WZSZ20]. Single- [HMR24]. single-image [WZSZ20]. Single-linkage [AN23]. Single-objective [HPCK22]. single-round [LYSC21]. single-sided [MHL^{+}20]. Singular [WZZ^{+}22, FQD^{+}23, HGNN22, KDL20, KSP21]. sink [ZGH^{+}22]. sink-based [ZGH^{+}22]. Sinkhorn [Pan24]. sinoatrial [MAG^{+}20]. site [KS22b]. situ [DPdS^{+}23]. situation [BSEN20, CZTC22, USK23, XLL^{+}23]. situation-aware [USK23]. situation/reaction [BSEN20]. situational [LHX^{+}23]. Size [BEÖ23, ECIB20, HXST22, KK22f, SK22c]. sizes [KMB23, LNI^{+}24]. sketch [ZZZ^{+}20, ZZZ^{+}20]. skew [GHT20]. skewed [RMR^{+}22]. skin [ESS23, GM22b, JM22, KIAA^{+}22, MGSB23, RMMK23, ZWY^{+}24]. sky [XXSL23]. skyline [ZWC^{+}22]. SLA [BF24, KB21, LMGG20, ZOS^{+}21]. SLA-guaranteed [KB21]. SLAC [BBP^{+}24]. SLAM [LXJ^{+}22]. SLChain [TXT24]. sleep [MTD^{+}20, MTT20, TAH22]. sleep-and-wake [MTT20]. SLEPc [BBB^{+}20c]. slice [RN22]. slices [WMY^{+}21]. slicing [DJF21]. sliding [NY22, ZTMC22, ZTMC23]. slime [SS23f]. slot [MHL^{+}20]. slots [SZJ21]. SM [Gho21]. SM-Detector [Gho21]. SMAC [LLC^{+}22a]. small [ASANR22, CÇ22c, FLP20, LWHW22, XLWX20]. small-file [FLP20]. Smart [AYH^{+}22, AAT21, CLDY21, FPÁ^{+}20, GLM^{+}22, HMB22, JR24, KTU^{+}21, KM21b, TA23a, XFPY24, dSNdL^{+}23, dRdSC^{+}21, AA23b, ALR22, Ano21i, AJAA21, BTDD20, BXH^{+}23, BS23b, BCC^{+}21, BTDJ21, BZK^{+}21, ÇKÇ22, CRB23, DKL21, DJ20, GLN23, GKLS23, GQ21, UZAA21, HT21, HLCH20, JDLP22, JP21, JHS^{+}21, KHEF22, KKAM21, Kat23, KAP20, KKE^{+}22, KS24, KCM^{+}22, LHWT20, Man21, NAR^{+}22, PK22c, RPMA22, SJA^{+}22, SAN24, SLJ23, SRRM23, SG24, SR23, SP22c, SK22d, SR20b, SB21, SJ23b, TY22, TAH22, TG23, VPQ22, WZ21, YLW^{+}21, ZBC^{+}21, ZOS^{+}21, ZUTK23, AAE23, Ano21f]. Smart-grid [AAT21, SG24]. smartphone [JPA^{+}23, RR23a]. smartphones [MGGA20]. smartwatches [GMP^{+}20]. smear [BP23a, PAN22]. smell [LXC^{+}22a]. SMiShing [Gho21]. smoke [ZXLD21]. smoothing [THT20]. SMOSA [ZGLS21]. SMS [AAMAA22]. SMSE [DS24a]. snapshot [NKKM21]. snowflake [JWT^{+}20]. SOA [LW20a]. SoC [CYK^{+}21, JK22a]. soccer [LJC23]. Social [EAS23, LC21, SSP24, YSP23, YL20c, YGZ^{+}21, APM^{+}21, Ano21-35, Ano21-39, ARS22b, AS23c, BV22a, BKM^{+}21, BWTJ20, BKLY20, CCM22a, CA23, DR22b, DBD22, DS23a, DGM21, EVVR21, GXL^{+}20, GlSL^{+}23, GM21, GK22a, HNS^{+}21, HAR24, HBB20, HGMK21, HJZ^{+}22, JS22c, JZB20, JD23, KM21a, KMR22, KUK^{+}24, KJHM21, KBS^{+}22, LK22a, LWZ^{+}22b, LDZ^{+}22, LX22, LSQW21, LHL^{+}22, LWZ^{+}20, MAR23b, MS22b, PS22a, RSJ21, RSJ22, SJ22, SVM24, SSN22, SLJ23, SKSB20, SS22d, SYRP22, SRIB23, WGZ^{+}20, XCZ^{+}21, XXD^{+}22, YYPR22, YB23b, ZR24, ZLM22, ZZLZ22b, uZKH^{+}20]. society [CPH20]. socio [SLL^{+}23]. socio-demographic [SLL^{+}23]. SoCs [MCL^{+}20]. Soft [ZX23a, CPH20, KCL^{+}20, KJ20, MSN22, SYRP22, VY23]. softmax [RP21b]. Software [BSN22, KT23, UWF^{+}21, AAMM23, AMR^{+}21, AA23d, Anb22, BMcKGK22, BSPM23, Boz22a, BPK24, CNG^{+}20, CCBA23, CTA^{+}23, CK24b, CD22, CÇ22c, Cud20, GYZ^{+}20, GM22c, HLL^{+}21, HLG24, IQS^{+}22, KT22a, KUK^{+}24, KKE^{+}22, LCZ^{+}20a, LZZ21, LSZ24, LLW^{+}22b, LGLZ20, LZXG22, MK22b, MS23, PA23, RRJ23, RZ21, RG23, SBB21, SKK22, SGHL20, SDR20, TTB^{+}22, VRV23, VNP^{+}23, WY20, WZZ^{+}22, WC22, XHZHXBQX22, YY20a, YLZC23, YLW^{+}22, YHL^{+}23, YFdP^{+}23, VK24, MV22]. software-defined [AAMM23, AA23d, CTA^{+}23, CD22, CÇ22c, LCZ^{+}20a, LLW^{+}22b, LZXG22, MK22b, RG23, SDR20, XHZHXBQX22, YY20a, VK24]. software-distributed [Cud20]. soil [AJS23, RM21, WKY22, LM20b]. solar [AAT21, DCT^{+}23, KNK22, NNVD22]. solar-powered [DCT^{+}23]. Solid [DZW23, Kum21, WWW^{+}23, ZTY^{+}22, ZCR23]. solid-state [WWW^{+}23]. Solidification [HSO^{+}21, HOS^{+}21, HMR24]. Solubility [PKSM24, AA23e]. solution [AFF22, CCBA23, DE20, GNMELC21, GADM20, KGGM22, KMZ^{+}20, KSA^{+}21, Le23, LLM^{+}22, Pan23]. solutions [ÇGB23, IA23, ORP21, OS21b, PUL20, ZS22]. solvers [KSA^{+}21]. Solving [DS22b, LGL^{+}22, ASMS21, BKD22, CD24a, CZ24, CMY21, DMS^{+}21, GFA21, MM21, PJKO24, XA22, Yil22b]. SOM [AAARR20]. soma [WYE23]. soma-synapses [WYE23]. some [KA21a]. Sommon [RT21]. Sorensen [SAPC21]. Sorensen-single [SAPC21]. Sorting [Yil21c, DCWM20, HXY^{+}22b, Wan20, yXlLyGX21]. sound [RGM22, RKL21, ZHJ20, ZHJW21]. sounds [Yil22a, YRSO23]. source [DS20a, DKH20, FZ23, JYW^{+}20, MT22, WWZ^{+}22, Zhu22a]. source-filter [JYW^{+}20]. sources [Ano21j, uHAU23]. SP [Bul22]. SP-search-based [Bul22]. Space [ZSC^{+}21, BBBC22, CD24a, GK22a, KMY24, LLC^{+}22a, MY24, SWZW20, SAF^{+}23, XLY^{+}23, ZLZ^{+}22b, ZWZ^{+}22]. Space-address [ZSC^{+}21]. spaces [OHFF20]. spam [AAMAA22, AMJK21]. spanning [MR23a, ST22a]. Spark [ABA22, Ano21j, KKR23, LM24, NWW^{+}22, dOPBdO21, ASL20, BVM22, DHSG23, GHT20, NMT^{+}23, RCS20, WWG^{+}20]. Spark-based [ABA22]. Sparrow [WWG21, Lia24]. Sparse [AAQOT23, LH21, YLGY20, AML^{+}22, AAG^{+}22, Anb22, BFM^{+}23, DE20, GCF^{+}20, HYG20, HCG21, KS21b, LHC^{+}23, LCL^{+}20, MIN^{+}23, wSYyCsD23, VVK23, XLL^{+}21, XWD^{+}22]. sparsification [JLZ^{+}24]. sparsity [GK23b, MLZ^{+}20]. sparsity-tolerant [MLZ^{+}20]. SPASC [XHZ^{+}21]. Spatial [LXKW23, Kal22, KS22d, PKB22, SSH22, XCY22, Yil21c, ZXLD21]. Spatial-temporal [LXKW23, ZXLD21]. Spatio [MS22a, ZK24, CLWX21, LHK22, TSD23, WHH^{+}23, XCD^{+}20]. spatio-clock [WHH^{+}23]. Spatio-temporal [MS22a, ZK24, CLWX21, LHK22, TSD23, XCD^{+}20]. spatiotemporal [LP21, WLZ^{+}21b]. SPCA [NWZ^{+}21]. speaker [SMM22]. Special [ABZS20, Ano21-41, Bad23, FZT22, IS23, JB20, JC21a, JAC^{+}21, JSLL20, KS21a, KSY23, KLL^{+}21, LMMS24, LG21, PJ21, RT20, SCH22a, SCH22b, SBG20, SKE22, TC22, Vin21, Wri22, WYZAD20, XZ20b, XZ20a, BI23, CSW20, FD20a, LXZ20, Liu24b, WJS21, Wu22]. Specializing [JSZS22]. speciation [PBD23]. speciation-based [PBD23]. specific [CDN^{+}21, RSJ22, SBGC21, ZHW^{+}20]. specification [MM22, PHZ^{+}22b]. speckle [MRS^{+}21]. SpecMap [RJH^{+}23]. Spectator [LSQW21]. Spectator-filter-spreader-stifler [LSQW21]. spectra [AKLK24, WPL20]. Spectral [PKB22, HVB22, JTY^{+}21, MLZ^{+}21a, RJH^{+}23, SHA^{+}22, Yil21c, ZQW^{+}21]. Spectrum [MSS^{+}20, BGG24, BB23, CÇM22b, IVP^{+}23, JPN21, WLX21, YY22, ZX23a]. Speech [JTY^{+}21, BMV22, CMJC23, RD23, SS21]. speed [JHG23, MKRK23, MS21a, NBK22, TT22, ZZZ^{+}20]. speeded [BEKS22]. speeded-up [BEKS22]. speeds [KMB23]. Speedup [ZSQ22]. spell [SS23b]. Spider [RK21b, ZGLS21, ASRN23, HGMK21, JD23, KK22d, KDS^{+}23, MK22b, SKSB20, ZZLZ22b]. SPIHT [KSK^{+}20]. SPIHT-based [KSK^{+}20]. spike [VRV23]. spiking [DLC^{+}21, EBDB22]. spillage [AZA20]. spin [MTD^{+}20]. spine [FMJ^{+}22, PSMM22]. spine-kernelled [PSMM22]. spiral [SRGK24]. spite [ADK^{+}23]. spleen [ZBC^{+}21]. split [LYBZ23, SV21]. splitting [FZZZ23, LZXG22, NKST24, SS23b]. SpMM [QKSK23]. SPoFC [YLZT23]. spoofing [LMM^{+}22]. sports [LXZ22]. spot [ÇG21]. spotted [NAK^{+}22]. SPPGKM [MG22]. spread [BA20, LDZ^{+}22]. spreader [LSQW21]. spreading [HLO^{+}21, LWZ^{+}22b]. spreadsheet [Boc21]. SpTRSV [Yil23]. SQL [IP20]. SQL-on-Hadoop [IP20]. Square [LWZC21, BEÖ23, CLT^{+}21, DCM21, NKY23, RM21]. squares [XCY22]. squeeze [ZKZ^{+}23]. squeeze-and-excitation-attention [ZKZ^{+}23]. Squirrel [PKSM24]. SR [RJ22]. sRSP [YM22]. SSAE [XLL^{+}21]. SSAE-MLP [XLL^{+}21]. SSD [LLW^{+}21]. SSDs [CPLX21, ZWX21]. SSHS [AA23a]. SSO [SKSB20]. stability [FCZ20, JZB20, LLO21, WWW24]. stabilizing [ADK^{+}23, KK24, KK21a, KKM21]. stable [HUJ22, JP21, KDS22]. stack [LDS^{+}23, AMR^{+}21]. Stacked [SS23g, XLL^{+}21, Anb22, KSVP22, MES23, SD23, SPK22b, TYL22, ZHX^{+}23, ZLW^{+}21, ZQW^{+}21]. stage [AAD20, CWH^{+}24, GHT20, WLL21a, XWC^{+}22, ZMLW23]. Staged [WLL21a, RMMK23]. stained [RK23b]. Stake [LK23]. Standard [TH22, SKPV22, ZWT22]. standardized [EGB21]. standards [vdSTC21]. star [AB20a, GQ24, STISM21, SRGK24]. starling [DCR23]. StarPU [DJST24]. start [YL20a]. start-up [YL20a]. starvation [SHBC20]. State [CFR^{+}21, ELIG23, AFF22, BBBC22, CI20, CON23, DZW23, HLCH20, KKC22, Kum21, SM22b, SBB^{+}20, SZZ^{+}24b, WCWG20, WWW^{+}23, ZTY^{+}22]. State-of-the [ELIG23]. state-of-the-art [CI20]. statement [YS22]. States [ST21, YYZS22]. Static [KHPH20, KRB^{+}20, LGW^{+}22b, AS22c, KRJS22, RJH^{+}23, RAK22, YZX^{+}22]. station [LWZ22a, NM23, WZX^{+}22, ZZLZ22b]. stationary [MNR^{+}22, SG22c]. stations [LLC21a]. statistic [ÖK22]. Statistical [SZS20, ZWJG21, DPS21, SR22a, SKB^{+}20, TK22a]. statistics [AI22]. stealing [DMPR24]. steel [HWBZ21, WJLC21]. steering [CCZ^{+}21, ZWC^{+}23]. steganalysis [SLHW20]. steganography [CCD^{+}20a, DG21, SKSP20, SS23g]. Stein [AAA^{+}22a]. STEM [KLJ21]. stemming [SG22b]. stencils [SMCM^{+}22]. stenographic [JT23]. step [KW21, NBK22, VBP^{+}24]. stepping [KBL^{+}21]. steps [OHFF20, YTN^{+}20]. stepwise [XLX^{+}21]. Stereo [LCZY20, CCZM23, HWY^{+}23, LZSC23, SKA23]. stiffness [XWD^{+}22]. stifler [LSQW21]. stimuli [JJZ^{+}21]. STLB [MSSM24]. Stochastic [AFK^{+}22, BDG^{+}23, JXL^{+}23, AV22b, CCCR21, CT22, DÖD22, GK23c, ÖK22, MR23a, WSM^{+}20, WHH^{+}23, WLDW22, WSL^{+}20]. stochastic-Variable [MR23a]. Stock [KSJ23b, SZGR21, TI22, AGM23, KKR23, NTB23, PJK23, XY21]. stomach [ZBC^{+}21]. stone [GNMELC21, VKSS23]. Storage [AS24, AMB23, CK23, FZA22, FXX22, GFQ20, GN21a, HCG21, JB24, KRK21, KS24, LLJR21, MDSG^{+}23, MJAM23, MTRA23, MCL^{+}20, MSB^{+}20, NA22a, NN23, PSB^{+}20, SS23e, SR22d, TZ23, TT23a, TXT24, WYL^{+}22, XWD^{+}22, YWQ^{+}21, YSS^{+}21, YZXL22, ZWO^{+}20, ZLO^{+}21, gZWfL^{+}20, ZWZ^{+}22]. storage/access [KS24]. stored [HMQO23]. stores [FCMM20, STISM21]. storing [Ano21-43, YSS^{+}21]. storm [GGJ^{+}24, HTZT23]. straightforward [QMC^{+}20]. strategies [AO23, CMT20, FDD^{+}24, HMK23, KLJ21, LLR^{+}21, LDH^{+}22, LYF^{+}23, NdSSSN20, Wan20]. Strategy [XHZ^{+}21, APM^{+}21, BF24, BZT23b, CH21, COZ21, CCZ^{+}21, CJC22, CWW^{+}23b, CZL^{+}22, Dah23, Den24, DCR23, DP22b, DQF^{+}23, FJCFB23, HAS^{+}22, HXZS23, IAQ20, JR24, KS22d, LGDW22, LFWJ22, LZZ^{+}20, LPZ^{+}22, LYG^{+}21, LZT^{+}23, MV24, NSPdO21, NA22c, OXBL23, PR24, PCK23, RF21, SLD24, SKSP23, SWS^{+}24, TDZ^{+}24, TAH22, WEH^{+}22, WJJ23, WJD22, XGCZ23, YSL^{+}21a, YJY^{+}21, Yil23, ZWCC23, ZDL^{+}22, Zhu22b]. Strategy-proof [XHZ^{+}21]. stratified [CS22]. stratum [PS22b]. stream [ACJ21a, ACJ21b, BA20, BCP^{+}23, DDB^{+}21, DS20a, ENB^{+}20, KRJS22, LLC^{+}22b, MB22, RSR^{+}22, SKH^{+}21, SWS^{+}24, VGDF22, YLZT23, ZHX^{+}23, SWS^{+}24]. streamed [SRG^{+}21]. streaming [BV22a, KR22c, LBA23, LLZ^{+}21a, LK22c, LLC^{+}21b, NLG^{+}20, PD20, TA23a, TWG^{+}21, WZY^{+}22, ZABT^{+}20, ZGWZ23]. Streamlining [WZB21]. Streams [CJP^{+}21, ARS22b, GlSL^{+}23, PS22a, SSSR20]. street [WNZ24]. strength [Kha22a, LLKS21, WBL22, ZWCS20]. stress [AV22b, SM23d, WBL22]. stress-strength [WBL22]. stretch [AP22]. Strictly [Ker22]. strider [PSK23]. string [KB22, MT22]. STRmt [SZZ^{+}24b]. stroke [BJK^{+}24, MMR^{+}22, VS21]. strong [CZCM23, ZRY^{+}22, ZYXX23]. Structural [DSS21, LSS24, Kha22a, RZ23, SGHL20, STH^{+}20]. Structure [DCWM20, Jeo20, AM20b, AQP^{+}22, FD20b, JW22, LSS^{+}21, LSSQ22, MS21a, PBD23, VCFZ20, YA22b, YOWY22, ZQW^{+}21, ZCH^{+}23]. structured [DWZ^{+}22, KBS^{+}22, MIN^{+}23, PMS^{+}21]. structured-based [KBS^{+}22]. structures [JSZS22]. student [CT22, CB22, FRS^{+}23, LGW^{+}22a, PVRM22, RM22, SRL23, ZHT^{+}23, uZKH^{+}20]. students [SV22e, TTZX22]. studies [KHK^{+}24, KCL^{+}20, ZCD^{+}22]. Studio [VZR^{+}21]. Study [HML19, xZlGCzJ20, Zha22, APP^{+}21, AK21, APA22, ASANR22, BANT20, BGPQ21, CCD^{+}20b, DM22c, ECIB20, FPHZ19, FSdP^{+}23, GKÇ22, GMSM21, GADM20, HPH^{+}20, HYT^{+}21, IP20, JK22a, JGJ^{+}21, KPM20, LMM23, LBFT22, LST22, LLW^{+}22a, LWCM21, LLA^{+}22, MPB^{+}22, MdAA^{+}21, OE22b, Pan20, PMS^{+}21, RZVC21, SZI^{+}23, SG22b, SSCN23, TTB^{+}22, TO22, WWW24, WHDS22, WWW^{+}23, XSZ^{+}23, ZBY22, ZDH^{+}22, ZWT22, uZKH^{+}20, FPHZ20]. Studying [HVT23]. style [SYRS^{+}22]. styled [HGNN22]. Stylized [LL23]. sub [KHHK21, KG23, LM21, SG21, VV23]. sub-events [VV23]. sub-harmonic [KG23]. subject [FJCFB23, ZGWZ23]. Submarine [ZABT^{+}20]. submodular [JXL^{+}23, ZGWZ23]. subpipeline [MS21a]. subscribe [NET20, XSGL20, YPO21, ZQX^{+}23]. subscription [ZABT^{+}20]. subscription-based [ZABT^{+}20]. subset [SM23d]. Subsets [SJ23b]. Subsets-based [SJ23b]. subsidized [BK22a]. subspace [LXT^{+}22]. subspaces [Eke22]. Substation [FHS^{+}22]. substitution [Yil21c]. substitution-based [Yil21c]. substrate [WZHL21]. substructure [ZRY^{+}22]. subtractive [MB22]. subtractor [GM22a]. SubtStream [MB22]. subtype [KMS^{+}22, ZZQ^{+}22]. subtypes [JSAA22]. subway [XSZ^{+}20]. success [CWDM^{+}21]. successful [KY23]. Sugeno [LSD21]. suicidal [SKCS23]. suitable [GZC^{+}21, yXlLyGX21]. sum [CC22b]. summarization [ACJ21b, CKL20, EA22, GSS^{+}23a]. summation [DS23b]. summer [BBD23]. sunflower [SEC22]. Super [Shi22a, LGDW22, PR22, SYT^{+}21, ZYX^{+}21]. Super-process [Shi22a]. super-resolution [SYT^{+}21, ZYX^{+}21]. super-resolutioned [PR22]. supercavitating [WWW24]. supercomputer [GHL^{+}23, LHPG21]. supercomputers [JOW23, MSPPD20, XZY^{+}22, XYY24]. SuperDome [RMS24]. supermodular [JXL^{+}23]. superpixels [ML20]. Supervised [CMK22, KBS^{+}22, AWS^{+}22, AEM22, LTLX22, SJ23b, WLL^{+}21b]. supervision [Liu21, ZML^{+}23]. supervisory [KNM21]. supplier [Faz23]. supply [FMNF22, LS23a, YYY^{+}23]. Support [KR22a, LJBS23, RG22, WSL^{+}20, AFK^{+}22, Ano21c, Ano21h, AYJ^{+}22, BS22, CK21a, CJP^{+}21, CMA^{+}21, DBD22, DP22a, DS20b, DCK21, EBLM22, FGJ^{+}21, GC20, Hem22, JMY21, JKS20, KPJ^{+}21, Kan22, KA22c, LZQF24, MTY21, PB22a, PKB^{+}23, RSR^{+}22, RRJ23, SS21, SAD^{+}21, SS22c, SR22c, SMAG22, SP21b, VPSM22, VNP^{+}23, WAY^{+}21, WJLC21, WLLZ20, Wu22, XCY22, YY22, ZJSJ20, ZJSJ21, IVP^{+}23]. support-vector [SS22c]. Supporting [MTRA23, Bai23, HKA20, MHPA21, WYL^{+}22, XZYW23, ZHXY23]. suppression [RRB22]. sure [LLO21]. surface [Aka22b, BZGM22, CLY^{+}21, GQ24, LLJ^{+}22b, NKST24, WLLX21, WZL^{+}22, WLX21]. surgery [Bai23, GCSZ24, TWW^{+}22]. surrogate [XW24]. surrogate-assisted [XW24]. surrounding [xZlGCzJ20]. surveillance [BRNR23, CLDY21, EKS20, JH21, JK22a, KR22b, NT23, SG22c, WMC^{+}23]. Survey [AKS^{+}22, FK20, XZ22, ZX23b, AMD20, BBB^{+}20b, ER22, GR22a, GKG^{+}20, HIEH22, HJT^{+}20, HZY^{+}21, HAR24, JK22b, JNS22, KSJ23a, KJMB22, LLJ^{+}20, Mit20, NNX^{+}23, ÖSTY22, SMD22b, SK22b, SKK21, SG24, VAB22, ZX21a, ZUTK23, SLL^{+}23]. Survival [RGM22, DBN^{+}22, PVRM22]. sustainability [FMNF22]. sustainable [AYH^{+}22, FSdP^{+}23, KDS^{+}20]. sustained [HPH^{+}20]. sustaining [SSM^{+}21]. SVD [AAQOT23, LYI^{+}20, SKS^{+}23]. SVM [SS21, AV22a, BDR23, FM22, TA22b]. swallow [Kot23]. swap [KB21]. Swarm [DSYF22, KSS21, PS23c, ZFH^{+}23, AS23a, ASL20, ASAAAA22, ABCP23, ASC22, ASMS21, AB21, BKK22, BKD22, BRS^{+}22, BMV22, CK23, CHMC21, CPCK23, DSSS22, DLC^{+}21, GNS22, GC20, HAR20, IAAA22, JAM24, Kot23, KWL^{+}24, KR23, KA22c, LG23, LGL^{+}21, LZC21, MT21, NM23, NSK^{+}24, ODK^{+}23, PJP21, RP22, SB23a, SA22a, SK22a, SBB21, SB23b, SDSW21, SKSP20, VS22b, VR22, VG21, WWL^{+}20, ZZY22a, ZX23a, ZZS^{+}21b]. swarm-based [SK22a]. swarming [ZSX21]. swarms [ADK^{+}23]. sweeping [ZCW^{+}23]. SWIR [LWZ^{+}22b]. switch [BJWY20, IQS^{+}22, YNK^{+}23]. switched [MHL^{+}20]. switches [HGX^{+}22]. switching [Ano21-38, LLO21, YLJ22]. SX [VMFL23]. SX-Aurora [VMFL23]. SYCL [GDGN23]. symbiotic [MNDK22, Pan23]. Symmetric [XZYW23, Ano21-38, HLS^{+}20, KHK^{+}24]. Symposium [BI23]. synapses [WYE23]. synchronization [BM22b, DE20, GCSZ24, KLH24, EGGG23, MAG^{+}20, NdMP22, SNGK21, ZQL^{+}21, ZYPD24]. synchronization-free [DE20]. synchronization-gossip [BM22b]. synchronous [CLE^{+}20, LZY^{+}20]. syncretic [CZG^{+}20]. syndrome [MRAM^{+}21]. synergic [LJC23]. Synergizing [TS24b]. syntactico [EGB21]. syntactico-semantic [EGB21]. Synthesis [APL^{+}21, GBB22, JNMG21, KWZ^{+}21, RK21b, RGM22, SYJL20]. synthesis-GPU [RGM22]. synthesized [LHK22]. synthetic [DBPC22, GDSS22, MRS^{+}21, SPHP21, XCX^{+}20]. synthetic-aperture [MRS^{+}21]. syscall [ST22d]. System [dCJBP20, KK22e, LGW^{+}22b, AYH^{+}22, Abb23, AA23a, AFF22, AKA^{+}22a, AV23, AA23d, AZA20, Ano21h, Ano21j, ARFA21, ALNJ21, BA20, BBF21, BZWH21, BKJ22, BM22c, BRNR23, BSPM23, BJS23, BKM^{+}24, BK22c, BBB22, CSL20, CK23, CC22b, CH21, CJY^{+}20, CLT^{+}21, CDR^{+}23, CK24b, CMA^{+}21, CDN^{+}21, DPFC20, DR22b, DS21, DBK21, DS20b, DD21, DCT^{+}23, EASN22, FPHZ19, FPÁ^{+}20, FDH22, FSWW21, GMM22, wGTC22, GLN23, GAS23, GRL^{+}22, GN21a, GCSZ24, GSS23b, GATK22, Gün23a, GM22c, GK22b, HT21, HNSS22, HLT23, HXZH21, HDS^{+}23, JPA^{+}23, JDLP22, JD22, JKP22, JZL21, JYC^{+}21, JZC^{+}23, KPJ^{+}21, KHEF22, KC20, Kan22, KB21, Ker22, KGE^{+}20, KCP23, KSTV21, KR22b, KDA^{+}22, KLK23, KCM^{+}22, KA22c, LA22, LW20a, LLX^{+}21, LLC21a, LH21, LZL^{+}22, LYC22, LM22b, LGL^{+}22, LSZ^{+}23, LMM^{+}22, LWW23, LSLY20, Liu21, LCW21, LCL^{+}22]. system [LYG^{+}21, LLYZ23, MR23b, MK22a, MV22, MSJ22, MNDK22, MES23, MKBB22, MHB^{+}22b, MTY21, MSB^{+}20, MMKA23, MLC^{+}21, NNVD22, NKY23, NS23, ND23, NKST24, NM24, NN23, NY22, NLG^{+}20, OBER22, PPA22, PS21, PM22, RSR^{+}22, RD23, RKC22, RT22c, SPA^{+}21, SJ22, SA22b, SMM22, SAQJ23, SLLMTS^{+}24, SEC22, SS22c, SS24, SZS20, SM23c, SP22b, SKK22, SR23, SJ23a, SR22c, SZL^{+}22, SK22d, SP22d, SS23g, SZZ^{+}24b, TAT^{+}23, TS21, TA22a, TBNPQ23, TP20, TLXX21, TH22, TPT^{+}22, TS22, TD21, TUD21, UE22, VS22b, VBP^{+}24, VKK^{+}24, VP22b, VPQ22, WHJ^{+}20, WGYZ22, WLG^{+}22, WZC^{+}22a, WQS^{+}23, WXLD21, WS21, XCJ22, XSZ^{+}20, XHM22, YKL23, YKW20, YL20a, YLLL20, YCY20, YÇC22, YPO21, YL20c, YWC^{+}23, YZXL22, ZWW^{+}21, ZMLW23, ZX21b, ZCD^{+}22, MT21, Pan20]. system-awareness [KGE^{+}20]. system-based [JKP22]. System-on-Chip [KK22e, CC22b]. Systematic [KSJ23a, KJ20, BKGC23, CBR22, DA22, GA23, Haj20, JZR22, KKAM21, KR22a, KKK21, KM23c, KHA22b, LLW^{+}22a, LCM22, NK22, SZI^{+}23, SRBH22, TSR22, VGDF22]. systematical [WZMJ23a]. Systems [CSW20, HFFA20, SKE22, AET^{+}22, AR23, ACJ21a, Ano21-43, ASA^{+}21, APL^{+}21, BT21, BEJD22, BSEN20, Ben22, BAR21, BSPM23, Boz22a, CdOO^{+}20, CYZ^{+}23, CWL^{+}20, CZTC22, CYW^{+}22, Cza23, DWZ^{+}22, DAT23, DSC^{+}21a, DJST24, DS24a, DS22b, DJ20, DE20, DHSG23, ELIG23, FD20a, GPR^{+}22, GS21, GKG^{+}20, Gün23a, HPH^{+}20, HMK23, UZAA21, HMB22, HPCK22, HXZS23, HXST22, KRKM22, Kat23, KGRG24, KA23, KAO22, KW21, KCL^{+}20, KSA^{+}21, KMB23, KR23, KST23, LBG^{+}20, LST22, LXC^{+}22b, LDCD22, LL23, LLO21, LGX^{+}23, LCM22, MBC23, MGS^{+}20, MYCH22, MTRA23, NET20, NKKM21, Ogi20b, ORP21, OAS23, PR24, PB22a, PRPD21, RS23, RSKA23, RMS24, RBWB21, SGS21a, SMD22b, SAD^{+}21, SVB23, Shi22a, SSI22, TS24b, TT23b, UPGCA22, VAB22, WHZL21, WJS21, WFY^{+}22a, WWA22, WGDP23, WSM^{+}24, Wri22, WWW^{+}23, WS21, XDH^{+}20, XYFZ23, YLM21]. systems [YSH^{+}22, YIB22b, YSS^{+}21, ZJ21, ZAB22, ZQW^{+}21, ZLTX21, ZX23a, ZBY^{+}23, ZALM23, ZQX^{+}23, ZUTK23, ZWZ^{+}22]. systems/leveling [KAO22].

T2FSM [DCR23]. table [VK24, NTK22, SBGC21]. table-centric [SBGC21]. tablets [MNYN21]. tabu [DD24, PD22b]. tabular [HD23]. tactics [LJC23, LSL20]. tag [Ano21h]. tagged [ML20]. tags [ZWO^{+}20]. Taguchi [RRIL22]. Tail [DGSB20]. Taiwanese [NTB23]. Takagi [LSD21]. Takagi-Sugeno [LSD21]. Take [WNZ24, WN23]. taken [TA22b]. takeout [PLX20]. Taking [WZ21, PLX20]. Tamil [SS23b]. Taming [WS21]. tampering [AK22b, CPA22]. Tanimoto [DT22a]. Tapis [CJP^{+}21]. tardiness [GMN21]. Target [LXJ^{+}22, BST^{+}22, CK21b, FHS^{+}22, FKY^{+}23, HAA23, KHK^{+}23, URK^{+}22, WHJ^{+}20, WXS^{+}23, YGS^{+}23]. target-based [CK21b]. Targeted [RT22a, AA23c]. TARNN [SM22d]. Task [DSC^{+}21a, DXXL20, DS22c, HAR20, LB21, MM21, NdSSSN20, QLL^{+}22, RCBTGE23, SM22d, AA21, ALBZ21, AB22a, BP24, CCCR21, CHI^{+}23, CWFK23, CLYG22, CMY21, DJST24, FSA^{+}22, FKO22, GBBS21, GWGR20, GSZ^{+}20, HLG24, JZC^{+}23, KRKM22, KKK21, KKS22, KV24, KT22b, LLX^{+}21, LK22c, LLH^{+}20, LS24, LALMGLG20, MHB22a, MSK22, MMS24, NAK^{+}22, NSK^{+}24, PK23b, PLZL24, RZCA21, RBWB21, SWCB20, TS24a, VG21, WWA22, WGDP23, YV23, YLZ^{+}21, ZLTX21, PB22a]. Task-allocation [LB21]. Task-aware [SM22d]. Task-based [MM21, NdSSSN20, RCBTGE23, CWFK23, CMY21, DJST24]. task-DAG [PB22a]. task-parallel [CHI^{+}23]. taskgraph [EK20]. tasks [Ano21g, CJC22, CHLD23, Cob22, Den24, DR20, FFG^{+}23, HK21, KE22, Ker22, LCL^{+}22, LDCD22, PB22a, SK20b, TS21, YS23, ZLW^{+}20]. Tatt [KXL^{+}21]. taxi [Kat23, LWCM21]. taxonomic [HLH^{+}20]. Taxonomy [GA23, KKK21, RSJ22, GQ21, HAR24, KK22a, MKS^{+}20, SJA^{+}22]. Taylor [BP23a, PKK23]. Tchebichef [AMBAJ22]. TCN [DTK24]. TCNN [Asl22]. TCNN-Bi-LSTM [Asl22]. TCP [AAN^{+}21]. TCSC [SYG22]. TDMA [KK22c]. teaching [Liq22, RSM21, RSM23, Shu22]. teaching-learning-based [RSM21, RSM23]. team [FGJ^{+}21, JR23a]. teams [LJC23]. technique [AR22a, AJK^{+}23, AADS21, AD22c, BV22b, BMV22, CDV23, CAAHC23, DDH^{+}20, EHST21, IAASK23, IDA22, JAR^{+}24, JM22, KNM21, KFKD22, KGGM22, KCP23, KR23, LA22, LWZC21, LNI^{+}24, LJB22, MSN22, MTT20, MRS^{+}21, Mon21, MSBR23, NA22a, PLP22, PR23, PS23c, RM23, RCK22, SS21, SA22a, SS23d, SS23e, SNGK21, SD22, SM22e, SS22g, SAGS24, TSG21, MKL21, ZHJ20]. Techniques [JB20, XZ20b, AAK^{+}21a, AKA22c, APA22, AOACAQ21, AS23b, Ano21c, ASA^{+}21, CO21, BGG24, BBD24, BFM^{+}21, CÇY22, CR23, CBK23, DA22, EUYY22, GCP22, GA23, GB23, HK22a, HJT^{+}20, HZY^{+}21, HPCK22, JKB22, JSYAA20, KK22a, KS23a, KJ20, KK23b, LK22b, MG23, NBS^{+}22, NAK^{+}22, NTB23, PMR^{+}21, PK22d, RDB22, SAQJ23, SS23d, SK20a, SET^{+}22, SRG^{+}21, TS21, TCW^{+}23, The21, TA23b, VY23, VB21, XZ20a, YMKH22, ZNDA22, ZLW^{+}20, dOdMC^{+}20]. Technologies [BZEM20, Liu24b, AA24, BTDD20, BP23b, CI20, HZZ^{+}23, MBM^{+}20b, RS24]. technology [AM20b, Ano21-43, CM20, JQ22, JGW20, Jia22, KLJ21, LS23a, LXZ22, MK22b, NY22]. Tegra [CYK^{+}21]. telecommunication [Ano21c]. telecommunications [BF22]. telehealth [ZAB22]. teleradiology [SN22b]. temperament [AAA22b]. temperature [OSK23, RS20, XLL^{+}21]. template [KTU^{+}21]. Temporal [MLZ^{+}20, VV23, MS22a, BM22c, CLWX21, KTK20, LHK22, LXKW23, RAK22, TSD23, XCD^{+}20, YW22, ZK24, ZLW^{+}23, ZXLD21]. Temporal-aware [MLZ^{+}20]. tenant [RAaB21, SAM^{+}23]. Tensor [LRC^{+}22, FQD^{+}23, XGX^{+}21, ZWY^{+}24, ZXLD21]. term [AKS^{+}22, ATC23, BKK23, CL23, CÇ22c, DR22c, DM22d, HR22a, IIK^{+}23, JR23b, KTM22, KH22, Kab22, Kar22, KDC22, LWLZ20, LLT21, PT22, PVRM22, SV22b, SMBM24, SVD^{+}22b, SHZY20, SMR23, SM23b, TLL20, XLZL22, YRV^{+}23, ZLCS21, ZCH^{+}23]. terminal [ZYY^{+}24]. terms [HC22]. ternary [FHH^{+}20]. terrain [Yil21b]. territory [JH21, PKVS21]. Tesia [LGZ^{+}22]. test [BFR24, MHPA21, MLZ21b, VNP^{+}23]. testing [DSS21, LHK22, LSLY20, LSS24, RRJ23, VBP^{+}24]. Text [MRKY22, TYA22, VS21, AMAT22, AMA22, CU22, CMK22, CPPP21, EA22, FEH22, GU23, HZY^{+}21, HGMK21, JLE22, KY22, KS22c, LWZ^{+}20, Par22, PKKL21, RSJ21, RSJ22, SAQJ23, ST23a, SM22b, TLL20, TWXL21, TSCM22, VSS23, WZ20a, XLZZ22, YHOY22, YÖT^{+}22, ZLCL21, ZLW^{+}20, vdSTC21]. text-to-image [AMAT22]. Textile [MWH^{+}23]. texts [BKK23, PP20a]. texture [AJS23, BCM22, CO21, GP22b, LGM21, LDS^{+}23, LS23b, PKSM24]. texture-based [LGM21, LS23b]. TF [SS22b]. TF-IDF [SS22b]. theater [LHWT20]. theft [AJAA21]. their [BFR24, HJT^{+}20]. theoretic [ACC^{+}20, JOW23]. theoretical [CWW^{+}23b]. Theory [MV22, Vin21, Cob22, DR21, FD20b, GKM22, HML21, K22, KGM23, LWCZ22, MBB22, MMS24, RNRK22, SRL23, ST22c, XCJ22, ZLCS21, ZOS^{+}21, TI22]. theory-based [Cob22, DR21, HML21, MBB22, ZOS^{+}21]. Thermal [GO22, HHP23, KSKS22, KCL^{+}20, STJ^{+}20]. theta [WWF^{+}22]. theta-join [WWF^{+}22]. thing [TWW^{+}22]. Things [CSW20, DKC^{+}21, GQ20, JPH23, JZR22, WJS21, AGSN23, CBK23, CÇM22b, DBB24, DLG24, IAASK23, JPAA21, KHEF22, LJ22, RAN21, RV23, UE22, WZMJ23a, YWC^{+}23, AYH^{+}22, Agr21, AS23a, AlS21, AADS21, AYB21, CM20, CYC21, CWW^{+}23b, CZTC22, DDH^{+}20, DPB22b, DKL21, DP22b, GMM22, GZC^{+}21, GXH^{+}21, HAR20, Hem22, HXZH21, JB21, JA23, JCG^{+}22, KH22, KB21, KK22c, KK22e, LGZ^{+}22, LYF^{+}23, MFE^{+}23, MK22b, Man21, MG23, NET21a, NET21b, NGXZ21, OO22, PPA22, PK23a, PZZ^{+}23, PJ21, PHDS22, RPMA22, RZCA21, RSMCP22, SSN22, SRRM23, SK22b, ST22c, ST23c, SS22e, SRBH22, TAH22, TT23b, VH22, VDL23, VRV23, VLVS22, VB22a, VPQ22, WGLL23, WKL^{+}22, XCY22, XHM22, YF20, ZCN22, ZZY^{+}22b, ZZK^{+}22, ZWT22, ZTP^{+}23, vdSTC21]. things-based [KHEF22, LJ22, DPB22b, DKL21, JA23]. Things-cloud [Hem22]. things-enabled [AGSN23, PPA22, VDL23]. thingsourcing [MFE^{+}23]. thinned [SYJL20]. third [SJ23a]. thoracic [SLL22]. thread [AÜ23, KHPH20]. thread-dense [KHPH20]. thread-to-core [AÜ23]. threaded [CHI^{+}23, DGSB20]. threading [LS22]. threat [CZTC22, KP21, LLC^{+}22b, OSMM23, QZB^{+}23, SRBH22, TT21]. threats [GA23, HJT^{+}20, PCG^{+}21, Tur23]. Three [IK22, SKS22, XWC^{+}22, ZHJ20, AM20b, GBBS21, GAMT22, KGRG24, KYP21, LSSQ22, LCW23b, SN22a, ZHJW21]. Three-class [IK22, SKS22]. Three-dimensional [ZHJ20, GBBS21, GAMT22, LCW23b, ZHJW21]. three-dimensional-cellular [KYP21]. three-factor [KGRG24]. three-input [AM20b, SN22a]. three-order [LSSQ22]. Three-stage [XWC^{+}22]. Threshold [CCM22a, CWW^{+}23b, DSS^{+}23, GZC^{+}21, JHZ20, JR23a, KYP21, LCZ^{+}20b, MSSM24, OS21a, SRGK24, SHBC20, SSDY20]. threshold-based [MSSM24]. threshold-moving [JHZ20]. thresholding [KK22f, SV22b]. thresholding-long [SV22b]. thresholds [CMK22]. Throttled [BM22a]. throughput [CTA^{+}23, KIN^{+}23, LGLZ20, MRAS^{+}23, SSSR20]. throwbox [CPCK23]. thru [SR22b]. thyroid [TZ22]. Tianzifang [WN23]. ticket [RT22b]. Tile [MTK^{+}21]. tiling [KW21, MT22]. Timbral [GSB21]. Time [CJP^{+}21, HK24, KMZ^{+}20, LYF22, LWHW22, NET21b, SLL^{+}23, WFY22b, WXCY20, ALBZ21, Ano21e, ARS22a, ACVK23, BM22a, BBD23, BS23a, BEJD22, BBP^{+}24, BVM22, BWS^{+}21, CD24a, CY22, CJ21b, CLX^{+}21, CJC22, ÇG21, DBD22, DWZ^{+}20b, EASN22, FJCFB23, FM20, GFPGT21, GCP22, GMK^{+}21, HSR23, HCH^{+}21, HHC^{+}22, HWG23, HGDD20, HLG24, IAO21, JS22b, JK22a, JAC^{+}21, KA23, Kay22a, Ker22, KLH24, KAR^{+}23, KSS24a, LZL^{+}20b, LLH^{+}20, LZL^{+}23, LLO21, PD20, PYC^{+}20, PMR^{+}21, RS23, RKuH^{+}20, RCS20, RZ24, SKA23, SKP22, SK21a, SMD^{+}21, Shi22b, SZJ21, SK20b, SK21b, SZZ^{+}24b, TS21, TSLL21, TYLY23, TTZX22, VKK^{+}24, WMC^{+}23, WSJ^{+}21, WS21, XLL^{+}23, XYFZ23, Yal22, YTW^{+}24, YÇC22, Yüc22, ZX21b, ZHT^{+}23, ZGH^{+}22]. time-aware [BVM22, FM20]. Time-based [HK24, NET21b, KSS24a]. Time-constrained [WFY22b, Ano21e]. Time-cost [LYF22]. Time-critical [KMZ^{+}20, Ker22]. time-delay [FJCFB23]. time-interleaved [ZX21b]. time-lock [BS23a]. time-of-use [KAR^{+}23]. time-sensitive [CJC22]. Time-series [CJP^{+}21]. Time-variant [WXCY20]. time-varying [DWZ^{+}20b, TSLL21]. times [BLB^{+}24, SWCB20]. timing [JCL^{+}20, MK22a]. Timo [ZLW^{+}23]. tiny [WXS^{+}23, ZWZ^{+}21]. tissue [TNP21]. TM [CMJC23]. TM-HOL [CMJC23]. TMFCC [SMM22]. to-go [DKA^{+}21]. TOALB [LS24]. tobacco [OSK23]. Tobit [ÖSTY22]. toddler [GATK22]. Toeplitz [DS22b]. together [HFFA20]. tokamak [DBPC22]. token [VCBB20]. tolerance [ACJ21a, AFNH21, BLT22, CLE^{+}20, FHH^{+}20, JAR^{+}24, SM23a, SKV22]. tolerant [AM20b, CPCK23, DR21, ITO21, Jia22, KMY24, MB21, MLZ^{+}20, QNZP22, VG20, ZLV22]. tomato [CÇ22a, KAAR23, MCT22b]. tomography [BS23c, DSSS22]. tongue [CWLL20, LLZ^{+}20, WLL^{+}21b, ZFZ^{+}20]. tool [AHT^{+}20, IAA20, KPJ^{+}21, KHK^{+}24, LSW21, MNDK22, OO22, SA22c, SI22, WKY22, LM20b, MCR^{+}23]. Tool-DEEP [LM20b]. toolbox [TSA21]. toolkit [KGW^{+}20, LP21, MC24]. Tools [LM20a, Bad23, CCPP21, IS23, KKK21, LP21, MNDK22, PUL20, SCKU23]. tooth [WLL^{+}21b]. tooth-mark [WLL^{+}21b]. Top [WSL^{+}20, TSCM22, ZLYS21]. top-down [TSCM22]. top-N [ZLYS21]. Topic [CMJC23, RA21, HBB20, MAR23b, SNET21, XSGL20]. topic-based [SNET21]. topic-centric [XSGL20]. topical [KXL^{+}21]. topological [GK23c]. Topology [AD22b, HZD^{+}22, Wan22a, WZHL21, CG22, HWG23, LB21, Wan20]. Topology-aware [WZHL21]. TOPSIS [HNG22, KT22a]. TOR [PCC^{+}21]. torque [MHL^{+}20]. TOSCA [DKA^{+}21]. touch [JPA^{+}23]. touchable [KPM20]. toughness [LYBZ23]. Tourism [Ano21f, UYÖ^{+}22, Wan24]. tourist [NWZ^{+}21]. toxin [KJS^{+}20]. trace [BEJD22, BJWY20, TK22a]. trace-based [BEJD22]. traceability [Ano21f]. traceable [LZL^{+}22]. traces [UMR23, YMZD21]. Tracing [DJST24, CZCM23, FD22, SSN21, WSM^{+}24]. tracking [CWL^{+}21, CWL^{+}22, ETKD23, GB20, GGS^{+}22, GATK22, JPK22, Pan24, RT22c, WLZ21a, ZLL^{+}22, ZTMC22, ZTMC23]. trade [JW21]. trade-off [JW21]. tradeoff [AAB23]. Trading [PSHJ20, AKA^{+}22a, KST23, LFX^{+}20, TCRP23]. traditional [CÇY22, DWDG20, JK22b, KIAA^{+}22, PD23]. Traffic [CLC^{+}20, QHL^{+}24, SMD22a, SKP23, DM22a, ZJSJ20, BS23b, BMJ^{+}20, CJ21a, CMJM22, DL23, GLN23, IAQ20, JK22a, JCL^{+}20, LBA23, LWCM21, MVR23, PK22c, SKP22, THT20, VPQ22, XLZL22, ZXL^{+}21a, ZJSJ21, ZJSJ21]. traffic-aware [IAQ20]. trained [KE22, ND23, Pan23, ZFF^{+}21]. training [AFK^{+}22, CLC^{+}23, Dah23, DCK21, FRS^{+}23, GRL^{+}22, GCSZ24, HXST22, JKKL21, NWX^{+}23, TWW^{+}22, TMB24, XLX^{+}21, YHL^{+}21]. traits [YB23a]. trajectories [LHK22, LWCM21, XCD^{+}20]. trajectory [ARHT21, CYZX23, KPF^{+}20, LXT^{+}22, TZ23, WMC21, ZTMC22, ZTMC23]. transaction [HXZH21, KR22c, PK22c, WYL^{+}22]. Transactional [LLP^{+}22, PKS22]. transactions [Ano21a, BAPS22, CY22, KSA22, RVF22]. transcriptional [LTL^{+}20]. transductive [DS20b]. Transfer [LPW^{+}21, LDH^{+}22, ÖTT23, SRG^{+}22, CÇY22, CXXW24, CYZ^{+}21, CON23, DSYF22, HHXH20, JS23, KE22, KBJ21, LLZ^{+}20, MBC23, MRK^{+}23, NKST24, PMP23, SS23e, SBGC21, SEC22, SKCS23, SYG22, UAS^{+}22, XPLX23, ZWL^{+}23]. transfer-learning [KE22]. Transferable [AAEA20]. transform [AS22d, AD22c, BEKS22, CCGN20, EKS20, FLG^{+}22, Gul22, JT23, JOW23, LGDW22, PSMM22, SO22, SA22d, VS21, ZSQ22, ACVK23]. Transformation [LLP^{+}22, AJK^{+}23, KKE^{+}22, KGE^{+}20, KSVP22, LPHK20, Yil23, ZSZ^{+}22]. transformations [DS20a, KHPH20]. Transformative [Ogi21, OLF21]. Transformer [YTW^{+}24, CXXW24, LCW^{+}23a, LC21, TYA22, Yil22b]. transformers [PLX^{+}24, SSP23]. transient [AFF22]. transit [DZLH20]. transition [SZZ^{+}24b]. Translation [TH22]. transmission [BZT^{+}23a, CL22a, SNET21, SM24a, SK22a, SGS21b, SEC22]. transmit [GMK^{+}21]. transonic [FZ23]. transparency [Ano21f, BTP^{+}21, MBO^{+}21]. Transparent [BK22a, BBB22, LCKJ21, CLYG22]. transplantation [ÇKÇ22]. transport [KA21b, UKS22]. transportation [Ano21-40, RSKA23, SVS22, TK22b]. transposed [KS21b]. travel [BMSD23, ZYXX23]. traveling [AAMM23, KWL^{+}24]. treatment [WLCW20]. treatment-oriented [WLCW20]. Tree [MB21, AB20a, BM22c, BBB22, DCK21, FRS^{+}23, FSdP^{+}23, HGM^{+}23, HWY^{+}23, IG24, JT23, JMV24, KMB23, MMR21, NWW^{+}22, PR24, MR23a, ST22a, SSS23a, SR22d, TS24a, TZ23, WWW^{+}20, YSH^{+}22, ZHW^{+}20, Zhu22a]. Tree-based [MB21, FRS^{+}23, MMR21, ST22a, TS24a]. tree-search [HGM^{+}23]. tree-shaped [KMB23]. trees [AD22c, ST23b, SNGK21, WL23]. trend [GD22, Kar22]. trend-based [GD22]. trends [ASA^{+}21, DJ20, HPCK22, PJ21]. trial [BBP^{+}24, INY^{+}23]. triangular [DE20]. triangulation [SAL22a]. trick [NMQ22]. tridiagonal [DS22b]. trigonometric [Bul22]. trip [DLY^{+}23]. trip-oriented [DLY^{+}23]. triplet [LHWT20, LLT21]. Trojan [CL22a]. truly [AB20b]. truncated [AAQOT23]. truncation [MKRK23, MRK^{+}23, SPSP23]. Trust [GB23, SPS23, AASPR22, AADS21, BKLY20, CYC21, GYZ^{+}20, GYL^{+}21, GZY^{+}22, JDG22, KA23, KLJ21, LSS^{+}21, LAH^{+}22, MKA23, RGKK21, SR22a, SYRP22, VH22, VD21, YY20a, YB23b, ZWW^{+}21, ZCN22]. trust-based [RGKK21, SR22a, VD21]. Trusted [Ano21-43, MA22, NGXZ21, HNSS22, HLT23, KUK^{+}24, LGZ^{+}22]. Trustful [AKA^{+}22a]. Trustworthy [DCM21, TWW^{+}22, HGHD22, ZAB22, ZOS^{+}21]. truthful [SZJ21]. TSCH [JCGT24]. TSet [WPK^{+}22]. TSS [TLX22]. Tucker [CB22]. tumbling [PS23a]. tumor [AJNS22, AMM^{+}20, BCK22, DBN^{+}22, IK22, KSJ23a, Kot23, KK22f, LJBS23, LLL^{+}24, ODK^{+}23, PS23b, SZqWZ20, SET^{+}22, SAMS23, VRR^{+}22, XLL^{+}20]. tunable [MLZ21b]. tune [DRM22b]. tuned [Dah23, ÖTT23, SBS22, YRV^{+}23]. Tunicate [AA23e, DSSS22, KV22a, NM23]. tuning [AML^{+}22, DCWM20, FJCFB23, JBBH21, KHK^{+}24, OHFF20, OAS23, QKSK23, RZVC21, VRB21, dOPBdO21]. Tunisian [TH22]. tunnel [XSZ^{+}23]. tuple [KMY24]. turbines [FDD^{+}24, XLL^{+}21]. Turkey [KA21b]. Turkish [AA24, EA22, KE21, Kab22, KY22]. turmeric [MCT22a]. Tversky [CB22]. tweets [Asl22, KIW^{+}22, MS22b]. twin [DCK21, LZQF24, MTY21, VPSM22, WNZ24]. Twister2 [KGW^{+}20, UGK^{+}22, WPK^{+}22]. Twitter [AMJK21, Aru22, HAA^{+}21, Kab22, KK22d, KG22, KJ20, RVJ^{+}22, VM23]. Two [KHK^{+}24, ÖSTY22, YZYT21, CD24a, CT21, CSWC20, CWH^{+}24, EAvM20, GMP^{+}20, HML20, HZQ^{+}24, ITO21, JS22a, KÖÖG22, LLH19, Mar23a, MAB22, MSA22, NKY23, RMMK23, SLLMTS^{+}24, SN22b, TÖK21, TLX22, MKL21, WLL21a, XYLW21, YYZS22, YTN^{+}20, ZMLW23]. two-bit [MKL21]. Two-dimensional [YZYT21, ITO21, NKY23]. two-layer [CSWC20]. two-level [SN22b]. Two-parameter [ÖSTY22, HML20, LLH19, TÖK21]. two-party [CT21, HZQ^{+}24, TLX22]. two-phase [EAvM20, XYLW21]. two-stage [CWH^{+}24, WLL21a, ZMLW23]. two-staged [RMMK23]. two-vector [JS22a]. twostep [ANAMSAR21]. type [AP22, AA22a, DCR23, HR22b, KSJK21, OAA22, SU23, ZBY22, ZJ21]. type-2 [DCR23, HR22b]. types [MCT22b, NKY23, OXBL23].

U [BDR23, BMG22, KK22b, KMD23, LJP^{+}21, XLL^{+}20]. U-Net [BDR23, BMG22, KK22b, KMD23, LJP^{+}21, XLL^{+}20]. UANET [XXSL23]. UAV [DSYF22, DTK24, LL24, QHL^{+}24, RC22]. UAV-assisted [QHL^{+}24]. UAV-MBN [RC22]. UAVs [SAGS24]. UBM [JYC^{+}21]. UChOA [RK23a]. UChOA-ANN [RK23a]. ulcer [DRM22a]. ultra [AM20b, KKKS23, RSKA23, Shi22a]. ultra-dense [AM20b]. ultra-large-scale [Shi22a]. ultra-low-dose [KKKS23]. ultrasound [YMKH22]. UML [CRB23]. Unauthorized [QZB^{+}23]. unbalanced [ZLW^{+}20]. unbiased [Cek22]. Uncertain [BSBF22, CDC20, GLW21, LDZ^{+}22, WQY^{+}22, YLZ^{+}21, YFF22]. Uncertainty [XYFZ23, GR22a, XCG^{+}22]. Uncertainty-aware [XYFZ23]. unclonable [NT23]. Unconstrained [GPDB20, KIN^{+}23, LNI^{+}24]. uncore [ADGT22]. Undefined [KTK20]. under-sampled [CNG^{+}20]. under-sampling [KSY24]. underground [DWZ^{+}20b, SRS23]. underlying [BJK^{+}24, RS21]. Undersampling [JYL^{+}23]. Understanding [SRS^{+}21, Ano21-42, JJZ^{+}21]. Underwater [GA23, ZLT21, CG22, PB22b, SA22a, SR20a, YZYT21]. undetectability [CCPP21]. undetectable [XXSL23]. undirected [HCH^{+}21, HHC^{+}22]. undulations [AKA22c]. Unequal [CLLB20, AP23]. Unequal-interval [CLLB20]. UNet [BCK22, WMS^{+}23]. uneven [SVB23]. unexpected [LFW20]. unfair [YB23a]. UNIConv [KK22b]. Unified [XLL^{+}23, CGW^{+}20, CYK^{+}21, PMC^{+}21, PVP^{+}20, YLZ20]. UniFlexView [YLZ20]. uniform [Kab23, ST23b]. unify [SPJI^{+}21]. Unigraphics [ZZL^{+}22]. UniIndex [CWL^{+}20]. unikernels [MK23]. union [ZHX^{+}23]. unique [CS23, SV22b]. unit [AS22c, Ano21-40, DBS^{+}22, FJCFB23, FRS^{+}23, HCG21, HZW^{+}23, HIN23, JPL22, KD22, MT22, OKJ^{+}21, SPC^{+}21, SCD^{+}23, TNP21, USI21, yXlLyGX21]. unit- [MT22]. unit-accelerated [HIN23]. unit-based [FJCFB23, MT22, USI21]. unit-recurrent [FRS^{+}23]. units [AAG^{+}22, GTM24, LZC21, Man21, NdMP22, PRS23, QPS20, SAF^{+}23, SRG^{+}21, SP22d, WGDP23]. units-based [LZC21]. univariate [SK21a]. universe [Yib22a]. university [DM22c, RM22]. unknown [RMR^{+}22]. Unmanned [WLLX21, BEKS22, ETKD23, FJCFB23, JYH^{+}24, KGK22, KS22b, LLX^{+}21, LGL^{+}22, LMM^{+}22, WZL^{+}22]. unnecessary [KGM23]. unobtrusive [dCJBP20]. unpaired [CXXW24]. unrelated [AE22b]. unrivaled [SR22b]. unsafe [SYRS^{+}22]. unsatisfiability [WRJ20]. unscented [JPK22]. unspecific [BWW^{+}20]. unstable [YLZ^{+}21]. unsteady [FZ23]. unstructured [AFG^{+}22, JLE22, OWB^{+}20, VSS23, ZS22]. Unsupervised [AS22b, XY21, YS22, PVP^{+}20, SK21a, WHJ^{+}20]. upcoming [CPH20]. update [MHPA21, SYL23]. updating [Lia24]. uploading [WBZ21]. upon [MSK22, SKHL22]. upsampling [KH22]. Urban [XCY22, XCD^{+}20, DZLH20, UZAA21, Liu21, MM24, MK22a, xZlGCzJ20, ZUTK23]. urbanization [LH21]. URL [GD22]. usability [KGK22, MGGA20, SGS21a]. Usage [GLM^{+}22, BBB^{+}20b, Boz22a, HZW^{+}23, PKKL21, WWL^{+}20]. Use [SLL^{+}23, Aka22b, DM22c, HUC^{+}22, KAR^{+}23, LM20b, PRS23, SRS^{+}21, WKY22]. used [CBR22, GKZ23, KE22, SN22b]. User [GSVS21, RT20, WGZ^{+}20, ZCD^{+}22, BWTJ20, CJ21b, EVVR21, EI22, GCSZ24, GK22b, GD22, IA23, KUK^{+}24, LBA23, LSS^{+}21, LWZY23, MG20, MG22, MS21b, MPLW24, NM23, OA22, SLJ23, SSSP21, SJ23b, SFJ^{+}21, TLQ21, TSA21, VDL23, VS22a, VG21, XZY^{+}22, YGZ^{+}21, ZWL^{+}20]. User-level [ZCD^{+}22]. user-profiling [MS21b]. users [FM20, JS22c, SVM24]. Using [Alm22, ASJ24, DE20, Eke22, KK23a, STH^{+}20, WLJ20, WKY22, AP22, AB22b, AB20a, AYH^{+}22, Abb23, AV22a, AK22a, AKA^{+}22a, ANP^{+}20, AS22a, AAN^{+}21, Aka22b, Akb22, ABB22, AS22c, AEMS24, AMV22, AMR^{+}21, AD22b, ABCP23, AAE23, AJNS22, AS23b, AB20b, AJS23, Ano21h, Ano21i, Ano21j, AS22d, AJAA21, AB21, AS20, AS23c, AR22b, AMM^{+}20, AHT^{+}20, AYJ^{+}22, AFM22, AA24, BCM22, BBA^{+}22, BMcKGK22, BLI20, CO21, BP23a, BBD23, BABS21, BEKS22, BD22, BKK22, BGG24, BMZ^{+}22, BKJ22, BF22, BABLH21, BSEN20, BM22b, BM22c, BD21, BMG22, BK22b, BK23, BCK22, BS23c, BVM22, BWS^{+}21, BBB^{+}20c, Bul22, ÇKÇ22, CK21a, CJ21b, CHMC21, CYZ^{+}23, CL22a, CR23, Cho20, CD22, CPCK23, CBK23, CPA22, DR21, DBS^{+}22, DÖD22, DR22b, DRV22, DBN^{+}22, DT22a, DK21, DKL21]. using [DA22, DPBP24, DP22b, DG21, DB23, DKK24, DS23b, DBB24, DTK24, DKH20, DS22c, DGP20, DM21, DJGF21, DPSJ22, ENB^{+}20, ESS23, ETKD23, EUYY22, EA22, EAvM20, FK24, FZ23, FFG^{+}23, FZA22, GDSS22, GPDB20, GS22, GDA22, GLW21, GLW22, GZC^{+}22, GM22a, GLN23, GA22, GDGN23, GDS23, GSTS22, GCP22, GP22a, GP22b, GVSS22, GM22b, Gul22, GSB21, GLA^{+}22, GBB22, Gün23a, GK22a, GK22b, GB23, GADM20, HK22a, HR22a, HK21, HAR20, HXY20, HYG^{+}23, HGNN22, HGM^{+}23, HK22b, HC22, HBB20, HE23, HTZ^{+}22, HGMK21, HHP23, HAAF22, HAA23, IAASK23, IAA20, IDA22, IIK^{+}23, IK22, JSS22, JDLP22, JH21, JS22c, JKB22, JS22b, JPL22, JSYAA20, JHG23, JA23, JSAA22, Jeo20, JAC^{+}21, JYW^{+}20, JYC^{+}21, JTY^{+}21, JNMG21, JR23a, JR23b, JKS20, JPO^{+}21, KDL20, KUK22, KH22, K22, KPJ^{+}21]. using [KHEF22, KKR23, KC20, KM21a, KSJK21, Kan22, KMR22, KKKS23, KNM21, KDC22, KK21b, KV22b, Kay22a, Kay22b, KV24, KSK^{+}22, KA22a, KHK^{+}23, KGGM22, KKC22, KEMZ22, KHHK21, KML21, KHK^{+}24, KLH24, KHY^{+}20, KA21b, KA22b, KMS^{+}22, KG22, KKS23, KK22e, KT22b, KJ20, KSP21, KPP^{+}22, KK22f, KDA^{+}22, KR22c, KS22d, KR23, KDS^{+}23, KSJ23b, KA22c, KBS^{+}22, KKM20, LA22, LMR22, LK22b, LG23, LGDW22, Lia24, LLKS21, LLW^{+}22b, LS23b, LM20b, LCL^{+}20, LPC^{+}21, LHL^{+}22, LXC^{+}22a, LXJ^{+}22, LLC^{+}22b, LM24, LLA^{+}22, LCZ^{+}20b, LF23, MRGP22, MNR^{+}22, MKRK23, MBM^{+}20b, MBB22, MS21a, MBC23, Man21, MSJ22, MNDK22, MNYN21, MAR23b, MES23, MTK^{+}21, MPV22, Mir22, MMS24, MS23, MG21b, MRKY22, Mon21, MH23, MCT22b, MV24, MSA22, MSBR23, NBK22, NSR22, NBPR22, NA22b]. using [NM24, NAK^{+}22, NSK^{+}24, NdMP22, NNJC23, NWZ^{+}21, NJ21, NJK22, NY22, ÖUG22, ODK^{+}23, OMA^{+}23, OE22a, PD20, PBK23, PPA22, Pal22, PSP22, PAN22, Pan23, PP21, PLP22, PSP24, PR22, MR23a, PLX20, PLZL24, PMR^{+}21, PCK23, PMP23, PVRM22, PB22b, PKR22, PA21, PD23, PK22d, AST22, RC22, RVF22, RR23a, RA21, RAN21, RSR^{+}22, RPPK23, RR23b, RF21, RM23, RMMK23, RK21a, RPMA22, RRIL22, RAaB21, RVJ^{+}22, RKuH^{+}20, RNRK22, RZCA21, RZ21, RG23, RKK23, RF23, RCS20, SR22a, SYRS^{+}22, SP23a, SPA^{+}21, SA22a, SA22b, SSMT22a, SSMT22b, SSSR20, SDR23, SK22a, SMD22a, SHA^{+}22, SGS21b, SKP22, SKP23, SS23d, SS23e, SS22a, SBA22, SV21, SV22b, SWK22, SPSP23, SCD^{+}23, SG21, SN22a, SEM^{+}20, SMD^{+}21, SM22b, SBB21, SAHAN22, SCP20, SRRM23, SS22c, SS24]. using [SX21, SYJL20, SM23c, SKS20, SKSB20, SYG22, SPS22, SPKK22, SGS^{+}23, SSP23, SKS^{+}23, SMKA22, SYRP22, SCM22, SM24b, SPHP21, SSP24, SKS22, SG22e, SP23b, SKSP23, SHVA23, SS23g, SPK22b, SSH22, SM22d, SS22g, SSS23b, SKÇA23, SBS22, SRS23, SKR^{+}24, DM22a, TI22, TTA20, TYA22, TB23, TA22a, The21, TT22, TO22, TAH22, TK24, TUD21, Tur23, TSD23, UWF^{+}21, UE22, VRR^{+}22, VH22, VY23, VPSM22, VPGK23, VLVS22, VRS22, VSS23, VB21, VR22, VPQ22, WMY^{+}21, WRJ20, WZZ^{+}20, WY20, WMC21, WWA22, WXS^{+}23, WWW^{+}20, WSM^{+}24, WSL^{+}20, WKB^{+}22, WTL23, WYLL24, XZXV21, XCJ22, XHZHXBQX22, XWD^{+}22, XGZ^{+}20, XPLX23, XR21, YAR22, YS22, Yal22, YZYT21, YHOY22, YB23b, YLW^{+}22, Yil22a, YÇC22, Yil22b, YMKH22, YY22, Yüc22, ZNDA22, ZHX^{+}23, ZR24, ZDH^{+}22, ZLT21, ZX23a, ZCZ^{+}22, ZHDW24]. using [ZWT22, ZOS^{+}21, ZZLZ22b, ZHW^{+}20, dOPBdO21, dRdSZ^{+}23, vdSTC21]. USVs [WZL^{+}22]. USVs-Sim [WZL^{+}22]. utilitarian [SM21]. Utility [KR22c, BVM22, CSV22, HZZ^{+}23, XSZ^{+}23]. utility-aware [CSV22]. utilization [AFBM^{+}23, AEM22, YTN^{+}20]. Utilizing [WT23, DS24b, Hem22, JR24, KK22f, SS22b]. UWB [CM20, ZXL^{+}21b].

V [VMFL23, XWQ^{+}23]. V-fused [XWQ^{+}23]. v3 [CJP^{+}21, SKÇA23]. vaccines [Asl22]. Vadis [MTSU22]. valence [ARS22b]. Validation [PBD^{+}21, HC22, TÖK21]. valuation [KBJ21]. value [FQD^{+}23, FCMM20, HGNN22, KDL20, KRSR23, KSP21, KTK20, WZZ^{+}22, WN21]. valued [FM20]. values [RRIL22, SK21a, SSS23a]. valve [Yil22a]. vampire [AP22, JDG22, JD23]. VANET [KDS22]. VANETS [AC23, GLN23]. Variable [MR23a, AP22, HMK24, HXST22, OO23b, AST22, WEH^{+}22, aWLY^{+}22, ZLTX21]. variables [SAHAN22]. variance [Cek22, NSSS22, YAR22]. variant [AAMM23, ASL20, FN23, Ker22, RSM21, WXCY20]. variation [CON23, SAHAN22]. variation-adapted [CON23]. variational [MLS^{+}24, NBK22]. Various [CZG^{+}20, BBD24, GQ21, NKY23]. varying [DWZ^{+}20b, TSLL21]. VCSNet [XWQ^{+}23]. Vector [IVP^{+}23, WSL^{+}20, AFK^{+}22, AML^{+}22, AAG^{+}22, AAQOT23, AYJ^{+}22, BS22, CK21a, DBD22, DP22a, DS20b, DCK21, GC20, GK22a, HCG21, Hem22, JS22a, JMY21, JKS20, KR22a, KS21b, LJBS23, LZQF24, RSR^{+}22, RRJ23, RG22, SS21, SS22c, SR22c, SP21b, VPSM22, VMFL23, WAY^{+}21, WJLC21, WLLZ20, XCY22, XWD^{+}22, YZ21, YY22, ZJSJ20, ZJSJ21]. vectorized [DS23b, GO22, SSS23a]. Vectorizing [JOW23]. vectors [SSP24]. vegetation [AJS23]. Vehicle [OO23b, AKS^{+}22, CMJM22, DSYF22, DM22c, ETKD23, FJCFB23, GZ20, HML21, JYH^{+}24, KAR^{+}23, LLX^{+}21, LGL^{+}22, LMM^{+}22, RSKA23, SWK22, SSRA23, TJZ23, TS24a, VPGK23, WWW24, Xia20, XL21, YZYT21, ZHJ20]. Vehicles [VBM^{+}21, BEKS22, DCZ^{+}22, KGK22, KS22b, KS23b, MNDK22, NM23, SSRA23, TZZ^{+}23, ZWC^{+}23, CYZX23, KGRG24, KJMB22]. Vehicular [TK22b, ABSS22, CD22, CPCK23, DKL21, FLB23, KGGM22, PDRM24, PA21, QLL^{+}22, SSW^{+}22]. Vein [ZLZ^{+}22a, DR22a]. ventricular [KNM21, LJB22]. venue [RKuH^{+}20]. veracity [AS23c]. verifiable [BPT^{+}23, HHYL22, KYP21, ZLO^{+}21, ZTP^{+}23]. verification [Ano21e, BEJD22, GN21a, UZAA21, DR22a, KQK^{+}20, WCWG20, YSK22]. verifier [PSP24]. verify [Ano21d, YR21]. verify-your-vote [Ano21d]. Verifying [FXX22, XZXV21]. veritable [OO22]. Vernam [DG21]. verse [ABCP23]. version [LM22b, NGB23]. versioning [YMZ^{+}20]. versus [AKA22c, KR22a, SGS21a]. Vertebrae [FMJ^{+}22]. vertex [DD24]. Vertical [Dev21, KRB^{+}20, KS23b, LZM^{+}23, SV21]. vertically [XZD^{+}21]. vertices [KSS24a]. Very [FSFM22, LAE^{+}22]. Vessel [DKC^{+}21, LS23b, RPPK23, WJL^{+}20, WLLX21]. vessels [WZL^{+}22]. VFLF [ZTP^{+}23]. via [Aka22b, AT22, CPQ^{+}22, CLX^{+}21, CÇ22c, DCK21, FZ23, GHRM21, GK23a, HYG^{+}23, HLH^{+}20, JHZ20, Kay22a, LXYY21, LLMZ21, LLC^{+}21b, MLZ^{+}21a, MRK^{+}23, ÖK22, QWL24, RR22, RBDS22, RBC20, SEMA^{+}22, SS23c, SSP22, WLZ21a, WKL^{+}22, XZG^{+}23, YYZS22, ZHX^{+}21, ZYZC23, ZWX21, ZXLD21]. vibration [YHL^{+}21]. Victim [HA21]. Video [RCK22, ZXLD21, Akb22, AS23b, Bai23, BRNR23, CC22b, EKS20, JT23, JB22a, KSP21, KR22b, LBA23, Mis22, RGPC23, SHL^{+}22, SSN23, TA23a, TJZ23, WMC^{+}23, YSP23]. videos [APM^{+}21, KSK^{+}22, SG22c]. view [CY22, DPFC20, Kab23, NNX^{+}23, RK21a, SZW^{+}22]. views [YLZ20]. Villa [WZ21]. viral [CÇ22a]. Virtual [ARC22, ACC^{+}23, WYL^{+}22, YLKC24, AWS^{+}22, ABC^{+}21, BRS^{+}22, BPM^{+}22, CWDM^{+}21, CCdCC21, DAdBM^{+}24, DS22c, FGZC23, GJK^{+}20, HAK22, HPS23, HAA23, KB21, KS20, KHA22b, LPZ^{+}22, LWCZ22, MR23b, MK23, MHPA21, NK22, PRS23, PS23c, RKL21, SA22a, SEM^{+}20, SM23a, SLC20, WZZ^{+}20, WGY20, WZHL21, WCZ^{+}23, XA22, YLJ22]. virtualization [EBLM22, IPRS21, RSV24, WYZAD20, Yu20]. virtualized [HTZ^{+}22, YLJ22]. virtue [LFW20]. virus [JS22b]. visibility [BTP^{+}21]. visible [SZG^{+}24, ZDJ^{+}21]. Vision [CMJM22, CWDM^{+}21, EASN22, HWY^{+}23, JKB22, JK22b, LCW^{+}23a, LCW21, LYG^{+}21, LG21, MCT22a, MSA21, MCT22b, PLX^{+}24, SKA23, TL21, YGS^{+}23]. Vision-based [CMJM22, JKB22, MSA21]. visionary [GGS^{+}22]. Visual [UYÖ^{+}22, MS22a, CLDY21, DAW22, JJZ^{+}21, Kum22, LH21, RKMS24, SBGC21, THW21, TL21, WLZ21a, WMS^{+}23, ZLL^{+}22]. Visualization [FD22, BSML21, Le23, MCD^{+}23, SC22b, SG22c, TL21, YA22a]. VLCC [FSFM22]. VLCC-Q [FSFM22]. VLSI [APA22, RK21b, SRGK24]. VM [VB24]. voice [EABZB21, IDA22, Kay22a]. VoIP [GMK^{+}21, Mon21]. volatile [AKZA22, KPA23]. voltage [FPHZ19, KK23a, Pan20]. volume [LP21, NBHN22, TK22a]. volumes [AAEA20, MBM^{+}20b]. voluminous [MRAS^{+}23]. volunteer [JR24]. volunteer-enabled [JR24]. Voronoi [DPBP24, SKR^{+}24]. vote [Ano21d]. VoterChoice [NRP^{+}20]. Voting [SDR20, NRP^{+}20, QWW^{+}22, ZLM22]. voting-attack-immune [ZLM22]. Voting-based [SDR20]. VR [GCSZ24]. vs [AKEC20, GGCGS20, HMK23]. VSLAM [Pan24]. Vulnerabilities [LGW^{+}22b, KSCL21]. Vulnerability [LYL20, MYCH22, AQP^{+}22, Gha20, SKK22, XSZ^{+}20]. vulture [KMR22].

wait [SWCB20]. wake [LSQW21, MTT20]. wake-up [LSQW21]. wakeup [HMK24]. walk [TT22]. wants [AYD21]. War [DCR23]. warehouse [JWT^{+}20]. warning [AC23]. Waste [AYH^{+}22, HHXH20, LKR^{+}22]. watching [QZB^{+}23]. Water [AJK^{+}23, AB22b, AFF22, CBR22, DP22b, DWZ^{+}20b, DWY^{+}21, HR22b, LG23, LM20b, NBPR22, PSK23, SRS23, WJLC21, WKY22]. watercolor [HIN23]. watermark [FLG^{+}22, Gul22]. watermarking [Gul22, KDL20, KKS23, KSK^{+}20, KSP21, RÖ22, SO22, SN22b, SKS^{+}23, SAL22a, TSG21]. Waters [BKM^{+}24]. wave [AB22b, AJK^{+}23, LML^{+}23]. wavelength [SA22b]. wavelet [AS22d, AS20, Gul22, JT23, KSP21, KSVP22, LMR22, RV23, SMD22a, SO22, WHZL21]. waves [RDB22]. way [CS23, NTB23]. ways [JHG23]. WBANs [PK22d]. Weak [KSD22]. weakly [HNOI24, WLL^{+}21b]. weapon [FZ23]. weapons [SJA^{+}22]. Wearable [BGPQ21, PYC^{+}20, WGYZ22]. wearing [HFH^{+}21]. Weather [BPAE20, Yal22, ARS22a, KCM^{+}22]. web [AV23, ASJ21, AQP^{+}22, AA24, BSBF22, CLLB20, CMA^{+}21, DPYS22, DAdBM^{+}24, GSS^{+}23a, GCP22, HGHD22, KQK^{+}20, KTU^{+}21, LmJdL^{+}22, MSN22, RV21, RBDS22, RDB22, SMR21, CPQ^{+}22, CHL^{+}24, FM20, GHRM21, HL20, KXL^{+}21, SM21, VSS23, YFF22, ZLCS21]. web-based [CMA^{+}21, SM21]. Weber [Fio20]. webpage [LZW^{+}21, YZX^{+}22]. websites [GB23, SA22c, ZLM22]. weed [RW24]. Weibull [WBL22]. Weight [OKJ^{+}21, AS23a, ANAMSAR21, BSN22, DWZ20a, DSC^{+}21b, DD24, Eke22, EBDB22, GC20, GSVS21, KHY^{+}20, LYF^{+}23, SDSW21, ZNLL22, DBD22]. weight-AdaBoost [DBD22]. weight-based [KHY^{+}20]. Weighted [BC22, LLR^{+}21, SSDY20, ZWZ^{+}21, DSS^{+}23, HNG22, JS22c, LJC23, LLW^{+}20, LLMX21, LWC^{+}21, LZQF24, Ona21, PLZL24, QWL24, SRL23, VH22, WZ20a, WHH^{+}20, YHOY22, ZZWZ21, ZCL21]. weighted-feature-based [WHH^{+}20]. weighting [TLL20]. weights [CCCR21, Dah23, SZZ^{+}24a]. Welding [CCGN20]. welfare [XCZ^{+}21]. Whale [CA23, DWZ20a, PD22b, ABESEh20, ABA22, DM22b, DS22a, JT23, JZC^{+}23, KS22d, RM23, ST22b, XLZL22, YZYT21]. wheel [CCZ^{+}21, HXY^{+}22a, TZZ^{+}23]. wheel-rail [TZZ^{+}23]. Wheeler [AD22c]. white [BABS21, BDR23, CÇY22]. Who [AYD21]. WHOG [EKS20]. Wi [MBM20a, MES23, NLG^{+}20, QZB^{+}23]. Wi-Fi [MES23, NLG^{+}20, QZB^{+}23]. Wi-Fi-based [MBM20a]. wide [BBD23]. widely [KE22]. widow [JR23b]. WiFi [TWQ^{+}21]. WiFi-aware [TWQ^{+}21]. wikis [IAO21]. Wild [JPO^{+}21, SHL^{+}22]. wildlife [RT22c]. will [KLJ21]. Wind [ZZLZ21, FDD^{+}24, KWL^{+}24, NBK22, XLL^{+}21]. window [KR22c, NY22]. Windows [LGW^{+}22b, RF23]. WinDRAM [KSD22]. Wireless [SEMA^{+}22, WXCY20, AP22, AM20a, ASC22, ASMS21, AGSN23, ASRN23, BK21, BLT22, BA24, BM22b, CSWZ22, CLC^{+}23, CL22b, CG22, DRV22, DK21, DPSJ22, ETKD23, FCX^{+}22, GJBM22, GSS23b, GNS22, HSL^{+}22, IA22, JH21, JB22b, JS22a, JCGT24, K22, KDS^{+}20, KGM23, KK22c, LCSR21, LPW^{+}21, LL21, Liq22, LCW23b, LAC21, LLL^{+}21, MES23, MKA23, MS21c, MG23, MP23, NVV24, NAR^{+}22, NKST24, Ngu21, PPA22, PB22b, PM23b, RJ22, RK23c, RP22, RT22b, SDR23, SGS21b, SV22b, SPS22, SS23f, SM24b, SP23b, SPK^{+}22a, SKR^{+}24, TA23a, TTM^{+}22, TKG^{+}23, URK^{+}22, WYZAD20, WJD22, WYLL24, YLGY20]. wisdom [TTZX22, ZHT^{+}23]. wise [Gün23a]. with/without [NR23]. within [CZTC22, LHWT20, MGS^{+}20, PCG^{+}21, SRIB23]. without [MNDK22, NR23, SK22c]. witnesses [ZOS^{+}21]. WM INPUT [LYL20]. Wolf [BJGF20, LSL20, AS23a, ASC22, AB21, CZ24, DS20b, DPSJ22, GP22a, JYH^{+}24, KV24, KGGM22, MCT22b, NNVD22, PVRM22, SFJ^{+}21, TB23, TS22, UDS21]. wolf-based [ASC22]. wolf-student [PVRM22]. wood [SAMS23]. woodpecker [GVSS22]. word [HLH^{+}20, JKKL21, NNJC23, Ona21, RR22, RP21b, YT21]. Word2Sent [KBBH21]. words [HV21, HH23]. work [APL^{+}21, LXT^{+}22, JB21]. Workflow [AB21, KSB23, RSM21, SAL^{+}22b, Wan22b, AA21, ANAMSAR21, BZL^{+}22, CDN^{+}21, GCQ^{+}24, GSVS21, KSSK22, LYF22, MSS22, MSK22, NWW^{+}22, QCC^{+}23, RSM22, RSM23, SKHL22, SK21b, TTB^{+}22, WCSG20, XCG^{+}22]. workflow-based [CDN^{+}21]. workflows [BMK^{+}20, BRL^{+}20, CCCR21, CHMC21, DPdS^{+}23, GSZ^{+}20, GR22b, KVV20, KMB23, NGOS22, NSPdO21, SB23a, XYFZ23, dOPBdO21, Wu22]. workgroup [KW21]. working [AV22b, SM23d]. workload [AFKS23, CMT20, HOS^{+}21, PKM21, SLC20, UMR23, YJY^{+}20]. workloads [CYK^{+}21, IPRS21, JSP20a, JSP20b, MK23, PMP23, RGPC23, TKS22, ZHWY22]. workplaces [HL23]. Workshop [LM20a, Bad23, IS23, LXZ20, SBG20]. world [SYT^{+}21, VP22a, WLJ20, XLWX20, BBBC22]. worldwide [JS22b]. wormhole [NM20]. Wrapper [SSMT22b, ÖUG22, SM22a]. wrapper-based [ÖUG22, SM22a]. Wrist [LLJ^{+}22b]. WSANs [ECIB20]. WSN [CSR24, DS22a, GJPS23, HLS^{+}20, LCW21, PCR21, PKR22, Sin23, TK24]. WSNs [AP23, JAR^{+}24, SS23d]. WSO [DCR23]. WSO-T2FSM [DCR23]. WVSN [LGM21].

X [Boz22b, GP22b, KCL^{+}20, KPP^{+}22, LLM^{+}24, MRGP22, WSM^{+}24, ZZL^{+}22]. X-ray [Boz22b, GP22b, KPP^{+}22, LLM^{+}24, WSM^{+}24, ZZL^{+}22]. X-rays [KCL^{+}20, MRGP22]. XC [KMRR20, MGS^{+}20]. XC40 [HPH^{+}20, SSSR20]. Xception [TTZX22]. Xeon [Mit20, HYT^{+}21]. Xevolver [KGE^{+}20]. XFEL [BBP^{+}24]. XGB [GRG^{+}22]. XGBoost [GRG^{+}22]. XII [Çet23]. XLM [Anb22]. XOR [SAGS24].

Yang [ADA22]. Yang-Chang [ADA22]. Yangtze [ZK24]. yard [gZWfL^{+}20]. year [PJ21]. yellow [MCT22a]. YHGSM [JGW20]. yield [DM22d, JR23b, SHA^{+}22]. YOLACT [ZTS^{+}22]. YOLO [FAM22]. Yu [WNZ24].

zero [AAA^{+}22a, BMZ^{+}22, CZTC22, FLG^{+}22, LAH^{+}22]. zero-day [BMZ^{+}22, CZTC22]. zero-inflated [AAA^{+}22a]. zero-trust [LAH^{+}22]. Zhejiang [LH21]. Zielonka [DMS^{+}21]. zone [SKS20].

References

Ahmad:2021:ELS

[AA21] Wakar Ahmad and Bashir Alam. An efficient list scheduling algorithm with task duplication for scientific big data workflow in heterogeneous computing environments. Concurrency and Computation: Practice and Experience, 33(5):e5987:1– e5987:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Algamal:2022:DLT

[AA22a] Zakariya Yahya Algamal and Mohamed R. Abonazel. Developing a Liu-type estimator in beta regression model. Concurrency and Computation: Practice and Experience, 34(5): e6685:1–e6685:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Altay:2022:CNB

[AA22b] Elif Varol Altay and Bilal Alatas. Chaos numbers based a new representation scheme for evolutionary computation: Applications in evolutionary association rule mining. Concurrency and Computation: Practice and Experience, 34(5):e6744:1– e6744:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abbas:2023:SSS

[AA23a] Shatha O. Abbas and Mohammed J. F. Alenazi. SSHS: SDN seamless handover system among LAN access points. Concurrency and Computation: Practice and Experience, 35 (23):e7821:1–e7821:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Almasri:2023:NCA

[AA23b] Marwah Mohammad Almasri and Abrar M. Alajlan. A novelcascaded ANFIS-based deep reinforcement learning for the detection of attack in cloud IoT-based smart city applications. Concurrency and Computation: Practice and Experience, 35 (22):e7738:1–e7738:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alrumaih:2023:EIN

[AA23c] Thuraya N. I. Alrumaih and Mohammed J. F. Alenazi. Evaluation of industrial network robustness against targeted attacks. Concurrency and Computation: Practice and Experience, 35(27):e7855:1–e7855:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alzahrani:2023:MIM

[AA23d] Abdulsalam O. Alzahrani and Mohammed J. F. Alenazi. ML-IDSDN: Machine learning based intrusion detection system for software-defined network. Concurrency and Computation: Practice and Experience, 35(1):e7438:1–e7438:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arumugasamy:2023:THG

[AA23e] Muthulakshmi Arumugasamy and Antonidoss Arokiasamy. Tunicate Henry gas solubility optimization-based deep residual network for fruit ripeness classification. Concurrency and Computation: Practice and Experience, 35(2):e7490:1– e7490:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aydin:2024:MBD

[AA24] Sahin Aydin and Dieaa Aldara. Microservices-based databank for Turkish hazelnut cultivars using IoT and semantic web technologies. Concurrency and Computation: Practice and Experience, 36(12):e8062:1–e8062:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Akram:2022:NSE

[AAA^{+}22a] Muhammad Nauman Akram, Mohamed R. Abonazel, Muhammad Amin, B. M. Golam Kibria, and Nimra Afzal. A new Stein estimator for the zero-inflated negative binomial regression model. Concurrency and Computation: Practice and Experience, 34(19):e7045:1–e7045:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Altun:2022:ADL

[AAA22b] Sinan Altun, Ahmet Alkan, and Hatice Altun. Application of deep learning and classical machine learning methods in the diagnosis of attention deficit hyperactivity disorder according to temperament features. Concurrency and Computation: Practice and Experience, 34(13):e6908:1–e6908:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Almiani:2020:CHI

[AAARR20] Muder Almiani, Alia AbuGhazleh, Amer Al-Rahayfeh, and Abdul Razaque. Cascaded hybrid intrusion detection model based on SOM and RBF neural networks. Concurrency and Computation: Practice and Experience, 32(21):e5233:1– e5233:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Altuntas:2023:AEG

[AAB23] Zuhal Altuntas, Sanem Arslan, and Betul Boz. Approximate execution and grouping of critical sections for performanceaccuracy tradeoff. Concurrency and Computation: Practice and Experience, 35(24):e7614:1–e7614:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

AhlertPinno:2020:CNS

[AAD20] Otto Julio Ahlert Pinno, André Ricardo Abed Grégio, and Luis C. E. De Bona. ControlChain: a new stage on the IoT access control authorization. Concurrency and Computation: Practice and Experience, 32(12):e5238:1–e5238:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anisetti:2021:TAT

[AADS21] Marco Anisetti, Claudio Agostino Ardagna, Ernesto Damiani, and Alessandro Sala. A trust assurance technique for Internet of Things based on human behavior compliance. Concurrency and Computation: Practice and Experience, 33(4):e5355:1– e5355:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Allegue:2023:SSB

[AAE23] Sahar Allegue, Takoua Abdellatif, and Houssem El Abed. SBM: a Smart Budget Manager in banking using machine learning, NLP, and NLU. Concurrency and Computation: Practice and Experience, 35(11):e6673:1–e6673:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

AlZubi:2020:THP

[AAEA20] Shadi AlZu’bi, Sokyna AlQatawneh, Mohammad ElBes, and Mohammad Alsmirat. Transferable HMM probability matrices in multi-orientation geometric medical volumes segmentation. Concurrency and Computation: Practice and Experience, 32(21):e5214:1–e5214:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aliaga:2022:CLB

[AAG^{+}22] José I. Aliaga, Hartwig Anzt, Thomas Grützmacher, Enrique S. Quintana-Ortí, and Andrés E. Tomás. Compression and load balancing for efficient sparse matrix-vector product on multicore processors and graphics processing units. Concurrency and Computation: Practice and Experience, 34(14): e6515:1–e6515:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aggarwal:2021:PCA

[AAK^{+}21a] Akarsh Aggarwal, Mohammed Alshehri, Manoj Kumar, Purushottam Sharma, Osama Alfarraj, and Vikas Deep. Principal component analysis, hidden Markov model, and artificial neural network inspired techniques to recognize faces. Concurrency and Computation: Practice and Experience, 33(9): e6157:1–e6157:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ahmed:2021:RHR

[AAK^{+}21b] Usman Ahmed, Muhammad Aleem, Yasir Noman Khalid, Muhammad Arshad Islam, and Muhammad Azhar Iqbal. RALB-HC: a resource-aware load balancer for heterogeneous cluster. Concurrency and Computation: Practice and Experience, 33(14):e5606:1–e5606:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amin:2021:NAL

[AAK21c] Muhammad Amin, Muhammad Nauman Akram, and B. M. Golam Kibria. A new adjusted Liu estimator for the Poisson regression model. Concurrency and Computation: Practice and Experience, 33(20):e6340:1–e6340:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abayomi-Alli:2022:DLM

[AAMAA22] Olusola Abayomi-Alli, Sanjay Misra, and Adebayo Abayomi-Alli. A deep learning method for automatic SMS spam classification: Performance of learning algorithms on indigenous dataset. Concurrency and Computation: Practice and Experience, 34(17):e6989:1–e6989:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdollahi:2023:FFM

[AAMM23] Sahar Abdollahi, Hamid Asadi, Ahmadreza Montazerolghaem, and Sayyed Majid Mazinani. FMap: a fuzzy map for scheduling elephant flows through jumping traveling salesman problem variant toward software-defined networking-based data center networks. Concurrency and Computation: Practice and Experience, 35(26):e7841:1–e7841:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ahmad:2021:AAR

[AAN^{+}21] Mudassar Ahmad, Usman Ahmad, Md Asri Ngadi, Muhammad Asif Habib, C. M. Nadeem Faisal, and Nasir Mahmood. ARFC: Advance response function of TCP CUBIC for IoTbased applications using big data. Concurrency and Computation: Practice and Experience, 33(4):e4927:1–e4927:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aliaga:2023:SMV

[AAQOT23] José I. Aliaga, Hartwig Anzt, Enrique S. Quintana-Ortí, and Andrés E. Tomás. Sparse matrix–vector and matrix– multivector products for the truncated SVD on graphics processors. Concurrency and Computation: Practice and Experience, 35(28):e7871:1–e7871:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ahmed:2022:QAT

[AASPR22] Usama Ahmed, Asma Al-Saidi, Ioan Petri, and Omer F. Rana. QoS-aware trust establishment for cloud federation. Concurrency and Computation: Practice and Experience, 34 (3):e6598:1–e6598:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdulkadir:2021:SGS

[AAT21] Abdulsalam Ahmed Abdulkadir and Fadi Al-Turjman. Smartgrid and solar energy harvesting in the IoT era: an overview. Concurrency and Computation: Practice and Experience, 33 (4):e4896:1–e4896:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aali:2020:DLS

[AB20a] Sahar Nikbakht Aali and Nader Bagherzadeh. Divisible load scheduling of image processing applications on the heterogeneous star and tree networks using a new genetic algorithm. Concurrency and Computation: Practice and Experience, 32 (10):e5498:1–e5498:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alzahrani:2020:NPA

[AB20b] Naif Alzahrani and Nirupama Bulusu. A new product anticounterfeiting blockchain using a truly decentralized dynamic consensus protocol. Concurrency and Computation: Practice and Experience, 32(12):e5232:1–e5232:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arora:2021:WSU

[AB21] Neeraj Arora and Rohitash K. Banyal. Workflow scheduling using particle swarm optimization and gray wolf optimization algorithm in cloud computing. Concurrency and Computation: Practice and Experience, 33(16):e6281:1–e6281:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aktan:2022:MTS

[AB22a] Merve Nur Aktan and Hasan Bulut. Metaheuristic task scheduling algorithms for cloud computing environments. Concurrency and Computation: Practice and Experience, 34 (9):e6513:1–e6513:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

AV:2022:LBC

[AB22b] Arulkumar V and N. Bhalaji. Load balancing in cloud computing using water wave algorithm. Concurrency and Computation: Practice and Experience, 34(8):e5492:1–e5492:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alshayeji:2022:SBP

[ABA22] Mohammad Alshayeji, Bader Behbehani, and Imtiaz Ahmad. Spark-based parallel processing whale optimization algorithm. Concurrency and Computation: Practice and Experience, 34(4):e6607:1–e6607:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Akkasaligar:2022:MSM

[ABB22] Prema T. Akkasaligar, Sumangala Biradar, and Sunanda Biradar. Multilevel security for medical image using heterogeneous chaotic map and deoxyribonucleic acid sequence operations. Concurrency and Computation: Practice and Experience, 34(24):e7222:1–e7222:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abbas:2023:IMI

[Abb23] Qaisar Abbas. An intelligent medical image classification system using few-shot learning. Concurrency and Computation: Practice and Experience, 35(2):e7451:1–e7451:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Assante:2021:RVR

[ABC^{+}21] Massimiliano Assante, Alice Boizet, Leonardo Candela, Donatella Castelli, Roberto Cirillo, Gianpaolo Coro, Enol Fernández, Matthias Filter, Luca Frosini, Teodor Georgiev, George Kakaletris, Panagis Katsivelis, Rob Knapen, Lucio Lelii, Rob M. Lokers, Francesco Mangiacrapa, Nikos Manouselis, Pasquale Pagano, Giancarlo Panichi, Lyubomir Penev, and Fabio Sinibaldi. Realizing virtual research environments for the agri-food community: The AGINFRA PLUS experience. Concurrency and Computation: Practice and Experience, 33(19):e6087:1–e6087:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Allada:2023:ADC

[ABCP23] Apparna Allada, Rajaram Bhavani, Kavitha Chaduvula, and Rajaram Priya. Alzheimer’s disease classification using competitive swarm multi-verse optimizer-based deep neuro-fuzzy network. Concurrency and Computation: Practice and Experience, 35(21):e7696:1–e7696:??, September 25, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdel-Basset:2020:MHW

[ABESEh20] Mohamed Abdel-Basset, Doaa El-Shahat, and Ibrahim Elhenawy. A modified hybrid whale optimization algorithm for the scheduling problem in multimedia data objects. Concurrency and Computation: Practice and Experience, 32(4): e5137:1–e5137:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Balasmeh:2022:FDP

[ABSS22] Hani Al-Balasmeh, Maninder Singh, and Raman Singh. Framework of data privacy preservation and location obfuscation in vehicular cloud networks. Concurrency and Computation: Practice and Experience, 34(5):e6682:1–e6682:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdel-Basset:2020:SIR

[ABZS20] Mohamed Abdel-Basset, Yongquan Zhou, and Florentin Smarandache. Special issue on recent advances in intelligent algorithms and its applications. Concurrency and Computation: Practice and Experience, 32(4):e5417:1–e5417:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aarthi:2022:DRN

[AC22] B. Aarthi and Balika J. Chelliah. Deep recurrent neural network-based Aquila optimization-based online shaming emotion analysis. Concurrency and Computation: Practice and Experience, 34(11):e6882:1–e6882:??, May 15, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arunkumar:2023:NNL

[AC23] R. S. Arunkumar and A. Chinnasamy. NLOS node localization for improving warning message distribution in VANETS. Concurrency and Computation: Practice and Experience, 35 (23):e7816:1–e7816:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anglano:2020:PAC

[ACC^{+}20] Cosimo Anglano, Massimo Canonico, Paolo Castagno, Marco Guazzone, and Matteo Sereno. Profit-aware coalition formation in fog computing providers: a game-theoretic approach. Concurrency and Computation: Practice and Experience, 32(21):e5220:1–e5220:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Assante:2023:VRE

[ACC^{+}23] Massimiliano Assante, Leonardo Candela, Donatella Castelli, Roberto Cirillo, Gianpaolo Coro, Andrea Dell’Amico, Luca Frosini, Lucio Lelii, Marco Lettere, Francesco Mangiacrapa, Pasquale Pagano, Giancarlo Panichi, Tommaso Piccioli, and Fabio Sinibaldi. Virtual research environments co-creation: The D4Science experience. Concurrency and Computation: Practice and Experience, 35(18):e6925:1–e6925:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Akber:2021:FFA

[ACJ21a] Syed Muhammad Abrar Akber, Hanhua Chen, and Hai Jin. FATM: a failure-aware adaptive fault tolerance model for distributed stream processing systems. Concurrency and Computation: Practice and Experience, 33(10):e6167:1–e6167:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aslam:2021:PFB

[ACJ21b] Adeel Aslam, Hanhua Chen, and Hai Jin. Pre-filtering based summarization for data partitioning in distributed stream processing. Concurrency and Computation: Practice and Experience, 33(20):e6338:1–e6338:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aseeri:2023:SPS

[ACVK23] Samar A. Aseeri, Anando Gopal Chatterjee, Mahendra K. Verma, and David E. Keyes. A scheduling policy to save 10% of communication time in parallel Fast Fourier Transform. Concurrency and Computation: Practice and Experience, 35 (15):e6508:1–e6508:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abonazel:2022:DRR

[AD22a] Mohamed R. Abonazel and Issam Dawoud. Developing robust ridge estimators for Poisson regression model. Concurrency and Computation: Practice and Experience, 34(15):e6979:1– e6979:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ali:2022:TBE

[AD22b] Farithkhan Abbas Ali and Kanmani Ruby Erode Dhanapal. Topology based energy efficient routing using integration of fuzzy based Markov chain cluster-optimized novel ant bee colony approach in FANET. Concurrency and Computation: Practice and Experience, 34(23):e7175:1–e7175:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arunpandian:2022:EIC

[AD22c] Sekar Arunpandian and Subbaiah S. Dhenakaran. An effective image compression technique based on Burrows–Wheeler transform with set partitioning in hierarchical trees. Concurrency and Computation: Practice and Experience, 34(5): e6705:1–e6705:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Awwad:2022:DRO

[ADA22] Fuad A. Awwad, Issam Dawoud, and Mohamed R. Abonazel. Development of robust Özkale–Kaçiranlar and Yang-Chang estimators for regression models in the presence of multicollinearity and outliers. Concurrency and Computation: Practice and Experience, 34(6):e6779:1–e6779:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Andre:2022:PDD

[ADGT22] Étienne André, Rémi Dulong, Amina Guermouche, and François Trahay. duf: Dynamic uncore frequency scaling to reduce power consumption. Concurrency and Computation: Practice and Experience, 34(3):e6580:1–e6580:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ashkenazi:2023:FFS

[ADK^{+}23] Yotam Ashkenazi, Shlomi Dolev, Sayaka Kamei, Fukuhito Ooshita, and Koichi Wada. Forgive and forget: Selfstabilizing swarms in spite of Byzantine robots. Concurrency and Computation: Practice and Experience, 35(11):e6123:1– e6123:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abhishek:2022:CIG

[AE22a] Kunal Abhishek and George Dharma Prakash Raj E. Computational investment in generation of elliptic curves randomly over large prime fields. Concurrency and Computation: Practice and Experience, 34(13):e6906:1–e6906:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agushaka:2022:IPD

[AE22b] Jeffrey O. Agushaka and Absalom E. Ezugwu. Influence of probability distribution initialization methods on the performance of advanced arithmetic optimization algorithm with application to unrelated parallel machine scheduling problem. Concurrency and Computation: Practice and Experience, 34 (22):e6871:1–e6871:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ata:2022:RFF

[AEM22] Mohamed Maher Ata, Khaled Mohammed Elgamily, and Mohamed A. Mohamed. Robust features fusion utilization for supervised palmprint recognition. Concurrency and Computation: Practice and Experience, 34(10):e6817:1–e6817:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alavizadeh:2024:DRC

[AEMS24] Alavieh Sadat Alavizadeh, Seyed Hossein Erfani, Meghdad Mirabi, and Amir Sahafi. A distributed reliable collusion-free algorithm for selecting multiple coordinators in IOTA using fog computing. Concurrency and Computation: Practice and Experience, 36(11):e8009:1–e8009:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdennebi:2022:MLB

[AET^{+}22] Anes Abdennebi, Anil Elakas, Fatih Tasyaran, Erdinç Öztürk, Kamer Kaya, and Sinan Yildirim. Machine learning-based load distribution and balancing in heterogeneous database management systems. Concurrency and Computation: Practice and Experience, 34(4):e6641:1–e6641:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Fadheeli:2023:PBE

[AFBM^{+}23] Mohammed Saud Ali Al-Fadheeli, Yassine Bouteraa, Adil Hussein Mohammed, Hayder Mahmood Salman, and Ghadir Ghasemi. Predicting building energy utilization with energy plus simulation and advanced sea lion optimization algorithm based on Elman neural network. Concurrency and Computation: Practice and Experience, 35(10):e7645:1–e7645:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdelhafidh:2022:GAB

[AFF22] Maroua Abdelhafidh, Mohamed Fourati, and Lamia Chaari Fourati. A genetic algorithm-based intelligent solution for water pipeline monitoring system in a transient state. Concurrency and Computation: Practice and Experience, 34(7): e5959:1–e5959:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alhaddad:2022:HFN

[AFG^{+}22] Samer Alhaddad, Jens Förstner, Stefan Groth, Daniel Grünewald, Yevgen Grynko, Frank Hannig, Tobias Kenter, Franz-Josef Pfreundt, Christian Plessl, Merlind Schotte, Thomas Steinke, Jürgen Teich, Martin Weiser, and Florian Wende. The HighPerMeshes framework for numerical algorithms on unstructured grids. Concurrency and Computation: Practice and Experience, 34(14):e6616:1–e6616:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abeykoon:2022:SGD

[AFK^{+}22] Vibhatha Abeykoon, Geoffrey Fox, Minje Kim, Saliya Ekanayake, Supun Kamburugamuve, Kannan Govindarajan, Pulasthi Wickramasinghe, Niranda Perera, Chathura Widanage, Ahmet Uyar, Gurhan Gunduz, and Selahatin Akkas. Stochastic gradient descent-based support vector machines training optimization on Big Data and HPC frameworks. Concurrency and Computation: Practice and Experience, 34 (8):e6292:1–e6292:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alanagh:2023:IAM

[AFKS23] Yoosef Alidoost Alanagh, Mojtaba Firouzi, Abdolreza Rasouli Kenari, and Mahboubeh Shamsi. Introducing an adaptive model for auto-scaling cloud computing based on workload classification. Concurrency and Computation: Practice and Experience, 35(22):e7720:1–e7720:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ata:2022:ROC

[AFM22] Mohamed Maher Ata, Mariam Labib Francies, and M. A. Mohamed. A robust optimized convolutional neural network model for human activity recognition using sensing devices. Concurrency and Computation: Practice and Experience, 34 (17):e6964:1–e6964:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Attallah:2021:PLB

[AFNH21] Salma M. A. Attallah, Magda B. Fayek, Salwa M. Nassar, and Elsayed E. Hemayed. Proactive load balancing fault tolerance algorithm in cloud computing. Concurrency and Computation: Practice and Experience, 33(10):e6172:1–e6172:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amirzadeh:2022:SCD

[AG22] Zaman Amirzadeh and Mohammad Gholami. Selective counter design in quantum-dot cellular automata nanotechnology. Concurrency and Computation: Practice and Experience, 34(10):e6798:1–e6798:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Asgarian:2023:GAN

[AGM23] Sepehr Asgarian, Rouzbeh Ghasemi, and Saeedeh Momtazi. Generative adversarial network for sentiment-based stock prediction. Concurrency and Computation: Practice and Experience, 35(2):e7467:1–e7467:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agrawal:2021:DLB

[Agr21] Neha Agrawal. Dynamic load balancing assisted optimized access control mechanism for edge-fog-cloud network in Internet of Things environment. Concurrency and Computation: Practice and Experience, 33(21):e6440:1–e6440:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aruchamy:2023:AIA

[AGSN23] Prasanth Aruchamy, Sabeena Gnanaselvi, Devi Sowndarya, and Pushpalatha Naveenkumar. An artificial intelligence approach for energy-aware intrusion detection and secure routing in internet of things-enabled wireless sensor networks. Concurrency and Computation: Practice and Experience, 35 (23):e7818:1–e7818:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Apel:2020:TKM

[AHS20] Sebastian Apel, Florian Hertrampf, and Steffen Späthe. Toward a knowledge model focusing on microservices and cloud computing. Concurrency and Computation: Practice and Experience, 32(13):e5414:1–e5414:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Asharindavida:2020:FTP

[AHT^{+}20] Fayas Asharindavida, M. Shamim Hossain, Azeemsha Thacham, Hedi Khammari, Irfan Ahmed, Fahad Alraddady, and Mehedi Masud. A forecasting tool for prediction of epileptic seizures using a machine learning approach. Concurrency and Computation: Practice and Experience, 32(1):e5111:1– e5111:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arman:2021:REA

[AHV21] Hosein Arman and Abdollah Hadi-Vencheh. The revised extent analysis method. Concurrency and Computation: Practice and Experience, 33(17):e6319:1–e6319:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Afzal:2022:APM

[AHW22] Ayesha Afzal, Georg Hager, and Gerhard Wellein. Analytic performance model for parallel overlapping memory-bound kernels. Concurrency and Computation: Practice and Experience, 34(10):e6816:1–e6816:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anandaraj:2021:ICS

[AI21] A. Peter Soosai Anandaraj and G. Indumathi. Improved cuckoo search load distribution (ICS-LD) and attack detection in cloud environment. Concurrency and Computation: Practice and Experience, 33(3):e5226:1–e5226:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Andhare:2022:ALB

[AI22] Chetan Vikram Andhare and Dayanand Ingle. Adaptive local binary descriptor statistics: a novel feature set for heart disease diagnosis. Concurrency and Computation: Practice and Experience, 34(27):e7339:1–e7339:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ajmal:2022:CCP

[AIA22] Abdullah Ajmal, Sundas Ibrar, and Rashid Amin. Cloud computing platform: Performance analysis of prominent cryptographic algorithms. Concurrency and Computation: Practice and Experience, 34(15):e6938:1–e6938:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

An:2021:HAM

[AJ21] Sojung An and Jason J. Jung. A heuristic approach on metadata recommendation for search engine optimization. Concurrency and Computation: Practice and Experience, 33(3): e5407:1–e5407:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arif:2021:BDA

[AJAA21] Arooj Arif, Nadeem Javaid, Abdulaziz Aldegheishem, and Nabil Alrajeh. Big data analytics for identifying electricity theft using machine learning approaches in microgrids for smart communities. Concurrency and Computation: Practice and Experience, 33(17):e6316:1–e6316:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amrita:2023:WWO

[AJK^{+}23] Amrita, Shivani Joshi, Rajiv Kumar, Avinash Dwivedi, Vipin Rai, and Sansar Singh Chauhan. Water wave optimized nonsubsampled shearlet transformation technique for multimodal medical image fusion. Concurrency and Computation: Practice and Experience, 35(7):e7591:1–e7591:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alnaggar:2022:BTD

[AJNS22] Omar Abdullah Murshed Farhan Alnaggar, Basavaraj N. Jagadale, Swaroopa H. Narayan, and Mufeed Ahmed Naji Saif. Brain tumor detection from 3D MRI using hyper-layer convolutional neural networks and hyper-heuristic extreme learning machine. Concurrency and Computation: Practice and Experience, 34(24):e7215:1–e7215:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arundhati:2022:AFM

[AJP22] Priya Arundhati, Sisir Kumar Jena, and Santosh Kumar Pani. Approximate function memoization. Concurrency and Computation: Practice and Experience, 34(23):e7204:1–e7204:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anand:2023:ODN

[AJS23] Mallekedi Anand, Anuj Jain, and Manoj Kumar Shukla. An optimized deep network based soil texture prediction and vegetation analysis using satellite image database. Concurrency and Computation: Practice and Experience, 35(7):e7611:1– e7611:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agrawal:2021:CSG

[AK21] Juhi Agrawal and Monit Kapoor. A comparative study on geographic-based routing algorithms for flying ad-hoc networks. Concurrency and Computation: Practice and Experience, 33(16):e6253:1–e6253:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abirami:2022:EED

[AK22a] A. Abirami and R. Kavitha. An efficient early detection of diabetic retinopathy using dwarf mongoose optimization based deep belief network. Concurrency and Computation: Practice and Experience, 34(28):e7364:1–e7364:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ahmad:2022:DLI

[AK22b] Mohassin Ahmad and Farida Khursheed. Detection and localization of image tampering in digital images with fused features. Concurrency and Computation: Practice and Experience, 34(23):e7191:1–e7191:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Akgungor:2022:BSD

[AK22c] Ali Payidar Akgüngör and Ersin Korkmaz. Bézier Search Differential Evolution algorithm based estimation models of delay parameter k for signalized intersections. Concurrency and Computation: Practice and Experience, 34(13):e6931:1– e6931:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abubaker:2022:TDT

[AKA^{+}22a] Zain Abubaker, Asad Ullah Khan, Ahmad Almogren, Shahid Abbas, Atia Javaid, Ayman Radwan, and Nadeem Javaid. Trustful data trading through monetizing IoT data using BlockChain based review system. Concurrency and Computation: Practice and Experience, 34(5):e6739:1–e6739:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Akar:2022:IAR

[Aka22b] Alper Akar. Improving the accuracy of random forest-based land-use classification using fused images and digital surface models produced via different interpolation methods. Concurrency and Computation: Practice and Experience, 34(6): e6787:1–e6787:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Akar:2022:PGU

[AKA22c] Alper Akar, Berkant Konakoglu, and Özlem Akar. Prediction of geoid undulations: Random forest versus classic interpolation techniques. Concurrency and Computation: Practice and Experience, 34(18):e7004:1–e7004:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Akbacak:2022:NMV

[Akb22] Enver Akbacak. A novel multilabel video retrieval method using multiple video queries and deep hash codes. Concurrency and Computation: Practice and Experience, 34(13):e6927:1– e6927:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdallah:2020:PAS

[AKEC20] Sarah Abdallah, Ayman Kayssi, Imad H. Elhajj, and Ali Chehab. Performance analysis of SDN vs OSPF in diverse network environments. Concurrency and Computation: Practice and Experience, 32(21):e5410:1–e5410:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Antkowiak:2024:ALE

[AKLK24] M. Antkowiak, L. Kucharski, R. Lemański, and G. Kamieniarz. Algorithms on low energy spectra of the Hubbard model pertinent to molecular nanomagnets. Concurrency and Computation: Practice and Experience, 36(4):e7931:1–e7931:??, February 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Artigues:2020:EPP

[AKRR20] Victor Artigues, Katharina Kormann, Markus Rampp, and Klaus Reuter. Evaluation of performance portability frameworks for the implementation of a particle-in-cell code. Concurrency and Computation: Practice and Experience, 32(11): e5640:1–e5640:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Allouch:2022:SRR

[AKS^{+}22] Mariem Allouch, Sondes Kallel, Ahmed Soua, Oyunchimeg Shagdar, and Samir Tohme. Survey on radio resource allocation in long-term evolution-vehicle. Concurrency and Computation: Practice and Experience, 34(7):e6228:1–e6228:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arfeen:2022:TAI

[AKUA22] Asad Arfeen, Zunair Ahmed Khan, Riaz Uddin, and Usama Ahsan. Toward accurate and intelligent detection of malware. Concurrency and Computation: Practice and Experience, 34 (4):e6652:1–e6652:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arfeen:2022:PBV

[AKZA22] Asad Arfeen, Muhammad Asim Khan, Obad Zafar, and Usama Ahsan. Process based volatile memory forensics for ransomware detection. Concurrency and Computation: Practice and Experience, 34(4):e6672:1–e6672:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Asahi:2020:OCG

[ALB^{+}20] Yuuichi Asahi, Guillaume Latu, Julien Bigot, Shinya Maeyama, Virginie Grandgirard, and Yasuhiro Idomura. Overlapping communications in gyrokinetic codes on acceleratorbased platforms. Concurrency and Computation: Practice and Experience, 32(5):e5551:1–e5551:??, March 10, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Akram:2021:RTP

[ALBZ21] Naveed Akram, Jianxin Li, Yan Bai, and Yangyang Zhang. Real-time partitioned scheduling: Exploiting the interresource affinity for task allocation on multiprocessors. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Almustafa:2022:CMP

[Alm22] Khaled Mohamad Almustafa. Covid19–Mexican-Patients’ Dataset (Covid19MPD) classification and prediction using feature importance. Concurrency and Computation: Practice and Experience, 34(4):e6675:1–e6675:??, February 15, 2022.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arvapally:2021:IOO

[ALNJ21] Ravi S. Arvapally, Xiaoqing Frank Liu, Fiona Fui-Hoon Nah, and Wei Jiang. Identifying outlier opinions in an online intelligent argumentation system. Concurrency and Computation: Practice and Experience, 33(8):e4107:1–e4107:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alti:2022:DAB

[ALR22] Adel Alti, Abderrahim Lakehal, and Philippe Roose. A decentralized agent-based semantic service control and selfadaptation in smart health mobile applications. Concurrency and Computation: Practice and Experience, 34(5):e6697:1– e6697:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

AlSuwaidan:2021:RDM

[AlS21] Lulwah AlSuwaidan. The role of data management in the industrial Internet of Things. Concurrency and Computation: Practice and Experience, 33(23):e6031:1–e6031:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdellatif:2020:EMI

[AM20a] Takoua Abdellatif and Mohamed Mosbah. Efficient monitoring for intrusion detection in wireless sensor networks. Concurrency and Computation: Practice and Experience, 32(15): e4907:1–e4907:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ahmadpour:2020:NUD

[AM20b] Seyed-Sajad Ahmadpour and Mohammad Mosleh. A novel ultra-dense and low-power structure for fault-tolerant threeinput majority gate in QCA technology. Concurrency and Computation: Practice and Experience, 32(5):e5548:1– e5548:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abbasi:2022:CBA

[AM22] Afsoon Abbasi and Behnaz Mohammadi. A clustering-based anonymization approach for privacy-preserving in the healthcare cloud. Concurrency and Computation: Practice and Experience, 34(1):e6487:1–e6487:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Asgarnezhad:2022:NHF

[AMA22] Razieh Asgarnezhad, S. Amirhassan Monadjemi, and Mohammadreza Soltan Aghaei. A new hierarchy framework for feature engineering through multi-objective evolutionary algorithm in text classification. Concurrency and Computation: Practice and Experience, 34(3):e6594:1–e6594:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abusukhon:2022:ASM

[AMAT22] Ahmad Abusukhon, Zeyad Mohammad, and Ali Al-Thaher. An authenticated, secure, and mutable multiple-session-keys protocol based on elliptic curve cryptography and text-to-image encryption algorithm. Concurrency and Computation: Practice and Experience, 34(4):e6649:1–e6649:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ahmad:2023:ESK

[AMB23] Shahnawaz Ahmad, Shabana Mehfuz, and Javed Beg. An efficient and secure key management with the extended convolutional neural network for intrusion detection in cloud storage. Concurrency and Computation: Practice and Experience, 35 (23):e7806:1–e7806:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdulhussain:2022:FAC

[AMBAJ22] Sadiq H. Abdulhussain, Basheera M. Mahmmod, Thar Baker, and Dhiya Al-Jumeily. Fast and accurate computation of high-order Tchebichef polynomials. Concurrency and Computation: Practice and Experience, 34(27):e7311:1–e7311:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Assi:2020:DLR

[AMD20] Ali Assi, Hamid Mcheick, and Wajdi Dhifli. Data linking over RDF knowledge graphs: a survey. Concurrency and Computation: Practice and Experience, 32(19):e5746:1–e5746:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abkenar:2021:HCM

[AMJK21] Sepideh Bazzaz Abkenar, Ebrahim Mahdipour, Seyed Mahdi Jameii, and Mostafa Haghi Kashani. A hybrid classification method for Twitter spam detection based on differential evolution and random forest. Concurrency and Computation: Practice and Experience, 33(21):e6381:1–e6381:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alappat:2022:ECM

[AML^{+}22] Christie Alappat, Nils Meyer, Jan Laukemann, Thomas Gruber, Georg Hager, Gerhard Wellein, and Tilo Wettig. Execution-cache-memory modeling and performance tuning of sparse matrix-vector multiplication and lattice quantum chromodynamics on A64FX. Concurrency and Computation: Practice and Experience, 34(20):e6512:1–e6512:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arunkumar:2020:FAM

[AMM^{+}20] N. Arunkumar, Mazin Abed Mohammed, Salama A. Mostafa, Dheyaa Ahmed Ibrahim, Joel J. P. C. Rodrigues, and Victor Hugo C. de Albuquerque. Fully automatic model-based segmentation and classification approach for MRI brain tumor using artificial neural networks. Concurrency and Computation: Practice and Experience, 32(1):e4962:1–e4962:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ali:2021:MSA

[AMR^{+}21] Mubashir Ali, Husnain Mushtaq, Muhammad B Rasheed, Anees Baqir, and Thamer Alquthami. Mining software architecture knowledge: Classifying Stack Overflow posts using machine learning. Concurrency and Computation: Practice and Experience, 33(16):e6277:1–e6277:??, August 25, 2021.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Afzal:2021:PAQ

[AMRH21] Humaira Afzal, Muhammad Rafiq Mufti, Aamir Raza, and Abbas Hassan. Performance analysis of QoS in IoT based cognitive radio ad hoc network. Concurrency and Computation: Practice and Experience, 33(23):e5853:1–e5853:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Albalawi:2022:CBC

[AMV22] Umar Albalawi, S Manimurugan, and R Varatharajan. Classification of breast cancer mammogram images using convolution neural network. Concurrency and Computation: Practice and Experience, 34(13):e5803:1–e5803:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Akathoott:2023:SLC

[AN23] Anju Mongandampulath Akathoott and Rupesh Nasre. Single-linkage clustering of dynamic data. Concurrency and Computation: Practice and Experience, 35(1):e7447:1– e7447:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Najjar:2021:SWJ

[ANAMSAR21] Hazem Al-Najjar, S. S. N. Alhady, Junita Mohamad-Saleh, and Nadia Al-Rousan. Scheduling of workflow jobs based on twostep clustering and lowest job weight. Concurrency and Computation: Practice and Experience, 33(20):e6336:1– e6336:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anbu:2022:IMO

[Anb22] M. Anbu. Improved mayfly optimization deep stacked sparse auto encoder feature selection scorched gradient descent driven dropout XLM learning framework for software defect prediction. Concurrency and Computation: Practice and Experience, 34(25):e7240:1–e7240:??, November 15, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIa

[Ano20a] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(1):e5371:1–e5371:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIb

[Ano20b] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(2):e5372:1–e5372:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIc

[Ano20c] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(3):e5373:1–e5373:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IId

[Ano20d] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(4):e5374:1–e5374:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIe

[Ano20e] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(5):e5375:1–e5375:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIf

[Ano20f] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(6):e5376:1–e5376:??, March 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIg

[Ano20g] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(7):e5377:1–e5377:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIh

[Ano20h] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(8):e5378:1–e5378:??, April 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIi

[Ano20i] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(9):e5379:1–e5379:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIj

[Ano20j] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(10):e5380:1–e5380:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIk

[Ano20k] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(11):e5381:1–e5381:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIl

[Ano20l] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(12):e5382:1–e5382:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIm

[Ano20m] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(13):e5383:1–e5383:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIn

[Ano20n] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(14):e5384:1–e5384:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIo

[Ano20o] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(15):e5385:1–e5385:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIp

[Ano20p] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(16):e5386:1–e5386:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIq

[Ano20q] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(17):e5387:1–e5387:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIr

[Ano20r] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(18):e5388:1–e5388:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIs

[Ano20s] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(19):e5389:1–e5389:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIt

[Ano20t] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(20):e5390:1–e5390:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIu

[Ano20u] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(21):e5391:1–e5391:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIv

[Ano20v] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(22):e5392:1–e5392:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIw

[Ano20w] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(23):e5393:1–e5393:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2020:IIx

[Ano20x] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 32(24):e5394:1–e5394:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:AMI

[Ano21a] Anonymous. Analysis of multi-input multi-output transactions in the Bitcoin network. Concurrency and Computation: Practice and Experience, 33(1):e5629:1–e5629:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:BAB

[Ano21b] Anonymous. Blockchain applications beyond the cryptocurrency casino: The Punishment not Reward blockchain architecture. Concurrency and Computation: Practice and Experience, 33(1):e5749:1–e5749:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:DSB

[Ano21c] Anonymous. Decision support based on optimized data mining techniques: Application to mobile telecommunication companies. Concurrency and Computation: Practice and Experience, 33(1):e5833:1–e5833:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:DPI

[Ano21d] Anonymous. Design and practical implementation of verifyyour-vote protocol. Concurrency and Computation: Practice and Experience, 33(1):e5813:1–e5813:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:DEM

[Ano21e] Anonymous. A discrete event model for analysis and verification of time-constrained business processes. Concurrency and Computation: Practice and Experience, 33(1):e5753:1– e5753:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:ETT

[Ano21f] Anonymous. Ensuring transparency and traceability of food local products: a blockchain application to a Smart Tourism Region. Concurrency and Computation: Practice and Experience, 33(1):e5857:1–e5857:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:FPP

[Ano21g] Anonymous. FuSTM: ProM plugin for fuzzy similar tasks mining based on entropy measure. Concurrency and Computation: Practice and Experience, 33(1):e5821:1–e5821:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:GBT

[Ano21h] Anonymous. Graph-based tag recommendations using clusters of patients in clinical decision support system. Concurrency and Computation: Practice and Experience, 33(1): e5624:1–e5624:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IES

[Ano21i] Anonymous. Implementation and evaluation of smart contracts using a hybrid on- and off-blockchain architecture. Concurrency and Computation: Practice and Experience, 33 (1):e5811:1–e5811:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:ISM

[Ano21j] Anonymous. IPDS: a semantic mediator-based system using spark for the integration of heterogeneous proteomics data sources. Concurrency and Computation: Practice and Experience, 33(1):e5814:1–e5814:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIa

[Ano21k] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(1):e5865:1–e5865:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIb

[Ano21l] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(2):e5866:1–e5866:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIc

[Ano21m] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(3):e5867:1–e5867:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IId

[Ano21n] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(4):e5868:1–e5868:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIe

[Ano21o] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(5):e5869:1–e5869:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIf

[Ano21p] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(6):e5870:1–e5870:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIg

[Ano21q] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(7):e5871:1–e5871:??, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIh

[Ano21r] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(8):e5872:1–e5872:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIi

[Ano21s] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(9):e5873:1–e5873:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIj

[Ano21t] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(10):e5874:1–e5874:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIk

[Ano21u] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(11):e5875:1–e5875:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIl

[Ano21v] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(12):e5876:1–e5876:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIm

[Ano21w] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(13):e5877:1–e5877:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIn

[Ano21x] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(14):e5878:1–e5878:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIo

[Ano21y] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(15):e5879:1–e5879:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIp

[Ano21z] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(16):e5880:1–e5880:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIq

[Ano21-27] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(17):e5881:1–e5881:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIr

[Ano21-28] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(18):e5882:1–e5882:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIs

[Ano21-29] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(19):e5883:1–e5883:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIt

[Ano21-30] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(20):e5884:1–e5884:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIu

[Ano21-31] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(21):e5885:1–e5885:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIv

[Ano21-32] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(22):e5886:1–e5886:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIw

[Ano21-33] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(23):e5887:1–e5887:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:IIx

[Ano21-34] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 33(24):e5888:1–e5888:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:MCD

[Ano21-35] Anonymous. Multidimensional community discovering in heterogeneous social networks. Concurrency and Computation: Practice and Experience, 33(1):e5809:1–e5809:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:NMH

[Ano21-36] Anonymous. A new method with hybrid direction for linear programming. Concurrency and Computation: Practice and Experience, 33(1):e5836:1–e5836:??, January 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:NLC

[Ano21-37] Anonymous. Novel low complexity decoding method for the binary coded frequency shift keying modulation. Concurrency and Computation: Practice and Experience, 33(1):e5760:1– e5760:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:NOS

[Ano21-38] Anonymous. A novel OFDM switching algorithm for symmetric and asymmetric MIMO architectures. Concurrency and Computation: Practice and Experience, 33(1):e5831:1– e5831:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:OLD

[Ano21-39] Anonymous. Opinion leaders’ detection in dynamic social networks. Concurrency and Computation: Practice and Experience, 33(1):e5692:1–e5692:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:RUD

[Ano21-40] Anonymous. A roadside unit deployment framework for enhancing transportation services in Maghrebian cities. Concurrency and Computation: Practice and Experience, 33(1): e5611:1–e5611:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:SIC

[Ano21-41] Anonymous. Special issue of concurrency and computation: Practice and experience “FPDAPP, Future Perspectives on Decentralized Applications”. Concurrency and Computation: Practice and Experience, 33(1):e6107:1–e6107:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:TUH

[Ano21-42] Anonymous. Towards understanding and harnessing the potential of Africa in digitalization. Concurrency and Computation: Practice and Experience, 33(1):e6127:1–e6127:??, January 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2021:TSR

[Ano21-43] Anonymous. Trusted systems of records based on blockchain technology — a prototype for mileage storing in the automotive industry. Concurrency and Computation: Practice and Experience, 33(1):e5630:1–e5630:??, January 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIa

[Ano22a] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(1):e6386:1–e6386:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIb

[Ano22b] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(2):e6387:1–e6387:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIc

[Ano22c] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(3):e6388:1–e6388:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IId

[Ano22d] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(4):e6389:1–e6389:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIe

[Ano22e] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(5):e6390:1–e6390:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIf

[Ano22f] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(6):e6391:1–e6391:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIg

[Ano22g] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(7):e6392:1–e6392:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIh

[Ano22h] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(8):e6393:1–e6393:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIi

[Ano22i] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(9):e6394:1–e6394:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIj

[Ano22j] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(10):e6395:1–e6395:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIk

[Ano22k] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(11):e6396:1–e6396:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIl

[Ano22l] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(12):e6397:1–e6397:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIm

[Ano22m] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(13):e6398:1–e6398:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIn

[Ano22n] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(14):e6399:1–e6399:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIo

[Ano22o] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(15):e6400:1–e6400:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIp

[Ano22p] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(16):e6401:1–e6401:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIq

[Ano22q] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(17):e6402:1–e6402:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIr

[Ano22r] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(18):e6403:1–e6403:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIs

[Ano22s] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(19):e6404:1–e6404:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIt

[Ano22t] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(20):e6405:1–e6405:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIu

[Ano22u] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(21):e6406:1–e6406:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIv

[Ano22v] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(22):e6407:1–e6407:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIw

[Ano22w] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(23):e6408:1–e6408:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIx

[Ano22x] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(24):e6409:1–e6409:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIy

[Ano22y] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(25):e6410:1–e6410:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIz

[Ano22z] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(26):e6411:1–e6411:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIba

[Ano22-27] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(27):e6412:1–e6412:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2022:IIbb

[Ano22-28] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 34(28):e6413:1–e6413:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:E

[Ano23a] Anonymous. Erratum. Concurrency and Computation: Practice and Experience, 35(14):e7777:1–e7777:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIa

[Ano23b] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(1):e7070:1–e7070:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIb

[Ano23c] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(2):e7071:1–e7071:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIc

[Ano23d] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(3):e7072:1–e7072:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IId

[Ano23e] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(4):e7073:1–e7073:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIe

[Ano23f] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(5):e7074:1–e7074:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIf

[Ano23g] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(6):e7075:1–e7075:??, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIg

[Ano23h] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(7):e7076:1–e7076:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIh

[Ano23i] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(8):e7077:1–e7077:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIi

[Ano23j] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(9):e7078:1–e7078:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIj

[Ano23k] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(10):e7079:1–e7079:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIk

[Ano23l] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(11):e7080:1–e7080:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIl

[Ano23m] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(12):e7081:1–e7081:??, May 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIm

[Ano23n] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(13):e7082:1–e7082:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIn

[Ano23o] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(14):e7083:1–e7083:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIo

[Ano23p] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(15):e7084:1–e7084:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIp

[Ano23q] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(16):e7085:1–e7085:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIq

[Ano23r] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(17):e7086:1–e7086:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIr

[Ano23s] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(18):e7087:1–e7087:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIs

[Ano23t] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(19):e7088:1–e7088:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIt

[Ano23u] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(20):e7089:1–e7089:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIu

[Ano23v] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(21):e7090:1–e7090:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIv

[Ano23w] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(22):e7091:1–e7091:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIw

[Ano23x] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(23):e7092:1–e7092:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIx

[Ano23y] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(24):e7093:1–e7093:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIy

[Ano23z] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(25):e7094:1–e7094:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIz

[Ano23-27] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(26):e7095:1–e7095:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIba

[Ano23-28] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(27):e7096:1–e7096:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:IIbb

[Ano23-29] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 35(28):e7097:1–e7097:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2023:R

[Ano23-30] Anonymous. Retraction. Concurrency and Computation: Practice and Experience, 35(16):e7747:1–e7747:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IIa

[Ano24a] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(1):e7778:1–e7778:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IIb

[Ano24b] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(2):e7779:1–e7779:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IIc

[Ano24c] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(3):e7780:1–e7780:??, February 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IId

[Ano24d] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(4):e7781:1–e7781:??, February 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IIe

[Ano24e] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(5):e7782:1–e7782:??, February 29, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IIf

[Ano24f] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(6):e7783:1–e7783:??, March 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IIg

[Ano24g] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(7):e7784:1–e7784:??, March 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IIh

[Ano24h] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(8):e7785:1–e7785:??, April 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IIi

[Ano24i] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(9):e7786:1–e7786:??, April 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IIj

[Ano24j] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(10):e7787:1–e7787:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IIk

[Ano24k] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(11):e7788:1–e7788:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IIl

[Ano24l] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(12):e7789:1–e7789:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:2024:IIm

[Ano24m] Anonymous. Issue information. Concurrency and Computation: Practice and Experience, 36(14):e7791:1–e7791:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agarwal:2020:AHM

[ANP^{+}20] Pratul K. Agarwal, Thomas Naughton, Byung H. Park, David E. Bernholdt, Joshua J. Hursey, and Al Geist. Application health monitoring for extreme-scale resiliency using cooperative fault management. Concurrency and Computation: Practice and Experience, 32(2):e5449:1–e5449:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aydemir:2023:MPA

[AO23] Salih Berkan Aydemir and Funda Kutlu Onay. Marine predator algorithm with elite strategies for engineering design problems. Concurrency and Computation: Practice and Experience, 35(7):e7612:1–e7612:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Omairi:2021:EAG

[AOACAQ21] Lamyaa J. Al-Omairi, Jemal Abawajy, Morshed U. Chowdhury, and Tahsien Al-Quraishi. An empirical analysis of graph-based linear dimensionality reduction techniques. Concurrency and Computation: Practice and Experience, 33(5): e5990:1–e5990:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

A:2022:MST

[AP22] Rajesh A. and Esther Rani P. Mitigation of stretch type vampire attack using probabilistic variable fuzzy rough set in wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(4):e6665:1–e6665:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agrawal:2023:FFB

[AP23] Deepika Agrawal and Sudhakar Pandey. FUCBR: Fuzzybased unequal clustering and bat-based routing protocol for WSNs assisted IoT. Concurrency and Computation: Practice and Experience, 35(26):e7833:1–e7833:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aklak:2022:SVI

[APA22] Annis Fathima Aklak, Murugesh Yadhav Pugazhenthi, and John Sahaya Rani Alex. A study on VLSI implementation of image enhancement techniques. Concurrency and Computation: Practice and Experience, 34(10):e6734:1–e6734:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Autili:2021:HCD

[APL^{+}21] Marco Autili, Alexander Perucci, Leonardo Leite, Massimo Tivoli, Fabio Kon, and Amleto Di Salle. Highly collaborative distributed systems: Synthesis and enactment at work. Concurrency and Computation: Practice and Experience, 33(6): e6039:1–e6039:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alqahtani:2021:HGB

[APM^{+}21] Abdullah Saleh Alqahtani, Saravanan Pandiaraj, Maheswari Murali, Sami Alshmrany, and Haytham Alsarrayrih. Hybrid grass bee optimization-multikernal extreme learning classifier: Multimodular fusion strategy and optimal feature selection for multimodal sentiment analysis in social media videos. Concurrency and Computation: Practice and Experience, 33 (16):e6259:1–e6259:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ahmadi:2022:FLB

[APM22] Zahra Ahmadi, Fereshteh-Azadi Parand, and Farzam Matinfar. A fuzzy logic-based approach for fuzzy queries over NoSQL graph database. Concurrency and Computation: Practice and Experience, 34(1):e6542:1–e6542:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Achary:2021:PSM

[APP^{+}21] Thimershen Achary, Shivani Pillay, Sarah M. Pillai, Malusi Mqadi, Emma Genders, and Absalom E. Ezugwu. A performance study of meta-heuristic approaches for quadratic assignment problem. Concurrency and Computation: Practice and Experience, 33(17):e6321:1–e6321:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Altinok:2022:GAE

[APTT22] Kaan Furkan Altinok, Afsin Peker, Cihangir Tezcan, and Alptekin Temizel. GPU accelerated 3DES encryption. Concurrency and Computation: Practice and Experience, 34(9): e6507:1–e6507:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anitha:2023:GRA

[APV23] S. Anitha, T. Padma, and V. Vallimayil. A generic resource augmentation architecture for efficient mobile communication. Concurrency and Computation: Practice and Experience, 35(21):e7703:1–e7703:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aliero:2022:DSQ

[AQP^{+}22] Muhammad Saidu Aliero, Kashif Naseer Qureshi, Muhammad Fermi Pasha, Awais Ahmad, and Gwanggil Jeon. Detection of structure query language injection vulnerability in web driven database application. Concurrency and Computation: Practice and Experience, 34(13):e5936:1–e5936:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

A:2022:EHP

[AR22a] Mohammed Hashim B A and Amutha R. Elderly Hajj pilgrims activity recognition based on candidate classification technique. Concurrency and Computation: Practice and Experience, 34(13):e6932:1–e6932:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arunachalam:2022:NAC

[AR22b] Saran Kumar Arunachalam and Rajagopal Rekha. A novel approach for cardiovascular disease prediction using machine learning algorithms. Concurrency and Computation: Practice and Experience, 34(19):e7027:1–e7027:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Adouth:2023:BBC

[AR23] Vamshi Adouth and Eswari Rajagopal. Blockchain-based certificateless public auditing with privacy-preserving for cloudbased cyber-physical systems. Concurrency and Computation: Practice and Experience, 35(12):e7690:1–e7690:??, May 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Almurshed:2022:GNH

[ARC22] Osama Almurshed, Omer Rana, and Kyle Chard. Greedy nominator heuristic: Virtual function placement on fog resources. Concurrency and Computation: Practice and Experience, 34(6):e6765:1–e6765:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Areed:2021:MSS

[ARFA21] Marwa F. Areed, Mohamed M. Rashed, Nehal Fayez, and Ehab H. Abdelhay. Modified SeDaSc system for efficient data sharing in the cloud. Concurrency and Computation: Practice and Experience, 33(21):e6377:1–e6377:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Adhinugraha:2021:DNC

[ARHT21] Kiki Adhinugraha, Wenny Rahayu, Takahiro Hara, and David Taniar. Dealing with noise in crowdsourced GPS human trajectory logging data. Concurrency and Computation: Practice and Experience, 33(19):e6139:1–e6139:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arora:2022:EAD

[ARS22a] Shruti Arora, Rinkle Rani, and Nitin Saxena. An efficient approach for detecting anomalous events in real-time weather datasets. Concurrency and Computation: Practice and Experience, 34(5):e6707:1–e6707:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arora:2022:MVA

[ARS22b] Shruti Arora, Rinkle Rani, and Nitin Saxena. Modified valence aware dictionary for sentiment reasoning classifier for detection and classification of Covid-19 related rumors from social media data streams. Concurrency and Computation: Practice and Experience, 34(21):e7124:1–e7124:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arumugam:2022:DAB

[Aru22] S. S. Arumugam. Development of argument based opinion mining model with sentimental data analysis from Twitter content. Concurrency and Computation: Practice and Experience, 34(15):e6956:1–e6956:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arumugam:2020:AIC

[AS20] Maheswari Arumugam and Arun Kumar Sangaiah. Arrhythmia identification and classification using wavelet centered methodology in ECG signals. Concurrency and Computation: Practice and Experience, 32(17):e5553:1–e5553:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agarwal:2022:SRD

[AS22a] Pratik Agarwal and Bam Bahadur Sinha. Scalable resource description framework clustering: a distributed approach for analyzing knowledge graphs using minHash locality sensitive hashing. Concurrency and Computation: Practice and Experience, 34(15):e6966:1–e6966:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ahmad:2022:UEB

[AS22b] Mir Shahnawaz Ahmad and Shahid Mehraj Shah. Unsupervised ensemble based deep learning approach for attack detection in IoT network. Concurrency and Computation: Practice and Experience, 34(27):e7338:1–e7338:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alavani:2022:PMG

[AS22c] Gargi Alavani and Santonu Sarkar. Performance modeling of graphics processing unit application using static and dynamic analysis. Concurrency and Computation: Practice and Experience, 34(3):e6602:1–e6602:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aravindan:2022:MID

[AS22d] T. E. Aravindan and R. Seshasayanan. Medical image DE-NOISING scheme using discrete wavelet transform and optimization with different noises. Concurrency and Computation: Practice and Experience, 34(8):e5540:1–e5540:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Agrawal:2023:HII

[AS23a] Akhileshwar Prasad Agrawal and Nanhay Singh. A hybrid intelligently initialized particle swarm optimizer with weight factored binary gray wolf optimizer for mitigation of security issues in Internet of Things and sensor nodes. Concurrency and Computation: Practice and Experience, 35(5):e7548:1– e7548:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alphonse:2023:EVD

[AS23b] Babitha Alphonse and Boyed Wesley Alfred Sylam. Enhancement of video demosaicing techniques using optimization algorithm. Concurrency and Computation: Practice and Experience, 35(8):e7616:1–e7616:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aruna:2023:EVM

[AS23c] Arulanandam Gnanaprakasam Aruna and Manickam Sangeetha. An effective veracity management for sentiment analysis in social media using crowdsourcing and multi-set feature learning. Concurrency and Computation: Practice and Experience, 35 (7):e7593:1–e7593:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Apirajitha:2024:DBC

[AS24] Pichai Subramanian Apirajitha and Godfrey Winster Sathianesan. On developing Block-Chain based Secure Storage Model (BSSM) with auditing and integrity analysis in the cloud. Concurrency and Computation: Practice and Experience, 36(7):e7970:1–e7970:??, March 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arfeen:2021:CRM

[ASA^{+}21] Zeeshan Ahmad Arfeen, Usman Ullah Sheikh, Mehreen Kausar Azam, Rabia Hassan, Hafiz Muhammad Faisal Shehzad, Shahzad Ashraf, Md Pauzi Abdullah, and Lubna Aziz. A comprehensive review of modern trends in optimization techniques applied to hybrid microgrid systems. Concurrency and Computation: Practice and Experience, 33(10):e6165:1– e6165:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Shammary:2022:EPS

[ASAAAA22] Dhiah Al-Shammary, Adil L. Albukhnefis, Ali Hakem Alsaeedi, and Muntasir Al-Asfoor. Extended particle swarm optimization for feature selection of high-dimensional biomedical data. Concurrency and Computation: Practice and Experience, 34(10):e6776:1–e6776:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ashutosh:2022:SSS

[ASANR22] Ashutosh Ashutosh, Usman Shahzad, Nadia H. Al-Noor, and Piyush Kant Rai. Simulation study of small domain with calibration approach. Concurrency and Computation: Practice and Experience, 34(27):e7323:1–e7323:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wakrime:2023:FIS

[ASB23] Abderrahim Ait Wakrime, Mohamed Sellami, and Riadh Ben Halima. Future internet services and applications. Concurrency and Computation: Practice and Experience, 35(11): e7329:1–e7329:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anitha:2022:MGW

[ASC22] S. Anitha, S. Saravanan, and A. Chandrasekar. A modified gray wolf-based chameleon swarm algorithm for minimizing energy consumption and enabling secure communication in wireless sensor network. Concurrency and Computation: Practice and Experience, 34(26):e7295:1–e7295:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aldwyan:2021:EDC

[ASJ21] Yasser Aldwyan, Richard O. Sinnott, and Glenn T. Jayaputera. Elastic deployment of container clusters across geographically distributed cloud data centers for web applications. Concurrency and Computation: Practice and Experience, 33(21): e6436:1–e6436:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ansari:2024:UPD

[ASJ24] Mohd Aquib Ansari, Dushyant Kumar Singh, and Ruchi Jayaswal. Using postural data and recurrent learning to monitor shoplifting activities in megastores. Concurrency and Computation: Practice and Experience, 36(2):e7900:1– e7900:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Al-Sawwa:2020:PPS

[ASL20] Jamil Al-Sawwa and Simone A. Ludwig. Parallel particle swarm optimization classification algorithm variant implemented with Apache Spark. Concurrency and Computation: Practice and Experience, 32(2):e5451:1–e5451:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aslan:2022:NTB

[Asl22] Serpil Aslan. A novel TCNN-bi-LSTM deep learning model for predicting sentiments of tweets about COVID-19 vaccines. Concurrency and Computation: Practice and Experience, 34(28):e7387:1–e7387:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Azumah:2021:PMC

[ASMK21] Kenneth K. Azumah, Lene T. Sørensen, Raffaele Montella, and Sokol Kosta. Process mining-constrained scheduling in the hybrid cloud. Concurrency and Computation: Practice and Experience, 33(4):e6025:1–e6025:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aravinth:2021:HSI

[ASMS21] S. S. Aravinth, J. Senthilkumar, V. Mohanraj, and Y. Suresh. A hybrid swarm intelligence based optimization approach for solving minimum exposure problem in wireless sensor networks. Concurrency and Computation: Practice and Experience, 33(3):e5370:1–e5370:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arunachalam:2023:CEE

[ASRN23] Gnana Soundari Arunachalam, Vimal S, Gomathi Ramalingam, and Rajendran Nanjappan. A classy energy efficient spider monkey optimization based clustering and data aggregation models for wireless sensor network. Concurrency and Computation: Practice and Experience, 35(2):e7492:1– e7492:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qaraghuli:2022:DAF

[AST22] Hawraa Al Qaraghuli, Reza Sheibani, and Hamid Tabatabaee. Detection of atrial fibrillation using variable length genetic algorithm and convolutional neural network. Concurrency and Computation: Practice and Experience, 34(10):e6789:1– e6789:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arslan:2022:ACA

[AT22] Recep Sinan Arslan and Murat Tasyurek. AMD-CNN: Android malware detection via feature graph and convolutional neural networks. Concurrency and Computation: Practice and Experience, 34(23):e7180:1–e7180:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Avci:2023:APL

[ATC23] Cigdem Avci, Bedir Tekinerdogan, and Cagatay Catal. Analyzing the performance of long short-term memory architectures for malware detection models. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aydemir:2023:CCP

[ATG23] Hamza Aydemir, Mehmet Tekerek, and Mehmet Gök. Complete coverage planning with clustering method for autonomous mobile robots. Concurrency and Computation: Practice and Experience, 35(26):e7830:1–e7830:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Akbar:2022:CAA

[ATS22] Sumaiya Begum Akbar, Kalaiselvi Thanupillai, and Suganthi Sundararaj. Combining the advantages of AlexNet convolutional deep neural network optimized with anopheles search algorithm based feature extraction and random forest classifier for COVID-19 classification. Concurrency and Computation: Practice and Experience, 34(15):e6958:1–e6958:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arum:2022:CPC

[AU22] Kingsley Chinedu Arum and Fidelis Ifeanyi Ugwuowo. Combining principal component and robust ridge estimators in linear regression model with multicollinearity and outlier. Concurrency and Computation: Practice and Experience, 34(10): e6803:1–e6803:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arslan:2023:ETC

[AÜ23] Sanem Arslan and Osman Ünsal. Efficient thread-to-core mapping alternatives for application-level redundant multithreading. Concurrency and Computation: Practice and Experience, 35(24):e7622:1–e7622:??, November 1, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Arif:2021:RBP

[AV21] Mahwish Arif and Hans Vandierendonck. Reducing the burden of parallel loop schedulers for many-core processors. Concurrency and Computation: Practice and Experience, 33(13): e6241:1–e6241:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abinash:2022:GDC

[AV22a] M. J. Abinash and V. Vasudevan. Gene data classification using map reduce based linear SVM. Concurrency and Computation: Practice and Experience, 34(8):e5497:1–e5497:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anitha:2022:NFS

[AV22b] Swaminathan Anitha and Muthuraman Vanitha. A novel feature selection with stochastic gradient descent logistic regression for multilabeled stress prediction in working employees. Concurrency and Computation: Practice and Experience, 34 (13):e6911:1–e6911:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alagappan:2023:DAE

[AV23] Jothi Kumar Alagappan and Savaridoss Paul Victor. Deep auto-encoder based clustering algorithm for graph-based web page recommendation system. Concurrency and Computation: Practice and Experience, 35(2):e7505:1–e7505:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:MCV

[aWLY^{+}22] Yun an Wang, Yaojin Lin, Xiehua Yu, Jie Wang, Zhisen Wei, Jianfeng Wu, and Shaozi Li. Multilabel causal variable discovery in multisource. Concurrency and Computation: Practice and Experience, 34(21):e7067:1–e7067:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

An:2022:AVA

[AWS^{+}22] Jing An, Kefan Wang, Hui Sun, Can Cui, Wei Li, and Chao Ma. Attention virtual adversarial based semi-supervised question generation. Concurrency and Computation: Practice and Experience, 34(10):e6797:1–e6797:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alaei:2021:HPF

[AY21] Mohammad Alaei and Fahimeh Yazdanpanah. A highperformance FPGA-based multicrossbar prioritized networkon-chip. Concurrency and Computation: Practice and Experience, 33(6):e6055:1–e6055:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ay:2023:PEB

[AY23] Ayse Nur Ay and Mustafa Zahid Yildiz. The performance of an electromyography-based deep neural network classifier for external and internal focus instructions. Concurrency and Computation: Practice and Experience, 35(2):e7470:1– e7470:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Awan:2021:CCI

[AYB21] Irfan Awan, Muhammad Younas, and Salima Benbernou. Convergence of cloud, Internet of Things, and big data: New platforms and applications. Concurrency and Computation: Practice and Experience, 33(23):e6667:1–e6667:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aydin:2021:CWW

[AYD21] Bahadir Ismail Aydin, Yavuz Selim Yilmaz, and Murat Demirbas. A crowdsourced “Who wants to be a millionaire?” player. Concurrency and Computation: Practice and Experience, 33(8):e4168:1–e4168:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

An:2021:CCD

[AYG^{+}21] Xianxi An, Sihan You, Ziyang Guo, Zeguang Lu, Bo Zheng, Shengfei Shi, and Yan Song. Column concept determination based on multiple evidences. Concurrency and Computation: Practice and Experience, 33(8):e5457:1–e5457:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Awan:2020:ECF

[AYH20] Irfan Awan, Muhammad Younas, and Farookh Hussain. Emerging challenges and frontiers in cloud computing. Concurrency and Computation: Practice and Experience, 32(1): e5500:1–e5500:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aarif:2022:SWS

[AYH^{+}22] K. O. Mohammed Aarif, Caffiyar Mohamed Yousuff, B. A. Mohammed Hashim, C. Mohamed Hashim, and P. Sivakumar. Smart bin: Waste segregation system using deep learning-Internet of Things for sustainable smart cities. Concurrency and Computation: Practice and Experience, 34(28):e7378:1–e7378:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aswini:2022:ECB

[AYJ^{+}22] J. Aswini, B. Yamini, Rajaram Jatothu, K. Sankara Nayaki, and M. Nalini. An efficient cloud-based healthcare services paradigm for chronic kidney disease prediction application using boosted support vector machine. Concurrency and Computation: Practice and Experience, 34(10):e6722:1–e6722:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aryan:2021:PAD

[AYKE21] Ramtin Aryan, Anis Yazidi, Øivind Kure, and Paal Einar Engelstad. A parallel approach for detecting OpenFlow rule anomalies based on a general formalism. Concurrency and Computation: Practice and Experience, 33(15):e5907:1– e5907:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

An:2020:CSO

[AZA20] Xiaogang An, Weiwei Zhu, and Na An. Control system optimization of spillage brake based on RBF neural network. Concurrency and Computation: Practice and Experience, 32(23): e5588:1–e5588:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ajarroud:2020:CBA

[AZI20] Ouafa Ajarroud, Ahmed Zellou, and Ali Idri. A coveragebased approach for filtering and prioritizing regions in a semantic cache. Concurrency and Computation: Practice and Experience, 32(9):e5639:1–e5639:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Aba:2020:EAS

[AZM20] Massinissa Ait Aba, Lilia Zaourar, and Alix Munier. Efficient algorithm for scheduling parallel applications on hybrid multicore machines with communications delays and energy constraint. Concurrency and Computation: Practice and Experience, 32(15):e5573:1–e5573:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

B:2020:GPR

[BA20] Sudeepa K. B. and Ganesh Aithal. Generation of pseudo random number sequence from discrete oscillating samples of equally spread objects and application for stream cipher system. Concurrency and Computation: Practice and Experience, 32(1):e5181:1–e5181:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bendaoud:2024:MLA

[BA24] Fayssal Bendaoud and Marwen Abdennebi. A machine learning access network selection in a heterogeneous wireless environment. Concurrency and Computation: Practice and Experience, 36(9):e7989:1–e7989:??, April 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ben-Abdallah:2021:CCA

[BABLH21] Emna Ben-Abdallah, Khouloud Boukadi, Jaime Lloret, and Mohamed Hammami. CROSA: Context-aware cloud service ranking approach using online reviews based on sentiment analysis. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baker:2021:AWB

[BABS21] Qanita Bani Baker, Mohammad A. Alsmirat, Khaled Balhaf, and Mohammed A. Shehab. Accelerating white blood cells image segmentation using GPUs. Concurrency and Computation: Practice and Experience, 33(2):e5133:1–e5133:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Badia:2023:SII

[Bad23] Rosa M. Badia. Special issue 19th international workshop on algorithms, models and tools for parallel computing on heterogeneous platforms (HeteroPar’21). Concurrency and Computation: Practice and Experience, 35(20):e7413:1–e7413:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bai:2023:DLB

[Bai23] Jiayuan Bai. Deep learning-based intraoperative video analysis for supporting surgery. Concurrency and Computation: Practice and Experience, 35(26):e7837:1–e7837:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bergman:2020:PBD

[BANT20] Sara Bergman, Mikael Asplund, and Simin Nadjm-Tehrani. Permissioned blockchains and distributed databases: a performance study. Concurrency and Computation: Practice and Experience, 32(12):e5227:1–e5227:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baheti:2022:DFD

[BAPS22] Shrey Baheti, Parwat Singh Anjana, Sathya Peri, and Yogesh Simmhan. DiPETrans: a framework for distributed parallel execution of transactions of blocks in blockchains. Concurrency and Computation: Practice and Experience, 34(10): e6804:1–e6804:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bernaschi:2021:BMG

[BAR21] Massimo Bernaschi, Elena Agostini, and Davide Rossetti. Benchmarking multi-GPU applications on modern multi-GPU integrated systems. Concurrency and Computation: Practice and Experience, 33(14):e5470:1–e5470:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bezerra:2023:PCS

[BB23] Rodrigo F. Bezerra and Jacir L. Bordim. Predicting consecutive spectrum opportunities with hidden Markov models. Concurrency and Computation: Practice and Experience, 35 (19):e7569:1–e7569:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bacha:2022:EDD

[BBA^{+}22] Sawcen Bacha, Khawla Ben Abdellafou, Ahamed Aljuhani, Okba Taouali, and Noureddine Liouane. Early detection of digital mammogram using kernel extreme learning machine. Concurrency and Computation: Practice and Experience, 34 (15):e6971:1–e6971:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benallal:2020:KBA

[BBB^{+}20a] Wehbi Benallal, Mahmoud Barhamgi, Djamal Benslimane, Zakaria Maamar, Noura Faci, and Ameni Bellaaj. A knowledge-based approach to manage configurable business processes. Concurrency and Computation: Practice and Experience, 32(15):e4920:1–e4920:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bernholdt:2020:SMU

[BBB^{+}20b] David E. Bernholdt, Swen Boehm, George Bosilca, Manjunath Gorentla Venkata, Ryan E. Grant, Thomas Naughton, Howard P. Pritchard, Martin Schulz, and Geoffroy R. Vallee. A survey of MPI usage in the US exascale computing project. Concurrency and Computation: Practice and Experience, 32 (3):e4851:1–e4851:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brown:2020:MEG

[BBB^{+}20c] Nick Brown, Brian Bainbridge, Ciarán Beggan, William Brown, Brian Hamilton, and Susan Macmillan. Modelling the Earth’s geomagnetic environment on Cray machines using PETSc and SLEPc. Concurrency and Computation: Practice and Experience, 32(20):e5660:1–e5660:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boruah:2022:RED

[BBB22] Arpita Nath Boruah, Saroj Kr. Biswas, and Sivaji Bandyopadhyay. Rule extraction from decision tree: Transparent expert system of rules. Concurrency and Computation: Practice and Experience, 34(15):e6935:1–e6935:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Badica:2022:EBW

[BBBC22] Amelia Badica, Costin Badica, Ion Buligiu, and Liviu Ion Ciora. Exploring the Blocks World state space. Concurrency and Computation: Practice and Experience, 34(14):e6472:1– e6472:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bajpai:2023:DWN

[BBD23] Vikas Bajpai, Anukriti Bansal, and Subrat Dash. A deep and wide neural network to predict summer monsoon rainfall using time series data. Concurrency and Computation: Practice and Experience, 35(8):e7626:1–e7626:??, April 10, 2023.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baruah:2024:RVF

[BBD24] Sangita Baruah, Dhruba Jyoti Borah, and Vaskar Deka. Reviewing various feature selection techniques in machine learning-based botnet detection. Concurrency and Computation: Practice and Experience, 36(12):e8076:1–e8076:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bahig:2021:FSA

[BBF21] Hazem M. Bahig, Hatem M. Bahig, and Khaled A. Fathy. Fast and scalable algorithm for product large data on multicore system. Concurrency and Computation: Practice and Experience, 33(2):e5259:1–e5259:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Blaschke:2024:RTX

[BBP^{+}24] Johannes P. Blaschke, Aaron S. Brewster, Daniel W. Paley, Derek Mendez, Asmit Bhowmick, Nicholas K. Sauter, Wilko Kröger, Murali Shankar, Bjoern Enders, and Deborah Bard. Real-time XFEL data analysis at SLAC and NERSC: a trial run of nascent exascale experimental data analysis. Concurrency and Computation: Practice and Experience, 36(12): e8019:1–e8019:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balaji:2021:CAR

[BC21] Mahesh Balaji and Aswani Kumar Ch. Context-aware resource management and alternative pricing model to improve enterprise cloud adoption. Concurrency and Computation: Practice and Experience, 33(6):e6056:1–e6056:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Babar:2022:WRL

[BC22] Iqra Babar and Sohail Chand. Weighted ridge and Liu estimators for linear regression model. Concurrency and Computation: Practice and Experience, 34(27):e7343:1–e7343:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Banerji:2023:AIS

[BC23a] Nandan Banerji and Subhrabrata Choudhury. Auction inspired service replication for context-aware IoT environment. Concurrency and Computation: Practice and Experience, 35 (9):e7641:1–e7641:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Beeharry:2023:PAR

[BC23b] Yogesh Beeharry and Dujaya R. Calchand. Performance analysis of regression-based machine learning models towards intelligent selection of MIMO configurations. Concurrency and Computation: Practice and Experience, 35(1):e7415:1– e7415:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bellandi:2021:DMM

[BCC^{+}21] Valerio Bellandi, Paolo Ceravolo, Alessia Cristiano, Ernesto Damiani, Alberto Sanna, and Diana Trojaniello. A design methodology for matching smart health requirements. Concurrency and Computation: Practice and Experience, 33(22): e6062:1–e6062:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boda:2022:ABT

[BCK22] Ravi Boda, Reni K. Cherian, and Vinit Kumar. An automated brain tumor segmentation framework using a novel fruit fly UNet. Concurrency and Computation: Practice and Experience, 34(22):e7171:1–e7171:??, October 10, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baarab:2022:CBI

[BCM22] Nissrine Baarab, Brahim E. L. Khalil Chaouki, and Lhoussaine Masmoudi. Content-based image retrieval using color and a novel texture descriptor: orientational-based local binary pattern. Concurrency and Computation: Practice and Experience, 34(26):e7302:1–e7302:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Braganca:2023:FFC

[BCP^{+}23] Lucas Bragança, Michael Canesche, Jeronimo Penha, Josué Campos, José Augusto M. Nacif, and Ricardo S. Ferreira. Fast flow cloud: a stream dataflow framework for cloud FPGA accelerator overlays at runtime. Concurrency and Computation: Practice and Experience, 35(17):e6977:1–e6977:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhat:2021:CAT

[BD21] Parnika Bhat and Kamlesh Dutta. CogramDroid — an approach towards malware detection in Android using opcode ngrams. Concurrency and Computation: Practice and Experience, 33(20):e6332:1–e6332:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balakrishnan:2022:FSH

[BD22] Kulanthaivel Balakrishnan and Ramasamy Dhanalakshmi. Feature selection in high-dimensional microarray cancer datasets using an improved equilibrium optimization approach. Concurrency and Computation: Practice and Experience, 34(28):e7381:1–e7381:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bistarelli:2023:SMA

[BDG^{+}23] Stefano Bistarelli, Rocco De Nicola, Letterio Galletta, Cosimo Laneve, Ivan Mercanti, and Adele Veschetti. Stochastic modeling and analysis of the bitcoin protocol in the presence of block communication delays. Concurrency and Computation: Practice and Experience, 35(16):e6749:1–e6749:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balakrishnan:2022:EMP

[BDK22] Kulanthaivel Balakrishnan, Ramasamy Dhanalakshmi, and Utkarsh Mahadeo Khaire. Excogitating marine predators algorithm based on random opposition-based learning for feature selection. Concurrency and Computation: Practice and Experience, 34(4):e6630:1–e6630:??, February 15, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balasubramanian:2023:CWB

[BDR23] Kishore Balasubramanian, K. Gayathri Devi, and K. Ramya. Classification of white blood cells based on modified U-Net and SVM. Concurrency and Computation: Practice and Experience, 35(28):e7862:1–e7862:??, December 25, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Beibei:2024:RDE

[Bei24] Wu Beibei. Research on denoising and enhancing methods of medical images based on convolutional neural networks. Concurrency and Computation: Practice and Experience, 36(11): e8016:1–e8016:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Beamonte:2022:ETB

[BEJD22] Raphael Beamonte, Naser Ezzati-Jivan, and Michel R. Dagenais. Execution trace-based model verification to analyze multicore and real-time systems. Concurrency and Computation: Practice and Experience, 34(17):e6974:1–e6974:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bal:2022:IBL

[BEKS22] Burak Bal, Tugba Erdem, Seda Kul, and Ahmet Sayar. Image-based locating and guiding for unmanned aerial vehicles using scale invariant feature transform, speeded-up robust features, and oriented fast and rotated brief algorithms. Concurrency and Computation: Practice and Experience, 34(9): e6766:1–e6766:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benmammar:2021:NFN

[Ben21] Badr Benmammar. A new fast k-nearest neighbor classification algorithm in cognitive radio networks based on parallel computing. Concurrency and Computation: Practice and Experience, 33(5):e6027:1–e6027:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benmammar:2022:QSO

[Ben22] Badr Benmammar. Quality of service optimization in orthogonal frequency division multiplexing-based cognitive radio systems based on shuffled frog leaping algorithm. Concurrency and Computation: Practice and Experience, 34(1): e6530:1–e6530:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Batyrow:2023:SDC

[BEÖ23] Merdan Batyrow, Ilknur Eruçar, and Hande Öztürk. Size dependent change of mean square displacement in gold nanocrystals: a molecular dynamics simulation. Concurrency and Computation: Practice and Experience, 35(24):e7566:1– e7566:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Beeharry:2022:HAU

[BF22] Yogesh Beeharry and Ristin Tsokizep Fokone. Hybrid approach using machine learning algorithms for customers’ churn prediction in the telecommunications industry. Concurrency and Computation: Practice and Experience, 34(4): e6627:1–e6627:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bouabdallah:2024:EOA

[BF24] Raouia Bouabdallah and Fairouz Fakhfakh. Extending OCCI for an autonomous SLA negotiation strategy in federated clouds. Concurrency and Computation: Practice and Experience, 36(1):e7880:1–e7880:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bentes:2021:NPP

[BFM^{+}21] Cristiana Barbosa Bentes, Felipe M. G. França, Leandro Augusto Justen Marzulo, Gabriele Mencagli, and Mauricio Lima Pilla. Novel parallel processing techniques for IoT-based machine learning applications. Concurrency and Computation: Practice and Experience, 33(11):e6255:1–e6255:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Berger:2023:AES

[BFM^{+}23] Gonzalo Berger, Manuel Freire, Renzo Marini, Ernesto Dufrechou, and Pablo Ezzatti. Advancing on an efficient sparse matrix multiplication kernel for modern GPUs. Concurrency and Computation: Practice and Experience, 35(20):e7271:1– e7271:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benet:2024:FTI

[BFR24] Luis Benet, Luca Ferranti, and Nathalie Revol. A framework to test interval arithmetic libraries and their IEEE 1788-2015 compliance. Concurrency and Computation: Practice and Experience, 36(1):e7856:1–e7856:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balasubramanian:2024:EAS

[BGG24] Jaishankar Balasubramanian, Bharathi Gururaj, and Nandam Gayatri. An effective autism spectrum disorder screening method using machine learning classification techniques. Concurrency and Computation: Practice and Experience, 36 (2):e7898:1–e7898:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bujari:2021:WSN

[BGPQ21] Armir Bujari, Ombretta Gaggi, Claudio E. Palazzi, and Giacomo Quadrio. Wearable sensor networks: a measurement study. Concurrency and Computation: Practice and Experience, 33(8):e5939:1–e5939:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bordin:2023:ISC

[BI23] Jacir Luiz Bordin and Yasuaki Ito. International Symposium on Computing and Networking (CANDAR 2019) special issue. Concurrency and Computation: Practice and Experience, 35(11):e6634:1–e6634:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhaskar:2020:MBM

[BJGF20] Bhattu Bhaskar, Chandrashekar Jatoth, G. R. Gangadharan, and Ugo Fiore. A MapReduce-based modified grey wolf optimizer for QoS-aware big service composition. Concurrency and Computation: Practice and Experience, 32(8):e5351:1– e5351:??, April 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhatt:2024:ANN

[BJK^{+}24] Harsh Bhatt, Nilesh Kumar Jadav, Aparna Kumari, Rajesh Gupta, Sudeep Tanwar, Zdzislaw Polkowski, Amr Tolba, and Azza S. Hassanein. Artificial neural network-driven federated learning for heart stroke prediction in healthcare 4.0 underlying 5G. Concurrency and Computation: Practice and Experience, 36(3):e7911:1–e7911:??, February 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bingu:2023:IMD

[BJS23] Rajesh Bingu, S. Jothilakshmi, and N. Srinivasu. An intelligent multiclass deep classifier-based intrusion detection system for cloud environment. Concurrency and Computation: Practice and Experience, 35(26):e7840:1–e7840:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bo:2020:BTS

[BJWY20] Lili Bo, Shujuan Jiang, Rongcun Wang, and Qiao Yu. A bidirectional trace simplification approach based on a context switch linked list for concurrent programs. Concurrency and Computation: Practice and Experience, 32(2):e5423:1– e5423:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Batiha:2021:DAE

[BK21] Tarek Batiha and Pavel Krömer. Design and analysis of efficient neural intrusion detection for wireless sensor networks. Concurrency and Computation: Practice and Experience, 33(23):e6152:1–e6152:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bala:2022:TSA

[BK22a] Kirti Bala and Pankaj Deep Kaur. Transparent subsidized agri-product distribution during pandemics with reputation based PoA blockchain. Concurrency and Computation: Practice and Experience, 34(22):e6863:1–e6863:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhushan:2022:NEL

[BK22b] Shashi Bhushan and Anoop Kumar. New efficient logarithmic estimators using multi-auxiliary information under ranked set sampling. Concurrency and Computation: Practice and Experience, 34(27):e7337:1–e7337:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bokade:2022:AMB

[BK22c] Gayatri Umakant Bokade and Rajendra D. Kanphade. An ArmurMimus multimodal biometric system for Khosher authentication. Concurrency and Computation: Practice and Experience, 34(18):e7011:1–e7011:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhushan:2023:NPE

[BK23] Shashi Bhushan and Anoop Kumar. Novel predictive estimators using ranked set sampling. Concurrency and Computation: Practice and Experience, 35(3):e7435:1–e7435:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Belkhiri:2022:EMS

[BKD22] Youcef Belkhiri, Nadjet Kamel, and Habiba Drias. An efficient multi-swarm elephant herding optimization for solving community detection problem in complex environment. Concurrency and Computation: Practice and Experience, 34 (3):e6590:1–e6590:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhagat:2023:LCA

[BKGC23] Vijesh Bhagat, Santosh Kumar, Sachin Kumar Gupta, and Mithilesh Kumar Chaube. Lightweight cryptographic algorithms based on different model architectures: a systematic review and futuristic applications. Concurrency and Computation: Practice and Experience, 35(1):e7425:1–e7425:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barbhuiya:2022:GRR

[BKJ22] Abul Abbas Barbhuiya, Ram Kumar Karsh, and Rahul Jain. Gesture recognition from RGB images using convolutional neural network-attention based system. Concurrency and Computation: Practice and Experience, 34(24):e7230:1– e7230:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balasubramanian:2022:OKO

[BKK22] Kishore Balasubramanian, Ramya Kishore, and Gayathri Devi Krishnamoorthy. Optimal knee osteoarthritis diagnosis using hybrid deep belief network based on Salp swarm optimization method. Concurrency and Computation: Practice and Experience, 34(13):e6913:1–e6913:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhumireddypalli:2023:ECR

[BKK23] Veera Sekhar Reddy Bhumireddypalli, Srinivas Rao Koppula, and Neeraja Koppula. Enhanced conditional random fieldlong short-term memory for name entity recognition in English texts. Concurrency and Computation: Practice and Experience, 35(9):e7640:1–e7640:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bok:2020:PCR

[BKLY20] Kyoungsoo Bok, Geonsik Ko, Jongtae Lim, and Jaesoo Yoo. Personalized content recommendation scheme based on trust in online social networks. Concurrency and Computation: Practice and Experience, 32(18):e5572:1–e5572:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Belcastro:2021:PER

[BKM^{+}21] Loris Belcastro, M. Tahar Kechadi, Fabrizio Marozzo, Luca Pastore, Domenico Talia, and Paolo Trunfio. Parallel extraction of regions-of-interest from social media data. Concurrency and Computation: Practice and Experience, 33(8): e5638:1–e5638:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bode:2024:BWS

[BKM^{+}24] Brett Bode, David King, Celso L. Mendes, William T. Kramer, Saurabh Jha, Roger Ford, Justin Davis, and Steven Dramstad. Blue Waters system and component reliability. Concurrency and Computation: Practice and Experience, 36 (8):e7978:1–e7978:??, April 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhavadharini:2022:EEP

[BKS22] Reddipalayam Murugeshan Bhavadharini, Subburathinam Karthik, and Rajagopal Sabitha. An energy-efficient prioritybased packet scheduling mechanism for enhancing quality of service in mobile ad hoc network. Concurrency and Computation: Practice and Experience, 34(6):e6784:1–e6784:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Booth:2022:ASS

[BL22] Joshua Dennis Booth and Phillip Allen Lane. An adaptive self-scheduling loop scheduler. Concurrency and Computation: Practice and Experience, 34(6):e6750:1–e6750:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balis:2024:IPC

[BLB^{+}24] Bartosz Balis, Tomasz Lelek, Jakub Bodera, Michal Grabowski, and Costin Grigoras. Improving prediction of computational job execution times with machine learning. Concurrency and Computation: Practice and Experience, 36(2):e7905:1– e7905:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bae:2020:RDU

[BLI20] Seong Il Bae, Gyu Bin Lee, and Eul Gyu Im. Ransomware detection using machine learning algorithms. Concurrency and Computation: Practice and Experience, 32(18):e5422:1– e5422:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Belkadi:2022:FTB

[BLT22] Khadidja Belkadi, Mohamed Lehsaini, and Mohammed Amin Tahraoui. Fault-tolerance based on augmenting approach in wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(28):e7359:1–e7359:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

B:2022:ETL

[BM22a] Ramana Reddy B and Indiramma M. Efficient throttled load balancing algorithm to improve the response time and processing time in data center. Concurrency and Computation: Practice and Experience, 34(23):e7208:1–e7208:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bethi:2022:ENN

[BM22b] Sravankumar Bethi and Nageswara Rao Moparthi. Effective neighbor node discovery using improved Bayesian clock synchronization-gossip routing protocol in mobile low duty cycle wireless sensor network. Concurrency and Computation: Practice and Experience, 34(11):e6880:1–e6880:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhargavi:2022:IEH

[BM22c] Vijayan Sharmila Bhargavi and Soundararajan Pavai Madheswari. IoT enabled healthcare system for predicting the diseases using feature optimization, decision tree, neural network, and fuzzy temporal rules. Concurrency and Computation: Practice and Experience, 34(27):e7327:1–e7327:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Badvath:2022:PSD

[BMcKGK22] Dhanalaxmi Badvath, Aruna Safali Miriyala, Sai chaitanya Kumar Gunupudi, and Parish Venkata Kumar Kuricheti. Prediction of software defects using deep learning with improved cuckoo search algorithm. Concurrency and Computation: Practice and Experience, 34(26):e7305:1–e7305:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhavani:2022:RUN

[BMG22] Murapaka Dhanalakshmi Bhavani, Raman Murugan, and Tripti Goel. Robust U-Net: Development of robust image enhancement model using modified U-Net architecture. Concurrency and Computation: Practice and Experience, 34(28): e7347:1–e7347:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bull:2020:FAP

[BMJ^{+}20] Peter Bull, Stephen Murphy, Nelson Bruno Junior, Ron Austin, and Mak Sharma. A flow analysis and preemption framework for periodic traffic in an SDN network. Concurrency and Computation: Practice and Experience, 32(1): e4531:1–e4531:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barreiros:2020:OPS

[BMK^{+}20] Willian Barreiros, Jr., Jeremias Moreira, Tahsin Kurc, Jun Kong, Alba C. M. A. Melo, Joel H. Saltz, and George Teodoro. Optimizing parameter sensitivity analysis of large-scale microscopy image analysis workflows with multilevel computation reuse. Concurrency and Computation: Practice and Experience, 32(2):e5403:1–e5403:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boulakbech:2023:CAP

[BMSD23] Marwa Boulakbech, Nizar Messai, Yacine Sam, and Thomas Devogele. Configuration approach for personalized travel mashup. Concurrency and Computation: Practice and Experience, 35(11):e6811:1–e6811:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Butta:2022:CSO

[BMV22] Rajasekhar Butta, Kamaraju Maddu, and Sumalatha Vangala. Cat swarm optimized ensemble technique for emotion recognition in speech signals. Concurrency and Computation: Practice and Experience, 34(27):e7319:1–e7319:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bami:2022:DZD

[BMZ^{+}22] Hamid Gavari Bami, Elaheh Moharamkhani, Behrouz Zadmehr, Vahid Najafpoor, and Mohammad Shokouhifar. Detection of zero-day attacks in computer networks using combined classification. Concurrency and Computation: Practice and Experience, 34(27):e7312:1–e7312:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bushehrian:2021:ACE

[BN21] Omid Bushehrian and Seyyed Yahya Nabavi. An adaptive and cost-efficient migration to cloud approach in dynamic environments. Concurrency and Computation: Practice and Experience, 33(17):e6309:1–e6309:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Burger:2022:SPR

[BNB22] Michael Burger, Giang Nam Nguyen, and Christian Bischof. SimAnMo-A parallelized runtime model generator. Concurrency and Computation: Practice and Experience, 34(20): e6771:1–e6771:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bock:2021:PSE

[Boc21] Alexander Asp Bock. Parallel spreadsheet evaluation and dynamic cycle detection. Concurrency and Computation: Practice and Experience, 33(13):e6218:1–e6218:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benassy:2023:ECD

[BOI23] Grégory Bénassy, Fukuhito Ooshita, and Michiko Inoue. Eventually consistent distributed ledger despite degraded atomic broadcast. Concurrency and Computation: Practice and Experience, 35(11):e6199:1–e6199:??, May 15, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bozkurt:2022:UCC

[Boz22a] Erkam Murat Bozkurt. The usage of cybernetic in complex software systems and its application to the deterministic multithreading. Concurrency and Computation: Practice and Experience, 34(28):e7375:1–e7375:??, December 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bozkurt:2022:DHF

[Boz22b] Ferhat Bozkurt. A deep and handcrafted features-based framework for diagnosis of COVID-19 from chest X-ray images. Concurrency and Computation: Practice and Experience, 34(5):e6725:1–e6725:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balaiah:2020:ACP

[BP20] Thanasekhar Balaiah and Ranjani Parthasarathi. Autotuning of configuration for program execution in GPUs. Concurrency and Computation: Practice and Experience, 32(9):e5635:1– e5635:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bai:2023:TPM

[BP23a] Gnana Prakasam Mercy Bai and Venkadesh Perumal. Taylor political monarch butterfly optimization driven deep learning model for acute lymphoblastic leukemia detection and severity classification using blood smear images. Concurrency and Computation: Practice and Experience, 35(2):e7455:1– e7455:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bondre:2023:RAA

[BP23b] Shweta Bondre and Dipti Patil. Recent advances in agricultural disease image recognition technologies: a review. Concurrency and Computation: Practice and Experience, 35(9): e7644:1–e7644:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baskar:2024:MCC

[BP24] Premalatha Baskar and Prakasam Periasamy. Minimal channel cost-based energy-efficient resource allocation algorithm for task offloading under FoG computing environment. Concurrency and Computation: Practice and Experience, 36(7): e7968:1–e7968:??, March 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Baya:2020:ACA

[BPAE20] Rodrigo Bayá, Martín Pedemonte, Alejandro Gutiérrez Arce, and Pablo Ezzatti. An asynchronous computation architecture for enhancing the performance of the weather research and forecasting model. Concurrency and Computation: Practice and Experience, 32(19):e5750:1–e5750:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brinkschulte:2021:IFO

[BPB21] Uwe Brinkschulte, Mathias Pacher, and Melanie Brinkschulte. Improving the fail-operational behavior of automotive applications by artificial DNA. Concurrency and Computation: Practice and Experience, 33(14):e5608:1–e5608:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bystrov:2024:CPA

[BPK24] Oleg Bystrov, Ruslan Pacevic, and Arnas Kaceniauskas. Costand performance-aware resource selection for parallel software on heterogeneous cloud. Concurrency and Computation: Practice and Experience, 36(10):e7877:1–e7877:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhattacharya:2022:PMI

[BPM^{+}22] Tathagata Bhattacharya, Xiaopu Peng, Jianzhou Mao, Chaowei Zhang, Taha Takreeti, Ye Wang, Ting Cao, and Xiao Qin. Performance modeling for I/O-intensive applications on virtual machines. Concurrency and Computation: Practice and Experience, 34(10):e6823:1–e6823:??, May 1, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barclay:2023:PAS

[BPT^{+}23] Iain Barclay, Alun Preece, Ian Taylor, Swapna Krishnakumar Radha, and Jarek Nabrzyski. Providing assurance and scrutability on shared data and machine learning models with verifiable credentials. Concurrency and Computation: Practice and Experience, 35(18):e6997:1–e6997:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bao:2020:PCM

[BPW^{+}20] Wei Bao, Chuang Ye Pan, Lu Wang, Qing Zhang, and Neng Gang Xie. Parallel computing modes of multiobjective game. Concurrency and Computation: Practice and Experience, 32(14):e5687:1–e5687:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barati:2021:PAC

[BR21] Masoud Barati and Omer Rana. Privacy-aware cloud ecosystems: Architecture and performance. Concurrency and Computation: Practice and Experience, 33(23):e5852:1–e5852:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brown:2020:MLC

[BRL^{+}20] Nick Brown, Anna Roubícková, Ioanna Lampaki, Lucy Mac-Gregor, Michelle Ellis, and Paola Vera de Newton. Machine learning on Crays to optimize petrophysical workflows in oil and gas exploration. Concurrency and Computation: Practice and Experience, 32(20):e5655:1–e5655:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhimavarapu:2023:CNN

[BRNR23] John Philip Bhimavarapu, Sriharsha Ramaraju, Dimmita Nagajyothi, and Inumula Veeraraghava Rao. Convolutional neural network based object detection system for video surveillance application. Concurrency and Computation: Practice and Experience, 35(3):e7461:1–e7461:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhagavathi:2022:IBS

[BRS^{+}22] Hariharan Bhagavathi, Siva Rathinavelayatham, Kaliraj Shanmugaiah, Kamaraj Kanagaraj, and Dinesh Elangovan. Improved beetle swarm optimization algorithm for energy efficient virtual machine consolidation on cloud environment. Concurrency and Computation: Practice and Experience, 34 (10):e6828:1–e6828:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bist:2022:ARA

[BS22] Uttam Singh Bist and Nanhay Singh. Analysis of recent advancements in support vector machine. Concurrency and Computation: Practice and Experience, 34(25):e7270:1– e7270:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barbara:2023:MHE

[BS23a] Fadi Barbàra and Claudio Schifanella. MP-HTLC: Enabling blockchain interoperability through a multiparty implementation of the hash time-lock contract. Concurrency and Computation: Practice and Experience, 35(9):e7656:1–e7656:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bawaneh:2023:NTC

[BS23b] Mohammad Bawaneh and Vilmos Simon. Novel traffic congestion detection algorithms for smart city applications. Concurrency and Computation: Practice and Experience, 35(5): e7563:1–e7563:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boyanapalli:2023:OCD

[BS23c] Arathi Boyanapalli and A. Shanthini. Ovarian cancer detection in computed tomography images using ensembled deep optimized learning classifier. Concurrency and Computation: Practice and Experience, 35(22):e7716:1–e7716:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Boulaares:2022:PAU

[BSBF22] Soura Boulaares, Salma Sassi, Djamal BenSlimane, and Sami Faiz. A probabilistic approach: Uncertain navigation of the uncertain web. Concurrency and Computation: Practice and Experience, 34(23):e7194:1–e7194:??, October 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benabidallah:2020:SSS

[BSEN20] Rymel Benabidallah, Salah Sadou, Armel Esnault, and Mohamed Ahmed Nacer. Simulating systems of systems using situation/reaction paradigm. Concurrency and Computation: Practice and Experience, 32(15):e4921:1–e4921:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Braines:2021:SLC

[BSML21] Dave Braines, Jane Stockdill-Mander, and Eunjin Lee. The science library: Curation and visualization of a science gateway repository. Concurrency and Computation: Practice and Experience, 33(19):e6100:1–e6100:??, October 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Babu:2022:MRS

[BSN22] Rathish Babu Thirukonda Krishnamoorthy Sivakumar Babu, Suresh Sivasubramanian, and Sankarram Natarajan. MLPNN-RF: Software fault prediction based on robust weight based optimization and Jacobian adaptive neural network. Concurrency and Computation: Practice and Experience, 34(21): e7122:1–e7122:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bilbao:2023:FSS

[BSPM23] Carlos Bilbao, Juan Carlos Saez, and Manuel Prieto-Matias. Flexible system software scheduling for asymmetric multicore systems with PMCSched: a case for Intel Alder Lake. Concurrency and Computation: Practice and Experience, 35(25): e7814:1–e7814:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bai:2022:ISS

[BST^{+}22] Dongxu Bai, Ying Sun, Bo Tao, Xiliang Tong, Manman Xu, Guozhang Jiang, Baojia Chen, Yongcheng Cao, Nannan Sun, and Zeshen Li. Improved single shot multibox detector target detection method based on deep feature fusion. Concurrency and Computation: Practice and Experience, 34(4):e6614:1– e6614:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bartuzi:2021:MHF

[BT21] Ewelina Bartuzi and Mateusz Trokielewicz. Multispectral hand features for secure biometric authentication systems. Concurrency and Computation: Practice and Experience, 33(18):e6471:1–e6471:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Banach:2020:NTS

[BTDD20] Marzena Banach, Tomasz Talaśka, Jakub Dalecki, and Rafał Długosz. New technologies for smart cities — high-resolution air pollution maps based on intelligent sensors. Concurrency and Computation: Practice and Experience, 32(13):e5179:1– e5179:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Borgaonkar:2021:ISG

[BTDJ21] Ravishankar Borgaonkar, Inger Anne Tøndel, Merkebu Zenebe Degefa, and Martin Gilje Jaatun. Improving smart grid security through 5G enabled IoT and edge computing. Concurrency and Computation: Practice and Experience, 33(18): e6466:1–e6466:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barclay:2021:FFT

[BTP^{+}21] Iain Barclay, Harrison Taylor, Alun Preece, Ian Taylor, Dinesh Verma, and Geeth de Mel. A framework for fostering transparency in shared artificial intelligence models by increasing visibility of contributions. Concurrency and Computation: Practice and Experience, 33(19):e6129:1–e6129:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Binh:2021:OCB

[BTT21] Nguyen Thanh Binh, Vo Thi Hong Tuyet, and Nguyen Chi Thanh. Object contour based on improving region growing method and context border condition. Concurrency and Computation: Practice and Experience, 33(2):e5472:1–e5472:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bulut:2022:SSB

[Bul22] Vahide Bulut. SP-search-based path planning algorithm for mobile robots using quintic trigonometric Bézier curves. Concurrency and Computation: Practice and Experience, 34(9): e6493:1–e6493:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Balakrishnan:2022:LRF

[BV22a] Mathiarasi Balakrishnan and Geetha Thekkumpurath Varieth. A learning to rank framework for future friendship ranking in streaming heterogeneous social networks. Concurrency and Computation: Practice and Experience, 34(3):e6592:1– e6592:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bandaru:2022:BCE

[BV22b] Venkata Naga Rani Bandaru and P Visalakshi. Block chain enabled auditing with optimal multi-key homomorphic encryption technique for public cloud computing environment. Concurrency and Computation: Practice and Experience, 34 (22):e7128:1–e7128:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brahmavar:2022:MHU

[BVM22] Anup Brahmavar, Harish Venkatarama, and Geetha Maiya. Mining high utility itemsets with time-aware scheduling using Apache Spark. Concurrency and Computation: Practice and Experience, 34(23):e7192:1–e7192:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhowmick:2022:SMN

[BVP22] Anwesha Bhowmick, Sathish Vadhiyar, and Varun Pv. Scalable multi-node multi-GPU Louvain community detection algorithm for heterogeneous architectures. Concurrency and Computation: Practice and Experience, 34(17):e6987:1– e6987:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bhatia:2020:TSI

[BVS20] Tarunpreet Bhatia, A. K. Verma, and Gaurav Sharma. Towards a secure incremental proxy re-encryption for ehealthcare data sharing in mobile cloud computing. Concurrency and Computation: Practice and Experience, 32(5): e5520:1–e5520:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brand:2021:PME

[BWS^{+}21] Marcel Brand, Michael Witterauf, Éricles Sousa, Alexandru Tanase, Frank Hannig, and Jürgen Teich. *-predictable MP-SoC execution of real-time control applications using invasive computing. Concurrency and Computation: Practice and Experience, 33(14):e5149:1–e5149:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Benkhelifa:2020:ANU

[BWTJ20] Elhadj Benkhelifa, Thomas Welsh, Lo’ai Tawalbeh, and Yaser Jararweh. Automated negotiated user profiling across distributed social mobile clouds for resource optimisation. Concurrency and Computation: Practice and Experience, 32(1): e4349:1–e4349:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bai:2020:NDL

[BWW^{+}20] Tian Bai, Chunyu Wang, Ye Wang, Lan Huang, and Fuyong Xing. A novel deep learning method for extracting unspecific biomedical relation. Concurrency and Computation: Practice and Experience, 32(1):e5005:1–e5005:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bao:2023:NNL

[BXH^{+}23] HaiYong Bao, Zheng Xu, HaiBo Hong, QingLei Kong, and HaiFeng Qian. NLSP: a novel lattice-based secure primitive for privacy-preserving smart grid communications. Concurrency and Computation: Practice and Experience, 35(19): e7406:1–e7406:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bo:2020:FCR

[BYL20] Wang Bo, Wang Ying, and Cui Lijie. Fuzzy clustering recognition algorithm of medical image with multi-resolution feature. Concurrency and Computation: Practice and Experience, 32 (1):e4886:1–e4886:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bae:2021:NDP

[BYPO21] Minho Bae, Sangho Yeo, Gyudong Park, and Sangyoon Oh. Novel data-placement scheme for improving the data locality of Hadoop in heterogeneous environments. Concurrency and Computation: Practice and Experience, 33(18):e5752:1– e5752:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bourjandi:2022:CDC

[BYTG22] Masoumeh Bourjandi, Meisam Yadollahzadeh-Tabari, and Mehdi Golsorkhtabaramiri. Combined deep centralized coordinate learning and hybrid loss for human activity recognition. Concurrency and Computation: Practice and Experience, 34(22):e6870:1–e6870:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bakhouya:2020:CCI

[BZEM20] Mohamed Bakhouya, Mostapha Zbakh, Mohamed Essaaidi, and Pierre Manneback. Cloud computing, IoT, and big data: Technologies and applications. Concurrency and Computation: Practice and Experience, 32(17):e5896:1–e5896:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bittibssi:2022:ISE

[BZGM22] Tarek M. Bittibssi, Abdelhalim Zekry, Mohamed A. Genedy, and Shady A. Maged. Implementation of surface electromyography controlled prosthetics limb based on recurrent neural network. Concurrency and Computation: Practice and Experience, 34(11):e6848:1–e6848:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

BrundoUriarte:2021:DSL

[BZK^{+}21] Rafael Brundo Uriarte, Huan Zhou, Kyriakos Kritikos, Zeshun Shi, Zhiming Zhao, and Rocco De Nicola. Distributed service-level agreement management with smart contracts and blockchain. Concurrency and Computation: Practice and Experience, 33(14):e5800:1–e5800:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bugingo:2022:DCC

[BZL^{+}22] Emmanuel Bugingo, Wei Zheng, Zhenfeng Lei, Defu Zhang, Samuel Rene Adolphe Sebakara, and Dongzhan Zhang. Deadline-constrained cost-energy aware workflow scheduling in cloud. Concurrency and Computation: Practice and Experience, 34(6):e6761:1–e6761:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bencheikh:2023:RAM

[BZT^{+}23a] Mohammed Bencheikh, Chakib Zouaoui, Nasreddine Taleb, Aoued Boukelif, Afaf Benaouda Chaht, and Mohamed Naimi. A reliable alternating multi-path transmission control protocol scheduling for bandwidth aggregation performance. Concurrency and Computation: Practice and Experience, 35(23): e7764:1–e7764:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bouhouch:2023:DDR

[BZT23b] Laila Bouhouch, Mostapha Zbakh, and Claude Tadonki. Dynamic data replication and placement strategy in geographically distributed data centers. Concurrency and Computation: Practice and Experience, 35(14):e6858:1–e6858:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bai:2021:OPA

[BZWH21] Enci Bai, Weizhe Zhang, Yulong Wu, and Hui He. Optimal priority assignment for messages on controller area network with maximum system robustness. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chinnaraj:2022:NMS

[CA22] Govindasamy Chinnaraj and Arokiasamy Antonidoss. A new methodology for secured inventory management by average fitness-based colliding bodies optimization integrated with block chain under cloud. Concurrency and Computation: Practice and Experience, 34(1):e6540:1–e6540:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cyril:2023:WSO

[CA23] Gnanasamy Lazar Sindhuraj Infant Cyril and John Patrick Ananth. Whale social optimization driven deep recurrent neural network for loan eligibility prediction. Concurrency and Computation: Practice and Experience, 35(3):e7510:1– e7510:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chehimi:2023:MLB

[CAAHC23] Mahdi Chehimi, Mohamad Khattar Awad, Mohammed Al-Husseini, and Ali Chehab. Machine learning-based antijamming technique at the physical layer. Concurrency and Computation: Practice and Experience, 35(9):e7649:1– e7649:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chandra:2024:ODP

[CAS24] Saurabh Chandra, Rajeev Arya, and Maheshwari Prasad Singh. Optimal D2D power for secure D2D communication with random eavesdropper in 5G-IoT networks. Concurrency and Computation: Practice and Experience, 36(9):e7993:1– e7993:??, April 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chinnasamy:2022:TCR

[CB22] Rajagopal Chinnasamy and Thangavel Balasubramanian. Tucker’s congruence regressive feature projected Tversky discriminant multiple instance learning boost data classification for school student dropout prediction. Concurrency and Computation: Practice and Experience, 34(18):e7021:1–e7021:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2023:MRA

[CBFS23] Liqiong Chen, Xiaoyu Bi, Guoqing Fan, and Huaiying Sun. A multitask recommendation algorithm based on DeepFM and graph convolutional network. Concurrency and Computation: Practice and Experience, 35(2):e7498:1–e7498:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chunka:2023:SCU

[CBK23] Chukhu Chunka, Subhasish Banerjee, and Gupta Sachin Kumar. A secure communication using multifactor authentication and key agreement techniques in internet of medical things for COVID-19 patients. Concurrency and Computation: Practice and Experience, 35(7):e7602:1–e7602:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chakravarthy:2022:SRM

[CBR22] S. R. Sannasi Chakravarthy, N. Bharanidharan, and Harikumar Rajaguru. A systematic review on machine learning algorithms used for forecasting lake-water level fluctuations. Concurrency and Computation: Practice and Experience, 34(24): e7231:1–e7231:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cengil:2022:HCN

[CÇ22a] Emine Cengil and Ahmet Çinar. Hybrid convolutional neural network based classification of bacterial, viral, and fungal diseases on tomato leaf images. Concurrency and Computation: Practice and Experience, 34(4):e6617:1–e6617:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chandran:2022:EES

[CC22b] K. R. Sarath Chandran and Premanand Venkatesh Chandramani. Energy-efficient system-on-chip reconfigurable architecture design for sum of absolute difference computation in motion estimation process of H.265/HEVC video encoding. Concurrency and Computation: Practice and Experience, 34(8): e5461:1–e5461:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cicioglu:2022:HMS

[CÇ22c] Murtaza Cicioglu and Ali Çalhan. Handover management in software-defined 5G small cell networks via long shortterm memory. Concurrency and Computation: Practice and Experience, 34(10):e6832:1–e6832:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carvalho:2023:DMI

[CCBA23] Ranyelson N. Carvalho, Lucas R. Costa, Jacir L. Bordim, and Eduardo A. P. Alchieri. DataPlane-ML: an integrated attack detection and mitigation solution for software defined networks. Concurrency and Computation: Practice and Experience, 35(19):e7434:1–e7434:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Caniou:2021:BAS

[CCCR21] Yves Caniou, Eddy Caron, Aurélie Kong Win Chang, and Yves Robert. Budget-aware scheduling algorithms for scientific workflows with stochastic task weights on infrastructure as a service cloud platforms. Concurrency and Computation: Practice and Experience, 33(17):e6065:1–e6065:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carpentieri:2020:CBS

[CCD^{+}20a] Bruno Carpentieri, Arcangelo Castiglione, Alfredo De Santis, Francesco Palmieri, and Raffaele Pizzolante. Compressionbased steganography. Concurrency and Computation: Practice and Experience, 32(8):e5322:1–e5322:??, April 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Clarizia:2020:MGA

[CCD^{+}20b] Fabio Clarizia, Francesco Colace, Massimo De Santo, Marco Lombardi, Francesco Pascale, Domenico Santaniello, and Allan Tuker. A multilevel graph approach for rainfall forecasting: a preliminary study case on London area. Concurrency and Computation: Practice and Experience, 32(8):e5289:1– e5289:??, April 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cuomo:2021:VAC

[CCdCC21] Salvatore Cuomo, Giovanni Colecchia, Vincenzo Schiano di Cola, and Ugo Chirico. A virtual assistant in cultural heritage scenarios. Concurrency and Computation: Practice and Experience, 33(3):e5331:1–e5331:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2020:WFD

[CCGN20] Weiming Chen, Siqi Chen, Hailin Guo, and Xiaoyang Ni. Welding flame detection based on color recognition and progressive probabilistic Hough transform. Concurrency and Computation: Practice and Experience, 32(19):e5815:1– e5815:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2022:GGB

[CCL^{+}22] Jinlin Chen, Jiannong Cao, Zhixuan Liang, Zhiqin Cheng, and Jia Wang. GraphWare: a graph-based middleware enabling multi-robot cooperation. Concurrency and Computation: Practice and Experience, 34(17):e6995:1–e6995:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2022:TIA

[CCM22a] Zhanwen Chen, Jiageng Chen, and Weizhi Meng. Threshold identity authentication signature: Impersonation prevention in social network services. Concurrency and Computation: Practice and Experience, 34(16):e5787:1–e5787:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cicioglu:2022:ERA

[CÇM22b] Murtaza Cicioglu, Ali Çalhan, and Md Sipon Miah. An effective routing algorithm for spectrum allocations in cognitive radio based internet of things. Concurrency and Computation: Practice and Experience, 34(28):e7368:1–e7368:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carpentieri:2021:UPG

[CCPP21] Bruno Carpentieri, Arcangelo Castiglione, Francesco Palmieri, and Raffaele Pizzolante. On the undetectability of payloads generated through automatic tools: a human-oriented approach. Concurrency and Computation: Practice and Experience, 33(19):e6351:1–e6351:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cengil:2022:HAE

[CÇY22] Emine Cengil, Ahmet Çinar, and Muhammed Yildirim. A hybrid approach for efficient multi-classification of white blood cells based on transfer learning techniques and traditional machine learning methods. Concurrency and Computation: Practice and Experience, 34(6):e6756:1–e6756:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2021:RFW

[CCZ^{+}21] Huipeng Chen, Sen Chen, Rougang Zhou, Xiaoyan Huang, and Shaopeng Zhu. Research on four-wheel independent steering intelligent control strategy based on minimum load. Concurrency and Computation: Practice and Experience, 33(9): e6145:1–e6145:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2023:DNN

[CCZM23] Juan Chen, Ying Chen, Han Zhao, and Tao Ma. Development of neural-network-based stereo bionic compound eyes with fiber bundles. Concurrency and Computation: Practice and Experience, 35(2):e7464:1–e7464:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Choudhary:2022:QSA

[CD22] Swapna Choudhary and Sanjay Dorle. A quality of serviceaware high-security architecture design for software-defined network powered vehicular ad-hoc networks using machine learning-based blockchain routing. Concurrency and Computation: Practice and Experience, 34(17):e6993:1–e6993:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cacciapuoti:2024:SAT

[CD24a] Rosalba Cacciapuoti and Luisa D’Amore. Scalability analysis of a two level domain decomposition approach in space and time solving data assimilation models. Concurrency and Computation: Practice and Experience, 36(10):e7937:1– e7937:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chauhan:2024:HOD

[CD24b] Nupa Ram Chauhan and Rakesh Kumar Dwivedi. Hybrid one-dimensional residual autoencoder and ensemble of gradient boosting for cloud IDS. Concurrency and Computation: Practice and Experience, 36(14):e8088:1–e8088:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2020:NIN

[CDC20] Yongqi Chen, Qinge Dai, and Yang Chen. A new interval native Bayes uncertain fault diagnosis method based on the firefly algorithm. Concurrency and Computation: Practice and Experience, 32(24):e5911:1–e5911:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cordasco:2021:TDS

[CDN^{+}21] Gennaro Cordasco, Matteo D’Auria, Alberto Negro, Vittorio Scarano, and Carmine Spagnuolo. Toward a domainspecific language for scientific workflow-based applications on multicloud system. Concurrency and Computation: Practice and Experience, 33(18):e5802:1–e5802:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cabral:2020:EMO

[CdOO^{+}20] Frederico L. Cabral, Sanderson L. Gonzaga de Oliveira, Carla Osthoff, Gabriel P. Costa, Diego N. Brandão, and Mauricio Kischinhevsky. An evaluation of MPI and OpenMP paradigms in finite-difference explicit methods for PDEs on shared-memory multi- and manycore systems. Concurrency and Computation: Practice and Experience, 32(20):e5642:1– e5642:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Conoci:2021:PCP

[CDP^{+}21] Stefano Conoci, Pierangelo Di Sanzo, Alessandro Pellegrini, Bruno Ciciani, and Francesco Quaglia. On power capping and performance optimization of multithreaded applications. Concurrency and Computation: Practice and Experience, 33 (13):e6205:1–e6205:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2023:ERI

[CDR^{+}23] Zhiwei Chen, Lunzhi Deng, Yu Ruan, Shuai Feng, Tao Wang, and Bo Wang. An efficient revocable identity-based encryption with ciphertext evolution in the cloud-assisted system. Concurrency and Computation: Practice and Experience, 35 (22):e7735:1–e7735:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Camacho:2023:POP

[CdRNB23] Thais Aparecida Silva Camacho, Vanderson Martins do Rosario, Otávio Oliveira Napoli, and Edson Borin. PB^{3} Opt: Profilebased biased Bayesian optimization to select computing clusters on the cloud. Concurrency and Computation: Practice and Experience, 35(18):e7540:1–e7540:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Champla:2023:DDD

[CDV23] Dharavath Champla, Sivakumar Dhandapani, and Nagarajan Velmurugan. DL-DC: Deep learning-based deadline constrained load balancing technique. Concurrency and Computation: Practice and Experience, 35(26):e7839:1–e7839:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cekim:2022:MUE

[Cek22] Hatice Oncel Cekim. Modified unbiased estimators for population variance: an application for COVID-19 deaths in Russia. Concurrency and Computation: Practice and Experience, 34(22):e7169:1–e7169:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chahardoli:2022:NBO

[CEN22] Meysam Chahardoli, Nafiseh Osati Eraghi, and Sara Nazari. Namib beetle optimization algorithm: a new meta-heuristic method for feature selection and dimension reduction. Concurrency and Computation: Practice and Experience, 34(1): e6524:1–e6524:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cetinkaya:2023:IBX

[Çet23]  Çagatay Çetinkaya. Inference of P (X > Y) for the Burr-

XII model under generalized progressive hybrid censored data with binomial removals. Concurrency and Computation: Practice and Experience, 35(7):e7609:1–e7609:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2021:CBD

[CF21] Feilong Cao and Fan Feng. Consensus-based distributed learning for robust convex optimization with a scenario approach. Concurrency and Computation: Practice and Experience, 33(8):e5049:1–e5049:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chalker:2021:OOS

[CFR^{+}21] Alan Chalker, Eric Franz, Morgan Rodgers, Trey Dockendorf, Doug Johnson, Doris Sajdak, Joseph P. White, Benjamin D. Plessinger, Mohammad Zia, Steven M. Gallo, Robert E. Settlage, and David E. Hudak. Open OnDemand: State of the platform, project, and the future. Concurrency and Computation: Practice and Experience, 33(19):e6114:1–e6114:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cicek:2021:FAR

[ÇG21] Emre Çiçek and Sezer Gören. Fully automated roadside parking spot detection in real time with deep learning. Concurrency and Computation: Practice and Experience, 33(23): e6006:1–e6006:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Choudhary:2022:DTC

[CG22] Monika Choudhary and Nitin Goyal. Dynamic topology control algorithm for node deployment in mobile underwater wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(15):e6942:1–e6942:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cimen:2023:CMO

[ÇGB23] Murat Erhan Çimen, Zeynep Garip, and Ali Fuat Boz. Comparison of metaheuristic optimization algorithms for numerical solutions of optimal control problems. Concurrency and Computation: Practice and Experience, 35(10):e7663:1– e7663:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2021:PME

[CGS^{+}21] Yongzhong Cao, Yalu Guo, Qiang She, Junwu Zhu, and Bin Li. Prediction of medical expenses for gastric cancer based on process mining. Concurrency and Computation: Practice and Experience, 34(1):e5694:1–e5694:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cai:2020:UHB

[CGW^{+}20] Xingjuan Cai, Shaojin Geng, Di Wu, Lei Wang, and Qidi Wu. A unified heuristic bat algorithm to optimize the LEACH protocol. Concurrency and Computation: Practice and Experience, 32(9):e5619:1–e5619:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chellouf:2021:PCA

[CH21] Maawya Chellouf and Tarek Hamrouni. Popularity and correlation aware data replication strategy based on half-life concept and clustering in cloud system. Concurrency and Computation: Practice and Experience, 33(10):e6159:1–e6159:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2023:R

[Che23] Yuan Cheng. Retraction. Concurrency and Computation: Practice and Experience, 30(24):e7748:1–e7748:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2024:MEA

[Che24] Lei Chen. Medical education and artificial intelligence: Question answering for medical questions based on intelligent interaction. Concurrency and Computation: Practice and Experience, 36(14):e8079:1–e8079:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Catalan:2023:FGT

[CHI^{+}23] Sandra Catalán, José R. Herrero, Francisco D. Igual, Enrique S. Quintana-Ortí, and Rafael Rodríguez-Sánchez. Finegrain task-parallel algorithms for matrix factorizations and inversion on many-threaded CPUs. Concurrency and Computation: Practice and Experience, 35(27):e6999:1–e6999:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2024:KDR

[CHL^{+}24] Buqing Cao, Hao Huang, Shanpeng Liu, Yizhi Liu, Yiping Wen, Dong Zhou, and Mingdong Tang. Knowledge distillation representation and DCNMIX quality prediction-based Web service recommendation. Concurrency and Computation: Practice and Experience, 36(12):e8038:1–e8038:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2023:SIT

[CHLD23] Jinchao Chen, Pengcheng Han, Yifan Liu, and Xiaoyan Du. Scheduling independent tasks in cloud environment based on modified differential evolution. Concurrency and Computation: Practice and Experience, 35(13):e6256:1–e6256:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2021:EDP

[CHMC21] Zheyi Chen, Jia Hu, Geyong Min, and Xing Chen. Effective data placement for scientific workflows in mobile edge computing using genetic particle swarm optimization. Concurrency and Computation: Practice and Experience, 33(8): e5413:1–e5413:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Choi:2020:DDD

[Cho20] Hyun-Sik Choi. Drowsy driving detection using neural network with backpropagation algorithm implemented by FPGA. Concurrency and Computation: Practice and Experience, 32(18):e5471:1–e5471:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Casalicchio:2020:SAC

[CI20] Emiliano Casalicchio and Stefano Iannucci. The state-of-the-art in container technologies: Application, orchestration and security. Concurrency and Computation: Practice and Experience, 32(17):e5668:1–e5668:??, September 10, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chehri:2021:OMB

[CJ21a] Abdellah Chehri and Gwanggil Jeon. Optimal matching between energy saving and traffic load for mobile multimedia communication. Concurrency and Computation: Practice and Experience, 33(4):e5035:1–e5035:??, February 25, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chehri:2021:RTM

[CJ21b] Abdellah Chehri and Gwanggil Jeon. Real-time multiuser scheduling based on end-user requirement using big data analytics. Concurrency and Computation: Practice and Experience, 33(4):e5021:1–e5021:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2022:AST

[CJC22] Mengpan Chen, Shunfu Jin, and Li Chen. Allocation strategy for time-sensitive tasks in mobile edge computing: an observable perspective. Concurrency and Computation: Practice and Experience, 34(24):e7225:1–e7225:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cleveland:2021:TVS

[CJP^{+}21] Sean B. Cleveland, Anagha Jamthe, Smruti Padhy, Joe Stubbs, Steven Terry, Julia Looney, Richard Cardone, Michael Packard, Maytal Dahan, and Gwen A. Jacobs. Tapis v3 Streams API: Time-series and data-driven event support in science gateway infrastructure. Concurrency and Computation: Practice and Experience, 33(19):e6103:1–e6103:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2020:REL

[CJY^{+}20] Xiaowu Chen, Guozhang Jiang, Le Yang, Gongfa Li, and Feng Xiang. Redesign of enterprise lean production system based on environmental dynamism. Concurrency and Computation: Practice and Experience, 32(14):e5706:1–e5706:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chanthiya:2021:FFD

[CK21a] P. Chanthiya and V. Kalaivani. Forest fire detection on LANDSAT images using support vector machine. Concurrency and Computation: Practice and Experience, 33(16): e6280:1–e6280:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2021:HNM

[CK21b] Ke Chen and Wende Ke. A hierarchical neural model for target-based sentiment analysis. Concurrency and Computation: Practice and Experience, 33(10):e6184:1–e6184:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chamkoori:2023:SSI

[CK23] Alireza Chamkoori and Serajdean Katebi. Security and storage improvement in distributed cloud data centers by increasing reliability based on particle swarm optimization and artificial immune system algorithms. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chaurasia:2024:MAM

[CK24a] Soni Chaurasia and Kamal Kumar. MH-ARO: Meta-heuristic based adaptive routing for large scale opportunistic networks. Concurrency and Computation: Practice and Experience, 36 (14):e8096:1–e8096:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chetouane:2024:RBI

[CK24b] Ameni Chetouane and Kamel Karoui. Risk based intrusion detection system in software defined networking. Concurrency and Computation: Practice and Experience, 36(9):e7988:1– e7988:??, April 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Calik:2022:NME

[ÇKÇ22] Elif Çalik, Hilal Kaya, and Fatih Vehbi Çelebi. A novel method to ensure the security of the shared medical data using smart contracts: Organ transplantation sample. Concurrency and Computation: Practice and Experience, 34(9): e6752:1–e6752:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chae:2020:IFS

[CKKK20] Dong-Kyu Chae, Bo-Kyum Kim, Seung-Ho Kim, and Sang-Wook Kim. Incremental feature selection for efficient classification of dynamic graph bags. Concurrency and Computation: Practice and Experience, 32(18):e5502:1–e5502:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Choi:2020:NAM

[CKL20] YunSeok Choi, Dahae Kim, and Jee-Hyong Lee. Neural attention model with keyword memory for abstractive document summarization. Concurrency and Computation: Practice and Experience, 32(18):e5433:1–e5433:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chhabra:2022:HOA

[CL22a] Surbhi Chhabra and Kusum Lata. Hardware obfuscation of AES IP core using combinational hardware Trojan circuit for secure data transmission in IoT applications. Concurrency and Computation: Practice and Experience, 34(21):e7058:1– e7058:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chikh:2022:CGC

[CL22b] Asma Chikh and Mohamed Lehsaini. Combination of greedy and compass approaches for efficient multipath geographic routing in wireless multimedia sensor networks. Concurrency and Computation: Practice and Experience, 34(5):e6703:1– e6703:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2023:CBL

[CL23] Ben Cao and Jianxun Liu. Combining bidirectional long short-term memory and self-attention mechanism for code search. Concurrency and Computation: Practice and Experience, 35(10):e7662:1–e7662:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2020:TMP

[CLC^{+}20] Xin Chen, Zhuo Li, Ying Chen, Bing Du, and Yongchao Zhang. Traffic modeling and performance evaluation of SDNbased NB-IoT access network. Concurrency and Computation: Practice and Experience, 32(16):e5145:1–e5145:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2023:ICA

[CLC^{+}23] Zhaohong Chen, Xin Long, Long Chen, Yalan Wu, Jigang Wu, and Shuangyin Liu. Intra-cluster aggregation aware routing for distributed training in wireless sensor networks. Concurrency and Computation: Practice and Experience, 35(17): e6795:1–e6795:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cai:2021:SCF

[CLDY21] Zhiming Cai, Daming Li, Lianbing Deng, and Xiang Yao. Smart city framework based on intelligent sensor network and visual surveillance. Concurrency and Computation: Practice and Experience, 33(12):e5301:1–e5301:??, June 25, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cai:2023:R

[CLDY23] Zhiming Cai, Daming Li, Lianbing Deng, and Xiang Yao. Retraction. Concurrency and Computation: Practice and Experience, 33(12):e7751:1–e7751:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chakraborty:2020:ESE

[CLE^{+}20] Sourav Chakraborty, Ignacio Laguna, Murali Emani, Kathryn Mohror, Dhabaleswar K. Panda, Martin Schulz, and Hari Subramoni. ER einit: Scalable and efficient fault-tolerance for bulk-synchronous MPI applications. Concurrency and Computation: Practice and Experience, 32(3):e4863:1–e4863:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cui:2021:RMF

[CLL^{+}21] Yixi Cui, Zixin Liu, Ziran Li, Kun Guo, and Yishan Sun. Research on multiscale features and integrated forecasting of the Belt and Road exchange rate index. Concurrency and Computation: Practice and Experience, 33(6):e6053:1–e6053:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cai:2020:UIB

[CLLB20] Zhicheng Cai, Duan Liu, Yifei Lu, and Rajkumar Buyya. Unequal-interval based loosely coupled control method for auto-scaling heterogeneous cloud resources for web applications. Concurrency and Computation: Practice and Experience, 32(23):e5926:1–e5926:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2021:PRI

[CLT^{+}21] Xin Chen, Zhuo Li, Chao Tang, Shenglong Xiao, and Ying Chen. Person re-identification in the edge computing system: a deep square similarity learning approach. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2021:DST

[CLWX21] Haitao Cheng, Peng Li, Ruchuan Wang, and He Xu. Dynamic spatio-temporal logic based on RCC-8. Concurrency and Computation: Practice and Experience, 33(22):e5900:1– e5900:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2021:RTA

[CLX^{+}21] Huafeng Chen, Hongyang Li, Zengmin Xu, Yunhong Zhao, and Tigang He. Real-time action feature extraction via fast PCA-Flow. Concurrency and Computation: Practice and Experience, 33(11):e5507:1–e5507:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2021:GRB

[CLY^{+}21] Yangwei Cheng, Gongfa Li, Mingchao Yu, Du Jiang, Juntong Yun, Ying Liu, Yibo Liu, and Disi Chen. Gesture recognition based on surface electromyography-feature image. Concurrency and Computation: Practice and Experience, 33(6): e6051:1–e6051:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ci:2022:ATT

[CLYG22] Qingyu Ci, Hourong Li, Shuwei Yang, and Jin Gao. Adaptive and transparent task scheduling of GPU-powered clusters. Concurrency and Computation: Practice and Experience, 34 (9):e5793:1–e5793:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chehri:2020:ITI

[CM20] Abdellah Chehri and Hussein T. Mouftah. Internet of Things — integrated IR-UWB technology for healthcare applications. Concurrency and Computation: Practice and Experience, 32 (2):e5454:1–e5454:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coti:2021:DPO

[CM21] Camille Coti and Allen D. Malony. DiPOSH: a portable Open-SHMEM implementation for short API-to-network path. Concurrency and Computation: Practice and Experience, 33 (11):e6179:1–e6179:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Conceicao:2021:WBG

[CMA^{+}21] Luís Conceição, Diogo Martinho, Rui Andrade, João Carneiro, Constantino Martins, Goreti Marreiros, and Paulo Novais. A web-based group decision support system for multicriteria problems. Concurrency and Computation: Practice and Experience, 33(2):e5298:1–e5298:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2023:THT

[CMJC23] Jing Chen, Kun Ma, Ke Ji, and Zhenxiang Chen. TM-HOL: Topic memory model for detection of hate speech and offensive language. Concurrency and Computation: Practice and Experience, 35(14):e6754:1–e6754:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chetouane:2022:VBV

[CMJM22] Ameni Chetouane, Sabra Mabrouk, Imen Jemili, and Mohamed Mosbah. Vision-based vehicle detection for road traffic congestion classification. Concurrency and Computation: Practice and Experience, 34(7):e5983:1–e5983:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cherif:2022:SCT

[CMK22] Walid Cherif, Abdellah Madani, and Mohamed Kissi. Supervised classification by thresholds: Application to automated text categorization and opinion mining. Concurrency and Computation: Practice and Experience, 34(4):e6613:1– e6613:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Choudhury:2021:CAB

[CMLL21] Khanjan Choudhury, R. Murugan, Mohammad Azharuddin Laskar, and Rabul Hussain Laskar. A comparative analysis between late fusion of features approach and ensemble of multiple classifiers approach for image classification. Concurrency and Computation: Practice and Experience, 33(20):e6371:1– e6371:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carracciuolo:2021:AGP

[CMS21] Luisa Carracciuolo, Valeria Mele, and Lukasz Szustak. About the granularity portability of block-based Krylov methods in heterogeneous computing environments. Concurrency and Computation: Practice and Experience, 33(4):e6008:1– e6008:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Carracciuolo:2024:TNL

[CMS24] Luisa Carracciuolo, Valeria Mele, and Gianluca Sabella. Toward a new linpack-like benchmark for heterogeneous computing resources. Concurrency and Computation: Practice and Experience, 36(10):e7962:1–e7962:??, May 1, 2024. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Calzarossa:2020:ECA

[CMT20] Maria Carla Calzarossa, Luisa Massari, and Daniele Tessera. Evaluation of cloud autoscaling strategies under different incoming workload patterns. Concurrency and Computation: Practice and Experience, 32(17):e5667:1–e5667:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coti:2021:TBA

[CMY21] Camille Coti, David Monniaux, and Hang Yu. A task-based approach to parallel parametric linear programming solving, and application to polyhedral computations. Concurrency and Computation: Practice and Experience, 33(6):e6050:1– e6050:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cai:2020:USS

[CNG^{+}20] Xingjuan Cai, Yun Niu, Shaojin Geng, Jiangjiang Zhang, Zhihua Cui, Jianwei Li, and Jinjun Chen. An under-sampled software defect prediction method based on hybrid multiobjective cuckoo search. Concurrency and Computation: Practice and Experience, 32(5):e5478:1–e5478:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bagavathi:2021:ETC

[CO21] Bagavathi C and Saraniya O. Enhanced texture classification through feature compaction using dihybrid bio-inspired computation techniques. Concurrency and Computation: Practice and Experience, 33(24):e6453:1–e6453:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coban:2022:NMA

[Cob22] Onder Coban. A new modification and application of item response theory-based feature selection for different machine learning tasks. Concurrency and Computation: Practice and Experience, 34(26):e7282:1–e7282:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chiba:2023:NBV

[CON23] Tairi Chiba, Ren Ohmura, and Junya Nakamura. Network bandwidth variation-adapted state transfer for geo-replicated state machines and its application to dynamic replica replacement. Concurrency and Computation: Practice and Experience, 35(19):e7408:1–e7408:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2021:COS

[COZ21] Weijun Chen, Urszula Ogiela, and Jindan Zhang. Customeroriented sales modeling strategy in a big data environment. Concurrency and Computation: Practice and Experience, 33 (19):e6366:1–e6366:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cristin:2022:ITD

[CPA22] Rajan Cristin, Sasi Padma Premnath, and John Patrick Ananth. Image tampering detection in image forensics using earthworm-rider optimization. Concurrency and Computation: Practice and Experience, 34(26):e7293:1–e7293:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chourasia:2023:IDR

[CPCK23] Vishakha Chourasia, Sudhakar Pandey, Rahul Chaurasiya, and Sanjay Kumar. Improving the delivery ratio in vehicular delay-tolerant network using customized binary particle swarm optimization-based optimized throwbox deployment. Concurrency and Computation: Practice and Experience, 35 (1):e7427:1–e7427:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Choi:2020:NDM

[CPH20] Chang Choi, Florin Pop, and Jun Huang. Novel data mining paradigms based on soft computing and machine learning in the current and upcoming information society revolution. Concurrency and Computation: Practice and Experience, 32(18):e5937:1–e5937:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2021:EEC

[CPLX21] Hao Chen, Yubiao Pan, Cheng Li, and Yinlong Xu. ECR: Eviction-cost-aware cache management policy for page-level flash-based SSDs. Concurrency and Computation: Practice and Experience, 33(15):e5395:1–e5395:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coro:2021:NIB

[CPPP21] Gianpaolo Coro, Giancarlo Panichi, Pasquale Pagano, and Erico Perrone. NLPHub: an e-infrastructure-based text mining hub. Concurrency and Computation: Practice and Experience, 33(5):e5986:1–e5986:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2022:WAR

[CPQ^{+}22] Buqing Cao, Mi Peng, Yueying Qing, Jianxun Liu, Guosheng Kang, Bing Li, and Kenneth K. Fletcher. Web API recommendation via combining graph attention representation and deep factorization machines quality prediction. Concurrency and Computation: Practice and Experience, 34(21):e7069:1– e7069:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2021:MSC

[CPYC21] Tinggui Chen, Lijuan Peng, Jianjun Yang, and Guodong Cong. Modeling, simulation, and case analysis of COVID-19 over network public opinion formation with individual internal factors and external information characteristics. Concurrency and Computation: Practice and Experience, 33(17): e6201:1–e6201:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chithanuru:2023:ADB

[CR23] Vasavi Chithanuru and Mangayarkarasi Ramaiah. An anomaly detection on blockchain infrastructure using artificial intelligence techniques: Challenges and future directions — a review. Concurrency and Computation: Practice and Experience, 35(22):e7724:1–e7724:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chaabane:2023:UDA

[CRB23] Mariam Chaabane, Ismael Bouassida Rodriguez, and Hamida Belkhiria. UML design for access control architecture in smart building application. Concurrency and Computation: Practice and Experience, 35(22):e7721:1–e7721:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cook:2021:ESH

[CS21] Steven Cook and Tamar Shinar. Enabling simulation of highdimensional micro-macro biophysical models through hybrid CPU and multi-GPU parallelism. Concurrency and Computation: Practice and Experience, 33(17):e6305:1–e6305:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2022:CHI

[CS22] Lu Cao and Hong Shen. CSS: Handling imbalanced data by improved clustering with stratified sampling. Concurrency and Computation: Practice and Experience, 34(2):e6071:1– e6071:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chakravarthy:2023:NNN

[CS23] Thota Seshu Chakravarthy and Lokesh Selvaraj. NNPEC: Neighborhood node propagation entropy centrality is a unique way to find the influential node in a complex network. Concurrency and Computation: Practice and Experience, 35(12): e7685:1–e7685:??, May 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chanu:2022:EMF

[CSC22] Usham Sanjota Chanu, Khundrakpam Johnson Singh, and Yambem Jina Chanu. An ensemble method for feature selection and an integrated approach for mitigation of distributed denial of service attacks. Concurrency and Computation: Practice and Experience, 34(13):e6919:1–e6919:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Catelan:2023:ECA

[CSD23] Daniela Catelan, Ricardo Santos, and Liana Duenha. Evaluation and characterization of approximate arithmetic circuits. Concurrency and Computation: Practice and Experience, 35 (17):e6865:1–e6865:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cai:2020:BHB

[CSL20] Jing Cai, Yimin Shi, and Ting Lin. E-Bayesian and hierarchical Bayesian estimations for parallel system model in the presence of masked data. Concurrency and Computation: Practice and Experience, 32(9):e5615:1–e5615:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cohen:2021:IMR

[CSL21] Fernand Cohen, Sowrirajan Sowmithran, and Chenxi Li. 3D iris model and reader for iris identification. Concurrency and Computation: Practice and Experience, 33(12):e5653:1– e5653:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chandan:2024:DEC

[CSR24] Madhavarapu Chandan, S. G. Santhi, and T. Srinivasa Rao. Deep ensemble of classifier for intrusion detection in WSN and improved attack mitigation process. Concurrency and Computation: Practice and Experience, 36(8):e7952:1–e7952:??, April 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Canbay:2022:NUA

[CSV22] Yavuz Canbay, Seref Sagiroglu, and Yilmaz Vural. A new utility-aware anonymization model for privacy preserving data publishing. Concurrency and Computation: Practice and Experience, 34(10):e6808:1–e6808:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cui:2020:FSI

[CSW20] Xiaohui Cui, Bilial Suman, and Zhibo Wang. Foreword to the special issue of the Intelligent Systems for the Internet of Things (ISIT2017). Concurrency and Computation: Practice and Experience, 32(13):e5618:1–e5618:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2020:ATL

[CSWC20] Huaming Chen, Jun Shen, Lei Wang, and Chi-Hung Chi. APEX2S: a two-layer machine learning model for discovery of host-pathogen protein-protein interactions on cloud-based multiomics data. Concurrency and Computation: Practice and Experience, 32(23):e5846:1–e5846:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2022:ADC

[CSWZ22] Zhansheng Chen, Hong Shen, Tingmei Wang, and Xiaofan Zhao. An adaptive on-demand charging scheme for rechargeable wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(2):e6136:1–e6136:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cai:2021:POT

[CT21] Yunlu Cai and Chunming Tang. Privacy of outsourced twoparty k-means clustering. Concurrency and Computation: Practice and Experience, 33(8):e5473:1–e5473:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chinnasamy:2022:RBS

[CT22] Rajagopal Chinnasamy and Balasubramanian T. Rank biserial stochastic feature embed bivariate kernelized regressive bootstrap aggregative classifier for school student dropout prediction. Concurrency and Computation: Practice and Experience, 34(21):e7133:1–e7133:??, September 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cassagne:2023:DHT

[CTA^{+}23] Adrien Cassagne, Romain Tajan, Olivier Aumage, Camille Leroux, Denis Barthou, and Christophe Jégo. A DSEL for high throughput and low latency software-defined radio on multicore CPUs. Concurrency and Computation: Practice and Experience, 35(23):e7820:1–e7820:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cekik:2022:NMF

[CU22] Rasim Cekik and Alper Kursat Uysal. A new metric for feature selection on short text datasets. Concurrency and Computation: Practice and Experience, 34(13):e6909:1–e6909:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cudennec:2020:AMP

[Cud20] Loïc Cudennec. Adaptive message passing polling for energy efficiency: Application to software-distributed shared memory over heterogeneous computing resources. Concurrency and Computation: Practice and Experience, 32(24):e5960:1– e5960:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2022:NAN

[CW22] Bin Chen and Jianping Wu. A new approach for node centrality evaluation based on residual network performance measurement. Concurrency and Computation: Practice and Experience, 34(3):e6560:1–e6560:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Calyam:2021:MSF

[CWDM^{+}21] Prasad Calyam, Nancy Wilkins-Diehr, Mark Miller, Emre H. Brookes, Ritu Arora, Amit Chourasia, Douglas M. Jennewein, Viswanath Nandigam, M. Drew LaMar, Sean B. Cleveland, Greg Newman, Shaowen Wang, Ilya Zaslavsky, Michael A. Cianfrocco, Kevin Ellett, David Tarboton, Keith G. Jeffery, Zhiming Zhao, Juan González-Aranda, Mark J. Perri, Greg Tucker, Leonardo Candela, Tamas Kiss, and Sandra Gesing.

Measuring success for a future vision: Defining impact in science gateways/virtual research environments. Concurrency and Computation: Practice and Experience, 33(19):e6099:1– e6099:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chung:2023:RPL

[CWFK23] Minh Thanh Chung, Josef Weidendorfer, Karl Fürlinger, and Dieter Kranzlmüller. From reactive to proactive load balancing for task-based parallel applications in distributed memory machines. Concurrency and Computation: Practice and Experience, 35(24):e7828:1–e7828:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cui:2024:TSP

[CWH^{+}24] Huanyu Cui, Nianbin Wang, Qilong Han, Ye Wang, and Jiahang Li. A two-stage parallel method on GPU based on hybrid-compression-format for diagonal matrix. Concurrency and Computation: Practice and Experience, 36(1):e7887:1– e7887:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2020:UIQ

[CWL^{+}20] Peng Cheng, Yong Wang, Yutong Lu, Yunfei Du, and Zhiguang Chen. UniIndex: an index and query middleware for parallel file systems. Concurrency and Computation: Practice and Experience, 32(9):e5609:1–e5609:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2021:RMF

[CWL^{+}21] Yuantao Chen, Jin Wang, Songjie Liu, Xi Chen, Jie Xiong, Jingbo Xie, and Kai Yang. Retracted: Multiscale fast correlation filtering tracking algorithm based on a feature fusion model. Concurrency and Computation: Practice and Experience, 33(15):e5533:1–e5533:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2022:RMF

[CWL^{+}22] Yuantao Chen, Jin Wang, Songjie Liu, Xi Chen, Jie Xiong, Jingbo Xie, and Kai Yang. Retracted: Multiscale fast correlation filtering tracking algorithm based on a feature fusion model. Concurrency and Computation: Practice and Experience, 34(1):e6589:1–e6589:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cai:2020:RII

[CWLL20] Yuanzheng Cai, Tao Wang, Wei Liu, and Zhiming Luo. A robust interclass and intraclass loss function for deep learning based tongue segmentation. Concurrency and Computation: Practice and Experience, 32(22):e5849:1–e5849:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2023:EEM

[CWW23a] Hong Chen, Yiping Wen, and Yuan Wang. An energy-efficient method of resource allocation based on request prediction in multiple cloud data centers. Concurrency and Computation: Practice and Experience, 35(9):e7636:1–e7636:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2023:DTS

[CWW^{+}23b] Lili Chen, Zhen Wang, Jintao Wu, Yunchuan Guo, Fenghua Li, and Zifu Li. Dynamic threshold strategy optimization for security protection in Internet of Things: an adversarial deep learning-based game-theoretical approach. Concurrency and Computation: Practice and Experience, 35(20):e6944:1– e6944:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2022:HSS

[CXX^{+}22] Peng Chen, Qi Xiao, Jian Xu, Xiaoli Dong, Linjun Sun, Weijun Li, Xin Ning, Guojun Wang, and Ziheng Chen. Harnessing semantic segmentation masks for accurate facial attribute editing. Concurrency and Computation: Practice and Experience, 34(12):e5798:1–e5798:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2024:RTG

[CXXW24] Yan Chen, Jiajian Xie, Jiajun Xue, and Hui Wang. A robust transformer GAN for unpaired data makeup transfer. Concurrency and Computation: Practice and Experience, 36(9): e7994:1–e7994:??, April 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chaudhry:2022:CCR

[CY22] Natalia Chaudhry and Muhammad Murtaza Yousaf. Concurrency control for real-time and mobile transactions: Historical view, challenges, and evolution of practices. Concurrency and Computation: Practice and Experience, 34(3):e6549:1– e6549:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chaudhry:2024:EHC

[CY24] Natalia Chaudhry and Muhammad Murtaza Yousaf. Exploiting hashing for concurrent query processing and indexing of current location of moving objects. Concurrency and Computation: Practice and Experience, 36(11):e8013:1–e8013:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2021:OCA

[CYC21] Hsing-Chung Chen, Wei-Ju Yang, and Chao-Lung Chou. An online cognitive authentication and trust evaluation application programming interface for cognitive security gateway based on distributed massive Internet of Things network. Concurrency and Computation: Practice and Experience, 33 (19):e6128:1–e6128:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cui:2020:SAR

[CYDW20] Hui Cui, Tsz Hon Yuen, Robert H. Deng, and Guilin Wang. Server-aided revocable attribute-based encryption for cloud computing services. Concurrency and Computation: Practice and Experience, 32(14):e5680:1–e5680:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Choi:2021:CUP

[CYK^{+}21] Jake Choi, Hojun You, Chongam Kim, Heon Young Yeom, and Yoonhee Kim. Comparing unified, pinned, and host/ device memory allocations for memory-intensive workloads on tegra SoC. Concurrency and Computation: Practice and Experience, 33(4):e6018:1–e6018:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2020:NRA

[CYQ^{+}20] Jiaying Chen, Jiong Yu, Yurong Qian, Ping Li, and Chen Bian. A novel recommender algorithm based on graph embedding and diffusion sampling. Concurrency and Computation: Practice and Experience, 32(17):e5664:1–e5664:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chi:2022:ALS

[CYW^{+}22] Xiaoxiao Chi, Chao Yan, Hao Wang, Wajid Rafique, and Lianyong Qi. Amplified locality-sensitive hashing-based recommender systems with privacy protection. Concurrency and Computation: Practice and Experience, 34(14):e5681:1– e5681:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2021:IPM

[CYZ^{+}21] Xianfu Cheng, Minhua You, Jian Zhou, Tian Yuan, Zhihu Guo, and Xiaotian Ma. An integrated product modularity method based on transfer network of failure mode-recycling decision for remanufacturing. Concurrency and Computation: Practice and Experience, 33(9):e6158:1–e6158:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2023:RPC

[CYZ^{+}23] Lin Chen, Xianzhong Yi, Yuanhua Zhou, Lijun Liu, Hangming Liu, and Saeid Razmjooy. Reactive power control of hybrid systems using improved coyote optimizer. Concurrency and Computation: Practice and Experience, 35(28):e7859:1– e7859:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2023:DBT

[CYZX23] Qiuling Chen, Ayong Ye, Ziwen Zhao, and Jinbo Xiong. The dummy-based trajectory privacy protection method to resist correlation attacks in Internet of Vehicles. Concurrency and Computation: Practice and Experience, 35(19):e7154:1– e7154:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cui:2021:HMO

[CZ21] Zhihua Cui and Jiangjiang Zhang. A hybrid many-objective optimization algorithm for coal green production problem. Concurrency and Computation: Practice and Experience, 33 (6):e6040:1–e6040:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2022:NIC

[CZ22] Jingqiang Chen and Hai Zhuge. A news image captioning approach based on multimodal pointer-generator network. Concurrency and Computation: Practice and Experience, 34(7): e5721:1–e5721:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2024:HGW

[CZ24] Shuilin Chen and Jianguo Zheng. Hybrid grey wolf optimizer for solving permutation flow shop scheduling problem. Concurrency and Computation: Practice and Experience, 36 (5):e7942:1–e7942:??, February 29, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Czarnul:2023:MCO

[Cza23] Paweł Czarnul. A multithreaded CUDA and OpenMP based power-aware programming framework for multi-node GPU systems. Concurrency and Computation: Practice and Experience, 35(25):e7897:1–e7897:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2023:MML

[CZCM23] Peng Chen, Jixin Zhang, Jiageng Chen, and Weizhi Meng. MLCT: a multi-level contact tracing scheme with strong privacy. Concurrency and Computation: Practice and Experience, 35(19):e6929:1–e6929:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2020:VSC

[CZG^{+}20] Meng Cao, Yonghua Zhu, Wenjing Gao, Mengyao Li, and Shaoxiu Wang. Various syncretic co-attention network for multimodal sentiment analysis. Concurrency and Computation: Practice and Experience, 32(24):e5954:1–e5954:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cui:2022:MDB

[CZL^{+}22] Zhihua Cui, Yaru Zhao, Hao Liu, Ling Wang, and Wenhui Fan. Malware detection based on multi-objective convolution restricted Boltzmann machine model and constraint-dividing crossover strategy algorithm. Concurrency and Computation: Practice and Experience, 34(19):e7030:1–e7030:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cheng:2022:CSP

[CZTC22] Xiang Cheng, Jiale Zhang, Yaofeng Tu, and Bing Chen. Cyber situation perception for Internet of Things systems based on zero-day attack activities recognition within advanced persistent threat. Concurrency and Computation: Practice and Experience, 34(16):e6001:1–e6001:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Can:2023:BDP

[CZUPA23] Ali Burak Can, Mounes Zaval, Meryem Uzun-Per, and Mehmet S. Aktas. On the big data processing algorithms for finding frequent sequences. Concurrency and Computation: Practice and Experience, 35(24):e7660:1–e7660:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chen:2022:CAA

[CZZ^{+}22] Genlang Chen, Jiajian Zhang, Zufang Zhu, Hao Wang, Hai Jiang, and Chaoyi Pang. CRAC: an automatic assistant compiler of checkpoint/restart for OpenCL program. Concurrency and Computation: Practice and Experience, 34(8):e6048:1– e6048:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dhanalakshmi:2022:SRB

[DA22] J. Dhanalakshmi and N. Ayyanathan. A systematic review of big data in energy analytics using energy computing techniques. Concurrency and Computation: Practice and Experience, 34(4):e6647:1–e6647:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Drummond:2024:DAW

[DAdBM^{+}24] Lúcia Maria A. Drummond, Luciano Andrade, Pedro de Brito Muniz, Matheus Marotti Pereira, Thiago do Prado Silva, and Luan Teylo. Design and analyses of web scraping on burstable virtual machines. Concurrency and Computation: Practice and Experience, 36(9):e7999:1–e7999:??, April 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dahiya:2023:HCS

[Dah23] Deepak Dahiya. Hybrid classifier strategy with tuned training weights for distributed denial of service attack detection. Concurrency and Computation: Practice and Experience, 35 (4):e7547:1–e7547:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dogra:2022:MMP

[DAK22] Ayush Dogra, Chirag Kamal Ahuja, and Sanjeev Kumar. A multi-modality paradigm for CT and MRI fusion with applications of quantum image processing. Concurrency and Computation: Practice and Experience, 34(20):e6610:1–e6610:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deebak:2023:LPA

[DAT23] Deebak Bakkiam Deebak and Fadi Al-Turjman. Lightweight privacy-aware secure authentication scheme for cyberphysical systems in the edge intelligence era. Concurrency and Computation: Practice and Experience, 35(13):e6510:1– e6510:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ding:2022:IRI

[DAW22] Nan Ding, Muaaz Awan, and Samuel Williams. Instruction roofline: an insightful visual performance model for GPUs. Concurrency and Computation: Practice and Experience, 34(20):e6591:1–e6591:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Diehl:2023:ICC

[DB23] Patrick Diehl and Steven R. Brandt. Interactive C++ code development using C++Explorer and GitHub classroom for educational purposes. Concurrency and Computation: Practice and Experience, 35(18):e6893:1–e6893:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Doostani:2024:LHM

[DBB24] Shadi Doostani, Hamid Barati, and Ali Barati. A lightweight hierarchical method for improving security in the internet of things using fuzzy logic. Concurrency and Computation: Practice and Experience, 36(6):e7959:1–e7959:??, March 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dangi:2022:SAS

[DBD22] Dharmendra Dangi, Amit Bhagat, and Dheeraj Kumar Dixit. Sentiment analysis of social media data based on chaotic coyote optimization algorithm based time weight-AdaBoost support vector machine approach. Concurrency and Computation: Practice and Experience, 34(3):e6581:1–e6581:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dixit:2023:AFN

[DBD23] Dheeraj Kumar Dixit, Amit Bhagat, and Dharmendra Dangi. An accurate fake news detection approach based on a Levy flight honey badger optimized convolutional neural network model. Concurrency and Computation: Practice and Experience, 35(1):e7382:1–e7382:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devarapalli:2021:NAI

[DBK21] Ramesh Devarapalli, Biplab Bhattacharyya, and Archana Kumari. A novel approach of intensified barnacles mating optimization for the mitigation of power system oscillations. Concurrency and Computation: Practice and Experience, 33(17):e6303:1–e6303:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Das:2022:BTS

[DBN^{+}22] Suchismita Das, Srijib Bose, Gopal K. Nayak, Suresh Chandra Satapathy, and Sanjay Saxena. Brain tumor segmentation and overall survival period prediction in glioblastoma multiforme using radiomic features. Concurrency and Computation: Practice and Experience, 34(20):e6501:1–e6501:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Delwadia:2022:PAS

[DBPC22] Kirtan Delwadia, Dhruvil Bhatt, Shishir Purohit, and Bhaskar Chaudhury. Parallel algorithm for synthetic image generation with application to tokamak plasma diagnostics. Concurrency and Computation: Practice and Experience, 34(24):e7217:1–e7217:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DaherDaibes:2022:EUG

[DBS^{+}22] João Victor Daher Daibes, Milton Brown Do Coutto Filho, Julio Cesar Stacchini de Souza, Esteban Walter Gonzalez Clua, and Rainer Zanghi. Experience of using graphical processing unit in power flow computation. Concurrency and Computation: Practice and Experience, 34(6):e6762:1– e6762:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Junior:2021:CCC

[dCJAAdOD21] Francisco Heron de Carvalho Junior, Wagner Guimarães Al-Alam, and Allberson B. de Oliveira Dantas. Contextual contracts for component-oriented resource abstraction in a cloud of high performance computing services. Concurrency and Computation: Practice and Experience, 33(18):e6225:1– e6225:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Junior:2020:TUP

[dCJBP20] Marcelo Antonio de Carvalho Junior and Paulo Bandiera-Paiva. Towards unobtrusive patient-centric access-control for health information system. Concurrency and Computation: Practice and Experience, 32(22):e5845:1–e5845:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dey:2021:SDT

[DCK21] Prasanjit Dey, Swades Kumar Chaulya, and Sanjay Kumar. Secure decision tree twin support vector machine training and classification process for encrypted IoT data via blockchain platform. Concurrency and Computation: Practice and Experience, 33(16):e6264:1–e6264:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DAmore:2021:SKF

[DCM21] Luisa D’Amore, Rosalba Cacciapuoti, and Valeria Mele. A scalable Kalman filter algorithm: Trustworthy analysis on constrained least square model. Concurrency and Computation: Practice and Experience, 33(4):e6022:1–e6022:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

deCarvalho:2023:AAF

[dCMA23] Leonardo Reboucas de Carvalho, Alba Cristina Alves Melo, and Aleteia Araujo. AFMC: an alignment framework for multiple computing services and providers. Concurrency and Computation: Practice and Experience, 35(18):e7595:1– e7595:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

deCamargo:2021:CAH

[dCMM21] Raphael Y. de Camargo, Fabrizio Marozzo, and Wellington Martins. Computer architecture and high performance computing. Concurrency and Computation: Practice and Experience, 33(18):e6526:1–e6526:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devi:2023:WTW

[DCR23] E. Ahila Devi, A. Chandrasekar, and S. Radhika. WSO-T2FSM: War strategy optimization-based type-2 fuzzy-based starling murmuration for addressing the routing problem in mobile ad hoc network. Concurrency and Computation: Practice and Experience, 35(21):e7684:1–e7684:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dikaiakos:2023:CPM

[DCT^{+}23] Marios D. Dikaiakos, Nikolas G. Chatzigeorgiou, Athanasios Tryfonos, Andreas Andreou, Nicholas Loulloudes, George Pallis, and George E. Georgiou. A cyber-physical management system for medium-scale solar-powered data centers. Concurrency and Computation: Practice and Experience, 35(10): e7658:1–e7658:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2020:STM

[DCWM20] Lei Du, Zhihua Cui, Lifang Wang, and Junming Ma. Structure tuning method on deep convolutional generative adversarial network with nondominated sorting genetic algorithm II. Concurrency and Computation: Practice and Experience, 32(14):e5688:1–e5688:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2022:MAB

[DCZ^{+}22] Tingting Dong, Lei Chen, Li Zhou, Fei Xue, and Huilin Qin. A multi-agent based mechanism for collaboratively detecting distributed denial of service attacks in internet of vehicles. Concurrency and Computation: Practice and Experience, 34 (13):e6904:1–e6904:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devi:2021:MNM

[DD21] Saravanan Durga Devi and Dhamotharan Rukmani Devi. Malicious node and malicious observer node detection system in MANETs. Concurrency and Computation: Practice and Experience, 33(3):e5241:1–e5241:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dulger:2024:NPT

[DD24] Özcan Dülger and Tansel Dökeroglu. A new parallel tabu search algorithm for the optimization of the maximum vertex weight clique problem. Concurrency and Computation: Practice and Experience, 36(2):e7891:1–e7891:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dautov:2021:DAT

[DDB^{+}21] Rustem Dautov, Salvatore Distefano, Dario Bruneo, Francesco Longo, Giovanni Merlino, and Antonio Puliafito. Data agility through clustered edge computing and stream processing. Concurrency and Computation: Practice and Experience, 33 (7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Darabian:2020:OBT

[DDH^{+}20] Hamid Darabian, Ali Dehghantanha, Sattar Hashemi, Sajad Homayoun, and Kim-Kwang Raymond Choo. An opcodebased technique for polymorphic Internet of Things malware detection. Concurrency and Computation: Practice and Experience, 32(6):e5173:1–e5173:??, March 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deshpande:2023:ISR

[DDUK23] Prasanna Deshpande, Sidharth Dinesh, Apurva Umredkar, and Ashwin G. Kothari. Improved search and rescue optimization based load balancing scheme of CCTV footage on edge computing environment. Concurrency and Computation: Practice and Experience, 35(5):e7556:1–e7556:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dai:2020:JMC

[DDZ^{+}20] Liang Dai, Guodong Du, Jia Zhang, Candong Li, Rong Wei, and Shaozi Li. Joint multilabel classification and feature selection based on deep canonical correlation analysis. Concurrency and Computation: Practice and Experience, 32(22): e5864:1–e5864:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dufrechou:2020:UAI

[DE20] Ernesto Dufrechou and Pablo Ezzatti. Using analysis information in the synchronization-free GPU solution of sparse triangular systems. Concurrency and Computation: Practice and Experience, 32(10):e5499:1–e5499:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dogan:2022:PHH

[DE22] Atakan Dogan and Kemal Ebcioglu. A parallel hardware hypervisor for hardware-accelerated cloud computing. Concurrency and Computation: Practice and Experience, 34(9): e6554:1–e6554:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deng:2024:AOD

[Den24] Bin Deng. Aquila optimizer with dynamic group strategy for global optimization tasks. Concurrency and Computation:

Practice and Experience, 36(7):e7971:1–e7971:??, March 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devi:2021:PDV

[Dev21] N Yamuna Devi. A parallel direct-vertical map reduce programming model for an effective frequent pattern mining in a dispersed environment. Concurrency and Computation: Practice and Experience, 33(24):e6470:1–e6470:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dhawan:2021:HQS

[DG21] Sachin Dhawan and Rashmi Gupta. High-quality steganography scheme using hybrid edge detector and Vernam algorithm based on hybrid fuzzy neural network. Concurrency and Computation: Practice and Experience, 33(24):e6448:1–e6448:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dorozhinskii:2024:FGT

[DGB24] Ravil Dorozhinskii, Gonzalo Brito Gadeschi, and Michael Bader. Fused GEMMs towards an efficient GPU implementation of the ADER-DG method in SeisSol. Concurrency and Computation: Practice and Experience, 36(12):e8037:1– e8037:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DeSalve:2021:EHC

[DGM21] Andrea De Salve, Barbara Guidi, and Andrea Michienzi. Exploiting homophily to characterize communities in online social networks. Concurrency and Computation: Practice and Experience, 33(8):e5929:1–e5929:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Durand:2020:DLI

[DGP20] Arnaud Durand, Pascal Gremaud, and Jacques Pasquier. Decentralized LPWAN infrastructure using blockchain and digital signatures. Concurrency and Computation: Practice and Experience, 32(12):e5352:1–e5352:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dosanjh:2020:TQM

[DGSB20] Matthew G. F. Dosanjh, Ryan E. Grant, Whit Schonbein, and Patrick G. Bridges. Tail queues: a multi-threaded matching architecture. Concurrency and Computation: Practice and Experience, 32(3):e5158:1–e5158:??, February 10, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dugre:2023:PCD

[DHSG23] Mathieu Dugré, Valérie Hayot-Sasson, and Tristan Glatard. Performance comparison of Dask and Apache Spark on HPC systems for neuroimaging. Concurrency and Computation: Practice and Experience, 35(21):e7635:1–e7635:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Drira:2020:NTA

[DJ20] Khalil Drira and Mohamed Jmaiel. New trends in adaptive smart systems, services, and architectures. Concurrency and Computation: Practice and Experience, 32(15):e5589:1– e5589:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Debbabi:2021:NSF

[DJF21] Fadoua Debbabi, Rihab Jmal, and Lamia Chaari Fourati. 5G network slicing: Fundamental concepts, architectures, algorithmics, projects practices, and open issues. Concurrency and Computation: Practice and Experience, 33(20):e6352:1– e6352:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dutta:2021:QAB

[DJGF21] Avik Dutta, Chandrashekar Jatoth, G. R. Gangadharan, and Ugo Fiore. QoS-aware big service composition using distributed co-evolutionary algorithm. Concurrency and Computation: Practice and Experience, 33(19):e6362:1–e6362:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Denis:2022:MAC

[DJJR22] Alexandre Denis, Julien Jaeger, Emmanuel Jeannot, and Florian Reynier. A methodology for assessing computation/communication overlap of MPI nonblocking collectives. Concurrency and Computation: Practice and Experience, 34 (22):e7168:1–e7168:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Denis:2024:TTB

[DJST24] Alexandre Denis, Emmanuel Jeannot, Philippe Swartvagher, and Samuel Thibault. Tracing task-based runtime systems: Feedbacks from the StarPU case. Concurrency and Computation: Practice and Experience, 36(3):e7920:1–e7920:??, February 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devi:2021:HCW

[DK21] V. Anusuya Devi and V. Kalaivani. Hybrid cryptosystem in wireless body area networks using message authentication code and modified and enhanced lattice-based cryptography (MAC-MELBC) in healthcare applications. Concurrency and Computation: Practice and Experience, 33(9):e6132:1– e6132:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dolek:2022:DLM

[DK22a] Ishak Dölek and Atakan Kurt. A deep learning model for Ottoman OCR. Concurrency and Computation: Practice and Experience, 34(20):e6937:1–e6937:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Doss:2022:HOB

[DK22b] Kingsleen Solomon Doss and Somasundaram Kamalakkannan. Hybrid optimization-based privacy preservation of database publishing in cloud environment. Concurrency and Computation: Practice and Experience, 34(11):e6844:1– e6844:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DesLauriers:2021:CAG

[DKA^{+}21] James DesLauriers, Tamas Kiss, Resmi C. Ariyattu, Hai-Van Dang, Amjad Ullah, James Bowden, Dagmar Krefting, Gabriele Pierantoni, and Gabor Terstyanszky. Cloud apps to-go: Cloud portability with TOSCA and MiCADO. Concurrency and Computation: Practice and Experience, 33 (19):e6093:1–e6093:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Diener:2020:HCO

[DKB20] Matthias Diener, Laxmikant V. Kale, and Daniel J. Bodony. Heterogeneous computing with OpenMP and Hydra. Concurrency and Computation: Practice and Experience, 32(20): e5728:1–e5728:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DiLuccio:2021:VSD

[DKC^{+}21] Diana Di Luccio, Sokol Kosta, Aniello Castiglione, Antonio Maratea, and Raffaele Montella. Vessel to shore data movement through the Internet of floating things: a microservice platform at the edge. Concurrency and Computation: Practice and Experience, 33(4):e5988:1–e5988:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Duracik:2020:SSC

[DKH20] Michal Duracík, Emil Krsák, and Patrik Hrkút. Searching source code fragments using incremental clustering. Concurrency and Computation: Practice and Experience, 32(13): e5416:1–e5416:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Diwan:2024:BAE

[DKK24] Princy Diwan, Ramgopal Kashyap, and Brijesh Khandelwal. Blockchain assisted encryption scheme for intellectual share estimation using medical research data. Concurrency and Computation: Practice and Experience, 36(2):e7896:1– e7896:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devi:2021:OIT

[DKL21] M. Ramya Devi, Sivakumar Krishnan, and S. Lokesh. An optimal Internet of Things-based smart cities using vehicular cloud for smart driving. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Daubner:2020:DQL

[DKP^{+}20] Jakub Daubner, Martin Klimo, Jozef Papán, Juraj Smiesko, and Ondrej Such. On the distribution of queue length in ideal links. Concurrency and Computation: Practice and Experience, 32(13):e5203:1–e5203:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dangi:2023:NHD

[DL23] Ramraj Dangi and Praveen Lalwani. A novel hybrid deep learning approach for 5G network traffic control and forecasting. Concurrency and Computation: Practice and Experience, 35(7):e7596:1–e7596:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Duan:2021:CEP

[DLC^{+}21] Lijuan Duan, Zhaoyang Lian, Juncheng Chen, Yuanhua Qiao, Jun Miao, and Mingai Li. Classification of epilepsy period based on combination feature extraction methods and spiking swarm intelligent optimization algorithm. Concurrency and Computation: Practice and Experience, 34(1):e5550:1– e5550:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2024:AFL

[DLG24] Ruizhong Du, Chuan Liu, and Yan Gao. Anonymous federated learning framework in the internet of things. Concurrency and Computation: Practice and Experience, 36(2): e7901:1–e7901:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2023:DPP

[DLY^{+}23] Danlei Du, Entao Luo, Yang Yi, Tao Peng, Xubin Li, Shaobo Zhang, Xu Jiang, and Tian Wang. Differential privacy protection method for trip-oriented shared data. Concurrency and Computation: Practice and Experience, 35(19):e7414:1– e7414:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dutta:2021:DPA

[DM21] Prarthana Dutta and Naresh Babu Muppalaneni. DigiNet: Prediction of Assamese handwritten digits using convolutional neural network. Concurrency and Computation: Practice and Experience, 33(24):e6451:1–e6451:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Suvitha:2022:TPU

[DM22a] Suvitha D and Vijayalakshmi M. Traffic prediction using MSSBiLS with self-attention model. Concurrency and Computation: Practice and Experience, 34(15):e6952:1–e6952:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Das:2022:REB

[DM22b] Ratnakar Das and Jibitesh Mishra. Rider exploitation based whale optimization algorithm for rainfall prediction from meteorological data. Concurrency and Computation: Practice and Experience, 34(19):e7026:1–e7026:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Demir:2022:AAG

[DM22c] Alparslan Serhat Demir and Ebru Mumcu. Analysis of automated guided vehicle use in health care by simulation: a case study in a university hospital. Concurrency and Computation: Practice and Experience, 34(18):e7019:1–e7019:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dwaram:2022:CYF

[DM22d] Jayanarayana Reddy Dwaram and Rudra Kumar Madapuri. Crop yield forecasting by long short-term memory network with Adam optimizer and Huber loss function in Andhra Pradesh, India. Concurrency and Computation: Practice and Experience, 34(27):e7310:1–e7310:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Das:2023:PBN

[DM23] Bappaditya Das and Chintan Kr Mandal. Prediction of behavior of a noncommunicative robot based on observations of its movement. Concurrency and Computation: Practice and Experience, 35(22):e7725:1–e7725:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Debauche:2020:CAP

[DMD^{+}20] Olivier Debauche, Sidi Ahmed Mahmoudi, Nicolas De Cock, Saïd Mahmoudi, Pierre Manneback, and Frédéric Lebeau. Cloud architecture for plant phenotyping research. Concurrency and Computation: Practice and Experience, 32(17): e5661:1–e5661:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2024:IBS

[DMPR24] Yishu Du, Loris Marchal, Guillaume Pallez, and Yves Robert. Improving batch schedulers with node stealing for failed jobs. Concurrency and Computation: Practice and Experience, 36 (12):e8043:1–e8043:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DAmore:2021:TMS

[DMS^{+}21] Luisa D’Amore, Aniello Murano, Loredana Sorrentino, Rossella Arcucci, and Giuliano Laccetti. Toward a multilevel scalable parallel Zielonka’s algorithm for solving parity games. Concurrency and Computation: Practice and Experience, 33(4):e6043:1–e6043:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dalcali:2022:PEE

[DÖD22] Adem Dalcali, Harun Özbay, and Serhat Duman. Prediction of electricity energy consumption including COVID-19 precautions using the hybrid MLR-FFANN optimized with the stochastic fractal search with fitness distance balance algorithm. Concurrency and Computation: Practice and Experience, 34(15):e6947:1–e6947:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

deOliveira:2020:MLT

[dOdMC^{+}20] Lilian M. de Oliveira, Fortunato S. de Menezes, Marcelo A. Cirillo, André V. Saúde, Flávio M. Borém, and Gilberto R. Liska. Machine learning techniques in muliclass problems with application in sensorial analysis. Concurrency and Computation: Practice and Experience, 32(7):e5579:1–e5579:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

deOliveira:2021:TOE

[dOPBdO21] Douglas de Oliveira, Fábio Porto, Cristina Boeres, and Daniel de Oliveira. Towards optimizing the execution of spark scientific workflows using machine learning-based parameter tuning. Concurrency and Computation: Practice and Experience, 33(5):e5972:1–e5972:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dennis:2022:DBN

[DP22a] J. Britto Dennis and M. Shanmuga Priya. Deep belief network and support vector machine fusion for distributed denial of service and economical denial of service attack detection in cloud. Concurrency and Computation: Practice and Experience, 34(1):e6543:1–e6543:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dhavamani:2022:EEP

[DP22b] Logeshwari Dhavamani and P. Prem Priya. Energy-efficient and privacy-preserving approach for Internet of Things nodes using a novel hybrid fuzzy water cycle and evaporation strategy and matrix-based Rivest–Shamir–Adleman encryption algorithm. Concurrency and Computation: Practice and Experience, 34(27):e7336:1–e7336:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dadure:2022:FEA

[DPB22a] Pankaj Dadure, Partha Pakray, and Sivaji Bandyopadhyay. A formula embedding approach for semantic similarity and relatedness between formulas. Concurrency and Computation: Practice and Experience, 34(22):e7146:1–e7146:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deshmukh:2022:MMB

[DPB22b] Varun Deshmukh, Sunil Pathak, and Santosh Bothe. MobEdge: Mobile blockchain-based privacy-edge scheme for healthcare Internet of Things-based ecosystems. Concurrency and Computation: Practice and Experience, 34(23):e7210:1– e7210:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dhar:2024:DGA

[DPBP24] Soumi Dhar, Shyamosree Pal, Gaurav Bhattacharjee, and Abhishek Pandey. A digital geometric approach for discrete Voronoi diagram construction using GPU. Concurrency and Computation: Practice and Experience, 36(11):e8018:1– e8018:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

M:2022:DLA

[DPD^{+}22] Mingcheng Xu M. D, Gaojian Xu Ph.D, Haoyu Xu M. D, Jiadong Zhou M. D, and Shaowen Li Ph.D. A decentralized lightweight authentication protocol under blockchain. Concurrency and Computation: Practice and Experience, 34(13): e6920:1–e6920:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Do:2023:PAE

[DPdS^{+}23] Tu Mai Anh Do, Loïc Pottier, Rafael Ferreira da Silva, Silvina Caíno-Lores, Michela Taufer, and Ewa Deelman. Performance assessment of ensembles of in situ workflows under resource constraints. Concurrency and Computation: Practice and Experience, 35(20):e7111:1–e7111:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DaCosta:2020:FMC

[DPFC20] Georges Da Costa, Jean-Marc Pierson, and Leandro Fontoura-Cupertino. Fast maximum coverage of system behavior from a performance and power point of view. Concurrency and Computation: Practice and Experience, 32(14):e5673:1–e5673:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Diamantini:2021:ACL

[DPS21] Claudia Diamantini, Domenico Potena, and Emanuele Storti. Analytics for citizens: a linked open data model for statistical data exploration. Concurrency and Computation: Practice and Experience, 33(8):e4186:1–e4186:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dwivedi:2022:LGL

[DPSJ22] Bhanu Dwivedi, Bachu Dushmanta Kumar Patro, Vivek Srivastava, and Shimpi Singh Jadon. LBR-GWO: Layered based routing approach using grey wolf optimization algorithm in wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(4):e6603:1–e6603:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deng:2022:FOH

[DPYS22] Weiping Deng, Yan Peng, Fan Yang, and Jun Song. Feature optimization and hybrid classification for malicious web page detection. Concurrency and Computation: Practice and Experience, 34(16):e5859:1–e5859:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2023:MSB

[DQF^{+}23] Zhiqiang Du, Zhi Qu, Yanfang Fu, Muhong Huang, and Liangxin Liu. Multi-strategy-based leader election mechanism for the Raft algorithm. Concurrency and Computation: Practice and Experience, 35(22):e7734:1–e7734:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qi:2022:ERP

[dQtZWS22] Rui dong Qi, Jian tao Zhou, Zhuowei Wang, and Xiaoyu Song. An elastic recommender process for cloud service recommendation scalability. Concurrency and Computation: Practice and Experience, 34(21):e7066:1–e7066:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dietze:2020:SBS

[DR20] Robert Dietze and Gudula Rünger. The search-based scheduling algorithm HP* for parallel tasks on heterogeneous platforms. Concurrency and Computation: Practice and Experience, 32(21):e5898:1–e5898:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

DV:2021:FTR

[DR21] Divya V and Leena Sri R. Fault tolerant resource allocation in fog environment using game theory-based reinforcement learning. Concurrency and Computation: Practice and Experience, 33(16):e6268:1–e6268:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Immanuel:2022:SFV

[DR22a] Jebakumar Immanuel D and Dhanapal R. A secure finger vein verification and authentication scheme for banking network. Concurrency and Computation: Practice and Experience, 34(26):e7281:1–e7281:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Daneshvar:2022:SHR

[DR22b] Hirad Daneshvar and Reza Ravanmehr. A social hybrid recommendation system using LSTM and CNN. Concurrency and Computation: Practice and Experience, 34(18):e7015:1– e7015:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Daniel:2022:NML

[DR22c] V. Antony Asir Daniel and Ravi Ramaraj. A novel modified long short term memory architecture for automatic liver disease prediction from patient records. Concurrency and Computation: Practice and Experience, 34(28):e7372:1–e7372:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dhandapani:2023:PRD

[DR23] Sudhagar Dhandapani and Arokia Renjit Jerald Rodriguez. Poor and rich dolphin optimization algorithm with modified deep fuzzy clustering for COVID-19 patient analysis. Concurrency and Computation: Practice and Experience, 35 (2):e7456:1–e7456:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

doRosario:2021:SSO

[dRdSC^{+}21] Vanderson Martins do Rosario, Anderson Faustino da Silva, Thais Aparecida Silva Camacho, Otávio O. Napoli, Mauricio Breternitz, and Edson Borin. Smart selection of optimizations in dynamic compilers. Concurrency and Computation: Practice and Experience, 33(18):e6089:1–e6089:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

doRosario:2023:FSC

[dRdSZ^{+}23] Vanderson Martins do Rosario, Anderson Faustino da Silva, André Felipe Zanella, Otávio O. Napoli, and Edson Borin. Fast selection of compiler optimizations using performance prediction with graph neural networks. Concurrency and Computation: Practice and Experience, 35(17):e6869:1– e6869:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Das:2022:FHD

[DRM22a] Sujit Kumar Das, Pinki Roy, and Arnab Kumar Mishra. Fusion of handcrafted and deep convolutional neural network features for effective identification of diabetic foot ulcer. Concurrency and Computation: Practice and Experience, 34(5): e6690:1–e6690:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Das:2022:OST

[DRM22b] Sujit Kumar Das, Pinki Roy, and Arnab Kumar Mishra. Oversample-select-tune: a machine learning pipeline for improving diabetes identification. Concurrency and Computation: Practice and Experience, 34(5):e6741:1–e6741:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Duraisamy:2022:CLR

[DRMA22] Abitha Kumari Duraisamy, Raja Guru Ramaraj, Mathankumar Manoharan, and Manjunathan Alagarsamy. Certificateless linkable ring signature-based blockchains for securing cognitive radio networks. Concurrency and Computation: Practice and Experience, 34(24):e7235:1–e7235:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devprasad:2022:CAE

[DRR22] Kayathri Devi Devprasad, Sukumar Ramanujam, and Suresh Babu Rajendran. Context adaptive ensemble classification mechanism with multi-criteria decision making for network intrusion detection. Concurrency and Computation: Practice and Experience, 34(21):e7110:1–e7110:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Daniel:2022:FOB

[DRV22] Jesline Daniel, J. T. Anita Rose, and F. Sangeetha Francelin Vinnarasi. Fuzzy optimization based detection of attacker nodes in wireless networks using deep neural network. Concurrency and Computation: Practice and Experience, 34(25): e7248:1–e7248:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dayarathna:2020:MSJ

[DS20a] Miyuru Dayarathna and Toyotaro Suzumura. Multiple stream job performance optimization with source operator graph transformations. Concurrency and Computation: Practice and Experience, 32(16):e5658:1–e5658:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devi:2020:ETS

[DS20b] E. M. Roopa Devi and R. C. Suganthe. Enhanced transductive support vector machine classification with grey wolf optimizer cuckoo search optimization for intrusion detection system. Concurrency and Computation: Practice and Experience, 32(4):e4999:1–e4999:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dashti:2021:HRS

[DS21] Seyed Ebrahim Dashti and Maryam Sarafraz. Hybrid recommender system based on fuzzy neural algorithm. Concurrency and Computation: Practice and Experience, 33(24):e6473:1– e6473:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devi:2022:EER

[DS22a] R. Renuga Devi and T. Sethukarasi. An energy-efficient routing based on a hybrid improved whale artificial ecosystem optimization algorithm in WSN. Concurrency and Computation: Practice and Experience, 34(4):e6639:1–e6639:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dmitruk:2022:STT

[DS22b] Beata Dmitruk and Przemysław Stpiczyński. Solving tridiagonal Toeplitz systems of linear equations on GPU-accelerated computers. Concurrency and Computation: Practice and Experience, 34(14):e6449:1–e6449:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Durairaj:2022:TSV

[DS22c] Selvam Durairaj and Rajeswari Sridhar. Task scheduling to a virtual machine using a multi-objective mayfly approach for a cloud environment. Concurrency and Computation: Practice and Experience, 34(24):e7236:1–e7236:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Daniya:2023:SSS

[DS23a] Thavasilingam Daniya and Vigneshwari Srinivasan. Shuffled shepherd social optimization based deep learning for rice leaf disease classification and severity percentage prediction. Concurrency and Computation: Practice and Experience, 35(4): e7523:1–e7523:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dmitruk:2023:IAS

[DS23b] Beata Dmitruk and Przemysław Stpiczyński. Improving accuracy of summation using parallel vectorized Kahan’s and Gill– Møller algorithms. Concurrency and Computation: Practice and Experience, 35(23):e7763:1–e7763:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Denninnart:2024:SSP

[DS24a] Chavit Denninnart and Mohsen Amini Salehi. SMSE: a serverless platform for multimedia cloud systems. Concurrency and Computation: Practice and Experience, 36(4): e7922:1–e7922:??, February 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dwivedi:2024:NAC

[DS24b] Pankaj Prasad Dwivedi and Dilip Kumar Sharma. A novel approach for constructing privacy-aware architecture utilizing Shannon’s entropy. Concurrency and Computation: Practice and Experience, 36(11):e8030:1–e8030:??, May 15, 2024. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deng:2021:TSH

[DSC^{+}21a] Zexi Deng, Hong Shen, Dunqian Cao, Zihan Yan, and Huimin Huang. Task scheduling on heterogeneous multiprocessor systems through coherent data allocation. Concurrency and Computation: Practice and Experience, 33(10):e6183:1– e6183:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Duan:2021:GRB

[DSC^{+}21b] Haojie Duan, Ying Sun, Wentao Cheng, Du Jiang, Juntong Yun, Ying Liu, Yibo Liu, and Dalin Zhou. Gesture recognition based on multi-modal feature weight. Concurrency and Computation: Practice and Experience, 33(5):e5991:1– e5991:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dewangan:2022:EDL

[DSJ22] Kranti Kumar Dewangan, Satya Prakash Sahu, and Rekhram Janghel. Enhanced deep learning frame model for an accurate segmentation of cancer affected part in breast. Concurrency and Computation: Practice and Experience, 34(28):e7379:1– e7379:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

daSilva:2023:SRA

[dSNdL^{+}23] Vinicius S. da Silva, Angelo G. D. Nogueira, Everton Camargo de Lima, Hiago M. G. de A. Rocha, Matheus S. Serpa, Marcelo C. Luizelli, Fábio D. Rossi, Philippe O. A. Navaux, Antonio Carlos S. Beck, and Arthur Francisco Lorenzon. Smart resource allocation of concurrent execution of parallel applications. Concurrency and Computation: Practice and Experience, 35(17):e6600:1–e6600:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Diaz:2021:STC

[DSS21] Silvia M. D. Diaz, Paulo S. L. Souza, and Simone R. S. Souza. Structural testing for communication events into loops of message-passing parallel programs. Concurrency and Computation: Practice and Experience, 33(18):e6082:1–e6082:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dubey:2023:DRS

[DSS^{+}23] Gaurav Dubey, Harivans Pratap Singh, Kavita Sheoran, Geetika Dhand, and Pooja Malik. Drug review sentimental analysis based on modular lexicon generation and a fusion of bidirectional threshold weighted mapping CNN-RNN. Concurrency and Computation: Practice and Experience, 35 (3):e7512:1–e7512:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Doraiswami:2022:JTS

[DSSS22] Palanivel Rajan Doraiswami, Velliangiri Sarveshwaran, Iwin Thanakumar Joseph Swamidason, and Sona Chandra Devadass Sorna. Jaya-tunicate swarm algorithm based generative adversarial network for COVID-19 prediction with chest computed tomography images. Concurrency and Computation: Practice and Experience, 34(23):e7211:1–e7211:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Darehnaei:2022:SES

[DSYF22] Zeinab Ghasemi Darehnaei, Mohammad Shokouhifar, Hossein Yazdanjouei, and Seyed Mohammad Jalal Rastegar Fatemi. SI-EDTL: Swarm intelligence ensemble deep transfer learning for multiple vehicle detection in UAV images. Concurrency and Computation: Practice and Experience, 34(5):e6726:1– e6726:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devaraj:2022:PDM

[DT22a] Kavitha Devaraj and Balasubramanain Thangavel. Pregnancy delivery mode prediction using dichotomous radial basis Tanimoto network. Concurrency and Computation: Practice and Experience, 34(13):e6928:1–e6928:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Devi:2022:OSN

[DT22b] Kalyanee Devi and Rohit Tripathi. Optimal seed node selection method for LTIS model. Concurrency and Computation: Practice and Experience, 34(15):e6982:1–e6982:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dudukcu:2024:UIP

[DTK24] Hatice Vildan Dudukcu, Murat Taskiran, and Nihan Kahraman. UAV instantaneous power consumption prediction using LR-TCN with simple moving average. Concurrency and Computation: Practice and Experience, 36(3):e7913:1–e7913:??, February 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dai:2020:MDA

[DWDG20] Yinglong Dai, Guojun Wang, Jianhua Dai, and Oana Geman. A multimodal deep architecture for traditional Chinese medicine diagnosis. Concurrency and Computation: Practice and Experience, 32(19):e5781:1–e5781:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2021:CBS

[DWY^{+}21] Yuanze Du, Qiang Wu, Yi Yao, Yingwang Zhao, Xiaoyan Zhang, Zhichao Hao, Hua Xu, and Zhili Du. A cloud-based service for emergency evacuation of mine water inrush accidents. Concurrency and Computation: Practice and Experience, 33(9):e6125:1–e6125:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ding:2020:WOA

[DWZ20a] Hangqi Ding, Zhiyong Wu, and Luchen Zhao. Whale optimization algorithm based on nonlinear convergence factor and chaotic inertial weight. Concurrency and Computation: Practice and Experience, 32(24):e5949:1–e5949:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Du:2020:PTV

[DWZ^{+}20b] Yuanze Du, Qiang Wu, Yingwang Zhao, Xiaoyan Zhang, Yi Yao, and Hua Xu. A parallel time-varying earliest arrival path algorithm for evacuation planning of underground mine water inrush accidents. Concurrency and Computation: Practice and Experience, 32(11):e5644:1–e5644:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dai:2022:PLR

[DWZ^{+}22] Xuefeng Dai, Jiazhi Wang, Jianqi Zhao, Dahui Li, and Zhifeng Yao. Predicate logic reasoning for exploration coordination of multi-robot systems in structured environments. Concurrency and Computation: Practice and Experience, 34(9):e5806:1– e5806:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2020:TSB

[DXXL20] Tingting Dong, Fei Xue, Chuangbai Xiao, and Juntao Li. Task scheduling based on deep reinforcement learning in a cloud manufacturing environment. Concurrency and Computation: Practice and Experience, 32(11):e5654:1–e5654:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2020:EIB

[DYF20] Fang Dong, Jianming Yong, and Xiang Fei. Emerging intelligent big data analytics for cloud and edge computing. Concurrency and Computation: Practice and Experience, 32(23): e5989:1–e5989:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Deng:2022:BSN

[DZCL22] Qi Deng, Shanshan Zhang, Gang Chen, and Huaxiang Lu. Blind separation of noncooperative paired carrier multiple access signals based on improved quantum-inspired evolutionary algorithm and receding horizon optimization. Concurrency and Computation: Practice and Experience, 34(12):e6119:1– e6119:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ding:2020:MUR

[DZLH20] Xiaobing Ding, Yuan Zhao, Zhigang Liu, and Hua Hu. The modeling of urban rail transit emergency delay propagation scope under network operation mode. Concurrency and Computation: Practice and Experience, 32(23):e5530:1–e5530:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dagnaw:2023:SSS

[DZW23] Girum Dagnaw, Ke Zhou, and Hua Wang. SACRO: Solid state drive-assisted chunk caching for restore optimization. Concurrency and Computation: Practice and Experience, 35 (18):e6162:1–e6162:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2022:CRS

[DZYY22] SongShou Dong, YiHua Zhou, YuGuang Yang, and Yanqing Yao. A certificateless ring signature scheme based on lattice. Concurrency and Computation: Practice and Experience, 34(28):e7385:1–e7385:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ertam:2022:ATS

[EA22] Fatih Ertam and Galip Aydin. Abstractive text summarization using deep learning with a new Turkish summarization benchmark dataset. Concurrency and Computation: Practice and Experience, 34(9):e6482:1–e6482:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

El-Ashri:2021:ACM

[EABZB21] Nader K. El-Ashri, Ehab F. Badran, Amira I. Zaki, and Waleed K. Badawi. Admission control mechanism for quality of service and security in H.323 voice gateway. Concurrency and Computation: Practice and Experience, 33(20):e6376:1– e6376:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Erol:2023:SBD

[EAS23] Salih Erdem Erol, Çagla Aksoy, and Seref Sagiroglu. Social big data applications and challenges. Concurrency and Computation: Practice and Experience, 35(5):e7567:1–e7567:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elshamy:2022:RTC

[EASN22] Mohamed R. Elshamy, Belal Abozalam, Amged Sayed, and Essam Nabil. Real-time control design and implementation of ball balancer system based on machine learning and machine vision. Concurrency and Computation: Practice and Experience, 34(27):e7317:1–e7317:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Esmat:2022:PHB

[EASR22] Aghaee-Meybodi Esmat, Nezarat Amin, Emadi Sima, and Ghaffari Mohammad Reza. A parallel hash-based method for local sequence alignment. Concurrency and Computation: Practice and Experience, 34(3):e6568:1–e6568:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ezhilchelvan:2020:NBT

[EAvM20] Paul Ezhilchelvan, Amjad Aldweesh, and Aad van Moorsel. Non-blocking two-phase commit using blockchain. Concurrency and Computation: Practice and Experience, 32(12): e5276:1–e5276:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elbez:2022:PCW

[EBDB22] Hammouda Elbez, Mohammed Kamel Benhaoua, Philippe Devienne, and Pierre Boulet. Progressive compression and weight reinforcement for spiking neural networks. Concurrency and Computation: Practice and Experience, 34(11): e6891:1–e6891:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eiling:2022:CVL

[EBLM22] Niklas Eiling, Jonas Baude, Stefan Lankes, and Antonello Monti. Cricket: a virtualization layer for distributed execution of CUDA applications with checkpoint/restart support. Concurrency and Computation: Practice and Experience, 34 (14):e6474:1–e6474:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elmazi:2020:ESG

[ECIB20] Donald Elmazi, Miralda Cuka, Makoto Ikeda, and Leonard Barolli. Effect of size of giant component for actor node selection in WSANs: a comparison study. Concurrency and Computation: Practice and Experience, 32(8):e5106:1–e5106:??, April 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elleuch:2021:LDM

[EGB21] Imen Elleuch, Bilel Gargouri, and Abdelmajid Ben Hamadou. Lexical data mining-based approach for the self-enrichment of LMF standardized dictionaries: Case of the syntacticosemantic knowledge. Concurrency and Computation: Practice and Experience, 33(17):e6312:1–e6312:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maarouf:2023:CRS

[EGGG23] Aboul-Karim Mohamed El Maarouf, Luc Giraud, Abdou Guermouche, and Thomas Guignon. Combining reduction with synchronization barrier on multi-core processors. Concurrency and Computation: Practice and Experience, 35(1): e7402:1–e7402:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elsayad:2021:PNC

[EHST21] Dina Elsayad, Safwat Hamad, Howida A. Shedeed, and Mohamed F. Tolba. Parallel network component analysis technique for gene regulatory network inference. Concurrency and Computation: Practice and Experience, 33(24):e6458:1– e6458:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ennahbaoui:2022:NAB

[EI22] Mohammed Ennahbaoui and Hind Idrissi. A new agent-based framework combining authentication, access control and user behavior analysis for secure and flexible cloud-based healthcare environment. Concurrency and Computation: Practice and Experience, 34(5):e6712:1–e6712:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eitschberger:2020:COH

[EK20] Patrick Eitschberger and Jörg Keller. Comparing optimal and heuristic taskgraph scheduling on parallel machines with frequency scaling. Concurrency and Computation: Practice and Experience, 32(10):e5396:1–e5396:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eken:2022:USW

[Eke22] Enes Eken. Using subspaces of weight matrix for evaluating generative adversarial networks with Fréchet distance. Concurrency and Computation: Practice and Experience, 34 (1):e6478:1–e6478:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Epsiba:2020:NDC

[EKS20] P. Epsiba, N. Kumaratharan, and G. Suresh. A novel discrete CURVELET transform and modified WHOG for video surveillance services. Concurrency and Computation: Practice and Experience, 32(4):e5046:1–e5046:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elen:2022:CDR

[Ele22] Abdullah Elen. Covid-19 detection from radiographs by feature-reinforced ensemble learning. Concurrency and Computation: Practice and Experience, 34(23):e7179:1–e7179:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elleuch:2023:MPB

[ELIG23] Marwa Elleuch, Nassim Laga, Oumaima Alaoui Ismaili, and Walid Gaaloul. Multi-perspective business process discovery from messaging systems: State-of-the art. Concurrency and Computation: Practice and Experience, 35(11):e6642:1– e6642:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elkhoukhi:2020:PAO

[ENB^{+}20] Hamza Elkhoukhi, Youssef NaitMalek, Mohamed Bakhouya, Anass Berouine, Abdelhak Kharbouch, Fadwa Lachhab, Majdoulayne Hanifi, Driss El Ouadghiri, and Mohamed Essaaidi. A platform architecture for occupancy detection using stream processing and machine learning approaches. Concurrency and Computation: Practice and Experience, 32(17):e5651:1– e5651:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eili:2022:SRP

[ER22] Mansoureh Yari Eili and Jalal Rezaeenour. A survey on recommendation in process mining. Concurrency and Computation: Practice and Experience, 34(26):e7304:1–e7304:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elloumi:2020:ARM

[ESB20] Farah Bellaaj Elloumi, Mohamed Sellami, and Sami Bhiri. Avoiding resource misallocations in business processes. Concurrency and Computation: Practice and Experience, 32(15): e4888:1–e4888:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Enturi:2023:ISC

[ESS23] Bala Krishna Manash Enturi, A. Suhasini, and Narayana Satyala. Intelligent skin cancer diagnosis using adaptive kmeans segmentation and deep learning models. Concurrency and Computation: Practice and Experience, 35(4):e7546:1– e7546:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ekane:2023:NNG

[ETH^{+}23] Brice Ekane, Alain Tchana, Daniel Hagimont, Boris Teabe, and Noel De Palma. Networking in next generation disaggregated datacenters. Concurrency and Computation: Practice and Experience, 35(21):e7702:1–e7702:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ergunsah:2023:EEA

[ETKD23] Senol Ergunsah, Vedat Tümen, Selahattin Kosunalp, and Kubilay Demir. Energy-efficient animal tracking with multiunmanned aerial vehicle path planning using reinforcement learning and wireless sensor networks. Concurrency and Computation: Practice and Experience, 35(4):e7527:1–e7527:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Erol:2022:AED

[EUYY22] Gizemnur Erol, Betül Uzbas, Cüneyt Yücelbas, and Sule Yücelbas. Analyzing the effect of data preprocessing techniques using machine learning algorithms on the diagnosis of COVID-19. Concurrency and Computation: Practice and Experience, 34(28):e7393:1–e7393:??, December 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Elangovan:2021:LBS

[EVVR21] Rajalakshmi Elangovan, Subramaniyaswamy Vairavasundaram, Vijayakumar Varadarajan, and Logesh Ravi. Locationbased social network recommendations with computational intelligence-based similarity computation and user check-in behavior. Concurrency and Computation: Practice and Experience, 33(22):e6106:1–e6106:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Francies:2022:RMO

[FAM22] Mariam L. Francies, Mohamed M. Ata, and Mohamed A. Mohamed. A robust multiclass 3D object recognition based on modern YOLO deep learning algorithms. Concurrency and Computation: Practice and Experience, 34(1):e6517:1– e6517:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fazlollahtabar:2023:RDR

[Faz23] Hamed Fazlollahtabar. Rough DT^{2}R^{2}ML for renewable energy supplier selection in the presence of big data. Concurrency and Computation: Practice and Experience, 35(23): e7773:1–e7773:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2021:ADP

[FC21] Tian Fan and Zhihua Cui. Adaptive differential privacy preserving based on multi-objective optimization in deep neural networks. Concurrency and Computation: Practice and Experience, 33(20):e6367:1–e6367:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Filho:2020:ARP

[FCMM20] José S. Costa Filho, Denis M. Cavalcante, Leonardo O. Moreira, and Javam C. Machado. An adaptive replica placement approach for distributed key-value stores. Concurrency and Computation: Practice and Experience, 32(11):e5675:1– e5675:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2022:BBA

[FCX^{+}22] Qing Fan, Jianhua Chen, Feng Xu, Li Li, and Min Luo. A biometrics-based anonymous authentication and key agreement scheme for wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(16):e6178:1– e6178:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fu:2020:MIS

[FCZ20] Hong Fu, Bo Chen, and Bin Zhang. A multi-Internet service provider game: Equilibrium, stability, and characteristics. Concurrency and Computation: Practice and Experience, 32 (14):e5700:1–e5700:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ferretti:2020:FSI

[FD20a] Stefano Ferretti and Gabriele D’Angelo. Foreword to the special issue on cryptocurrencies and blockchains for distributed systems. Concurrency and Computation: Practice and Experience, 32(12):e5539:1–e5539:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ferretti:2020:EBS

[FD20b] Stefano Ferretti and Gabriele D’Angelo. On the Ethereum blockchain structure: a complex networks theory perspective. Concurrency and Computation: Practice and Experience, 32 (12):e5493:1–e5493:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fiorini:2022:VPT

[FD22] Arnaud Fiorini and Michel R. Dagenais. Visualization of profiling and tracing in CPU-GPU programs. Concurrency and Computation: Practice and Experience, 34(23):e7188:1– e7188:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Feng:2024:FDW

[FDD^{+}24] Rongqiang Feng, Hongwei Du, Tianchi Du, Xueqiong Wu, Haiping Yu, Kexin Zhang, Chenxi Huang, and Lianlian Cao.

Fault diagnosis for wind turbines based on LSTM and feature optimization strategies. Concurrency and Computation: Practice and Experience, 36(1):e7886:1–e7886:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fehri:2022:AQA

[FDH22] Hela Fehri, Sondes Dardour, and Kais Haddar. ARmed question answering system. Concurrency and Computation: Practice and Experience, 34(21):e7054:1–e7054:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Feng:2021:MFP

[FDY21] Hao Feng, Yuhui Deng, and Liang Yu. Modeling the failures of power-aware data centers by leveraging heat recirculation. Concurrency and Computation: Practice and Experience, 33 (7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Farghaly:2022:NFS

[FEH22] Heba Mamdouh Farghaly and Tarek Abd El-Hafeez. A new feature selection method based on frequent and associated itemsets for text classification. Concurrency and Computation: Practice and Experience, 34(25):e7258:1–e7258:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fakih:2020:GDE

[FEK20] Bilal Fakih, Didier El Baz, and Igor Kotenko. GRIDHPC: a decentralized environment for high performance computing. Concurrency and Computation: Practice and Experience, 32(10):e5320:1–e5320:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Faverge:2023:PHA

[FFG^{+}23] Mathieu Faverge, Nathalie Furmento, Abdou Guermouche, Gwenolé Lucas, Raymond Namyst, Samuel Thibault, and Pierre-André Wacrenier. Programming heterogeneous architectures using hierarchical tasks. Concurrency and Computation: Practice and Experience, 35(25):e7811:1–e7811:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fortin:2021:HPS

[FFLM21] Pierre Fortin, Ambroise Fleury, François Lemaire, and Michael Monagan. High-performance SIMD modular arithmetic for polynomial evaluation. Concurrency and Computation: Practice and Experience, 33(16):e6270:1–e6270:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fischer:2021:LLI

[FGJ^{+}21] Joel E. Fischer, Chris Greenhalgh, Wenchao Jiang, Sarvapali D. Ramchurn, Feng Wu, and Tom Rodden. In-the-loop or on-the-loop? interactional arrangements to support team coordination with a planning agent. Concurrency and Computation: Practice and Experience, 33(8):e4082:1–e4082:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ferreira:2020:HMM

[FGL^{+}20] Kurt Ferreira, Ryan E. Grant, Michael J. Levenhagen, Scott Levy, and Taylor Groves. Hardware MPI message matching: Insights into MPI matching behavior to inform design. Concurrency and Computation: Practice and Experience, 32 (3):e5150:1–e5150:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2023:MOO

[FGZC23] Tian Fan, Wanwan Guo, Zhixia Zhang, and Zhihua Cui. A many-objective optimization based intelligent algorithm for virtual machine migration in mobile edge computing. Concurrency and Computation: Practice and Experience, 35(23): e7770:1–e7770:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2020:RFT

[FHH^{+}20] Weibei Fan, Jing He, Zhijie Han, Peng Li, and Ruchuan Wang. Reconfigurable fault-tolerance mapping of ternary Ncubes onto chips. Concurrency and Computation: Practice and Experience, 32(11):e5659:1–e5659:??, June 10, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Feng:2022:SIT

[FHS^{+}22] Qiaosheng Feng, Li Huang, Ying Sun, Xiliang Tong, Xin Liu, Yuanmin Xie, Jun Li, Hanwen Fan, and Baojia Chen. Substation instrumentation target detection based on multi-scale feature fusion. Concurrency and Computation: Practice and Experience, 34(23):e7177:1–e7177:??, October 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fiore:2020:MOB

[Fio20] Ugo Fiore. Minority oversampling based on the attractionrepulsion Weber problem. Concurrency and Computation: Practice and Experience, 32(18):e5601:1–e5601:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fagundes-Junior:2023:HPG

[FJCFB23] Leonardo Fagundes-Junior, Michael Canesche, Ricardo Ferreira, and Alexandre Brandão. High-performance graphics processing unit-based strategy for tuning a unmanned aerial vehicle controller subject to time-delay constraints. Concurrency and Computation: Practice and Experience, 35(24): e7767:1–e7767:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fenil:2020:SDD

[FK20] E. Fenil and P. Mohan Kumar. Survey on DDoS defense mechanisms. Concurrency and Computation: Practice and Experience, 32(4):e5114:1–e5114:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

F:2024:ECS

[FK24] Margret Sharmila F and K. Karuppasamy. An effective cryptanalysis of DES for secure communication using hybrid cryptanalysis and deep neural network. Concurrency and Computation: Practice and Experience, 36(1):e7872:1–e7872:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fisne:2022:EHP

[FKGO22] Alparslan Fisne, Berkan Kilic, Alper Güngör, and Adnan Ozsoy. Efficient heterogeneous parallel programming for compressed sensing based direction of arrival estimation. Concurrency and Computation: Practice and Experience, 34(9): e6490:1–e6490:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Funika:2023:CSA

[FKK23] Włodzimierz Funika, Paweł Koperek, and Jacek Kitowski. Continuous self-adaptation of control policies in automatic cloud management. Concurrency and Computation: Practice and Experience, 35(20):e7371:1–e7371:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Finnerty:2022:SAT

[FKO22] Patrick Finnerty, Tomio Kamada, and Chikara Ohta. A selfadjusting task granularity mechanism for the Java lifelinebased global load balancer library on many-core clusters. Concurrency and Computation: Practice and Experience, 34 (2):e6224:1–e6224:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Finnerty:2023:ABR

[FKO23] Patrick Finnerty, Tomio Kamada, and Chikara Ohta. Automatically balancing relocatable distributed collections. Concurrency and Computation: Practice and Experience, 35(27): e7717:1–e7717:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fisne:2023:EEC

[FKY^{+}23] Alparslan Fisne, Alperen Kalay, Faruk Yavuz, Cagri Cetintepe, and Adnan Ozsoy. Energy-efficient computing for machine learning based target detection. Concurrency and Computation: Practice and Experience, 35(24):e7582:1– e7582:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Farias:2023:MMD

[FLB23] Paulo V. G. Farias, Tulio A. Lima, and Jacir L. Bordim. Mitigating message dissemination issues in safety applications for vehicular ad hoc networks. Concurrency and Computation: Practice and Experience, 35(11):e6034:1–e6034:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2022:NZW

[FLG^{+}22] Di Fan, Yongyi Li, Shang Gao, Wanda Chi, and Changzhi Lv. A novel zero watermark optimization algorithm based on Gabor transform and discrete cosine transform. Concurrency and Computation: Practice and Experience, 34(14):e5689:1– e5689:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fragalla:2020:NLF

[FLP20] John Fragalla, Bill Loewe, and Torben Kling Petersen. New Lustre features to improve Lustre metadata and small-file performance. Concurrency and Computation: Practice and Experience, 32(20):e5649:1–e5649:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fayala:2020:WSR

[FM20] Mayssa Fayala and Haithem Mezni. Web service recommendation based on time-aware users clustering and multi-valued QoS prediction. Concurrency and Computation: Practice and Experience, 32(9):e5603:1–e5603:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Faris:2022:IMD

[FM22] Najlaa Nsrulaah Faris and Firsas Saber Miften. An intelligence model for detection of PCOS based on k-means coupled with LS-SVM. Concurrency and Computation: Practice and Experience, 34(21):e7139:1–e7139:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fatima:2022:VLS

[FMJ^{+}22] Joddat Fatima, Mashood Mohsan, Amina Jameel, Muhammad Usman Akram, and Adeel Muzaffar Syed. Vertebrae localization and spine segmentation on radiographic images for feature-based curvature classification for scoliosis. Concurrency and Computation: Practice and Experience, 34(26): e7300:1–e7300:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fishani:2022:MOL

[FMNF22] Behzad Aghaei Fishani, Ali Mahmoodirad, Sadegh Niroomand, and Mohammad Fallah. Multi-objective locationallocation-routing problem of perishable multi-product supply chain with direct shipment and open routing possibilities under sustainability. Concurrency and Computation: Practice and Experience, 34(11):e6860:1–e6860:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2023:BFV

[FN23] Xing Fan and Baoning Niu. A Bloom filter variant for blockchain. Concurrency and Computation: Practice and Experience, 35(16):e6643:1–e6643:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fecilak:2020:NIS

[FPÁ^{+}20] Peter Fecilák, Adrián Pekár, Norbert Ádám, Anton Baláz, and Eva Chovancová. A non-intrusive smart metering system: Analytics and simulation of power consumption. Concurrency and Computation: Practice and Experience, 32(13):e5250:1– e5250:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2019:RME

[FPHZ19] Linjing Fan, Xuchao Pan, Zhengxiang Huang, and Xudong Zu. Retracted: The mechanism and experimental study on the interference of high voltage lines to navigation system. Concurrency and Computation: Practice and Experience, 31 (12):e4718:1–e4718:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See retraction notice [Pan20].

Fan:2020:RSM

[FPHZ20] Linjing Fan, Xuchao Pan, Zhengxiang Huang, and Xudong Zu. Retracted: Study on measurement reliability based on Liu estimator. Concurrency and Computation: Practice and Experience, 31(12):e5560:1–e5560:??, January 25, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See [HML19].

Filelis-Papadopoulos:2021:SAA

[FPXM21] Christos Filelis-Papadopoulos, Huanhuan Xiong, and John Morrison. A scalable and adaptable allocation framework for heterogeneous resources in a large cluster environment. Concurrency and Computation: Practice and Experience, 33(14): e5564:1–e5564:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fang:2023:PTD

[FQD^{+}23] Zisen Fang, Fumin Qi, Yichuan Dong, Yong Zhang, and Shengzhong Feng. Parallel tensor decomposition with distributed memory based on hierarchical singular value decomposition. Concurrency and Computation: Practice and Experience, 35(17):e6656:1–e6656:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fernandez:2023:POD

[FRS^{+}23] F. Mary Harin Fernandez, T. Venkata Ramana, Mahammad Shabana, V. Kannagi, and M. Nalini. Personalized ontology and deep training tree-based optimal gated recurrent unitrecurrent neural network for prediction of students’ behavior. Concurrency and Computation: Practice and Experience, 35 (1):e7420:1–e7420:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fahad:2022:MQP

[FSA^{+}22] Muhammad Fahad, Mohammad Shojafar, Mubashir Abbas, Israr Ahmed, and Humaira Ijaz. A multi-queue prioritybased task scheduling algorithm in fog computing environment. Concurrency and Computation: Practice and Experience, 34(28):e7376:1–e7376:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ferro:2023:TSA

[FSdP^{+}23] Mariza Ferro, Gabrieli D. Silva, Felipe B. de Paula, Vitor Vieira, and Bruno Schulze. Towards a sustainable artificial intelligence: a case study of energy efficiency in decision tree algorithms. Concurrency and Computation: Practice and Experience, 35(17):e6815:1–e6815:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fayyaz:2022:VQV

[FSFM22] Mohsin Fayyaz, Yasir Ali Shah, Ahmed Fayyaz, and Ghulam Mujtaba. VLCC-Q: Very low computational complexity optical interconnect architecture with queueing for reducing delay and back pressure probability in data center networks. Concurrency and Computation: Practice and Experience, 34(18): e7018:1–e7018:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Feng:2021:GIF

[FSWW21] Xi Feng, Huanping Shi, Jian Wang, and Shaoguang Wang. Green intelligent financial system construction paradigm based on deep learning and concurrency models. Concurrency and Computation: Practice and Experience, 33(12):e5784:1– e5784:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Feng:2023:R

[FSWW23] Xi Feng, Huanping Shi, Jian Wang, and Shaoguang Wang. Retraction. Concurrency and Computation: Practice and Experience, 33(12):e7755:1–e7755:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fu:2022:VMR

[FXX22] Wei Fu, Zhenjie Xie, and Jianqiao Xu. Verifying multiple replica possession with public key segment under public cloud storage environment. Concurrency and Computation: Practice and Experience, 34(5):e6699:1–e6699:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fan:2021:DPC

[FYH^{+}21] Tanghuai Fan, Zhanfeng Yao, Longzhe Han, Baohong Liu, and Li Lv. Density peaks clustering based on k-nearest neighbors sharing. Concurrency and Computation: Practice and Experience, 33(5):e5993:1–e5993:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fadil:2023:PAC

[FZ23] Ali Can Fadil and Baha Zafer. Parallel aeroacoustic computation of unsteady transonic weapon bay using detached-eddy simulation via open computational fluid dynamics source codes. Concurrency and Computation: Practice and Experience, 35(24):e7675:1–e7675:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fu:2022:SES

[FZA22] Shaojing Fu, Chao Zhang, and Weijun Ao. Searchable encryption scheme for multiple cloud storage using double-layer blockchain. Concurrency and Computation: Practice and Experience, 34(16):e5860:1–e5860:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Furtak:2020:PSN

[FZC20] Janusz Furtak, Zbigniew Zieliński, and Jan Chudzikiewicz. Procedures for sensor nodes operation in the secured domain. Concurrency and Computation: Practice and Experience, 32 (13):e5183:1–e5183:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fernandes:2022:SIR

[FZT22] Steven Fernandes, Yu-Dong Zhang, and João Manuel R. S. Tavares. Special issue: Recent advances in quantum computing and quantum neural networks. Concurrency and Computation: Practice and Experience, 34(20):e7203:1–e7203:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Feng:2023:IDR

[FZZZ23] Junkai Feng, Haibin Zhang, Kaili Zhang, and Pengfei Zhao. An inertial Douglas-Rachford splitting algorithm for nonconvex and nonsmooth problems. Concurrency and Computation: Practice and Experience, 35(17):e6343:1–e6343:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ghalan:2022:DLH

[GA22] Mamta Ghalan and Rajesh Kumar Aggarwal. Daily life human activities recognition using a novel AR-DenseNet. Concurrency and Computation: Practice and Experience, 34(22): e7142:1–e7142:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gola:2023:UAS

[GA23] Kamal Kumar Gola and Shikha Arya. Underwater acoustic sensor networks: Taxonomy on applications, architectures, localization methods, deployment techniques, routing techniques, and threats: a systematic review. Concurrency and Computation: Practice and Experience, 35(23):e7815:1– e7815:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gutierrez:2020:MAP

[GADM20] Samuel K. Gutiérrez, Dorian C. Arnold, Kei Davis, and Patrick McCormick. On the memory attribution problem: a solution and case study using MPI. Concurrency and Computation: Practice and Experience, 32(3):e5159:1–e5159:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gafsi:2022:ICR

[GAH^{+}22] Mohamed Gafsi, Rim Amdouni, Mohamed Ali Hajjaji, Jihene Malek, and Abdellatif Mtibaa. Improved chaos-RSA-based hybrid cryptosystem for image encryption and authentication. Concurrency and Computation: Practice and Experience, 34 (23):e7187:1–e7187:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guttula:2023:HOA

[GAM23] Ramakrishna Guttula, Ramesh Adireddy, and Venkateswarlu Mannepally. Hybrid optimization algorithm for optimal designing of microstrip patch antenna. Concurrency and Computation: Practice and Experience, 35(7):e7603:1–e7603:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gastineau:2022:LEL

[GAMT22] Nicolas Gastineau, Wahabou Abdou, Nader Mbarek, and Olivier Togni. Leader election and local identifiers for threedimensional programmable matter. Concurrency and Computation: Practice and Experience, 34(7):e6067:1–e6067:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Garg:2023:IAD

[GAS23] Aakansha Garg, Rajeev Arya, and Maheshwari Prasad Singh. An integrated approach for dual resource optimization of relay-based mobile edge computing system. Concurrency and Computation: Practice and Experience, 35(10):e7682:1– e7682:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guney:2022:DNN

[GATK22] Hanife Guney, Melek Aydin, Murat Taskiran, and Nihan Kahraman. A deep neural network based toddler tracking system. Concurrency and Computation: Practice and Experience, 34(14):e6636:1–e6636:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ghorbani:2020:RDT

[GB20] Mehrdad Ghorbani and Seyed Morteza Babamir. Runtime deadlock tracking and prevention of concurrent multithreaded programs: a learning-based approach. Concurrency and Computation: Practice and Experience, 32(10):e5324:1–e5324:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2023:TEH

[GB23] Sarika Gupta and Himani Bansal. Trust evaluation of health websites by eliminating phishing websites and using similarity techniques. Concurrency and Computation: Practice and Experience, 35(21):e7695:1–e7695:??, September 25, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gunen:2022:KSM

[GBB22] Mehmet Akif Günen, Pinar Çivicioglu Besdok, and Erkan Besdok. Keypose synthesis from 3D motion capture data by using evolutionary clustering. Concurrency and Computation: Practice and Experience, 34(1):e6485:1–e6485:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gaffour:2021:NEM

[GBBS21] Khadidja Gaffour, Mohammed Kamel Benhaoua, Abou El Hassan Benyamina, and Amit Kumar Singh. A new efficient multi-task applications mapping for three-dimensional Network-on-Chip based MPSoC. Concurrency and Computation: Practice and Experience, 33(10):e6194:1–e6194:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gholamnezhad:2024:IMB

[GBS24] Pezhman Gholamnezhad, Ali Broumandnia, and Vahid Seydi. An improved model-based evolutionary algorithm for multiobjective optimization. Concurrency and Computation: Practice and Experience, 36(10):e6566:1–e6566:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guo:2020:ISV

[GC20] Chengjun Guo and Yongqi Chen. An interval support vector domain description based on the dynamic decreasing inertia weight particle swarm optimization. Concurrency and Computation: Practice and Experience, 32(23):e5591:1–e5591:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grutzmacher:2020:CPF

[GCF^{+}20] Thomas Grützmacher, Terry Cojean, Goran Flegar, Fritz Göbel, and Hartwig Anzt. A customized precision format based on mantissa segmentation for accelerating sparse linear algebra. Concurrency and Computation: Practice and Experience, 32(15):e5418:1–e5418:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gnanasekaran:2022:AQP

[GCP22] Amirthayogam Gnanasekaran, Anbu Ananth Chinnasamy, and Elango Parasuraman. Analyzing the QoS prediction for web service recommendation using time series forecasting with deep learning techniques. Concurrency and Computation: Practice and Experience, 34(28):e7356:1–e7356:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gu:2024:MMA

[GCQ^{+}24] Yang Gu, Jian Cao, Shiyou Qian, Nengjun Zhu, and Wei Guan. MANSOR: a module alignment method based on neighbor information for scientific workflow. Concurrency and Computation: Practice and Experience, 36(10):e7736:1– e7736:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gil-Costa:2020:HPC

[GCS20] Veronica Gil-Costa and Hermes Senger. High-performance computing for computational science. Concurrency and Computation: Practice and Experience, 32(20):e5904:1–e5904:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ganji:2023:SSP

[GCS23] Venkata Ratnam Ganji, Aparna Chaparala, and Radhika Sajja. Shuffled shepherd political optimization-based deep learning method for credit card fraud detection. Concurrency and Computation: Practice and Experience, 35(10):e7666:1– e7666:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gong:2024:RHS

[GCSZ24] Ao Gong, Yun Cheng, Jing Su, and Li Zhang. Research on hybrid synchronization methods in multi-user collaborative VR simulation medical surgery training system. Concurrency and Computation: Practice and Experience, 36(11):e8008:1– e8008:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2022:NUT

[GD22] Ashlesha Gupta and Ashutosh Dixit. A novel user trend-based priority assigner and URL scheduler for dynamic incremental crawling. Concurrency and Computation: Practice and Experience, 34(3):e6555:1–e6555:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Garriga:2021:BCC

[GDA^{+}21] Martin Garriga, Stefano Dalla Palma, Maxmiliano Arias, Alan De Renzis, Remo Pareschi, and Damian Andrew Tamburri. Blockchain and cryptocurrencies: a classification and comparison of architecture drivers. Concurrency and Computation: Practice and Experience, 33(8):e5992:1–e5992:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ganasigamony:2022:EMB

[GDA22] Wiselin Jiji Ganasigamony, Johnson Durairaj, and Rajesh Athiswamy. An empirical model based environmental pollution level analysis in coastal area of Gulf of Mannar using remote sensing data. Concurrency and Computation: Practice and Experience, 34(11):e6859:1–e6859:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2022:FFC

[GDCW22] Zicong Gao, Weiyu Dong, Rui Chang, and Yisen Wang. Fwfuzz: a code coverage-guided fuzzing framework for network protocols on firmware. Concurrency and Computation: Practice and Experience, 34(16):e5756:1–e5756:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gholami-Dastgerdi:2022:NED

[GDFDF22] Pejman Gholami-Dastgerdi, Mohammad-Reza Feizi-Derakhshi, and Aynaz Forouzandeh. Named entities detection by beam search algorithm. Concurrency and Computation: Practice and Experience, 34(27):e7325:1–e7325:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ghiglio:2023:IPS

[GDGN23] Pietro Ghiglio, Uwe Dolinsky, Mehdi Goli, and Kumudha Narasimhan. Improving performance of SYCL applications on CPU architectures using LLVM-directed compilation flow. Concurrency and Computation: Practice and Experience, 35(27):e7810:1–e7810:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gicic:2023:ICS

[GDS23] Adaleta Gicić, Dzenana Donko, and Abdulhamit Subasi. Intelligent credit scoring using deep learning methods. Concurrency and Computation: Practice and Experience, 35(9): e7637:1–e7637:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

G:2022:CII

[GDSS22] Srinitya G, Sharmila D, Logeswari S, and Daniel Madan Raja S. Certain investigations on image segmentation algorithms on synthetic aperture radar images and classification using convolution neural network. Concurrency and Computation: Practice and Experience, 34(10):e6733:1–e6733:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gharehchopogh:2021:MFF

[GFA21] Farhad Soleimanian Gharehchopogh, Behnam Farnad, and Ali Alizadeh. A modified farmland fertility algorithm for solving constrained engineering problems. Concurrency and Computation: Practice and Experience, 33(17):e6310:1–e6310:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Giannoutakis:2021:ESO

[GFPGT21] Konstantinos M. Giannoutakis, Christos K. Filelis-Papadopoulos, George A. Gravvanis, and Dimitrios Tzovaras. Evaluation of self-organizing and self-managing heterogeneous high performance computing clouds through discrete-time simulation. Concurrency and Computation: Practice and Experience, 33(17):e6326:1–e6326:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2020:ECP

[GFQ20] Gaimei Gao, Hongxia Fei, and Zefeng Qin. An efficient certificateless public auditing scheme in cloud storage. Concurrency and Computation: Practice and Experience, 32(24):e5924:1– e5924:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gokgoz:2024:OMB

[GGA24] Baki Gökgöz, Fatih Gül, and Tolga Aydin. An overview memristor based hardware accelerators for deep neural network. Concurrency and Computation: Practice and Experience, 36 (9):e7997:1–e7997:??, April 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gaioso:2020:PES

[GGCGS20] Roussian Gaioso, Veronica Gil-Costa, Helio Guardia, and Hermes Senger. Performance evaluation of single vs. batch of queries on GPUs. Concurrency and Computation: Practice and Experience, 32(20):e5474:1–e5474:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guo:2024:IBS

[GGJ^{+}24] Jinglei Guo, Chen Gao, Shouyong Jiang, Wei Xie, and Zhijian Wu. An improved brain storm optimization with chunkinggrouping method. Concurrency and Computation: Practice and Experience, 36(4):e7924:1–e7924:??, February 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Goyal:2022:PDE

[GGS^{+}22] Pankhuri Goyal, Nupur Goyal, Pavneet Singh, Nischay Mittal, Neeru Jindal, and Kanwarpreet Kaur. Pharmaceutical drugs expiry date tracking: a visionary approach. Concurrency and Computation: Practice and Experience, 34(28): e7358:1–e7358:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ghasemi:2020:RCV

[Gha20] Reza Ghasemi. Resolving a common vulnerability in secret sharing scheme-based data outsourcing schemes. Concurrency and Computation: Practice and Experience, 32(2):e5363:1– e5363:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Grant:2023:EPM

[GHL^{+}23] Ryan E. Grant, Simon D. Hammond, James H. Laros III, Michael Levenhagen, Stephen L. Olivier, Kevin Pedretti, Lee Ward, and Andrew J. Younge. Enabling power measurement and control on Astra: The first petascale Arm supercomputer. Concurrency and Computation: Practice and Experience, 35 (15):e7303:1–e7303:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ghourabi:2021:SDS

[Gho21] Abdallah Ghourabi. SM-Detector: a security model based on BERT to detect SMiShing messages in mobile environments. Concurrency and Computation: Practice and Experience, 33(24):e6452:1–e6452:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ghafouri:2021:WSQ

[GHRM21] Seyyed Hamid Ghafouri, Seyyed Mohsen Hashemi, Mohammad Reza Razzazi, and Ali Movaghar. Web service quality of service prediction via regional reputation-based matrix factorization. Concurrency and Computation: Practice and Experience, 33(17):e6318:1–e6318:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guo:2020:HDS

[GHT20] Wenxia Guo, Chaojie Huang, and Wenhong Tian. Handling data skew at reduce stage in Spark by ReducePartition. Concurrency and Computation: Practice and Experience, 32(9): e5637:1–e5637:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ghrab:2022:CAR

[GJBM22] Dhouha Ghrab, Imen Jemili, Abdelfettah Belghith, and Mohamed Mosbah. Context-aware routing framework for dutycycled wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(7):e5958:1–e5958:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gilesh:2020:OLM

[GJK^{+}20] Malayam Parambath Gilesh, Subham Jain, S. D. Madhu Kumar, Lillykutty Jacob, and Umesh Bellur. Opportunistic live migration of virtual machines. Concurrency and Computation: Practice and Experience, 32(5):e5477:1–e5477:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ganeshan:2023:SIK

[GJPS23] Arulkumaran Ganeshan, Santhosh Jayagopalan, Balamurugan Perumal, and Velliangiri Sarveshwaran. Secure identity key and blockchain-based authentication approach for secure data communication in multi-WSN. Concurrency and Computation: Practice and Experience, 35(28):e7861:1–e7861:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2022:DEM

[GK22a] Aakansha Gupta and Rahul Katarya. Deep embedding for mental health content on social media using vector space model with feature clusters. Concurrency and Computation: Practice and Experience, 34(13):e6930:1–e6930:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2022:MBA

[GK22b] Shweta Gupta and Vibhor Kant. A model-based approach to user preference discovery in multi-criteria recommender system using genetic programming. Concurrency and Computation: Practice and Experience, 34(11):e6899:1–e6899:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gorikapudi:2023:NCM

[GK23a] Sateesh Gorikapudi and Hari Kishan Kondaveeti. A novel clustering model via optimized fuzzy C-means algorithm and sandpiper optimization with cycle crossover process in IoT. Concurrency and Computation: Practice and Experience, 35 (23):e7776:1–e7776:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2023:NAA

[GK23b] Garima Gupta and Rahul Katarya. A novel approach to alleviate data sparsity and generate dynamic fruit recommendations from point-of-sale data. Concurrency and Computation: Practice and Experience, 35(1):e7423:1–e7423:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guzel:2023:CSP

[GK23c] Ismail Güzel and Atabey Kaygun. Classification of stochastic processes with topological data analysis. Concurrency and Computation: Practice and Experience, 35(24):e7732:1– e7732:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guzel:2020:ADL

[GKAO20] Metehan Guzel, Ibrahim Kok, Diyar Akay, and Suat Ozdemir. ANFIS and deep learning based missing sensor data prediction in IoT. Concurrency and Computation: Practice and Experience, 32(2):e5400:1–e5400:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Garip:2022:SPP

[GKÇ22] Zeynep Garip, Durmus Karayel, and Murat Erhan Çimen. A study on path planning optimization of mobile robots based on hybrid algorithm. Concurrency and Computation: Practice and Experience, 34(5):e6721:1–e6721:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Graja:2020:CSM

[GKG^{+}20] Imen Graja, Slim Kallel, Nawal Guermouche, Saoussen Cheikhrouhou, and Ahmed Hadj Kacem. A comprehensive survey on modeling of cyber-physical systems. Concurrency and Computation: Practice and Experience, 32(15):e4850:1– e4850:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gec:2023:MPS

[GKLS23] Sandi Gec, Petar Kochovski, Dejan Lavbic, and Vlado Stankovski. Multi-party smart contract for an AI services ecosystem: an application to smart construction. Concurrency and Computation: Practice and Experience, 35(18): e6895:1–e6895:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gangadhari:2022:ARS

[GKM22] Rajan Kumar Gangadhari, Vivek Khanzode, and Shankar Murthy. Application of rough set theory and machine learning algorithms in predicting accident outcomes in the Indian petroleum industry. Concurrency and Computation: Practice and Experience, 34(26):e7277:1–e7277:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gunes:2023:CML

[GKZ23] Oguzhan Mehmet Günes, Pelin Kasap, and Burcin Seyda Çorba Zorlu. The comparison of machine learning classification algorithms used to diagnose liver cirrhosis disease and a brief review. Concurrency and Computation: Practice and Experience, 35(8):e7628:1–e7628:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gulzar:2022:EER

[GLA^{+}22] Muhammad Awais Gulzar, Waqas Latif, Muhammad Abid, Hafiz Zafar Nazir, and Muhammad Riaz. On enhanced exponential-cum-ratio estimators using robust measures of location. Concurrency and Computation: Practice and Experience, 34(6):e6763:1–e6763:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guerrero:2020:OPF

[GLJ20] Carlos Guerrero, Isaac Lera, and Carlos Juiz. Optimization policy for file replica placement in fog domains. Concurrency and Computation: Practice and Experience, 32(21):e5343:1– e5343:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Giorgi:2022:EIT

[GLM^{+}22] Giacomo Giorgi, Antonio La Marra, Fabio Martinelli, Paolo Mori, Athanasios Rizos, and Andrea Saracino. Exploiting if this then that and usage control obligations for smart home security and management. Concurrency and Computation: Practice and Experience, 34(16):e6189:1–e6189:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gaouar:2023:CCB

[GLN23] Nihal Gaouar, Mohamed Lehsaini, and Tawfiq Nebbou. CC-ITL: a cloud-based smart traffic management protocol using intelligent traffic light system in VANETs. Concurrency and Computation: Practice and Experience, 35(12):e7686:1– e7686:??, May 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guimaraes:2021:HPI

[GLRB21] Antonio Guimarães, Luis Lacalle, Charles B. Rodamilans, and Edson Borin. High-performance IO for seismic processing on the cloud. Concurrency and Computation: Practice and Experience, 33(18):e6250:1–e6250:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ge:2023:EIE

[GlSL^{+}23] Jun Ge, Lei lei Shi, Lu Liu, Hongwei Shi, and John Panneerselvam. Edge intelligence-enabled dynamic overlapping community discovery and evolution prediction in social media data streams. Concurrency and Computation: Practice and Experience, 35(13):e6786:1–e6786:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2021:UFP

[GLW21] Jianqi Gao, Xiangfeng Luo, and Hao Wang. An uncertain future: Predicting events using conditional event evolutionary graph. Concurrency and Computation: Practice and Experience, 33(9):e6164:1–e6164:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2022:CCE

[GLW22] Jianqi Gao, Xiangfeng Luo, and Hao Wang. Chinese causal event extraction using causality-associated graph neural network. Concurrency and Computation: Practice and Experience, 34(3):e6572:1–e6572:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Graffi:2021:LPP

[GM21] Kalman Graffi and Newton Masinde. LibreSocial: a peerto-peer framework for online social networks. Concurrency and Computation: Practice and Experience, 33(8):e6150:1– e6150:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2022:NAM

[GM22a] Yuan Gao and Bayan Omar Mohammed. A new applicable and multilayer design of nanoscale adder-subtractor using quantum-dots. Concurrency and Computation: Practice and Experience, 34(21):e7050:1–e7050:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Govindaswamy:2022:DSM

[GM22b] Bharathi Govindaswamy and Malleswaran Mallappa. Detection and segmentation of melanoma skin cancer in dermoscopy images using modified Alexnet convolutional neural network-morphological methodology. Concurrency and Computation: Practice and Experience, 34(25):e7266:1–e7266:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2022:ASE

[GM22c] Neha Gupta and Rajendra Prasad Mahapatra. Automated software effort estimation for agile development system by heuristically improved hybrid learning. Concurrency and Computation: Practice and Experience, 34(25):e7267:1– e7267:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ghazimirsaeed:2020:CAM

[GMA20] S. Mahdieh Ghazimirsaeed, Seyed H. Mirsadeghi, and Ahmad Afsahi. Communication-aware message matching in MPI. Concurrency and Computation: Practice and Experience, 32 (3):e4862:1–e4862:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2021:PMH

[GMK^{+}21] Nishant Gupta, Nitish Mahajan, Sakshi Kaushal, Naresh Kumar, Harish Kumar, and Arun Kumar Sangaiah. Predictive model for hardware calibration to transmit real-time applications in VoIP networks. Concurrency and Computation: Practice and Experience, 33(4):e5190:1–e5190:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gangula:2022:NID

[GMM22] Rekha Gangula, Murali Mohan V, and Ranjeeth Kumar M. Network intrusion detection system for Internet of Things based on enhanced flower pollination algorithm and ensemble classifier. Concurrency and Computation: Practice and Experience, 34(21):e7103:1–e7103:??, September 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Goudarzi:2021:MFH

[GMN21] Farzaneh Khalili Goudarzi, Hamid Reza Maleki, and Sadegh Niroomand. Mathematical formulation and hybrid metaheuristic algorithms for multiproduct oil pipeline scheduling problem with tardiness penalties. Concurrency and Computation: Practice and Experience, 33(17):e6299:1–e6299:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guerar:2020:SPB

[GMP^{+}20] Meriem Guerar, Mauro Migliardi, Francesco Palmieri, Luca Verderame, and Alessio Merlo. Securing PIN-based authentication in smartwatches with just two gestures. Concurrency and Computation: Practice and Experience, 32(18):e5549:1– e5549:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Girelli:2021:IMP

[GMS^{+}21] Valéria S. Girelli, Francis B. Moreira, Matheus S. Serpa, Danilo Carastan-Santos, and Philippe O. A. Navaux. Investigating memory prefetcher performance over parallel applications: From real to simulated. Concurrency and Computation: Practice and Experience, 33(18):e6207:1–e6207:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gomes:2021:MAA

[GMSM21] Antônio Tadeu A. Gomes, Enzo Molion, Roberto P. Souto, and Jean-François Méhaut. Memory allocation anomalies in high-performance computing applications: a study with numerical simulations. Concurrency and Computation: Practice and Experience, 33(18):e6094:1–e6094:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

George:2021:IMP

[GN21a] Anju Susan George and A. Shajin Nargunam. Improved multi-party verification protocol with reduced computational overhead in cloud storage system. Concurrency and Computation: Practice and Experience, 33(21):e6427:1–e6427:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2021:QCM

[GN21b] Mayank Gupta and Manisha J. Nene. Quantum computing: a measurement and analysis review. Concurrency and Computation: Practice and Experience, 33(20):e6344:1–e6344:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guillen-Navarro:2021:HPI

[GNMELC21] Miguel A. Guillén-Navarro, Raquel Martínez-España, Belén López, and José M. Cecilia. A high-performance IoT solution to reduce frost damages in stone fruits. Concurrency and Computation: Practice and Experience, 33(2):e5299:1– e5299:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gunasekar:2022:MOM

[GNS22] Manikandan Gunasekar, Gobalakrishnan Natesan, and Duraimurugan Samiayya. Multi-objective moth swarm based sailfish model for optimal routing in wireless sensor network. Concurrency and Computation: Practice and Experience, 34(21):e7125:1–e7125:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guermouche:2022:TDP

[GO22] Amina Guermouche and Anne-Cécile Orgerie. Thermal design power and vectorized instructions behavior. Concurrency and Computation: Practice and Experience, 34(2):e6261:1– e6261:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gohil:2022:LBC

[GP22a] Bhavesh N. Gohil and Dhiren R. Patel. Load balancing in cloud using improved gray wolf optimizer. Concurrency and Computation: Practice and Experience, 34(11):e6888:1– e6888:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gopatoti:2022:MTF

[GP22b] Anandbabu Gopatoti and Vijayalakshmi P. Multi-texture features and optimized DeepNet for COVID-19 detection using chest X-ray images. Concurrency and Computation: Practice and Experience, 34(22):e7157:1–e7157:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ganapathi:2020:UED

[GPDB20] Iyyakutti Iyappan Ganapathi, Surya Prakash, Ishan R. Dave, and Sambit Bakshi. Unconstrained ear detection using ensemble-based convolutional neural network model. Concurrency and Computation: Practice and Experience, 32(1): e5197:1–e5197:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gayathri:2022:CFB

[GPR^{+}22] A. Gayathri, A. V. Prabu, S. Rajasoundaran, Sidheswar Routray, P. Narayanasamy, Naween Kumar, and Yinan Qi. Cooperative and feedback based authentic routing protocol for energy efficient IoT systems. Concurrency and Computation: Practice and Experience, 34(11):e6886:1–e6886:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2020:OIT

[GQ20] B. B. Gupta and Megha Quamara. An overview of Internet of Things (IoT): Architectural aspects, challenges, and protocols. Concurrency and Computation: Practice and Experience, 32(21):e4946:1–e4946:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2021:TVA

[GQ21] B. B. Gupta and Megha Quamara. A taxonomy of various attacks on smart card-based applications and countermeasures. Concurrency and Computation: Practice and Experience, 33 (7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gibbons:2024:RSA

[GQ24] Ethan Gibbons and Ke Qiu. A recurrence for the surface area of the (n, k)-star graph. Concurrency and Computation: Practice and Experience, 36(12):e8059:1–e8059:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2022:CGU

[GR22a] Yuan Gao and Muhammad Adnan Rafi. Combination of graphics, uncertainty, and semantics: a survey. Concurrency and Computation: Practice and Experience, 34(7):e6711:1– e6711:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gholami:2022:CMA

[GR22b] Hadi Gholami and Mohammad Taghi Rezvan. A cooperative multi-agent offline learning algorithm to scheduling IoT workflows in the cloud computing environment. Concurrency and Computation: Practice and Experience, 34(22):e7148:1– e7148:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Groves:2023:AAB

[GRC^{+}23] Taylor Groves, Naveen Ravichandrasekaran, Brandon Cook, Noel Keen, David Trebotich, Nicholas J. Wright, Bob Alverson, Duncan Roweth, and Keith Underwood. Not all applications have boring communication patterns: Profiling message matching with BMM. Concurrency and Computation: Practice and Experience, 35(15):e6380:1–e6380:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2022:NIX

[GRG^{+}22] Aditya Gupta, Ishwari Singh Rajput, Gunjan, Vibha Jain, and Soni Chaurasia. NSGA-II-XGB: Meta-heuristic feature selection with XGBoost framework for diabetes prediction. Concurrency and Computation: Practice and Experience, 34(21):e7123:1–e7123:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ge:2022:BEG

[GRL^{+}22] Keshi Ge, Zhejiang Ran, Zhiquan Lai, Lizhi Zhang, and Dongsheng Li. BRGraph: an efficient graph neural network training system by reusing batch data on GPU. Concurrency and Computation: Practice and Experience, 34(15):e6961:1– e6961:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Governor:2023:NDD

[GRMP23] Kalaiarasi Governor, Padmavathy Ramanujam, Suja Cherukullapurath Mana, and Geetha Perumal. Near duplicate detection of images with area and proposed pixel-based feature extraction. Concurrency and Computation: Practice and Experience, 35(2):e7477:1–e7477:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gillan:2021:CIE

[GS21] Charles J. Gillan and Ivor Spence. Computing integrals for electron molecule scattering on heterogeneous accelerator systems. Concurrency and Computation: Practice and Experience, 33(5):e5984:1–e5984:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ganasigamony:2022:CAD

[GS22] Wiselin Jiji Ganasigamony and Muthuraj Antony Arul Selvaraj. Computer assisted diagnosis of bipolar disorder using invariant features. Concurrency and Computation: Practice and Experience, 34(17):e6984:1–e6984:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gulhane:2021:ICM

[GSB21] Sushen Rameshpant Gulhane, Suresh Damodar Shirbahadurkar, and Sanjay Shrikrushna Badhe. Indian classical musical instrument classification using timbral features. Concurrency and Computation: Practice and Experience, 33(21): e6418:1–e6418:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Garcia:2020:PNP

[GSG20] Adriano Marques Garcia, Claudio Schepke, and Alessandro Girardi. PAMPAR: a new parallel benchmark for performance and energy consumption evaluation. Concurrency and Computation: Practice and Experience, 32(20):e5504:1–e5504:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gola:2023:MOH

[GSG^{+}23] Kamal Kumar Gola, Brij Mohan Singh, Bhumika Gupta, Nishant Chaurasia, and Shikha Arya. Multi-objective hybrid capuchin search with genetic algorithm based hierarchical resource allocation scheme with clustering model in cloud computing environment. Concurrency and Computation: Practice and Experience, 35(7):e7606:1–e7606:??, March 25, 2023.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gangathimmappa:2023:DLE

[GSS^{+}23a] Mahesh Gangathimmappa, Neelakandan Subramani, Velmurugan Sambath, Rengaraj Alias Muralidharan Ramanujam, Naresh Sammeta, and Maheswari Marimuthu. Deep learning enabled cross-lingual search with metaheuristic web based query optimization model for multi-document summarization. Concurrency and Computation: Practice and Experience, 35 (2):e7476:1–e7476:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Govardhan:2023:AIB

[GSS23b] Praveena Nalamani Govardhan, Srigitha Surendranath, and Manju Sundararajan. Artificial intelligence-based remote monitoring system for automated anomaly detection in wireless sensor networks. Concurrency and Computation: Practice and Experience, 35(2):e7462:1–e7462:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Giveki:2022:NMI

[GSTS22] Davar Giveki, Ashkan Shakarami, Hadis Tarrah, and Mohammad Ali Soltanshahi. A new method for image classification and image retrieval using convolutional neural networks. Concurrency and Computation: Practice and Experience, 34 (1):e6533:1–e6533:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gupta:2021:UDW

[GSVS21] Swati Gupta, Ravi S. Singh, Umare D. Vasant, and Vijit Saxena. User defined weight based budget and deadline constrained workflow scheduling in cloud. Concurrency and Computation: Practice and Experience, 33(24):e6454:1–e6454:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gerhards:2020:NAA

[GSZ^{+}20] Michael Gerhards, Volker Sander, Miroslav Zivković, Adam Belloum, and Marian Bubak. New approach to allocation planning of many-task workflows on clouds. Concurrency and Computation: Practice and Experience, 32(2):e5404:1– e5404:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gonzalez-Tallada:2024:CUO

[GTM24] Marc Gonzàlez-Tallada and Enric Morancho. Compute units in OpenMP: Extensions for heterogeneous parallel programming. Concurrency and Computation: Practice and Experience, 36(1):e7885:1–e7885:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gurbuz:2023:RAC

[GU23] Tugba Gürbüz and Çelebi Uluyol. Research article classification with text mining method. Concurrency and Computation: Practice and Experience, 35(1):e7437:1–e7437:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gui:2022:CDI

[Gui22] Dawei Gui. The computer desktop image compression based on clustering algorithm. Concurrency and Computation: Practice and Experience, 34(13):e5892:1–e5892:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gul:2022:BRC

[Gul22] Ertugrul Gul. A blind robust color image watermarking method based on discrete wavelet transform and discrete cosine transform using grayscale watermark image. Concurrency and Computation: Practice and Experience, 34(22):e6884:1– e6884:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guney:2023:ANA

[Gün23a] Hüseyin Güney. AWC-NIDS: Attack-wise customized network intrusion detection system using machine learning, concurrency, and distributed systems. Concurrency and Computation: Practice and Experience, 35(26):e7832:1–e7832:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guney:2023:FSI

[Gün23b] Hüseyin Güney. Feature selection-integrated classifier optimisation algorithm for network intrusion detection. Concurrency and Computation: Practice and Experience, 35(23):e7807:1– e7807:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Govindaraju:2022:IDF

[GVSS22] Shanthi Govindaraju, Wilson Vimala Rani Vinisha, Francis H. Shajin, and D. Adhimuga Sivasakthi. Intrusion detection framework using auto-metric graph neural network optimized with hybrid woodpecker mating and capuchin search optimization algorithm in IoT network. Concurrency and Computation: Practice and Experience, 34(24):e7197:1– e7197:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gu:2023:OSA

[GWA^{+}23] Jiayan Gu, Yan Wu, Ashiq Anjum, John Panneerselvam, Yao Lu, and Bo Yuan. Optimization of service addition in multilevel index model for edge computing. Concurrency and Computation: Practice and Experience, 35(13):e6626:1–e6626:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gu:2024:AKS

[GWA^{+}24] Jiayan Gu, Yan Wu, Ashiq Anjum, Lu Liu, John Panneerselvam, and Yao Lu. An adaptive key selection method for the multilevel index model for effective service management in the cloud. Concurrency and Computation: Practice and Experience, 36(10):e7768:1–e7768:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2020:MNG

[GWGR20] Ziyan Gao, Yong Wang, Yifan Gao, and Xingtian Ren. Multiobjective noncooperative game model for cost-based task scheduling in cloud computing. Concurrency and Computation: Practice and Experience, 32(7):e5570:1–e5570:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guan:2021:PAP

[GXH^{+}21] Zheng Guan, Lei Xiong, Min He, Wen-Hua Qian, and Zhaoxu Zhou. Performance analysis of polling-based MAC protocol with retrial for Internet of Things. Concurrency and Computation: Practice and Experience, 33(4):e5544:1–e5544:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2020:IMB

[GXL^{+}20] Min Gao, Li Xu, Limei Lin, Yanze Huang, and Xinxin Zhang. Influence maximization based on activity degree in mobile social networks. Concurrency and Computation: Practice and Experience, 32(11):e5677:1–e5677:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guo:2021:TMM

[GYL^{+}21] Liangmin Guo, Hao Yang, Kaixuan Luan, Yonglong Luo, Liping Sun, and Xiaoyao Zheng. A trust management model based on mutual trust and a reward-with-punishment mechanism for cloud environments. Concurrency and Computation: Practice and Experience, 33(16):e6283:1–e6283:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gong:2020:ITM

[GYZ^{+}20] Changqing Gong, Delong Yu, Liang Zhao, Xiguang Li, and Xianwei Li. An intelligent trust model for hybrid DDoS detection in software defined networks. Concurrency and Computation: Practice and Experience, 32(16):e5264:1–e5264:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gan:2020:EVO

[GZ20] Haiqing Gan and Chuiyong Zheng. An electric vehicle operation optimization method based on demand-side management. Concurrency and Computation: Practice and Experience, 32(23):e5532:1–e5532:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gong:2021:TGS

[GZC^{+}21] Bei Gong, Xinyi Zhang, Yang Cao, Zheng Li, Jia Yang, and Wei Wang. A threshold group signature scheme suitable for the Internet of Things. Concurrency and Computation: Practice and Experience, 33(13):e6243:1–e6243:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2022:CER

[GZC^{+}22] Jie Gao, Zuping Zhang, Ping Cao, Wei Huang, and Fangfang Li. Citation entity recognition method using multi-feature semantic fusion based on deep learning. Concurrency and Computation: Practice and Experience, 34(6):e6770:1–e6770:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2020:RBS

[gZWfL^{+}20] Hong gen Zhou, Zhen Wang, Jin feng Liu, Gui zhong Tian, Xu wen Jing, and Lei Li. Research on block storage yard scheduling according to grid method. Concurrency and Computation: Practice and Experience, 32(23):e5576:1–e5576:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Guo:2022:NNQ

[GZY^{+}22] Liangmin Guo, Ying Zhu, Hao Yang, Yonglong Luo, Liping Sun, and Xiaoyao Zheng. A k-nearest neighbor query method based on trust and location privacy protection. Concurrency and Computation: Practice and Experience, 34(16):e5766:1– e5766:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hong:2021:VDP

[HA21] Minsung Hong and Rajendra Akerkar. Victim detection platform in IoT paradigm. Concurrency and Computation: Practice and Experience, 33(3):e5254:1–e5254:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hadaya:2022:PRR

[HA22] Nariman Najeeb Hadaya and Salah Abdulghani Alabady. Proposed RPL routing protocol in the IoT applications. Concurrency and Computation: Practice and Experience, 34(10): e6805:1–e6805:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haider:2021:IAA

[HAA^{+}21] Sajjad Haider, Muhammad Tanvir Afzal, Muhammad Asif, Hermann Maurer, Awais Ahmad, and Abdelrahman Abuarqoub. Impact analysis of adverbs for sentiment classification on Twitter product reviews. Concurrency and Computation: Practice and Experience, 33(4):e4956:1–e4956:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hyder:2023:TDI

[HAA23] Muhammad Faraz Hyder, Waqas Ahmed, and Maaz Ahmed. Toward deceiving the intrusion attacks in containerized cloud environment using virtual private cloud-based moving target defense. Concurrency and Computation: Practice and Experience, 35(5):e7549:1–e7549:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hyder:2022:PPM

[HAAF22] Muhammad Faraz Hyder, Saadia Arshad, Asad Arfeen, and Tasbiha Fatima. Privacy preserving mobile forensic framework using role-based access control and cryptography. Concurrency and Computation: Practice and Experience, 34(23): e7178:1–e7178:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hajiali:2020:BDS

[Haj20] Mahdi Hajiali. Big data and sentiment analysis: a comprehensive and systematic literature review. Concurrency and Computation: Practice and Experience, 32(14):e5671:1–e5671:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Henry:2022:HMH

[HAK22] Niroshini Infantia Henry, C Anbuananth, and S Kalarani. Hybrid meta-heuristic algorithm for optimal virtual machine placement and migration in cloud computing. Concurrency and Computation: Practice and Experience, 34(28):e7353:1– e7353:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hasan:2020:TSI

[HAR20] Mohammed Zaki Hasan and Hussain Al-Rizzo. Task scheduling in Internet of Things cloud environment using a robust particle swarm optimization. Concurrency and Computation: Practice and Experience, 32(2):e5442:1–e5442:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hemmati:2024:TSB

[HAR24] Atefeh Hemmati, Hanieh Mohammadi Arzanagh, and Amir Masoud Rahmani. A taxonomy and survey of big data in social media. Concurrency and Computation: Practice and Experience, 36(1):e7875:1–e7875:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haidri:2022:DAL

[HAS^{+}22] Raza A. Haidri, Mahfooz Alam, Mohammad Shahid, Shiv Prakash, and Mohammad Sajid. A deadline aware load balancing strategy for cloud computing. Concurrency and Computation: Practice and Experience, 34(1):e6496:1–e6496:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ho:2020:DSN

[HBB20] Thi Kim Thoa Ho, Quang Vu Bui, and Marc Bui. Dynamic social network analysis: a novel approach using agent-based model, author-topic model, and pretopology. Concurrency and Computation: Practice and Experience, 32(13):e5321:1– e5321:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hemavathy:2022:PVC

[HC22] P. Hemavathy and M. Chinnadurai. Performance validation of clustering algorithms using selection of attributes and application of filters in terms of data reduction. Concurrency and Computation: Practice and Experience, 34(8):e5364:1– e5364:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2021:NDS

[HCG21] Guixia He, Qi Chen, and Jiaquan Gao. A new diagonal storage for efficient implementation of sparse matrix-vector multiplication on graphics processing unit. Concurrency and Computation: Practice and Experience, 33(13):e6230:1–e6230:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2021:PTA

[HCH^{+}21] Jing He, Jinjun Chen, Guangyan Huang, Jie Cao, Zhiwang Zhang, Hui Zheng, Peng Zhang, Roozbeh Zarei, Ferry Sansoto, Ruchuan Wang, Yimu Ji, Weibei Fan, Zhijun Xie, Xiancheng Wang, Mengjiao Guo, Chi-Hung Chi, Paulo A. de Souza, Jiekui Zhang, Youtao Li, Xiaojun Chen, Yong Shi, David Green, Taraporewalla Kersi, and André Van Zundert. A polynomial-time algorithm for simple undirected graph isomorphism. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hira:2023:AHB

[HD23] Swati Hira and Parag S. Deshpande. Automated heuristic based context dependent ETL process to generate multidimensional model for tabular data. Concurrency and Computation: Practice and Experience, 35(2):e7459:1–e7459:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2023:DPL

[HDS^{+}23] Hongsheng Hu, Gillian Dobbie, Zoran Salcic, Meng Liu, Jianbing Zhang, Lingjuan Lyu, and Xuyun Zhang. Differentially private locality sensitive hashing based federated recommender system. Concurrency and Computation: Practice and Experience, 35(14):e6233:1–e6233:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2020:IPL

[HDXH20] Guang-Li Huang, Ke Deng, Zhijun Xie, and Jing He. Intelligent pseudo-location recommendation for protecting personal location privacy. Concurrency and Computation: Practice and Experience, 32(2):e5435:1–e5435:??, January 25, 2020.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hormozi:2023:SBD

[HE23] Mohammad Hormozi and S. Hossein Erfani. An SDN-based DDoS defense approach using route obfuscation. Concurrency and Computation: Practice and Experience, 35(1):e7439:1– e7439:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hemalatha:2022:OAK

[Hem22] M. Hemalatha. Optimal AdaBoost kernel support vector machine for monitoring arrhythmia patients utilizing Internet of Things-cloud environment. Concurrency and Computation: Practice and Experience, 34(27):e7298:1–e7298:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hodon:2020:EIN

[HFFA20] Michal Hodon, Janusz Furtak, Günter Fahrnberger, and Ali I. Awad. Editorial on innovative network systems and applications together with the conference on information systems innovations for community services. Concurrency and Computation: Practice and Experience, 32(13):e5698:1–e5698:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2021:DAS

[HFH^{+}21] Li Huang, Qiaobo Fu, Meiling He, Du Jiang, and Zhiqiang Hao. Detection algorithm of safety helmet wearing based on deep learning. Concurrency and Computation: Practice and Experience, 33(13):e6234:1–e6234:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Herskind:2020:BER

[HGDD20] Lasse Herskind, Alberto Giaretta, Michele De Donno, and Nicola Dragoni. BitFlow: Enabling real-time cash-flow evaluations through blockchain. Concurrency and Computation: Practice and Experience, 32(12):e5333:1–e5333:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hashemi:2022:NMT

[HGHD22] Seyyed Mohsen Hashemi, Seyyed Hamid Ghafouri, Patrick C. K. Hung, and Chen Ding. A new model for trustworthy web service QoS prediction. Concurrency and Computation: Practice and Experience, 34(6):e6778:1–e6778:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Helbecque:2023:PDP

[HGM^{+}23] Guillaume Helbecque, Jan Gmys, Nouredine Melab, Tiago Carneiro, and Pascal Bouvry. Parallel distributed productivityaware tree-search using Chapel. Concurrency and Computation: Practice and Experience, 35(27):e7874:1–e7874:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hosseinalipour:2021:TTP

[HGMK21] Ali Hosseinalipour, Farhad Soleimanian Gharehchopogh, Mohammad Masdari, and Ali Khademi. Toward text psychology analysis using social spider optimization algorithm. Concurrency and Computation: Practice and Experience, 33(17): e6325:1–e6325:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Heidari:2022:DSI

[HGNN22] Marzi Heidari, Mehdi Ghatee, Ahmad Nickabadi, and Arash Pourhasan Nezhad. Diverse and styled image captioning using singular value decomposition-based mixture of recurrent experts. Concurrency and Computation: Practice and Experience, 34(22):e6866:1–e6866:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Han:2023:ODC

[HGW^{+}23] Xinxin Han, Guichen Gao, Yang Wang, Hing-Fung Ting, Ilsun You, and Yong Zhang. Online data caching in edge computing. Concurrency and Computation: Practice and Experience, 35(17):e6468:1–e6468:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haghi:2022:RSH

[HGX^{+}22] Pouya Haghi, Anqi Guo, Qingqing Xiong, Chen Yang, Tong Geng, Justin T. Broaddus, Ryan Marshall, Derek Schafer, Anthony Skjellum, and Martin C. Herbordt. Reconfigurable switches for high performance and flexible MPI collectives. Concurrency and Computation: Practice and Experience, 34 (6):e6769:1–e6769:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Han:2023:ARW

[HH23] Nguyen Han and Nguyen Cong Hao. Algorithm for reasoning with words based on linguistic fuzzy cognitive map. Concurrency and Computation: Practice and Experience, 35(15): e6415:1–e6415:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2022:PTA

[HHC^{+}22] Jing He, Guangyan Huang, Jie Cao, Zhiwang Zhang, Hui Zheng, Peng Zhang, Roozbeh Zarei, Ferry Susanto, Ruchuan Wang, Yimu Ji, Weibei Fan, Zhijun Xie, Xiancheng Wang, Mengjiao Guo, Chi-Hung Chi, Jiekui Zhang, Youtao Li, Xiaojun Chen, Yong Shi, and Andre Van Zundert. A polynomialtime algorithm for simple undirected graph isomorphism. Concurrency and Computation: Practice and Experience, 34 (3):e6599:1–e6599:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2023:TCE

[HHP23] Guiqiu Hu, Xiji Hu, and Somayeh Pouramini. Thermal comfort and energy performance improvement by optimization of shading devices using improved Battle Royal algorithm. Concurrency and Computation: Practice and Experience, 35(9): e7646:1–e7646:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2020:CMB

[HHXH20] Guang-Li Huang, Jing He, Zenglin Xu, and Guangyan Huang. A combination model based on transfer learning for waste classification. Concurrency and Computation: Practice and Experience, 32(19):e5751:1–e5751:??, October 10, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2022:VRC

[HHYL22] Baohua Huang, Pirong Huang, Hong Yuan, and Sheng Liang. A verifiable ranked ciphertext retrieval scheme based on bilinear mapping. Concurrency and Computation: Practice and Experience, 34(12):e5829:1–e5829:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hamed:2022:SIP

[HIEH22] Belal A. Hamed, Osman Ali Sadek Ibrahim, and Tarek Abd El-Hafeez. A survey on improving pattern matching algorithms for biological sequences. Concurrency and Computation: Practice and Experience, 34(26):e7292:1–e7292:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2023:GPU

[HIN23] Jiamian Huang, Yasuaki Ito, and Koji Nakano. Graphics processing unit-accelerated high-quality watercolor painting image generation. Concurrency and Computation: Practice and Experience, 35(19):e7471:1–e7471:??, August 30, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Han:2020:CSS

[HJT^{+}20] Tao Han, Syed Rooh Ullah Jan, Zhiyuan Tan, Muhammad Usman, Mian Ahmad Jan, Rahim Khan, and Yongzhao Xu. A comprehensive survey of security threats and their mitigation techniques for next-generation SDN controllers. Concurrency and Computation: Practice and Experience, 32(16):e5300:1– e5300:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2022:PPA

[HJZ^{+}22] Xiangyu Hu, Zhiping Jin, Lefeng Zhang, Andi Zhou, and Dayong Ye. Privacy preservation auction in a dynamic social network. Concurrency and Computation: Practice and Experience, 34(16):e6058:1–e6058:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hariharan:2021:ARR

[HK21] I. Hariharan and M. Kannan. Algorithms for reducing reconfiguration overheads using prefetch, reuse, and optimal mapping of tasks. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Habib:2022:PEM

[HK22a] Beenish Habib and Farida Khursheed. Performance evaluation of machine learning models for distributed denial of service attack detection using improved feature selection and hyper-parameter optimization techniques. Concurrency and Computation: Practice and Experience, 34(26):e7299:1– e7299:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hema:2022:ASH

[HK22b] Sri Vigna Hema V and Ramesh K. Augmented security for healthcare data using obfuscation and elliptic curve cryptography algorithm in health cloud environment. Concurrency and Computation: Practice and Experience, 34(26):e7275:1– e7275:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Habib:2024:TBD

[HK24] Beenish Habib and Farida Khursheed. Time-based DDoS attack detection through hybrid LSTM-CNN model architectures: an investigation of many-to-one and many-to-many approaches. Concurrency and Computation: Practice and Experience, 36(9):e7996:1–e7996:??, April 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Handa:2020:EPP

[HKA20] Rohit Handa, C. Rama Krishna, and Naveen Aggarwal. Efficient privacy-preserving scheme supporting disjunctive multikeyword search with ranking. Concurrency and Computation: Practice and Experience, 32(2):e5450:1–e5450:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hong:2021:CIE

[HKMS21] Jong Youl Hong, Hoon Ko, Libor Mesicek, and MoonBae Song. Cultural intelligence as education contents: Exploring the pedagogical aspects of effective functioning in higher education. Concurrency and Computation: Practice and Experience, 33(2):e5489:1–e5489:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hwang:2021:MDA

[HKP21] Syehyun Hwang, Kyeong-Ri Ko, and Sung Bum Pan. Motion data acquisition method for motion analysis in golf. Concurrency and Computation: Practice and Experience, 33 (2):e5215:1–e5215:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2020:MPD

[HL20] Min Hu and Yunru Liu. E-maintenance platform design for public infrastructure maintenance based on IFC ontology and Semantic Web services. Concurrency and Computation: Practice and Experience, 32(6):e5204:1–e5204:??, March 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2023:DCM

[HL23] Ying Huang and Wei Li. Dynamic cognitive maps for robot route planning in complex workplaces represented by abstract mazes. Concurrency and Computation: Practice and Experience, 35(1):e7426:1–e7426:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2021:OAB

[HLC^{+}21] Zhenjie Huang, Zhiwei Lin, Qunshan Chen, Yuping Zhou, and Hui Huang. Outsourced attribute-based signatures with perfect privacy for circuits in cloud computing. Concurrency and Computation: Practice and Experience, 33(10):e6173:1– e6173:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hao:2023:NFF

[HLC23] Run Fang Hao, Chuang Liu, and Yong Qiang Cheng. New fusion features convolutional neural network with high generalization ability on rolling bearing fault diagnosis. Concurrency and Computation: Practice and Experience, 35(13):e7600:1– e7600:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2020:ASC

[HLCH20] Yao-Chieh Hu, Ting-Ting Lee, Dimitris Chatzopoulos, and Pan Hui. Analyzing smart contract interactions and contract level state consensus. Concurrency and Computation: Practice and Experience, 32(12):e5228:1–e5228:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huynh:2023:MPD

[HLDT^{+}23] Hiep Xuan Huynh, Be Ut Lai, Nghia Duong-Trung, Hai Thanh Nguyen, and Thuong-Cang Phan. Modeling population dynamics for information dissemination through Facebook. Concurrency and Computation: Practice and Experience, 35(15):e6333:1–e6333:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2024:MTD

[HLG24] Qiguo Huang, Zhengliang Li, and Qing Gu. Multi-task deep neural networks for just-in-time software defect prediction on mobile apps. Concurrency and Computation: Practice and Experience, 36(10):e7664:1–e7664:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2020:ITR

[HLH^{+}20] Subin Huang, Xiangfeng Luo, Jing Huang, Hao Wang, Shengwei Gu, and Yike Guo. Improving taxonomic relation learning via incorporating relation descriptions into word embeddings. Concurrency and Computation: Practice and Experience, 32 (14):e5696:1–e5696:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2021:JOE

[HLL^{+}21] Yihao He, Zebin Lu, Junru Lei, Shuhua Deng, and Xieping Gao. Joint optimization of energy saving and load balancing for data center networks based on software defined networks. Concurrency and Computation: Practice and Experience, 33(9):e6134:1–e6134:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huynh:2021:SSB

[HLO^{+}21] Hiep Xuan Huynh, Quy Thanh Lu, Linh My Thi Ong, Huong Hoang Luong, and Lan Phuong Phan. Simulating the spreading of brown plant hoppers based on cellular automata. Concurrency and Computation: Practice and Experience, 33 (2):e5261:1–e5261:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2020:SEW

[HLS^{+}20] Yi-Li Huang, Fang-Yie Leu, Ruey-Kai Sheu, Jung-Chun Liu, and Heru Susanto. A secure and efficient WSN by employing symmetric key matrix and rectangular frame scheme. Concurrency and Computation: Practice and Experience, 32(18): e5568:1–e5568:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hong:2023:NTD

[HLT23] Yu Hong, Yi Li, and Haining Tan. A noninterference trusted dual system security guarantee method based on secure memory. Concurrency and Computation: Practice and Experience, 35(2):e7463:1–e7463:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2021:FSI

[HLW^{+}21] Jiayan Huang, Zuoyong Li, Chuansheng Wang, Zhaochai Yu, and Xinrong Cao. FFNet: a simple image dedusting network with feature fusion. Concurrency and Computation: Practice and Experience, 33(24):e6462:1–e6462:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2023:MSA

[HLZ23] Haize Hu, Jianxun Liu, and Xiangping Zhang. Multilayer self-attention residual network for code search. Concurrency and Computation: Practice and Experience, 35(9):e7650:1– e7650:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hedhli:2020:DBA

[HM20] Ameni Hedhli and Haithem Mezni. A DFA-based approach for the deployment of BPaaS fragments in the cloud. Concurrency and Computation: Practice and Experience, 32(14): e5075:1–e5075:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Harichane:2022:KRS

[HMB22] Ishak Harichane, Sid Ahmed Makhlouf, and Ghalem Belalem. KubeSC-RTP: Smart scheduler for Kubernetes platform on CPU–GPU heterogeneous systems. Concurrency and Computation: Practice and Experience, 34(21):e7108:1–e7108:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hamrouni:2023:RDR

[HMK23] Tarek Hamrouni, Riad Mokadem, and Amel Khelifa. Review on data replication strategies in single vs. interconnected cloud systems: Focus on data correlation-aware strategies. Concurrency and Computation: Practice and Experience, 35 (22):e7758:1–e7758:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hamead:2024:SWP

[HMK24] Haja Moinudeen Shahul Hamead, Thanga Nadar Thangathai Mirnalinee, and Dhandapani Kavipriya. Selective wakeup for parallel forward engines under variable load towards energy conservation in IP routers. Concurrency and Computation: Practice and Experience, 36(1):e7881:1–e7881:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2019:SMR

[HML19] Jiewu Huang, Lie Ma, and Rong Li. Study on measurement reliability based on Liu estimator. Concurrency and Computation: Practice and Experience, 31(12):e4709:1–e4709:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See retraction notice [FPHZ20].

Huang:2020:REP

[HML20] Jiewu Huang, Lie Ma, and Rong Li. Retracted: Evaluation of the predictive performance of the principal component two-parameter estimator. Concurrency and Computation: Practice and Experience, 31(12):e5562:1–e5562:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See [LLH19].

Hadded:2021:GTB

[HML21] Mohamed Hadded, Pascale Minet, and Jean-Marc Lasgouttes. A game theory-based route planning approach for automated vehicle collection. Concurrency and Computation: Practice and Experience, 33(16):e6246:1–e6246:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Heavner:2023:CRR

[HMQO23] N. Heavner, P. G. Martinsson, and G. Quintana-Ortí. Computing rank-revealing factorizations of matrices stored outof-core. Concurrency and Computation: Practice and Experience, 35(22):e7726:1–e7726:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Halbiniak:2024:SMG

[HMR24] Kamil Halbiniak, Norbert Meyer, and Krzysztof Rojek. Single- and multi-GPU computing on NVIDIA- and AMDbased server platforms for solidification modeling application. Concurrency and Computation: Practice and Experience, 36 (9):e8000:1–e8000:??, April 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Heydari:2022:IWP

[HNG22] Ahmad Heydari, Sadegh Niroomand, and Harish Garg. An improved weighted principal component analysis integrated with TOPSIS approach for global financial development ranking problem of Middle East countries. Concurrency and Computation: Practice and Experience, 34(13):e6923:1–e6923:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hirose:2024:FGD

[HNOI24] Jion Hirose, Junya Nakamura, Fukuhito Ooshita, and Michiko Inoue. Fast gathering despite a linear number of weakly Byzantine agents. Concurrency and Computation: Practice and Experience, 36(14):e8055:1–e8055:??, June 25, 2024. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hasal:2021:CSP

[HNS^{+}21] Martin Hasal, Jana Nowaková, Khalifa Ahmed Saghair, Hussam Abdulla, Václav Snásel, and Lidia Ogiela. Chatbots: Security, privacy, data protection, and social aspects. Concurrency and Computation: Practice and Experience, 33(19): e6426:1–e6426:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hong:2022:DST

[HNSS22] Yu Hong, Zhenhu Ning, Yu Sun, and Changxiang Shen. A dual-system trusted computing node construction method based on ARM multi-core CPU architecture. Concurrency and Computation: Practice and Experience, 34(19):e7037:1– e7037:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Halbiniak:2021:DWP

[HOS^{+}21] Kamil Halbiniak, Tomasz Olas, Lukasz Szustak, Adam Kulawik, and Marco Lapegna. Dynamic workload prediction and distribution in numerical modeling of solidification on multi-/manycore architectures. Concurrency and Computation: Practice and Experience, 33(11):e5905:1–e5905:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hayyolalam:2022:SOS

[HPCK22] Vahideh Hayyolalam, Behrouz Pourghebleh, Mohammad Reza Chehrehzad, and Ali Asghar Pourhaji Kazem. Singleobjective service composition methods in cloud manufacturing systems: Recent techniques, classification, and future trends. Concurrency and Computation: Practice and Experience, 34(5):e6698:1–e6698:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hadri:2020:PSS

[HPH^{+}20] Bilel Hadri, Matteo Parsani, Maxwell Hutchinson, Alexander Heinecke, Lisandro Dalcin, and David Keyes. Performance study of sustained petascale direct numerical simulation on Cray XC40 systems. Concurrency and Computation: Practice and Experience, 32(20):e5725:1–e5725:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hummaida:2023:HDA

[HPS23] Abdul R. Hummaida, Norman W. Paton, and Rizos Sakellariou. A hierarchical decentralized architecture to enable adaptive scalable virtual machine migration. Concurrency and Computation: Practice and Experience, 35(2):e7487:1– e7487:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hameed:2022:BPP

[HR22a] Aarif Ahamed Shahul Hameed and Chandrasekar Ravi. Bitcoin price prediction using optimized multiplicative long short term memory with attention mechanism using modified cuckoo search optimization. Concurrency and Computation: Practice and Experience, 34(28):e7384:1–e7384:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hooda:2022:ITF

[HR22b] Diksha Hooda and Rinkle Rani. An interval type-2 fuzzy ontological model: Predicting water quality from sensory data. Concurrency and Computation: Practice and Experience, 34(28):e7377:1–e7377:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2022:RKM

[HSL^{+}22] Lu Huang, Jeho Song, Xuejiao Lin, Yachao Du, and Sidi Yang. Research on kick motion before Sanda based on 3D wireless sensor network image. Concurrency and Computation: Practice and Experience, 34(13):e5894:1–e5894:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Halbiniak:2021:EOH

[HSO^{+}21] Kamil Halbiniak, Lukasz Szustak, Tomasz Olas, Roman Wyrzykowski, and Pawel Gepner. Exploration of OpenCL heterogeneous programming for porting solidification modeling to CPU-GPU platforms. Concurrency and Computation: Practice and Experience, 33(4):e6011:1–e6011:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hafizpour:2023:NAB

[HSR23] Hadis Hafizpour, Mohammad Ebrahim Shiri, and Amir Masoud Rahmani. New attribute-based encryption schemes with anonymous authentication and time limitation in fog computing. Concurrency and Computation: Practice and Experience, 35(21):e7681:1–e7681:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Harimoorthy:2021:CAP

[HT21] Karthikeyan Harimoorthy and Menakadevi Thangavelu. Cloud-assisted Parkinson disease identification system for remote patient monitoring and diagnosis in the smart healthcare applications. Concurrency and Computation: Practice and Experience, 33(21):e6419:1–e6419:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2022:RMP

[HTVL22] Yu Huang, Yufei Tang, James VanZwieten, and Jianxun Liu. Reliable machine prognostic health management in the presence of missing data. Concurrency and Computation: Practice and Experience, 34(12):e5762:1–e5762:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

HoseinyFarahabady:2022:EDI

[HTZ^{+}22] MohammadReza HoseinyFarahabady, Javid Taheri, Albert Y. Zomaya, Zahir Tari, and Wei Bao. Enhancing disk input output performance in consolidated virtualized cloud platforms using a randomized approximation scheme. Concurrency and Computation: Practice and Experience, 34(2):e6247:1– e6247:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

HoseinyFarahabady:2023:EER

[HTZT23] MohammadReza HoseinyFarahabady, Javid Taheri, Albert Y. Zomaya, and Zahir Tari. Energy efficient resource controller for Apache Storm. Concurrency and Computation: Practice and Experience, 35(17):e6799:1–e6799:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2020:PIM

[Hua20] Fang Huang. Parallelization implementation of the multiscale Retinex image-enhancement algorithm based on a many integrated core platform. Concurrency and Computation: Practice and Experience, 32(22):e5832:1–e5832:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hussain:2022:EDF

[HUC^{+}22] Abid Hussain, Kalim Ullah, Salman A. Cheema, Akbar Ali Khan, and Zawar Hussain. Empirical distribution function based dual use of auxiliary information for the improved estimation of finite population mean. Concurrency and Computation: Practice and Experience, 34(27):e7346:1–e7346:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hemalatha:2022:ESN

[HUJ22] R. Hemalatha, R. Umamaheswari, and S. Jothi. An efficient stable node selection based on Garson’s pruned recurrent neural network and MSO model for multipath routing in MANET. Concurrency and Computation: Practice and Experience, 34(21):e7105:1–e7105:??, September 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Han:2021:RWH

[HV21] Nguyen Van Han and Phan Cong Vinh. Reasoning with words: a hedge algebra linguistic cognitive map approach. Concurrency and Computation: Practice and Experience, 33 (2):e5711:1–e5711:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2022:PAM

[HVB22] Guanlin He, Stephane Vialle, and Marc Baboulin. Parallel and accurate k-means algorithm on CPU–GPU architectures for spectral clustering. Concurrency and Computation: Practice and Experience, 34(14):e6621:1–e6621:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hagerty:2023:SPP

[HVT23] Nick Hagerty, Verónica G. Melesse Vergara, and Arnold Tharrington. Studying performance portability of LAMMPS across diverse GPU-based platforms. Concurrency and Computation: Practice and Experience, 35(28):e7895:1–e7895:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hao:2021:TSP

[HWBZ21] Zhiqiang Hao, Zhigang Wang, Dongxu Bai, and Shiyang Zhou. Towards the steel plate defect detection: Multidimensional feature information extraction and fusion. Concurrency and Computation: Practice and Experience, 33(21):e6384:1– e6384:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2023:NTG

[HWG23] Feng He, Zhiyu Wang, and Xiaoyan Gu. Network topology generation based on eigenvector centrality with real-time guarantee. Concurrency and Computation: Practice and Experience, 35(21):e6955:1–e6955:??, September 25, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2023:OPE

[HWY^{+}23] Li Huang, Cheng Wang, Juntong Yun, Bo Tao, Jinxian Qi, Ying Liu, Hongjie Ma, and Hui Yu. Object pose estimation based on stereo vision with improved K-D tree ICP algorithm. Concurrency and Computation: Practice and Experience, 35(21):e7714:1–e7714:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2022:FAV

[HXST22] Zhongzhe Hu, Junmin Xiao, Ninghui Sun, and Guangming Tan. Fast and accurate variable batch size convolution neural network training on large scale distributed systems. Concurrency and Computation: Practice and Experience, 34(21): e7119:1–e7119:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haugerud:2020:ACB

[HXY20] Hårek Haugerud, Noha Xue, and Anis Yazidi. On automated cloud bursting and hybrid cloud setups using Apache Mesos. Concurrency and Computation: Practice and Experience, 32(17):e5662:1–e5662:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2022:HAB

[HXY^{+}22a] Bo Hu, Hui Xiao, Nan Yang, Hao Jin, and Lei Wang. A hybrid approach based on double roulette wheel selection and quadratic programming for cardinality constrained portfolio optimization. Concurrency and Computation: Practice and Experience, 34(10):e6818:1–e6818:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2022:FND

[HXY^{+}22b] Bo Hu, Hui Xiao, Nan Yang, Lei Wang, and Hao Jin. Fast non-dominated sorting evolutionary algorithm II based on relative non-dominance matrix for portfolio optimization. Concurrency and Computation: Practice and Experience, 34(1): e6518:1–e6518:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2021:DDL

[HXZH21] Gang Hu, Zhuoyuan Xiang, Yin Zhang, and M. Shamim Hossain. DLIFT: a deep-learning-based intelligent fund transaction system for financial Internet of Things. Concurrency and Computation: Practice and Experience, 33(22):e5982:1– e5982:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2023:AOS

[HXZS23] Gaofeng He, Bingfeng Xu, Junjie Zhao, and Yadong Shi. Architecture-oriented security strategy determination for cyber-physical systems. Concurrency and Computation: Practice and Experience, 35(13):e6781:1–e6781:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2024:NME

[HYC^{+}24] Yulin He, Jin Yang, Yingchao Cheng, Xueqin Du, and Joshua Zhexue Huang. A novel member enhancement-based clustering ensemble algorithm. Concurrency and Computation: Practice and Experience, 36(10):e7992:1–e7992:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2020:ESA

[HYG20] Guixia He, Renjie Yin, and Jiaquan Gao. An efficient sparse approximate inverse preconditioning algorithm on GPU. Concurrency and Computation: Practice and Experience, 32(7): e5598:1–e5598:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

He:2023:QCE

[HYG^{+}23] Zhenzhen He, Jiong Yu, Tiquan Gu, Zhe Li, Xusheng Du, and Ping Li. Query cost estimation in graph databases via emphasizing query dependencies by using a neural reasoning network. Concurrency and Computation: Practice and Experience, 35(23):e7817:1–e7817:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hamd:2021:MLB

[HYK21] Ravyar Jasim Hamd, Tara Ali Yahiya, and Pinar Kirci. Multicellular 4G and load balancing over cloud computing. Concurrency and Computation: Practice and Experience, 33(23): e6294:1–e6294:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2021:PSA

[HYT^{+}21] Fang Huang, Hao Yang, Jian Tao, Jian Wang, and Xicheng Tan. Preliminary study on the automatic parallelism optimization model for image enhancement algorithms based on Intel’s(R) Xeon Phi. Concurrency and Computation: Practice and Experience, 33(16):e6260:1–e6260:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Han:2022:TAR

[HZD^{+}22] Zhijie Han, Qingfang Zhang, Xiaoyu Du, Kun Guo, and Mingshu He. Topology analysis and routing algorithms design for PTNet network. Concurrency and Computation: Practice and Experience, 34(8):e5767:1–e5767:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huo:2024:PQS

[HZQ^{+}24] Yachao Huo, Zongqu Zhao, Panke Qin, Shujing Wang, and Chengfu Zheng. Post-quantum secure two-party computing protocols against malicious adversaries. Concurrency and Computation: Practice and Experience, 36(4):e7923:1– e7923:??, February 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hu:2023:RGR

[HZW^{+}23] Yonghua Hu, Xin Zhang, Shuying Wang, Wei Liang, and Kuan-Ching Li. Research on global register allocation for code containing array-unit dual-usage register names. Concurrency and Computation: Practice and Experience, 35(19): e7519:1–e7519:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Han:2021:STA

[HZY^{+}21] Mengting Han, Xuan Zhang, Xin Yuan, Jiahao Jiang, Wei Yun, and Chen Gao. A survey on the techniques, applications, and performance of short text semantic similarity. Concurrency and Computation: Practice and Experience, 33(5): e5971:1–e5971:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2023:DPR

[HZZ^{+}23] Wen Huang, Ming Zhuo, Tianqing Zhu, Shijie Zhou, and Yongjian Liao. Differential privacy: Review of improving utility through cryptography-based technologies. Concurrency and Computation: Practice and Experience, 35(5):e7565:1– e7565:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2022:MNN

[HZZC22] Zhaohong Huang, Xiangchen Zhang, Guowei Zhang, and Guorong Cai. MSSA-Net: a novel multi-scale feature fusion and global self-attention network for lesion segmentation. Concurrency and Computation: Practice and Experience, 34(21):e7060:1–e7060:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ibrahim:2022:EAI

[IA22] Mostafa E. A. Ibrahim and Alaa E. S. Ahmed. Energyaware intelligent hybrid routing protocol for wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(3):e6601:1–e6601:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ikeya:2023:DCC

[IA23] Toshi Ikeya and Masaki Aida. Derivation and characteristics of closed-form solutions of the fundamental equations for online user dynamics. Concurrency and Computation: Practice and Experience, 35(14):e6619:1–e6619:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Imdoukh:2020:OSD

[IAA20] Mahmoud Imdoukh, Imtiaz Ahmad, and Mohammad Alfailakawi. Optimizing scheduling decisions of container management tool using many-objective genetic algorithm. Concurrency and Computation: Practice and Experience, 32(5): e5536:1–e5536:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Isuwa:2022:HPS

[IAAA22] Jeremiah Isuwa, Mohammed Abdullahi, Yusuf Sahabi Ali, and Abdulrazaq Abdulrahim. Hybrid particle swarm optimization with sequential one point flipping algorithm for feature selection. Concurrency and Computation: Practice and Experience, 34(25):e7239:1–e7239:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ibaida:2023:ECT

[IAASK23] Ayman Ibaida, Sharif Abuadbba, Dhiah Al-Shammary, and Ibrahim Khalil. ECG compression technique using fast fractals in the Internet of medical things. Concurrency and Computation: Practice and Experience, 35(23):e7812:1–e7812:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ignat:2021:ERC

[IAO21] Claudia-Lavinia Ignat, Luc André, and Gérald Oster. Enhancing rich content wikis with real-time collaboration. Concurrency and Computation: Practice and Experience, 33(8): e4110:1–e4110:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Iqbal:2020:AIA

[IAQ20] Saleem Iqbal, Abdul Hanan Abdullah, and Kashif Naseer Qureshi. An adaptive interference-aware and traffic-aware channel assignment strategy for backhaul networks. Concurrency and Computation: Practice and Experience, 32(11): e5650:1–e5650:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Iqbal:2023:PEA

[IAT^{+}23] Tariq Iqbal, Nada M. Alfaer, Muhammad H. Tahir, Hassan M. Aljohani, Farrukh Jamal, and Ahmed Z. Afify. Properties and estimation approaches of the odd JCA family with applications. Concurrency and Computation: Practice and Experience, 35(1):e7417:1–e7417:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Iqbal:2022:OAD

[IDA22] Mohammed Asim Iqbal, Krishnamoorthy Devarajan, and Syed Musthak Ahmed. An optimal asthma disease detection technique for voice signal using hybrid machine learning technique. Concurrency and Computation: Practice and Experience, 34(11):e6856:1–e6856:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Immanuel:2024:OLE

[IG24] Paul Immanuel and A. Berin Greeni. Optimization of layout for embedding half hypercube into conventional tree architectures. Concurrency and Computation: Practice and Experience, 36(12):e8058:1–e8058:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Incardona:2023:PCL

[IGYS23] Pietro Incardona, Aryaman Gupta, Serhii Yaskovets, and Ivo F. Sbalzarini. A portable C++ library for memory and compute abstraction on multi-core CPUs and GPUs. Concurrency and Computation: Practice and Experience, 35(25): e7870:1–e7870:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ishigami:2023:ICU

[IIK^{+}23] Tomohiro Ishigami, Teijiro Isokawa, Naotake Kamiura, Hiroki Masumoto, and Hitoshi Tabuchi. Instillation checking using long short-term memories for ophthalmology patients. Concurrency and Computation: Practice and Experience, 35 (19):e7466:1–e7466:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Isunuri:2022:TCB

[IK22] Bala Venkateswarlu Isunuri and Jagadeesh Kakarla. Threeclass brain tumor classification from magnetic resonance images using separable convolution based neural network. Concurrency and Computation: Practice and Experience, 34(1): e6541:1–e6541:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Iqbal:2022:IGL

[IMNC22] Anam Iqbal, Tahir Mahmood, Hafiz Zafar Nazir, and Niladri Chakraborty. On the improved generalized linear model-based monitoring methods for Poisson distributed processes. Concurrency and Computation: Practice and Experience, 34(11): e6889:1–e6889:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Imanaga:2023:SIT

[INY^{+}23] Tomohiro Imanaga, Koji Nakano, Ryota Yasudo, Yasuaki Ito, Yuya Kawamata, Ryota Katsuki, Yusuke Tabata, Takashi Yazane, and Kenichiro Hamano. Simple iterative trial search for the maximum independent set problem optimized for the GPUs. Concurrency and Computation: Practice and Experience, 35(14):e6681:1–e6681:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ivanov:2020:ICF

[IP20] Todor Ivanov and Matteo Pergolesi. The impact of columnar file formats on SQL-on-Hadoop engine performance: a study on ORC and Parquet. Concurrency and Computation: Practice and Experience, 32(5):e5523:1–e5523:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Iserte:2021:IME

[IPRS21] Sergio Iserte, Javier Prades, Carlos Reaño, and Federico Silla. Improving the management efficiency of GPU workloads in data centers through GPU virtualization. Concurrency and Computation: Practice and Experience, 33(2):e5275:1– e5275:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Iqbal:2022:MDS

[IQS^{+}22] Saleem Iqbal, Kashif Naseer Qureshi, Faisal Shoaib, Awais Ahmad, and Gwanggil Jeon. Minimize the delays in software defined network switch controller communication. Concurrency and Computation: Practice and Experience, 34(13): e5940:1–e5940:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ilic:2023:SII

[IS23] Aleksandar Ilic and Leonel Sousa. Special issue: 20th international workshop on algorithms, models and tools for parallel computing on heterogeneous platforms (HeteroPar’22). Concurrency and Computation: Practice and Experience, 35(25): e7915:1–e7915:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ipek:2021:HHA

[ITO21] Anil Ipek, Suleyman Tosun, and Suat Ozdemir. HAFTA: Highly adaptive fault-tolerant routing algorithm for twodimensional network-on-chips. Concurrency and Computation: Practice and Experience, 33(21):e6378:1–e6378:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Iyer:2023:SVM

[IVP^{+}23] Srikrishna Iyer, Thangappa Velmurugan, Periasamy Prakasam, Durairaj Sumathi, and Thengalpalayam Rajamanickam Suresh Kumar. Support Vector Machine based spectrum handoff scheme for seamless handover in Cognitive Radio Networks. Concurrency and Computation: Practice and Experience, 35 (4):e7534:1–e7534:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jayapalan:2023:ITB

[JA23] Daniel Francis Selvaraj Jayapalan and John Patrick Ananth. Internet of Things-based root disease classification in alfalfa plants using hybrid optimization-enabled deep convolutional neural network. Concurrency and Computation: Practice and Experience, 35(3):e7504:1–e7504:??, February 1, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jeon:2021:SIR

[JAC^{+}21] Gwanggil Jeon, Chehri Abdellah, Salvatore Cuomo, Sadia Din, and Sohail Jabbar. Special issue on real-time behavioral monitoring in IoT applications using big data analytics. Concurrency and Computation: Practice and Experience, 33 (4):e5529:1–e5529:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jafari:2024:DPA

[JAM24] Somayyeh Jafari and Nasrin Aghaee-Maybodi. Detection of phishing addresses and pages with a data set balancing approach by generative adversarial network (GAN) and convolutional neural network (CNN) optimized with swarm intelligence. Concurrency and Computation: Practice and Experience, 36(11):e8033:1–e8033:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jalili:2024:MCB

[JAR^{+}24] Ahmad Jalili, Jafar A. Alzubi, Roghayeh Rezaei, Julian L. Webber, Christian Fernández-Campusano, Mehdi Gheisari, Rashid Amin, and Abolfazl Mehbodniya. Markov chainbased analysis and fault tolerance technique for enhancing chain-based routing in WSNs. Concurrency and Computation: Practice and Experience, 36(12):e8032:1–e8032:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jeon:2020:SIC

[JB20] Gwanggil Jeon and Valerio Bellandi. Special issue on computational intelligence techniques for industrial and medical applications. Concurrency and Computation: Practice and Experience, 32(17):e5717:1–e5717:??, September 10, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jayabal:2021:PAD

[JB21] Chandra Priya Jayabal and Ponsy R. K. Sathia Bhama. Performance analysis on Diversity Mining-based Proof of Work in bifolded consortium blockchain for Internet of Things consensus. Concurrency and Computation: Practice and Experience, 33(16):e6285:1–e6285:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jagadeesh:2022:DFD

[JB22a] M. Jagadeesh and B. Baranidharan. Dynamic FERNet: Deep learning with optimal feature selection for face expression recognition in video. Concurrency and Computation: Practice and Experience, 34(28):e7373:1–e7373:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jain:2022:PEO

[JB22b] Abhishek Jain and Amit Kumar Bhardwaj. Power-efficient optimized clustering method with intelligent fog computing for wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(15):e6983:1–e6983:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jadhav:2024:QIR

[JB24] Poonam Samir Jadhav and Gautam M. Borkar. Quasiidentifier recognition with echo chamber optimization-based anonymization for privacy preservation of cloud storage. Concurrency and Computation: Practice and Experience, 36(2): e7906:1–e7906:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Janjic:2021:RIT

[JBBH21] Vladimir Janjic, Christopher Brown, Adam Barwell, and Kevin Hammond. Refactoring for introducing and tuning parallelism for heterogeneous multicore machines in Erlang. Concurrency and Computation: Practice and Experience, 33(14): e5420:1–e5420:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jung:2020:PIM

[JBCI20] ByeongHo Jung, Seong Il Bae, Chang Choi, and Eul Gyu Im. Packer identification method based on byte sequences. Concurrency and Computation: Practice and Experience, 32(8): e5082:1–e5082:??, April 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jeon:2021:SIA

[JC21a] Gwanggil Jeon and Abdellah Chehri. Special issue on applied computational intelligence. Concurrency and Computation: Practice and Experience, 33(22):e6632:1–e6632:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Joseph:2021:NIR

[JC21b] Christina Terese Joseph and Kandasamy Chandrasekaran. Nature-inspired resource management and dynamic rescheduling of microservices in cloud datacenters. Concurrency and Computation: Practice and Experience, 33(17):e6290:1– e6290:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jerbi:2022:BBS

[JCG^{+}22] Wassim Jerbi, Omar Cheikhrouhou, Abderrahmen Guermazi, Mohamed Baz, and Hafedh Trabelsi. BSI: Blockchain to secure routing protocol in Internet of Things. Concurrency and Computation: Practice and Experience, 34(10):e6794:1– e6794:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jerbi:2024:EMT

[JCGT24] Wassim Jerbi, Omar Cheikhrouhou, Abderrahmen Guermazi, and Hafedh Trabelsi. An enhanced MSU-TSCH scheduling algorithms for industrial wireless sensor networks. Concurrency and Computation: Practice and Experience, 36(5):e7938:1– e7938:??, February 29, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2020:CEA

[JCL^{+}20] Changbing Jiang, Tinggui Chen, Ruolan Li, Liang Li, Gongfa Li, Chonghuan Xu, and Shufang Li. Construction of extended ant colony labor division model for traffic signal timing and its application in mixed traffic flow model of single intersection. Concurrency and Computation: Practice and Experience, 32(7):e5592:1–e5592:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jayaswal:2022:FMD

[JD22] Ruchi Jayaswal and Manish Dixit. A face mask detection system: an approach to fight with COVID-19 scenario. Concurrency and Computation: Practice and Experience, 34(28): e7394:1–e7394:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Juneja:2023:PVA

[JD23] Vikas Juneja and Shail Kumar Dinkar. A predictive vampire attack detection by social spider optimized Gaussian mixture model clustering. Concurrency and Computation: Practice and Experience, 35(2):e7481:1–e7481:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Juneja:2022:ACO

[JDG22] Vikas Juneja, Shail Kumar Dinkar, and Dharam Vir Gupta. An anomalous co-operative trust and PG-DRL based vampire attack detection and routing. Concurrency and Computation: Practice and Experience, 34(3):e6557:1–e6557:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jagannath:2022:IES

[JDLP22] Duraiswamy Jothinath Jagannath, Raveena Judie Dolly, Gunamony Shine Let, and James Dinesh Peter. An IoT enabled smart healthcare system using deep reinforcement learning. Concurrency and Computation: Practice and Experience, 34(28):e7403:1–e7403:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jeon:2020:SGB

[Jeo20] Gwanggil Jeon. Structure and gradient-based industrial interpolation using computational intelligence. Concurrency and Computation: Practice and Experience, 32(17):e5439:1– e5439:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jha:2021:SEH

[JGJ^{+}21] Devki Nandan Jha, Saurabh Garg, Prem Prakash Jayaraman, Rajkumar Buyya, Zheng Li, Graham Morgan, and Rajiv Ranjan. A study on the evaluation of HPC microservices in containerized environment. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2020:OSD

[JGW20] Tao Jiang, Peiming Guo, and Jianping Wu. One-sided ondemand communication technology for the semi-Lagrange scheme in the YHGSM. Concurrency and Computation: Practice and Experience, 32(7):e5586:1–e5586:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jain:2021:SMM

[JH21] Usha Jain and Muzzammil Hussain. Security mechanism for maritime territory and frontier surveillance in naval operations using wireless sensor networks. Concurrency and Computation: Practice and Experience, 33(17):e6300:1–e6300:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jin:2021:PAC

[JHD^{+}21] Xin Jin, Rui Han, Yuwei Duan, Ning Ning, and Xiaodong Li. AR CAPTCHA: Recognizing robot by augmented reality. Concurrency and Computation: Practice and Experience, 34(1): e5585:1–e5585:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Javadian:2023:ISA

[JHG23] Mohammad Javadian, Sajad Hayati, and Vahid Ghasemi. An intelligent signboard to assign the speed limit on ways using ANN. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ji:2021:CSS

[JHS^{+}21] Sai Ji, Rui Huang, Jian Shen, Xin Jin, and Youngju Cho. A certificateless signcryption scheme for smart home networks. Concurrency and Computation: Practice and Experience, 33 (7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2020:CED

[JHZ20] Tianming Jiang, Ping Huang, and Ke Zhou. Cost-efficiency disk failure prediction via threshold-moving. Concurrency and Computation: Practice and Experience, 32(14):e5669:1– e5669:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiao:2022:ADN

[Jia22] Yuhan Jiao. Analysis and design of a new fault-tolerant digital comparator based on nano-scale quantum technology. Concurrency and Computation: Practice and Experience, 34(20): e7158:1–e7158:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2021:CME

[JJZ^{+}21] Dazhi Jiang, Donghui Jin, Jiaxi Zhuang, Daqiang Tan, Dicheng Chen, and Yujie Liang. A computational model of emotion based on audio-visual stimuli understanding and personalized regulation with concurrency. Concurrency and Computation: Practice and Experience, 33(17):e6269:1– e6269:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jaiswal:2022:CSS

[JK22a] Deepak Jaiswal and Praveen Kumar. A comparative study on SoC embedded low power GPUs for real-time edge-based automated traffic surveillance. Concurrency and Computation: Practice and Experience, 34(10):e6736:1–e6736:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jaiswal:2022:SPC

[JK22b] Deepak Jaiswal and Praveen Kumar. A survey on parallel computing for traditional computer vision. Concurrency and Computation: Practice and Experience, 34(4):e6638:1– e6638:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jain:2022:LRV

[JKB22] Rahul Jain, Ram Kumar Karsh, and Abul Abbas Barbhuiya. Literature review of vision-based dynamic gesture recognition using deep learning techniques. Concurrency and Computation: Practice and Experience, 34(22):e7159:1–e7159:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jo:2021:ECA

[JKKL21] Jinhyuk Jo, Song Ko, Dae-Won Kim, and Jaesung Lee. Effective computer-assisted pronunciation training based on phone-sensitive word recommendation. Concurrency and Computation: Practice and Experience, 33(2):e5151:1– e5151:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jegatheesan:2022:ANF

[JKP22] A Jegatheesan, N Sathish Kumar, and C Anna Palagan. Adaptive neuro-fuzzy inference system-based energy conservation system for performance enhancement of MANET. Concurrency and Computation: Practice and Experience, 34 (8):e5348:1–e5348:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ju:2020:DMC

[JKS20] YeongJi Ju, MinGu Kim, and JuHyun Shin. Detection of malicious code using the direct hashing and pruning and support vector machine. Concurrency and Computation: Practice and Experience, 32(18):e5483:1–e5483:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jyothi:2022:IDL

[JLE22] Bankapalli Jyothi, Sumalatha Lingamgunta, and Suneetha Eluri. Intelligent deep learning-based hierarchical clustering for unstructured text data. Concurrency and Computation: Practice and Experience, 34(28):e7388:1–e7388:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jia:2024:EAF

[JLZ^{+}24] Juncheng Jia, Ji Liu, Chendi Zhou, Hao Tian, Mianxiong Dong, and Dejing Dou. Efficient asynchronous federated learning with sparsification and quantization. Concurrency and Computation: Practice and Experience, 36(9):e8002:1– e8002:??, April 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jha:2022:EAB

[JM22] Saurabh Jha and Ashok Kumar Mehta. An evolutionary algorithm based feature selection and fuzzy rule reduction technique for the prediction of skin cancer. Concurrency and Computation: Practice and Experience, 34(5):e6694:1–e6694:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

J:2023:AFR

[JM23] Jeyshri J. and Kowsigan M. Adaptive fuzzy-region growing fusion and improved CNN-ANFIS-based automated segmentation and classification of cervical cancer. Concurrency and Computation: Practice and Experience, 35(3):e7493:1– e7493:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jackulin:2024:IDN

[JMV24] C. Jackulin, S. Murugavalli, and K. Valarmathi. IFATA-Deep net: Improved invasive feedback artificial tree algorithm with deep quantum neural network for root disease classification. Concurrency and Computation: Practice and Experience, 36 (6):e7946:1–e7946:??, March 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2021:AMC

[JMY21] Rong Jiang, Zifei Ma, and Juan Yang. An assessment model for cloud service security risk based on entropy and support vector machine. Concurrency and Computation: Practice and Experience, 33(21):e6423:1–e6423:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jaishankar:2022:MPI

[JNM^{+}22] B. Jaishankar, J. Naveen, Balamurugan Marimuthu, Bhuvana Jayabalan, and M. Mathivanan. A multi-preference integrated algorithm for deep learning based recommender framework. Concurrency and Computation: Practice and Experience, 34(25):e7241:1–e7241:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jorge:2021:CFH

[JNMG21] Carlos A. C. Jorge, Alexandre S. Nery, Alba C. M. A. Melo, and Alfredo Goldman. A CPU-FPGA heterogeneous approach for biological sequence comparison using high-level synthesis. Concurrency and Computation: Practice and Experience, 33(4):e6007:1–e6007:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jena:2022:CNN

[JNS22] Biswajit Jena, Gopal Krishna Nayak, and Sanjay Saxena. Convolutional neural network and its pretrained models for image classification and object detection: a survey. Concurrency and Computation: Practice and Experience, 34(6): e6767:1–e6767:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jesus:2023:VDN

[JOW23] Ricardo Jesus, Tomás Oliveira e Silva, and Michèle Weiland. Vectorizing and distributing number-theoretic transform to count Goldbach partitions on Arm-based supercomputers. Concurrency and Computation: Practice and Experience, 35(28):e7882:1–e7882:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

James:2021:ICC

[JP21] K. Immanuvel Arokia James and R. Prabakaran. Intracluster cooperative communication in smart home scenario through stable election protocol. Concurrency and Computation: Practice and Experience, 33(3):e5365:1–e5365:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jacob:2023:SAS

[JPA^{+}23] Susmi Jacob, Vinod Puthuvath, Muralidharan Akarsh, Jackson George, Jewel Joseph, and Jigil Joy. A smartphone authentication system based on touch gesture dynamics. Concurrency and Computation: Practice and Experience, 35(1): e7449:1–e7449:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jeon:2021:IEE

[JPAA21] Gwanggil Jeon, Kitsuchart Pasupa, Marco Anisetti, and Awais Ahmad. Image enhancement in embedded devices for internet of things. Concurrency and Computation: Practice and Experience, 33(3):e5398:1–e5398:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jain:2023:LSE

[JPH23] Usha Jain, Saied Pirasteh, and Muzzammil Hussain. Lightweight, secure, efficient, and dynamic scheme for mutual authentication of devices in Internet-of-things-fog environment. Concurrency and Computation: Practice and Experience, 35(1):e7428:1–e7428:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jaganathan:2022:ODM

[JPK22] Thirumalai Jaganathan, Anandan Panneerselvam, and Senthil Kumar Kumaraswamy. Object detection and multiobject tracking based on optimized deep convolutional neural network and unscented Kalman filtering. Concurrency and Computation: Practice and Experience, 34(25):e7245:1– e7245:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Janssen:2022:GPU

[JPL22] Dylan M. Janssen, Wayne Pullan, and Alan Wee-Chung Liew. Graphics processing unit acceleration of the island model genetic algorithm using the CUDA programming platform. Concurrency and Computation: Practice and Experience, 34 (2):e6286:1–e6286:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jacob:2021:ESS

[JPN21] T. Prem Jacob, A. Pravin, and G. Nagarajan. Efficient spectrum sensing framework for cognitive networks. Concurrency and Computation: Practice and Experience, 33(3):e5187:1– e5187:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Junior:2021:WBR

[JPO^{+}21] Luís Carlos Silva Júnior, Maricéia B. S. Pádua, Leonardo M. Ogusuku, Marcelo Keese Albertini, Renato Pimentel, and André R. Backes. Wild boar recognition using convolutional neural networks. Concurrency and Computation: Practice and Experience, 33(22):e6010:1–e6010:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jian:2022:ABD

[JQ22] Zhichao Jian and Lichun Qin. The application of big data network crawler technology for architectural culture and environment protection. Concurrency and Computation: Practice and Experience, 34(9):e5769:1–e5769:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2020:LDU

[JQGL20] Hua Jiang, Weidi Qiu, Weixia Gui, and Jiarong Liang. Local diagnosability under the PMC model with application to matching composition networks. Concurrency and Computation: Practice and Experience, 32(6):e5175:1–e5175:??, March 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jacob:2022:FAB

[JR22] Minu Susan Jacob and P. Selvi Rajendran. Fuzzy artificial bee colony-based CNN-LSTM and semantic feature for fake product review classification. Concurrency and Computation: Practice and Experience, 34(1):e6539:1–e6539:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Joseph:2023:DJA

[JR23a] Dani Reagan Vivek Joseph and Shantha Selvakumari Ramapackiyam. Distributed job allocation using response threshold for heterogeneous robot team under deadline constraints. Concurrency and Computation: Practice and Experience, 35 (8):e7623:1–e7623:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Joseph:2023:HBL

[JR23b] Linu Joseph and Dhanapal Ramasamy. Hybrid bidirectional long short term memory with black widow optimization for crop yield prediction using data mining. Concurrency and Computation: Practice and Experience, 35(23):e7775:1– e7775:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jangu:2024:SAC

[JR24] Nupur Jangu and Zahid Raza. Smart admission control strategy utilizing volunteer-enabled fog–cloud computing. Concurrency and Computation: Practice and Experience, 36(2): e7908:1–e7908:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jain:2022:TVD

[JS22a] Khushboo Jain and Akansha Singh. A two-vector dataprediction model for energy-efficient data-aggregation in wireless sensor network. Concurrency and Computation: Practice and Experience, 34(11):e6898:1–e6898:??, May 15, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jain:2022:PFW

[JS22b] Priyank Jain and Shriya Sahu. Prediction and forecasting of worldwide corona virus (COVID-19) outbreak using time series and machine learning. Concurrency and Computation: Practice and Experience, 34(26):e7286:1–e7286:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jain:2022:DIU

[JS22c] Somya Jain and Adwitiya Sinha. Discovering influential users in social network using weighted cumulative centrality. Concurrency and Computation: Practice and Experience, 34(1): e6521:1–e6521:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jia:2023:CDP

[JS23] Meng Jia and Zheren Sun. Computer data processing method of map manifold based on derivative transfer approach. Concurrency and Computation: Practice and Experience, 35(23): e7808:1–e7808:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jamil:2020:JSA

[JSA^{+}20] Bushra Jamil, Mohammad Shojafar, Israr Ahmed, Atta Ullah, Kashif Munir, and Humaira Ijaz. A job scheduling algorithm for delay and performance optimization in fog computing. Concurrency and Computation: Practice and Experience, 32(7):e5581:1–e5581:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jenefer:2022:DCA

[JSAA22] Balraj M. Monica Jenefer, K. Senathipathi, Aarthi, and Annapandi. Detection and categorization of acute intracranial hemorrhage subtypes using a multilayer DenseNet-ResNet architecture with improved random forest classifier. Concurrency and Computation: Practice and Experience, 34(22): e7167:1–e7167:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jin:2020:SIG

[JSLL20] Hai Jin, Xipeng Shen, Robert Lovas, and Xiaofei Liao. Special issue: Graph computing. Concurrency and Computation: Practice and Experience, 32(3):e5452:1–e5452:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jeyaraj:2020:IMSa

[JSP20a] Rathinaraja Jeyaraj, Ananthanarayana V S, and Anand Paul. Improving MapReduce scheduler for heterogeneous workloads in a heterogeneous environment. Concurrency and Computation: Practice and Experience, 32(7):e5558:1–e5558:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jeyaraj:2020:IMSb

[JSP20b] Rathinaraja Jeyaraj, Ananthanarayana V S, and Anand Paul. Improving MapReduce scheduler for heterogeneous workloads in a heterogeneous environment. Concurrency and Computation: Practice and Experience, 32(17):e5978:1–e5978:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jaber:2022:MPC

[JSS22] Murtadha Jaber, Reza Sheibani, and Hassan Shakeri. A model for predicting crimes using big data and neural-fuzzy networks. Concurrency and Computation: Practice and Experience, 34(17):e6985:1–e6985:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jararweh:2020:ICC

[JSYAA20] Yaser Jararweh, Mohammed A. Shehab, Qussai Yaseen, and Mahmoud Al-Ayyoub. Improving classification and clustering techniques using GPUs. Concurrency and Computation: Practice and Experience, 32(21):e5538:1–e5538:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jordan:2022:SPD

[JSZS22] Herbert Jordan, Pavle Subotić, David Zhao, and Bernhard Scholz. Specializing parallel data structures for Datalog. Concurrency and Computation: Practice and Experience, 34 (2):e5643:1–e5643:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

J:2023:ARS

[JT23] Jerisha Liby J and Jaya T. Advanced reversible stenographic method for image and video frames by adapting dual tree complex wavelet transform and chaotic whale optimization algorithm. Concurrency and Computation: Practice and Experience, 35(22):e7719:1–e7719:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2021:SER

[JTY^{+}21] Lin Jiang, Ping Tan, Junfeng Yang, Xingbao Liu, and Chao Wang. Speech emotion recognition using emotion perception spectral feature. Concurrency and Computation: Practice and Experience, 33(11):e5427:1–e5427:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2021:MRA

[JW21] Suhan Jiang and Jie Wu. Multi-resource allocation in cloud data centers: a trade-off on fairness and efficiency. Concurrency and Computation: Practice and Experience, 33(6): e6061:1–e6061:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jing:2022:DDB

[JW22] Hengchang Jing and Jian Wang. DDoS detection based on graph structure features and non-negative matrix factorization. Concurrency and Computation: Practice and Experience, 34(9):e5783:1–e5783:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2020:IKB

[JWL20] Jialin Jiang, Xinzhi Wang, and Xiangfeng Luo. Interdisciplinary knowledge-based implicit emotion recognition. Concurrency and Computation: Practice and Experience, 32(22): e5838:1–e5838:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jianmin:2020:IJF

[JWT^{+}20] Wang Jianmin, Zhao Wenbin, Fan Tongrang, Yang Shilong, and Lv Hongwei. An improved join-free snowflake schema for ETL and OLAP of data warehouse. Concurrency and Computation: Practice and Experience, 32(23):e5519:1–e5519:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ji:2023:SGA

[JXL^{+}23] Sai Ji, Dachuan Xu, Min Li, Yishui Wang, and Dongmei Zhang. Stochastic greedy algorithms for maximizing constrained submodular + supermodular functions. Concurrency and Computation: Practice and Experience, 35(17):e6575:1– e6575:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jia:2020:RJR

[JY20] Hailong Jia and Jie Yang. Research on joint ranking recommendation model based on Markov chain. Concurrency and Computation: Practice and Experience, 32(6):e5191:1– e5191:??, March 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2021:NIC

[JYC^{+}21] Lin Jiang, Yumei Yi, Defeng Chen, Ping Tan, and Xingbao Liu. A novel infant cry recognition system using auditory model-based robust feature and GMM-UBM. Concurrency and Computation: Practice and Experience, 33(11):e5405:1– e5405:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2024:PPM

[JYH^{+}24] Hao Jiang, Qizhou Yu, Dan Han, Yaqing Chen, and Zejun Li. A path planning method for unmanned aerial vehicle based on improved wolf pack algorithm. Concurrency and Computation: Practice and Experience, 36(14):e8095:1–e8095:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2023:UAC

[JYL^{+}23] Lei Jiang, Peng Yuan, Jing Liao, Qiongbing Zhang, Jianxun Liu, and Keqin Li. Undersampling of approaching the classification boundary for imbalance problem. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2020:NSF

[JYW^{+}20] Lin Jiang, Shaoqian Yu, Xiaochen Wang, Chao Wang, and Tonghan Wang. A new source-filter model audio bandwidth extension using high frequency perception feature for IoT communications. Concurrency and Computation: Practice and Experience, 32(13):e4638:1–e4638:??, July 10, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2020:SSA

[JZB20] Fengzhen Jiang, Xu Zhao, and Qingguo Bai. Simulation and stability analysis of conflict events between employees and organization based on the social network. Concurrency and Computation: Practice and Experience, 32(6):e5097:1– e5097:??, March 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jiang:2023:QST

[JZC^{+}23] Weijin Jiang, Wanqing Zhang, Pingping Chen, Junpeng Chen, Yongxia Sun, and Ying Yang. Quantity sensitive task allocation based on improved whale optimization algorithm in crowdsensing system. Concurrency and Computation: Practice and Experience, 35(20):e6637:1–e6637:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jia:2021:QER

[JZL21] Menghan Jia, Yiming Zhang, and Dongsheng Li. QRDF: an efficient RDF graph processing system for fast query. Concurrency and Computation: Practice and Experience, 33(24): e6441:1–e6441:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ji:2022:PPD

[JZL22] Shuo Ji, Yinliang Zhao, and Yuxiang Li. Performance prediction for distributed graph computing. Concurrency and Computation: Practice and Experience, 34(12):e5961:1–e5961:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jameii:2022:IFT

[JZR22] Seyed Mahdi Jameii, Romina Sadat Zamirnaddafi, and Reza Rezabakhsh. Internet of flying things security: a systematic review. Concurrency and Computation: Practice and Experience, 34(24):e7213:1–e7213:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

K:2022:MSD

[K22] Mohaideen Pitchai K. Modeling and simulation of defense game model for jamming attack in wireless sensor networks using evolutionary game theory. Concurrency and Computation: Practice and Experience, 34(5):e6742:1–e6742:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karakaplan:2021:CSC

[KA21a] Mustafa Karakaplan and Fatih Mehmet Avcu. Classification of some chemical drugs by genetic algorithm and deep neural network hybrid method. Concurrency and Computation: Practice and Experience, 33(13):e6242:1–e6242:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Korkmaz:2021:FAT

[KA21b] Ersin Korkmaz and Ali Payidar Akgüngör. The forecasting of air transport passenger demands in Turkey by using novel meta-heuristic algorithms. Concurrency and Computation: Practice and Experience, 33(16):e6263:1–e6263:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khan:2022:EDD

[KA22a] Angshuman Khan and Rajeev Arya. Efficient design of dualmode nano counter: an approach using quantum dot cellular automata. Concurrency and Computation: Practice and Experience, 34(13):e6910:1–e6910:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Korkmaz:2022:OCL

[KA22b] Ersin Korkmaz and Ali Payidar Akgüngör. Optimum cycle length models using atom search optimization and grasshopper optimization algorithms. Concurrency and Computation: Practice and Experience, 34(5):e6732:1–e6732:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumari:2022:DSS

[KA22c] Nancy Kumari and Debi Prasanna Acharjya. A decision support system for diagnosis of hepatitis disease using an integrated rough set and fish swarm algorithm. Concurrency and Computation: Practice and Experience, 34(21):e7107:1– e7107:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaur:2023:RTT

[KA23] Amanjot Kaur and Nitin Auluck. Real-time trust aware scheduling in fog-cloud systems. Concurrency and Computation: Practice and Experience, 35(10):e7680:1–e7680:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2023:NSC

[KAAR23] R. Raja Kumar, Jegadeesh Athimoolam, Ahilan Appathurai, and Surendiran Rajendiran. Novel segmentation and classification algorithm for detection of tomato leaf disease. Concurrency and Computation: Practice and Experience, 35(12): e7674:1–e7674:??, May 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kabakus:2022:NCS

[Kab22] Abdullah Talha Kabakus. A novel COVID-19 sentiment analysis in Turkish based on the combination of convolutional neural network and bidirectional long-short term memory on Twitter. Concurrency and Computation: Practice and Experience, 34(22):e6883:1–e6883:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kabakus:2023:NRC

[Kab23] Abdullah Talha Kabakus. A novel robust convolutional neural network for uniform resource locator classification from the view of cyber security. Concurrency and Computation: Practice and Experience, 35(3):e7517:1–e7517:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khosrowshahi-Asl:2020:DFS

[KABE^{+}20] Ehsan Khosrowshahi-Asl, Jamal Bentahar, Rebeca Estrada, Hadi Otrok, Rabeb Mizouni, and Babak Khosravifar. Dynamic formation of service communities in the cloud under distribution and incomplete information settings. Concurrency and Computation: Practice and Experience, 32(1): e4338:1–e4338:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kalay:2022:NAB

[Kal22] Mustafa Utku Kalay. Novel approaches on bulk-loading of large scale spatial datasets. Concurrency and Computation: Practice and Experience, 34(9):e6596:1–e6596:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kannan:2022:ACD

[Kan22] Sridharan Kannan. An automated clinical decision support system for predicting cardiovascular disease using ensemble learning approach. Concurrency and Computation: Practice and Experience, 34(18):e7007:1–e7007:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Konakoglu:2022:AMG

[KAO22] Berkant Konakoglu, Salih Berkan Aydemir, and Funda Kutlu Onay. Application of a metaheuristic gradient-based optimizer algorithm integrated into artificial neural network model in a local geoid modeling with global navigation satellite systems/leveling measurements. Concurrency and Computation: Practice and Experience, 34(18):e7017:1–e7017:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khan:2020:BEC

[KAP20] Zaheer Khan, Abdul Ghafoor Abbasi, and Zeeshan Pervez. Blockchain and edge computing-based architecture for participatory smart city applications. Concurrency and Computation: Practice and Experience, 32(12):e5566:1–e5566:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khani:2023:MLB

[KAQ23] Mohammadreza Khani and Ikhlas Abdel-Qader. Machine learning based asynchronous computational framework for generalized Kalman filter. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kara:2022:DLF

[Kar22] Ahmet Kara. A deep learning framework with convolutional long short-term memory for influenza-like illness trend estimation. Concurrency and Computation: Practice and Experience, 34(17):e6988:1–e6988:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2023:MHA

[KAR^{+}23] Ashwani Kumar, Ashutosh Aggarwal, Anu Rani, Jatin Bedi, and Ravinder Kumar. A mat-heuristics approach for electric vehicle route optimization under multiple recharging options and time-of-use energy prices. Concurrency and Computation: Practice and Experience, 35(27):e7854:1–e7854:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Katarya:2023:TST

[Kat23] Rahul Katarya. Towards the significance of taxi recommender systems in smart cities. Concurrency and Computation: Practice and Experience, 35(2):e7475:1–e7475:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaya:2022:AGP

[Kay22a] Duygu Kaya. Automated gender–Parkinson’s disease detection at the same time via a hybrid deep model using human voice. Concurrency and Computation: Practice and Experience, 34(26):e7289:1–e7289:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kayhan:2022:CPD

[Kay22b] Gökhan Kayhan. Comparison of the performance of different learning algorithms in leaf feature extraction and recognition using convolution neural network. Concurrency and Computation: Practice and Experience, 34(26):e7294:1–e7294:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karthikeyan:2021:EAS

[KB21] Ramamoorthy Karthikeyan and Venkatachalam Balamurugan. Energy-aware and SLA-guaranteed optimal virtual machine swap and migrate system in cloud-Internet of Things. Concurrency and Computation: Practice and Experience, 33 (10):e6171:1–e6171:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karcioglu:2022:QFH

[KB22] Abdullah Ammar Karcioglu and Hasan Bulut. q-frame hash comparison based exact string matching algorithms for DNA sequences. Concurrency and Computation: Practice and Experience, 34(9):e6505:1–e6505:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kasri:2021:WNL

[KBBH21] Mohammed Kasri, Marouane Birjali, and Abderrahim Beni-Hssane. Word2Sent: a new learning sentiment-embedding model with low dimension for sentence level sentiment classification. Concurrency and Computation: Practice and Experience, 33(9):e6149:1–e6149:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2021:DDE

[KBJ21] Yunhu Kim, Khac-Hoai Nam Bui, and Jason J. Jung. Datadriven exploratory approach on player valuation in football transfer market. Concurrency and Computation: Practice and Experience, 33(3):e5353:1–e5353:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaushik:2022:NML

[KBK^{+}22] Keshav Kaushik, Akashdeep Bhardwaj, Manoj Kumar, Sachin Kumar Gupta, and Abhishek Gupta. A novel machine learning-based framework for detecting fake Instagram profiles. Concurrency and Computation: Practice and Experience, 34(28):e7349:1–e7349:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ke:2021:CSD

[KBL^{+}21] Wende Ke, Yan Bai, Huazhong Li, Ke Chen, and Quande Yuan. Control of stepping downstairs for humanoid robot based on dynamic multi-objective optimization. Concurrency and Computation: Practice and Experience, 33(5):e5999:1– e5999:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaur:2024:BHO

[KBR24] Prabhjot Kaur, Shalini Batra, and Prashant Singh Rana. Border-hunting optimization for mobile agent-based intrusion detection with deep convolutional neural network. Concurrency and Computation: Practice and Experience, 36(1): e7876:1–e7876:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumari:2022:SLP

[KBS^{+}22] Anisha Kumari, Ranjan Kumar Behera, Kshira Sagar Sahoo, Anand Nayyar, Ashish Kumar Luhach, and Satya Prakash Sahoo. Supervised link prediction using structured-based feature extraction in social network. Concurrency and Computation: Practice and Experience, 34(13):e5839:1–e5839:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kandasamy:2020:QEU

[KC20] Saravanakumar Kandasamy and Aswani Kumar Cherukuri. Query expansion using named entity disambiguation for a question-answering system. Concurrency and Computation: Practice and Experience, 32(4):e5119:1–e5119:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kavaklioglu:2022:SGB

[KC22] Can Kavaklioglu and Mucahit Cevik. Scalable grid-based approximation algorithms for partially observable Markov decision processes. Concurrency and Computation: Practice and Experience, 34(5):e6743:1–e6743:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krawczyk:2020:PCS

[KCL^{+}20] Rafal Krawczyk, Tomasz Czarski, Pawel Linczuk, Andrzej Wojenski, Maryna Chernyshova, Krzysztof Pozniak, Didier Mazon, Piotr Kolasinski, Grzegorz Kasprowicz, Wojciech Zabolotny, Michal Gaska, Ewa Kowalsaka-Strzeciwilk, Karol Malinowski, Axel Jardin, and Philippe Malard. Parallel computing in soft X-rays plasma diagnostic systems for thermal fusion reactors — feasibility studies for GPUs. Concurrency and Computation: Practice and Experience, 32(10):e5235:1– e5235:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumari:2022:AIB

[KCM^{+}22] Sushma Kumari, Monika Choudhary, Richa Mishra, Swades Kumar Chaulya, Girendra Mohan Prasad, Sujit Kumar Mandal, and Gautam Banerjee. Artificial intelligent based smart system for safe mining during foggy weather. Concurrency and Computation: Practice and Experience, 34(4):e6631:1– e6631:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ke:2022:AAM

[KCP^{+}22] Wende Ke, Yuhan Chen, Yijie Pang, Qingfeng Li, and Dongxin Lu. An activate appearance model-based algorithm for ear characteristic points positioning. Concurrency and Computation: Practice and Experience, 34(27):e7315:1– e7315:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kota:2023:IHH

[KCP23] Prabhakar N. Kota, Ashok S. Chandak, and B. P. Patil. IOT-HML: a hybrid machine learning technique for IoT enabled industrial monitoring and control system. Concurrency and Computation: Practice and Experience, 35(3):e7458:1– e7458:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kokangul:2022:CPG

[KD22] Ali Kokangul and Cansu Dagsuyu. Catheter planning with goal programming method in neonatal intensive care unit. Concurrency and Computation: Practice and Experience, 34 (1):e6492:1–e6492:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2022:EID

[KDA^{+}22] Chandra Umakantham Om Kumar, Jeyakumar Durairaj, Samsu Aliar Ahamed Ali, Y. Justindhas, and Suguna Marappan. Effective intrusion detection system for IoT using optimized capsule auto encoder model. Concurrency and Computation: Practice and Experience, 34(13):e6918:1–e6918:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karuppannan:2022:HAR

[KDC22] Kiruba Karuppannan, Shiloah Elizabeth Darmanayagam, and Sunil Retmin Raj Cyril. Human action recognition using fusion-based discriminative features and long short term memory classification. Concurrency and Computation: Practice and Experience, 34(25):e7250:1–e7250:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kummerlander:2023:IPD

[KDFK23] Adrian Kummerländer, Márcio Dorn, Martin Frank, and Mathias J. Krause. Implicit propagation of directly addressed grids in lattice Boltzmann methods. Concurrency and Computation: Practice and Experience, 35(8):e7509:1–e7509:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

K:2020:DBQ

[KDL20] Vivekananda Bhat K, Ashok Kumar Das, and Jong-Hyouk Lee. Design of a blind quantization-based audio watermarking scheme using singular value decomposition. Concurrency and Computation: Practice and Experience, 32(18):e5253:1– e5253:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kabbinale:2020:BES

[KDS^{+}20] Aniruddh Rao Kabbinale, Emmanouil Dimogerontakis, Mennan Selimi, Anwaar Ali, Leandro Navarro, Arjuna Sathiaseelan, and Jon Crowcroft. Blockchain for economically sustainable wireless mesh networks. Concurrency and Computation: Practice and Experience, 32(12):e5349:1–e5349:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2022:MHC

[KDS22] Parveen Kumar, Pawan Kumar Dahiya, and Bijay Kumar Singh. Modularized hypergraph clustering scheme model for stable VANET. Concurrency and Computation: Practice and Experience, 34(26):e7283:1–e7283:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2023:EPA

[KDS^{+}23] Mohit Kumar, Kalka Dubey, Samayveer Singh, Jitendra Kumar Samriya, and Sukhpal Singh Gill. Experimental performance analysis of cloud resource allocation framework using spider monkey optimization algorithm. Concurrency and Computation: Practice and Experience, 35(2):e7469:1– e7469:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kabakus:2021:NHT

[KE21] Abdullah Talha Kabakus and Pakize Erdogmus. A novel handwritten Turkish letter recognition model based on convolutional neural network. Concurrency and Computation: Practice and Experience, 33(21):e6429:1–e6429:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kabakus:2022:ECW

[KE22] Abdullah Talha Kabakus and Pakize Erdogmus. An experimental comparison of the widely used pre-trained deep neural networks for image classification tasks towards revealing the promise of transfer-learning. Concurrency and Computation: Practice and Experience, 34(24):e7216:1–e7216:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

K:2020:OFL

[KEK^{+}20] Shankar K, Mohamed Elhoseny, Lakshmanaprabu S K, Ilayaraja M, Vidhyavathi Rm, Mohamed A. Elsoud, and Majid Alkhambashi. Optimal feature level fusion based ANFIS classifier for brain MRI image classification. Concurrency and Computation: Practice and Experience, 32(1):e4887:1– e4887:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khosravian:2022:SFC

[KEMZ22] Pouya Khosravian, Sima Emadi, Ghasem Mirjalily, and Behzad Zamani. Service function chain composition and placement using grammar-based genetic algorithm. Concurrency and Computation: Practice and Experience, 34(3):e6587:1– e6587:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kermia:2022:SPF

[Ker22] Omar Kermia. Strictly periodic first: an optimal variant of LLF for scheduling tasks in a time-critical cyber-physical system. Concurrency and Computation: Practice and Experience, 34(7):e5908:1–e5908:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Keskes:2022:HPO

[KFKD22] Nesrine Keskes, Sameh Fakhfakh, Olfa Kanoun, and Nabil Derbel. High performance oversampling technique considering intra-class and inter-class distances. Concurrency and Computation: Practice and Experience, 34(6):e6753:1–e6753:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khaleghzadeh:2020:NDP

[KFML20] Hamidreza Khaleghzadeh, Muhammad Fahad, Ravi Reddy Manumachu, and Alexey Lastovetsky. A novel data partitioning algorithm for dynamic energy optimization on heterogeneous high-performance computing platforms. Concurrency and Computation: Practice and Experience, 32(21):e5928:1– e5928:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kour:2022:PLD

[KG22] Harnain Kour and Manoj Kumar Gupta. Predicting the language of depression from multivariate Twitter data using a feature-rich hybrid deep learning model. Concurrency and Computation: Practice and Experience, 34(24):e7224:1– e7224:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumaraswamy:2023:ISH

[KG23] Balachandra Kumaraswamy and Poonacha P G. An improved sub-harmonic to harmonic ratio method for pitch estimation and Shadja detection. Concurrency and Computation: Practice and Experience, 35(7):e7604:1–e7604:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Komatsu:2020:XCT

[KGE^{+}20] Kazuhiko Komatsu, Ayumu Gomi, Ryusuke Egawa, Daisuke Takahashi, Reiji Suda, and Hiroyuki Takizawa. Xevolver: a code transformation framework for separation of systemawareness from application codes. Concurrency and Computation: Practice and Experience, 32(7):e5577:1–e5577:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kheradmand:2022:CBR

[KGGM22] Behbod Kheradmand, Ali Ghaffari, Farhad Soleimanian Gharehchopogh, and Mohammad Masdari. Clustering-based routing protocol using gray wolf optimization and technique for order of preference by similarity to ideal solution algorithms in the vehicular ad hoc networks. Concurrency and Computation: Practice and Experience, 34(23):e7209:1– e7209:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kadi:2022:COU

[KGK22] Ayse Divit Kadi, Ertan Gokalp, and Fatih Kadi. Creating orthophotos with unmanned aerial vehicles and examining its accuracy and usability in geodetic applications. Concurrency and Computation: Practice and Experience, 34(21):e7132:1– e7132:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaur:2023:RUH

[KGM23] Gaganpreet Kaur, Raman Kumar Goyal, and Rajesh Mehta. Reducing unnecessary handovers and improving ranking abnormality based on multi-attribute decision making graph theory and matrix approach with Euclidean distance in heterogeneous wireless networks. Concurrency and Computation: Practice and Experience, 35(22):e7715:1–e7715:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Katyal:2024:FDT

[KGRG24] Siddharth Katyal, Shashank Gupta, Oshin Rawlley, and Debjani Ghosh. A fog-driven three-factor authentication protocol for secure data sharing in Internet of Vehicles cyber-physical systems. Concurrency and Computation: Practice and Experience, 36(8):e7981:1–e7981:??, April 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kamburugamuve:2020:TDB

[KGW^{+}20] Supun Kamburugamuve, Kannan Govindarajan, Pulasthi Wickramasinghe, Vibhatha Abeykoon, and Geoffrey Fox. Twister2: Design of a big data toolkit. Concurrency and Computation: Practice and Experience, 32(3):e5189:1– e5189:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

K:2022:EIT

[KH22] Geetha K and Brahmananda S H. Effective Internet of Things botnet classification by data upsampling using generative adversarial network and scale fused bidirectional long short term memory attention model. Concurrency and Computation: Practice and Experience, 34(28):e7380:1–e7380:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khajehzadeh:2022:CSP

[Kha22a] Mohammad Khajehzadeh. Compressive strength prediction of structural lightweight concrete by applying the recurrent neural network-fractional order bat optimization algorithm. Concurrency and Computation: Practice and Experience, 34(15): e6975:1–e6975:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kong:2022:NIV

[KHA22b] Yuqiang Kong, Yaoping He, and Karlo Abnoosian. Natureinspired virtual machine placement mechanisms: a systematic review. Concurrency and Computation: Practice and Experience, 34(11):e6900:1–e6900:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kamal:2022:FCE

[KHEF22] Randa Kamal, Ezz El-Din Hemdan, and Nawal El-Fishway. Forensics chain for evidence preservation system: an evidence preservation forensics framework for internet of things-based smart city security using blockchain. Concurrency and Computation: Practice and Experience, 34(21):e7062:1–e7062:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2021:CAP

[KHHK21] Joohyun Kim, Seohee Hong, Sengphil Hong, and Jaehoon Kim. Context-aware pub/sub control method using reinforcement learning. Concurrency and Computation: Practice and Experience, 34(1):e5727:1–e5727:??, August 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khan:2023:RPU

[KHK^{+}23] Muhammad Mubashir Khan, Muhammad Faraz Hyder, Shariq Mahmood Khan, Junaid Arshad, and Muhammad M. Khan. Ransomware prevention using moving target defense based approach. Concurrency and Computation: Practice and Experience, 35(7):e7592:1–e7592:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kobayashi:2024:APT

[KHK^{+}24] Masato Kobayashi, Yusuke Hirota, Shuhei Kudo, Takeo Hoshi, and Yusaku Yamamoto. Automatic performance tuning using the ATMathCoreLib tool: Two experimental studies related to dense symmetric eigensolvers. Concurrency and Computation: Practice and Experience, 36(10):e7849:1– e7849:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2020:SCT

[KHPH20] Hyunjun Kim, Sungin Hong, Jeonghwan Park, and Hwansoo Han. Static code transformations for thread-dense memory accesses in GPU computing. Concurrency and Computation: Practice and Experience, 32(5):e5512:1–e5512:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Koochaksaraei:2022:BDA

[KHR22] Mohammad Hossein Ghasemian Koochaksaraei, Abolfazl Toroghi Haghighat, and Mohammad Hossein Rezvani. A bartering double auction resource allocation model in cloud environments. Concurrency and Computation: Practice and Experience, 34(19):e7024:1–e7024:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kongshavn:2020:MDU

[KHY^{+}20] Madeleine Kongshavn, Hårek Haugerud, Anis Yazidi, Torleiv Maseng, and Hugo Hammer. Mitigating DDoS using weightbased geographical clustering. Concurrency and Computation: Practice and Experience, 32(11):e5679:1–e5679:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khan:2022:ELD

[KIAA^{+}22] Adil H. Khan, Dayang NurFatimah Awang Iskandar, Jawad F. Al-Asad, Hiren Mewada, and Muhammad Abid Sherazi. Ensemble learning of deep learning and traditional machine learning approaches for skin lesion segmentation and classification. Concurrency and Computation: Practice and Experience, 34(13):e6907:1–e6907:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2021:AEL

[Kim21a] Pyoung Won Kim. Assessing engagement levels in a nonface-to-face learning environment through facial expression analysis. Concurrency and Computation: Practice and Experience, 33(22):e6182:1–e6182:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2021:BSP

[Kim21b] Pyoung Won Kim. Bio-signal-processing-based convolutional neural networks model for music program scene editing. Concurrency and Computation: Practice and Experience, 33(22): e6257:1–e6257:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kagawa:2023:HTF

[KIN^{+}23] Hiroshi Kagawa, Yasuaki Ito, Koji Nakano, Ryota Yasudo, Yuya Kawamata, Ryota Katsuki, Yusuke Tabata, Takashi Yazane, and Kenichiro Hamano. High-throughput FPGA implementation for quadratic unconstrained binary optimization. Concurrency and Computation: Practice and Experience, 35(14):e6565:1–e6565:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kayesh:2022:DLM

[KIW^{+}22] Humayun Kayesh, Md. Saiful Islam, Junhu Wang, A. S. M. Kayes, and Paul A. Watters. A deep learning model for mining and detecting causally related events in tweets. Concurrency and Computation: Practice and Experience, 34(2):e5938:1– e5938:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2020:SLR

[KJ20] Akshi Kumar and Arunima Jaiswal. Systematic literature review of sentiment analysis on Twitter using soft computing techniques. Concurrency and Computation: Practice and Experience, 32(1):e5107:1–e5107:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kolodziej:2021:SMC

[KJ21] Joanna Kołodziej and Martin Gilje Jaatun. Secure mobile cloud computing. Concurrency and Computation: Practice and Experience, 33(18):e6528:1–e6528:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2021:KNS

[KJHM21] Bongjae Kim, Jinman Jung, Junyoung Heo, and Hong Min. Key node selection based on a genetic algorithm for fast patching in social networks. Concurrency and Computation: Practice and Experience, 33(2):e5194:1–e5194:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ksouri:2022:TGI

[KJMB22] Chahrazed Ksouri, Imen Jemili, Mohamed Mosbah, and Abdelfettah Belghith. Towards general Internet of Vehicles networking: Routing protocols survey. Concurrency and Computation: Practice and Experience, 34(7):e5994:1–e5994:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2020:ARB

[KJS^{+}20] HyoJoon Kim, SangHui Jeong, JiHyeon Seo, InSeok Park, Hoon Ko, and Seong Yong Moon. Augmented reality for botulinum toxin injection. Concurrency and Computation: Practice and Experience, 32(18):e5526:1–e5526:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kamei:2021:SSD

[KK21a] Sayaka Kamei and Hirotsugu Kakugawa. A self-stabilizing distributed algorithm for the local (1, |N_{i}|)-critical section problem. Concurrency and Computation: Practice and Experience, 33(12):e5628:1–e5628:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kasiselvanathan:2021:BER

[KK21b] M. Kasiselvanathan and N. Sathish Kumar. Bit error rate aware accurate prediction of original signals with low distortion using low complexity detection algorithms. Concurrency and Computation: Practice and Experience, 33(3):e5192:1– e5192:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kashyap:2022:LBT

[KK22a] Vijaita Kashyap and Ashok Kumar. Load balancing techniques for fog computing environment: Comparison, taxonomy, open issues, and challenges. Concurrency and Computation: Practice and Experience, 34(23):e7183:1–e7183:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaur:2022:UEU

[KK22b] Jaspreet Kaur and Prabhpreet Kaur. UNIConv: an enhanced U-Net based InceptionV3 convolutional model for DR semantic segmentation in retinal fundus images. Concurrency and Computation: Practice and Experience, 34(21):e7138:1– e7138:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kosunalp:2022:ITP

[KK22c] Selahattin Kosunalp and Yunus Kaya. IoT-TDMA: a performance evaluation of TDMA scheme for wireless sensor networks with Internet of Things. Concurrency and Computation: Practice and Experience, 34(21):e7063:1–e7063:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kothamasu:2022:SAT

[KK22d] Lakshmi Anusha Kothamasu and E. Kannan. Sentiment analysis on Twitter data based on spider monkey optimization and deep learning for future prediction of the brands. Concurrency and Computation: Practice and Experience, 34(21): e7104:1–e7104:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krishnamoorthy:2022:OSE

[KK22e] Ramesh Krishnamoorthy and Kalimuthu Krishnan. An optimum security enabled system-on-chip selection for Internet of Things applications using genetic algorithm modeling. Concurrency and Computation: Practice and Experience, 34(4): e6635:1–e6635:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2022:DBT

[KK22f] Aarthi Ramesh Kumar and Helenprabha Kuttiappan. Detection of brain tumor size using modified deep learning and multilevel thresholding utilizing modified dragonfly optimization algorithm. Concurrency and Computation: Practice and Experience, 34(18):e7016:1–e7016:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khatibi:2023:UDV

[KK23a] Arezoo Khatibi and Omid Khatibi. Using dynamic voltage frequency scaling and service-level agreement to reduce energy consumption in cloud datacenters based on distance. Concurrency and Computation: Practice and Experience, 35(8): e7631:1–e7631:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumari:2023:DLT

[KK23b] Monika Kumari and Ajay Kaul. Deep learning techniques for remote sensing image scene classification: a comprehensive review, current challenges, and future directions. Concurrency and Computation: Practice and Experience, 35(22):e7733:1– e7733:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kakugawa:2024:SSD

[KK24] Hirotsugu Kakugawa and Sayaka Kamei. A self-stabilizing distributed algorithm for the bounded lattice domination problems under the distance-2 model. Concurrency and Computation: Practice and Experience, 36(3):e7902:1–e7902:??, February 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karimi:2021:LBD

[KKAM21] Yaghoob Karimi, Mostafa Haghi Kashani, Mohammad Akbari, and Ebrahim Mahdipour. Leveraging big data in smart cities: a systematic review. Concurrency and Computation: Practice and Experience, 33(21):e6379:1–e6379:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khosla:2022:NME

[KKC22] Ashima Khosla, Padmavati Khandnor, and Trilok Chand. A novel method for EEG based automated eyes state classification using recurrence plots and machine learning approach. Concurrency and Computation: Practice and Experience, 34 (13):e6912:1–e6912:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khlif:2022:MTA

[KKE^{+}22] Ilhem Khlif, Mohamed Hadj Kacem, Cédric Eichler, Khalil Drira, and Ahmed Hadj Kacem. A model transformation approach for multiscale modeling of software architectures applied to smart cities. Concurrency and Computation: Practice and Experience, 34(7):e6298:1–e6298:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaur:2021:SRT

[KKK21] Navjeet Kaur, Ashok Kumar, and Rajesh Kumar. A systematic review on task scheduling in fog computing: Taxonomy, tools, challenges, and future directions. Concurrency and Computation: Practice and Experience, 33(21):e6432:1– e6432:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kannan:2023:DCP

[KKKS23] G. Kannan, Karunambiga K, P. J. Sathish Kumar, and Francis H. Shajin. Detection of COVID-19 patient based on attention segmental recurrent neural network (ASRNN) Archimedes optimization algorithm using ultra-low-dose CT images. Concurrency and Computation: Practice and Experience, 35(21):e7705:1–e7705:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kvet:2020:LAL

[KKM20] Michal Kvet, Emil Krsak, and Karol Matiasko. Locating and accessing large datasets using Flower Index Approach. Concurrency and Computation: Practice and Experience, 32(13): e5209:1–e5209:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2021:SSA

[KKM21] Yonghwan Kim, Yoshiaki Katayama, and Toshimitsu Masuzawa. A self-stabilizing algorithm for constructing a maximal (σ, τ)-directed acyclic mixed graph. Concurrency and Computation: Practice and Experience, 33(12):e5812:1– e5812:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2020:CNC

[KKP20] Dae-Young Kim, Seokhoon Kim, and Jong Hyuk Park. A combined network control approach for the edge cloud and LPWAN-based IoT services. Concurrency and Computation: Practice and Experience, 32(1):e4406:1–e4406:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaur:2024:KDS

[KKP24] Amandeep Kaur, C. Rama Krishna, and Nilesh Vishwasrao Patil. K-DDoS-SDN: a distributed DDoS attacks detection approach for protecting SDN environment. Concurrency and Computation: Practice and Experience, 36(3):e7912:1– e7912:??, February 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kanchanamala:2023:HOE

[KKR23] Pendela Kanchanamala, Ramesh Karnati, and Palagiri Vijaya Bhaskar Reddy. Hybrid optimization enabled deep learning and spark architecture using big data analytics for stock market forecasting. Concurrency and Computation: Practice and Experience, 35(8):e7618:1–e7618:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaur:2022:OOC

[KKS22] Akashdeep Kaur, Rajesh Kumar, and Sharad Saxena. OCTRA-5G: Osmotic computing based task scheduling and resource allocation framework for 5G. Concurrency and Computation: Practice and Experience, 34(28):e7369:1–e7369:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kricha:2023:RIW

[KKS23] Anis Kricha, Zied Kricha, and Anis Sakly. Robust image watermarking in DCT domain using optimal inter-block difference. Concurrency and Computation: Practice and Experience, 35(28):e7857:1–e7857:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kobori:2024:PEP

[KLH24] Kazuto Kobori, Chunghan Lee, and Toshio Hirotsu. Performance evaluation of portable time synchronization method using eBPF. Concurrency and Computation: Practice and Experience, 36(8):e7957:1–e7957:??, April 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kee:2021:IYB

[KLJ21] Kerk F. Kee, Bethanie Le, and Kulsawasd Jitkajornwanich. If you build it, promote it, and they trust you, then they will come: Diffusion strategies for science gateways and cyberinfrastructure adoption to harness big data in the science, technology, engineering, and mathematics (STEM) community. Concurrency and Computation: Practice and Experience, 33 (19):e6192:1–e6192:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2023:ARM

[KLK23] Ashok Kumar, Madan Lal, and Sumandeep Kaur. An autonomic resource management system for energy efficient and quality of service aware resource scheduling in cloud environment. Concurrency and Computation: Practice and Experience, 35(21):e7699:1–e7699:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kosta:2021:SIH

[KLL^{+}21] Sokol Kosta, Giuliano Laccetti, Marco Lapegna, Valeria Mele, and Raffaele Montella. Special issue on High-end Heterogeneous Architectures, Methodologies, and Algorithms (HHAMA20). Concurrency and Computation: Practice and Experience, 33(4):e6108:1–e6108:??, February 25, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kandasamy:2021:DFR

[KM21a] Nithya Kandasamy and Krishnamoorthi Murugasamy. Detecting and filtering rumor in social media using news media event. Concurrency and Computation: Practice and Experience, 33(20):e6329:1–e6329:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ko:2021:SMA

[KM21b] Hoon Ko and Goreti Marreiros. Smart media and application. Concurrency and Computation: Practice and Experience, 33 (2):e5491:1–e5491:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karthikeyan:2023:MPP

[KM23a] K. Karthikeyan and P. Madhavan. Multilayer privacypreserving model for security and privacy access control in edge computing. Concurrency and Computation: Practice and Experience, 35(10):e7677:1–e7677:??, May 1, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khanum:2023:LBB

[KM23b] Saba Khanum and Khurram Mustafa. A lattice-based blind ring signature scheme for sensitive data protection in blockchain applications. Concurrency and Computation: Practice and Experience, 35(26):e7835:1–e7835:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khanum:2023:SLR

[KM23c] Saba Khanum and Khurram Mustafa. A systematic literature review on sensitive data protection in blockchain applications. Concurrency and Computation: Practice and Experience, 35 (1):e7422:1–e7422:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kulagina:2023:MTS

[KMB23] Svetlana Kulagina, Henning Meyerhenke, and Anne Benoit. Mapping tree-shaped workflows on systems with different memory sizes and processor speeds. Concurrency and Computation: Practice and Experience, 35(25):e7842:1–e7842:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kirola:2023:OUN

[KMD23] Madhu Kirola, Minakshi Memoria, and Ankur Dumka. Optimized U-Net convolutional neural network based breast cancer prediction for accuracy increment in big data. Concurrency and Computation: Practice and Experience, 35(9): e7652:1–e7652:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Klarqvist:2021:ECP

[KML21] Marcus D. R. Klarqvist, Wojciech Muła, and Daniel Lemire. Efficient computation of positional population counts using SIMD instructions. Concurrency and Computation: Practice and Experience, 33(17):e6304:1–e6304:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khaleghzadeh:2023:EEA

[KML23] Hamidreza Khaleghzadeh, Ravi Reddy Manumachu, and Alexey Lastovetsky. Efficient exact algorithms for continuous bi-objective performance-energy optimization of applications with linear energy and monotonically increasing performance profiles on heterogeneous high performance computing platforms. Concurrency and Computation: Practice and Experience, 35(20):e7285:1–e7285:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kannan:2022:NCB

[KMR22] Sankara Nayaki Kannan, Sudheep Elayidom Mannathazhathu, and Rajesh Raghavan. A novel compression based community detection approach using hybrid honey badger African vulture optimization for online social networks. Concurrency and Computation: Practice and Experience, 34(23): e7205:1–e7205:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kommaraju:2020:SRL

[KMRR20] Ananda V. Kommaraju, Kristyn J. Maschhoff, Michael F. Ringenburg, and Benjamin Robbins. Scalable reinforcement learning on Cray XC. Concurrency and Computation: Practice and Experience, 32(20):e5636:1–e5636:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kallel:2021:RBF

[KMS^{+}21] Slim Kallel, Zakaria Maamar, Mohamed Sellami, Noura Faci, Ahmed Ben Arab, Walid Gaaloul, and Thar Baker. Restriction-based fragmentation of business processes over the cloud. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Korra:2022:IHS

[KMS^{+}22] Sampath Korra, Ravikanth Mamidi, Narasimha Reddy Soora, Kotte Vinay Kumar, and Naliganti Cornel Santosh Kumar. Intracranial hemorrhage subtype classification using learned fully connected separable convolutional network. Concurrency and Computation: Practice and Experience, 34(24):e7218:1– e7218:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kirti:2024:FTM

[KMY24] Medha Kirti, Ashish Kumar Maurya, and Rama Shankar Yadav. A fault-tolerant model for tuple space coordination in distributed environments. Concurrency and Computation: Practice and Experience, 36(1):e7884:1–e7884:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Koulouzis:2020:TCD

[KMZ^{+}20] Spiros Koulouzis, Paul Martin, Huan Zhou, Yang Hu, Junchao Wang, Thierry Carval, Baptiste Grenier, Jani Heikkinen, Cees de Laat, and Zhiming Zhao. Time-critical data management in clouds: Challenges and a dynamic real-time infrastructure planner (DRIP) solution. Concurrency and Computation: Practice and Experience, 32(16):e5269:1–e5269:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2022:HPI

[KNK22] Mantosh Kumar, Kumari Namrata, and Neha Kumari. Hyper-parametric improved machine learning models for solar radiation forecasting. Concurrency and Computation: Practice and Experience, 34(23):e7190:1–e7190:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karnan:2021:HMI

[KNM21] Hemalatha Karnan, Sivakumaran Natarajan, and Rajajeyakumar Manivel. Human machine interfacing technique for diagnosis of ventricular arrhythmia using supervisory machine learning algorithms. Concurrency and Computation: Practice and Experience, 33(4):e5001:1–e5001:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kara:2022:PNS

[KÖ22] Gökçehan Kara and Can Özturan. Parallel network simplex algorithm for the minimum cost flow problem. Concurrency and Computation: Practice and Experience, 34(4):e6659:1– e6659:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaciranlar:2022:CLT

[KÖÖG22] Selahattin Kaçiranlar, Nimet Özbay, Ecem Özkan, and Hüseyin Güler. Comparison of Liu and two parameter principal component estimator to combat multicollinearity. Concurrency and Computation: Practice and Experience, 34(5): e6737:1–e6737:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kothandaraman:2023:BSS

[Kot23] Vigneshwari Kothandaraman. Binary swallow swarm optimization with convolutional neural network brain tumor classifier for magnetic resonance imaging images. Concurrency and Computation: Practice and Experience, 35(10):e7661:1– e7661:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krishnaveni:2021:EAN

[KP21] S. Krishnaveni and S. Prabakaran. Ensemble approach for network threat detection and classification on cloud computing. Concurrency and Computation: Practice and Experience, 33(3):e5272:1–e5272:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kruger:2023:DEM

[KPA23] Kleber Kruger, Ricardo Pannain, and Rodolfo Azevedo. DONUTS: an efficient method for checkpointing in nonvolatile memories. Concurrency and Computation: Practice and Experience, 35(18):e7574:1–e7574:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khoshlessan:2020:PPM

[KPF^{+}20] Mahzad Khoshlessan, Ioannis Paraskevakos, Geoffrey C. Fox, Shantenu Jha, and Oliver Beckstein. Parallel performance of molecular dynamics trajectory analysis. Concurrency and Computation: Practice and Experience, 32(19):e5789:1– e5789:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kalantary:2021:DDS

[KPJ^{+}21] Saba Kalantary, Reza Pourbabaki, Ali Jahani, Mohsen Sadeghi Yarandi, Sajjad Samiei, and Reza Jahani. Development of a decision support system tool to predict the pulmonary function using artificial neural network approach. Concurrency and Computation: Practice and Experience, 33(16):e6258:1– e6258:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ko:2020:PIS

[KPM20] Hoon Ko, Sung Bum Pan, and Libor Mesícek. Personal identification study for touchable devices with ECG. Concurrency and Computation: Practice and Experience, 32(8):e5169:1– e5169:??, April 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2022:CEA

[KPP^{+}22] L. Sathish Kumar, A. V. Prabu, V. Pandimurugan, S. Rajasoundaran, Prince Priya Malla, and Sidheswar Routray.

A comparative experimental analysis and deep evaluation practices on human bone fracture detection using X-ray images. Concurrency and Computation: Practice and Experience, 34(26):e7307:1–e7307:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kazmi:2020:FMV

[KQK^{+}20] Syed Asad Raza Kazmi, Awais Qasim, Adnan Khalid, Ruttaba Assad, and Muhammad Shahbaz. Formal modeling and verification of cloud-based web service composition. Concurrency and Computation: Practice and Experience, 32(21): e5249:1–e5249:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaul:2022:SVM

[KR22a] Ajay Kaul and Sneha Raina. Support vector machine versus convolutional neural network for hyperspectral image classification: a systematic review. Concurrency and Computation: Practice and Experience, 34(15):e6945:1–e6945:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2022:CAP

[KR22b] Kadari Kishore Kumar and Husnabad Venkateswara Reddy. Crime activities prediction system in video surveillance by an optimized deep learning framework. Concurrency and Computation: Practice and Experience, 34(11):e6852:1–e6852:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2022:HMU

[KR22c] Ponna Mahesh Kumar and P. Srinivasa Rao. High Median Utility Itemset Mining for recovering streaming window transaction using novel Modified Heap-based Optimization. Concurrency and Computation: Practice and Experience, 34(18): e7009:1–e7009:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2023:EEP

[KR23] G. Harish Kumar and P. Trinatha Rao. An energy efficiency perceptive on MIMO-OFDM systems using hybrid fruit fly-based Salp swarm optimization technique. Concurrency and Computation: Practice and Experience, 35(1): e7416:1–e7416:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kloh:2020:SJS

[KRB^{+}20] Henrique Kloh, Vinod E. F. Rebello, Cristina Boeres, Bruno Schulze, and Mariza Ferro. Static job scheduling for environments with vertical elasticity. Concurrency and Computation: Practice and Experience, 32(19):e5761:1–e5761:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Katebi:2022:RRA

[KRJS22] Matin Katebi, Afshin RezaKhani, Saba Joudaki, and Mohammad Ebrahim Shiri. RAPSAMS: Robust affinity propagation clustering on static Android malware stream. Concurrency and Computation: Practice and Experience, 34(15):e6980:1– e6980:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaur:2021:BBF

[KRK21] Jasleen Kaur, Rinkle Rani, and Nidhi Kalra. Blockchainbased framework for secured storage, sharing, and querying of electronic healthcare records. Concurrency and Computation: Practice and Experience, 33(20):e6369:1–e6369:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

K:2022:NFR

[KRKM22] Rajchandar K, Baskaran R, Padmanabhan Panchu K, and Rajmohan M. A novel fuzzy and reverse auction-based algorithm for task allocation with optimal path cost in multi-robot systems. Concurrency and Computation: Practice and Experience, 34(5):e6716:1–e6716:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krishnasamy:2023:ARB

[KRSR23] Srinivasan Krishnasamy, Murugasan Rajiah, Krishna Kumari SenthilKumar, and Shanker Nagalingam Rajendiran. Association rule-based multilevel regression pricing and artificial neural networks based land selling price prediction based on market value. Concurrency and Computation: Practice and Experience, 35(5):e7550:1–e7550:??, February 28, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kavitha:2020:PAA

[KS20] Kadarla Kavitha and S. C. Sharma. Performance analysis of ACO-based improved virtual machine allocation in cloud for IoT-enabled healthcare. Concurrency and Computation: Practice and Experience, 32(21):e5613:1–e5613:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kalyanam:2021:ESG

[KS21a] Rajesh Kalyanam and Vlado Stankovski. Editorial: Science gateways special issue 2020. Concurrency and Computation: Practice and Experience, 33(19):e6335:1–e6335:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kozicky:2021:JDT

[KS21b] Claudio Kozický and Ivan Simecek. Joint direct and transposed sparse matrix-vector multiplication for multithreaded CPUs. Concurrency and Computation: Practice and Experience, 33(13):e6236:1–e6236:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karat:2022:ORA

[KS22a] Chitharanjan Karat and Radha Senthilkumar. Optimal resource allocation with deep reinforcement learning and greedy adaptive firefly algorithm in cloud computing. Concurrency and Computation: Practice and Experience, 34(4):e6657:1– e6657:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khullar:2022:PPI

[KS22b] Vikas Khullar and Harjit Pal Singh. Privacy protected internet of unmanned aerial vehicles for disastrous site identification. Concurrency and Computation: Practice and Experience, 34(19):e7040:1–e7040:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2022:LSF

[KS22c] K. Suresh Kumar and C. Helen Sulochana. Local search five-element cycle optimized reLU-BiLSTM for multilingual aspect-based text classification. Concurrency and Computation: Practice and Experience, 34(28):e7374:1–e7374:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2022:HOQ

[KS22d] Ram Kumar and S. C. Sharma. Hybrid optimized query expansion strategy for semantic information retrieval using spatial bound whale and binary moth flame optimization algorithm. Concurrency and Computation: Practice and Experience, 34(27):e7320:1–e7320:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaur:2023:ROT

[KS23a] Palwinder Kaur and Rajesh Kumar Singh. A review on optimization techniques for medical image analysis. Concurrency and Computation: Practice and Experience, 35(1):e7443:1– e7443:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2023:RLN

[KS23b] P. Pramod Kumar and K. Sagar. Reinforcement learning and neuro-fuzzy GNN-based vertical handover decision on internet of vehicles. Concurrency and Computation: Practice and Experience, 35(12):e7688:1–e7688:??, May 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kuppusamy:2023:NHD

[KS23c] Mouthami Kuppusamy and Anandamurugan Selvaraj. A novel hybrid deep learning model for aspect based sentiment analysis. Concurrency and Computation: Practice and Experience, 35(4):e7538:1–e7538:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2024:BFE

[KS24] Mandeep Kumar and Amritpal Singh. Bloom filter empowered smart storage/access in IoMT [edge–fog–cloud] hierarchy for health-care data ingestion. Concurrency and Computation: Practice and Experience, 36(11):e8012:1–e8012:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kruse:2021:PSS

[KSA^{+}21] Carola Kruse, Masha Sosonkina, Mario Arioli, Nicolas Tardieu, and Ulrich Rüde. Parallel solution of saddle point systems with nested iterative solvers based on the Golub– Kahan bidiagonalization. Concurrency and Computation: Practice and Experience, 33(11):e5914:1–e5914:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khan:2022:GCA

[KSA22] Muhammad Milhan Afzal Khan, Hafiz Muhammad Azeem Sarwar, and Muhammad Awais. Gas consumption analysis of Ethereum blockchain transactions. Concurrency and Computation: Practice and Experience, 34(4):e6679:1–e6679:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumari:2023:WAA

[KSB23] Anisha Kumari, Bibhudatta Sahoo, and Ranjan Kumar Behera. Workflow aware analytical model to predict performance and cost of serverless execution. Concurrency and Computation: Practice and Experience, 35(22):e7743:1– e7743:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kwon:2021:PFF

[KSCL21] Soonhong Kwon, Sang-Jin Son, Yangseo Choi, and Jong-Hyouk Lee. Protocol fuzzing to find security vulnerabilities of RabbitMQ. Concurrency and Computation: Practice and Experience, 33(23):e6012:1–e6012:??, December 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2022:WWR

[KSD22] Sudershan Kumar, Prabuddha Sinha, and Shirshendu Das. WinDRAM: Weak rows as in-DRAM cache. Concurrency and Computation: Practice and Experience, 34(28):e7350:1– e7350:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaur:2023:SSG

[KSJ23a] Jatinder Kaur, Ashutosh Kumar Singh, and Neeru Jindal. Systematic survey on generative adversarial networks for brain tumor segmentation and classification. Concurrency and Computation: Practice and Experience, 35(27):e7850:1– e7850:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2023:SPF

[KSJ23b] Gourav Kumar, Uday Pratap Singh, and Sanjeev Jain. Stock price forecasting based on the relationship among Asian stock markets using deep learning. Concurrency and Computation: Practice and Experience, 35(28):e7864:1–e7864:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kang:2021:EPA

[KSJK21] Kyeung Ho Kang, Si Ho Shin, Jaehyo Jung, and Youn Tae Kim. Estimation of a physical activity energy expenditure with a patch-type sensor module using artificial neural network. Concurrency and Computation: Practice and Experience, 33(2):e5455:1–e5455:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2020:SBM

[KSK^{+}20] Chandan Kumar, Ak Singh, P Kumar, Rajiv Singh, and Siddharth Singh. SPIHT-based multiple image watermarking in NSCT domain. Concurrency and Computation: Practice and Experience, 32(1):e4912:1–e4912:??, January 10, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khan:2022:CHA

[KSK^{+}22] Amjad Rehman Khan, Tanzila Saba, Muhammad Zeeshan Khan, Suliman Mohamed Fati, and Muhammad Usman Ghani Khan. Classification of human’s activities from gesture recognition in live videos using deep learning. Concurrency and Computation: Practice and Experience, 34(10): e6825:1–e6825:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kushwaha:2023:HAR

[KSK23] Arati Kushwaha, Prashant Srivastava, and Ashish Khare. Human activity recognition based on integration of multilayer information of convolutional neural network architecture. Concurrency and Computation: Practice and Experience, 35(5): e7571:1–e7571:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2022:TAL

[KSKS22] Manoj Kumar, Sukhwinder Singh Sran, Lakhwinder Kaur, and Jagpreet Singh. Thermal aware learning based CPU governor. Concurrency and Computation: Practice and Experience, 34(11):e6862:1–e6862:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2021:NVW

[KSP21] Pala Mahesh Kumar, Kalyanapu Srinivas, and Bitla Prabhakar. A new video watermarking using redundant discrete wavelet in singular value decomposition domain with multiobjective optimization. Concurrency and Computation: Practice and Experience, 33(13):e6217:1–e6217:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

K:2021:PSO

[KSS21] Pravin Kumar S. K., M. G. Sumithra, and N. Saranya. Particle swarm optimization (PSO) with fuzzy c means (PSO-FCM)-based segmentation and machine learning classifier for leaf diseases prediction. Concurrency and Computation: Practice and Experience, 33(3):e5312:1–e5312:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2024:IIV

[KSS24a] Pankaj Kumar, Anurag Singh, and Ajay K Sharma. Identification of influential vertices in complex networks: a hitting time-based approach. Concurrency and Computation: Practice and Experience, 36(11):e8031:1–e8031:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumari:2024:MCM

[KSS24b] Twinkle Kumari, Damanpreet Singh, and Birmohan Singh. Multi-chaotic maps and blockchain based image encryption. Concurrency and Computation: Practice and Experience, 36 (14):e8092:1–e8092:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumari:2022:EAA

[KSSK22] Monika Kumari, Gadadhar Sahoo, Kishore Kumar Senapati, and Gaurav Kumar. Energy-aware autoscaling for scientific workflow in cloud environment. Concurrency and Computation: Practice and Experience, 34(3):e6576:1–e6576:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krishnaveni:2022:NID

[KSSP22] Sivamohan Krishnaveni, Sivanandam Sivamohan, Subramanian Sridhar, and Subramani Prabhakaran. Network intrusion detection based on ensemble classification and feature selection method for cloud computing. Concurrency and Computation: Practice and Experience, 34(11):e6838:1–e6838:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2023:HMF

[KST23] Deepak Kumar, Kartik Sahoo, and Manoj Thakur. A hybrid multicategory framework for generating automated trading systems. Concurrency and Computation: Practice and Experience, 35(22):e7746:1–e7746:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2021:IDS

[KSTV21] K. Pradeep Mohan Kumar, M. Saravanan, M. Thenmozhi, and K. Vijayakumar. Intrusion detection system based on GA-fuzzy classifier for detecting malicious attacks. Concurrency and Computation: Practice and Experience, 33(3): e5242:1–e5242:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2022:BAG

[KSVP22] S. Suresh Kumar, R. Sarankumar, O. Vignesh, and A. Prakash. Bilateral anisotropic Gabor wavelet transformation based deep stacked auto encoding for lossesless image compression. Concurrency and Computation: Practice and Experience, 34(28):e7383:1–e7383:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaya:2023:SIH

[KSY23] Kamer Kaya, Cevat Sener, and Hüsnü Yenigün. Special issue on high-performance computing conference (BASARIM 2022). Concurrency and Computation: Practice and Experience, 35(24):e7926:1–e7926:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2024:EIK

[KSY24] Anil Kumar, Dinesh Singh, and Rama Shankar Yadav. Entropy and improved k-nearest neighbor search based undersampling (ENU) method to handle class overlap in imbalanced datasets. Concurrency and Computation: Practice and Experience, 36(2):e7894:1–e7894:??, January 25, 2024. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kalaiselvi:2020:GES

[KT20] B. Kalaiselvi and M. Thangamani. Gibbs entropy simulated annealing based Edman firefly optimization for big data protein sequencing. Concurrency and Computation: Practice and Experience, 32(4):e5056:1–e5056:??, February 25, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kannan:2022:EBT

[KT22a] Durkadevi Kannan and Revathi Thiyagarajan. Entropy based TOPSIS method for controller selection in software defined networking. Concurrency and Computation: Practice and Experience, 34(1):e6499:1–e6499:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Krishnan:2022:MOT

[KT22b] Malathy Navaneetha Krishnan and Revathi Thiyagarajan. Multi-objective task scheduling in fog computing using improved gaining sharing knowledge based algorithm. Concurrency and Computation: Practice and Experience, 34(24): e7227:1–e7227:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karakati:2023:SCR

[KT23] Chitti Babu Karakati and Sethukarasi Thirumaaran. Software code refactoring based on deep neural network-based fitness function. Concurrency and Computation: Practice and Experience, 35(4):e7531:1–e7531:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kvet:2020:CTD

[KTK20] Michal Kvet, Stefan Toth, and Emil Krsak. Concept of temporal data retrieval: Undefined value management. Concurrency and Computation: Practice and Experience, 32(13):e5399:1– e5399:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

K:2022:ARF

[KTM22] Ananthajothi K, Karthick T, and Amanullah M. Automated rain fall prediction enabled by optimized convolutional neural network-based feature formation with adaptive long shortterm memory framework. Concurrency and Computation: Practice and Experience, 34(11):e6868:1–e6868:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khan:2021:SAB

[KTU^{+}21] Fazal Qudus Khan, Georgios Tsaramirsis, Naimat Ullah, Mohamed Nazmudeen, Sadeeq Jan, and Awais Ahmad. Smart algorithmic based web crawling and scraping with template autoupdate capabilities. Concurrency and Computation: Practice and Experience, 33(22):e6042:1–e6042:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

K:2022:BDB

[KUK22] Srikanth K, S. Zahoor Ul Huq, and A. P. Siva Kumar. Big data based analytic model to predict and classify breast cancer using improved fractional rough fuzzy K-means clustering and labeled ensemble classifier algorithm. Concurrency and Computation: Practice and Experience, 34(10):e6715:1– e6715:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khan:2024:CEU

[KUK^{+}24] Javed Ali Khan, Tahir Ullah, Arif Ali Khan, Affan Yasin, Muhammad Azeem Akbar, and Khursheed Aurangzeb. Can end-user feedback in social media be trusted for software evolution: Exploring and analyzing fake reviews. Concurrency and Computation: Practice and Experience, 36(10):e7990:1– e7990:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:2021:SSD

[Kum21] Manish Kumar. Solid state drive forensics analysis-challenges and recommendations. Concurrency and Computation: Practice and Experience, 33(24):e6442:1–e6442:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumaraswamy:2022:OCM

[Kum22] Balachandra Kumaraswamy. Optimal classification of music genres based on acoustic and visual features. Concurrency and Computation: Practice and Experience, 34(23):e7174:1– e7174:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kanimozhi:2022:OTF

[KV22a] P. Kanimozhi and T. Aruldoss Albert Victoire. Oppositional tunicate fuzzy c-means algorithm and logistic regression for intrusion detection on cloud. Concurrency and Computation: Practice and Experience, 34(4):e6624:1–e6624:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kavitha:2022:SHM

[KV22b] Ravichandran Kavitha and AngaThevar Valarmathi. A security hybrid mechanism using fuzzy with adaptive ElGamal approach in cloud. Concurrency and Computation: Practice and Experience, 34(26):e7276:1–e7276:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Keshri:2024:MBT

[KV24] Rashmi Keshri and Deo Prakash Vidyarthi. An ML-based task clustering and placement using hybrid Jaya–gray wolf optimization in fog–cloud ecosystem. Concurrency and Computation: Practice and Experience, 36(14):e8109:1–e8109:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Koens:2021:BAD

[KVP21] Tommy Koens, Pol Van Aubel, and Erik Poll. Blockchain adoption drivers: The rationality of irrational choices. Concurrency and Computation: Practice and Experience, 33(8): e5843:1–e5843:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kathiravelu:2020:INA

[KVV20] Pradeeban Kathiravelu, Peter Van Roy, and Luís Veiga. Interoperable and network-aware service workflows for big data executions at Internet scale. Concurrency and Computation: Practice and Experience, 32(21):e5212:1–e5212:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Korch:2021:DIM

[KW21] Matthias Korch and Tim Werner. An in-depth introduction of multi-workgroup tiling for improving the locality of explicit one-step methods for ODE systems with limited access distance on GPUs. Concurrency and Computation: Practice and Experience, 33(11):e6016:1–e6016:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kou:2024:ODP

[KWL^{+}24] Lei Kou, Junhe Wan, Hailin Liu, Wende Ke, Hui Li, Jie Chen, Zhen Yu, and Quande Yuan. Optimized design of patrol path for offshore wind farms based on genetic algorithm and particle swarm optimization with traveling salesman problem. Concurrency and Computation: Practice and Experience, 36 (2):e7907:1–e7907:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kong:2021:IDD

[KWZ^{+}21] Di Kong, Lili Wan, Zhizhuo Zhang, Wanru Xu, and Shenghui Wang. Interactive deformation-driven person silhouette image synthesis. Concurrency and Computation: Practice and Experience, 33(22):e5951:1–e5951:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kang:2021:TBW

[KXL^{+}21] Guosheng Kang, Yong Xiao, Jianxun Liu, Yingcheng Cao, Buqing Cao, Xiangping Zhang, and Linghang Ding. Tatt– BiLSTM: Web service classification with topical attentionbased BiLSTM. Concurrency and Computation: Practice and Experience, 33(16):e6287:1–e6287:??, August 25, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Koksal:2022:IAT

[KY22] Ömer Köksal and Eyüp Halit Yilmaz. Improving automated Turkish text classification with learning-based algorithms. Concurrency and Computation: Practice and Experience, 34 (11):e6874:1–e6874:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kaushal:2023:LSI

[KY23] Vaishali Kaushal and Rajan Yadav. Learning successful implementation of Chatbots in businesses from B2B customer experience perspective. Concurrency and Computation: Practice and Experience, 35(1):e7450:1–e7450:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Koikara:2021:PVT

[KYP21] Rosemary Koikara, Eun-Jun Yoon, and Anand Paul. Publicly verifiable threshold secret sharing based on threedimensional-cellular automata. Concurrency and Computation: Practice and Experience, 33(22):e6146:1–e6146:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kang:2020:TCM

[KYZ20] Guosheng Kang, Liqin Yang, and Liang Zhang. Toward configurable modeling for artifact-centric business processes. Concurrency and Computation: Practice and Experience, 32 (2):e5367:1–e5367:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lakshmanan:2022:EDS

[LA22] Karthikeyan Lakshmanan and Samydurai Arumugam. An efficient data science technique for IoT assisted healthcare monitoring system using cloud computing. Concurrency and Computation: Practice and Experience, 34(11):e6857:1–e6857:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Loganathan:2021:EEC

[LAC21] Sathyapriya Loganathan, Jawahar Arumugam, and Vinothkumar Chinnababu. An energy-efficient clustering algorithm with self-diagnosis data fault detection and prediction for wireless sensor networks. Concurrency and Computation: Practice and Experience, 33(17):e6288:1–e6288:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Laban:2022:TOA

[LAE^{+}22] Noureldin Laban, Bassam Abdellatif, Hala M. Ebeid, Howida A. Shedeed, and Mohamed F. Tolba. Toward object alphabet augmentation for object detection in very highresolution satellite images. Concurrency and Computation: Practice and Experience, 34(6):e6785:1–e6785:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:IAE

[LAH^{+}22] Haiqing Liu, Ming Ai, Rong Huang, Rixuan Qiu, and Yuancheng Li. Identity authentication for edge devices based on zero-trust architecture. Concurrency and Computation: Practice and Experience, 34(23):e7198:1–e7198:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lukman:2022:KEL

[LAK22] Adewale F. Lukman, Muhammad Amin, and B. M. Golam Kibria. K-L estimator for the linear mixed models: Computation and simulation. Concurrency and Computation: Practice and Experience, 34(6):e6780:1–e6780:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lukman:2021:KEI

[LAKA21] Adewale F. Lukman, Zakariya Y. Algamal, B. M. Golam Kibria, and Kayode Ayinde. The KL estimator for the inverse Gaussian regression model. Concurrency and Computation:

Practice and Experience, 33(13):e6222:1–e6222:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lopez-Albelda:2020:HCT

[LALMGLG20] B. López-Albelda, A. J. Lázaro-Muñoz, J. M. González-Linares, and Nicolas Guil. Heuristics for concurrent task scheduling on GPUs. Concurrency and Computation: Practice and Experience, 32(20):e5571:1–e5571:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lukman:2023:RBE

[LAP23] Adewale F. Lukman, Mohammad Arashi, and Vilmos Prokaj. Robust biased estimators for Poisson regression model: Simulation and applications. Concurrency and Computation: Practice and Experience, 35(7):e7594:1–e7594:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lund:2021:TAL

[LB21] Andreas Lund and Uwe Brinkschulte. Task-allocation in a large-scaled hierarchical many-core topology. Concurrency and Computation: Practice and Experience, 33(14):e5731:1– e5731:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Laiche:2023:QAT

[LBA23] Fatima Laiche, Asma Ben Letaifa, and Taoufik Aguili. QoEaware traffic monitoring based on user behavior in video streaming services. Concurrency and Computation: Practice and Experience, 35(11):e6678:1–e6678:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:FDD

[LBFT22] Wenhao Li, Niranjan Bidargaddi, John Fouyaxis, and Dan Thorpe. A framework for developing digital health applications with efficiency and reuse principles: DHLink case study. Concurrency and Computation: Practice and Experience, 34(27):e7322:1–e7322:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Larrea:2020:EPM

[LBG^{+}20] Verónica G. Vergara Larrea, Reuben D. Budiardja, Rahulkumar Gayatri, Christopher Daley, Oscar Hernandez, and Wayne Joubert. Experiences in porting mini-applications to OpenACC and OpenMP on heterogeneous systems. Concurrency and Computation: Practice and Experience, 32(20): e5780:1–e5780:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:PPP

[LBZ^{+}22] Fang (Cherry) Liu, Mehmet Belgin, Nuyun Zhang, Kevin Manalo, Ruben Lara, Christopher P. Stone, and Paul Manno. ProvBench: a performance provenance capturing framework for heterogeneous research computing environments. Concurrency and Computation: Practice and Experience, 34(10): e6820:1–e6820:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2020:IQA

[LC20] Li-Hui Lin and Tzong-Jer Chen. Image quality assessment by an efficient correlation-based metric. Concurrency and Computation: Practice and Experience, 32(19):e5794:1–e5794:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2021:SRD

[LC21] Lijun Lin and Zhiyun Chen. Social rumor detection based on multilayer transformer encoding blocks. Concurrency and Computation: Practice and Experience, 33(6):e6083:1– e6083:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liang:2022:ERS

[LCCT22] Ying Liang, Tianxu Cui, Yang Cao, and Hengliang Tang. An effective resource scheduling model for edge cloud oriented AIoT. Concurrency and Computation: Practice and Experience, 34(5):e6720:1–e6720:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2021:TMC

[LCKJ21] Chan-Gyu Lee, Joong-Yeon Cho, Jooho Kim, and Hyun-Wook Jin. Transparent many-core partitioning for highperformance big data I/O. Concurrency and Computation: Practice and Experience, 33(18):e6017:1–e6017:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2020:MIF

[LCL^{+}20] Feiqiang Liu, Lihui Chen, Lu Lu, Awais Ahmad, Gwanggil Jeon, and Xiaomin Yang. Medical image fusion method by using Laplacian pyramid and convolutional sparse representation. Concurrency and Computation: Practice and Experience, 32(17):e5632:1–e5632:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:KFG

[LCL^{+}22] Zijie Liu, Can Chen, Junjiang Li, Yi Cheng, Yingjie Kou, and Dengyin Zhang. KubFBS: a fine-grained and balanceaware scheduling system for deep learning tasks based on Kubernetes. Concurrency and Computation: Practice and Experience, 34(11):e6836:1–e6836:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2022:SLR

[LCM22] Gang Luo, Zhiyuan Chen, and Bayan Omar Mohammed. A systematic literature review of intrusion detection systems in the cloud-based IoT environments. Concurrency and Computation: Practice and Experience, 34(10):e6822:1–e6822:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Leach:2024:ASM

[LCR^{+}24] Kieran Leach, Philip Cass, Steven Robson, Eimantas Kazakevicius, Martin Lafferty, Andrew Turner, and Alan Simpson. Automated service monitoring in the deployment of ARCHER2. Concurrency and Computation: Practice and Experience, 36(2):e7892:1–e7892:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lazrag:2021:ESR

[LCSR21] Hilmi Lazrag, Abdellah Chehri, Rachid Saadane, and Moulay Driss Rahmani. Efficient and secure routing protocol based on blockchain approach for wireless sensor networks. Concurrency and Computation: Practice and Experience, 33(22):e6144:1–e6144:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2021:IMI

[LCW21] Feng Liu, Zhigang Chen, and Jie Wang. Intelligent medical IoT system based on WSN with computer vision platforms. Concurrency and Computation: Practice and Experience, 33 (12):e5036:1–e5036:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2023:MFA

[LCW^{+}23a] Zuhe Li, Yuhao Cui, Fengqin Wang, Weihua Liu, Yongshuang Yang, Zeqi Yu, Bin Jiang, and Hui Chen. A multimodal face antispoofing method based on multifeature vision transformer and multirank fusion. Concurrency and Computation: Practice and Experience, 35(23):e7824:1–e7824:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:EEO

[LCW23b] Dakun Liu, Guifen Chen, and Yijun Wang. Energy efficiency optimization routing decision scheme in three-dimensional wireless ad hoc network. Concurrency and Computation: Practice and Experience, 35(8):e7624:1–e7624:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:Ra

[LCW23c] Feng Liu, Zhigang Chen, and Jie Wang. Retraction. Concurrency and Computation: Practice and Experience, 33(12): e7749:1–e7749:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:LES

[LCZ^{+}20a] Qiliang Li, Jie Cui, Hong Zhong, Yichao Du, Yonglong Luo, and Lu Liu. LBBESA: an efficient software-defined networking load-balancing scheme based on elevator scheduling algorithm. Concurrency and Computation: Practice and Experience, 32(16):e5222:1–e5222:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2020:NID

[LCZ^{+}20b] Jun Luo, Senchun Chai, Baihai Zhang, Yuanqing Xia, Jianlei Gao, and Guoqiang Zeng. A novel intrusion detection method based on threshold modification using receiver operating characteristic curve. Concurrency and Computation: Practice and Experience, 32(14):e5690:1–e5690:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2020:SMM

[LCZY20] Huaiguang Liu, Yu Cai, Shiyang Zhou, and Jintang Yang. Stereo matching with multi-scale based on anisotropic match cost. Concurrency and Computation: Practice and Experience, 32(24):e5918:1–e5918:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:SEC

[LDCD22] Yifan Liu, Chenglie Du, Jinchao Chen, and Xiaoyan Du. Scheduling energy-conscious tasks in distributed heterogeneous computing systems. Concurrency and Computation: Practice and Experience, 34(1):e6520:1–e6520:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:TLM

[LDH^{+}22] Yong Liu, Pavel Dmitriev, Yifei Huang, Andrew Brooks, Li Dong, Mengyue Liang, Zvi Boshernitzan, Jiwei Cao, and Bobby Nguy. Transfer learning meets sales engagement email classification: Evaluation, analysis, and strategies. Concurrency and Computation: Practice and Experience, 34(8): e5759:1–e5759:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2023:MAH

[LDS^{+}23] Shuang Li, Xiaoli Dong, Yuan Shi, Baoli Lu, Linjun Sun, and Wenfa Li. Multi-angle head pose classification with masks based on color texture analysis and stack generalization. Concurrency and Computation: Practice and Experience, 35(18): e6331:1–e6331:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:LSK

[LDW^{+}23] Ji Liu, Daxiang Dong, Xi Wang, An Qin, Xingjian Li, Patrick Valduriez, Dejing Dou, and Dianhai Yu. Large-scale knowledge distillation with elastic heterogeneous computing resources. Concurrency and Computation: Practice and Experience, 35(26):e7272:1–e7272:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:MSN

[LDZ^{+}22] Weimin Li, Zhibin Deng, Xiaokang Zhou, Qun Jin, and Bin Sheng. Modeling social network behavior spread based on group cohesion under uncertain environment. Concurrency and Computation: Practice and Experience, 34(21):e7101:1– e7101:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Le:2023:AEE

[Le23] Truong Xuan Le. An approach to evaluating the efficiency of a solution for visualization problem. Concurrency and Computation: Practice and Experience, 35(15):e6322:1–e6322:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luqman:2023:AFA

[LF23] Mohammad Luqman and Arman Rasool Faridi. Authentication of fog-assisted IoT networks using Advanced Encryption credibility approach with modified Diffie–Hellman encryption. Concurrency and Computation: Practice and Experience, 35 (22):e7742:1–e7742:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:DAO

[LFG^{+}22] Juan Li, Xianwen Fang, Xin Guo, Yuzhou Liu, and George K. Agordzo. Data attribute oriented business process effective infrequency behavior mining method. Concurrency and Computation: Practice and Experience, 34(25):e7265:1–e7265:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Levy:2020:UVA

[LFW20] Scott Levy, Kurt B. Ferreira, and Patrick Widener. The unexpected virtue of almost: Exploiting MPI collective operations to approximately coordinate checkpoints. Concurrency and Computation: Practice and Experience, 32(3):e4890:1– e4890:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:DPR

[LFWJ22] Kangshun Li, Xiangzheng Fu, Feng Wang, and Hassan Jalil. A dynamic population reduction differential evolution algorithm combining linear and nonlinear strategy piecewise functions. Concurrency and Computation: Practice and Experience, 34 (6):e6773:1–e6773:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:DSB

[LFX^{+}20] Ya-Nan Li, Xiaotao Feng, Jan Xie, Hanwen Feng, Zhenyu Guan, and Qianhong Wu. A decentralized and secure blockchain platform for open fair data trading. Concurrency and Computation: Practice and Experience, 32(7):e5578:1– e5578:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2021:SIC

[LG21] Huimin Lu and Joze Guna. Special issue on cognitive computing for robotic vision. Concurrency and Computation: Practice and Experience, 33(22):e6633:1–e6633:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lekshmanan:2023:LIU

[LG23] Lijesh Lekshmanan and Arockia Selva Saroja George. Landslide identification using water cycle particle swarm optimization-based deep generative adversarial network. Concurrency and Computation: Practice and Experience, 35(8): e7617:1–e7617:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Latif:2020:CAB

[LGA^{+}20] Sajid Latif, S. Mushhad M. Gilani, Liaqat Ali, Saleem Iqbal, and Misbah Liaqat. Characterizing the architectures and brokering protocols for enabling clouds interconnection. Concurrency and Computation: Practice and Experience, 32(21): e5676:1–e5676:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lepcha:2022:EMI

[LGDW22] Dawa Chyophel Lepcha, Bhawna Goyal, Ayush Dogra, and Shui-Hua Wang. An efficient medical image super resolution based on piecewise linear regression strategy using domain transform filtering. Concurrency and Computation: Practice and Experience, 34(20):e6644:1–e6644:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:MSC

[LGL^{+}21] Wuzhao Li, Weian Guo, Yongmei Li, Lei Wang, and Qidi Wu. Multi-swarm competitive swarm optimizer for large-scale optimization by entropy-assisted diversity measurement and management. Concurrency and Computation: Practice and Experience, 33(9):e6126:1–e6126:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:SLM

[LGL^{+}22] Xuejun Li, Lina Gong, Xiao Liu, Frank Jiang, Wenyu Shi, Lingmin Fan, Han Gao, Rui Li, and Jia Xu. Solving the last mile problem in logistics: a mobile edge computing and blockchain-based unmanned aerial vehicle delivery system. Concurrency and Computation: Practice and Experience, 34 (7):e6068:1–e6068:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2020:PBO

[LGLZ20] Liang Liu, Songtao Guo, Guiyan Liu, and Yue Zeng. Prioritybased online flow scheduling for network throughput maximization in software defined networking. Concurrency and Computation: Practice and Experience, 32(9):e5633:1– e5633:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lakshmi:2021:ECT

[LGM21] T. C. Subbu Lakshmi, D. Gnanadurai, and I. Muthulakshmi. Energy conserving texture-based adaptable compressive sensing scheme for WVSN. Concurrency and Computation: Practice and Experience, 33(3):e5178:1–e5178:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:NAF

[LGT^{+}23] Yi Liu, Xiaozhou Guo, Kaijun Tan, Guoliang Gong, and Huaxiang Lu. Novel activation function with pixelwise modeling capacity for lightweight neural network design. Concurrency and Computation: Practice and Experience, 35(18): e6350:1–e6350:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2022:GCN

[LGW^{+}22a] Jinjiao Lin, Tianqi Gao, Yuhua Wen, Xianmiao Yu, Bizhen You, Yanfang Yin, Yanze Zhao, and Haitao Pu. Graph convolutional neural networks-based assessment of students’ collaboration ability. Concurrency and Computation: Practice and Experience, 34(28):e7395:1–e7395:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2022:SDF

[LGW^{+}22b] Jiadong Lu, Fangming Gu, Yiqi Wang, Jiahui Chen, Zhiniang Peng, and Sheng Wen. Static detection of file access control vulnerabilities on windows system. Concurrency and Computation: Practice and Experience, 34(16):e6004:1–e6004:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luan:2023:LBN

[LGX^{+}23] Siyu Luan, Zonghua Gu, Rui Xu, Qingling Zhao, and Gang Chen. LRP-based network pruning and policy distillation of robust and non-robust DRL agents for embedded systems. Concurrency and Computation: Practice and Experience, 35 (19):e7351:1–e7351:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:TTE

[LGZ^{+}22] Fengyin Li, Rui Ge, Huiyu Zhou, Yilei Wang, Zhongxing Liu, and Xiaomei Yu. Tesia: a trusted efficient service evaluation model in Internet of Things based on improved aggregation signature. Concurrency and Computation: Practice and Experience, 34(16):e5739:1–e5739:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:SDI

[LH21] Jie Li and Bin Hu. Sparse data inspired depth of urbanization in Zhejiang based on geographic information system and visual system. Concurrency and Computation: Practice and Experience, 33(12):e5704:1–e5704:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2023:R

[LH23] Jie Li and Bin Hu. Retraction. Concurrency and Computation: Practice and Experience, 33(12):e7754:1–e7754:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lei:2023:CED

[LHC^{+}23] Dajiang Lei, Jie Huang, Hao Chen, Jie Li, and Yu Wu. Communication-efficient distributed large-scale sparse multinomial logistic regression. Concurrency and Computation: Practice and Experience, 35(18):e6148:1–e6148:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lai:2022:IAI

[LHJ22] Zhaolin Lai, Xiaochun Hu, and Caoqing Jiang. An intelligent algorithm with interactive learning mechanism for highdimensional optimization problem based on improved animal migration optimization. Concurrency and Computation: Practice and Experience, 34(12):e5774:1–e5774:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2022:GST

[LHK22] Sanghyun Lee, Bonghee Hong, and Woochan Kim. A generation of synthesized trajectories of free moving objects for testing a set of continuous spatio-temporal queries. Concurrency and Computation: Practice and Experience, 34(20): e7147:1–e7147:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lv:2020:NLA

[LHL^{+}20] Yali Lv, Weixin Hu, Jiye Liang, Yuhua Qian, and Junzhong Miao. A naive learning algorithm for class-bridge-decomposable multidimensional Bayesian network classifiers.

Concurrency and Computation: Practice and Experience, 32 (19):e5778:1–e5778:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:CDU

[LHL^{+}22] Changzheng Liu, Fengling Huang, Ruixuan Li, Qi Yang, Yuhua Li, and Shui Yu. Community detection using multitopology and attributes in social networks. Concurrency and Computation: Practice and Experience, 34(12):e6028:1– e6028:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lirkov:2021:PAP

[LHPG21] Ivan Lirkov, Stanislav Harizanov, Marcin Paprzycki, and Maria Ganzha. Performance analysis of parallel highresolution image restoration algorithms on Intel supercomputer. Concurrency and Computation: Practice and Experience, 33(4):e5996:1–e5996:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:BRL

[LHWT20] Dengshi Li, Ruimin Hu, Xiaochen Wang, and Weiping Tu. Book review: Loudspeaker triplet selection based on low distortion within head for multichannel conversion of smart 3D home theater. Concurrency and Computation: Practice and Experience, 32(13):e4796:1–e4796:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2023:ADM

[LHX^{+}23] Di Li, Yikun Hu, Guoqing Xiao, Mingxing Duan, and Kenli Li. An active defense model based on situational awareness and firewalls. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:FIL

[Li20] Feng Li. On forwarding indices of lexicographic product networks. Concurrency and Computation: Practice and Experience, 32(23):e5487:1–e5487:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:FER

[Li21] Hongbin Li. Feature extraction, recognition, and matching of damaged fingerprint: Application of deep learning network. Concurrency and Computation: Practice and Experience, 33 (6):e6057:1–e6057:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lian:2024:ISS

[Lia24] Lian Lian. An improved sparrow search algorithm using chaotic opposition-based learning and hybrid updating rules. Concurrency and Computation: Practice and Experience, 36 (14):e8101:1–e8101:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liqiang:2022:RSC

[Liq22] Guo Liqiang. Research on single chip microcomputer teaching platform based on wireless communication. Concurrency and Computation: Practice and Experience, 34(13):e5895:1– e5895:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2021:CUA

[Liu21] Zhanyu Liu. Construction of urban agricultural health informatics safety supervision system based on imaging and deep learning. Concurrency and Computation: Practice and Experience, 33(12):e5834:1–e5834:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:Rb

[Liu23] Zhanyu Liu. Retraction. Concurrency and Computation: Practice and Experience, 33(12):e7823:1–e7823:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2024:MOF

[Liu24a] Lei Liu. Multi-objective firefly algorithm for enhanced balanced exploitation and exploration capabilities. Concurrency and Computation: Practice and Experience, 36(7):e7973:1– e7973:??, March 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2024:ESI

[Liu24b] Xiao Liu. Editorial for special issue on “Edge computing accelerated deep learning: Technologies and applications”. Concurrency and Computation: Practice and Experience, 36(10): e7345:1–e7345:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lalitha:2022:IMT

[LJ22] Priya Raghavan Nair Lalitha and S. Vinila Jinny. Internet of medical things-based multitiered and hybrid architectural framework for effective heart disease prediction model. Concurrency and Computation: Practice and Experience, 34(15): e6953:1–e6953:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lilda:2022:EML

[LJB22] S Daphin Lilda, R Jayaparvathy, and Akshaya Balaji. An efficient machine learning based ventricular late potential detection and classification technique for cardiac healthcare. Concurrency and Computation: Practice and Experience, 34(26): e7279:1–e7279:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lavanya:2023:SVM

[LJBS23] S. Lavanya, S. V. Annlin Jeba, P. Bhuvaneswari, and Francis H. Shajin. Support vector machine classifier optimized with seagull optimization algorithm for brain tumor classification. Concurrency and Computation: Practice and Experience, 35(1):e7396:1–e7396:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2023:EWA

[LJC23] Geon Ju Lee, Jason J. Jung, and David Camacho. Exploiting weighted association rule mining for indicating synergic formation tactics in soccer teams. Concurrency and Computation: Practice and Experience, 35(16):e6221:1–e6221:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2021:DAE

[LJP^{+}21] Jinzhao Lin, Meiqiu Jiang, Yu Pang, Huiqian Wang, Zhu Chen, Chongyuan Yan, Qinghui Liu, and Yuanfa Wang. A desmoking algorithm for endoscopic images based on improved U-Net model. Concurrency and Computation: Practice and Experience, 33(22):e6320:1–e6320:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:AIM

[LJZ21] Fachao Li, Chenxia Jin, and Xiao Zhang. Attribute importance measurement based on the data effect and its performance analysis in computation practice. Concurrency and Computation: Practice and Experience, 33(8):e5038:1– e5038:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lakshmi:2022:ARM

[LK22a] N. Lakshmi and M. Krishnamurthy. Association rule mining based fuzzy manta ray foraging optimization algorithm for frequent itemset generation from social media. Concurrency and Computation: Practice and Experience, 34(10):e6790:1– e6790:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lalithadevi:2022:DDR

[LK22b] Balakrishnan Lalithadevi and Sivamohan Krishnaveni. Detection of diabetic retinopathy and related retinal disorders using fundus images based on deep learning and image processing techniques: a comprehensive review. Concurrency and Computation: Practice and Experience, 34(19):e7032:1–e7032:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Litzinger:2022:CGE

[LK22c] Sebastian Litzinger and Jörg Keller. Code generation for energy-efficient execution of dynamic streaming task graphs on parallel and heterogeneous platforms. Concurrency and Computation: Practice and Experience, 34(2):e6072:1– e6072:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2023:SAP

[LK23] Suhyeon Lee and Seungjoo Kim. Shorting attack: Predatory, destructive short selling on Proof-of-Stake cryptocurrencies. Concurrency and Computation: Practice and Experience, 35 (16):e6585:1–e6585:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Latha:2022:DCN

[LKR^{+}22] C. Jenifa Latha, K. Kalaiselvi, S. Ramanarayan, R. Srivel, S. Vani, and T. V. M. Sairam. Dynamic convolutional neural network based e-waste management and optimized collection planning. Concurrency and Computation: Practice and Experience, 34(17):e6941:1–e6941:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liazid:2021:BRI

[LL21] Hidaya Liazid and Mohamed Lehsaini. A brief review on integration between wireless sensor networks and cloud. Concurrency and Computation: Practice and Experience, 33(20): e6328:1–e6328:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:SIR

[LL23] Xiaozhu Liu and Bang Li. Stylized image resolution enhancement scheme based on an improved convolutional neural network in cyber-physical systems. Concurrency and Computation: Practice and Experience, 35(13):e6615:1–e6615:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2024:ROD

[LL24] Jing Lu and Yongyi Li. Research on optimal deep learningbased formation flight positioning of UAV. Concurrency and Computation: Practice and Experience, 36(14):e8099:1– e8099:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Louati:2022:PFR

[LLA^{+}22] Ali Louati, Rahma Lahyani, Abdulaziz Aldaej, Abdullah Aldumaykhi, and Saad Otai. Price forecasting for real estate using machine learning: a case study on Riyadh City. Concurrency and Computation: Practice and Experience, 34(6): e6748:1–e6748:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Laxman:2022:PAR

[LLAV22] Raghavender Reddy Kothi Laxman, Amit Lathigara, Rajanikanth Aluvalu, and Uma Maheswari Viswanadhula. PGWO-AVS-RDA: an intelligent optimization and clustering based load balancing model in cloud. Concurrency and Computation: Practice and Experience, 34(21):e7136:1–e7136:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:EMS

[LLC21a] Wei Li, Guogang Liu, and Junho Choi. Environmental monitoring system for intelligent stations. Concurrency and Computation: Practice and Experience, 33(2):e5131:1–e5131:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lv:2021:FIB

[LLC^{+}21b] Yan Lv, Yaojin Lin, Xiangyan Chen, Chenxi Wang, and Shaozi Li. Feature interaction based online streaming feature selection via buffer mechanism. Concurrency and Computation: Practice and Experience, 33(21):e6435:1–e6435:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:PSS

[LLC^{+}22a] Hui Li, Qingqing Liang, Mei Chen, Zhenyu Dai, Huanjun Li, and Ming Zhu. Pruning SMAC search space based on key hyperparameters. Concurrency and Computation: Practice and Experience, 34(9):e5805:1–e5805:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lobato:2022:FAT

[LLC^{+}22b] Antonio G. Pastana Lobato, Martin Andreoni Lopez, Alvaro A. Cardenas, Otto Carlos M. B. Duarte, and Guy Pujolle. A fast and accurate threat detection and prevention architecture using stream processing. Concurrency and Computation: Practice and Experience, 34(3):e6561:1–e6561:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2019:EPP

[LLH19] Rong Li, Fei Li, and Jiewu Huang. Evaluation of the predictive performance of the principal component two-parameter estimator. Concurrency and Computation: Practice and Experience, 31(12):e4710:1–e4710:??, June 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See retraction notice [HML20].

Liu:2020:EAT

[LLH^{+}20] Xing Liu, Panwen Liu, Lun Hu, Chengming Zou, and Zhangyu Cheng. Energy-aware task scheduling with time constraint for heterogeneous cloud datacenters. Concurrency and Computation: Practice and Experience, 32(18):e5437:1– e5437:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:DLE

[LLJ^{+}20] Gen Li, Chang Ha Lee, Jason J. Jung, Young Chul Youn, and David Camacho. Deep learning for EEG data analytics: a survey. Concurrency and Computation: Practice and Experience, 32(18):e5199:1–e5199:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lei:2022:GCN

[LLJ^{+}22a] Fangyuan Lei, Xun Liu, Jianjian Jiang, Liping Liao, Jun Cai, and Huimin Zhao. Graph convolutional networks with higherorder pooling for semisupervised node classification. Concurrency and Computation: Practice and Experience, 34(16): e5695:1–e5695:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:WAP

[LLJ^{+}22b] Yibo Liu, Chengcheng Li, Du Jiang, Baojia Chen, Nannan Sun, Yongcheng Cao, Bo Tao, and Gongfa Li. Wrist angle prediction under different loads based on GA-ELM neural network and surface electromyography. Concurrency and Computation: Practice and Experience, 34(3):e6574:1–e6574:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2021:MRI

[LLJR21] Yangfei Lin, Jie Li, Xiaohua Jia, and Kui Ren. Multiplereplica integrity auditing schemes for cloud data storage. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2021:DLB

[LLKS21] Szu-Yin Lin, Fang-Yie Leu, Chia-Yin Ko, and Ming-Chien Shih. A deep learning-based indoor-positioning approach using received strength signal indication and carrying mode information. Concurrency and Computation: Practice and Experience, 33(23):e6135:1–e6135:??, December 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2021:EMC

[LLL^{+}21] Feng Lu, Ruoxue Liu, Wei Li, Hai Jin, and Albert Y. Zomaya. Efficient mobile code offloading in heterogeneous wireless networks. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:MDM

[LLL^{+}22] Yanhua Liu, Jiaqi Li, Baoxu Liu, Xiaoling Gao, and Ximeng Liu. Malware detection method based on image analysis and generative adversarial networks. Concurrency and Computation: Practice and Experience, 34(22):e7170:1–e7170:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lou:2023:PGI

[LLL23] Qihe Lou, Yanbin Li, and Ning Liu. Power grid investment forecasting method based on PSO-GRU neural network. Concurrency and Computation: Practice and Experience, 35(27): e7846:1–e7846:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2024:LMO

[LLL^{+}24] Lei Li, Sheng Lian, Dazhen Lin, Zhiming Luo, Beizhan Wang, and Shaozi Li. Learning multi-organ and tumor segmentation from partially labeled datasets by a conditional dynamic attention network. Concurrency and Computation: Practice and Experience, 36(1):e7869:1–e7869:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2020:RAI

[LLLX20] Shouqiang Liu, Miao Li, Min Li, and Qingzhen Xu. Research of animals image semantic segmentation based on deep learning. Concurrency and Computation: Practice and Experience, 32(1):e4892:1–e4892:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:BAS

[LLM^{+}22] Zhenpeng Liu, Qiannan Liu, Dewei Miao, Lele Ren, and Yonggang Zhao. A blockchain anonymity solution to prevent location homogeneity attacks. Concurrency and Computation: Practice and Experience, 34(27):e7326:1–e7326:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liang:2024:DOD

[LLM^{+}24] Zhihao Liang, Huijuan Lu, Zhendong Ming, Zhuijun Chai, and Yudong Yao. DML-OFA: Deep mutual learning with online feature alignment for the detection of COVID-19 from chest X-ray images. Concurrency and Computation: Practice and Experience, 36(11):e8023:1–e8023:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:PCB

[LLMX21] Lei-Jun Li, Mei-Zheng Li, Ju-Sheng Mi, and Bin Xie. Prediction confidence-based dynamic selection and weighted integration. Concurrency and Computation: Practice and Experience, 33(8):e5055:1–e5055:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2021:LLR

[LLMZ21] Dezheng Liu, Zhongyu Li, Yuanyuan Ma, and Yulong Zhang. Listwise learning to rank with extreme order sensitive constraint via cross-correntropy. Concurrency and Computation: Practice and Experience, 33(22):e5796:1–e5796:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2023:BIQ

[LLN^{+}23] Yaxuan Lu, Weijun Li, Xin Ning, Xiaoli Dong, Liping Zhang, Linjun Sun, and Chuantong Cheng. Blind image quality assessment based on the multiscale and dual-domains features fusion. Concurrency and Computation: Practice and Experience, 35(18):e6177:1–e6177:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Long:2021:ASS

[LLO21] Fei Long, Cai Liu, and Weihua Ou. Almost sure stability for a class of dual switching linear discrete-time systems. Concurrency and Computation: Practice and Experience, 34(1): e5666:1–e5666:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

LaBorde:2022:DTT

[LLP^{+}22] Pierre LaBorde, Lance Lebanoff, Christina Peterson, Deli Zhang, and Damian Dechev. Dynamic transactional transformation. Concurrency and Computation: Practice and Experience, 34(2):e5732:1–e5732:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Laso:2021:LIW

[LLR^{+}21] R. Laso, O. G. Lorenzo, F. F. Rivera, J. C. Cabaleiro, T. F. Pena, and J. A. Lorenzo. LBMA and IMAR^{2}: Weighted lottery based migration strategies for NUMA multiprocessing servers. Concurrency and Computation: Practice and Experience, 33(11):e5950:1–e5950:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:RLB

[LLS22] Zhixiang Liu, Yu Li, and Wei Song. Regularized lattice Boltzmann method parallel model on heterogeneous platforms. Concurrency and Computation: Practice and Experience, 34 (22):e6875:1–e6875:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:EGA

[LLSC22] Bowen Li, Hua Li, Qiubai Sun, and Xuebo Chen. Evolutionary game analysis between businesses and consumers under the background of Internet rumors. Concurrency and Computation: Practice and Experience, 34(13):e5897:1–e5897:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2021:DHB

[LLT21] Zhuotong Liu, Chen Li, and Lihua Tian. Deep hashing based on triplet labels and quantitative regularization term with exponential convergence. Concurrency and Computation: Practice and Experience, 34(1):e5583:1–e5583:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2024:SAB

[LLTZ24] Chen Li, Ziniu Lin, Lihua Tian, and Bin Zhang. A scheduling algorithm based on critical factors for heterogeneous multicore processors. Concurrency and Computation: Practice and Experience, 36(7):e7969:1–e7969:??, March 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:CWC

[LLW^{+}20] Xiaoyan Li, Zhigang Lv, Peng Wang, Mengyu Sun, and Mengyu Qiao. Combination weighted clustering algorithms in cognitive radio networks. Concurrency and Computation: Practice and Experience, 32(23):e5516:1–e5516:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liao:2021:OGR

[LLW^{+}21] Shangchun Liao, Gongfa Li, Hao Wu, Du Jiang, Ying Liu, Juntong Yun, Yibo Liu, and Dalin Zhou. Occlusion gesture recognition based on improved SSD. Concurrency and Computation: Practice and Experience, 33(6):e6063:1–e6063:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:SMS

[LLW^{+}22a] Peng Li, Kang Li, Yilei Wang, Ying Zheng, Dongfeng Wang, Guoyu Yang, and Xiaomei Yu. A systematic mapping study for blockchain based on complex network. Concurrency and Computation: Practice and Experience, 34(14):e5712:1– e5712:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2022:SCN

[LLW^{+}22b] Jiashuo Lin, Liping Liao, Tao Wang, Jun Zhang, and Lianglun Cheng. SDCCP: Control the network using software-defined networking and end-to-end congestion control. Concurrency and Computation: Practice and Experience, 34(16):e5716:1– e5716:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2023:DMM

[LLW^{+}23] Yujin Li, Bo Liu, Enju Wu, Jianqiang Li, Zhangbing Zhou, and Wenbo Zhang. DRA-MQoS: an MQoS scheduling algorithm based on resource feature matching in federated edge cloud. Concurrency and Computation: Practice and Experience, 35(2):e7478:1–e7478:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:EAD

[LLX^{+}21] Rui Li, Xuejun Li, Jia Xu, Frank Jiang, Zhaohong Jia, Di Shao, Lei Pan, and Xiao Liu. Energy-aware decisionmaking for dynamic task migration in MEC-based unmanned aerial vehicle delivery system. Concurrency and Computation: Practice and Experience, 33(22):e6092:1–e6092:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2024:PND

[LLX^{+}24] Qiong Lu, Huaizhong Lin, Wei Xing, Lei Zhao, and Jingjing Chen. PKD-Net: Distillation of prior knowledge for image completion by multi-level semantic attention. Concurrency and Computation: Practice and Experience, 36(10):e7960:1– e7960:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2023:LDL

[LLY^{+}23] Yulin Li, Yaojin Lin, Xiehua Yu, Lei Guo, and Shaozi Li. Label distribution learning with high-order label correlations. Concurrency and Computation: Practice and Experience, 35(21):e7710:1–e7710:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2023:DDL

[LLYZ23] Qianqian Luo, Zhenzhou Lin, Guohua Yang, and Xiaofeng Zhao. DEC: a deep-learning based edge-cloud orchestrated system for recyclable garbage detection. Concurrency and Computation: Practice and Experience, 35(13):e6661:1– e6661:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:ITL

[LLZ^{+}20] Lei Li, Zhiming Luo, Mengting Zhang, Yuanzheng Cai, Candong Li, and Shaozi Li. An iterative transfer learning framework for cross-domain tongue segmentation. Concurrency and Computation: Practice and Experience, 32(14):e5714:1– e5714:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:CBO

[LLZ^{+}21a] Longzhu Li, Yaojin Lin, Hong Zhao, Jinkun Chen, and Shaozi Li. Causality-based online streaming feature selection. Concurrency and Computation: Practice and Experience, 33(20): e6347:1–e6347:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:PDE

[LLZ^{+}21b] Xiaole Li, Yingji Luo, Wenyin Zhang, Deqian Fu, Hua Wang, and Linbo Zhai. Progressive disaster evacuation in cloud datacenter network. Concurrency and Computation: Practice and Experience, 33(6):e6060:1–e6060:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lastovetsky:2020:IHC

[LM20a] Alexey L. Lastovetsky and Ravi Reddy Manumachu. The 27th International Heterogeneity in Computing Workshop and the 16th International Workshop on Algorithms, Models and Tools for Parallel Computing on Heterogeneous Platforms. Concurrency and Computation: Practice and Experience, 32(15):e5736:1–e5736:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2020:ECC

[LM20b] Hongxia Liu and Scott Mizzi. Evaluating climate changes and land use changes on water resources using hybrid Soil and Water Assessment Tool-DEEP optimized by metaheuristics. Concurrency and Computation: Practice and Experience, 32(24):e5945:1–e5945:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lakshmi:2021:CMR

[LM21] K Lakshmi and T Meyyappan. Compact in-memory representation of large graph databases for efficient mining of maximal frequent sub graphs. Concurrency and Computation: Practice and Experience, 33(3):e5243:1–e5243:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:GLP

[LM22a] Bingchan Li and Bo Mao. A grain loss prediction method based on integration of multiple classification models. Concurrency and Computation: Practice and Experience, 34(8): e6116:1–e6116:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:ODP

[LM22b] Xiaoyan Li and Saleh Mobayen. Optimal design of a PEMFCbased combined cooling, heating and power system based on an improved version of Aquila optimizer. Concurrency and Computation: Practice and Experience, 34(15):e6976:1– e6976:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Longkumer:2024:ICP

[LM24] Imtisenla Longkumer and Dilwar Hussain Mazumder. Improving cancer prediction using feature selection in spark environment. Concurrency and Computation: Practice and Experience, 36(2):e7903:1–e7903:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2023:AAP

[LMBK23] Minho Lee, Libor Mesicek, Kitae Bae, and Hoon Ko. AI advisor platform for disaster response based on big data. Concurrency and Computation: Practice and Experience, 35(16): e6215:1–e6215:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Labidi:2020:CSN

[LMGG20] Taher Labidi, Achraf Mtibaa, Walid Gaaloul, and Faiez Gargouri. Cloud SLA negotiation and re-negotiation: an ontology-based context-aware approach. Concurrency and Computation: Practice and Experience, 32(15):e5315:1– e5315:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:DLW

[LmJdL^{+}22] Kui Li, Yi mu Ji, Shang dong Liu, Fei Wu, Hai chang Yao, Jing He, Qiang Liu, Yan lan Liu, Si si Shao, and Shuai You. DVO + LCLMF: a web service recommendation mechanism with QoS privacy preservation. Concurrency and Computation: Practice and Experience, 34(18):e7003:1–e7003:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2023:SWF

[LML^{+}23] Gang Li, Yuan Mengge, Yu Li, Jingang Huang, Ling Zhang, Haixuan Zhang, and Zhenhua Han. Seismic P -wave firstarrival picking model based on EQK-IncResNet. Concurrency and Computation: Practice and Experience, 35(13):e7392:1– e7392:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liang:2022:DGP

[LMM^{+}22] Chen Liang, Meixia Miao, Jianfeng Ma, Hongyang Yan, Qun Zhang, and Xinghua Li. Detection of global positioning system spoofing attack on unmanned aerial vehicle system. Concurrency and Computation: Practice and Experience, 34(7): e5925:1–e5925:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lerat:2023:SND

[LMM23] Jean-Sébastien Lerat, Sidi Ahmed Mahmoudi, and Saïd Mahmoudi. Single node deep learning frameworks: Comparative study and CPU/GPU performance analysis. Concurrency and Computation: Practice and Experience, 35(14):e6730:1– e6730:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lapegna:2024:SIP

[LMMS24] Marco Lapegna, Valeria Mele, Raffaele Montella, and Lukasz Szustak. Special issue on the pervasive nature of HPC (PN-HPC). Concurrency and Computation: Practice and Experience, 36(10):e8003:1–e8003:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lakshmanan:2022:ESD

[LMR22] Selvam Lakshmanan, Braveen Manimozhi, and Venkatesan Ramachandran. An efficient and secure data sharing scheme for cloud data using hash based quadraplet wavelet permuted cryptography approach. Concurrency and Computation: Practice and Experience, 34(27):e7324:1–e7324:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:FEB

[LN20] Xiaolin Li and Haitao Niu. Feature extraction based on deepconvolutional neural network for face recognition. Concurrency and Computation: Practice and Experience, 32(22):1, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:PRS

[LNC^{+}20] Weihao Li, Ben Niu, Jin Cao, Yurong Luo, and Hui Li. A personalized range-sensitive privacy-preserving scheme in LBSs. Concurrency and Computation: Practice and Experience, 32 (5):e5462:1–e5462:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2024:BDT

[LNI^{+}24] Xiaotian Li, Koji Nakano, Yasuaki Ito, Daisuke Takafuji, Takashi Yazane, Junko Yano, Takumi Kato, Shiro Ozaki, Rie Mori, and Ryota Katsuki. Bit duplication technique to generate hard quadratic unconstrained binary optimization problems with adjustable sizes. Concurrency and Computation: Practice and Experience, 36(10):e7967:1–e7967:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2021:SDA

[LP21] Seong-Hun Lee and Jaehwa Park. A spatiotemporal data acquisition toolkit for volume estimation tools in precision agriculture. Concurrency and Computation: Practice and Experience, 33(3):e5258:1–e5258:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2021:SBP

[LPC^{+}21] Cong Liu, Yulong Pei, Long Cheng, Qingtian Zeng, and Hua Duan. Sampling business process event logs using graph-based ranking model. Concurrency and Computation: Practice and Experience, 33(5):e5974:1–e5974:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2020:DNN

[LPHK20] Keon Myung Lee, Ki-Sun Park, Kyung-Soon Hwang, and Kwang-Il Kim. Deep neural network model construction with interactive code reuse and automatic code transformation. Concurrency and Computation: Practice and Experience, 32(18):e5480:1–e5480:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:TLB

[LPW^{+}21] Sijia Li, Jiali Pang, Qiang Wu, Na Yao, and Weiwei Yuan. Transfer learning based attack detection for wireless communication networks. Concurrency and Computation: Practice and Experience, 33(24):e6461:1–e6461:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:AFB

[LPZ^{+}22] Danxu Liu, Junjie Peng, Xin Zhang, Yongkang You, and Bo Ning. Application features-based virtual machine deployment strategy in cloud environment. Concurrency and Computation: Practice and Experience, 34(5):e6691:1–e6691:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2022:TMI

[LRC^{+}22] Liangfu Lu, Xiaohan Ren, Chenming Cui, Yun Luo, and Maolin Huang. Tensor mutual information and its applications. Concurrency and Computation: Practice and Experience, 34(14):e5686:1–e5686:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:GBO

[LS20] Shengpu Li and Yize Sun. Garden balsam optimization algorithm. Concurrency and Computation: Practice and Experience, 32(2):e5456:1–e5456:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Langr:2022:CPQ

[LS22] Daniel Langr and Klára Schovánková. CPP11sort: a parallel quicksort based on C++ threading. Concurrency and Computation: Practice and Experience, 34(4):e6606:1–e6606:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lahane:2023:SIS

[LS23a] Priti S. Lahane and Manish Sharma. Secured information sharing for supply chain management based on blockchain technology and optimal key generation process. Concurrency and Computation: Practice and Experience, 35(3):e7522:1– e7522:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lisha:2023:HAB

[LS23b] L. B. Lisha and C. Helen Sulochana. Highly accurate blood vessel segmentation using texture-based modified K-means clustering with deep learning model. Concurrency and Computation: Practice and Experience, 35(7):e7590:1–e7590:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lone:2024:TET

[LS24] Kalimullah Lone and Shabir Ahmad Sofi. e-TOALB: an efficient task offloading in IoT–fog networks. Concurrency and Computation: Practice and Experience, 36(6):e7951:1– e7951:??, March 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:HMG

[LSD21] Xiaoli Li, Guomei Song, and Zhenlong Du. Hybrid model of generative adversarial network and Takagi-Sugeno for multidimensional incomplete hydrological big data prediction.

Concurrency and Computation: Practice and Experience, 34 (1):e5713:1–e5713:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2020:TSC

[LSL20] Hongdan Liu, Rong Sun, and Qi Liu. The tactics of ship collision avoidance based on Quantum-behaved Wolf Pack Algorithm. Concurrency and Computation: Practice and Experience, 32(6):e5196:1–e5196:??, March 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2020:ACS

[LSLY20] Xiaolong Liu, Ruey-Kai Sheu, Win-Tsung Lo, and Shyan-Ming Yuan. Automatic cloud service testing and bottleneck detection system with scaling recommendation. Concurrency and Computation: Practice and Experience, 32(1):e5161:1– e5161:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2021:SFS

[LSQW21] Shuqi Liu, Chengai Sun, Liqing Qiu, and Moji Wei. Spectatorfilter-spreader-stifler information propagation model with wake-up mechanism in social networks. Concurrency and Computation: Practice and Experience, 33(5):e5973:1– e5973:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lei:2021:RAB

[LSS^{+}21] Guo Lei, Yang Sheng, Li Shaozi, Wu Qingshou, and Yu Wensen. Recommendation algorithm based on community structure and user trust. Concurrency and Computation: Practice and Experience, 33(20):e6375:1–e6375:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luz:2024:STC

[LSS24] Helder J. F. Luz, Paulo S. L. Souza, and Simone R. S. Souza. Structural testing for CUDA programming model. Concurrency and Computation: Practice and Experience, 36(14): e8105:1–e8105:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lei:2022:HCD

[LSSQ22] Guo Lei, Yang Sheng, Li Shaozi, and Wu Qingshou. Hierarchical community-discovery algorithm combining core nodes and three-order structure model. Concurrency and Computation: Practice and Experience, 34(4):e6669:1–e6669:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:OEL

[LST22] Bo Li, Hongbin Sun, and Milad Teimourian. Optimal electric load forecasting for systems by an adaptive Crow Search Algorithm: a case study. Concurrency and Computation: Practice and Experience, 34(21):e7120:1–e7120:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2020:LEN

[LSW^{+}20] Ruizhen Liu, Zhiyi Sun, Anhong Wang, Kai Yang, Yin Wang, and Qianlai Sun. Lightweight efficient network for defect classification of polarizers. Concurrency and Computation: Practice and Experience, 32(11):e5663:1–e5663:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2021:DMB

[LSW21] XingHai Lu, XiaoHua Shi, and WenRu Wang. DATAM: a model-based tool for dependability analysis. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2021:NED

[LSX21] Zhiyi Lin, Qing Su, and Guobo Xie. NMIEDA: Estimation of distribution algorithm based on normalized mutual information. Concurrency and Computation: Practice and Experience, 33(6):e6074:1–e6074:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2023:LMS

[LSZ^{+}23] Wei Li, Guanxi Shen, Jingfang Zhang, Dong Liu, and Chang Choi. A LoRaWAN monitoring system for large buildings based on embedded edge computing in indoor environment.

Concurrency and Computation: Practice and Experience, 35 (16):e6306:1–e6306:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2024:NCL

[LSZ24] Li Li, Renjia Su, and Xin Zhao. Neighbor cleaning learning based cost-sensitive ensemble learning approach for software defect prediction. Concurrency and Computation: Practice and Experience, 36(12):e8017:1–e8017:??, May 30, 2024. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lei:2023:EED

[LSZL23] Dajiang Lei, Ling Shen, Liping Zhang, and Weisheng Li. Extraction-and-excitation deep neural network for pansharpening. Concurrency and Computation: Practice and Experience, 35(18):e6098:1–e6098:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2020:PGM

[LTL^{+}20] Shuai Liu, Na Ta, Mengye Lu, Gaocheng Liu, Weiling Bai, and Wenhui Li. Parallel generated method of transcriptional regulatory networks. Concurrency and Computation: Practice and Experience, 32(1):e4938:1–e4938:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:EIS

[LTLX22] Wenjuan Li, Fei Tian, Jin Li, and Yang Xiang. Evaluating intrusion sensitivity allocation with supervised learning in collaborative intrusion detection. Concurrency and Computation: Practice and Experience, 34(16):e5957:1–e5957:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2024:DND

[LTW^{+}24] Jinchao Li, Yuxuan Tian, Jindong Wang, Jinling Yang, Yinfang Zhu, Xuemin Cheng, and Qun Hao. Design of novel deformable mirror with large displacement. Concurrency and Computation: Practice and Experience, 36(8):e7984:1– e7984:??, April 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:CNS

[LW20a] Bo Li and Shi Wang. Cutting and nesting system of cloud manufacturing based on SOA. Concurrency and Computation: Practice and Experience, 32(6):e5088:1–e5088:??, March 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:MIS

[LW20b] Yuanmu Li and Zhanqing Wang. A medical image segmentation method based on hybrid active contour model with global and local features. Concurrency and Computation: Practice and Experience, 32(19):e5763:1–e5763:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2021:DWF

[LWC^{+}21] Jiaxiang Lin, Liping Wu, Riqing Chen, Jianwei Wu, and Xueping Wang. Double-weighted fuzzy clustering with samples and generalized entropy features. Concurrency and Computation: Practice and Experience, 33(8):e5758:1–e5758:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2021:DAM

[LWCM21] Chun Liu, Shuangyan Wang, Salvatore Cuomo, and Gang Mei. Data analysis and mining of traffic features based on taxi GPS trajectories: a case study in Beijing. Concurrency and Computation: Practice and Experience, 33(3):e5332:1– e5332:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:EAV

[LWCZ22] Xialin Liu, Junsheng Wu, Lijun Chen, and Lili Zhang. Energy-aware virtual machine consolidation based on evolutionary game theory. Concurrency and Computation: Practice and Experience, 34(10):e6830:1–e6830:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:TSD

[LWHW22] Dongsheng Liu, Yuting Wu, Deyan Hong, and Siting Wang. Time series data augmentation method of small sample based on optimized generative adversarial network. Concurrency and Computation: Practice and Experience, 34(27):e7331:1– e7331:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:EDA

[LWL^{+}23] Tonglai Liu, Jigang Wu, Jiaxing Li, Jingyi Li, and Yidong Li. Efficient decentralized access control for secure data sharing in cloud computing. Concurrency and Computation: Practice and Experience, 35(17):e6383:1–e6383:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2024:OAD

[LWL^{+}24] Zuoyong Li, Weice Wang, Taotao Lai, Haiping Xu, and Pantea Keikhosrokiani. Object-aware deep feature extraction for feature matching. Concurrency and Computation: Practice and Experience, 36(5):e7932:1–e7932:??, February 29, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:EDL

[LWLZ20] Ning Li, Lu Wang, Xinquan Li, and Qing Zhu. An effective deep learning neural network model for short-term load forecasting. Concurrency and Computation: Practice and Experience, 32(7):e5595:1–e5595:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liao:2023:MFH

[LWW23] Junguo Liao, Haifeng Wang, and June Wu. A multikey fully homomorphic encryption privacy protection protocol based on blockchain for edge computing system. Concurrency and Computation: Practice and Experience, 35(4):e7539:1– e7539:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2020:CDS

[LWZ^{+}20] Nijia Lu, Guohua Wu, Zhen Zhang, Yitao Zheng, Yizhi Ren, and Kim-Kwang Raymond Choo. Cyberbullying detection in social media text based on character-level convolutional neural network with shortcuts. Concurrency and Computation: Practice and Experience, 32(23):e5627:1–e5627:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:SET

[LWZ22a] Aihua Li, Diwen Wang, and Meihong Zhu. Service evaluation through FH-entropy method: a framework for the elderly care station. Concurrency and Computation: Practice and Experience, 34(8):e6045:1–e6045:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:FSR

[LWZ^{+}22b] Weimin Li, Dingmei Wei, Xiaokang Zhou, Shaohua Li, and Qun Jin. F-SWIR: Rumor fick-spreading model considering fusion information decay in social networks. Concurrency and Computation: Practice and Experience, 34(22):e7166:1– e7166:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:CCM

[LWZC21] Dengao Li, Xiaoyu Wang, Jumin Zhao, and Junbing Cheng. Channel compensation multipath mitigation technique for Kalman-based least mean square based on Kalman estimation. Concurrency and Computation: Practice and Experience, 33(9):e6133:1–e6133:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2023:AUE

[LWZY23] Qingpeng Li, Gang Wang, Yang Zhang, and Qin Yang. Analysis of user electricity consumption behavior based on density peak clustering with shared neighbors and attractiveness. Concurrency and Computation: Practice and Experience, 35 (3):e7518:1–e7518:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2022:TPD

[LX22] Sin Hong Lin and Ruliang Xiao. Towards publishing directed social network data with k-degree anonymization. Concurrency and Computation: Practice and Experience, 34(24): e7226:1–e7226:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:FBM

[LXC^{+}22a] Zhiqiang Liu, Bo Xu, Bo Cheng, Xiaomei Hu, and Karlo Abnoosian. A fuzzy-based method for cloud service migration using a shark smell optimization algorithm. Concurrency and Computation: Practice and Experience, 34(15):e6970:1– e6970:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:IDS

[LXC^{+}22b] Zhiqiang Liu, Bo Xu, Bo Cheng, Xiaomei Hu, and Mehdi Darbandi. Intrusion detection systems in the cloud computing: a comprehensive and deep literature review. Concurrency and Computation: Practice and Experience, 34(4):e6646:1– e6646:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:TLL

[LXJ^{+}22] Yuting Liu, Manman Xu, Guozhang Jiang, Xiliang Tong, Juntong Yun, Ying Liu, Baojia Chen, Yongcheng Cao, Nannan Sun, and Zeshen Li. Target localization in local dense mapping using RGBD SLAM and object detection. Concurrency and Computation: Practice and Experience, 34(4):e6655:1– e6655:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:STA

[LXKW23] Jianxun Liu, Runbin Xie, Guosheng Kang, and Yiping Wen. Spatial-temporal aware service composition for production factors under industrial internet. Concurrency and Computation: Practice and Experience, 35(12):e7689:1–e7689:??, May 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:MTP

[LXT^{+}22] Yibo Liu, Fan Xiao, Xiliang Tong, Bo Tao, Manman Xu, Guozhang Jiang, Baojia Chen, Yongcheng Cao, and Nannan Sun. Manipulator trajectory planning based on work subspace division. Concurrency and Computation: Practice and Experience, 34(5):e6710:1–e6710:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2023:CEQ

[LXW^{+}23] Hongjian Li, Chen Xu, Tiantian Wang, Jingjing Wang, Peng Zheng, Tongming Liu, and Libo Tang. A cost-efficient and QoS-aware adaptive placement of applications in fog computing. Concurrency and Computation: Practice and Experience, 35(21):e7701:1–e7701:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lan:2021:SRD

[LXYY21] Sun Lan, Huang Xin, Wu Yingjie, and Guo Yongyi. Sensitivity reduction of degree histogram publication under node differential privacy via mean filtering. Concurrency and Computation: Practice and Experience, 33(8):e5621:1–e5621:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:FSI

[LXZ20] Tonglin Li, Bing Xie, and Boyu Zhang. Foreword to the special issue of the workshop on data-intensive computing in the clouds. Concurrency and Computation: Practice and Experience, 32(16):e5735:1–e5735:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:RIP

[LXZ22] Dongmei Li, Feng Xiao, and Yanjun Zheng. Research on the inheritance and protection of data mining technology in national sports. Concurrency and Computation: Practice and Experience, 34(13):e5893:1–e5893:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2023:PAC

[LYBZ23] Fengwei Li, Qingfang Ye, Hajo Broersma, and Xiaoyan Zhang. Polynomial algorithms for computing the isolated toughness of interval and split graphs. Concurrency and Computation: Practice and Experience, 35(17):e6345:1–e6345:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:NIE

[LYC22] Tu Li, Wang Yan, and Zhang Chi. A new image encryption algorithm based on optimized Lorenz chaotic system. Concurrency and Computation: Practice and Experience, 34(13): e5902:1–e5902:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:TCE

[LYF22] Zengpeng Li, Huiqun Yu, and Guisheng Fan. Time-cost efficient memory configuration for serverless workflow applications. Concurrency and Computation: Practice and Experience, 34(27):e7308:1–e7308:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:AML

[LYF^{+}23] Xingchun Liu, Jingjing Yu, Zhipeng Feng, Hongxv Wang, and Hui Tian. Adaptive multi-layer clustering strategies based on capacity weight for Internet of Things. Concurrency and Computation: Practice and Experience, 35(17):e7243:1– e7243:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2021:IAB

[LYG^{+}21] Rong Lu, Zeyu Yang, Chuyi Gao, Maolong Xi, Yang Zhang, Jian Xiong, Chi-Man Pun, and Hao Gao. An improved artificial bee colony algorithm based on elite search strategy with segmentation application on robot vision system. Concurrency and Computation: Practice and Experience, 33(22): e5745:1–e5745:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2020:RAC

[LYI^{+}20] Yuechao Lu, Ichitaro Yamazaki, Fumihiko Ino, Yasuyuki Matsushita, Stanimire Tomov, and Jack Dongarra. Reducing the amount of out-of-core data access for GPU-accelerated randomized SVD. Concurrency and Computation: Practice and Experience, 32(19):e5754:1–e5754:??, October 10, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2020:VAI

[LYL20] Kyungroul Lee, Kangbin Yim, and Sun-Young Lee. Vulnerability analysis on the image-based authentication: Through the WM INPUT message. Concurrency and Computation: Practice and Experience, 32(18):e5596:1–e5596:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:MMD

[LYL21] Chong Li, Yingjie Yang, and Sifeng Liu. Micro-macro dynamics of the online opinion evolution: an asynchronous network model approach. Concurrency and Computation: Practice and Experience, 33(5):e5981:1–e5981:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:IEB

[LYSC21] Yinghua Li, He Yu, Bin Song, and Jinjun Chen. Image encryption based on a single-round dictionary and chaotic sequences in cloud computing. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:PMR

[LYW^{+}21] Qilei Li, Xiaomin Yang, Wei Wu, Kai Liu, and Gwanggil Jeon. Pansharpening multispectral remote-sensing images with guided filter for monitoring impact of human behavior on environment. Concurrency and Computation: Practice and Experience, 33(4):e5074:1–e5074:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:OSA

[LZC^{+}20] Dan Li, Lichao Zhao, Qingfeng Cheng, Ning Lu, and Wenbo Shi. Opcode sequence analysis of Android malware by a convolutional neural network. Concurrency and Computation: Practice and Experience, 32(18):e5308:1–e5308:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2021:GPU

[LZC21] Min Lin, Yiwen Zhong, and Riqing Chen. Graphic process units-based chicken swarm optimization algorithm for function optimization problems. Concurrency and Computation: Practice and Experience, 33(8):e5953:1–e5953:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:MNM

[LZF20] Bin Li, Jingjuan Zhao, and Hui Fu. MFDC-net: Multifeaturebased deep convolutional neural network for single image haze removal. Concurrency and Computation: Practice and Experience, 32(9):e5631:1–e5631:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2020:IGA

[LZL^{+}20a] Ling Li, Xiangbing Zhou, Yiping Li, Jiangang Gu, and Shaopeng Shen. An improved genetic algorithm with Lagrange and density method for clustering. Concurrency and Computation: Practice and Experience, 32(24):e5969:1– e5969:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2020:DMD

[LZL^{+}20b] Bo Liu, Huaipu Zhao, Yinxing Liu, Suyu Wang, Jianqiang Li, Yong Li, Jianlei Lang, and Rentao Gu. Discovering multidimensional motifs from multi-dimensional time series for air pollution control. Concurrency and Computation: Practice and Experience, 32(11):e5645:1–e5645:??, June 10, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2022:BBT

[LZL^{+}22] Jinjiang Li, Zhihua Zheng, Zhi Li, Ziyu Niu, Hong Qin, and Hao Wang. A blockchain-based traceable group loan system. Concurrency and Computation: Practice and Experience, 34 (16):e5741:1–e5741:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:RTP

[LZL^{+}23] Haoyu Liu, Shiwen Zhang, Mengling Li, Wei Liang, and Voundi Koe Arthur Sandor. A real-time privacy-preserving scheme based on grouping queries for continuous locationbased services. Concurrency and Computation: Practice and Experience, 35(19):e7404:1–e7404:??, August 30, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:FGM

[LZLZ23] Wei Liu, Ying-Hui Zhang, Yi-Fei Li, and Dong Zheng. A fine-grained medical data sharing scheme based on federated learning. Concurrency and Computation: Practice and Experience, 35(20):e6847:1–e6847:??, September 10, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:DDV

[LZM^{+}23] Ji Liu, Xuehai Zhou, Lei Mo, Shilei Ji, Yuan Liao, Zheng Li, Qin Gu, and Dejing Dou. Distributed and deep vertical federated learning with big data. Concurrency and Computation: Practice and Experience, 35(21):e7697:1–e7697:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Long:2022:SLA

[LZQ^{+}22] Xiaojun Long, Jingtao Zhang, Xing Qi, Wenlong Xu, Tianguo Jin, and Kai Zhou. A self-learning artificial bee colony algorithm based on reinforcement learning for a flexible job-shop scheduling problem. Concurrency and Computation: Practice and Experience, 34(4):e6658:1–e6658:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lv:2024:AKD

[LZQF24] Li Lv, Faying Zhang, Shenyu Qiu, and Tanghuai Fan. Adaptive kernel density estimation weighted twin support vector machine and its sample screening method. Concurrency and Computation: Practice and Experience, 36(6):e7958:1– e7958:??, March 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2023:ANA

[LZSC23] Yujun Liu, Xiangchen Zhang, Jinhe Su, and Guorong Cai. AAEE-Net: Attention-guided aggregation and error-aware enhancement network for accurate and efficient stereo matching. Concurrency and Computation: Practice and Experience, 35(22):e7744:1–e7744:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lv:2023:MOF

[LZT^{+}23] Li Lv, Xiaodong Zhou, Dekun Tan, Ping Kang, and Runxiu Wu. Multi-objective firefly algorithm with multi-strategy integration. Concurrency and Computation: Practice and Experience, 35(2):e7496:1–e7496:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2021:MEM

[LZW^{+}21] Yizhi Liu, Chaoqun Zhu, Yadi Wu, Heng Xu, and Jun Song. MMWD: an efficient mobile malicious webpage detection framework based on deep learning and edge cloud. Concurrency and Computation: Practice and Experience, 33(18): e6191:1–e6191:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2022:FSS

[LZXG22] Lin Liu, Jian-Tao Zhou, Hai-Feng Xing, and Xiao-Yong Guo. Flow splitting scheme over link-disjoint multiple paths in software-defined networking. Concurrency and Computation: Practice and Experience, 34(10):e6793:1–e6793:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2020:EIG

[LZY^{+}20] Chao Liu, Deze Zeng, Hong Yao, Xuesong Yan, Linchen Yu, and Zhangjie Fu. An efficient iterative graph data processing framework based on bulk synchronous parallel model. Concurrency and Computation: Practice and Experience, 32(3): e4432:1–e4432:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liu:2021:GCR

[LZY^{+}21] Bo Liu, Yelong Zhao, Bin Yang, Shuangtao Zhao, Rentao Gu, and Mark Gahegan. A gastric cancer recognition algorithm on gastric pathological sections based on multistage attention — DenseNet. Concurrency and Computation: Practice and Experience, 33(10):e6188:1–e6188:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liao:2020:DCS

[LZZ^{+}20] Xiaofei Liao, Long Zheng, Binsheng Zhang, Yu Zhang, Hai Jin, Xuanhua Shi, and Yi Lin. Dynamic cluster strategy for hierarchical rollback-recovery protocols in MPI HPC applications. Concurrency and Computation: Practice and Experience, 32(3):e4173:1–e4173:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:2021:CED

[LZZ21] Yuepeng Li, Deze Zeng, and Long Zheng. On communication efficient dataflow computing in software defined networking enabled cloud. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Liang:2023:EHA

[LZZ^{+}23] Xiubo Liang, Qian Zhao, Yanyu Zhang, Hongyu Liu, and Qifei Zhang. EduChain: a highly available education consortium blockchain platform based on Hyperledger Fabric. Concurrency and Computation: Practice and Experience, 35(18): e6330:1–e6330:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Muralidharan:2022:TCB

[MA22] C. Muralidharan and R. Anitha. Trusted cloud broker for estimating the reputation of cloud providers in federated cloud environment. Concurrency and Computation: Practice and Experience, 34(1):e6537:1–e6537:??, January 10, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mustafa:2022:PLF

[MAAA22] Sidra Mustafa, Muhammad Amin, Muhammad Nauman Akram, and Nimra Afzal. On the performance of link functions in the beta ridge regression model: Simulation and application. Concurrency and Computation: Practice and Experience, 34(18):e7005:1–e7005:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Majid:2023:RKL

[MAAK23] Abdul Majid, Shakeel Ahmad, Muhammad Aslam, and Muhammad Kashif. A robust Kibria–Lukman estimator for linear regression model to combat multicollinearity and outliers. Concurrency and Computation: Practice and Experience, 35(4):e7533:1–e7533:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mohsen:2022:SCR

[MAB22] Fadi Mohsen, Hamed Abdelhaq, and Halil Bisgin. Securitycentric ranking algorithm and two privacy scores to mitigate intrusive apps. Concurrency and Computation: Practice and Experience, 34(14):e6571:1–e6571:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mata:2020:PSS

[MAG^{+}20] Aurelio Nicolás Mata, Graciela Román Alonso, Gabriel López Garza, José Rafael Godinez Fernández, Miguel Alfonso Castro García, and Norma Pilar Castellanos Ábrego. Parallel simulation of the synchronization of heterogeneous cells in the sinoatrial node. Concurrency and Computation: Practice and Experience, 32(10):e5317:1–e5317:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Manoharan:2021:OCR

[Man21] Hariprasath Manoharan. An operative constellation rate for smart safety units using Internet of Things. Concurrency and Computation: Practice and Experience, 33(6):e6085:1– e6085:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marowka:2023:CTP

[Mar23a] Ami Marowka. A comparison of two performance portability metrics. Concurrency and Computation: Practice and Experience, 35(25):e7868:1–e7868:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mashayekhi:2023:MTD

[MAR23b] Hoda Mashayekhi, Elham Sadat Alavi, and Behzad Rezaie. Microblog topic detection using evolutionary clustering and social network information. Concurrency and Computation: Practice and Experience, 35(27):e7847:1–e7847:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mehta:2022:CFO

[MAW^{+}22] Kshitij Mehta, Bryce Allen, Matthew Wolf, Jeremy Logan, Eric Suchyta, Swati Singhal, Jong Y. Choi, Keichi Takahashi, Kevin Huck, Igor Yakushin, Alan Sussman, Todd Munson, Ian Foster, and Scott Klasky. A codesign framework for online data analysis and reduction. Concurrency and Computation: Practice and Experience, 34(14):e6519:1–e6519:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Margolin:2021:TBF

[MB21] Alexander Margolin and Amnon Barak. Tree-based faulttolerant collective operations for MPI. Concurrency and Computation: Practice and Experience, 33(14):e5826:1–e5826:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Milli:2022:SOS

[MB22] Musa Milli and Hasan Bulut. SubtStream: Online subtractive stream clustering algorithm. Concurrency and Computation: Practice and Experience, 34(15):e6968:1–e6968:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Malouche:2022:NFC

[MBB22] Hela Malouche, Youssef Ben Halima, and Henda Ben Ghezala. A negotiation framework for the cloud using rough set theorybased preference prediction. Concurrency and Computation: Practice and Experience, 34(22):e7149:1–e7149:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mankodi:2023:PPS

[MBC23] Amit Mankodi, Amit Bhatt, and Bhaskar Chaudhury. Performance prediction from simulation systems to physical systems using machine learning with transfer learning and scaling. Concurrency and Computation: Practice and Experience, 35 (18):e6433:1–e6433:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Machaj:2020:CIH

[MBM20a] Juraj Machaj, Peter Brida, and Norbert Majer. Challenges introduced by heterogeneous devices for Wi-Fi-based indoor localization. Concurrency and Computation: Practice and Experience, 32(13):e5198:1–e5198:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mahmoudi:2020:MPU

[MBM^{+}20b] Sidi Ahmed Mahmoudi, Mohammed Amin Belarbi, Saïd Mahmoudi, Ghalem Belalem, and Pierre Manneback. Multimedia processing using deep learning technologies, high-performance computing cloud resources, and Big Data volumes. Concurrency and Computation: Practice and Experience, 32(17): e5699:1–e5699:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mokhtari:2021:MFT

[MBO^{+}21] Karima Aslaoui Mokhtari, Salima Benbernou, Mourad Ouziri, Hakim Lahmar, and Muhammad Younas. A monitoring framework for transparency and fairness in big data platform. Concurrency and Computation: Practice and Experience, 33(23):e6069:1–e6069:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mustafee:2020:CCA

[MBT^{+}20] Navonil Mustafee, Nik Bessis, Simon J. E. Taylor, Jianhua Hou, and Peter Matthew. Co-citation analysis of literature in e-science and e-infrastructures. Concurrency and Computation: Practice and Experience, 32(9):e5620:1–e5620:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mohammed:2020:SSA

[MC20] Ali Mohammed and Florina M. Ciorba. SimAS: a simulationassisted approach for the scheduling algorithm selection under perturbations. Concurrency and Computation: Practice and Experience, 32(15):e5648:1–e5648:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Munhoz:2024:EEH

[MC24] Vanderlei Munhoz and Márcio Castro. Enabling the execution of HPC applications on public clouds with HPC@Cloud toolkit. Concurrency and Computation: Practice and Experience, 36(8):e7976:1–e7976:??, April 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McLean:2023:BPC

[MCD^{+}23] Jared McLean, Sean B. Cleveland, Michael Dodge II, Matthew P. Lucas, Ryan J. Longman, Thomas W. Giambelluca, and Gwen A. Jacobs. Building a portal for climate datamapping automation, visualization, and dissemination. Concurrency and Computation: Practice and Experience, 35(18): e6727:1–e6727:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Morganti:2020:LPS

[MCL^{+}20] Lucia Morganti, Elena Corni, Luca Lama, Carmelo Pellegrino, Francieli Zanon Boito, Ivan Merelli, Daniele D’Agostino, and Daniele Cesini. On low-power SoCs as storage bricks for bioinformatics. Concurrency and Computation: Practice and Experience, 32(10):e5415:1–e5415:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Munoz-Caro:2020:FGL

[MCNR20] Camelia Muñoz-Caro, Alfonso Niño, and Sebastián Reyes. A fine-grained loop-level parallel approach to efficient fuzzy community detection in complex networks. Concurrency and Computation: Practice and Experience, 32(5):e5537:1– e5537:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McLean:2023:HGR

[MCR^{+}23] Jared McLean, Sean B. Cleveland, Kolja Rotzoll, Scot K. Izuka, Jason Leigh, Gwen A. Jacobs, and Ryan Theriot. The Hawai‘i Groundwater Recharge Tool. Concurrency and Computation: Practice and Experience, 35(18):e6843:1–e6843:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mandal:2022:DNN

[MCT22a] Dipankar Mandal, Arpitam Chatterjee, and Bipan Tudu. A deep neural network and random forests driven computer vision framework for identification and prediction of metanil yellow adulteration in turmeric powder. Concurrency and Computation: Practice and Experience, 34(1):e6500:1– e6500:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mukherjee:2022:ITD

[MCT22b] Gunjan Mukherjee, Arpitam Chatterjee, and Bipan Tudu. Identification of the types of disease for tomato plants using a modified gray wolf optimization optimized MobileNetV2 convolutional neural network architecture driven computer vision framework. Concurrency and Computation: Practice and Experience, 34(22):e7161:1–e7161:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mello:2021:GGA

[MdAA^{+}21] Rui R. Mello, Jr., Leandro S. de Araújo, Tiago A. O. Alves, Leandro A. J. Marzulo, Gabriel A. L. Paillard, and Felipe M. G. França. Gamma — General Abstract Model for Multiset mAnipulation and dynamic dataflow model: an equivalence study. Concurrency and Computation: Practice and Experience, 33(11):e6176:1–e6176:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moori:2023:IEG

[MdARS^{+}23] Marcelo K. Moori, Hiago Mayk G. de A. Rocha, Janaina Schwarzrock, Arthur F. Lorenzon, and Antonio Carlos S. Beck. Improving the efficiency of graph algorithm executions on high-performance computing. Concurrency and Computation: Practice and Experience, 35(18):e7419:1–e7419:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mellone:2023:NAL

[MDSG^{+}23] Gennaro Mellone, Ciro Giuseppe De Vita, Dante D. Sánchez-Gallegos, Genaro Sánchez-Gallegos, Catherine A. Torres-Charles, Javier Garcia-Blas, Jesús Carretero Pérez, J. L. Gonzalez-Compean, and Giuliano Laccetti. A novel approach for large-scale environmental data partitioning on cloud and on-premises storage for compute continuum applications. Concurrency and Computation: Practice and Experience, 35(25):e7893:1–e7893:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mathappan:2023:HII

[MES23] Nivaashini Mathappan, Suganya Elavarasan, and Sountharrajan Sehar. Hybrid intelligent intrusion detection system for multiple Wi-Fi attacks in wireless networks using stacked restricted Boltzmann machine and deep belief networks. Concurrency and Computation: Practice and Experience, 35(23): e7769:1–e7769:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Muller:2021:ADC

[MFA^{+}21] Juliane Müller, Boris Faybishenko, Deborah Agarwal, Stephen Bailey, Chongya Jiang, Youngryel Ryu, Craig Tull, and Lavanya Ramakrishnan. Assessing data change in scientific datasets. Concurrency and Computation: Practice and Experience, 33(16):e6245:1–e6245:??, August 25, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maamar:2023:STI

[MFE^{+}23] Zakaria Maamar, Noura Faci, Said Elnaffar, Fadwa Yahya, Khouloud Boukadi, and Djamal Benslimane. Semantic thingsourcing for the Internet of Things. Concurrency and Computation: Practice and Experience, 35(11):e6684:1–e6684:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mamta:2020:EKD

[MG20] Mamta and B. B. Gupta. An efficient KP design framework of attribute-based searchable encryption for user level revocation in cloud. Concurrency and Computation: Practice and Experience, 32(18):e5291:1–e5291:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mirtaheri:2021:OLB

[MG21a] Seyedeh Leili Mirtaheri and Lucio Grandinetti. Optimized load balancing in high-performance computing for big data analytics. Concurrency and Computation: Practice and Experience, 33(16):e6265:1–e6265:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mohan:2021:NIM

[MG21b] Vijay Mohan and Indumathi Ganesan. A nature-inspired meta-heuristic paradigm for person identification using multimodal biometrics. Concurrency and Computation: Practice and Experience, 33(21):e6420:1–e6420:??, November 10, 2021.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Manikandan:2022:SSP

[MG22] S. Periasamy Manikandan and Shanmugam Milton Ganesh. SPPGKM: a secure polynomial function powered group key management scheme for dynamic user environments in cloud. Concurrency and Computation: Practice and Experience, 34(21):e7053:1–e7053:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mudasser:2023:SIT

[MG23] Abdul Wasay Mudasser and Shah Aqueel Ahmed Abdul Gafoor. Secure Internet of Things based hybrid optimization techniques for optimal centroid routing protocol in wireless sensor network. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mirebeau:2023:MPC

[MGB^{+}23] Jean-Marie Mirebeau, Lionel Gayraud, Remi Barrère, Da Chen, and François Desquilbet. Massively parallel computation of globally optimal shortest paths with curvature penalization. Concurrency and Computation: Practice and Experience, 35(2):e7472:1–e7472:??, January 25, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mesfin:2020:RFE

[MGGA20] Gebremariam Mesfin, Gheorghita Ghinea, Tor-Morten Grønli, and Sahel Alouneh. REST4Mobile: a framework for enhanced usability of REST services on smartphones. Concurrency and Computation: Practice and Experience, 32 (1):e4174:1–e4174:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mu:2022:IPS

[MGN^{+}22] Rui Mu, Bei Gong, Zhenhu Ning, Jiangjiang Zhang, Yang Cao, Zheng Li, Wei Wang, and Xiaoping Wang. An identity privacy scheme for blockchain-based on edge computing. Concurrency and Computation: Practice and Experience, 34 (1):e6545:1–e6545:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Martinasso:2020:CPE

[MGS^{+}20] Maxime Martinasso, Miguel Gila, William Sawyer, Rafael Sarmiento, Guilherme Peretti-Pezzi, and Vasileios Karakasis. Cray programming environments within containers on Cray XC systems. Concurrency and Computation: Practice and Experience, 32(20):e5543:1–e5543:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Majji:2023:HOB

[MGSB23] Ramachandro Majji, Om Prakash P. G, Vairamuthu S, and Selva Rani B. Hybrid optimization based deep neuro fuzzy network for skin cancer detection. Concurrency and Computation: Practice and Experience, 35(3):e7521:1–e7521:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mudrakola:2023:RNM

[MH23] Swapna Mudrakola and Nagaratna Hegde. Removal of noise on mammogram breast images using filtering methods. Concurrency and Computation: Practice and Experience, 35(1): e7444:1–e7444:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Malti:2022:MOT

[MHB22a] Arslan Nedhir Malti, Mourad Hakem, and Badr Benmammar. Multi-objective task scheduling in cloud computing. Concurrency and Computation: Practice and Experience, 34(25): e7252:1–e7252:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moharamkhani:2022:IDS

[MHB^{+}22b] Elaheh Moharamkhani, Mohammad Yahyaei Feriz Hendi, Eisa Bandar, Amir Izadkhasti, and Rzgar Sirwan Raza. Intrusion detection system based firefly algorithm-random forest for cloud computing. Concurrency and Computation: Practice and Experience, 34(24):e7220:1–e7220:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2020:ESS

[MHL^{+}20] Hongzhong Ma, Chaozhi Huang, Xiping Liu, Wangfeng Shi, and Wei Liu. The effect of a single-sided pole shoe and slot on reducing torque ripple in a switched reluctance motor. Concurrency and Computation: Practice and Experience, 32 (19):e5810:1–e5810:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Millar:2021:OUS

[MHPA21] Jeremy Robert Millar, Douglas Hodson, Gilbert Peterson, and Darryl Ahner. Optimizing update scheduling parameters for distributed virtual environments supporting operational test. Concurrency and Computation: Practice and Experience, 33 (8):e4156:1–e4156:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Matsumura:2023:NSS

[MIN^{+}23] Naoki Matsumura, Yasuaki Ito, Koji Nakano, Akihiko Kasagi, and Tsuguchika Tabaru. A novel structured sparse fully connected layer in convolutional neural networks. Concurrency and Computation: Practice and Experience, 35(11):e6213:1– e6213:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mirzaei:2022:NAQ

[Mir22] Abbas Mirzaei. A novel approach to QoS-aware resource allocation in NOMA cellular HetNets using multi-layer optimization. Concurrency and Computation: Practice and Experience, 34(21):e7068:1–e7068:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mishra:2022:HFE

[Mis22] Ravi Mishra. Hybrid feature extraction and optimized deep convolutional neural network based video shot boundary detection. Concurrency and Computation: Practice and Experience, 34(25):e7256:1–e7256:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mittal:2020:SEO

[Mit20] Sparsh Mittal. A survey on evaluating and optimizing performance of Intel Xeon Phi. Concurrency and Computation: Practice and Experience, 32(19):e5742:1–e5742:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mishra:2023:EPP

[MJAM23] Alok Mishra, Thr Satar Jabar, Yehia Ibrahim Alzoubi, and Kamta Nath Mishra. Enhancing privacy-preserving mechanisms in cloud storage: a novel conceptual framework. Concurrency and Computation: Practice and Experience, 35(26): e7831:1–e7831:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Malathi:2022:NOI

[MK22a] Narra Malathi and Molugaram Kumar. A novel optimized intelligent green signal timing system for urban corridor. Concurrency and Computation: Practice and Experience, 34(26): e7284:1–e7284:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Manocha:2022:ISM

[MK22b] Prabhjot Singh Manocha and Rajiv Kumar. Improved spider monkey optimization-based multi-objective software-defined networking routing with block chain technology for Internet of Things security. Concurrency and Computation: Practice and Experience, 34(11):e6861:1–e6861:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mavridis:2023:OSC

[MK23] Ilias Mavridis and Helen Karatza. Orchestrated sandboxed containers, unikernels, and virtual machines for isolationenhanced multitenant workloads and serverless computing in cloud. Concurrency and Computation: Practice and Experience, 35(11):e6365:1–e6365:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Meenakshi:2023:BEJ

[MKA23] B. Meenakshi, D. Karunkuzhali, and Syed Muqthadar Ali. Blockchain-enabled joint trust (RF-AOA-HTSA) algorithmbased multiobjective clustering protocol for wireless sensor network. Concurrency and Computation: Practice and Experience, 35(28):e7860:1–e7860:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mechti:2022:DSC

[MKBB22] Seifeddine Mechti, Moez Krichen, Dhouha Ben Noureddine, and Lamia H. Belguith. A decision system for computational authors profiling: From machine learning to deep learning. Concurrency and Computation: Practice and Experience, 34 (7):e5985:1–e5985:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Varalakshmi:2021:HPI

[MKL21] Varalakshmi M, Amit P. Kesarkar, and Daphne Lopez. Highperformance implementation of a two-bit geohash coding technique for nearest neighbor search. Concurrency and Computation: Practice and Experience, 33(5):e6029:1–e6029:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mahendran:2023:DAP

[MKRK23] P. Mahendran, M. S. Kavitha, R. Radhika, and C. Kotteeswaran. Design of all pass make over based capricious digital filter using eminent speed dual carry select adder and truncation and rounding approximate multiplier for image processing application. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moin:2020:GSE

[MKS^{+}20] Sana Moin, Ahmad Karim, Kalsoom Safdar, Iqra Iqbal, Zanab Safdar, V. Vijayakumar, K. T. Ahmed, and S. A. Abid. GREEN SDN — an enhanced paradigm of SDN: Review, taxonomy, and future directions. Concurrency and Computation: Practice and Experience, 32(21):e5086:1–e5086:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mao:2020:COD

[ML20] Bo Mao and Bingchan Li. City object detection from airborne Lidar data with OpenStreetMap-tagged superpixels. Concurrency and Computation: Practice and Experience, 32(23): e6026:1–e6026:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mun:2021:DRS

[MLC^{+}21] Jonghyeok Mun, Sanghwan Lee, Jongsun Choi, Jaeyoung Choi, and Kitae Bae. Dataset retrieval system based on automation of data preparation with dataset description model. Concurrency and Computation: Practice and Experience, 33 (2):e5288:1–e5288:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Merelli:2020:LAP

[MLKD20] Ivan Merelli, Pietro Liò, Igor Kotenko, and Daniele D’Agostino. Latest advances in parallel, distributed, and network-based processing. Concurrency and Computation: Practice and Experience, 32(10):e5683:1–e5683:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2024:PPG

[MLS^{+}24] Liya Ma, Chen Li, Siqi Sun, Shu Guo, Lihong Wang, and Jianxin Li. Privacy-preserving graph publishing with disentangled variational information bottleneck. Concurrency and Computation: Practice and Experience, 36(10):e7963:1– e7963:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mikkelsen:2024:NMR

[MLVGR24] Carl Christian Kjelgaard Mikkelsen, Lorién López-Villellas, and Pablo García-Risueño. Newton’s method revisited: How accurate do we have to be? Concurrency and Computation: Practice and Experience, 36(10):e7853:1–e7853:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Meng:2020:TAS

[MLZ^{+}20] Shunmei Meng, Qianmu Li, Jing Zhang, Wenmin Lin, and Wanchun Dou. Temporal-aware and sparsity-tolerant hybrid collaborative recommendation method with privacy preservation. Concurrency and Computation: Practice and Experience, 32(2):e5447:1–e5447:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ma:2021:MSC

[MLZ^{+}21a] Shuangxun Ma, Yuehu Liu, Qinghai Zheng, Yaochen Li, and Zhichao Cui. Multiview spectral clustering via complementary information. Concurrency and Computation: Practice and Experience, 34(1):e5701:1–e5701:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mo:2021:KET

[MLZ21b] Hongqiang Mo, Zhong Li, and Chunmei Zhu. A kind of epistasis-tunable test functions for genetic algorithms. Concurrency and Computation: Practice and Experience, 33(8): e5030:1–e5030:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Myllykoski:2021:TBG

[MM21] Mirko Myllykoski and Carl Christian Kjelgaard Mikkelsen. Task-based, GPU-accelerated and robust library for solving dense nonsymmetric eigenvalue problems. Concurrency and Computation: Practice and Experience, 33(11):e5915:1– e5915:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mishra:2022:RSR

[MM22] Aditya Dev Mishra and Khurram Mustafa. A review on security requirements specification by formal methods. Concurrency and Computation: Practice and Experience, 34(5): e6702:1–e6702:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mahenthiran:2024:HPA

[MM24] Prabha Mahenthiran and Dinakaran Muruganadam. Handoff performance analysis of PMIPv6-based distributed mobility management protocol-urban scenario. Concurrency and Computation: Practice and Experience, 36(11):e8010:1–e8010:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mukhtar:2023:EED

[MMKA23] Naila Mukhtar, Ali Mehrabi, Yinan Kong, and Ashiq Anjum. Edge enhanced deep learning system for IoT edge device security analytics. Concurrency and Computation: Practice and Experience, 35(13):e6764:1–e6764:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mazidi:2021:ADT

[MMR21] Arash Mazidi, Mehregan Mahdavi, and Fahimeh Roshanfar. An autonomic decision tree-based and deadline-constraint resource provisioning in cloud applications. Concurrency and Computation: Practice and Experience, 33(10):e6196:1– e6196:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Melingi:2022:BIA

[MMR^{+}22] Sunil Babu Melingi, Ramesh Kumar Mojjada, Vanga Karunakar Reddy, Ch. V. M. S. N. Pavan Kumar, and K. Ashok Kumar. A bio-inspired AlexNet-DrpXLm archetype for an effective brain stroke lesion detection and classification. Concurrency and Computation: Practice and Experience, 34(21):e7100:1– e7100:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mishra:2024:CTB

[MMS24] Ashis Kumar Mishra, Subasish Mohapatra, and Pradip Kumar Sahu. Coalition theory based task scheduling algorithm using DLFC-NN model. Concurrency and Computation: Practice and Experience, 36(10):e8005:1–e8005:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mareddy:2022:OSM

[MNDK22] Padma Lalitha Mareddy, Sivarami Reddy Narapureddy, Venkata Ramamurthy Dwivedula, and Prahlada Rao Karanam. Optimum scheduling of machines, automated guided vehicles and tools without tool delay in a multi-machine flexible manufacturing system using symbiotic organisms search algorithm. Concurrency and Computation: Practice and Experience, 34 (15):e6950:1–e6950:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mahajan:2022:GPB

[MNR^{+}22] Shubham Mahajan, Anand Nayyar, Akshay Raina, Samreen J. Singh, Ashutosh Vashishtha, and Amit Kant Pandit. A Gaussian process-based approach toward credit risk modeling using stationary activations. Concurrency and Computation: Practice and Experience, 34(5):e6692:1–e6692:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Martinez-Noriega:2021:COE

[MNYN21] Edgar Josafat Martinez-Noriega, Syunji Yazaki, and Tetsu Narumi. CUDA offloading for energy-efficient and high-frame-rate simulations using tablets. Concurrency and Computation: Practice and Experience, 33(2):e5488:1–e5488:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Montazerolghaem:2021:OVS

[Mon21] Ahmadreza Montazerolghaem. Optimizing VoIP server resources using linear programming model and autoscaling technique: an SDN approach. Concurrency and Computation: Practice and Experience, 33(21):e6424:1–e6424:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Murugaveni:2023:LEN

[MP23] S. Murugaveni and B. Priyalakshmi. Layering of edge node for jamming attack detection and elimination in wireless sensor network. Concurrency and Computation: Practice and Experience, 35(22):e7737:1–e7737:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Martinez:2022:AIO

[MPB^{+}22] Pablo Antonio Martínez, Biagio Peccerillo, Sandro Bartolini, José Manuel García, and Gregorio Bernabé. Applying Intel’s oneAPI to a machine learning case study. Concurrency and Computation: Practice and Experience, 34(13):e6917:1– e6917:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Menezo:2020:RCC

[MPG20] Lucia G. Menezo, Valentin Puente, and Jose A Gregorio. Rainbow: a composable coherence protocol for multi-chip servers. Concurrency and Computation: Practice and Experience, 32(24):e5947:1–e5947:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Michels:2024:ENF

[MPLW24] Alexander Michels, Anand Padmanabhan, Zhiyu Li, and Shaowen Wang. EasyScienceGateway: a new framework for providing reproducible user environments on science gateways. Concurrency and Computation: Practice and Experience, 36(4):e7929:1–e7929:??, February 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Merugula:2022:ROH

[MPV22] Suneetha Merugula, Buddi Padmaja, and Ragavi Veerubommu. RideNN-OptDRN: Heart disease detection using RideNN based feature fusion and optimized deep residual network. Concurrency and Computation: Practice and Experience, 34(28):e7355:1–e7355:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Manhrawy:2021:HFS

[MQEK21] Ibrahim I. M. Manhrawy, Mohammed Qaraad, and Passent El-Kafrawy. Hybrid feature selection model based on reliefbased algorithms and regulizer algorithms for cancer classification. Concurrency and Computation: Practice and Experience, 33(17):e6200:1–e6200:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pavithra:2023:OMS

[MR23a] Pavithra M and Parvathi Rms. Optimizing minimum spanning tree using stochastic-variable neighborhood search for efficient clustering of cancer gene data. Concurrency and Computation: Practice and Experience, 35(5):e7573:1–e7573:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Madireddy:2023:DVM

[MR23b] Amarendhar Reddy Madireddy and Kongara Ravindranath. Dynamic virtual machine relocation system for energyefficient resource management in the cloud. Concurrency and Computation: Practice and Experience, 35(3):e7520:1– e7520:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moreira:2021:NFM

[MRAM^{+}21] Mário W. L. Moreira, Joel J. P. C. Rodrigues, Jalal Al-Muhtadi, Valery V. Korotaev, and Victor Hugo C. de Albuquerque. Neuro-fuzzy model for HELLP syndrome prediction in mobile cloud computing environments. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mitra:2023:RHT

[MRAS^{+}23] Saptashwa Mitra, Maxwell Roselius, Pedro Andrade-Sanchez, John K. McKay, and Sangmi Lee Pallickara. Radix+: Highthroughput georeferencing and data ingestion over voluminous and fast-evolving phenotyping sensor data. Concurrency and Computation: Practice and Experience, 35(8):e7484:1– e7484:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mahajan:2022:CDU

[MRGP22] Shubham Mahajan, Akshay Raina, Xiao-Zhi Gao, and Amit Kant Pandit. COVID-19 detection using hybrid deep learning model in chest X-rays images. Concurrency and Computation: Practice and Experience, 34(5):e6747:1– e6747:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mukoya:2023:ADL

[MRK^{+}23] Esther Mukoya, Richard Rimiru, Michael Kimwele, Consolata Gakii, and Grace Mugambi. Accelerating deep learning inference via layer truncation and transfer learning for fingerprint classification. Concurrency and Computation: Practice and Experience, 35(8):e7619:1–e7619:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mohana:2022:TSA

[MRKY22] R. S. Mohana, K. Rajathi, K. Kousalya, and T. Yuvaraja. Text sentiment analysis on e-shopping product reviews using chaotic coyote optimized deep belief network approach. Concurrency and Computation: Practice and Experience, 34(19): e7039:1–e7039:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mohan:2021:DNN

[MRS^{+}21] Ellappan Mohan, Arunachalam Rajesh, Gurram Sunitha, Reddy Madhavi Konduru, Janagaraj Avanija, and Loganathan Ganesh Babu. A deep neural network learningbased speckle noise removal technique for enhancing the quality of synthetic-aperture radar images. Concurrency and Computation: Practice and Experience, 33(13):e6239:1– e6239:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maniam:2021:AEH

[MS21a] Senthil Murugan Maniam and T. Sasilatha. Area-efficient and high-speed hardware structure of hybrid cryptosystem (AES-RC4) for maximizing key lifetime using parallel subpipeline architecture. Concurrency and Computation: Practice and Experience, 33(3):e5287:1–e5287:??, February 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Manju:2021:DDS

[MS21b] Anamala Balaji Manju and Subramanian Sumathy. Dispersed dummy selection approach for location-based services to preempt user-profiling. Concurrency and Computation: Practice and Experience, 33(20):e6361:1–e6361:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mehta:2021:LBN

[MS21c] Deepak Mehta and Sharad Saxena. Load-based node ranked low-energy adaptive clustering hierarchy: an enhanced energy-efficient algorithm for cluster head selection in wireless sensor networks. Concurrency and Computation: Practice and Experience, 33(21):e6416:1–e6416:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amaresh:2022:STF

[MS22a] Amaresh M and Chitrakala S. Spatio-temporal feature fusion based correlative binary relevance for visual object detection. Concurrency and Computation: Practice and Experience, 34 (5):e6709:1–e6709:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Meesala:2022:FBO

[MS22b] Shobha Rani Meesala and Sumathy Subramanian. Feature based opinion analysis on social media tweets with association rule mining and multi-objective evolutionary algorithms. Concurrency and Computation: Practice and Experience, 34 (3):e6586:1–e6586:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Modi:2023:ERU

[MS23] Tejas M. Modi and Pravati Swain. Enhanced routing using recurrent neural networks in software defined-data center network. Concurrency and Computation: Practice and Experience, 35(5):e7557:1–e7557:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moslemi:2021:CVB

[MSA21] Negar Moslemi, Mohsen Soryani, and Reza Azmi. Computer vision-based recognition of driver distraction: a review. Concurrency and Computation: Practice and Experience, 33(24): e6475:1–e6475:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mushtaq:2022:ECR

[MSA22] Nadia Mushtaq, Iram Saleem, and Mustansar Aatizaz Amjad. Evaluating COVID-19 risk under the estimation of population mean using two attributes. Concurrency and Computation: Practice and Experience, 34(28):e7386:1–e7386:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mu:2020:IHS

[MSB^{+}20] Jingqing Mu, Jerome Soumagne, Suren Byna, Quincey Koziol, Houjun Tang, and Richard Warren. Interfacing HDF5 with a scalable object-centric storage system on hierarchical storage. Concurrency and Computation: Practice and Experience, 32(20):e5715:1–e5715:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Musthafa:2023:HML

[MSBR23] A. Syed Musthafa, K. Sankar, T. Benil, and Yamarthi Narasimha Rao. A hybrid machine learning technique for early prediction of lung nodules from medical images using a learning-based neural network classifier. Concurrency and Computation: Practice and Experience, 35(3):e7488:1–e7488:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Manoharan:2022:OML

[MSJ22] Saravanapriya Manoharan, Radha Senthilkumar, and Saktheeswaran Jayakumar. Optimized multi-label convolutional neural network using modified genetic algorithm for popularity based personalized news recommendation system. Concurrency and Computation: Practice and Experience, 34(19): e7033:1–e7033:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Miglani:2022:MOR

[MSK22] Neha Miglani, Gaurav Sharma, and Savita Khurana. Multiobjective reliability-based workflow scheduler: an elastic and persuasive task scheduler based upon modified-flower pollination algorithm in cloud environment. Concurrency and Computation: Practice and Experience, 34(22):e7150:1–e7150:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mumin:2022:EAL

[MSL22] Diyawu Mumin, Lei-Lei Shi, and Lu Liu. An efficient algorithm for link prediction based on local information: Considering the effect of node degree. Concurrency and Computation: Practice and Experience, 34(7):e6289:1–e6289:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Maheswari:2022:HSC

[MSN22] K. G. Maheswari, C. Siva, and G. Nalinipriya. A hybrid soft computing technique for intrusion detection in web and cloud environment. Concurrency and Computation: Practice and Experience, 34(22):e7046:1–e7046:??, October 10, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McIntosh-Smith:2020:BFG

[MSPPD20] Simon McIntosh-Smith, James Price, Andrei Poenaru, and Tom Deakin. Benchmarking the first generation of production quality Arm-based supercomputers. Concurrency and Computation: Practice and Experience, 32(20):e5569:1–e5569:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Musa:2020:SMS

[MSS^{+}20] Ahmed Musa, Haythem Bany Salameh, Nusseibeh Abu Sannad, Rami Halloush, and Khalid Darabkh. Spectrum management with simultaneous power-controlled assignment decisions in cognitive radio networks. Concurrency and Computation: Practice and Experience, 32(21):e5224:1–e5224:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Medara:2022:ECA

[MSS22] Rambabu Medara, Ravi Shankar Singh, and Mahesh Sompalli. Energy and cost aware workflow scheduling in clouds with deadline constraint. Concurrency and Computation: Practice and Experience, 34(13):e6922:1–e6922:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Malla:2024:EES

[MSSM24] Parvaz Ahmad Malla, Sophiya Sheikh, Mohammad Shahid, and Sheikh Umar Mushtaq. Energy-efficient sender-initiated threshold-based load balancing (e-STLB) in cloud computing environment. Concurrency and Computation: Practice and Experience, 36(5):e7943:1–e7943:??, February 29, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mahil:2021:CPS

[MT21] Mahil M and Jayasree T. Combined particle swarm optimization and Ant Colony System for energy efficient cloud data centers. Concurrency and Computation: Practice and Experience, 33(10):e6195:1–e6195:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Misic:2022:CPC

[MT22] Marko J. Misić and Milo V. Tomasević. Comparison of parallel central processing unit- and graphics processing unit-based implementations of greedy string tiling algorithm for source code plagiarism detection. Concurrency and Computation:

Practice and Experience, 34(21):e7135:1–e7135:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marotta:2020:MLC

[MTD^{+}20] Romolo Marotta, Davide Tiriticco, Pierangelo Di Sanzo, Alessandro Pellegrini, Bruno Ciciani, and Francesco Quaglia. Mutable locks: Combining the best of spin and sleep locks. Concurrency and Computation: Practice and Experience, 32(22):e5858:1–e5858:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Matsumura:2021:TAI

[MTK^{+}21] Naoki Matsumura, Hiroki Tokura, Yuki Kuroda, Yasuaki Ito, and Koji Nakano. Tile art image generation using parallel greedy algorithm on the GPU and its approximation with machine learning. Concurrency and Computation: Practice and Experience, 33(12):e5623:1–e5623:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Monniot:2023:SDA

[MTRA23] Julien Monniot, François Tessier, Matthieu Robert, and Gabriel Antoniu. Supporting dynamic allocation of heterogeneous storage resources on HPC systems. Concurrency and Computation: Practice and Experience, 35(28):e7890:1– e7890:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mahmood:2022:CML

[MTSU22] Atif Mahmood, Amod Kumar Tiwari, Sanjay Kumar Singh, and Sandeep S. Udmale. Contemporary machine learning applications in agriculture: Quo vadis? Concurrency and Computation: Practice and Experience, 34(15):e6940:1–e6940:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marozzo:2020:SWT

[MTT20] Fabrizio Marozzo, Domenico Talia, and Paolo Trunfio. A sleep-and-wake technique for reducing energy consumption in BitTorrent networks. Concurrency and Computation: Practice and Experience, 32(14):e5723:1–e5723:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mostafa:2021:EAD

[MTY21] Fahed Mostafa, Longquan Tao, and Wenjin Yu. An effective architecture of digital twin system to support human decision making and AI-driven autonomy. Concurrency and Computation: Practice and Experience, 33(19):e6111:1–e6111:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Malavika:2022:LBB

[MV22] Rajadurai Malavika and Muniappan Lakshapalam Valarmathi. Load Balancing Based on Closed Loop Control Theory (LBBCLCT): a Software Defined Networking (SDN) powered server load balancing system based on closed loop control theory. Concurrency and Computation: Practice and Experience, 34(11):e6854:1–e6854:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Muniswamy:2024:JOL

[MV24] Saravanan Muniswamy and Radhakrishnan Vignesh. Joint optimization of load balancing and resource allocation in cloud environment using optimal container management strategy. Concurrency and Computation: Practice and Experience, 36(12):e8035:1–e8035:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mangayarkarasi:2023:RMT

[MVR23] R. Mangayarkarasi, C. Vanmathi, and Vinayakumar Ravi. A robust malware traffic classifier to combat security breaches in industry 4.0 applications. Concurrency and Computation: Practice and Experience, 35(23):e7772:1–e7772:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Murray:2021:DAM

[MW21] Charles D. Murray and Tobias Weinzierl. Delayed approximate matrix assembly in multigrid with dynamic precisions. Concurrency and Computation: Practice and Experience, 33 (11):e5941:1–e5941:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mao:2023:TPR

[MWH^{+}23] Kuang Mao, Sai Wu, Jiajia He, Haichao Huang, Yanlong Yin, and Zujie Ren. Textile pattern recommendations with convolutional neural networks and autoencoder. Concurrency and Computation: Practice and Experience, 35(18):e6113:1– e6113:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Marts:2023:DPI

[MWS^{+}23] W. Pepper Marts, Andrew Worley, Prema Soundarajan, Derek Schafer, Matthew G. F. Dosanjh, Ryan E. Grant, Purushotham V. Bangalore, Anthony Skjellum, and Sheikh Ghafoor. Design of a portable implementation of partitioned point-to-point communication primitives. Concurrency and Computation: Practice and Experience, 35(20):e7655:1– e7655:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Menguc:2024:IMP

[MY24] Kenan Menguc and Alper Yilmaz. Improving model performance of shortest-path-based centrality measures in network models through scale space. Concurrency and Computation: Practice and Experience, 36(14):e8082:1–e8082:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Meng:2022:VCN

[MYCH22] Lingpeng Meng, Xiaobo Yao, Qian Chen, and Chuanfeng Han. Vulnerability cloud: a novel approach to assess the vulnerability of critical infrastructure systems. Concurrency and Computation: Practice and Experience, 34(21):e7131:1– e7131:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

N:2022:EAB

[NA22a] Saravanan N and Umamakeswari A. Enhanced attribute based encryption technique for secured access in cloud storage for personal health records. Concurrency and Computation: Practice and Experience, 34(11):e6890:1–e6890:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nanjappan:2022:HBN

[NA22b] Manikandan Nanjappan and Pravin Albert. Hybrid-based novel approach for resource scheduling using MCFCM and PSO in cloud computing environment. Concurrency and Computation: Practice and Experience, 34(7):e5517:1– e5517:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nazir:2022:DEE

[NA22c] Babar Nazir and Zulfiqar Ahmad. Dynamic energy efficient load balancing strategy for computational grid. Concurrency and Computation: Practice and Experience, 34(1):e6484:1– e6484:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Natesan:2022:OTT

[NAK^{+}22] Gobalakrishnan Natesan, Javid Ali, Pradeep Krishnadoss, Raman Chidambaram, and Manikandan Nanjappan. Optimization techniques for task scheduling criteria in IaaS cloud computing atmosphere using nature inspired hybrid spotted hyena optimization algorithm. Concurrency and Computation: Practice and Experience, 34(24):e7228:1–e7228:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Naresh:2022:PSS

[NAR^{+}22] Vankamamidi S. Naresh, V. V. L. Divakar Allavarpu, Sivaranjani Reddi, Pilla Sita Rama Murty, N. V. S. Lakshmipathi Raju, and R. N. V. Jagan Mohan. A provably secure sharding based blockchain smart contract centric hierarchical group key agreement for large wireless ad-hoc networks. Concurrency and Computation: Practice and Experience, 34 (3):e6553:1–e6553:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Notash:2022:EEF

[NBHN22] Anaram Yaghoobi Notash, Peyman Bayat, Shahpar Haghighat, and Ali Yaghoobi Notash. Evolutionary ensemble feature selection learning for image-based assessment of lymphedema arm volume. Concurrency and Computation: Practice and Experience, 34(1):e6334:1–e6334:??, January 10, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nachimuthu:2022:MSW

[NBK22] Deepa Subramaniam Nachimuthu, Abhik Banerjee, and Jayakumar Karuppaiah. Multi-step wind speed and wind power forecasting using variational momentum factor and deep learning based intelligent neural network models. Concurrency and Computation: Practice and Experience, 34(6): e6772:1–e6772:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nalinipriya:2022:RRB

[NBPR22] Ganapathi Nalinipriya, Maram Balajee, Chittibabu Priya, and Cristin Rajan. Ransomware recognition in blockchain network using water moth flame optimization-aware DRNN. Concurrency and Computation: Practice and Experience, 34 (19):e7047:1–e7047:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nagaraju:2022:AFM

[NBS^{+}22] Melam Nagaraju, B. Sobhan Babu, Meduri V. N. S. S. R. K. Sai Somayajulu, K. Subrahmanya Kousik Sarma, and Advaitha Vetagiri. An accurate foreground moving object detection based on segmentation techniques and optimal classifier. Concurrency and Computation: Practice and Experience, 34 (5):e6689:1–e6689:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Narengbam:2023:HHO

[ND23] Lenin Narengbam and Shouvik Dey. Harris hawk optimization trained artificial neural network for anomaly based intrusion detection system. Concurrency and Computation: Practice and Experience, 35(23):e7771:1–e7771:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nelson:2022:DFA

[NdMP22] Jacob Nelson, dePaul Miller, and Roberto Palmieri. Don’t forget about synchronization! guidelines for using locks on graphics processing units. Concurrency and Computation: Practice and Experience, 34(2):e5757:1–e5757:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nesi:2020:TBP

[NdSSSN20] Lucas Leandro Nesi, Matheus da Silva Serpa, Lucas Mello Schnorr, and Philippe Olivier Alexandre Navaux. Task-based parallel strategies for computational fluid dynamic application in heterogeneous CPU/GPU resources. Concurrency and Computation: Practice and Experience, 32(20):e5772:1– e5772:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nakamura:2020:IFC

[NET20] Shigenari Nakamura, Tomoya Enokido, and Makoto Takizawa. Information flow control in object-based peer-to-peer publish/subscribe systems. Concurrency and Computation: Practice and Experience, 32(8):e5118:1–e5118:??, April 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nakamura:2021:IEI

[NET21a] Shigenari Nakamura, Tomoya Enokido, and Makoto Takizawa. Implementation and evaluation of the information flow control for the Internet of Things. Concurrency and Computation: Practice and Experience, 33(19):e6311:1–e6311:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nakamura:2021:TBL

[NET21b] Shigenari Nakamura, Tomoya Enokido, and Makoto Takizawa. Time-based legality of information flow in the capability-based access control model for the Internet of Things. Concurrency and Computation: Practice and Experience, 33(23):e5944:1–e5944:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nasirian:2021:DEL

[NFF21] Sara Nasirian, Farhad Faghani, and Mahmoud Daneshvar Farzanegan. Doughnutie: an efficient and low-latency cloud data center network architecture. Concurrency and Computation: Practice and Experience, 33(20):e6337:1–e6337:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nowicki:2023:PEJ

[NGB23] Marek Nowicki, Lukasz Górski, and Piotr Bała. Performance evaluation of Java/PCJ implementation of parallel algorithms on the cloud (extended version). Concurrency and Computation: Practice and Experience, 35(15):e6536:1–e6536:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ning:2022:CGA

[NGD^{+}22] Xin Ning, Duoduo Gou, Xiaoli Dong, Weijuan Tian, Lina Yu, and Chuansheng Wang. Conditional generative adversarial networks based on the principle of homologycontinuity for face aging. Concurrency and Computation: Practice and Experience, 34(12):e5792:1–e5792:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

NTakpe:2022:DAS

[NGOS22] Tchimou N’Takpé, Jean Edgard Gnimassoun, Souleymane Oumtanaga, and Frédéric Suter. Data-aware and simulationdriven planning of scientific workflows on IaaS clouds. Concurrency and Computation: Practice and Experience, 34(14): e6719:1–e6719:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nguyen:2021:MRP

[Ngu21] Tung T. Nguyen. Multihop routing protocols on wireless ad hoc sensor networks with modified MAC layer and broadcasting scheme. Concurrency and Computation: Practice and Experience, 33(2):e5720:1–e5720:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ning:2021:TDC

[NGXZ21] Zhenhu Ning, Bei Gong, Meng Xu, and Caiqiu Zhou. Trusted data collection for Internet of Things. Concurrency and Computation: Practice and Experience, 33(10):e6166:1–e6166:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ni:2021:MCD

[NJ21] Tianquan Ni and Yi Jiang. Multiplatform collaborative detection resource scheduling method using K-means clustering algorithm and Hungarian algorithm. Concurrency and Computation: Practice and Experience, 33(22):e6075:1–e6075:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nirmal:2022:EOS

[NJK22] K. Nirmal, B. Janet, and R. Kumar. Enhancing online security using selective DOM approach to counter phishing attacks. Concurrency and Computation: Practice and Experience, 34(7):e5525:1–e5525:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Najjari:2022:SOL

[NK22] Ahmadreza Hassannezhad Najjari and Ali Asghar Pourhaji Kazem. A systematic overview of live virtual machine migration methods. Concurrency and Computation: Practice and Experience, 34(17):e6915:1–e6915:??, August 1, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nakamura:2021:CPS

[NKKM21] Junya Nakamura, Yonghwan Kim, Yoshiaki Katayama, and Toshimitsu Masuzawa. A cooperative partial snapshot algorithm for checkpoint-rollback recovery of large-scale and dynamic distributed systems and experimental evaluations. Concurrency and Computation: Practice and Experience, 33 (12):e5647:1–e5647:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Naresh:2024:JOB

[NKST24] M. Naresh, G. V. Pradeep Kumar, V. Sireesha, and V. V. Satyanarayana Tallapragada. Joint optimal beam forming and resource allocation in intelligent reflecting surface aided wireless power transfer rate splitting multiple access system. Concurrency and Computation: Practice and Experience, 36(14):e8098:1–e8098:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nakamura:2023:EVT

[NKY23] Junya Nakamura, Sayaka Kamei, and Yukiko Yamauchi. Evacuation from various types of finite two-dimensional square grid fields by a metamorphic robotic system. Concurrency and Computation: Practice and Experience, 35(14): e6628:1–e6628:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nanehkaran:2022:NNA

[NLB22] Farimah Houshmand Nanehkaran, Seyed Mohammadreza Lajevardi, and Mahmoud Mahlouji Bidgholi. Nearest neighbors algorithm and genetic-based collaborative filtering. Concurrency and Computation: Practice and Experience, 34(1): e6538:1–e6538:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Noh:2020:CAH

[NLG^{+}20] Hyunmin Noh, Hwanwook Lee, Yunmin Go, Hyun Park, Jongman Lee, Jaewoo Kim, and Hwangjun Song. Congestionaware HTTP adaptive streaming system over SDN-enabled Wi-Fi network. Concurrency and Computation: Practice and Experience, 32(21):e5310:1–e5310:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Narayanan:2020:MSA

[NM20] S. Sankara Narayanan and G. Murugaboopathi. Modified secure AODV protocol to prevent wormhole attack in MANET. Concurrency and Computation: Practice and Experience, 32 (4):e5017:1–e5017:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nair:2023:NMT

[NM23] Lakshmi M. Nair and Mary Synthia Regis Prabha D. M. A novel modified tunicate swarm algorithm for optimal charge scheduling and pricing of electric vehicles with user prioritybased charging station selection. Concurrency and Computation: Practice and Experience, 35(3):e7494:1–e7494:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nasreen:2024:CFC

[NM24] Shaiqa Nasreen and Ajaz Hussain Mir. Cloud forensics: a centralized cloud provenance investigation system using MECC. Concurrency and Computation: Practice and Experience, 36 (6):e7949:1–e7949:??, March 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nedelec:2021:SSE

[NMM21] Brice Nédelec, Pascal Molli, and Achour Mostéfaoui. A scalable sequence encoding for collaborative editing. Concurrency and Computation: Practice and Experience, 33(8):e4108:1– e4108:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nazir:2022:DIE

[NMQ22] Azra Nazir, Roohie Naaz Mir, and Shaima Qureshi. Deflateinflate: Exploiting hashing trick for bringing inference to the edge with scalable convolutional neural networks. Concurrency and Computation: Practice and Experience, 34(3): e6593:1–e6593:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Novais:2021:OCL

[NMS^{+}21] João P. P. Novais, Lucas A. Maciel, Matheus A. Souza, Mark A. J. Song, and Henrique C. Freitas. An open computing language-based parallel brute force algorithm for formal concept analysis on heterogeneous architectures. Concurrency and Computation: Practice and Experience, 33(18):e6220:1– e6220:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nguyen:2022:FBH

[NMS^{+}22] Tan Nguyen, Colin MacLean, Marco Siracusa, Douglas Doerfler, Nicholas J. Wright, and Samuel Williams. FPGA-based HPC accelerators: an evaluation on performance and energy efficiency. Concurrency and Computation: Practice and Experience, 34(20):e6570:1–e6570:??, September 10, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nunes:2023:OCC

[NMT^{+}23] Alan L. Nunes, Alba Melo, Claude Tadonki, Cristina Boeres, Daniel de Oliveira, and Lúcia Maria de Assumpção. Optimizing computational costs of Spark for SARS-CoV-2 sequences comparisons on a commercial cloud. Concurrency and Computation: Practice and Experience, 35(18):e7678:1–e7678:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nataraj:2023:NHM

[NN23] N. Nataraj and R. V. Nataraj. A novel hybrid meta-heuristic-oriented latency sensitive cloud object storage system. Concurrency and Computation: Practice and Experience, 35(21): e7672:1–e7672:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nguyen:2023:ECF

[NNJC23] Luong Vuong Nguyen, Tri-Hai Nguyen, Jason J. Jung, and David Camacho. Extending collaborative filtering recommendation using word embedding: a hybrid approach. Concurrency and Computation: Practice and Experience, 35(16): e6232:1–e6232:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nagadurga:2022:GWO

[NNVD22] T. Nagadurga, P. V. R. L. Narasimham, V. S. Vakula, and Ramesh Devarapalli. Gray wolf optimization-based optimal grid connected solar photovoltaic system with enhanced power quality features. Concurrency and Computation: Practice and Experience, 34(5):e6696:1–e6696:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ning:2023:MVF

[NNX^{+}23] Xin Ning, Fangzhe Nan, Shaohui Xu, Lina Yu, and Liping Zhang. Multi-view frontal face image generation: a survey. Concurrency and Computation: Practice and Experience, 35 (18):e6147:1–e6147:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Niasari:2021:ESP

[NQ21] Mehrdad Arabpour Niasari and Ke Qiu. An efficient shortest path routing on the hypercube with blocking/faulty nodes. Concurrency and Computation: Practice and Experience, 33 (12):e6124:1–e6124:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Naresh:2022:SLM

[NR22] Vankamamidi S. Naresh and Sivaranjani Reddi. Secure lightweight multi-party key agreement based on hyperelliptic curve Diffie–Hellman for constraint networks. Concurrency and Computation: Practice and Experience, 34(13):e6921:1– e6921:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nosrati:2023:HFH

[NR23] Vahid Nosrati and Mohsen Rahmani. HMDE-FS: a homogeneous distributed ensemble feature selection framework based on resampling with/without replacement. Concurrency and Computation: Practice and Experience, 35(7):e7613:1– e7613:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Naranjo:2023:SGE

[NRMB23] Diana M. Naranjo, Sebastián Risco, Germán Moltó, and Ignacio Blanquer. A serverless gateway for event-driven machine learning inference in multiple clouds. Concurrency and Computation: Practice and Experience, 35(18):e6728:1–e6728:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ng:2020:VRD

[NRP^{+}20] Chee Keong Ng, Sutharshan Rajasegarar, Lei Pan, Frank Jiang, and Leo Yu Zhang. VoterChoice: a ransomware detection honeypot with multiple voting framework. Concurrency and Computation: Practice and Experience, 32(14):e5726:1– e5726:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Narahari:2023:CAP

[NS23] Narasimhaiah Narahari and Rachapudi Praveen Sam. Canny aspiration paraphernalia framework based healthcare monitoring system and secure medical interoperability. Concurrency and Computation: Practice and Experience, 35(22): e7722:1–e7722:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Natarajan:2021:ASA

[NSBT21] Deepa Natarajan, Esakkirajan Sankaralingam, Keerthiveena Balraj, and Veerakumar Thangaraj. Automated segmentation algorithm with deep learning framework for early detection of glaucoma. Concurrency and Computation: Practice and Experience, 33(10):e6181:1–e6181:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nayak:2024:ETS

[NSK^{+}24] Pritam Kumar Nayak, Ravi Shankar Singh, Shweta Kushwaha, Prasanth Kumar Bevara, Vinod Kumar, and Rambabu Medara. Efficient task scheduling on the cloud using artificial neural network and particle swarm optimization. Concurrency and Computation: Practice and Experience, 36(6): e7954:1–e7954:??, March 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nawrocki:2021:ACA

[NSKS21] Piotr Nawrocki, Bartlomiej Sniezynski, Joanna Kolodziej, and Pawel Szynkiewicz. Adaptive context-aware service optimization in mobile cloud computing accounting for security aspects. Concurrency and Computation: Practice and Experience, 33(18):e6070:1–e6070:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nascimento:2021:IRL

[NSPdO21] André Nascimento, Vítor Silva, Aline Paes, and Daniel de Oliveira. An incremental reinforcement learning scheduling strategy for data-intensive scientific workflows in the cloud. Concurrency and Computation: Practice and Experience, 33 (11):e6193:1–e6193:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nadendla:2022:FDI

[NSR22] Hanumantha Rao Nadendla, Atluri Srikrishna, and Kancherla Gangadhara Rao. FSOA-DNFNet: Incremental indexing and image classification using hybrid optimization-based deep neuro fuzzy network. Concurrency and Computation: Practice and Experience, 34(19):e7049:1–e7049:??, August 30, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Niaz:2022:IEC

[NSSS22] Iqra Niaz, Aamir Sanaullah, Iram Saleem, and Javid Shabbir. An improved efficient class of estimators for the population variance. Concurrency and Computation: Practice and Experience, 34(4):e6620:1–e6620:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nair:2023:LAP

[NT23] Aiswarya S. Nair and Sabu M. Thampi. A location-aware physical unclonable function and Chebyshev map-based mutual authentication mechanism for internet of surveillance drones. Concurrency and Computation: Practice and Experience, 35(19):e7564:1–e7564:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nguyen:2023:EWT

[NTB23] Hai T. Nguyen, Toan B. Tran, and Phuong H. D. Bui. An effective way for Taiwanese stock price prediction: Boosting the performance with machine learning techniques. Concurrency and Computation: Practice and Experience, 35(15):e6437:1– e6437:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nallusamy:2022:AAE

[NTK22] Priyanka Nallusamy, Reshmi Tr, and Murugan Krishnan. AGFT: Adaptive entries aggregation scheme to prevent overflow in multiple flow table environment. Concurrency and Computation: Practice and Experience, 34(1):e6491:1– e6491:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nagasundharamoorthi:2024:HMA

[NVV24] Insozhan Nagasundharamoorthi, Prabhu Venkatesan, and Parthasarathy Velusamy. Hash message authentication codes for securing data in wireless body area networks. Concurrency and Computation: Practice and Experience, 36(5):e7934:1– e7934:??, February 29, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nguyen:2021:EMA

[NWT21] Truong Thao Nguyen, Mohamed Wahib, and Ryousei Takano. Efficient MPI-AllReduce for large-scale deep learning on GPU-clusters. Concurrency and Computation: Practice and Experience, 33(12):e5574:1–e5574:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Niu:2022:PSW

[NWW^{+}22] Xinzheng Niu, Peng Wu, Chase Q. Wu, Aiqin Hou, and Mideng Qian. On a parallel spark workflow for frequent itemset mining based on array prefix-tree. Concurrency and Computation: Practice and Experience, 34(14):e6313:1–e6313:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ning:2023:RRC

[NWX^{+}23] Xin Ning, Xinran Wang, Shaohui Xu, Weiwei Cai, Liping Zhang, Lina Yu, and Wenfa Li. A review of research on cotraining. Concurrency and Computation: Practice and Experience, 35(18):e6276:1–e6276:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ni:2021:DTF

[NWZ^{+}21] Tian Ni, Lei Wang, Pengchao Zhang, Bin Wang, and Wei Li. Daily tourist flow forecasting using SPCA and CNN-LSTM neural network. Concurrency and Computation: Practice and Experience, 33(5):e5980:1–e5980:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Niu:2022:MDS

[NY22] Xin Niu and Xiaomo Yu. Mobile device security defense method and system based on address jump using sliding window technology. Concurrency and Computation: Practice and Experience, 34(14):e5625:1–e5625:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nie:2024:ETE

[NZD^{+}24] Chenfei Nie, Zhipeng Zhou, Mianxiong Dong, Kaoru Ota, and Qiang Li. EPIDL: Towards efficient and privacy-preserving inference in deep learning. Concurrency and Computation: Practice and Experience, 36(14):e8110:1–e8110:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Olmezogullari:2022:PRC

[OA22] Erdi Olmezogullari and Mehmet S. Aktas. Pattern2Vec: Representation of clickstream data sequences for learning user navigational behavior. Concurrency and Computation: Practice and Experience, 34(9):e6546:1–e6546:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Owolabi:2022:NRT

[OAA22] Abiola T. Owolabi, Kayode Ayinde, and Olusegun O. Alabi. A new ridge-type estimator for the linear regression model with correlated regressors. Concurrency and Computation: Practice and Experience, 34(15):e6933:1–e6933:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ouared:2023:EAD

[OAS23] Abdelkader Ouared, Moussa Amrani, and Pierre-Yves Schobbens. Explainable AI for DBA: Bridging the DBA’s experience and machine learning in tuning database systems. Concurrency and Computation: Practice and Experience, 35(21):e7698:1–e7698:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ouedrhiri:2022:IRS

[OBER22] Oumayma Ouedrhiri, Oumayma Banouar, Salah El Hadaj, and Said Raghay. Intelligent recommender system based on quantum clustering and matrix completion. Concurrency and Computation: Practice and Experience, 34(15):e6943:1– e6943:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Omezzine:2020:TGM

[OBTC20] Aya Omezzine, Narjes Bellamine Ben Saoud, Said Tazi, and Gene Cooperman. Towards a generic multilayer negotiation framework for efficient application provisioning in the cloud. Concurrency and Computation: Practice and Experience, 32 (1):e4182:1–e4182:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ouared:2022:DDN

[OCD22] Abdelkader Ouared, Abdelhafid Chadli, and Mohamed Amine Daoud. DeepCM: Deep neural networks to improve accuracy prediction of database cost models. Concurrency and Computation: Practice and Experience, 34(10):e6724:1–e6724:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oliveira:2023:HRA

[OCR^{+}23] Westerley C. Oliveira, Michael Canesche, Lucas Reis, José Augusto M. Nacif, and Ricardo S. Ferreira. Heterogeneous reconfigurable architectures for machine learning dataflows. Concurrency and Computation: Practice and Experience, 35 (17):e6939:1–e6939:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Omana:2023:HGB

[ODK^{+}23] Sinciya Ponnupilla Omana, Jawad Ahmad Dar, Thevasigamani Rajesh Kumar, Arpakkam Karuppan Sampath, and Sudhir Sharma. Henry gas bird swarm optimization algorithm-based deep learning for brain tumor classification using magnetic resonance imaging. Concurrency and Computation: Practice and Experience, 35(4):e7541:1–e7541:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oyelade:2022:CAB

[OE22a] Olaide N. Oyelade and Absalom E. Ezugwu. Characterization of abnormalities in breast cancer images using natureinspired metaheuristic optimized convolutional neural networks model. Concurrency and Computation: Practice and Experience, 34(4):e6629:1–e6629:??, February 15, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oyelade:2022:CPS

[OE22b] Olaide N. Oyelade and Absalom E. Ezugwu. A comparative performance study of random-grid model for hyperparameters selection in detection of abnormalities in digital breast images. Concurrency and Computation: Practice and Experience, 34 (13):e6914:1–e6914:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Orhean:2022:ESD

[OGA^{+}22] Alexandru Iulian Orhean, Anna Giannakou, Katie Antypas, Ioan Raicu, and Lavanya Ramakrishnan. Evaluation of a scientific data search infrastructure. Concurrency and Computation: Practice and Experience, 34(27):e7261:1–e7261:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2020:CIC

[Ogi20a]  Lidia Ogiela.  Cognitive and innovative computation

paradigms for big data and cloud computing applications. Concurrency and Computation: Practice and Experience, 32 (8):e5495:1–e5495:??, April 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2020:AAI

[Ogi20b] Marek R. Ogiela. Advanced approaches for information processing in multimedia, decision-making, and security systems. Concurrency and Computation: Practice and Experience, 32(18):e5665:1–e5665:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2020:CCD

[Ogi20c] Urszula Ogiela. Cognitive cryptography for data security in cloud computing. Concurrency and Computation: Practice and Experience, 32(18):e5557:1–e5557:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2021:TCA

[Ogi21] Urszula Ogiela. Transformative and cognitive approaches to information retrieval and security procedures. Concurrency and Computation: Practice and Experience, 33(23):e5890:1– e5890:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ozeloglu:2022:DRL

[ÖGS22] Alican Özeloglu, Ismihan Gül Gürbüz, and Ismail San. Deep reinforcement learning-based autonomous parking design with neural network compute accelerators. Concurrency and Computation: Practice and Experience, 34(9):e6670:1– e6670:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oh:2021:PPM

[Oh21] Sangyoon Oh. Parallel programming models in highperformance cloud (ParaMo 2019). Concurrency and Computation: Practice and Experience, 33(18):e6527:1–e6527:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Olha:2020:EHD

[OHFF20] Jaroslav Olha, Jana Hozzová, Jan Fousek, and Jirí Filipovic. Exploiting historical data: Pruning autotuning spaces and estimating the number of tuning steps. Concurrency and Computation: Practice and Experience, 32(21):e5962:1–e5962:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oswald:2021:END

[OHRS21] C. Oswald, E. Haritha, A. Akash Raja, and B. Sivaselvan. An efficient and novel data clustering and run length encoding approach to image compression. Concurrency and Computation: Practice and Experience, 33(10):e6185:1–e6185:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ozkale:2022:MSC

[ÖK22] M. Revan Özkale and Özge Kuran. Model selection via conditional conceptual predictive statistic for mixed and stochastic restricted ridge estimators in linear mixed models. Concurrency and Computation: Practice and Experience, 34(28): e7366:1–e7366:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Olimov:2021:WIB

[OKJ^{+}21] Bekhzod Olimov, Sanjar Karshiev, Eungyeong Jang, Sadia Din, Anand Paul, and Jeonghong Kim. Weight initialization based-rectified linear unit activation function to improve the performance of a convolutional neural network model. Concurrency and Computation: Practice and Experience, 33(22): e6143:1–e6143:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2021:TCS

[OLF21] Lidia Ogiela, Fang-Yie Leu, and Ugo Fiore. Transformative computing in security, big data analysis, and cloud computing applications. Concurrency and Computation: Practice and Experience, 33(23):e6654:1–e6654:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oruche:2023:SGA

[OMA^{+}23] Roland Oruche, Eric Milman, Mauro Lemus Alarcon, Xiyao Cheng, Ashish Pandey, Songjie Wang, Prasad Calyam, and Kerk Kee. Science gateway adoption using plug-in middleware for evidence-based healthcare data management. Concurrency and Computation: Practice and Experience, 35(18): e7195:1–e7195:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Onan:2021:SAP

[Ona21] Aytug Onan. Sentiment analysis on product reviews based on weighted word embeddings and deep neural networks. Concurrency and Computation: Practice and Experience, 33(23): e5909:1–e5909:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2020:CSP

[OO20] Lidia Ogiela and Marek R. Ogiela. Cognitive security paradigm for cloud computing applications. Concurrency and Computation: Practice and Experience, 32(8):e5316:1– e5316:??, April 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2021:NCS

[OO21] Lidia Ogiela and Marek R. Ogiela. New cognitive sharing algorithms for cloud service management. Concurrency and Computation: Practice and Experience, 33(23):e6354:1– e6354:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Olakanmi:2022:SRM

[OO22] Oladayo Olufemi Olakanmi and Kehinde Oluwasesan Odeyemi. Secure reputation and morphism-based offloading scheme: a veritable tool for multi-party computation in Industrial Internet of Things. Concurrency and Computation: Practice and Experience, 34(21):e7116:1–e7116:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ouared:2023:CDC

[OO23a] Abdelkader Ouared and Yassine Ouhammou. Capitalizing the database cost models process through a service-based pipeline. Concurrency and Computation: Practice and Experience, 35(11):e6463:1–e6463:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ozcetin:2023:VNS

[OO23b] Erdener Ozcetin and Gurkan Ozturk. A variable neighborhood search for Open Vehicle Routing Problem. Concurrency and Computation: Practice and Experience, 35(7):e7598:1– e7598:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2021:HOS

[ORP21] Marek R. Ogiela, Wenny Rahayu, and Francesco Palmieri. Human oriented solutions for intelligent analysis, multimedia and communication systems. Concurrency and Computation: Practice and Experience, 33(19):e6532:1–e6532:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2021:IST

[OS21a] Lidia Ogiela and Václav Snásel. Intelligent and semantic threshold schemes for security in cloud computing. Concurrency and Computation: Practice and Experience, 33(2): e5247:1–e5247:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2021:THO

[OS21b] Lidia Ogiela and Václav Snásel. Towards human-oriented solutions for deep semantic data analysis. Concurrency and Computation: Practice and Experience, 33(19):e6252:1– e6252:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Odabas:2023:DQG

[OSK23] Mehmet Serhat Odabas, Nurettin Senyer, and Dursun Kurt. Determination of quality grade of tobacco leaf by image processing on correlated color temperature. Concurrency and Computation: Practice and Experience, 35(2):e7506:1– e7506:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Osliak:2023:CTI

[OSMM23] Oleksii Osliak, Andrea Saracino, Fabio Martinelli, and Paolo Mori. Cyber threat intelligence for critical infrastructure security. Concurrency and Computation: Practice and Experience, 35(23):e7759:1–e7759:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ozbay:2022:TPE

[ÖSTY22] Nimet Özbay, Gülesen Üstündag Siray, Selma Toker, and Ismail Yenilmez. Two-parameter estimation for Tobit model: an application to national health and nutrition examination survey dataset. Concurrency and Computation: Practice and Experience, 34(18):e7014:1–e7014:??, August 15, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ogiela:2020:NPC

[OT20] Lidia Ogiela and Makoto Takizawa. New protocols of cognitive data management and sharing in cloud computing. Concurrency and Computation: Practice and Experience, 32(18): e5546:1–e5546:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ozturk:2023:TLF

[ÖTT23] Celal Öztürk, Murat Tasyürek, and Mehmet Ugur Türkdamar. Transfer learning and fine-tuned transfer learning methods’ effectiveness analyse in the CNN-based deep learning models. Concurrency and Computation: Practice and Experience, 35 (4):e7542:1–e7542:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oranye:2022:MJK

[OU22] Henrietta Ebele Oranye and Fidelis Ifeanyi Ugwuowo. Modified jackknife Kibria–Lukman estimator for the Poisson regression model. Concurrency and Computation: Practice and Experience, 34(6):e6757:1–e6757:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Oksuz:2022:CDS

[ÖUG22] Cosku Öksüz, Oguzhan Urhan, and Mehmet Kemal Güllü. COVID-19 detection with severity level analysis using the deep features, and wrapper-based selection of ranked features. Concurrency and Computation: Practice and Experience, 34(20):e6802:1–e6802:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Owenson:2020:UCM

[OWB^{+}20] A. M. B. Owenson, S. A. Wright, R. A. Bunt, Y. K. Ho, M. J. Street, and S. A. Jarvis. An unstructured CFD miniapplication for the performance prediction of a production CFD code. Concurrency and Computation: Practice and Experience, 32(10):e5443:1–e5443:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ouyang:2023:CDS

[OXBL23] Mingxue Ouyang, Jianqing Xi, Weihua Bai, and Keqin Li. A container deployment strategy for server clusters with different resource types. Concurrency and Computation: Practice and Experience, 35(10):e7665:1–e7665:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pulagara:2021:IRC

[PA21] Seshu Babu Pulagara and P. J. A. Alphonse. An intelligent and robust conditional privacy preserving authentication and group-key management scheme for vehicular ad hoc networks using elliptic curve cryptosystem. Concurrency and Computation: Practice and Experience, 33(3):e5153:1–e5153:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perumal:2023:ODN

[PA23] Karthika Perumal and Karmel Arockiasamy. Optimized deep neural network based DDoS attack detection and bait mitigation process in software defined network. Concurrency and Computation: Practice and Experience, 35(12):e7692:1– e7692:??, May 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Palancioglu:2022:HMB

[Pal22] Haci Mustafa Palancioglu. Histogram modification based pansharpening by using differential evolution algorithm. Concurrency and Computation: Practice and Experience, 34(27): e7335:1–e7335:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pan:2020:RME

[Pan20] Xuchao Pan. Retracted: The mechanism and experimental study on the interference of high voltage lines to navigation system. Concurrency and Computation: Practice and Experience, 32(2):e5513:1–e5513:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See [FPHZ19].

Palanisamy:2022:PSB

[PAN22] Vijayanand Sellamuthu Palanisamy, Rajiv Kannan Athiappan, and Thirugnanasambandan Nagalingam. Pap smear based cervical cancer detection using residual neural networks deep learning architecture. Concurrency and Computation: Practice and Experience, 34(4):e6608:1–e6608:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Panda:2023:DAS

[Pan23] Arnapurna Panda. Determining approximate solution of partial differential equations using radial basis function model trained with parallel symbiotic organisms search algorithm. Concurrency and Computation: Practice and Experience, 35 (5):e7558:1–e7558:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pan:2024:LFM

[Pan24] Pan Pan. Local feature matching and tracking optimization of machine VSLAM based on AG Sinkhorn. Concurrency and Computation: Practice and Experience, 36(11):e8022:1– e8022:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Parlak:2022:CIC

[Par22] Bekir Parlak. Class-index corpus-index measure: a novel feature selection method for imbalanced text data. Concurrency and Computation: Practice and Experience, 34(21):e7140:1– e7140:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Priya:2021:NGF

[PB21] R. Vishnu Priya and Rawal Bharat. A novel geometric fuzzy membership functions for mouth and eye brows to recognize emotions. Concurrency and Computation: Practice and Experience, 33(14):e5610:1–e5610:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peccerillo:2022:FTD

[PB22a] Biagio Peccerillo and Sandro Bartolini. Flexible task-DAG management in PHAST library: Data-parallel tasks and orchestration support for heterogeneous systems. Concurrency and Computation: Practice and Experience, 34(2):e5842:1– e5842:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pradeep:2022:EEC

[PB22b] S. Pradeep and B. R. Tapas Bapu. An efficient energy consumption and delay aware autonomous data gathering routing protocol scheme using a deep learning mobile edge model and beetle antennae search algorithm for underwater wireless sensor network. Concurrency and Computation: Practice and Experience, 34(15):e6946:1–e6946:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Penha:2023:GRA

[PBC^{+}23] Jeronimo Costa Penha, Lucas Bragança, Michael Canesche, Dener Ribeiro, José Augusto M. Nacif, and Ricardo S. Ferreira. Gene regulatory accelerators on cloud FPGA. Concurrency and Computation: Practice and Experience, 35(24): e7822:1–e7822:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Paciorek:2021:VSP

[PBD^{+}21] Mateusz Paciorek, Jakub Bujas, Dawid Dworak, Wojciech Turek, and Aleksander Byrski. Validation of signal propagation modeling for highly scalable simulations. Concurrency and Computation: Practice and Experience, 33(14):e5718:1– e5718:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Parpinelli:2023:MPS

[PBD23] Rafael Stubs Parpinelli, Mateus Boiani, and André E. P. Dias. A massively parallel speciation-based differential evolution algorithm applied to the 3D-AB protein structure prediction. Concurrency and Computation: Practice and Experience, 35 (17):e6745:1–e6745:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

P:2023:KHO

[PBK23] Pratheep Kumar P, Mary Amala Bai V, and Ram P. Krish. Krill herd optimization algorithm with deep convolutional neural network fostered breast cancer classification using mammogram images. Concurrency and Computation: Practice and Experience, 35(7):e7605:1–e7605:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pizzolante:2021:MLB

[PCC^{+}21] Raffaele Pizzolante, Arcangelo Castiglione, Bruno Carpentieri, Roberto Contaldo, Gianni D’Angelo, and Francesco Palmieri. A machine learning-based memory forensics methodology for TOR browser artifacts. Concurrency and Computation: Practice and Experience, 33(23):e5935:1– e5935:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perez:2021:MCT

[PCG^{+}21] Manuel Gil Pérez, Alberto Huertas Celdrán, Pietro G. Giardina, Giacomo Bernini, Simone Pizzimenti, Félix J. García Clemente, Gregorio Martínez Perez, Giovanni Festa, and Fabio Paglianti. Mitigation of cyber threats: Protection mechanisms in federated SDN/NFV infrastructures for 5G within FIRE+. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perumalla:2023:SCU

[PCK23] Subhadra Perumalla, Santanu Chatterjee, and A. P. Siva Kumar. Secure communication using multilevel authentication strategy in Internet of Drones. Concurrency and Computation: Practice and Experience, 35(12):e7667:1–e7667:??, May 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Prakash:2021:HHH

[PCR21] B. Guru Prakash, C. Balasubramanian Chelliah, and R. Sukumar Ramanujam. HHFDS: Heterogeneous hybridized fuzzybased dijkstra’s multitask scheduling in WSN. Concurrency and Computation: Practice and Experience, 33(3):e5354:1– e5354:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

P:2020:RTA

[PD20] Chellammal P and Sheba Kezia Malarchelvi P D. Real-time anomaly detection using parallelized intrusion detection architecture for streaming data. Concurrency and Computation: Practice and Experience, 32(4):e5013:1–e5013:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peterson:2022:CEM

[PD22a] Christina Peterson and Damian Dechev. The CAS-extended model. Concurrency and Computation: Practice and Experience, 34(2):e5656:1–e5656:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ponmalar:2022:HWT

[PD22b] A. Ponmalar and V. Dhanakoti. Hybrid Whale Tabu algorithm optimized convolutional neural network architecture for intrusion detection in big data. Concurrency and Computation: Practice and Experience, 34(19):e7038:1–e7038:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pullagura:2023:BHA

[PD23] Joshua Reginald Pullagura and Venkata Rao Dhulipalla. Black-hole attack and counter measure in ad hoc networks using traditional routing optimization. Concurrency and Computation: Practice and Experience, 35(9):e7643:1–e7643:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Priyadharshini:2024:EKA

[PDRM24] A. Priyadharshini, Suresh Dannana, Arun Sekar Rajasekaran, and Azees Maria. An efficient key agreement and anonymous privacy preserving scheme for vehicular ad-hoc networks with handover authentication. Concurrency and Computation: Practice and Experience, 36(7):e7979:1–e7979:??, March 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Plauth:2023:IAN

[PEGP23] Max Plauth, Felix Eberhardt, Andreas Grapentin, and Andreas Polze. Improving the accessibility of NUMA-aware C++ application development based on the PGASUS framework. Concurrency and Computation: Practice and Experience, 35(14):e6887:1–e6887:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Petrovic:2022:RDL

[PGD^{+}22] Veljko B. Petrović, Gorana Gojić, Dinu Dragan, Dusan B. Gajić, Nebojsa Horvat, Radovan Turović, and Ana Oros. Robustness of deep learning methods for ocular fundus segmentation: Evaluation of blur sensitivity. Concurrency and Computation: Practice and Experience, 34(14):e6809:1–e6809:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2023:BBM

[PGL^{+}23] Tao Peng, Kejian Guan, Jierong Liu, Jianer Chen, Guojun Wang, and Jiawei Zhu. A blockchain-based mobile crowdsensing scheme with enhanced privacy. Concurrency and Computation: Practice and Experience, 35(19):e6664:1–e6664:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pourghebleh:2022:RTE

[PHDS22] Behrouz Pourghebleh, Negin Hekmati, Zahra Davoudnia, and Mehrdad Sadeghi. A roadmap towards energy-efficient data fusion methods in the Internet of Things. Concurrency and Computation: Practice and Experience, 34(15):e6959:1– e6959:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2022:ESRb

[PHZ^{+}22a] Zhiguo Peng, Meifa Huang, Yanru Zhong, Leilei Chen, and Guanghao Liu. Enhanced semantic representation of coaxiality with double material requirements. Concurrency and Computation: Practice and Experience, 34(9):e5744:1– e5744:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2022:ESRa

[PHZ^{+}22b] Zhiguo Peng, Meifa Huang, Yanru Zhong, Leilei Chen, and Guanghao Liu. Explicitly semantic representation of pattern and combined geometrical specification. Concurrency and Computation: Practice and Experience, 34(9):e5743:1– e5743:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Punitha:2021:CCI

[PI21] A. Amali Angel Punitha and G. Indumathi. Centralized cloud information accountability integrity with Firefly Key Generation Algorithm (CCIAI-FKGA) for cloud environment. Concurrency and Computation: Practice and Experience, 33(3): e5223:1–e5223:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Piccialli:2021:SIT

[PJ21] Francesco Piccialli and Gwanggil Jeon. Special issue on toward the Internet of Things of year 2020: Applications and future trends. Concurrency and Computation: Practice and Experience, 33(3):e5733:1–e5733:??, February 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pattewar:2023:BBB

[PJK23] Tareek Pattewar, Dinesh Jain, and B. V. Kiranmayee. A Buffalo-based bi-directional recurrent paradigm for Indian stock market prediction. Concurrency and Computation: Practice and Experience, 35(9):e7642:1–e7642:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pantekis:2024:OMP

[PJKO24] Filippos Pantekis, Phillip James, Oliver Kullmann, and Liam O’Reilly. Optimized massively parallel solving of N-Queens on GPGPUs. Concurrency and Computation: Practice and Experience, 36(10):e8004:1–e8004:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Potu:2021:ORS

[PJP21] Narayana Potu, Chandrashekar Jatoth, and Premchand Parvataneni. Optimizing resource scheduling based on extended particle swarm optimization in fog computing environments. Concurrency and Computation: Practice and Experience, 33(23):e6163:1–e6163:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pandian:2022:IDC

[PK22a] J. Arun Pandian and K. Kanchanadevi. An improved deep convolutional neural network for detecting plant leaf diseases. Concurrency and Computation: Practice and Experience, 34(28):e7357:1–e7357:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Polat:2022:GAM

[PK22b] Önder Polat and Sema Koç Kayhan. GPU-accelerated and mixed norm regularized online extreme learning machine. Concurrency and Computation: Practice and Experience, 34 (15):e6967:1–e6967:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pon:2022:BBC

[PK22c] Partheeban Pon and Kavitha V. Blockchain based cloud service security architecture with distributed machine learning for smart device traffic record transaction. Concurrency and Computation: Practice and Experience, 34(3):e683:1–e683:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Punj:2022:DDA

[PK22d] Roopali Punj and Rakesh Kumar. Data dissemination approach using machine learning techniques for WBANs. Concurrency and Computation: Practice and Experience, 34(5): e6688:1–e6688:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Park:2023:DNA

[PK23a] Kicheol Park and Bongjae Kim. Dynamic neuromorphic architecture selection scheme for intelligent Internet of Things services. Concurrency and Computation: Practice and Experience, 35(16):e6357:1–e6357:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Patil:2023:MLB

[PK23b] Bhushan Patil and Satish Ket. Machine learning based secure and efficient task allocation in multi-cloud. Concurrency and Computation: Practice and Experience, 35(27):e7848:1– e7848:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pathak:2022:SSJ

[PKB22] Diganta Kumar Pathak, Sanjib Kumar Kalita, and Dhruba Kumar Bhattacharya. Spectral spatial joint feature based convolution neural network for hyperspectral image classification. Concurrency and Computation: Practice and Experience, 34 (3):e6547:1–e6547:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pierantoni:2023:TRA

[PKB^{+}23] Gabriele Pierantoni, Tamas Kiss, Alexander Bolotov, Dimitrios Kagialis, James DesLauriers, Amjad Ullah, Huankai Chen, David Chan You Fee, Hai-Van Dang, Jozsef Kovacs, Anna Belehaki, Themistocles Herekakis, Ioanna Tsagouri, and Sandra Gesing. Toward a reference architecture based science gateway framework with embedded e-learning support. Concurrency and Computation: Practice and Experience, 35 (18):e6872:1–e6872:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Papcun:2020:EEI

[PKC^{+}20] Peter Papcun, Erik Kajati, Dominika Cupkova, Jozef Mocnej, Martin Miskuf, and Iveta Zolotova. Edge-enabled IoT gateway criteria selection and evaluation. Concurrency and Computation: Practice and Experience, 32(13):e5219:1–e5219:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Patil:2021:DFD

[PKK21] Nilesh Vishwasrao Patil, C. Rama Krishna, and Krishan Kumar. Distributed frameworks for detecting distributed denial of service attacks: a comprehensive review, challenges and future directions. Concurrency and Computation: Practice and Experience, 33(10):e6197:1–e6197:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Poongavanam:2023:TAL

[PKK23] Elumalaivasan Poongavanam, Padmanathan Kasinathan, and Kulothungan Kanagasabai. Taylor ant lion optimizationbased generative adversarial networks for forecasting electricity consumption. Concurrency and Computation: Practice and Experience, 35(7):e7607:1–e7607:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pintye:2021:BDM

[PKKL21] István Pintye, Eszter Kail, Péter Kacsuk, and Róbert Lovas. Big data and machine learning framework for clouds and its usage for text classification. Concurrency and Computation: Practice and Experience, 33(19):e6164:1–e6164:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ponto:2021:CWC

[PKM21] René Ponto, Gábor Kecskeméti, and Zoltán Á. Mann. Comparison of workload consolidation algorithms for cloud data centers. Concurrency and Computation: Practice and Experience, 33(9):e6138:1–e6138:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Prakash:2022:SSR

[PKR22] Perumalla Suman Prakash, Dwaram Kavitha, and Pakanati Chenna Reddy. Safe and secured routing using multiobjective fractional artificial lion algorithm in WSN. Concurrency and Computation: Practice and Experience, 34(21): e7098:1–e7098:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Puri:2022:KMA

[PKS22] Vartika Puri, Parmeet Kaur, and Shelly Sachdeva. (k, m, t)- anonymity: Enhanced privacy for transactional data. Concurrency and Computation: Practice and Experience, 34(18): e7020:1–e7020:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Priya:2024:SHG

[PKSM24] G. Nirmala Priya, K. Suresh Kumar, N. Suganthi, and Satish Muppidi. Squirrel Henry Gas Solubility Optimization driven Deep Maxout Network with multi-texture feature descriptors for digital image forgery detection. Concurrency and Computation: Practice and Experience, 36(8):e7965:1–e7965:??, April 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Platos:2021:PDM

[PKVS21] Jan Platos, Pavel Kromer, Miroslav Voznak, and Vaclav Snasel. Population data mobility retrieval at territory of Czechia in pandemic COVID-19 period. Concurrency and Computation: Practice and Experience, 33(23):e6105:1– e6105:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Periasamy:2021:EHF

[PL21] J. K. Periasamy and B. Latha. Efficient hash function-based duplication detection algorithm for data deduplication deduction and reduction. Concurrency and Computation: Practice and Experience, 33(3):e5213:1–e5213:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Parimanam:2022:HOB

[PLP22] Kavitha Parimanam, Latha Lakshmanan, and Thangam Palaniswamy. Hybrid optimization based learning technique for multi-disease analytics from healthcare big data using optimal pre-processing, clustering and classifier. Concurrency and Computation: Practice and Experience, 34(17):e6986:1– e6986:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pei:2020:ECE

[PLX20] Yi-Lei Pei, Dan-Dan Li, and Wan-Xin Xue. The evaluation of customer experience using BP neural network-taking catering O2O takeout. Concurrency and Computation: Practice and Experience, 32(23):e5515:1–e5515:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2024:PIL

[PLX^{+}24] Houfu Peng, Xing Lu, Lili Xu, Daoxun Xia, and Xiaoyao Xie. Parameter instance learning with enhanced vision transformers for aerial person re-identification. Concurrency and Computation: Practice and Experience, 36(12):e8045:1–e8045:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2024:FAT

[PLZL24] Kai Peng, Chengfang Ling, Bohai Zhao, and Victor C. M. Leung. A fairness-aware task offloading method in edgeenabled IIoT with multi-constraints using AGE-MOEA and weighted MMF. Concurrency and Computation: Practice and Experience, 36(10):e7858:1–e7858:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Prasad:2022:EFS

[PM22] Vedavathi Nagendra Prasad and Anil Kumar Kureekatil Muthappa. An efficient framework for the similarity prediction with query recommendation in e-learning system. Concurrency and Computation: Practice and Experience, 34(22): e7145:1–e7145:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pareek:2023:EDK

[PM23a] Gaurav Pareek and Sumana Maiti. Efficient dynamic keyaggregate cryptosystem for secure and flexible data sharing. Concurrency and Computation: Practice and Experience, 35 (19):e7553:1–e7553:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Prashanth:2023:CBE

[PM23b] G. S. Prashanth and P. Manjunatha. Cluster based energy efficient routing protocol for heterogeneous wireless sensor networks. Concurrency and Computation: Practice and Experience, 35(21):e7693:1–e7693:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Park:2021:MGC

[PMC^{+}21] Yoosang Park, Jonghyeok Mun, Jongsun Choi, Jaeyoung Choi, and Hoda Kim. A method to generate context information sets from analysis results with a unified abstraction model based on an extension of data enrichment scheme. Concurrency and Computation: Practice and Experience, 33 (19):e6117:1–e6117:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Plauth:2023:IDT

[PMP23] Max Plauth, Joan Bruguera Micó, and Andreas Polze. Improved data transfer efficiency for scale-out heterogeneous workloads using on-the-fly I/O link compression. Concurrency and Computation: Practice and Experience, 35(11):e6101:1– e6101:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perumal:2021:RTI

[PMR^{+}21] Varalakshmi Perumal, SureshKumar Murugaiyan, Pavithran Ravichandran, R. Venkatesan, and R. Sundar. Real time identification of anomalous events in coastal regions using deep learning techniques. Concurrency and Computation: Practice and Experience, 33(21):e6421:1–e6421:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Park:2021:CIR

[PMS^{+}21] Seong Wan Park, Libor Mesicek, Joohyun Shin, Kitae Bae, Kyungjin An, and Hoon Ko. Customizing intelligent recommendation study with multiple advisors based on hierarchy structured fuzzy-analytic hierarchy process. Concurrency and Computation: Practice and Experience, 33(23):e5930:1– e5930:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2020:RAM

[PP20a] Zong Peng and Beth Plale. Reliable access to massive restricted texts: Experience-based evaluation. Concurrency and Computation: Practice and Experience, 32(16):e5255:1– e5255:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pham:2020:NGL

[PP20b] Hoang Pham and David H. Pham. A novel generalized logistic dependent model to predict the presence of breast cancer based on biomarkers. Concurrency and Computation: Practice and Experience, 32(1):e5467:1–e5467:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pandian:2021:EAR

[PP21] Senthilkumar Kuppusamy Pandian and Sivakumar Poruran. An efficient approach for removing haze from single image using Gaussian pyramidal decomposition. Concurrency and Computation: Practice and Experience, 33(16):e6279:1– e6279:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Padmanaban:2022:EEA

[PPA22] Paruthi Ilam Vazhuthi Padmanaban, Manikandan Shanmugaperumal Periasamy, and Prasanth Aruchamy. An energyefficient auto clustering framework for enlarging quality of service in Internet of Things-enabled wireless sensor networks using fuzzy logic system. Concurrency and Computation: Practice and Experience, 34(25):e7269:1–e7269:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Parlavantzas:2020:SBF

[PPM^{+}20] Nikos Parlavantzas, Linh Manh Pham, Christine Morin, Sandie Arnoux, Gaël Beaunée, Luyuan Qi, Philippe Gontier, and Pauline Ezanno. A service-based framework for building and executing epidemic simulation applications in the cloud. Concurrency and Computation: Practice and Experience, 32 (5):e5554:1–e5554:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pattnaik:2022:MOA

[PPM22] Sanat K. Pattnaik, Sumanta Panda, and Debadutta Mishra. A multi-objective approach for local path planning of autonomous mobile robot based on metaheuristics. Concurrency and Computation: Practice and Experience, 34(10):e6801:1– e6801:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pham-Quoc:2021:HPF

[PQKDT21] Cuong Pham-Quoc, Binh Kieu-Do, and Tran Ngoc Thinh. A high-performance FPGA-based BWA-MEM DNA sequence alignment. Concurrency and Computation: Practice and Experience, 33(2):e5328:1–e5328:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pathak:2022:CBI

[PR22] Debanjan Pathak and Undi Surya Narayana Raju. Contentbased image retrieval for super-resolutioned images using feature fusion: Deep learning and hand crafted. Concurrency and Computation: Practice and Experience, 34(22):e6851:1– e6851:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Parmar:2023:PPE

[PR23] Dilay Parmar and Udai Pratap Rao. Privacy-preserving enhanced dummy-generation technique for location-based services. Concurrency and Computation: Practice and Experience, 35(2):e7501:1–e7501:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Passos:2024:APS

[PR24] Fernanda G. O. Passos and Vinod E. F. Rebello. An autonomic parallel strategy for exhaustive search tree algorithms on shared or heterogeneous systems. Concurrency and Computation: Practice and Experience, 36(6):e7955:1–e7955:??, March 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Platini:2021:COP

[PRPD21] Marc Platini, Thomas Ropars, Benoit Pelletier, and Noel De Palma. CPU overheating prediction in HPC systems. Concurrency and Computation: Practice and Experience, 33(13): e6231:1–e6231:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Penaranda:2023:EUD

[PRS23] Cristian Peñaranda, Carlos Reaño, and Federico Silla. Exploring the use of data compression for accelerating machine learning in the edge with remote virtual graphics processing units. Concurrency and Computation: Practice and Experience, 35(20):e7328:1–e7328:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perente:2021:MPC

[PS21] Osman Kerem Perente and Tacha Serif. Multienvironment performance comparison of robot-assisted indoor location estimation system. Concurrency and Computation: Practice and Experience, 33(23):e6073:1–e6073:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

P:2022:NMM

[PS22a] Kumaran P. and Chitrakala S. A novel mathematical modeling in shift in emotion for gauging the social influential in big data streams with hybrid sarcasm detection. Concurrency and Computation: Practice and Experience, 34(3):e6597:1– e6597:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pandey:2022:MMK

[PS22b] Kamlesh Kumar Pandey and Diwakar Shukla. Min-max kurtosis stratum mean: an improved K-means cluster initialization approach for microarray gene clustering on multidimensional big data. Concurrency and Computation: Practice and Experience, 34(23):e7185:1–e7185:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pawar:2022:CFA

[PS22c] Vijayant Pawar and Shelly Sachdeva. CovidBChain: Framework for access-control, authentication, and integrity of Covid-19 data. Concurrency and Computation: Practice and Experience, 34(28):e7397:1–e7397:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pathan:2023:NBT

[PS23a] Summiya A. Pathan and Noorullah Shariff. A novel blockchain tumbling scalability framework in a bitcoin network. Concurrency and Computation: Practice and Experience, 35(21):e7634:1–e7634:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Priya:2023:DAE

[PS23b] M. Mary Adline Priya and Joseph Jawhar S. Dropout AlexNet-extreme learning optimized with fast gradient descent optimization algorithm for brain tumor classification. Concurrency and Computation: Practice and Experience, 35 (6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pushpa:2023:FAB

[PS23c] Ramaiah Pushpa and Maadappa Siddappa. Fractional artificial bee chicken swarm optimization technique for QoS aware virtual machine placement in cloud. Concurrency and Computation: Practice and Experience, 35(4):e7532:1–e7532:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Patonico:2020:ECB

[PSB^{+}20] Simone Patonico, Placide Shabisha, An Braeken, Abdellah Touhafi, and Kris Steenhaut. Elliptic curve-based proxy re-signcryption scheme for secure data storage on the cloud. Concurrency and Computation: Practice and Experience, 32(17):e5657:1–e5657:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Promberger:2023:CDC

[PSF23] Laura Promberger, Rainer Schwemmer, and Holger Fröning. Characterization of data compression across CPU platforms and accelerators. Concurrency and Computation: Practice and Experience, 35(20):e6465:1–e6465:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Perez-Sola:2020:BTD

[PSHJ20] Cristina Pérez-Solà and Jordi Herrera-Joancomartí. BArt: Trading digital contents through digital assets. Concurrency and Computation: Practice and Experience, 32(12):e5490:1– e5490:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Preethi:2023:CCG

[PSK23] Bc Preethi, G Sugitha, and G Kavitha. Cycle-consistent generative adversarial network optimized with water strider optimization algorithm fostered intrusion detection framework for securing cloud computing environment. Concurrency and Computation: Practice and Experience, 35(5):e7552:1– e7552:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pandi:2022:OSA

[PSMM22] S. Senthil Pandi, A. Senthilselvi, M. Maragatharajan, and I. Manju. An optimal self adaptive deep neural network and spine-kernelled chirplet transform for image registration. Concurrency and Computation: Practice and Experience, 34(27): e7297:1–e7297:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Papan:2020:NMR

[PSP^{+}20] Jozef Papan, Pavel Segec, Peter Paluch, Jana Uramova, and Marek Moravcik. The new Multicast Repair (M-REP) IP fast reroute mechanism. Concurrency and Computation: Practice and Experience, 32(13):e5105:1–e5105:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Palanisamy:2022:NNB

[PSP22] Tamilselvi Palanisamy, Geetha Sadayan, and Nagasankar Pathinetampadiyan. Neural network-based leaf classification using machine learning. Concurrency and Computation: Practice and Experience, 34(8):e5366:1–e5366:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Patel:2024:NPS

[PSP24] Khushboo A. Patel, Shivangi Shukla, and Sankita J. Patel. A novel and provably secure mutual authentication protocol for cloud environment using elliptic curve cryptography and fuzzy verifier. Concurrency and Computation: Practice and Experience, 36(2):e7889:1–e7889:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Patra:2024:FFP

[PST24] Manoj Kumar Patra, Bibhudatta Sahoo, and Ashok Kumar Turuk. FPHO: Fractional Pelican Hawks optimization based container consolidation in CaaS cloud. Concurrency and Computation: Practice and Experience, 36(12):e8052:1– e8052:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Palanisamy:2022:EHG

[PT22] Gajalakshmi Palanisamy and Sree Sharmila Thangaswamy. An efficient hand gesture recognition based on optimal deep embedded hybrid convolutional neural network-long short term memory network model. Concurrency and Computation: Practice and Experience, 34(21):e7109:1–e7109:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peddoju:2020:FIM

[PUL20] Suresh K. Peddoju, Himanshu Upadhyay, and Leonel Lagos. File integrity monitoring tools: Issues, challenges, and solutions. Concurrency and Computation: Practice and Experience, 32(22):e5825:1–e5825:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pisani:2020:UMA

[PVP^{+}20] Flávia Pisani, Lucas Pascotti Valem, Daniel Carlos Guimarães Pedronette, Ricardo da S. Torres, Edson Borin, and Mauricio Breternitz Jr. A unified model for accelerating unsupervised iterative re-ranking algorithms. Concurrency and Computation: Practice and Experience, 32(14):e5702:1–e5702:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Prabhakar:2022:GWS

[PVRM22] Telagarapu Prabhakar, Subbiah Vairamuthu, Balasubramaniam Selva Rani, and Balajee Maram. Gray wolf-student psychology optimization-based deep long short term memory for survival prediction using cancer gene-expression data. Concurrency and Computation: Practice and Experience, 34 (23):e7206:1–e7206:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pullaiah:2022:BDE

[PVVS22] Nagaraja Rao Pamula Pullaiah, Dorai Venkatasekhar, Padarthi Venkatramana, and Balaraj Sudhakar. Binary differential evolution with self learning and deep neural network for breast cancer classification. Concurrency and Computation: Practice and Experience, 34(15):e6951:1–e6951:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pei:2021:GSP

[PWJ^{+}21] Songwen Pei, Jinkai Wang, Linhua Jiang, Naixue Xiong, and Jean-Luc Gaudiot. Genetic scheduling policy on codelet model. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Park:2020:EAS

[PYC^{+}20] Yoosang Park, Jaehyung Ye, Jongsun Choi, Jaeyoung Choi, and Hoon Ko. Extraction of abstracted sensory data to reduce the execution time of context-aware services in wearable computing environments. Concurrency and Computation: Practice and Experience, 32(18):e5286:1–e5286:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2021:CFA

[PZD^{+}21] Hu Peng, Wenhua Zhu, Changshou Deng, Kun Yu, and Zhijian Wu. Composite firefly algorithm for breast cancer recognition. Concurrency and Computation: Practice and Experience, 33(5):e6032:1–e6032:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Peng:2023:SFH

[PZZ^{+}23] Zhiniang Peng, Wei Zhou, Xiaogang Zhu, Youke Wu, and Sheng Wen. On the security of fully homomorphic encryption for data privacy in Internet of Things. Concurrency and Computation: Practice and Experience, 35(19):e7330:1–e7330:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiao:2023:OWS

[QCC^{+}23] Qihang Qiao, Lan Chen, Hong Cai, He Zhang, and Zhenjie Yao. An online workflow scheduling algorithm considering license limitation in heterogeneous environment. Concurrency and Computation: Practice and Experience, 35(2):e7452:1– e7452:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiu:2024:TOP

[QHL^{+}24] Wanyu Qiu, Chuanhe Huang, Zhengfa Li, M. Wasim Abbas Ashraf, and Danxin Wang. Traffic offloading for postdisaster rescue in UAV-assisted networks: a coalition formation game approach. Concurrency and Computation: Practice and Experience, 36(10):e7998:1–e7998:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qi:2023:DPT

[QKSK23] Bin Qi, Kazuhiko Komatsu, Masayuki Sato, and Hiroaki Kobayashi. A dynamic parameter tuning method for SpMM parallel execution. Concurrency and Computation: Practice and Experience, 35(17):e6755:1–e6755:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qiao:2022:TMC

[QLL^{+}22] Bingxue Qiao, Chubo Liu, Jing Liu, Yikun Hu, Kenli Li, and Keqin Li. Task migration computation offloading with low delay for mobile edge computing in vehicular networks. Concurrency and Computation: Practice and Experience, 34 (1):e6494:1–e6494:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qin:2020:BRC

[QMC^{+}20] Jiayu Qin, Gang Mei, Salvatore Cuomo, Sixu Guo, and Yixuan Li. Book review: CudaCHPre2D: a straightforward preprocessing approach for accelerating 2D convex hull computations on the GPU. Concurrency and Computation: Practice and Experience, 32(10):e5229:1–e5229:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Quan:2022:EFT

[QNZP22] DongXiao Quan, Li Niu, LiLi Zhu, and ChangXing Pei. Efficient fault-tolerant logical Hadamard gates implementation in Reed–Muller quantum codes. Concurrency and Computation: Practice and Experience, 34(2):e6079:1–e6079:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Quach:2020:GSA

[QPS20] Dara Q. Quach, Daniel P. Playne, and Chris J. Scogings. GPUAnimats — simulating animats, an agent-based, artificial life model on graphical processing units. Concurrency and Computation: Practice and Experience, 32(24):e5968:1– e5968:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qin:2024:ODE

[QWL24] Wei Qin, Hao Wang, and Xiangfeng Luo. Open-domain event schema induction via weighted attentive hypergraph neural network. Concurrency and Computation: Practice and Experience, 36(12):e8029:1–e8029:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qu:2022:EVP

[QWW^{+}22] Wenlei Qu, Lei Wu, Wei Wang, Zhaoman Liu, and Hao Wang. A electronic voting protocol based on blockchain and homomorphic signcryption. Concurrency and Computation: Practice and Experience, 34(16):e5817:1–e5817:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Qi:2023:UPI

[QZB^{+}23] Fang Qi, Yingkai Zhao, Md Zakirul Alam Bhuiyan, Tao Hai, Monirul Islam, Shaobo Zhang, and Zhe Tang. Unauthorized and privacy-intrusive human activity watching through Wi-Fi signals: an emerging cybersecurity threat. Concurrency and Computation: Practice and Experience, 35(19):e7313:1– e7313:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rajangam:2021:TEU

[RA21] Engels Rajangam and Chitra Annamalai. Topic extraction using local graph centrality and semantic similarity. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Raouf:2021:PRM

[RAaB21] Ahmed E. Abdel Raouf, Alshaimaa Abo-alian, and Nagwa L. Badr. A predictive replication for multi-tenant databases using deep learning. Concurrency and Computation: Practice and Experience, 33(13):e6226:1–e6226:??, July 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Renukadevi:2021:EPP

[RAG21] M. Renukadevi, E. A. Mary Anita, and D. Mohana Geetha. An efficient privacy-preserving model based on OMFTSA for query optimization in crowdsourcing. Concurrency and Computation: Practice and Experience, 33(24):e6447:1–e6447:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rehman:2022:PPI

[RAK22] Akif Rehman, Masab Ahmad, and Omer Khan. A performance predictor for implementation selection of parallelized static and temporal graph algorithms. Concurrency and Computation: Practice and Experience, 34(2):e6267:1–e6267:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rajesh:2021:LPD

[RAN21] R. Rajesh, C. Annadurai, and K. Nirmaladevi. Low power device coordination in internet of things environment using analytic hierarchy process model. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramkumar:2022:IBC

[RAN22] D. Ramkumar, C. Annadurai, and I. Nelson. Iris-based continuous authentication in mobile ad hoc network. Concurrency and Computation: Practice and Experience, 34(8): e5542:1–e5542:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Raviprabhakaran:2023:QSC

[Rav23] Vijay Raviprabhakaran. Quorum sensing centered bacterial horde algorithm for global optimization. Concurrency and Computation: Practice and Experience, 35(8):e7627:1– e7627:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

R:2022:BSB

[RB22] Naga Priyadarsini R and Ponsy Rk Sathia Bhama. B-SCORE — a blockchain based hybrid chaotic signatures for medical image encryption in an IoT environment. Concurrency and Computation: Practice and Experience, 34(21):e7115:1– e7115:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Reis:2020:CMC

[RBC20] Luís Reis, João Bispo, and João M. P. Cardoso. Compilation of MATLAB computations to CPU/GPU via C/OpenCL generation. Concurrency and Computation: Practice and Experience, 32(22):e5854:1–e5854:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rani:2022:BAR

[RBDS22] Neetu Rani, Amit Kumar Bhardwaj, Prasenjit Das, and Anju Sharma. Bibliometric analysis of rumor detection via web of science from 1989 to 2021. Concurrency and Computation: Practice and Experience, 34(25):e7260:1–e7260:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rjoub:2021:DRL

[RBWB21] Gaith Rjoub, Jamal Bentahar, Omar Abdel Wahab, and Ahmed Saleh Bataineh. Deep and reinforcement learning for automated task scheduling in large-scale cloud computing systems. Concurrency and Computation: Practice and Experience, 33(23):e5919:1–e5919:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

R:2022:BDA

[RC22] Mahaveerakannan R and Suresh Gnana Dhas C. Big data analytics for large-scale UAV-MBN in quantum networks using efficient hybrid GKM. Concurrency and Computation: Practice and Experience, 34(7):e5559:1–e5559:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rodriguez-Canal:2023:TBP

[RCBTGE23] Gabriel Rodriguez-Canal, Nick Brown, Yuri Torres, and Arturo Gonzalez-Escribano. Task-based preemptive scheduling on FPGAs leveraging partial reconfiguration. Concurrency and Computation: Practice and Experience, 35(25):e7867:1– e7867:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rao:2022:MEV

[RCK22] B. Janardhana Rao, Y. Chakrapani, and S. Srinivas Kumar. MABC-EPF: Video in-painting technique with enhanced priority function and optimal patch search algorithm. Concurrency and Computation: Practice and Experience, 34(11): e6840:1–e6840:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rong:2020:PTS

[RCS20] Chuitian Rong, Lili Chen, and Yasin N. Silva. Parallel time series join using Spark. Concurrency and Computation: Practice and Experience, 32(9):e5622:1–e5622:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramay:2021:HAB

[RCYuRH21] Waheed Yousuf Ramay, Xu Cheng-Yin, Shams ur Rahman, and Muhammad Asif Habib. Hybrid approach for big data localization and semantic annotation. Concurrency and Computation: Practice and Experience, 33(4):e4955:1–e4955:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Robert:2023:RCM

[RD23] Eunice Jennifer Robert and Hemanth Jude Duraisamy. A review on computational methods based automated sign language recognition system for hearing and speech impaired community. Concurrency and Computation: Practice and Experience, 35(9):e7653:1–e7653:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rani:2022:RMA

[RDB22] Neetu Rani, Prasenjit Das, and Amit Kumar Bhardwaj. Rumor, misinformation among web: a contemporary review of rumor detection techniques during different web waves. Concurrency and Computation: Practice and Experience, 34 (1):e6479:1–e6479:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rakib:2021:MCO

[RF21] Abdur Rakib and Rokan Uddin Faruqui. Model checking ontology-driven reasoning agents using strategy and abstraction. Concurrency and Computation: Practice and Experience, 33(2):e5205:1–e5205:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rizvi:2023:RMC

[RF23] Syed Khurram Jah Rizvi and Muhammad Moazam Fraz. Robust malware clustering of windows portable executables using ensemble latent representation and distribution modeling. Concurrency and Computation: Practice and Experience, 35(8):e7621:1–e7621:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rao:2022:SVM

[RG22] K. Venkateswara Rao and L. Mary Gladence. Support vector machine based disease classification model employing hasten eagle Cuculidae search optimization. Concurrency and Computation: Practice and Experience, 34(25):e7259:1–e7259:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rezaei:2023:QBR

[RG23] Hesam Rezaei and Ali Ghaffari. QoS-based routing scheme in software-defined networks using fuzzy analytic hierarchy process. Concurrency and Computation: Practice and Experience, 35(9):e7638:1–e7638:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rajeswari:2021:ETB

[RGKK21] Alagan Ramasamy Rajeswari, Sannasi Ganapathy, Kanagasabai Kulothungan, and Arputharaj Kannan. An efficient trust-based secure energy-aware clustering to mitigate trust distortion attack in mobile ad-hoc network. Concurrency and Computation: Practice and Experience, 33(13):e6223:1– e6223:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Renney:2022:SSG

[RGM22] Harri Renney, Benedict Gaster, and Thomas J. Mitchell. Survival of the synthesis-GPU accelerating evolutionary sound matching. Concurrency and Computation: Practice and Experience, 34(10):e6824:1–e6824:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rivas:2023:PCV

[RGPC23] Daniel Rivas, Francesc Guim, Jordà Polo, and David Carrera. Performance characterization of video analytics workloads in heterogeneous edge infrastructures. Concurrency and Computation: Practice and Experience, 35(14):e6317:1–e6317:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rajesh:2022:ECC

[RJ22] Dennison Rajesh and Thirasamma Jaya. Energy competent cluster-based secured CH routing EC^{2} SR protocol for mobile wireless sensor network. Concurrency and Computation: Practice and Experience, 34(1):e6525:1–e6525:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Raj:2023:SES

[RJH^{+}23] R. S. Reshma Raj, Athul Joseph, Akhil Sarwar T. H., Saidalavi Kalady, and Jayaraj P. B. SpecMap: an efficient spectral partitioning based static application mapping algorithm for 2D mesh NoCs. Concurrency and Computation: Practice and Experience, 35(26):e7838:1–e7838:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramani:2021:NKR

[RK21a] R. Geetha Ramani and S. Suresh Kumar. Nonvolatile kernel rootkit detection using cross-view clean boot in cloud computing. Concurrency and Computation: Practice and Experience, 33(3):e5239:1–e5239:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rani:2021:SMO

[RK21b] M. Thillai Rani and S. Suresh Kumar. Spider monkey optimization-based high-level synthesis in VLSI circuits for runtime adaptability. Concurrency and Computation: Practice and Experience, 33(3):e5280:1–e5280:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramakrishna:2023:UAB

[RK23a] Malige Mahadevaiah Ramakrishna and Vedanandam Karthikeyan. UChOA-ANN based effective diagnostic model for early detection and accurate prediction of dengue affected blood samples. Concurrency and Computation: Practice and Experience, 35(8):e7597:1–e7597:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramasamy:2023:HCS

[RK23b] Sathya Ramasamy and Vaidehi Kaliyaperumal. A hybridized channel selection approach with deep convolutional neural network for effective ovarian cancer prediction in periodic acid-Schiff-stained images. Concurrency and Computation: Practice and Experience, 35(5):e7568:1–e7568:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Raut:2023:IMP

[RK23c] Archana R. Raut and S. P. Khandait. An intelligent MAC protocol for mission critical applications in wireless sensor networks. Concurrency and Computation: Practice and Experience, 35(23):e7813:1–e7813:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Roy:2022:ANF

[RKC22] Kallol Roy, Mandal Kamal Krishna, and Mandal Atis Chandra. Adaptive neuro fuzzy inference system with elephant herding optimization based energy management scheme. Concurrency and Computation: Practice and Experience, 34(21): e7061:1–e7061:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rim:2023:GDQ

[RKK23] KwangCheol Rim, Pankoo Kim, and Hoon Ko. Gradient descent for quadratic functions using geometric mean and the Kai Fang method. Concurrency and Computation: Practice and Experience, 35(16):e6605:1–e6605:??, July 25, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ryu:2021:EVE

[RKL21] Seokhoon Ryu, Seonghyeon Kim, and Young-Sup Lee. Effect of virtual error sensor location for active sound quality control in a car cabin. Concurrency and Computation: Practice and Experience, 33(22):e6005:1–e6005:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rathi:2024:QSS

[RKMS24] Deepa Rathi, Sanjeev Kumar, Farhan Musanna, and Deepak Singh. Quantum secret sharing protocol through noisy channel with application in visual cryptography. Concurrency and Computation: Practice and Experience, 36(2):e7899:1– e7899:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

R:2022:IDC

[RKR22] Dhivyaa C. R, Nithya Kandasamy, and Sudhakar Rajendran. Integration of dilated convolution with residual dense block network and multi-level feature detection network for cassava plant leaf disease identification. Concurrency and Computation: Practice and Experience, 34(11):e6879:1–e6879:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rehman:2020:SMR

[RKuH^{+}20] Faisal Rehman, Osman Khalid, Nuhman ul Haq, Tahir Maqsood, Mazhar Ali, Raja Wasim Ahmad, Junaid Shuja, Shahzad Sarwar, and Sajjad Ahmad Madani. A scalable model for real-time venue recommendations using MapReduce. Concurrency and Computation: Practice and Experience, 32(21):e5597:1–e5597:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ribeiro:2020:OPQ

[RLdO20] Marcello W. M. Ribeiro, Alexandre A. B. Lima, and Daniel de Oliveira. OLAP parallel query processing in clouds with C-ParGRES. Concurrency and Computation: Practice and Experience, 32(7):e5590:1–e5590:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rajamanickam:2021:KCS

[RM21] Jayalakshmi Rajamanickam and Savitha Devi Mani. Kullback chi square and Gustafson Kessel probabilistic neural network based soil fertility prediction. Concurrency and Computation: Practice and Experience, 33(24):e6460:1–e6460:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ranjan:2022:PUS

[RM22] Sakshi Ranjan and Subhankar Mishra. Perceiving university students’ opinions from Google app reviews. Concurrency and Computation: Practice and Experience, 34(10):e6800:1– e6800:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramachandran:2023:AAB

[RM23] Arun Kumar Ramachandran and Vijay Franklin John Bosco Martin. Adaptive autism behavior prediction using improved binary whale optimization technique. Concurrency and Computation: Practice and Experience, 35(3):e7511:1–e7511:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramamurthy:2023:NTS

[RMMK23] Karthik Ramamurthy, Anirudh Muthuswamy, Nagharjun Mathimariappan, and Gugan S. Kathiresan. A novel twostaged network for skin disease detection using atrous residual convolutional networks. Concurrency and Computation: Practice and Experience, 35(26):e7834:1–e7834:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Riaz:2022:MCC

[RMR^{+}22] Muhammad Riaz, Rashid Mehmood, Masooda Rehman, Iftikhar Ali, and Babar Zaman. Mean control chart based on ranked set schemes for unknown skewed probability distribution and parameters. Concurrency and Computation: Practice and Experience, 34(21):e7065:1–e7065:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Retnamony:2022:EGL

[RMS22] Jeen Retna Kumar Retnamony, Sundaram Muniasamy, and Berakhah Florence Stanley. Enhanced global and local face feature extraction for effective recognition of facial emotions. Concurrency and Computation: Practice and Experience, 34 (5):e6701:1–e6701:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rickett:2024:OCG

[RMS24] Christopher D. Rickett, Kristyn J. Maschhoff, and Sreenivas R. Sukumar. Optimizing the Cray Graph Engine for performant analytics on cluster, SuperDome Flex, Shasta systems and cloud deployment. Concurrency and Computation: Practice and Experience, 36(10):e7982:1–e7982:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Raju:2022:DDK

[RN22] N. V. S. Lakshmipathi Raju and Vankamamidi S. Naresh. Dynamic distributed KC_{i}-slice data publishing model with multiple sensitive attributes. Concurrency and Computation: Practice and Experience, 34(21):e7064:1–e7064:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Remani:2022:CDO

[RNRK22] Naga Venkata Jagan Mohan Remani, Vankamamidi S. Naresh, Sivaranjani Reddi, and Kadali Dileep Kumar. Crime data optimization using neutrosophic logic based game theory. Concurrency and Computation: Practice and Experience, 34(15):e6973:1–e6973:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Renklier:2022:NFC

[RÖ22] Ayhan Renklier and Serkan Öztürk. A novel Frei–Chen based fragile watermarking method for authentication of an image. Concurrency and Computation: Practice and Experience, 34 (22):e6897:1–e6897:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rajesh:2021:HAB

[RP21a] Kolluri Rajesh and Sumanta Pyne. A hybrid artificial bee colony algorithm for scheduling of digital microfluidic biochip operations. Concurrency and Computation: Practice and Experience, 33(13):e6223:1–e6223:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rangriz:2021:LMF

[RP21b] Emad Rangriz and Vahid Pourahmadi. Language mapping functions: Improving softmax estimation and word embedding quality. Concurrency and Computation: Practice and Experience, 33(24):e6464:1–e6464:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rani:2021:RFO

[RP21c] Antony Jaya Mabel Rani and Albert Pravin. Rainfall flow optimization based K-means clustering for medical data. Concurrency and Computation: Practice and Experience, 33(17): e6308:1–e6308:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Reddy:2022:RBG

[RP22] Vandana Reddy and Gayathri P. Revised beaconing glowworm swarm optimization ant colony optimization algorithm to localize nodes and optimize the energy consumed by nodes in wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(13):e6013:1–e6013:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ren:2022:SNC

[RPM22] Yuxin Ren, Gabriel Parmer, and Dejan Milojicic. Sharing non-cache-coherent memory with bounded incoherence. Concurrency and Computation: Practice and Experience, 34 (2):e6414:1–e6414:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramkumar:2022:DMN

[RPMA22] Muthuperumal Periyaperumal Ramkumar, Pauliah David Mano Paul, Balajee Maram, and John Patrick Ananth. Deep Maxout network for lung cancer detection using optimization algorithm in smart Internet of Things. Concurrency and Computation: Practice and Experience, 34(25):e7264:1–e7264:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rajeswari:2023:BVD

[RPPK23] Ramasamy Rajeswari, Manickam Prabhakar, Govindhan Padmapriya, and Balan Santhosh Kumar. Blood vessel detection using enhanced DeepJoint fuzzy clustering algorithm with deep Maxout network for glaucoma detection. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramathulasi:2022:ALD

[RR22] Thota Ramathulasi and M. Rajasekharababu. Augmented latent Dirichlet allocation model via word embedded clusters for mashup service clustering. Concurrency and Computation: Practice and Experience, 34(15):e6896:1–e6896:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rajamoney:2023:RDD

[RR23a] Jansi Rajamoney and Amutha Ramachandran. Representativediscriminative dictionary learning algorithm for human action recognition using smartphone sensors. Concurrency and Computation: Practice and Experience, 35(2):e7468:1–e7468:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Raju:2023:LCC

[RR23b] Mallela Siva Naga Raju and Battula Srinivasa Rao. Lung and colon cancer classification using hybrid principle component analysis network-extreme learning machine. Concurrency and Computation: Practice and Experience, 35(1):e7361:1– e7361:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rahmati:2022:ACS

[RRB22] Mohammad Rahmati, Farbod Razzazi, and Alireza Behrad. Adaptive content suppression based on shared-hidden-layer convolutional autoencoder for blind image forensics. Concurrency and Computation: Practice and Experience, 34(26): e7278:1–e7278:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rankovic:2022:IIV

[RRIL22] Nevena Rankovic, Dragica Rankovic, Mirjana Ivanovic, and Ljubomir Lazic. Influence of input values on the prediction model error using artificial neural network based on Taguchi’s orthogonal array. Concurrency and Computation: Practice and Experience, 34(20):e6831:1–e6831:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramesh:2023:HSV

[RRJ23] Lilly Ramesh, S. Radhika, and S. Jothi. Hybrid support vector machine and K-nearest neighbor-based software testing for educational assistant. Concurrency and Computation: Practice and Experience, 35(1):e7433:1–e7433:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramesh:2020:CSD

[RS20] T. Ramesh and R. M. Suresh. Co-scheduling of data intensive jobs and processor redistribution under temperature constraints. Concurrency and Computation: Practice and Experience, 32(4):e5033:1–e5033:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Reano:2021:RRC

[RS21] Carlos Reaño and Federico Silla. Redesigning the rCUDA communication layer for a better adaptation to the underlying hardware. Concurrency and Computation: Practice and Experience, 33(14):e5481:1–e5481:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rajput:2023:MAA

[RS23] Pushpendra Kumar Rajput and Geeta Sikka. Multi-agent architecture approach for self-healing systems: Run-time recovery with case-based reasoning. Concurrency and Computation: Practice and Experience, 35(1):e7442:1–e7442:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Reshi:2024:BCD

[RS24] Iraq Ahmad Reshi and Sahil Sholla. The blockchain conundrum: an in-depth examination of challenges, contributing technologies, and alternatives. Concurrency and Computation: Practice and Experience, 36(8):e7987:1–e7987:??, April 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Renjith:2021:SEE

[RSJ21] Shini Renjith, A. Sreekumar, and M. Jathavedan. Sem-Rec — an efficient ensemble recommender with sentiment based clustering for social media text corpus. Concurrency and Computation: Practice and Experience, 33(20):e6359:1– e6359:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Renjith:2022:TGA

[RSJ22] Shini Renjith, A. Sreekumar, and M. Jathavedan. Taxonomy grooming algorithm — an autodidactic domain specific dimensionality reduction approach for fast clustering of social media text data. Concurrency and Computation: Practice and Experience, 34(11):e6837:1–e6837:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramasamy:2023:HLD

[RSKA23] Sathya Ramasamy, Ananthi Selvarajan, Vaidehi Kaliyaperumal, and Prasanth Aruchamy. A hybrid location-dependent ultra convolutional neural network-based vehicle number plate recognition approach for intelligent transportation systems. Concurrency and Computation: Practice and Experience, 35(8):e7615:1–e7615:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ram:2021:VTL

[RSM21] Satya Deo K. Ram, Shashank Srivastava, and Krishn Kumar Mishra. A variant of teaching-learning-based optimization and its application for minimizing the cost of workflow execution in the cloud computing. Concurrency and Computation: Practice and Experience, 33(21):e6425:1–e6425:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ram:2022:NMH

[RSM22] Satya Deo Kumar Ram, Shashank Srivastava, and Krishn Kumar Mishra. A new meta-heuristic approach for load awarecost effective workflow scheduling. Concurrency and Computation: Practice and Experience, 34(21):e7112:1–e7112:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ram:2023:RLM

[RSM23] Satya Deo Kumar Ram, Shashank Srivastava, and Krishn Kumar Mishra. Redefining the learning mechanism in teachinglearning-based optimization and its applications for flowtimeaware-cost minimizing of the workflow in cloud. Concurrency and Computation: Practice and Experience, 35(23):e7762:1– e7762:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rizzardi:2022:SAC

[RSMCP22] Alessandra Rizzardi, Sabrina Sicari, Daniele Miorandi, and Alberto Coen-Porisini. Securing the access control policies to the Internet of Things resources through permissioned blockchain. Concurrency and Computation: Practice and Experience, 34(15):e6934:1–e6934:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

raj:2021:CSB

[rSN21] T. Reena raj, K. Sakthidasan @ Sankaran, and V. Nagarajan. Chaotic sequence-based MC-CDMA for 5G. Concurrency and Computation: Practice and Experience, 33(4):e4992:1– e4992:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rajeswari:2022:EID

[RSR^{+}22] Poleboina Venkata Naga Rajeswari, Mogalla Shashi, Tiyyabindi Kameswara Rao, Matta Rajya Lakshmi, and Lanka Venkateswara Kiran. Effective intrusion detection system using concept drifting data stream and support vector machine. Concurrency and Computation: Practice and Experience, 34(21):e7118:1– e7118:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rajkumar:2020:ECB

[RSS20] M. Rajkumar, R. M. Suresh, and R. Sasikumar. An effective cluster based data dissemination in a hybrid cellular ad hoc network. Concurrency and Computation: Practice and Experience, 32(4):e5125:1–e5125:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Reano:2024:AV

[RSV24] Carlos Reaño, Federico Silla, and Blesson Varghese. Accelerator virtualization. Concurrency and Computation: Practice and Experience, 36(10):e6254:1–e6254:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Robinson:2020:FSI

[RT20] Timothy W. Robinson and Abhinav Thota. Foreword to the special issue of the Cray User Group. Concurrency and Computation: Practice and Experience, 32(20):e5755:1–e5755:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramanathan:2021:MSF

[RT21] Kaviyarasi Ramanathan and Balasubramanian Thangavel. Minkowski sommon feature map-based densely connected deep convolution network with LSTM for academic performance prediction. Concurrency and Computation: Practice and Experience, 33(13):e6244:1–e6244:??, July 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramanathan:2022:TPP

[RT22a] Kaviyarasi Ramanathan and Balasubramanian Thangavel. Targeted projection pursuit similarity based attribute selection for academic performance prediction. Concurrency and Computation: Practice and Experience, 34(21):e7102:1– e7102:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Roberts:2022:OTM

[RT22b] Michaelraj Kingston Roberts and Jayapratha Thangavel. An optimized ticket manager based energy-aware multipath routing protocol design for IoT based wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(28):e7398:1–e7398:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rubinpur:2022:SPR

[RT22c] Yaniv Rubinpur and Sivan Toledo. Signal processing for a reverse-GPS wildlife tracking system: CPU and GPU implementation experiences. Concurrency and Computation: Practice and Experience, 34(14):e6506:1–e6506:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ramamurthy:2023:NDL

[RTBC23] Karthik Ramamurthy, Rashmi Dinesh Thekkath, Shivam Batra, and Sreejan Chattopadhyay. A novel deep learning architecture for disease classification in Arabica coffee plants. Concurrency and Computation: Practice and Experience, 35(8): e7625:1–e7625:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Regunta:2023:EPA

[RTSK23] Sai Charan Regunta, Sai Harsh Tondomker, Kshitij Shukla, and Kishore Kothapalli. Efficient parallel algorithms for dynamic closeness- and betweenness centrality. Concurrency and Computation: Practice and Experience, 35(17):e6650:1– e6650:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

R:2021:MWD

[RV21] Menaha R and Jayanthi Ve. Mining the web to discover acronym-definitions based on sequence labeling and iterative query expansion model. Concurrency and Computation: Practice and Experience, 33(17):e6291:1–e6291:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Reddy:2023:EDW

[RV23] Tatiparti B. Prasad Reddy and D Vydeki. Ebola deep wavelet extreme learning machine based chronic kidney disease prediction on the internet of medical things platform. Concurrency and Computation: Practice and Experience, 35(1): e7446:1–e7446:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rasouli:2021:PDI

[RVAE21] Hatef Rasouli, Changiz Valmohammadi, Naser Azad, and Ghanbar Abbaspour Esfeden. Proposing a digital identity management framework: a mixed-method approach. Concurrency and Computation: Practice and Experience, 33(17): e6271:1–e6271:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rahmani:2022:DFT

[RVF22] Farzaneh Rahmani, Changiz Valmohammadi, and Kiamars Fathi. Detecting fraudulent transactions in banking cards using scale-free graphs. Concurrency and Computation: Practice and Experience, 34(19):e7028:1–e7028:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rathi:2022:PPI

[RVJ^{+}22] Shubhangi Rathi, Jai Prakash Verma, Rachna Jain, Anand Nayyar, and Narina Thakur. Psychometric profiling of individuals using Twitter profiles: a psychological Natural Language Processing based approach. Concurrency and Computation: Practice and Experience, 34(19):e7029:1–e7029:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rathod:2024:FCS

[RW24] Nilesh Rathod and Sunil Wankhade. Fractional cuckoo search invasive weed optimized neural network for data classification. Concurrency and Computation: Practice and Experience, 36 (6):e7948:1–e7948:??, March 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ru:2021:EDC

[RYG^{+}21] Jia Ru, Yun Yang, John Grundy, Jacky Keung, and Li Hao. An efficient deadline constrained and data locality aware dynamic scheduling framework for multitenancy clouds. Concurrency and Computation: Practice and Experience, 33(5): e6037:1–e6037:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ren:2021:NSD

[RZ21] Junhua Ren and Qing Zhang. A novel software defect prediction approach using modified objective cluster analysis. Concurrency and Computation: Practice and Experience, 33(9): e6112:1–e6112:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rong:2023:DSC

[RZ23] Chuitian Rong and Jinyu Zhou. Distributed structural clustering on large graph. Concurrency and Computation: Practice and Experience, 35(22):e7756:1–e7756:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Roshan:2024:AIR

[RZ24] Khushnaseeb Roshan and Aasim Zafar. AE-Integrated: Realtime network intrusion detection with Apache Kafka and autoencoder. Concurrency and Computation: Practice and Experience, 36(11):e8034:1–e8034:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ren:2021:EAM

[RZCA21] Xiaojun Ren, Zhijun Zhang, Shaochun Chen, and Karlo Abnoosian. An energy-aware method for task allocation in the Internet of Things using a hybrid optimization algorithm. Concurrency and Computation: Practice and Experience, 33 (6):e5967:1–e5967:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Robert:2021:CSB

[RZVC21] Sophie Robert, Soraya Zertal, Grégory Vaumourin, and Philippe Couvée. A comparative study of black-box optimization heuristics for online tuning of high performance computing I/O accelerators. Concurrency and Computation:

Practice and Experience, 33(16):e6274:1–e6274:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saha:2022:AVA

[SA22a] Souvik Saha and Rajeev Kumar Arya. Adaptive virtual anchor node based underwater localization using improved shortest path algorithm and particle swarm optimization (PSO) technique. Concurrency and Computation: Practice and Experience, 34(3):e6552:1–e6552:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saifuddin:2022:IPD

[SA22b] Shameem Saifuddin and Sivasubramanian Arunagiri. Investigations on the performance of distributed Raman amplifier in dense wavelength division multiplexed communication system using different modulation formats. Concurrency and Computation: Practice and Experience, 34(10):e6713:1–e6713:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seetha:2022:PEA

[SA22c] Jayaraman Seetha and Maruthu Ayyadurai. Performance evaluation of accessibility checker tool for educational websites. Concurrency and Computation: Practice and Experience, 34(24):e7237:1–e7237:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sinhal:2022:MTB

[SA22d] Rishi Sinhal and Irshad Ahmad Ansari. A multiple transform based approach for robust and blind image copyright protection. Concurrency and Computation: Practice and Experience, 34(28):e7362:1–e7362:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2022:DDA

[SAAAA22] Sonam Sharma, Izzat Alsmadi, Rami S. Alkhawaldeh, and Bilal Al-Ahmad. Data-driven analysis and predictive modeling on COVID-19. Concurrency and Computation: Practice and Experience, 34(28):e7390:1–e7390:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sami:2022:REC

[SAB22] Faiza Sami, Muhammad Amin, and Muhammad Moeen Butt. On the ridge estimation of the Conway-Maxwell Poisson regression model with multicollinearity: Methods and applications. Concurrency and Computation: Practice and Experience, 34(1):e6477:1–e6477:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sarkar:2022:SLA

[SAC22] Paramita Sarkar, Ayush Aryan, and Rituparna Chaki. Sea lion attacking-based deer hunting optimization algorithm for dynamic nurse scheduling in health care sector contribution of hybrid algorithm in cloud. Concurrency and Computation: Practice and Experience, 34(25):e7249:1–e7249:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shafi:2021:ANN

[SAD^{+}21] Imran Shafi, Sana Ansari, Sadia Din, Gwanggil Jeon, and Anand Paul. Artificial neural networks as clinical decision support systems. Concurrency and Computation: Practice and Experience, 33(22):e6342:1–e6342:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sonohata:2023:PPG

[SAF^{+}23] Rhayssa Sonohata, Danillo Christi A. Arigoni, Eraldo Rezende Fernandes, Ricardo Ribeiro dos Santos, and Liana Dessandre Duenha. Performance predictors for graphics processing units applied to dark-silicon-aware design space exploration. Concurrency and Computation: Practice and Experience, 35(17): e6877:1–e6877:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Syed:2024:LXT

[SAGS24] Farheen Syed, Saeed Hamood Alsamhi, Sachin Kumar Gupta, and Abdu Saif. LSB-XOR technique for securing captured images from disaster by UAVs in B5G networks. Concurrency and Computation: Practice and Experience, 36(12):e8061:1– e8061:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shahzad:2022:ECV

[SAHAN22] Usman Shahzad, Ishfaq Ahmad, Muhammad Hanif, and Nadia H. Al-Noor. Estimation of coefficient of variation using linear moments and calibration approach for nonsensitive and sensitive variables. Concurrency and Computation: Practice and Experience, 34(18):e7006:1–e7006:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Soualmi:2022:NBM

[SAL22a] Abdallah Soualmi, Adel Alti, and Lamri Laouamer. A novel blind medical image watermarking scheme based on Schur triangulation and chaotic sequence. Concurrency and Computation: Practice and Experience, 34(1):e6480:1–e6480:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Souza:2022:WPL

[SAL^{+}22b] Renan Souza, Leonardo G. Azevedo, Vítor Lourenço, Elton Soares, Raphael Thiago, Rafael Brandão, Daniel Civitarese, Emilio Vital Brazil, Marcio Moreno, Patrick Valduriez, Marta Mattoso, Renato Cerqueira, and Marco A. S. Netto. Workflow provenance in the lifecycle of scientific machine learning. Concurrency and Computation: Practice and Experience, 34(14): e6544:1–e6544:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saif:2023:MOC

[SAM^{+}23] Mufeed Ahmed Naji Saif, Sk Niranjan Aradhya, Belal Abdullah Hezam Murshed, Omar Abdullah Murshed Farhan Alnaggar, and Issa Mohammed Saeed Ali. Multi-objective container scheduling and multi-path routing for elastic business process management in autonomic multi-tenant cloud. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Suganthy:2023:RER

[SAMS23] M. Suganthy, S. Ashok, A. Uma Maheswari, and T. D. Subha. Recalling-enhanced recurrent neural network optimized with wood pecker mating algorithm for brain tumor classification. Concurrency and Computation: Practice and Experience, 35 (22):e7729:1–e7729:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saidi:2024:SMA

[SAN24] Ahmed Saidi, Abdelouahab Amira, and Omar Nouali. A secure multi-authority attribute based encryption approach for robust smart grids. Concurrency and Computation: Practice and Experience, 36(7):e7972:1–e7972:??, March 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saoucha:2022:BJB

[Sao22] Naziha Ali Saoucha. Binary Jaya-based channel allocation in cognitive radio networks. Concurrency and Computation: Practice and Experience, 34(26):e7365:1–e7365:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

S:2021:ESS

[SAPC21] Sathish S, Lakshmanan Ar, Karuppuswamy P, and Bhagyanathan C. An effective Sorensen-single linkage clustering hybrid algorithm for cell formation problems in cellular manufacturing industry. Concurrency and Computation: Practice and Experience, 33(3):e5211:1–e5211:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saleem:2023:CAT

[SAQJ23] Zeeshan Saleem, Adi Alhudhaif, Kashif Naseer Qureshi, and Gwanggil Jeon. Context-aware text classification system to improve the quality of text: a detailed investigation and techniques. Concurrency and Computation: Practice and Experience, 35(15):e6489:1–e6489:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Staifi:2021:ASH

[SB21] Noureddine Staifi and Meriem Belguidoum. Adapted smart home services based on smart contracts and service level agreements. Concurrency and Computation: Practice and Experience, 33(23):e6208:1–e6208:??, December 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

S:2023:OMO

[SB23a] Nivethithai S and Hariharan B. Optimized multi-objective Qlearning with enhanced beetle swarm optimization based scientific workflows scheduling on cloud computing environment. Concurrency and Computation: Practice and Experience, 35 (1):e7409:1–e7409:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2023:NEH

[SB23b] Sonia Sharma and Rajendra Kumar Bharti. New efficient Hadoop scheduler: Generalized particle swarm optimization and simulated annealing-dominant resource fairness. Concurrency and Computation: Practice and Experience, 35(4): e7528:1–e7528:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Selcuk:2022:ADE

[SBA22] Turab Selçuk, Abdullah Beyoglu, and Ahmet Alkan. Automatic detection of exudates and hemorrhages in low-contrast color fundus images using multi semantic convolutional neural network. Concurrency and Computation: Practice and Experience, 34(6):e6768:1–e6768:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Souahlia:2020:ESN

[SBB^{+}20] Abdelkerim Souahlia, Ammar Belatreche, Abdelkader Benyettou, Zoubir Ahmed-Foitih, Elhadj Benkhelifa, and Kevin Curran. Echo state network-based feature extraction for efficient color image segmentation. Concurrency and Computation: Practice and Experience, 32(21):e5719:1–e5719:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shahpar:2021:PAB

[SBB21] Zahra Shahpar, Vahid Khatibi Bardsiri, and Amid Khatibi Bardsiri. Polynomial analogy-based software development effort estimation using combined particle swarm optimization and simulated annealing. Concurrency and Computation: Practice and Experience, 33(20):e6358:1–e6358:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Skjellum:2020:FSI

[SBG20] Anthony Skjellum, Purushotham V. Bangalore, and Ryan E. Grant. Foreword to the special issue of the Workshop on Exascale MPI (ExaMPI 2017). Concurrency and Computation: Practice and Experience, 32(3):e5459:1–e5459:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Scott:2021:IDS

[SBGC21] Stacey D. Scott, Guillaume Besacier, Nippun Goyal, and Frank Cento. Investigating device-specific visual feedback for cross-device transfer in table-centric multisurface environments. Concurrency and Computation: Practice and Experience, 33(8):e4084:1–e4084:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Suresh:2022:MLB

[SBS22] T. Suresh, Z. Brijet, and T. D. Subha. Modified local binary patterns based feature extraction and hyper parameters tuned attention segmental recurrent neural network classifier using flamingo search optimization algorithm for disease diagnosis model. Concurrency and Computation: Practice and Experience, 34(23):e7182:1–e7182:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2022:MFO

[SBSK22] Priyanka Singh, Samir Kumar Borgohain, Lakhan Dev Sharma, and Jayendra Kumar. Minimized feature overhead malware detection machine learning model employing MRMR-based ranking. Concurrency and Computation: Practice and Experience, 34(17):e6992:1–e6992:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saleem:2022:LAD

[SC22a] Tausifa Jan Saleem and Mohammad Ahsan Chishti. LPLB: an approach for the design of a lightweight convolutional neural network. Concurrency and Computation: Practice and Experience, 34(22):e7143:1–e7143:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sassi:2022:STB

[SC22b] Salma Sassi and Richard Chbeir. SIMCard: Toward better connected electronic health record visualization. Concurrency and Computation: Practice and Experience, 34(28):e7399:1– e7399:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Serpa:2023:MEU

[SCD^{+}23] Matheus S. Serpa, Eduardo H. M. Cruz, Matthias Diener, Arthur F. Lorenzon, Antonio C. S. Beck, and Philippe O. A. Navaux. Mitigating execution unit contention in parallel applications using instruction-aware mapping. Concurrency and Computation: Practice and Experience, 35(17):e6819:1– e6819:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stramondo:2020:DBA

[SCdLV20] Giulio Stramondo, Catalin Bogdan Ciobanu, Cees de Laat, and Ana Lucia Varbanescu. Designing and building application-centric parallel memories. Concurrency and Computation: Practice and Experience, 32(15):e5485:1–e5485:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Si:2022:SIPa

[SCH22a] Min Si, Quan Chen, and Zhiyi Huang. Special issue on programming models and applications for multicores and manycores 2019-2020. Concurrency and Computation: Practice and Experience, 34(2):e6677:1–e6677:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Si:2022:SIPb

[SCH22b] Min Si, Quan Chen, and Zhiyi Huang. Special issue on programming models and applications for multicores and manycores 2020. Concurrency and Computation: Practice and Experience, 34(2):e6457:1–e6457:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2024:LAL

[SCH^{+}24] Guohua Shen, Mengnan Cai, Zhiqiu Huang, Yang Yang, Feifei Guo, and Linlin Wei. LoHDP: Adaptive local differential privacy for high-dimensional data publishing. Concurrency and Computation: Practice and Experience, 36(11):e8039:1– e8039:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singhal:2023:OPD

[SCK^{+}23] Sharad Singhal, Clarete Riana Crasta, Mashood Abdulla K, Faizan Barmawer, Gautham Bhat, Ramya Ahobala Rao, Soumya P N, and Rishi Kesh K Rajak. OpenFAM: Programming disaggregated memory. Concurrency and Computation: Practice and Experience, 35(28):e7910:1–e7910:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sasongko:2023:PES

[SCKU23] Muhammad Aditya Sasongko, Milind Chabbi, Paul H. J. Kelly, and Didem Unat. Precise event sampling-based data locality tools for AMD multicore architectures. Concurrency and Computation: Practice and Experience, 35(24):e7707:1– e7707:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2020:PPD

[SCL^{+}20] Liping Sun, Shang Ci, Xiaoqing Liu, Xiaoyao Zheng, Qingying Yu, and Yonglong Luo. A privacy-preserving density peak clustering algorithm in cloud computing. Concurrency and Computation: Practice and Experience, 32(11):e5641:1– e5641:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sivalingam:2022:ECL

[SCM22] Nithya Priya Sivalingam, Sundar Chinnasamy, and Thanabal Suruli Muniyandi. An effective chronic lymphocytic leukemia detection method using hybrid optimization aware random multimodal deep learning. Concurrency and Computation: Practice and Experience, 34(18):e7012:1–e7012:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shang:2020:CEC

[SCP20] Qianyi Shang, Lijun Chen, and Peng Peng. On-chip evolution of combinational logic circuits using an improved geneticsimulated annealing algorithm. Concurrency and Computation: Practice and Experience, 32(23):e5486:1–e5486:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Slalmi:2021:NIE

[SCS^{+}21] Ahmed Slalmi, Hasna Chaibi, Rachid Saadane, Abdellah Chehri, and Gwanggil Jeon. 5G NB-IoT: Efficient network call admission control in cellular networks. Concurrency and Computation: Practice and Experience, 33(22):e6047:1– e6047:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Skorych:2022:PCG

[SD22] Vasyl Skorych and Maksym Dosta. Parallel CPU-GPU computing technique for discrete element method. Concurrency and Computation: Practice and Experience, 34(11):e6839:1– e6839:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Srinivasan:2023:IMD

[SD23] Sathiyandrakumar Srinivasan and P. Deepalakshmi. An innovative malware detection methodology employing the amalgamation of stacked BiLSTM and CNN+LSTM-based classification networks with the assistance of Mayfly metaheuristic optimization algorithm in cyber-attack. Concurrency and Computation: Practice and Experience, 35(10):e7679:1–e7679:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Swami:2020:VBI

[SDR20] Rochak Swami, Mayank Dave, and Virender Ranga. Votingbased intrusion detection framework for securing softwaredefined networks. Concurrency and Computation: Practice and Experience, 32(24):e5927:1–e5927:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sampradeepraj:2023:SMW

[SDR23] T Sampradeepraj, V Anusuya Devi, and S. P. Raja. Secure multicasting in wireless sensor networks using identity based cryptography. Concurrency and Computation: Practice and Experience, 35(1):e7430:1–e7430:??, January 10, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2021:CDE

[SDSW21] Chengzhuo Shi, Zheng Dou, Arun Kumar Sangaiah, and Jin Wang. Cognitive decision engine based on binary particles swarm optimization with non-linear decreasing inertia weight. Concurrency and Computation: Practice and Experience, 33 (12):e4975:1–e4975:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sekar:2022:ECS

[SEC22] Jai Ganesh Sekar, Periyathambi Ezhumalai, and Arun Chokkalingam. An efficient chaotic system based hybrid radiation heat transfer sunflower optimization algorithm for securing digital image transmission. Concurrency and Computation: Practice and Experience, 34(10):e6814:1–e6814:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sha:2020:MVM

[SEM^{+}20] Jing Sha, Abdol Ghaffar Ebadi, Dinesh Mavaluru, Mohmmed Alshehri, Osama Alfarraj, and Lila Rajabion. A method for virtual machine migration in cloud computing using a collective behavior-based metaheuristics algorithm. Concurrency and Computation: Practice and Experience, 32(2):e5441:1– e5441:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saad:2022:WLS

[SEMA^{+}22] Mohamed Saad, Ali El-Moursy, Oruba Alfawaz, Khawla Alnajjar, and Saeed Abdallah. Wireless link scheduling via parallel genetic algorithm. Concurrency and Computation: Practice and Experience, 34(6):e6783:1–e6783:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shoaib:2022:EDL

[SET^{+}22] Mohamed R. Shoaib, Mohamed R. Elshamy, Taha E. Taha, Adel S. El-Fishawy, and Fathi E. Abd El-Samie. Efficient deep learning models for brain tumor detection with segmentation and data augmentation techniques. Concurrency and Computation: Practice and Experience, 34(21):e7031:1–e7031:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2021:RMU

[SFJ^{+}21] Xiyan Sun, Zhuo Fan, Yuanfa Ji, Shouhua Wang, Suqing Yan, Sunyong Wu, Qiang Fu, and Kamarul Hawari Ghazali. Robust multi-user detection based on hybrid grey wolf optimization. Concurrency and Computation: Practice and Experience, 33(15):e5273:1–e5273:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Settipalli:2021:HFD

[SG21] Lavanya Settipalli and G. R. Gangadharan. Healthcare fraud detection using primitive sub peer group analysis. Concurrency and Computation: Practice and Experience, 33(23): e6275:1–e6275:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Samanta:2022:FFE

[SG22a] Riya Samanta and Soumya K. Ghosh. FogiRecruiter: a fogenabled selection mechanism of crowdsourcing for disaster management. Concurrency and Computation: Practice and Experience, 34(23):e7207:1–e7207:??, October 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Senturk:2022:FIS

[SG22b] Fatmana Sentürk and Gurhan Gunduz. A framework for investigating search engines’ stemming mechanisms: a case study on Bing. Concurrency and Computation: Practice and Experience, 34(9):e6562:1–e6562:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2022:ABD

[SG22c] Preeti Sharma and Mandlem Gangadharappa. Abnormal behavior detection of stationary objects in surveillance videos with visualization and classification. Concurrency and Computation: Practice and Experience, 34(21):e7056:1–e7056:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sivasakthi:2022:QAM

[SG22d] Dharmalingam Adhimuga Sivasakthi and Raja Gunasekaran. QoE-aware mobile computation offloading in mobile edge computing. Concurrency and Computation: Practice and Experience, 34(11):e6853:1–e6853:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Srinath:2022:EDD

[SG22e] Rajagopalan Srinath and Rajagopal Gayathri. Epilepsy disorder detection and diagnosis using empirical mode decomposition and deep learning architecture. Concurrency and Computation: Practice and Experience, 34(11):e6903:1–e6903:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shekhawat:2024:SLB

[SG24] Hema Shekhawat and Daya Sagar Gupta. A survey on latticebased security and authentication schemes for smart-grid networks in the post-quantum era. Concurrency and Computation: Practice and Experience, 36(14):e8080:1–e8080:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shumeng:2023:SSM

[SGH23] He Shumeng, Xu Gaodi, and Yang Houqun. A semantic segmentation method for remote sensing images based on multiple contextual feature extraction. Concurrency and Computation: Practice and Experience, 35(2):e7483:1–e7483:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stepanovic:2020:EMQ

[SGHL20] Srboljub Stepanovic, Georgios Georgakarakos, Simon Holmbacka, and Johan Lilius. An efficient model for quantifying the interaction between structural properties of software and hardware in the ARM big.LITTLE architecture. Concurrency and Computation: Practice and Experience, 32(10):e5230:1– e5230:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sagar:2021:MVA

[SGS21a] Kalpna Sagar, Deepak Gupta, and Arun K. Sangaiah. Manual versus automated qualitative usability assessment of interactive systems. Concurrency and Computation: Practice and Experience, 33(12):e5091:1–e5091:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sathyasri:2021:EPT

[SGS21b] B. Sathyasri, E. N. Ganesh, and P. Senthilkumar. Enhance packet transmission using improved channel assignment in wireless mess network. Concurrency and Computation: Practice and Experience, 33(3):e5357:1–e5357:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2023:APS

[SGS^{+}23] Upendra Singh, Puja Gupta, Mukul Shukla, Varsha Sharma, Sunita Varma, and Sumit Kumar Sharma. Acknowledgment of patient in sense behaviors using bidirectional ConvLSTM. Concurrency and Computation: Practice and Experience, 35(28):e7819:1–e7819:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shankhpal:2022:KBQ

[SH22] Shilpa V. Shankhpal and Brahmananda Savadatti Hanumantha. KMFA2 based QoS improvement for multi-channel IoT networks. Concurrency and Computation: Practice and Experience, 34(15):e6949:1–e6949:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sathiyamoorthi:2022:EMP

[SHA^{+}22] Velayutham Sathiyamoorthi, Pon Harshavardhanan, Hussain Azath, Marimuthu Senbagavalli, Ayyappillai Mallika Viswa Bharathy, and Bala Subramanian Chokkalingam. An effective model for predicting agricultural crop yield on remote sensing hyper-spectral images using adaptive logistic regression classifier. Concurrency and Computation: Practice and Experience, 34(25):e7242:1–e7242:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Semmoud:2020:LBC

[SHBC20] Abderraziq Semmoud, Mourad Hakem, Badr Benmammar, and Jean-Claude Charr. Load balancing in cloud computing environments based on adaptive starvation threshold. Concurrency and Computation: Practice and Experience, 32(11): e5652:1–e5652:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shirazi:2022:SPI

[Shi22a] Babak Shirazi. Super-process interoperability optimization architecture in healthcare ultra-large-scale systems: a graphbased multi-objective approach. Concurrency and Computation: Practice and Experience, 34(3):e6595:1–e6595:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shivani:2022:EFB

[Shi22b] Shivendra Shivani. An efficient front-to-back depth-buffer algorithm for real-time rendering of partially occluded game objects in absence of GPU. Concurrency and Computation: Practice and Experience, 34(15):e6962:1–e6962:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Samadiani:2022:MFF

[SHL^{+}22] Najmeh Samadiani, Guangyan Huang, Wei Luo, Chi-Hung Chi, Yanfeng Shu, Rui Wang, and Tuba Kocaturk. A multiple feature fusion framework for video emotion recognition in the wild. Concurrency and Computation: Practice and Experience, 34(8):e5764:1–e5764:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shukla:2022:CTL

[Shu22] Alok Kumar Shukla. Chaos teaching learning based algorithm for large-scale global optimization problem and its application. Concurrency and Computation: Practice and Experience, 34(1):e6514:1–e6514:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sree:2024:FSE

[SHV24] Thankaraja Raja Sree, R. Harish, and T. Veni. FogSec: a secure and effective mutual authentication scheme for fog computing. Concurrency and Computation: Practice and Experience, 36(12):e8054:1–e8054:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stanislaus:2023:NFB

[SHVA23] Oswalt Manoj Stanislaus, Pon Harshavardhanan, Akila Victor, and Sajeev Ram Arumugam. A novel fuzzy based deep neural network for rain fall prediction using cloud images. Concurrency and Computation: Practice and Experience, 35 (1):e7412:1–e7412:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2020:AEH

[SHZY20] Peiying Shi, Feng Hou, Xiangwei Zheng, and Feng Yuan. Analysis of electronic health records based on long short-term memory. Concurrency and Computation: Practice and Experience, 32(14):e5684:1–e5684:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sowa:2022:DOM

[SI22] Piotr Sowa and Jacek Izydorczyk. Darknet on OpenCL: a multiplatform tool for object detection and classification. Concurrency and Computation: Practice and Experience, 34 (15):e6936:1–e6936:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2023:OCA

[Sin23] Amit Singh. Opportunistic coding across on-route dominating nodes in a flow-oriented SDN enabled WSN. Concurrency and Computation: Practice and Experience, 35(10):e7659:1– e7659:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Safavi:2022:DHA

[SJ22] Sadaf Safavi and Mehrdad Jalali. DeePOF: a hybrid approach of deep convolutional neural network and friendship to Pointof-Interest (POI) recommendation system in location-based social networks. Concurrency and Computation: Practice and Experience, 34(15):e6981:1–e6981:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2023:ONP

[SJ23a] Kishan Kumar Singh and Vijay Kumar Jha. An optimized novel public cloud system to secure the medical record from third parties. Concurrency and Computation: Practice and Experience, 35(19):e6926:1–e6926:??, August 30, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sudha:2023:DFR

[SJ23b] K. Swapna Sudha and N. Jeyanthi. Dynamic forest of random subsets-based supervised hash signature scheme for secure user authentication in smart home environment. Concurrency and Computation: Practice and Experience, 35(27): e6302:1–e6302:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saeedvand:2022:DLT

[SJA^{+}22] Saeed Saeedvand, Masoumeh Jafari, Hadi S. Aghdasi, Jacky Baltes, and Amir Masoud Rahmani. Deep learning: a taxonomy of modern weapons to combat Covid-19 similar pandemics in smart cities. Concurrency and Computation: Practice and Experience, 34(27):e7314:1–e7314:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Savitha:2020:ADM

[SK20a] G. Savitha and S. Karthikeyan. Architecture for diverse modulation techniques in SDR. Concurrency and Computation: Practice and Experience, 32(4):e5129:1–e5129:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stavrinides:2020:SRT

[SK20b] Georgios L. Stavrinides and Helen D. Karatza. Scheduling real-time bag-of-tasks applications with approximate computations in SaaS clouds. Concurrency and Computation: Practice and Experience, 32(1):e4208:1–e4208:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Savarimuthu:2021:UNN

[SK21a] Nickolas Savarimuthu and Shobha Karesiddaiah. An unsupervised neural network approach for imputation of missing values in univariate time series data. Concurrency and Computation: Practice and Experience, 33(9):e6156:1–e6156:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stavrinides:2021:CAC

[SK21b] Georgios L. Stavrinides and Helen D. Karatza. Cost-aware cloud bursting in a fog-cloud environment with real-time workflow applications. Concurrency and Computation: Practice and Experience, 33(23):e5850:1–e5850:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sundharamurthy:2021:CBO

[SK21c] Gnanamurthy Sundharamurthy and Vishnu Kumar Kaliappan. Cloud-based onboard prediction and diagnosis of diabetic retinopathy. Concurrency and Computation: Practice and Experience, 33(24):e6444:1–e6444:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Santhalakshmi:2022:SMD

[SK22a] M. Santhalakshmi and P. Kavitha. Sensitive medical data transmission and maintaining data quality using bacterial bee swarm-based hybrid lifetime maximization large-scale ad hoc routing protocol. Concurrency and Computation: Practice and Experience, 34(7):e5509:1–e5509:??, March 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shanmugapriya:2022:CSD

[SK22b] Rathinam Shanmugapriya and Sripathi Venkata Naga Santhosh Kumar. Comprehensive survey on data dissemination protocols for efficient reprogramming in Internet of Things. Concurrency and Computation: Practice and Experience, 34(26):e7280:1–e7280:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sinha:2022:EPM

[SK22c] Raghaw Raman Sinha and Bharti Khanna. Estimation of population mean under probability proportional to size sampling with and without measurement errors. Concurrency and Computation: Practice and Experience, 34(18):e7023:1– e7023:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sonya:2022:EBB

[SK22d] Ansar Sonya and Ganesh Kavitha. An effective blockchainbased smart contract system for securing electronic medical data in smart healthcare application. Concurrency and Computation: Practice and Experience, 34(28):e7363:1–e7363:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sahana:2023:CRS

[SK23] Devarayapattana Siddalingaiah Sahana and Chandrasekaran Gopala Krishnan. Certificate revocation schemes for providing secure QoS in MANETs. Concurrency and Computation: Practice and Experience, 35(4):e7524:1–e7524:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saidi:2023:IRT

[SKA23] Taki Eddine Saidi, Abdelhakim Khouas, and Abbes Amira. Implementation of a real-time stereo vision algorithm on a cost-effective heterogeneous multicore platform. Concurrency and Computation: Practice and Experience, 35(1):e7411:1– e7411:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Samuel:2020:ARR

[SKB^{+}20] David Samuel, Syeduzzaman Khan, Cody J. Balos, Zachariah Abuelhaj, Anthony D. Dutoi, Chadi Kari, David Mueller, and Vivek K. Pallipuram. A2Cloud-RF: a random forest based statistical framework to guide resource selection for high-performance scientific computing on the cloud. Concurrency and Computation: Practice and Experience, 32(24): e5942:1–e5942:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sunnetci:2023:CPG

[SKÇA23] Kubilay Muhammed Sunnetci, Esat Kaba, Fatma Beyazal Çeliker, and Ahmet Alkan. Comparative parotid gland segmentation by using ResNet-18 and MobileNetV2 based DeepLab v3+ architectures from magnetic resonance images. Concurrency and Computation: Practice and Experience, 35 (1):e7405:1–e7405:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2023:CED

[SKCS23] Akshita Sharma, Baijnath Kaushik, Akshma Chadha, and Reya Sharma. Comparative evaluation of deep dense sequential and deep dense transfer learning models for suicidal emotion prediction. Concurrency and Computation: Practice and Experience, 35(22):e7745:1–e7745:??, October 10, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Solak:2022:SII

[SKE22] Serdar Solak, Zeynep Hilal Kilimci, and Süleyman Eken. Special issue on international conference on innovations in intelligent systems and applications (INISTA 2021). Concurrency and Computation: Practice and Experience, 34(20):e6948:1– e6948:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sarathchandra:2021:RAS

[SKH^{+}21] Madushi Sarathchandra, Chulani Karandana, Winma Heenatigala, Miyuru Dayarathna, and Sanath Jayasena. Resource aware scheduler for distributed stream processing in cloud native environments. Concurrency and Computation: Practice and Experience, 33(20):e6373:1–e6373:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2022:CIH

[SKHL22] Gaurav Sharma, Savita Khurana, Shilpi Harnal, and Showkat Ahmad Lone. CSFPA: an intelligent hybrid workflow scheduling algorithm based upon global and local optimization approach in cloud. Concurrency and Computation: Practice and Experience, 34(23):e7176:1–e7176:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2021:SAD

[SKK21] Shweta Sharma, Rakesh Kumar, and C. Rama Krishna. A survey on analysis and detection of Android ransomware. Concurrency and Computation: Practice and Experience, 33 (16):e6272:1–e6272:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2022:EVA

[SKK22] Avtar Singh, Navjot Kaur, and Harpreet Kaur. An extensive vulnerability assessment and countermeasures in open network operating system software defined networking controller. Concurrency and Computation: Practice and Experience, 34(15):e6978:1–e6978:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Satti:2022:RIR

[SKP22] Satish Kumar Satti, Suganya Devi K, and Srinivasan P. R-ICTS: Recognize the Indian cautionary traffic signs in realtime using an optimized adaptive boosting cascade classifier and a convolutional neural network. Concurrency and Computation: Practice and Experience, 34(10):e6796:1–e6796:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Satti:2023:RIC

[SKP23] Satish Kumar Satti, Suganya Devi K, and Srinivasan P. Recognizing the Indian cautionary traffic signs using GAN, improved mask R-CNN, and Grab Cut. Concurrency and Computation: Practice and Experience, 35(2):e7453:1–e7453:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saini:2023:HEM

[SKPD23] Neeraj Saini, Vivekananda Bhat Kasaragod, Krishna Prakasha, and Ashok Kumar Das. A hybrid ensemble machine learning model for detecting APT attacks based on network behavior anomaly detection. Concurrency and Computation: Practice and Experience, 35(28):e7865:1–e7865:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saravanan:2022:RBR

[SKPV22] Mohan Saravanan, Rajakani Kalidoss, Bactavatchalame Partibane, and Kuttathati Srinivasan Vishvaksenan. Reservation based resource allocation in 5G new radio standard. Concurrency and Computation: Practice and Experience, 34(8): e5496:1–e5496:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Suseela:2024:EAO

[SKR^{+}24] Sellamuthu Suseela, Ravi Krithiga, Muthusamy Revathi, Gajendran Sudhakaran, and Reddiyapalayam Murugeshan Bhavadharini. Energy aware optimal routing model for wireless multimedia sensor networks using modified Voronoi assisted prioritized double deep Q-learning. Concurrency and Computation: Practice and Experience, 36(6):e7945:1– e7945:??, March 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stefanic:2021:QSA

[SKRS21] Polona Stefanic, Petar Kochovski, Omer F. Rana, and Vlado Stankovski. Quality of service-aware matchmaking for adaptive microservice-based applications. Concurrency and Computation: Practice and Experience, 33(19):e6120:1–e6120:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2020:ECR

[SKS20] Simar Preet Singh, Rajesh Kumar, and Anju Sharma. Efficient content retrieval in fog zone using Nano-Caches. Concurrency and Computation: Practice and Experience, 32(2): e5438:1–e5438:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Soulami:2022:BCT

[SKS22] Khaoula Belhaj Soulami, Naima Kaabouch, and Mohamed Nabil Saidi. Breast cancer: Three-class masses classification in mammograms using apriori dynamic selection. Concurrency and Computation: Practice and Experience, 34(24): e7233:1–e7233:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2023:FBS

[SKS^{+}23] Om Prakash Singh, Chandan Kumar, Amit Kumar Singh, Maheshwari Prasad Singh, and Hoon Ko. Fuzzy-based secure exchange of digital data using watermarking in NSCT-RDWT-SVD domain. Concurrency and Computation: Practice and Experience, 35(16):e6251:1–e6251:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2020:ISM

[SKSB20] Shashank Sheshar Singh, Ajay Kumar, Kuldeep Singh, and Bhaskar Biswas. IM-SSO: Maximizing influence in social networks using social spider optimization. Concurrency and Computation: Practice and Experience, 32(2):e5421:1– e5421:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2022:LEE

[SKSN22] Simar Preet Singh, Rajesh Kumar, Anju Sharma, and Anand Nayyar. Leveraging energy-efficient load balancing algorithms in fog computing. Concurrency and Computation: Practice and Experience, 34(13):e5913:1–e5913:??, June 10, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Snasel:2020:JSP

[SKSP20] Vaclav Snasel, Pavel Kromer, Jakub Safarik, and Jan Platos. JPEG steganography with particle swarm optimization accelerated by AVX. Concurrency and Computation: Practice and Experience, 32(8):e5448:1–e5448:??, April 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Srivel:2023:AQE

[SKSP23] R. Srivel, K. Kalaiselvi, S. Shanthi, and Uma Perumal. An automation query expansion strategy for information retrieval by using fuzzy based grasshopper optimization algorithm on medical datasets. Concurrency and Computation: Practice and Experience, 35(3):e7418:1–e7418:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2022:IBF

[SKV22] Jaya Singh, Aditya Kumawat, and Subramanian Venkatesan. Improved Byzantine fault tolerance with fast consensus. Concurrency and Computation: Practice and Experience, 34(10): e6813:1–e6813:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sekar:2020:LRS

[SL20] S. Sekar and B. Latha. Lightweight reliable and secure multicasting routing protocol based on cross-layer for MANET. Concurrency and Computation: Practice and Experience, 32 (4):e5025:1–e5025:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shirinbab:2020:PEC

[SLC20] Sogand Shirinbab, Lars Lundberg, and Emiliano Casalicchio. Performance evaluation of containers and virtual machines when running Cassandra workload concurrently. Concurrency and Computation: Practice and Experience, 32(17):e5693:1– e5693:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2024:EDP

[SLD24] Zhijie Shen, Bowen Liu, and Wanchun Dou. An effective data placement strategy for IIoT applications. Concurrency and Computation: Practice and Experience, 36(10):e7977:1– e7977:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shankar:2020:AOM

[SLG^{+}20] K. Shankar, S. K. Lakshmanaprabu, Deepak Gupta, Ashish Khanna, and Victor Hugo C. de Albuquerque. Adaptive optimal multi key based encryption for digital image security. Concurrency and Computation: Practice and Experience, 32 (4):e5122:1–e5122:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2020:ISB

[SLHW20] Zhanquan Sun, Feng Li, Huifen Huang, and Jian Wang. Image steganalysis based on convolutional neural network and feature selection. Concurrency and Computation: Practice and Experience, 32(5):e5469:1–e5469:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shan:2023:SAC

[SLJ23] Fangfang Shan, Fuyang Li, and Peiyu Ji. A smart access control mechanism based on user preference in online social networks. Concurrency and Computation: Practice and Experience, 35(20):e6864:1–e6864:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Su:2022:CRA

[SLL22] Jiawei Su, Zhiming Luo, and Shaozi Li. Consistent response for automated multilabel thoracic disease classification. Concurrency and Computation: Practice and Experience, 34(23): e7201:1–e7201:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2023:EHB

[SLL^{+}23] Ling Song, Hongxin Liu, Shunming Lyu, Xiaofei Niu, Tie Hou, Yuling Ma, and Qiang Lyu. Exploring human behavior patterns and socio-demographic factors based on American Time Use Survey. Concurrency and Computation: Practice and Experience, 35(20):e6878:1–e6878:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sanchez-Lara:2024:IET

[SLLMTS^{+}24] Rafael Sánchez-Lara, José L. López-Martínez, Joel A. Trejo-Sánchez, Herman L. Offerhaus, and José A. Álvarez-Chávez. Implementation and evaluation of two parallel computational models for the simulation of a long-haul DWDM system limited by FWM. Concurrency and Computation: Practice and Experience, 36(7):e7964:1–e7964:??, March 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Srinivasan:2021:WBG

[SM21] K. Srinivasan and R. Murugasan. Web-based GIS for live GV of land generation through utilitarian association rule mining and multiple regression analysis for chennai district. Concurrency and Computation: Practice and Experience, 33 (3):e5154:1–e5154:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sankar:2022:DWB

[SM22a] K. Sankar and P. Uma Maheswari. A dynamic wrapper-based feature selection for improved precision in content-based image retrieval. Concurrency and Computation: Practice and Experience, 34(8):e5368:1–e5368:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shah:2022:ESR

[SM22b] Vandana Shah and Meet Mehta. Emotional state recognition from text data using machine learning and deep learning algorithm. Concurrency and Computation: Practice and Experience, 34(17):e7001:1–e7001:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shingade:2022:HDQ

[SM22c] Sachin Dattatraya Shingade and Rohini Prashant Mudhalwadkar. Hybrid deep-Q Elman neural network for crop prediction and recommendation based on environmental changes. Concurrency and Computation: Practice and Experience, 34 (17):e6991:1–e6991:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sujaudeen:2022:TTA

[SM22d] N. Sujaudeen and T. T. Mirnalinee. TARNN: Task-aware autonomic resource management using neural networks in cloud environment. Concurrency and Computation: Practice and Experience, 34(8):e5463:1–e5463:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sumathi:2022:MLB

[SM22e] A. Sumathi and S. Meganathan. Machine learning based pattern detection technique for diabetes mellitus prediction. Concurrency and Computation: Practice and Experience, 34(6): e6751:1–e6751:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sheeba:2023:EFT

[SM23a] Adlin Sheeba and B. Uma Maheswari. An efficient fault tolerance scheme based enhanced firefly optimization for virtual machine placement in cloud computing. Concurrency and Computation: Practice and Experience, 35(7):e7610:1– e7610:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shingade:2023:HEL

[SM23b] Sachin Dattatraya Shingade and Rohini Prashant Mudhalwadkar. Hybrid extreme learning machine based bidirectional long short-term memory for crop prediction. Concurrency and Computation: Practice and Experience, 35(2):e7482:1– e7482:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shingade:2023:SIB

[SM23c] Sachin Dattatraya Shingade and Rohini Prashant Mudhalwadkar. Sensor information-based crop recommendation system using machine learning for the fertile regions of Maharashtra. Concurrency and Computation: Practice and Experience, 35(23):e7774:1–e7774:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Swaminathan:2023:MOB

[SM23d] Anitha Swaminathan and Vanitha Muthuraman. Metaheuristic optimization based feature subset selection with deep belief network for stress prediction in working employees. Concurrency and Computation: Practice and Experience, 35(1): e7431:1–e7431:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sakthivel:2024:ASC

[SM24a] G. Sakthivel and P. Madhubala. Advanced set containment deep learned Rabin certificateless signcryption for secured transmission with big data in cloud. Concurrency and Computation: Practice and Experience, 36(1):e7883:1–e7883:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Siyad:2024:PMB

[SM24b] B. Mohammed Siyad and Ramasundaram Mohan. Processingin-memory based multilateration localization in wireless sensor networks using memristor crossbar arrays. Concurrency and Computation: Practice and Experience, 36(12):e8047:1– e8047:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Srivastava:2024:FBM

[SM24c] Ankita Srivastava and Pramod Kumar Mishra. Fuzzy based multi-criteria based cluster head selection for enhancing network lifetime and efficient energy consumption. Concurrency and Computation: Practice and Experience, 36(4):e7921:1– e7921:??, February 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sonmezer:2022:MML

[SMAG22] Hasan Berk Sonmezer, Nitel Muhtaroglu, Ismail Ari, and Deniz Gokcin. MaLeFICE: Machine learning support for continuous performance improvement in computational engineering. Concurrency and Computation: Practice and Experience, 34(9):e6674:1–e6674:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Selvi:2024:AGS

[SMBM24] M. Chengathir Selvi, J. Maruthupandi, T. Bhuvaneswari, and T. Manonmani. An adaptive gravitational search optimization (AGSO)-based convolutional neural network-long shortterm memory (CNN-LSTM) approach for face recognition and classification. Concurrency and Computation: Practice and Experience, 36(3):e7916:1–e7916:??, February 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sai:2022:AHO

[SMCM^{+}22] Ryuichi Sai, John Mellor-Crummey, Xiaozhu Meng, Keren Zhou, Mauricio Araya-Polo, and Jie Meng. Accelerating high-order stencils on GPUs. Concurrency and Computation: Practice and Experience, 34(20):e6467:1–e6467:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shafi:2021:CIN

[SMD^{+}21] Imran Shafi, Zeeshan Malik, Sadia Din, Gwanggil Jeon, and Jamil Ahmad. A computationally intelligent neural networkbased nonlinear autoregressive exogenous balancing approach for real-time processing in industrial applications using big data. Concurrency and Computation: Practice and Experience, 33(22):e6382:1–e6382:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sanyal:2022:TSR

[SMD22a] Banhi Sanyal, Ramesh Kumar Mohapatra, and Ratnakar Dash. Traffic sign recognition on Indian database using wavelet descriptors and convolutional neural network ensemble. Concurrency and Computation: Practice and Experience, 34(10):e6827:1–e6827:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sarkar:2022:SAC

[SMD22b] Samriddhi Sarkar, Mariana Curado Malta, and Animesh Dutta. A survey on applications of coalition formation in multi-agent systems. Concurrency and Computation: Practice and Experience, 34(11):e6876:1–e6876:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sinthia:2022:ILC

[SMKA22] P. Sinthia, M. Malathi, Anitha K., and M. Suresh Anand. Improving lung cancer detection using faster region-based convolutional neural network aided with fuzzy butterfly optimization algorithm. Concurrency and Computation: Practice and Experience, 34(27):e7251:1–e7251:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sailaja:2022:HHT

[SMM22] C Sailaja, Suman Maloji, and Kasiprasad Mannepalli. A hybrid HXPLS-TMFCC parameterization and DCNN-SFO clustering based speaker diarization system. Concurrency and Computation: Practice and Experience, 34(15):e6954:1– e6954:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seshadri:2021:SPA

[SMR21] Karthick Seshadri, Aswin Maruthappan, and Mukunthapriya Sundar Raman. A scalable parallel algorithm for building web directories. Concurrency and Computation: Practice and Experience, 33(9):e6121:1–e6121:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2023:STL

[SMR23] Huifeng Shi, Kai Miao, and Xiaochen Ren. Short-term load forecasting based on CNN-BiLSTM with Bayesian optimization and attention mechanism. Concurrency and Computation: Practice and Experience, 35(17):e6676:1–e6676:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seyedi:2022:DML

[SN22a] Saeid Seyedi and Nima Jafari Navimipour. Designing a multi-layer full-adder using a new three-input majority gate based on quantum computing. Concurrency and Computation: Practice and Experience, 34(4):e6653:1–e6653:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shanmugam:2022:SBT

[SN22b] Karthiyayini Shanmugam and Karthikeyan Nagarajan. Security-based teleradiology in DICOM used two-level DWT based optimization watermarking. Concurrency and Computation: Practice and Experience, 34(15):e6963:1–e6963:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Santos-Neto:2024:DAD

[SNBAI24] Marcos J. Santos-Neto, Jacir L. Bordim, Eduardo A. P. Alchieri, and Edison Ishikawa. DDoS attack detection in SDN: Enhancing entropy-based detection with machine learning. Concurrency and Computation: Practice and Experience, 36(11):e8021:1–e8021:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saito:2021:PTB

[SNET21] Takumi Saito, Shigenari Nakamura, Tomoya Enokido, and Makoto Takizawa. Probability and topic-based data transmission protocol. Concurrency and Computation: Practice and Experience, 33(19):e6372:1–e6372:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Siakavaras:2021:RHG

[SNGK21] Dimitrios Siakavaras, Konstantinos Nikas, Georgios Goumas, and Nectarios Koziris. RCU-HTM: a generic synchronization technique for highly efficient concurrent search trees. Concurrency and Computation: Practice and Experience, 33(10): e6174:1–e6174:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saritas:2022:CCM

[SO22] Omer Faruk Saritas and Serkan Ozturk. A color channel multiplexing approach for robust discrete wavelet transform based image watermarking. Concurrency and Computation: Practice and Experience, 34(25):e7255:1–e7255:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Samir:2021:ARA

[SP21a] Areeg Samir and Claus Pahl. Autoscaling recovery actions for container-based clusters. Concurrency and Computation: Practice and Experience, 33(23):e5955:1–e5955:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2021:HDD

[SP21b] Yujia Sun and Jan Platos. High-dimensional data classification model based on random projection and bagging-support vector machine. Concurrency and Computation: Practice and Experience, 33(9):e6095:1–e6095:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2022:AFB

[SP22a] Leyan Shi and Somayeh Pouramini. Adaptive façade for building energy efficiency improvement by arithmetical optimization algorithm. Concurrency and Computation: Practice and Experience, 34(22):e7152:1–e7152:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shokouhifar:2022:CAN

[SP22b] Mohammad Shokouhifar and Nazanin Pilevari. Combined adaptive neuro-fuzzy inference system and genetic algorithm for e-learning resilience assessment during COVID-19 pandemic. Concurrency and Computation: Practice and Experience, 34(10):e6791:1–e6791:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sivaraj:2022:DNN

[SP22c] Aparna Sivaraj and Valarmathie Palanisamy. Deep neural network based interactive fuzzy Bayesian search algorithm for low-cost smart farming automation model. Concurrency and Computation: Practice and Experience, 34(27):e7296:1– e7296:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Suchithra:2022:LRB

[SP22d] Marathezhath Subramanian Suchithra and Maya L. Pai. Label ranking-based recommendation system to rank crops for agroecological units. Concurrency and Computation: Practice and Experience, 34(5):e6695:1–e6695:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sadhukhan:2023:OMC

[SP23a] Payel Sadhukhan and Sarbani Palt. Oversampling the minority class using a dedicated fitness function and genetic algorithmic progression. Concurrency and Computation: Practice and Experience, 35(18):e6648:1–e6648:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Srivastava:2023:CHA

[SP23b] Abhishek Srivastava and Rajeev Paulus. Coverage hole aware optimal cluster based routing for wireless sensor network assisted IoT using hybrid deep recurrent neural network. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saeed:2021:IIR

[SPA^{+}21] Faisal Saeed, Anand Paul, Muhammad Jamal Ahmed, Malik Junaid Jami Gul, Won-Hwa Hong, and Hyuncheol Seo. Intelligent implementation of residential demand response using multiagent system and deep neural networks. Concurrency and Computation: Practice and Experience, 33(22):e6168:1– e6168:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Serpa:2021:EEP

[SPC^{+}21] Matheus S. Serpa, Pablo J. Pavan, Eduardo H. M. Cruz, Rodrigo L. Machado, Jairo Panetta, Antônio Azambuja, Alexandre S. Carissimi, and Philippe O. A. Navaux. Energy efficiency and portability of oil and gas simulations on multicore and graphics processing unit architectures. Concurrency and Computation: Practice and Experience, 33(18):e6212:1– e6212:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2021:SID

[SPHP21] Lan Song, Dennis K. Peters, Weimin Huang, and Desmond Power. Ship–iceberg discrimination from Sentinel-1 synthetic aperture radar data using parallel convolutional neural network. Concurrency and Computation: Practice and Experience, 33(17):e6297:1–e6297:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sperhac:2021:GBG

[SPJI^{+}21] Jeanette M. Sperhac, Kristin Poinar, Renette Jones-Ivey, Jason Briner, Beata Csatho, Sophie Nowicki, Erika Simon, Eric Larour, Justin Quinn, and Abani Patra. GHub: Building a glaciology gateway to unify a community. Concurrency and Computation: Practice and Experience, 33(19):e6130:1– e6130:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Subramani:2022:CEA

[SPK^{+}22a] Neelakandan Subramani, Santhosh Kumar Perumal, Jagadish Shivappa Kallimani, Sakthi Ulaganathan, Sanjay Bhargava, and Sangeetha Meckanizi. Controlling energy aware clustering and multihop routing protocol for IoT assisted wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(21):e7106:1–e7106:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sugitha:2022:IDF

[SPK22b] G. Sugitha, B. C. Preethi, and G. Kavitha. Intrusion detection framework using stacked auto encoder based deep neural network in IOT network. Concurrency and Computation: Practice and Experience, 34(28):e7401:1–e7401:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2022:PAF

[SPKK22] Astha Singh, Shiv Prakash, Ankit Kumar, and Divya Kumar. A proficient approach for face detection and recognition using machine learning and high-performance computing. Concurrency and Computation: Practice and Experience, 34 (3):e6582:1–e6582:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sangaiah:2020:CDM

[SPQM20] Arun Kumar Sangaiah, Hoang Pham, Tie Qiu, and Khan Muhammad. Convergence of deep machine learning and parallel computing environment for bio-engineering applications. Concurrency and Computation: Practice and Experience, 32 (1):e5424:1–e5424:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2022:ORR

[SPS22] Ankita Singh, Shiv Prakash, and Sudhakar Singh. Optimization of reinforcement routing for wireless mesh network using machine learning and high-performance computing. Concurrency and Computation: Practice and Experience, 34(15): e6960:1–e6960:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saravanan:2023:TAA

[SPS23] S. Saravanan, D. Prabakar, and S. S. Sathya. Trust aware ad hoc routing protocol with key management based mechanism and optimal energy-efficient cluster head selection in mobile ad hoc networks. Concurrency and Computation: Practice and Experience, 35(7):e7599:1–e7599:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Senthilkumar:2023:DRF

[SPSP23] S. Senthilkumar, V. Samuthira Pandi, T. Sripriya, and N. Pragadish. Design of recustomize finite impulse response filter using truncation based scalable rounding approximate multiplier and error reduced carry prediction approximate adder for image processing application. Concurrency and Computation: Practice and Experience, 35(8):e7629:1– e7629:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2023:MLP

[SPWX23] Yuanyuan Shen, Manman Peng, Qiang Wu, and Guoqi Xie. Multigraph learning for parallelism discovery in sequential programs. Concurrency and Computation: Practice and Experience, 35(9):e7648:1–e7648:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sivakumar:2020:QAE

[SR20a] V. Sivakumar and D. Rekha. A QoS-aware energy-efficient memetic flower pollination routing protocol for underwater acoustic sensor network. Concurrency and Computation: Practice and Experience, 32(4):e5166:1–e5166:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Srivatchan:2020:NLC

[SR20b] N. S. Srivatchan and P. Rangarajan. A novel low-cost smart energy meter based on IoT for developing countries’ micro grids. Concurrency and Computation: Practice and Experience, 32(4):e5042:1–e5042:??, February 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seshadri:2021:FIF

[SR21] R. Seshadri and S. Ramakrishnan. FPGA implementation of fast digital FIR and IIR filters. Concurrency and Computation: Practice and Experience, 33(3):e5246:1–e5246:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sadayan:2022:EDS

[SR22a] Geetha Sadayan and Karthiyayini Ramaiah. Enhanced data security in MANET using trust-based Bayesian statistical model with RSSI by AOMDV. Concurrency and Computation: Practice and Experience, 34(8):e5397:1–e5397:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shelar:2022:DMH

[SR22b] Manoj D. Shelar and S. Srinivasa Rao. Deep malware hunter based unrivaled malware detection schema thru cache retrospective empiricism. Concurrency and Computation: Practice and Experience, 34(19):e7025:1–e7025:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sokkalingam:2022:IID

[SR22c] Sumathi Sokkalingam and Rajesh Ramakrishnan. An intelligent intrusion detection system for distributed denial of service attacks: a support vector machine with hybrid optimization algorithm based approach. Concurrency and Computation: Practice and Experience, 34(27):e7334:1–e7334:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sujatha:2022:EEP

[SR22d] G Sujatha and R Jeberson Retnaraj. An efficient enhanced prefix hash tree model for optimizing the storage and image deduplication in cloud. Concurrency and Computation: Practice and Experience, 34(23):e7199:1–e7199:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2023:FIB

[SR23] Akhilesh Kumar Singh and Zahid Raza. A framework for IoT and blockchain based smart food chain management system. Concurrency and Computation: Practice and Experience, 35 (4):e7526:1–e7526:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shelar:2024:ECN

[SR24] Manoj D. Shelar and S. Srinivasa Rao. Enhanced capsule network-based executable files malware detection and classification-deep learning approach. Concurrency and Computation: Practice and Experience, 36(4):e7928:1–e7928:??, February 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Soltani:2022:AIE

[SRBH22] Nasim Soltani, Amir Masoud Rahmani, Mahdi Bohlouli, and Mehdi Hosseinzadeh. Artificial intelligence empowered threat detection in the Internet of Things: a systematic review. Concurrency and Computation: Practice and Experience, 34 (22):e6894:1–e6894:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Stein:2021:LAA

[SRG^{+}21] Charles M. Stein, Dinei A. Rockenbach, Dalvan Griebler, Massimo Torquati, Gabriele Mencagli, Marco Danelutto, and Luiz G. Fernandes. Latency-aware adaptive micro-batching techniques for streamed data compression on graphics processing units. Concurrency and Computation: Practice and Experience, 33(11):e5786:1–e5786:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sasikala:2022:TLB

[SRG^{+}22] S. Sasikala, S Ramesh, S Gomathi, S Balambigai, and V Anbumani. Transfer learning based recurrent neural network algorithm for linguistic analysis. Concurrency and Computation: Practice and Experience, 34(5):e6708:1–e6708:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Satheeskumar:2024:EVA

[SRGK24] R. Satheeskumar, Bonam Raghavaiah, S. Geerthik, and T. Senthil Kumar. An efficient VLSI architecture for fast motion estimation by hybridizing parallel spiral and adaptive threshold star diamond search algorithm-based search algorithm in HEVC. Concurrency and Computation: Practice and Experience, 36(1):e7873:1–e7873:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Souabi:2023:RAB

[SRIB23] Sonia Souabi, Asmaâ Retbi, Mohammed Khalidi Idrissi, and Samir Bennani. A recommendation approach based on correlation and co-occurrence within social learning network. Concurrency and Computation: Practice and Experience, 35(14): e6618:1–e6618:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sateesh:2023:OEL

[SRL23] Nayani Sateesh, Pasupureddy Srinivasa Rao, and Davuluri Rajya Lakshmi. Optimized ensemble learning-based student’s performance prediction with weighted rough set theory enabled feature mining. Concurrency and Computation: Practice and Experience, 35(7):e7601:1–e7601:??, March 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shanmugam:2023:ITB

[SRRM23] Vimala Shanmugam, Telu Venkata Madhusudhana Rao, Hanumantu Joga Rao, and Balajee Maram. Internet of Things based smart application for rice leaf disease classification using optimization integrated deep Maxout network. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sultana:2021:UUM

[SRS^{+}21] Nawrin Sultana, Martin Rüfenacht, Anthony Skjellum, Purushotham Bangalore, Ignacio Laguna, and Kathryn Mohror. Understanding the use of message passing interface in exascale proxy applications. Concurrency and Computation: Practice and Experience, 33(14):e5901:1–e5901:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sureshkumar:2023:EUW

[SRS23] V. Sureshkumar, S. Rajasomashekar, and B. Sarala. An efficient underground water prediction using optimal deep neural network. Concurrency and Computation: Practice and Experience, 35(1):e7421:1–e7421:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

S:2021:EPK

[SS21] Pavithra S and Janakiraman S. Enhanced polynomial kernel (EPK)-based support vector machine (SVM) (EPK-SVM) classification technique for speech recognition in hearingimpaired listeners. Concurrency and Computation: Practice and Experience, 33(3):e5210:1–e5210:??, February 10, 2021.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Seetharaman:2022:GSC

[SS22a] Akila Seetharaman and Allin Christe Sundersingh. Gene selection and classification using correlation feature selection based binary bat algorithm with greedy crossover. Concurrency and Computation: Practice and Experience, 34(5): e6718:1–e6718:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2022:IAP

[SS22b] Bhawna Sharma and Parvinder Singh. An improved antiphishing model utilizing TF-IDF and AdaBoost. Concurrency and Computation: Practice and Experience, 34(26):e7287:1– e7287:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2022:ECP

[SS22c] Durgesh M. Sharma and Shishir Kumar Shandilya. An efficient cyber-physical system using hybridized enhanced support-vector machine with Ada-boost classification algorithm. Concurrency and Computation: Practice and Experience, 34(21):e7134:1–e7134:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2022:GSA

[SS22d] Ankita Singh and Nanhay Singh. Gravitational search algorithm-driven missing links prediction in social networks. Concurrency and Computation: Practice and Experience, 34 (11):e6901:1–e6901:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2022:MAP

[SS22e] Khushal Singh and Nanhay Singh. Multilevel authentication protocol for enabling secure communication in Internet of Things. Concurrency and Computation: Practice and Experience, 34(3):e6578:1–e6578:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Suganthi:2022:IHF

[SS22f] Sumi Murugan Suganthi and Ganesh Ramaswamy Shunmugavel. Improved hard fusion methods for enhancing detection and energy efficiency in cognitive radio networks. Concurrency and Computation: Practice and Experience, 34(5): e6686:1–e6686:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sundaram:2022:ODE

[SS22g] Aarthi Sundaram and Chitrakala Sakthivel. Object detection and estimation: a hybrid image segmentation technique using convolutional neural network model. Concurrency and Computation: Practice and Experience, 34(21):e7114:1–e7114:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sabah:2023:NFE

[SS23a] Levent Sabah and Mehmet Simsek. A new fast entropy-based method to generate composite centrality measures in complex networks. Concurrency and Computation: Practice and Experience, 35(10):e7657:1–e7657:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sampath:2023:HTS

[SS23b] Anbukkarasi Sampath and Varadhaganapathy Shanmugavel. Hybrid Tamil spell checker with combined character splitting. Concurrency and Computation: Practice and Experience, 35 (1):e7440:1–e7440:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saravanakumar:2023:SPA

[SS23c] Selvaraj Saravanakumar and Thangavel Saravanan. Secure personal authentication in fog devices via multimodal ranklevel fusion. Concurrency and Computation: Practice and Experience, 35(10):e7673:1–e7673:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Satyavathi:2023:OIO

[SS23d] D. Monica Satyavathi and A. Ch. Sudhir. OQ-IICA: Optimal QoS-aware intra-inter cluster data aggregation technique for IoT-assisted WSNs using hybrid optimization techniques. Concurrency and Computation: Practice and Experience, 35 (22):e7723:1–e7723:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saviour:2023:IBF

[SS23e] Mariyaprincy Antony Saviour and Dhandapani Samiappan. IPFS based file storage access control and authentication model for secure data transfer using block chain technique. Concurrency and Computation: Practice and Experience, 35 (2):e7485:1–e7485:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2023:NSA

[SS23f] Supreet Singh and Urvinder Singh. A novel self-adaptive hybrid slime mould naked mole-rat algorithm for numerical optimization and energy-efficient wireless sensor network. Concurrency and Computation: Practice and Experience, 35 (23):e7809:1–e7809:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Suganya:2023:SML

[SS23g] S. Suganya and S. Selvamuthukumaran. Stacked multilayer security for Hadoop distributed file system using HSCT steganography. Concurrency and Computation: Practice and Experience, 35(21):e7711:1–e7711:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2024:FFD

[SS24] Hidangmayum Satyajeet Sharma and Khundrakpam Johnson Singh. A feed forward deep neural network model using feature selection for cloud intrusion detection system. Concurrency and Computation: Practice and Experience, 36(9): e8001:1–e8001:??, April 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Spalazzi:2023:BBC

[SSCN23] Luca Spalazzi, Francesco Spegni, Alessandra Corneli, and Berardo Naticchia. Blockchain based choreographies: The construction industry case study. Concurrency and Computation: Practice and Experience, 35(16):e6740:1–e6740:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2020:WRT

[SSDY20] Yan Song, Weiyun Si, Feifan Dai, and Guisong Yang. Weighted ReliefF with threshold constraints of feature selection for imbalanced data classification. Concurrency and Computation: Practice and Experience, 32(14):e5691:1– e5691:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sugitha:2022:QAR

[SSH22] George Sugitha, Thimmakkondu Babuji Sivakumar, and Shahul Hameed Hasan Hussain. QoS aware routing protocol using robust spatial Gabriel graph based clustering scheme for ad hoc network. Concurrency and Computation: Practice and Experience, 34(27):e7309:1–e7309:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sliem:2022:DMF

[SSI22] Mehdi Sliem, Nabila Salmi, and Malika Ioualalen. Designing a methodological framework for modeling and performance forecasting of self-adaptive cloud systems. Concurrency and Computation: Practice and Experience, 34(3): e6569:1–e6569:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shelke:2023:MSL

[SSK^{+}23a] Mahesh B. Shelke, Daivat D. Sawant, Chatrabhuj B. Kadam, Kailas Ambhure, and Sachin N. Deshmukh. Marathi Senti-WordNet: a lexical resource for sentiment analysis of Marathi. Concurrency and Computation: Practice and Experience, 35 (2):e7497:1–e7497:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2023:LFD

[SSK23b] Harpreet Singh, Birmohan Singh, and Manpreet Kaur. Lévy flight and disrupt operator-based elephant herding optimization for global optimization problems and feature selection to classify medical data. Concurrency and Computation: Practice and Experience, 35(23):e7766:1–e7766:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2021:JDM

[SSM^{+}21] Kaumudi Singh, Pratyush Shukla, Sachin S. M., Nithish K. Gnani, Prabhakar T. V., and Joy Kuri. Judicious data management for sustaining an energy harvesting sensor node. Concurrency and Computation: Practice and Experience, 33(23):e5997:1–e5997:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sakthivel:2022:BLI

[SSMT22a] N. K. Sakthivel, S. Subasree, Pachhaiammal Alias Priya M, and Amit Kumar Tyagi. Breast lesion identification and categorization using mammography screening based on combined convolutional recursive neural network framework with parameters optimized using multi-objective seagull optimization algorithm. Concurrency and Computation: Practice and Experience, 34(28):e7348:1–e7348:??, December 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sakthivel:2022:WBF

[SSMT22b] N. K. Sakthivel, S. Subasree, Shaveta Malik, and Amit Kumar Tyagi. A wrapper based feature extraction framework based on AlexNet deep convolutional neural network parameters optimized using gradient-based optimizer for mammogram images. Concurrency and Computation: Practice and Experience, 34(18):e7008:1–e7008:??, August 15, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sampaio:2021:EHR

[SSN21] Adrianno A. Sampaio, Alexandre C. Sena, and Alexandre S. Nery. Enabling heterogeneous ray-tracing acceleration in edge/cloud architectures. Concurrency and Computation: Practice and Experience, 33(11):e5822:1–e5822:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shaji:2022:NDN

[SSN22] B. Shaji, R. Lal Raja Singh, and K. L. Nisha. A novel deep neural network based marine predator model for effective classification of big data from social Internet of Things. Concurrency and Computation: Practice and Experience, 34(25): e7244:1–e7244:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singla:2023:RFO

[SSN23] Neetu Singla, Jyotsna Singh, and Sushama Nagpal. Raven finch optimized deep convolutional neural network model for intra-frame video forgery detection. Concurrency and Computation: Practice and Experience, 35(3):e7516:1–e7516:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2022:CDO

[SSP22] Rahul Kumar Singh, Manoj Kumar Sachan, and Ram Bahadur Patel. Cross-domain opinion classification via aspect analysis and attention sharing mechanism. Concurrency and Computation: Practice and Experience, 34(15):e6957:1– e6957:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2023:CDS

[SSP23] Rahul Kumar Singh, Manoj Kumar Sachan, and Ram Bahadur Patel. Cross-domain sentiment classification using decoding-enhanced bidirectional encoder representations from transformers with disentangled attention. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sonti:2024:SNB

[SSP24] Nagaraju Sonti, Rukmini Mulpuri Santhi Sri, and Venkatappa Reddy Pamulapati. Social network based link correlation using graph neural network with deep learning architectures for feature vectors prediction and classification. Concurrency and Computation: Practice and Experience, 36(14): e8090:1–e8090:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sundriyal:2020:RPA

[SSPG20] Vaibhav Sundriyal, Masha Sosonkina, David Poole, and Mark S. Gordon. Runtime power allocation approach for GAMESS hybrid CPU–GPU implementation. Concurrency and Computation: Practice and Experience, 32(24):e5917:1– e5917:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Suresh:2023:HOE

[SSRA23] Pandian Suresh, Selvaraj Shobana, Ganesan Ramya, and Manasea Selvin Belsam Jeba Ananth. Hybrid optimization enabled multi-aggregator-based charge scheduling of electric vehicle in internet of electric vehicles. Concurrency and Computation: Practice and Experience, 35(9):e7654:1–e7654:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Santos:2020:MLA

[SSS^{+}20] Reneilson Santos, Danilo Souza, Walter Santo, Admilson Ribeiro, and Edward Moreno. Machine learning algorithms to detect DDoS attacks in SDN. Concurrency and Computation: Practice and Experience, 32(16):e5402:1–e5402:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2023:MBS

[SSS23a] Durgesh M. Sharma, Shishir Kumar Shandilya, and Suresh Chandra Satapathy. Maximizing blockchain security: Merkle tree hash values generated through advanced vectorized elliptic curve cryptography mechanisms. Concurrency and Computation: Practice and Experience, 35(23):e7829:1–e7829:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sundaram:2023:PDD

[SSS23b] Mohana Ramaiyar Sundaram, Iwin Thanakumar Joseph Swamidason, and Velliangiri Sarveshwaran. Parkinson’s disease detection using sea lion shuffled shepherd optimization assisted deep Maxout network. Concurrency and Computation: Practice and Experience, 35(4):e7537:1–e7537:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2021:EDG

[SSSP21] Doohee Song, Moonbae Song, Vladimir Shakhov, and Kwangjin Park. Efficient dummy generation for considering obstacles and protecting user location. Concurrency and Computation: Practice and Experience, 33(2):e5146:1– e5146:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Salem:2020:SDS

[SSSR20] Farouk Salem, Florian Schintke, Thorsten Schütt, and Alexander Reinefeld. Scheduling data streams for low latency and high throughput on a Cray XC40 using Libfabric. Concurrency and Computation: Practice and Experience, 32 (20):e5563:1–e5563:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2024:MDM

[SSV24] Vikramjit Singh, Krishna Pal Sharma, and Harsh Kumar Verma. MF-DLB: Multimetric forwarding and directed acyclic graph-based load balancing for geographic routing in FANETs. Concurrency and Computation: Practice and Experience, 36(11):e8025:1–e8025:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shit:2022:PPC

[SSW^{+}22] Rathin Chandra Shit, Suraj Sharma, Paul Watters, Kumar Yelamarthi, Biswajeet Pradhan, Richard Davison, Graham Morgan, and Deepak Puthal. Privacy-preserving cooperative localization in vehicular edge computing infrastructure. Concurrency and Computation: Practice and Experience, 34(14): e5827:1–e5827:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Smyk:2021:DAG

[ST21] Adam Smyk and Marek Tudruj. Distributed application global states monitoring in PEGASUS DA applied to parallel graph partitioning. Concurrency and Computation: Practice and Experience, 33(11):e6052:1–e6052:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saar:2022:MST

[ST22a] Fadi Saar and Ahmet E. Topcu. Minimum spanning treebased cluster analysis: a new algorithm for determining inconsistent edges. Concurrency and Computation: Practice and Experience, 34(9):e6717:1–e6717:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Saminathan:2022:EED

[ST22b] Karunakaran Saminathan and Renukadevi Thangavel. Energy efficient and delay aware clustering in mobile adhoc network: a hybrid fruit fly optimization algorithm and whale optimization algorithm approach. Concurrency and Computation: Practice and Experience, 34(11):e6867:1–e6867:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sharma:2022:DDM

[ST22c] Ankur Sharma and Veni Thangaraj. DMAP: a decentralized matching game theory based optimized Internet of Things application placement in fog computing environment. Concurrency and Computation: Practice and Experience, 34(23): e7189:1–e7189:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Surendran:2022:DMA

[ST22d] Roopak Surendran and Tony Thomas. Detection of malware applications from centrality measures of syscall graph. Concurrency and Computation: Practice and Experience, 34(10): e6835:1–e6835:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Selvaraj:2023:OIG

[ST23a] Selvin Ebenezer Selvaraj and Raghuveera Tripuraribhatla. Optimization integrated generative adversarial network for occluded text recognition with language modeling. Concurrency and Computation: Practice and Experience, 35(8): e7630:1–e7630:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shibata:2023:SUD

[ST23b] Masahiro Shibata and Sébastien Tixeuil. Semi-uniform deployment of mobile robots in perfect ℓ-ary trees. Concurrency and Computation: Practice and Experience, 35(19):e7432:1– e7432:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Siddharthan:2023:NFA

[ST23c] Hariprasad Siddharthan and Deepa Thangavel. A novel framework approach for intrusion detection based on improved critical feature selection in Internet of Things networks. Concurrency and Computation: Practice and Experience, 35(1):e7445:1–e7445:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sundfeld:2020:UGA

[STH^{+}20] Daniel Sundfeld, George Teodoro, Jakob H. Havgaard, Jan Gorodkin, and Alba C. M. A. Melo. Using GPU to accelerate the pairwise structural RNA alignment with base pair probabilities. Concurrency and Computation: Practice and Experience, 32(10):e5468:1–e5468:??, May 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sangat:2021:NJP

[STISM21] Prajwol Sangat, David Taniar, Maria Indrawan-Santiago, and Christopher Messom. Nimble join: a parallel star join for main memory column-stores. Concurrency and Computation: Practice and Experience, 33(8):e5616:1–e5616:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2020:NST

[STJ^{+}20] Ying Sun, Jinrong Tian, Du Jiang, Bo Tao, Ying Liu, Juntong Yun, and Disi Chen. Numerical simulation of thermal insulation and longevity performance in new lightweight ladle. Concurrency and Computation: Practice and Experience, 32(22): e5830:1–e5830:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shewa:2023:KLT

[SU23] Gladys Amos Shewa and Fidelis Ifeayi Ugwuowo. Kibria– Lukman type estimator for gamma regression model. Concurrency and Computation: Practice and Experience, 35(1): e7441:1–e7441:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Selvi:2021:EDS

[SV21] R. Senthamil Selvi and M. L. Valarmathi. Enabling data security in data using vertical split with parallel feature selection using meta heuristic algorithms. Concurrency and Computation: Practice and Experience, 33(3):e5248:1–e5248:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

S:2022:GCR

[SV22a] Sheela S. and Sathyanarayana S. V. Generation of chaotic random binary sequences for cryptographic applications. Concurrency and Computation: Practice and Experience, 34(1): e6497:1–e6497:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Selvi:2022:DUD

[SV22b] Seruvan Tamil Selvi and Pandikkannu Visalakshi. Detection of unique delete attack in wireless sensor network using gradient thresholding-long short-term memory algorithm. Concurrency and Computation: Practice and Experience, 34(27): e7332:1–e7332:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sham:2022:IAC

[SV22c] Eht E Sham and Deo Prakash Vidyarthi. Intelligent admission control manager for fog-integrated cloud: a hybrid machine learning approach. Concurrency and Computation: Practice and Experience, 34(10):e6687:1–e6687:??, May 1, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Siddaramanna:2022:KSB

[SV22d] Sheela Siddaramanna and Sathyanarayana Sarapady Venkatramanayya. Key sequences based on cyclic elliptic curves over GF(2^{8}) with logistic map for cryptographic applications. Concurrency and Computation: Practice and Experience, 34 (11):e6849:1–e6849:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sumalakshmi:2022:FDL

[SV22e] Chundakath House Sumalakshmi and Perumal Vasuki. Fused deep learning based facial expression recognition of students in online learning mode. Concurrency and Computation: Practice and Experience, 34(21):e7137:1–e7137:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shir:2023:MPU

[SVB23] Bhargav Shir, Jai Prakash Verma, and Pronaya Bhattacharya. Mobility prediction for uneven distribution of bikes in bike sharing systems. Concurrency and Computation: Practice and Experience, 35(2):e7465:1–e7465:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Selvi:2022:IMH

[SVD22a] Ramar Senthamil Selvi, Muniyappan Lakshapalam Valarmathi, and Prathima Devadas. Improved meta-heuristic algorithm for selecting optimal features: a big data classification model. Concurrency and Computation: Practice and Experience, 34(17):e7000:1–e7000:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shanmugam:2022:ESO

[SVD^{+}22b] Padmapriya Shanmugam, Balajivijayan Venkateswarulu, Rajalakshmi Dharmadurai, Thiagarajan Ranganathan, Mohan Indiran, and Manikandan Nanjappan. Electro search optimization based long short-term memory network for mobile malware detection. Concurrency and Computation: Practice and Experience, 34(19):e7044:1–e7044:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shah:2024:DNM

[SVM24] Akash Shah, Sapna Varshney, and Monica Mehrotra. Deep-MUI: a novel method to identify malicious users on online social network platforms. Concurrency and Computation: Practice and Experience, 36(3):e7917:1–e7917:??, February 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sankari:2022:CML

[SVS22] Subbiah Sankari, Subramaniam Sankaran Varshini, and Savvas Mohamed Aafia Shifana. COVID-19: Machine learning for safe transportation. Concurrency and Computation: Practice and Experience, 34(19):e7041:1–e7041:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Samfass:2020:LTO

[SWCB20] Philipp Samfass, Tobias Weinzierl, Dominic E. Charrier, and Michael Bader. Lightweight task offloading exploiting MPI wait times for parallel adaptive mesh refinement. Concurrency and Computation: Practice and Experience, 32(24):e5916:1– e5916:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

SenthamilSelvan:2022:ICA

[SWK22] R. SenthamilSelvan, R. S. D. Wahidabanu, and B. Karthik. Intersection collision avoidance in dedicated short-range communication using vehicle ad hoc network. Concurrency and Computation: Practice and Experience, 34(13):e5856:1– e5856:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2024:LSE

[SWS^{+}24] Dawei Sun, Yueru Wang, Jialiang Sui, Shang Gao, Jia Rong, and Rajkumar Buyya. Lc-Stream: an elastic scheduling strategy with latency constraints in geo-distributed stream computing environments. Concurrency and Computation: Practice and Experience, 36(14):e8085:1–e8085:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2020:ISM

[SWZW20] Yongqiang Song, Hong Wang, Changyong Zhang, and Lutong Wang. Impression space model for the evaluation of Internet advertising effectiveness. Concurrency and Computation: Practice and Experience, 32(11):e5678:1–e5678:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2021:RCC

[SX21] Xiao Shen and Chonghuan Xu. Research on children’s cognitive development for learning disabilities using recommendation method. Concurrency and Computation: Practice and Experience, 33(9):e6097:1–e6097:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2023:BLD

[SXC^{+}23] Daoyuan Song, Xinghua Xu, Xiaopeng Cui, Yangbin Ou, and Weiming Chen. Bolt looseness detection based on Canny edge detection algorithm. Concurrency and Computation: Practice and Experience, 35(21):e7713:1–e7713:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sadiq:2023:SPA

[SY23] Muhammad Umair Sadiq and Muhammad Murtaza Yousaf. A scalable parallel algorithm for global sequence alignment with customizable scoring scheme. Concurrency and Computation: Practice and Experience, 35(25):e7888:1–e7888:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Singh:2022:EAT

[SYG22] Joginder Singh, Naresh Kumar Yadav, and Saral Kumar Gupta. Enhancement of available transfer capability using TCSC with hybridized model: Combining lion and moth flame algorithms. Concurrency and Computation: Practice and Experience, 34(21):e7052:1–e7052:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2020:PST

[SYJL20] Guimin Shi, Sheng Yang, Yi Jiang, and Zhiming Luo. Pattern synthesis of thinned multi-input multi-output radar using difference set and differential evolution. Concurrency and Computation: Practice and Experience, 32(22):e5847:1–e5847:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2023:TDP

[SYL23] Shuaiyong Shen, Yang Yang, and Ximeng Liu. Toward data privacy preservation with ciphertext update and key rotation for IoT. Concurrency and Computation: Practice and Experience, 35(20):e6729:1–e6729:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sirisala:2022:NTR

[SYRP22] NageswaraRao Sirisala, Anitha Yarava, Y. C. A. Padmanabha Reddy, and Veeresh Poola. A novel trust recommendation model in online social networks using soft computing methods. Concurrency and Computation: Practice and Experience, 34(22):e7153:1–e7153:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sadeghi-Yarandi:2022:MOU

[SYRS^{+}22] Mohsen Sadeghi-Yarandi, Ehsan Rastegarzadeh, Ahmad Soltanzadeh, Mostafa Mohammad-Ghasemi, Shahram Arsang-Jang, Sevda Panahi, and Ali Karimi. Modeling the occurrence of unsafe behaviors based on safety climate and organizational leadership style dimensions in oil platforms using the fuzzy logic approach. Concurrency and Computation: Practice and Experience, 34(4):e6612:1–e6612:??, February 15, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2021:MGA

[SYT^{+}21] Ying Sun, Zhiwen Yang, Bo Tao, Guozhang Jiang, Zhiqiang Hao, and Baojia Chen. Multiscale generative adversarial network for real-world super-resolution. Concurrency and Computation: Practice and Experience, 33(21):e6430:1–e6430:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2024:SKA

[SZ24] Xiaoping Sun and Hai Zhuge. Semantics, knowledge and advanced cyber-infrastructure for intelligent applications. Concurrency and Computation: Practice and Experience, 36(10): e7022:1–e7022:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2024:MAI

[SZG^{+}24] Wenhao Song, Wenzhe Zhai, Mingliang Gao, Qilei Li, Abdellah Chehri, and Gwanggil Jeon. Multiscale aggregation and illumination-aware attention network for infrared and visible image fusion. Concurrency and Computation: Practice and Experience, 36(10):e7712:1–e7712:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2021:SMP

[SZGR21] Yong Shi, Yuanchun Zheng, Kun Guo, and Xinyue Ren. Stock movement prediction with sentiment analysis based on deep learning networks. Concurrency and Computation: Practice and Experience, 33(6):e6076:1–e6076:??, March 25, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sakhdari:2023:ECS

[SZI^{+}23] Jalal Sakhdari, Behrooz Zolfaghari, Shaghayegh Izadpanah, Samaneh Mahdizadeh Zargar, Mahla Rahati Quchani, Mahsa Shadi, Saeid Abrishami, and Abbas Rasoolzadegan. Edge computing: a systematic mapping study. Concurrency and Computation: Practice and Experience, 35(22):e7741:1– e7741:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2021:TAM

[SZJ21] Heng Song, Junwu Zhu, and Yi Jiang. On truthful auction mechanism for cloud resources allocation and consumption shifting with different time slots. Concurrency and Computation: Practice and Experience, 33(22):e6122:1–e6122:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2022:HAM

[SZL^{+}22] Hui Song, Wenli Zhang, Ke Liu, Yifan Shen, and Mingyu Chen. HCMonitor: an accurate measurement system for high concurrent network services. Concurrency and Computation: Practice and Experience, 34(12):e6081:1–e6081:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2020:MLT

[SZqWZ20] Hongyan Shi, NanDong Zhang, Xiao qiang Wu, and Yu-Dong Zhang. Multimodal lung tumor image recognition algorithm based on integrated convolutional neural network. Concurrency and Computation: Practice and Experience, 32(21): e4965:1–e4965:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2020:SIH

[SZS20] Xiaolin Shi, Pei Zhan, and Yimin Shi. Statistical inference for a hybrid system model with incomplete observed data under adaptive progressive hybrid censoring. Concurrency and Computation: Practice and Experience, 32(14):e5708:1– e5708:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2022:DML

[SZW^{+}22] Sheng Shen, Tianqing Zhu, Di Wu, Wei Wang, and Wanlei Zhou. From distributed machine learning to federated learning: In the view of data privacy and security. Concurrency and Computation: Practice and Experience, 34(16):e6002:1– e6002:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2022:NDP

[SZY^{+}22] Sheng Shen, Tianqing Zhu, Dayong Ye, Minghao Wang, Xuhan Zuo, and Andi Zhou. A novel differentially private advising framework in cloud server environment. Concurrency and Computation: Practice and Experience, 34(7):e5932:1– e5932:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2021:IAH

[SZZ^{+}21] Peng Shi, Zhen Zhao, Huaqiang Zhong, Hangyu Shen, and Lianhong Ding. An improved agglomerative hierarchical clustering anomaly detection method for scientific data. Concurrency and Computation: Practice and Experience, 33(6): e6077:1–e6077:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shen:2022:DCA

[SZZ^{+}22] Ziyu Shen, Qing Zhou, Xusheng Zhang, Bin Xia, Zheng Liu, and Yun Li. Data characteristics aware prediction model for power consumption of data center servers. Concurrency and Computation: Practice and Experience, 34(11):e6902:1– e6902:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shang:2024:AHC

[SZZ^{+}24a] Jiangwei Shang, Zhan Zhang, Kun Zhang, Chuanyou Li, Lei Qian, and Hongwei Liu. An algorithm/hardware co-optimized method to accelerate CNNs with compressed convolutional weights on FPGA. Concurrency and Computation: Practice and Experience, 36(11):e8011:1–e8011:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2024:SST

[SZZ^{+}24b] Li Sun, Juanjuan Zhao, Jun Zhang, Fan Zhang, Kejiang Ye, and Chengzhong Xu. STRmt: a state transition based model for real-time crowd counting in a metro system. Concurrency and Computation: Practice and Experience, 36(14):e8086:1– e8086:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thati:2022:GSE

[TA22a] Jagadeesh Thati and Samit Ari. GLeSI: a system for extraction of glacial lakes using satellite imagery. Concurrency and Computation: Practice and Experience, 34(23):e7184:1– e7184:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tuncer:2022:CES

[TA22b] Seda Arslan Tuncer and Ahmet Alkan. Classification of EMG signals taken from arm with hybrid CNN-SVM architecture. Concurrency and Computation: Practice and Experience, 34 (5):e6746:1–e6746:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Taha:2023:SAW

[TA23a] Miran Taha and Aree Ali. Smart algorithm in wireless networks for video streaming based on adaptive quantization. Concurrency and Computation: Practice and Experience, 35 (9):e7633:1–e7633:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tufek:2023:PET

[TA23b] Alper Tufek and Mehmet S. Aktas. On the provenance extraction techniques from large scale log files. Concurrency and Computation: Practice and Experience, 35(15):e6559:1– e6559:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Teja:2021:EFP

[TAB21] Nandireddygari Ramya Teja, S. Arunmetha, and Srinivas Bachu. An efficient field programmable gate array based hardware architecture for efficient motion estimation with parallel implemented genetic algorithm. Concurrency and Computation: Practice and Experience, 33(24):e6459:1–e6459:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Touati:2022:NDN

[TAH22] Haifa Touati, Ahmed Aboud, and Brahim Hnich. Named data networking-based communication model for Internet of Things using energy aware forwarding strategy and smart sleep mode. Concurrency and Computation: Practice and Experience, 34(3):e6584:1–e6584:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Terzi:2022:FPR

[TAT22] Ramazan Terzi, Nuh Azginoglu, and Duygu Sinanc Terzi. False positive repression: Data centric pipeline for object detection in brain MRI. Concurrency and Computation: Practice and Experience, 34(20):e6821:1–e6821:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tajima:2023:NNL

[TAT^{+}23] Yoshiharu Tajima, Masahiro Asaoka, Akihiro Tabuchi, Akihiko Kasagi, and Tsuguchika Tabaru. NEEBS: Nonexpert large-scale environment building system for deep neural network. Concurrency and Computation: Practice and Experience, 35(19):e7499:1–e7499:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thanupillai:2023:PCN

[TB23] Kalai Selvi Thanupillai and Rahimunnisa Kamal Basha. Pulse coupled neural network optimized with chaotic grey wolf algorithm for breast cancer classification using mammogram images. Concurrency and Computation: Practice and Experience, 35(1):e7448:1–e7448:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thinh:2023:HPA

[TBNPQ23] Tran Ngoc Thinh, Tran Hoang Quoc Bao, Duc-Minh Ngo, and Cuong Pham-Quoc. High-performance anomaly intrusion detection system with ensemble neural networks on reconfigurable hardware. Concurrency and Computation: Practice and Experience, 35(15):e6370:1–e6370:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thamsen:2021:MHH

[TBT^{+}21] Lauritz Thamsen, Jossekin Beilharz, Vinh Thuy Tran, Sasho Nedelkoski, and Odej Kao. Mary, Hugo, and Hugo*: Learning to schedule distributed data-parallel processing jobs on shared clusters. Concurrency and Computation: Practice and Experience, 33(18):e5823:1–e5823:??, September 25, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Topcu:2022:SIH

[TC22] Ahmet E. Topcu and Ali Osman Cibikdiken. Special issue on high performance computing conference (BASARIM-2020). Concurrency and Computation: Practice and Experience, 34 (9):e6972:1–e6972:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Toor:2024:CMC

[TC24] Rashmeet Toor and Inderveer Chana. Cloud menu: Cloud based network analysis for disease-diet associations and recommendations. Concurrency and Computation: Practice and Experience, 36(14):e8065:1–e8065:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tsai:2023:PPP

[TCA23] Yu-Hsiang M. Tsai, Terry Cojean, and Hartwig Anzt. Providing performance portable numerics for Intel GPUs. Concurrency and Computation: Practice and Experience, 35(20): e7400:1–e7400:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tan:2023:MET

[TCRP23] Chaoyue Tan, Yuling Chen, Xiaojun Ren, and Changgen Peng. A mobile energy trading scheme based on Lightning Network. Concurrency and Computation: Practice and Experience, 35(20):e6623:1–e6623:??, September 10, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thabit:2023:DST

[TCW^{+}23] Fursan Thabit, Ozgu Can, Rizwan Uz Zaman Wani, Mohammed Ali Qasem, S. B. Thorat, and Hoda A. Alkhzaimi. Data security techniques in cloud computing based on machine learning algorithms and cryptographic algorithms: Lightweight algorithms and genetics algorithms. Concurrency and Computation: Practice and Experience, 35(21):e7691:1– e7691:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tung:2021:DIB

[TD21] Nguyen Thanh Tung and Nguyen Huu Duc. Design and implementing Big Data system for cardiovascular data. Concurrency and Computation: Practice and Experience, 33(2): e5068:1–e5068:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2024:RCR

[TDZ^{+}24] Yongwei Tang, Wei Ding, Jun Zhou, Yonghao Yu, and Huijuan Hao. Research on cloud robot security strategy based on chaos encryption. Concurrency and Computation: Practice and Experience, 36(10):e8007:1–e8007:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tan:2023:IDM

[TFZC23] Yumeng Tan, Xupeng Fu, Jianbo Zhu, and Lina Chen. A improved detection method for lung nodule based on multi-scale 3dconvolutional neural network. Concurrency and Computation: Practice and Experience, 35(13):e7034:1–e7034:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tyagi:2023:BBS

[TG23] Nitin K. Tyagi and Mukta Goyal. Blockchain-based smart contract for issuance of country of origin certificate for Indian Customs Exports Clearance. Concurrency and Computation: Practice and Experience, 35(16):e6249:1–e6249:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Torjmen:2022:TST

[TH22] Roua Torjmen and Kais Haddar. Translation system from Tunisian dialect to Modern Standard Arabic. Concurrency and Computation: Practice and Experience, 34(6):e6788:1– e6788:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Theerthagiri:2021:FHH

[The21] Prasannavenkatesan Theerthagiri. Forecasting hyponatremia in hospitalized patients using multilayer perceptron and multivariate linear regression techniques. Concurrency and Computation: Practice and Experience, 33(16):e6248:1–e6248:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tran:2020:CRE

[THT20] Quang Thanh Tran, Li Hao, and Quang Khai Trinh. A comprehensive research on exponential smoothing methods in modeling and forecasting cellular traffic. Concurrency and Computation: Practice and Experience, 32(23):e5602:1– e5602:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2021:IPM

[THW21] Tiantian Tang, Pei Hu, and Ge Wu. Influence of promotion mode on purchase decision based on multilevel psychological distance dimension of visual attention model and data mining. Concurrency and Computation: Practice and Experience, 33 (12):e5587:1–e5587:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2023:R

[THW23] Tiantian Tang, Pei Hu, and Ge Wu. Retraction. Concurrency and Computation: Practice and Experience, 33(12):e7753:1– e7753:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2022:FLM

[THX22] Zhongyun Tang, Haiyang Hu, and Chonghuan Xu. A federated learning method for network intrusion detection. Concurrency and Computation: Practice and Experience, 34(10): e6812:1–e6812:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tastan:2022:ACC

[TI22] Birhan Tastan and Hatice Imamoglu. The analysis of crosscorrelation between Istanbul Stock Exchange and major stock markets and indices: an empirical analysis using Random Matrix Theory. Concurrency and Computation: Practice and Experience, 34(21):e7113:1–e7113:??, September 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tan:2024:PSL

[TJC24] Zuowen Tan, Jintao Jiao, and Yanliang Chen. Provably secure lightweight certificateless lattice-based authenticated key exchange scheme for IIoT. Concurrency and Computation: Practice and Experience, 36(8):e7983:1–e7983:??, April 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tan:2023:MLE

[TJZ23] Chao Tan, Genlin Ji, and Xiaoqian Zeng. Multi-label enhancement manifold learning algorithm for vehicle video. Concurrency and Computation: Practice and Experience, 35 (13):e6660:1–e6660:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thangavel:2022:DTB

[TK22a] Subburaj Thangavel and Suthendran Kannan. Detection and trace back of low and high volume of distributed denial-of-service attack based on statistical measures. Concurrency and Computation: Practice and Experience, 34(8):e5428:1– e5428:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Theerthagiri:2022:VMI

[TK22b] Prasannavenkatesan Theerthagiri and Chandrasekaran Gopala Krishnan. Vehicular multihop intelligent transportation framework for effective communication in vehicular ad-hoc networks. Concurrency and Computation: Practice and Experience, 34(10):e6833:1–e6833:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tripathy:2024:IAL

[TK24] Pujasuman Tripathy and P. M. Khilar. Improvement of amorphous localization algorithm in WSN using ALO and GWO. Concurrency and Computation: Practice and Experience, 36 (11):e8036:1–e8036:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tyagi:2023:SPP

[TKG^{+}23] Pooja Tyagi, Saru Kumari, Mridul Kumar Gupta, Chien-Ming Chen, Tsu-Yang Wu, and Sachin Kumar. A secure protocol for patient monitoring in wireless body area networks. Concurrency and Computation: Practice and Experience, 35(10):e7676:1–e7676:??, May 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thiyyakat:2022:NNS

[TKS22] Meghana Thiyyakat, Subramaniam Kalambur, and Dinkar Sitaram. Niyama: Node scheduling for cloud workloads with resource isolation. Concurrency and Computation: Practice and Experience, 34(23):e7196:1–e7196:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tran:2021:AHV

[TL21] Phuoc Vinh Tran and Truong Xuan Le. Approaching human vision perception to designing visual graph in data visualization. Concurrency and Computation: Practice and Experience, 33(2):e5722:1–e5722:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tao:2022:RBP

[TLH^{+}22] Bo Tao, Yao Liu, Licheng Huang, Guanjun Chen, and Baojia Chen. 3D reconstruction based on photoelastic fringes. Concurrency and Computation: Practice and Experience, 34 (1):e6481:1–e6481:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2020:ITW

[TLL20] Zhong Tang, Wenqiang Li, and Yan Li. An improved term weighting scheme for text classification. Concurrency and Computation: Practice and Experience, 32(9):e5604:1– e5604:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2021:CBK

[TLQ21] Hao Tang, Baisong Liu, and Jiangbo Qian. Content-based and knowledge graph-based paper recommendation: Exploring user preferences with the knowledge graphs for scientific paper recommendation. Concurrency and Computation:

Practice and Experience, 33(13):e6227:1–e6227:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2022:APD

[TLS22] Hui Tian, Alan Wee-Chung Liew, and Hong Shen. Advances in parallel and distributed computing and its applications. Concurrency and Computation: Practice and Experience, 34 (2):e6667:1–e6667:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2022:TTP

[TLX22] Wenlong Tian, Ruixuan Li, and Zhiyong Xu. TSS: a twoparty secure server-aid chunking algorithm. Concurrency and Computation: Practice and Experience, 34(12):e6577:1– e6577:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2021:SDE

[TLXX21] Wenlong Tian, Ruixuan Li, Cheng-Zhong Xu, and Zhiyong Xu. Sed-dedup: an efficient secure deduplication system with data modifications. Concurrency and Computation: Practice and Experience, 33(15):e5350:1–e5350:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tesser:2024:LPP

[TMB24] Rafael Keller Tesser, Alvaro Marques, and Edson Borin. A lightweight performance proxy for deep-learning model training on Amazon SageMaker. Concurrency and Computation: Practice and Experience, 36(14):e8104:1–e8104:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Takafuji:2023:EPI

[TNI23] Daisuke Takafuji, Koji Nakano, and Yasuaki Ito. Efficient parallel implementations to compute the diameter of a graph. Concurrency and Computation: Practice and Experience, 35 (11):e5963:1–e5963:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Takafuji:2023:GID

[TNIK23] Daisuke Takafuji, Koji Nakano, Yasuaki Ito, and Akihiko Kasagi. GPU implementations of deflate encoding and decoding. Concurrency and Computation: Practice and Experience, 35(19):e7454:1–e7454:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thiruvenkadam:2021:ASI

[TNP21] Kalaiselvi Thiruvenkadam, Kalaichelvi Nagarajan, and Sriramakrishnan Padmanaban. An automatic self-initialized clustering method for brain tissue segmentation and pathology detection from magnetic resonance human head scans with graphics processing unit machine. Concurrency and Computation: Practice and Experience, 33(6):e6084:1–e6084:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Turk:2021:CNN

[TÖ21] Ömer Türk and Mehmet Siraç Özerdem. The convolutional neural network approach from electroencephalogram signals in emotional detection. Concurrency and Computation: Practice and Experience, 33(20):e6356:1–e6356:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Topcu:2022:PMM

[TO22] Ahmet E. Topcu and Isameddin Omak. Parallelization of a meteorological model using message passing interface and CUDA: a case study with the inversion estimation algorithm. Concurrency and Computation: Practice and Experience, 34 (9):e6738:1–e6738:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tekeli:2021:IVN

[TÖK21] Erkut Tekeli, Nimet Özbay, and Selahattin Kaçiranlar. Implementation and validation of new optimization methods by genetic algorithm for two-parameter ridge estimator. Concurrency and Computation: Practice and Experience, 33(9): e6088:1–e6088:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2020:ACC

[TP20] Junfeng Tian and Yanan Pang. Adjoin: a causal consistency model based on the adjacency list in a distributed system. Concurrency and Computation: Practice and Experience, 32(22):e5835:1–e5835:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tran:2022:DFD

[TPT^{+}22] Nam N. Tran, Hemanshu R. Pota, Quang N. Tran, Xuefei Yin, and Jiankun Hu PhD. Designing false data injection attacks penetrating AC-based bad data detection system and FDI dataset generation. Concurrency and Computation: Practice and Experience, 34(7):e5956:1–e5956:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tran:2024:ACM

[TRM^{+}24] Tu Tran, Bharath Ramesh, Benjamin Michalowicz, Mustafa Abduljabbar, Hari Subramoni, Aamir Shafi, and Dhabaleswar K. Panda. Accelerating communication with multi-HCA aware collectives in MPI. Concurrency and Computation: Practice and Experience, 36(1):e7879:1–e7879:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Talmale:2021:CFT

[TS21] Girish Talmale and Urmila Shrawankar. Cluster formation techniques for hierarchical real time tasks allocation on multiprocessor system. Concurrency and Computation: Practice and Experience, 33(21):e6438:1–e6438:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tripathi:2022:EMB

[TS22] Vijay Kumar Tripathi and Mayank Singh. An efficient metrics based self-adaptive design model by multiobjective gray wolf optimization with extreme learning machine for autonomic computing system application. Concurrency and Computation: Practice and Experience, 34(4):e6609:1–e6609:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tay:2024:DTB

[TS24a] Muhammet Tay and Arafat Sentürk. Decision tree-based task offloading in vehicle edge computing. Concurrency and Computation: Practice and Experience, 36(12):e8026:1–e8026:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thakur:2024:SEC

[TS24b] Payal Thakur and Vivek Kumar Sehgal. Synergizing edge computing and blockchain for cyber-physical systems. Concurrency and Computation: Practice and Experience, 36(12): e8066:1–e8066:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Teimouri:2021:PFE

[TSA21] Mehdi Teimouri, Zahra Seyedghorban, and Fatemeh Amirjani. Fragments-Expert: a graphical user interface MAT-LAB toolbox for classification of file fragments. Concurrency and Computation: Practice and Experience, 33(9):e6154:1– e6154:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Turkyilmaz:2023:EML

[TSB23] Yasin Turkyilmaz, Arafat Senturk, and Muhammed Enes Bayrakdar. Employing machine learning based malicious signal detection for cognitive radio networks. Concurrency and Computation: Practice and Experience, 35(2):e7457:1– e7457:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tissaoui:2022:TEA

[TSCM22] Anis Tissaoui, Salma Sassi, Richard Chbeir, and Ameni Mechergui. A top-down enriching approach for ontology learning from text. Concurrency and Computation: Practice and Experience, 34(19):e7036:1–e7036:??, August 30, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tyagi:2023:PST

[TSD23] Anshuman Tyagi, Pawan Singh, and Harsh Dev. Proposed spatio-temporal features for human activity classification using ensemble classification model. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thakur:2021:NDB

[TSG21] S. Thakur, A. K. Singh, and S. P. Ghrera. NSCT domainbased secure multiple-watermarking technique through lightweight encryption for medical images. Concurrency and Computation: Practice and Experience, 33(2):e5108:1– e5108:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tan:2021:DIC

[TSLL21] Manchun Tan, Zhiqiang Song, Yunfeng Liu, and Zhong Li. Dynamics of interval Cohen–Grossberg neural networks with time-varying delays based on LMI computation. Concurrency and Computation: Practice and Experience, 33(8):e5023:1– e5023:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tehrani:2022:FPM

[TSR22] Ahmad Fadaei Tehrani, Mahdi Sharifi, and Amir Masoud Rahmani. Frequent pattern mining algorithms in fog computing environments: a systematic review. Concurrency and Computation: Practice and Experience, 34(24):e7229:1– e7229:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tuli:2022:PNC

[TSV^{+}22] Rohan Tuli, Hitesh Soneji, Sahil Vahora, Prathamesh Churi, and Nagachetan M. Bangalore. PixJS: a novel chaos-based approach for image encryption. Concurrency and Computation: Practice and Experience, 34(17):e6990:1–e6990:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Toz:2021:NAN

[TT21] Metin Toz and Güliz Toz. A novel approach for naturebased optimization algorithms: The threat factor approach. Concurrency and Computation: Practice and Experience, 33 (20):e6341:1–e6341:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Theerthagiri:2022:MSP

[TT22] Prasannavenkatesan Theerthagiri and Menakadevi Thangavelu. Mobility speed prediction using ARIMA and RNN for random walk mobility model in mobile ad hoc networks. Concurrency and Computation: Practice and Experience, 34(4): e6625:1–e6625:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2023:FGA

[TT23a] Junfeng Tian and Naxi Tian. Fine-grained assured insertion and deletion scheme based on onion encryption in cloud storage. Concurrency and Computation: Practice and Experience, 35(27):e7852:1–e7852:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tung:2023:DCF

[TT23b] Nguyen Thanh Tung and Le Duc Thinh. Design cloud-fog systems based on the energy of Internet of Things devices. Concurrency and Computation: Practice and Experience, 35(15): e6284:1–e6284:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tawalbeh:2020:IIP

[TTA20] Lo’ai Tawalbeh, Mohammad A. Tawalbeh, and Monther Aldwairi. Improving the impact of power efficiency in mobile cloud applications using cloudlet model. Concurrency and Computation: Practice and Experience, 32(21):e5709:1– e5709:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tasgetiren:2022:DSA

[TTB^{+}22] Nail Tasgetiren, Umit Tigrak, Erdal Bozan, Guven Gul, Emir Demirci, Hakan Saribiyik, and Mehmet S. Aktas. On the distributed software architecture of a data analysis workflow: a case study. Concurrency and Computation: Practice and Experience, 34(9):e6522:1–e6522:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Taktak:2022:HAS

[TTM^{+}22] Emna Taktak, Mohamed Tounsi, Mohamed Mosbah, Abdessalem Mnif, and Ahmed Hadj Kacem. A handshake algorithm for scheduling communications in wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(7):e6038:1–e6038:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2022:RTS

[TTZX22] Xingyu Tian, Shengnan Tang, Huihui Zhu, and Daoxun Xia. Real-time sentiment analysis of students based on minixception architecture for wisdom classroom. Concurrency and Computation: Practice and Experience, 34(21):e7059:1– e7059:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Turan:2021:NMR

[TUD21] Erhan Turan, Ferhat Ucar, and Besir Dandil. A novel marble recognition system using extreme learning machine with LBP and histogram features. Concurrency and Computation: Practice and Experience, 33(21):e6428:1–e6428:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Turukmane:2023:FIB

[Tur23] Anil V. Turukmane. Forecasting the IoT-based cyber threats using the hybrid forage dependent ensemble classifier. Concurrency and Computation: Practice and Experience, 35(2): e7460:1–e7460:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2021:NRP

[TWG^{+}21] Yayuan Tang, Hao Wang, Kehua Guo, Tao Luo, and Tao Chi. A new replica placement mechanism for mobile media streaming in edge computing. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Thoman:2022:MGR

[TWH^{+}22] Peter Thoman, Markus Wippler, Robert Hranitzky, Philipp Gschwandtner, and Thomas Fahringer. Multi-GPU room response simulation with hardware raytracing. Concurrency and Computation: Practice and Experience, 34(4):e6663:1– e6663:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tan:2020:LEB

[TWL^{+}20] Yejin Tan, Wei Wu, Jian Liu, Huimei Wang, and Ming Xian. Lightweight edge-based kNN privacy-preserving classification scheme in cloud computing circumstance. Concurrency and Computation: Practice and Experience, 32(19):e5804:1– e5804:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2021:WAM

[TWQ^{+}21] Wenda Tang, Chaobing Wu, Lianyong Qi, Xuyun Zhang, Xiaolong Xu, and Wanchun Dou. A WiFi-aware method for mobile data offloading with deadline constraints. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tang:2021:NFC

[TWW^{+}21] Leiyu Tang, Chenxi Wang, Shudong Wang, Guoliang Dong, and Jiancong Fan. A novel fuzzy clustering algorithm based on rough set and inhibitive factor. Concurrency and Computation: Practice and Experience, 33(6):e6078:1–e6078:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tai:2022:TBB

[TWW^{+}22] Yonghang Tai, Yinjia Wang, Zhifeng Wang, Feiyan Li, Lei Wei, Lei Pan, Jun Zhang, and Junsheng Shi. Trustworthy blockchain-based medical Internet of thing for minimal invasive surgery training simulator. Concurrency and Computation: Practice and Experience, 34(16):e5816:1–e5816:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2021:FST

[TWXL21] Zhiqiang Tian, Chunhui Wang, Youzi Xiao, and Yuping Lin. Flexible scene text recognition based on dual attention mechanism. Concurrency and Computation: Practice and Experience, 33(22):e5863:1–e5863:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2024:SSL

[TXT24] Junfeng Tian, Hongwei Xu, and Jin Tian. SLChain: a secure and low-storage pressure sharding blockchain. Concurrency and Computation: Practice and Experience, 36(3):e7918:1– e7918:??, February 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2022:AMR

[TY22] Junfeng Tian and Qian Yang. An arbitrable multi-replica data auditing scheme based on smart contracts. Concurrency and Computation: Practice and Experience, 34(22):e7164:1– e7164:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tezgider:2022:TCU

[TYA22] Murat Tezgider, Beytullah Yildiz, and Galip Aydin. Text classification using improved bidirectional transformer. Concurrency and Computation: Practice and Experience, 34(9): e6486:1–e6486:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Teng:2022:RDP

[TYL22] Qi Teng, Jianjun Yi, and Huazhong Li. Research of 6-DOF pose estimation in stacked scenes. Concurrency and Computation: Practice and Experience, 34(3):e6563:1–e6563:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tan:2023:FEE

[TYLY23] Sixiang Tan, Wenzhong Yang, JianZhuang Lin, and Weijie Yu. Feature extraction and enhancement for real-time semantic segmentation. Concurrency and Computation: Practice and Experience, 35(17):e6573:1–e6573:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

T:2022:OAC

[TZ22] Suresh T and Brijet Z. Optimization assisted convolutional neural network for detection of thyroid. Concurrency and Computation: Practice and Experience, 34(25):e7268:1– e7268:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2023:DPT

[TZ23] Junfeng Tian and Qi Zhu. A differential privacy trajectory data storage and publishing scheme based on radix tree. Concurrency and Computation: Practice and Experience, 35 (22):e7731:1–e7731:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Teng:2023:LMW

[TZZ^{+}23] Fei Teng, Rui Zhu, Yabo Zhou, Maoru Chi, and Haibo Zhang. A lightweight model of wheel-rail force inversion for railway vehicles. Concurrency and Computation: Practice and Experience, 35(14):e6443:1–e6443:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Umer:2022:IFC

[UAS^{+}22] Muhammad Junaid Umer, Javeria Amin, Muhammad Sharif, Muhammad Almas Anjum, Faisal Azam, and Jamal Hussain Shah. An integrated framework for COVID-19 classification based on classical and quantum transfer learning from a chest radiograph. Concurrency and Computation: Practice and Experience, 34(20):e6434:1–e6434:??, September 10, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ucan:2021:EAS

[UDS21] Alaettin Uçan, Murat Dörterler, and Ebru Akçapinar Sezer. An emotion analysis scheme based on Gray Wolf optimization and deep learning. Concurrency and Computation: Practice and Experience, 33(13):e6204:1–e6204:??, July 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Uma:2022:MFD

[UE22] S. Uma and R. Eswari. A multimodel fire detection and alarm system for automobiles using internet of things and machine learning. Concurrency and Computation: Practice and Experience, 34(21):e7117:1–e7117:??, September 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Uyar:2022:TCP

[UGK^{+}22] Ahmet Uyar, Gurhan Gunduz, Supun Kamburugamuve, Pulasthi Wickramasinghe, Chathura Widanage, Kannan Govindarajan, Niranda Perera, Vibhatha Abeykoon, Selahattin Akkas, and Geoffrey Fox. Twister2 cross-platform resource scheduler for big data. Concurrency and Computation: Practice and Experience, 34(9):e6502:1–e6502:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

ulHaque:2023:CDR

[uHAU23] Hafiz Mahfooz ul Haque, Salwa Muhammad Akhtar, and Ijaz Uddin. Contextual defeasible reasoning framework for heterogeneous knowledge sources. Concurrency and Computation: Practice and Experience, 35(15):e6446:1–e6446:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

UlIslam:2023:CAA

[UK23] Mir Salim Ul Islam and Ashok Kumar. Context-aware application scheduling in fog computing environment. Concurrency and Computation: Practice and Experience, 35(26):e7836:1– e7836:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ucak:2022:SPB

[UKS22] Ege Uçak, Elif Karagümüs, and Cevat Sener. A scalable platform for big data analysis in public transport. Concurrency and Computation: Practice and Experience, 34(9):e6534:1– e6534:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Umer:2023:PMB

[UMR23] Adnan Umer, Adnan Noor Mian, and Omer Rana. Predicting machine behavior from Google cluster workload traces. Concurrency and Computation: Practice and Experience, 35 (5):e7559:1–e7559:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Uzun-Per:2022:SRS

[UPGCA22] Meryem Uzun-Per, Ahmet Volkan Gurel, Ali Burak Can, and Mehmet S. Aktas. Scalable recommendation systems based on finding similar items and sequences. Concurrency and Computation: Practice and Experience, 34(20):e6841:1–e6841:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Upreti:2022:APL

[URK^{+}22] Ramesh Upreti, Ashish Rauniyar, Jeevan Kunwar, Hårek Haugerud, Paal Engelstad, and Anis Yazidi. Adaptive pursuit learning for energy-efficient target coverage in wireless sensor networks. Concurrency and Computation: Practice and Experience, 34(7):e5975:1–e5975:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Uysal:2021:GPU

[USI21] Furkan Uysal, Rifat Sonmez, and Selcuk Kursat Isleyen. A graphical processing unit-based parallel hybrid genetic algorithm for resource-constrained multi-project scheduling problem. Concurrency and Computation: Practice and Experience, 33(16):e6266:1–e6266:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haque:2023:MBD

[USK23] Hafiz Mahfooz Ul Haque, Kiran Saleem, and Ahmad Salman Khan. Modeling belief-desire-intention reasoning agents for situation-aware formalisms. Concurrency and Computation: Practice and Experience, 35(15):e6417:1–e6417:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Uller:2023:LML

[USP^{+}23] João Fellipe Uller, João Vicente Souto, Pedro Henrique Penna, Márcio Castro, Henrique Freitas, and Jean-François Méhaut. LWMPI: an MPI library for NoC-based lightweight manycore processors with on-chip memory constraints. Concurrency and Computation: Practice and Experience, 35(17): e6693:1–e6693:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ullah:2021:SPD

[UWF^{+}21] Farhan Ullah, Junfeng Wang, Muhammad Farhan, Masood Habib, and Shehzad Khalid. Software plagiarism detection in multiprogramming languages using machine learning approach. Concurrency and Computation: Practice and Experience, 33(4):e5000:1–e5000:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Uguz:2022:EPT

[UYÖ^{+}22] Sezer Uguz, Mete Yaganoglu, Baris Özyer, Gülsah Tümüklü Özyer, and Gül Tokdemir. The effect of population and tourism factors on Covid-19 cases in Italy: Visual data analysis and forecasting approach. Concurrency and Computation: Practice and Experience, 34(6):e6774:1–e6774:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haque:2021:CAF

[UZAA21] Hafiz Mahfooz Ul Haque, Haidar Zulfiqar, Abrar Ahmed, and Yasir Ali. A context-aware framework for modelling and verification of smart parking systems in urban cities. Concurrency and Computation: Practice and Experience, 33(2):e5401:1– e5401:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

uzZaman:2020:POC

[uZKH^{+}20] S. Khaliq uz Zaman, Iftikhar Ahmed Khan, Syed Sajid Hussain, Tassawar Iqbal, Juanid Shuja, Syed Faraz Ahmed, Yaser Jararweh, and Kwangman Ko. PreDiKT-OnOff: a complex adaptive approach to study the impact of digital social networks on Pakistani students’ personal and social life. Concurrency and Computation: Practice and Experience, 32(21): e5121:1–e5121:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vultureanu-Albisi:2022:SEA

[VAB22] Alexandra Vultureanu-Albisi and Costin Badica. A survey on effects of adding explanations to recommender systems. Concurrency and Computation: Practice and Experience, 34(20): e6834:1–e6834:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vigil:2021:DPB

[VB21] M. S. Antony Vigil and V. Subbiah Bharathi. Detection of periodontal bone loss in mandibular area from dental panoramic radiograph using image processing techniques. Concurrency and Computation: Practice and Experience, 33(17):e6323:1– e6323:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Verma:2022:SLA

[VB22a] Upendra Verma and Diwakar Bhardwaj. A secure lightweight anonymous elliptic curve cryptography-based authentication and key agreement scheme for fog assisted-Internet of Things enabled networks. Concurrency and Computation: Practice and Experience, 34(23):e7172:1–e7172:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vhatkar:2022:SIM

[VB22b] Kapil Netaji Vhatkar and Girish P. Bhole. Self-improved moth flame for optimal container resource allocation in cloud. Concurrency and Computation: Practice and Experience, 34 (23):e7200:1–e7200:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Verma:2024:EAS

[VB24] Shveta Verma and Anju Bala. Efficient auto-scaling for host load prediction through VM migration in cloud. Concurrency and Computation: Practice and Experience, 36(4):e7925:1– e7925:??, February 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vijayarangam:2021:ESP

[VBM^{+}21] S. Vijayarangam, Gokulnath Chandra Babu, S. Ananda Murugan, N. Kalpana, and Priyan Malarvizhi Kumar. Enhancing the security and performance of nodes in Internet of vehicles. Concurrency and Computation: Practice and Experience, 33 (7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vergara:2024:STF

[VBP^{+}24] Verónica G. Melesse Vergara, Reuben Budiardja, Paul Peltz, Jeffery Niles, Christopher Zimmer, Daniel Dietz, Christopher Fuson, Hong Liu, Paul Newman, James Simmons, and Christopher Muzyn. A step towards the final frontier: Lessons learned from acceptance testing of the first HPE/Cray EX 3000 system at ORNL. Concurrency and Computation: Practice and Experience, 36(3):e7914:1–e7914:??, February 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vinh:2020:PBT

[VCBB20] Thinh Le Vinh, Hervé Cagnon, Samia Bouzefrane, and Soumya Banerjee. Property-based token attestation in mobile computing. Concurrency and Computation: Practice and Experience, 32(1):e4350:1–e4350:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vasconcellos:2020:ASC

[VCFZ20] Eduardo C. Vasconcellos, Esteban W. G. Clua, Flavio H. Fenton, and Marcelo Zamith. Accelerating simulations of cardiac electrical dynamics through a multi-GPU platform and an optimized data structure. Concurrency and Computation: Practice and Experience, 32(5):e5528:1–e5528:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vekaria:2021:RSC

[VCS^{+}21] Komal Vekaria, Prasad Calyam, Sai Swathi Sivarathri, Songjie Wang, Yuanxun Zhang, Ashish Pandey, Cong Chen, Dong Xu, Trupti Joshi, and Satish Nair. Recommenderas-a-service with chatbot guided domain-science knowledge discovery in a science gateway. Concurrency and Computation: Practice and Experience, 33(19):e6080:1–e6080:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vetha:2021:TBH

[VD21] S. Vetha and K. Vimala Devi. A trust-based hypervisor framework for preventing DDoS attacks in cloud. Concurrency and Computation: Practice and Experience, 33(3):e5279:1– e5279:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vangala:2023:PSS

[VDL23] Anusha Vangala, Ashok Kumar Das, and Jong-Hyouk Lee. Provably secure signature-based anonymous user authentication protocol in an Internet of Things-enabled intelligent precision agricultural environment. Concurrency and Computation: Practice and Experience, 35(16):e6187:1–e6187:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanderSchaaf:2021:FBA

[vdSTC21] Koen van der Schaaf, Bedir Tekinerdogan, and Cagatay Catal. A feature-based approach for guiding the selection of Internet of Things cybersecurity standards using text mining. Concurrency and Computation: Practice and Experience, 33(21): e6385:1–e6385:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vizier:2020:CBB

[VG20] Guillaume Vizier and Vincent Gramoli. ComChain: a blockchain with Byzantine fault-tolerant reconfiguration. Concurrency and Computation: Practice and Experience, 32 (12):e5494:1–e5494:??, June 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vivekanandan:2021:HHH

[VG21] Nivethitha Vivekanandan and Aghila Gnanasekaran. Hybrid Harris Hawk–Salp swarm optimization algorithm-based integrated optimal data placement and task scheduling for improving the user experience in edge computing. Concurrency and Computation: Practice and Experience, 33(24):e6455:1– e6455:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vogel:2022:SAP

[VGDF22] Adriano Vogel, Dalvan Griebler, Marco Danelutto, and Luiz Gustavo Fernandes. Self-adaptation on parallel stream processing: a systematic review. Concurrency and Computation: Practice and Experience, 34(6):e6759:1–e6759:??, March 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vairagade:2022:SIT

[VH22] Rupali Sachin Vairagade and Brahmananda Savadatti Hanumantha. Secure Internet of Things network using lightweighted trust and blockchain-powered PoW framework. Concurrency and Computation: Practice and Experience, 34(21):e7057:1–e7057:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vinh:2021:SIC

[Vin21] Phan Cong Vinh. Special issue on context-aware computing: Theory and applications. Concurrency and Computation: Practice and Experience, 33(2):e5824:1–e5824:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Varghese:2022:ORB

[VJ22] Meble Varghese and M. Victor Jose. An optimized radial bias function neural network for intrusion detection of distributed denial of service attack in the cloud. Concurrency and Computation: Practice and Experience, 34(27):e7321:1–e7321:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Verma:2023:NQK

[VK23a] Garima Verma and Ashutosh Kumar. Novel quantum key distribution and attribute based encryption for cloud data security. Concurrency and Computation: Practice and Experience, 35(21):e7700:1–e7700:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vinayaki:2023:EES

[VK23b] V. Desika Vinayaki and R. Kalaiselvi. ESLO: Enhanced sea lion optimization based bi-directional CNN-RNN for accurate detection of diabetic retinopathy. Concurrency and Computation: Practice and Experience, 35(1):e7391:1–e7391:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Karthikeyan:2024:NML

[VK24] Karthikeyan V and Murugan K. A novel machine learningbased classification approach to prevent flow table overflow attack in Software-Defined Networking. Concurrency and Computation: Practice and Experience, 36(1):e7878:1–e7878:??, January 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vibhute:2024:LBN

[VKK^{+}24] Amol D. Vibhute, Minhaj Khan, Anuradha Kanade, Chandrashekhar H. Patil, Sandeep V. Gaikwad, Kanubhai K. Patel, and Jatinderkumar R. Saini. An LSTM-based novel nearreal-time multiclass network intrusion detection system for complex cloud environments. Concurrency and Computation: Practice and Experience, 36(11):e8024:1–e8024:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

V:2023:SAB

[VKSS23] Brindha Devi V., Johny Elma K., Rooban S., and Francis H. Shajin. Self-attention based progressive generative adversarial network optimized with arithmetic optimization algorithm for kidney stone detection. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Venkatesan:2022:CDD

[VLVS22] Muthulakshmi Venkatesan, Priya Lakshmipathy, Vani Vijayan, and Ramesh Sundar. Cardiac disease diagnosis using feature extraction and machine learning based classification with Internet of Things(IoT). Concurrency and Computation: Practice and Experience, 34(4):e6622:1–e6622:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Venkataraman:2023:FRF

[VM23] Jayalakshmi Venkataraman and Lakshmi Mohandoss. FBO-RNN: Fuzzy butterfly optimization-based RNN-LSTM for extracting sentiments from Twitter Emoji database. Concurrency and Computation: Practice and Experience, 35(12): e7683:1–e7683:??, May 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vizcaino:2023:ALV

[VMFL23] Pablo Vizcaino, Filippo Mantovani, Roger Ferrer, and Jesus Labarta. Acceleration with long vector architectures: Implementation and evaluation of the FFT kernel on NEC SX-Aurora and RISC-V vector extension. Concurrency and Computation: Practice and Experience, 35(20):e7424:1–e7424:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vilela:2023:BIO

[VNP^{+}23] Ricardo Ferreira Vilela, João Choma Neto, Victor Hugo Santiago Costa Pinto, Paulo Sérgio Lopes de Souza, and Simone do Rocio Senger de Souza. Bio-inspired optimization to support the test data generation of concurrent software. Concurrency and Computation: Practice and Experience, 35 (2):e7489:1–e7489:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Verma:2022:AHA

[VP22a] Pooja Verma and Raghav Prasad Parouha. An advanced hybrid algorithm for nonlinear function optimization with real world applications. Concurrency and Computation: Practice and Experience, 34(3):e6551:1–e6551:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vignesh:2022:BBS

[VP22b] Ramamoorthi Vignesh and Kulasekaran Mohana Prasad. Blockchain-based security enhancement mechanism for employee performance assessment system. Concurrency and Computation: Practice and Experience, 34(27):e7318:1– e7318:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Verner:2023:BOD

[VPB^{+}23] Eric Verner, Helen Petropoulos, Bradley Baker, Henry Jeremy Bockholt, Jill Fries, Anastasia Bohsali, Rajikha Raja, Duc Hoai Trinh, and Vince Calhoun. BrainForge: an online data analysis platform for integrative neuroimaging acquisition, analysis, and sharing. Concurrency and Computation: Practice and Experience, 35(18):e6855:1–e6855:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Velusamy:2023:MNH

[VPGK23] Praveen Velusamy, Keerthika Periasamy, Sivapriya Gurusamy, and Sangeetha Kuppusamy. Modeling a novel heuristic model for handling vehicle routing using multiobjective optimization. Concurrency and Computation: Practice and Experience, 35(21):e7694:1–e7694:??, September 25, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vu:2022:BST

[VPQ22] Tuan-Anh Vu and Cuong Pham-Quoc. Building a smart traffic light system based on Internet of Things using π-calculus. Concurrency and Computation: Practice and Experience, 34 (10):e6731:1–e6731:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Velliangiri:2022:MRA

[VPSM22] S. Velliangiri, Shanthini Pandiaraj, Iwin Thanakumar Joseph S, and S. Muthubalaji. Multiclass recognition of AD neurological diseases using a bag of deep reduced features coupled with gradient descent optimized twin support vector machine classifier for early diagnosis. Concurrency and Computation: Practice and Experience, 34(21):e7099:1–e7099:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vijayalakshmi:2022:HBD

[VR22] J. Vijayalakshmi and E. Ramaraj. A Hadoop-big data analytic model to predict and classify chronic kidney diseases using improved fractional rough fuzzy K-means clustering and extreme gradient boost rat swarm optimizer. Concurrency and Computation: Practice and Experience, 34(28):e7354:1– e7354:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vysocky:2021:AIP

[VRB21] Ondrej Vysocký, Lubomír Ríha, and Andrea Bartolini. Application instrumentation for performance analysis and tuning with focus on energy efficiency. Concurrency and Computation: Practice and Experience, 33(11):e5966:1–e5966:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vadlamudi:2022:ALT

[VRR^{+}22] Sree Harsha Vadlamudi, Yerrabapu Sai Souhith Reddy, Polu Ajith Sai Kumar Reddy, Prakasam Periasamy, and Noor Mohammed Vali Mohamad. Automatic liver tumor segmentation and identification using fully connected convolutional neural network from CT images. Concurrency and Computation: Practice and Experience, 34(24):e7212:1–e7212:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Venugopal:2022:IBC

[VRS22] Sabaresan Venugopal, Ramesh Rengaswamy, and Godfrey Winster Sathianesan. IoT based cyber forensics in big data optimization and privacy using deep neural anomaly detection with Hadoop clustering and convolution based Adam optimizer. Concurrency and Computation: Practice and Experience, 34(11):e6881:1–e6881:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Varma:2023:EES

[VRV23] P. Ravi Kiran Varma, Sathiya R R, and M. Vanitha. Enhanced Elman spike neural network based intrusion attack detection in software defined Internet of Things network. Concurrency and Computation: Practice and Experience, 35 (2):e7503:1–e7503:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vidhyalakshmi:2021:TDN

[VS21] M. Vidhyalakshmi and S. Sudha. Text detection in natural images with hybrid stroke feature transform and high performance deep Convnet computing. Concurrency and Computation: Practice and Experience, 33(3):e5271:1–e5271:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Varalakshmi:2022:OSM

[VS22a] Perumal Varalakshmi and Sankari Subbiah. Optimized scheduling of multi-user map-reduce jobs in heterogeneous environment. Concurrency and Computation: Practice and Experience, 34(27):e7316:1–e7316:??, December 10, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Venkatesh:2022:MSC

[VS22b] Maganti Venkatesh and S. Sathyalaksmi. Memetic swarm clustering with deep belief network model for e-learning recommendation system to improve learning performance. Concurrency and Computation: Practice and Experience, 34(18): e7010:1–e7010:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vishwakarma:2022:CEM

[VS22c] Gajendra Vishwakarma and Abhishek Singh. Computing the effect of measurement errors on ranked set sampling estimators of the population mean. Concurrency and Computation: Practice and Experience, 34(27):e7333:1–e7333:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vishwakarma:2022:CFE

[VSK22] Gajendra K. Vishwakarma, Neha Singh, and Neelesh Kumar. A computational framework for estimation of mean in presence of observational error. Concurrency and Computation: Practice and Experience, 34(11):e6842:1–e6842:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Verma:2022:MLI

[VSP22] Navneet Verma, Sukhdip Singh, and Devendra Prasad. Machine learning and IoT-based model for patient monitoring and early prediction of diabetes. Concurrency and Computation: Practice and Experience, 34(24):e7219:1–e7219:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Venugopal:2023:WIM

[VSS23] Madamanchi Venugopal, Virendra K. Sharma, and Kalpana Sharma. Web information mining and semantic analysis in heterogeneous unstructured text data using enhanced latent Dirichlet allocation. Concurrency and Computation: Practice and Experience, 35(1):e7410:1–e7410:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vijayakumar:2023:TIM

[VV23] Angel Petricia Vijayakumar and Mary Anita Rajam V. Temporal intensity modeling of sub-events in microblogs. Concurrency and Computation: Practice and Experience, 35(3): e7473:1–e7473:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vooturi:2023:RBG

[VVK23] Dharma Teja Vooturi, Girish Varma, and Kishore Kothapalli. Ramanujan bipartite graph products for efficient block sparse neural networks. Concurrency and Computation: Practice and Experience, 35(14):e6363:1–e6363:??, June 25, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Varchagall:2023:EDL

[VY23] Manjunath Varchagall and Prasad Adaguru Yogegowda. Early detection of liver disorders using hybrid soft computing techniques for optimal feature selection and classification. Concurrency and Computation: Practice and Experience, 35 (6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Valentine:2021:PED

[VZR^{+}21] David Valentine, Ilya Zaslavsky, Stephen Richard, Ouida Meier, Gary Hudman, Bernhard Peucker-Ehrenbrink, and

Karen Stocks. EarthCube Data Discovery Studio: a gate-

way into geoscience data discovery and exploration with Jupyter notebooks. Concurrency and Computation: Practice and Experience, 33(19):e6086:1–e6086:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:NSG

[Wan20] Junyan Wang. Nondominated sorting genetic algorithm II with merged strategies for industrial network topology optimization. Concurrency and Computation: Practice and Experience, 32(19):e5768:1–e5768:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wan:2022:THR

[Wan22a] Shaohua Wan. Topology hiding routing based on learning with errors. Concurrency and Computation: Practice and Experience, 34(14):e5740:1–e5740:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:WOP

[Wan22b] Jin Wang. Workflow offloading with privacy preservation in a cloud-edge environment. Concurrency and Computation:

Practice and Experience, 34(18):e7002:1–e7002:??, August 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2024:RAB

[Wan24] Qi Wang. Research on the application of bilateral collaborative algorithm in hotel tourism consumption preference recommendation. Concurrency and Computation: Practice and Experience, 36(11):e8028:1–e8028:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:KLO

[WAY^{+}21] Kefan Wang, Jing An, Zibo Yu, Xingshu Yin, and Chao Ma. Kernel local outlier factor-based fuzzy support vector machine for imbalanced classification. Concurrency and Computation: Practice and Experience, 33(13):e6235:1–e6235:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:MGC

[WBL22] Zhanzhong Wang, Xuchao Bai, and Jiajun Li. Modified generalized confidence interval for the stress-strength reliability from exponentiated Weibull distribution. Concurrency and Computation: Practice and Experience, 34(15):e6969:1– e6969:??, July 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:AAD

[WBZ21] Ji Wang, Weidong Bao, and Xiaomin Zhu. ADAPT: Adaptive distributed optimization approach for uploading data with redundancy in cooperative mobile cloud. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wyrzykowski:2022:ASD

[WC22] Roman Wyrzykowski and Florina M. Ciorba. Algorithmic and software development advances for next-generation heterogeneous platforms. Concurrency and Computation: Practice and Experience, 34(14):e7013:1–e7013:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wen:2020:PEI

[WCCC20] Yean-Fu Wen, Yu-Fang Chen, Tse Kai Chiu, and Yen-Chou Chen. Performance enhancement for iterative data computing with in-memory concurrent processing. Concurrency and Computation: Practice and Experience, 32(7):e5593:1– e5593:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wan:2024:HDB

[WCL^{+}24] Lanjun Wan, Xueyan Cui, Yuanyuan Li, Weihua Zheng, and Xinpan Yuan. HeteroPP: a directive-based heterogeneous cooperative parallel programming framework. Concurrency and Computation: Practice and Experience, 36(11):e8014:1– e8014:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2020:CMS

[WCSG20] Hao Wu, Xin Chen, Xiaoyu Song, and He Guo. Cost minimization of scheduling scientific workflow applications on clouds. Concurrency and Computation: Practice and Experience, 32(5):e5503:1–e5503:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:ESS

[WCWG20] Xiaoyan Wang, Xingshu Chen, Yitong Wang, and Long Ge. An efficient scheme for SDN state consistency verification in cloud computing environment. Concurrency and Computation: Practice and Experience, 32(2):e5440:1–e5440:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2021:LPA

[WCWG21] Qingshou Wu, Rongwang Chen, Lijin Wang, and Kun Guo. A label propagation algorithm for community detection on high-mixed networks. Concurrency and Computation: Practice and Experience, 33(9):e6141:1–e6141:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2023:LPC

[WCZ^{+}23] Hao Wu, Yuqi Chen, Chi Zhang, Jiangchao Dong, and Yuxin Wang. Loads prediction and consolidation of virtual machines in cloud. Concurrency and Computation: Practice and Experience, 35(23):e7760:1–e7760:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2022:PPC

[WDW^{+}22] Jiashu Wu, Hao Dai, Yang Wang, Shigen Shen, and Chengzhong Xu. PECCO: a profit and cost-oriented computation offloading scheme in edge-cloud environment with improved moth-flame optimization. Concurrency and Computation: Practice and Experience, 34(22):e7163:1–e7163:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:BCS

[WEH^{+}22] Song Wang, Haroon Elahi, Yanzhu Hu, Yixin Zhang, and Jiao Wang. A botnets control strategy based on variable forgetting rate of control commands. Concurrency and Computation: Practice and Experience, 34(7):e6118:1–e6118:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:MLF

[WER^{+}23] Chenxi Wang, Chen E, Mengli Ren, Lei Guo, Xiehua Yu, Yaojin Lin, and Shaozi Li. Multi-label feature selection based on relative entropy and fuzzy neighborhood mutual discrimination index. Concurrency and Computation: Practice and Experience, 35(27):e7845:1–e7845:??, December 10, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:SAD

[WFHC21] Xianghui Wang, Zhiyong Feng, Keman Huang, and Shizhan Chen. Self-adaptation and distributed knowledge-based service ecosystem evolution. Concurrency and Computation: Practice and Experience, 33(24):e6469:1–e6469:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:CER

[WFY^{+}22a] Chunzhi Wang, Yanlin Fu, Junyi Yan, Xinyun Wu, Yucheng Zhang, Huiling Xia, and Ye Yuan. A cost-efficient resemblance detection scheme for post-deduplication delta compression in backup systems. Concurrency and Computation:

Practice and Experience, 34(3):e6558:1–e6558:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:TCR

[WFY22b] Zhaorui Wang, Guisheng Fan, and Huiqun Yu. Timeconstrained and reliability aware energy minimization scheduling algorithm for heterogeneous multiprocessor environments. Concurrency and Computation: Practice and Experience, 34(26):e7306:1–e7306:??, November 30, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wilhelm:2023:CMA

[WGDP23] Martin Wilhelm, Hanna Geppert, Anna Drewes, and Thilo Pionteck. A comprehensive modeling approach for the task mapping problem in heterogeneous systems with dataflow processing units. Concurrency and Computation: Practice and Experience, 35(25):e7909:1–e7909:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:KAB

[WGLL23] Bo Wang, Yina Guo, Hongtao Li, and Zhiying Li. kanonymity based location privacy protection method for location-based services in Internet of Things. Concurrency and Computation: Practice and Experience, 35(20):e6760:1– e6760:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gao:2022:PCA

[wGTC22] Shi wei Gao, Ran Tian, and Peng Chen. Principal component analysis for process monitoring in distributed system environment. Concurrency and Computation: Practice and Experience, 34(8):e5309:1–e5309:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:OVR

[WGY20] Xiaohui Wang, Haoran Gu, and YuXian Yue. The optimization of virtual resource allocation in cloud computing based on RBPSO. Concurrency and Computation: Practice and Experience, 32(16):e5113:1–e5113:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:CAE

[WGYZ22] Zhihui Wang, Liheng Gong, Jingjing Yang, and Xiao Zhang. Cloud-assisted elliptic curve password authenticated key exchange protocol for wearable healthcare monitoring system. Concurrency and Computation: Practice and Experience, 34 (9):e5734:1–e5734:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:ULA

[WGZ^{+}20] Zhibo Wang, Yuechuan Guo, Senzhe Zheng, Wei Xu, Lin Liu, Zixin Liu, and Xiaohui Cui. Users’ location analysis based on Chinese mobile social media. Concurrency and Computation: Practice and Experience, 32(13):e4669:1–e4669:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:GEC

[WHC^{+}20] Penghong Wang, Jianrou Huang, Zhihua Cui, Liping Xie, and Jinjun Chen. A Gaussian error correction multi-objective positioning model with NSGA-II. Concurrency and Computation: Practice and Experience, 32(5):e5464:1–e5464:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2022:MLD

[WHDS22] Amin Wu, Fouzi Harrou, Abdelkader Dairi, and Ying Sun. Machine learning and deep learning-driven methods for predicting ambient particulate matters levels: a case study. Concurrency and Computation: Practice and Experience, 34(19): e7035:1–e7035:??, August 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wasti:2020:ASS

[WHH^{+}20] Shahbaz Hassan Wasti, Muhammad Jawad Hussain, Guangjian Huang, Aftab Akram, Yuncheng Jiang, and Yong Tang. Assessing semantic similarity between concepts: a weightedfeature-based approach. Concurrency and Computation: Practice and Experience, 32(7):e5594:1–e5594:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:APA

[WHH^{+}23] Jinyong Wang, Zhiqiu Huang, Xiaowei Huang, Tiexin Wang, Guohua Shen, and Jian Xie. An accident prediction architecture based on spatio-clock stochastic and hybrid model for autonomous driving safety. Concurrency and Computation: Practice and Experience, 35(17):e6550:1–e6550:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:CUS

[WHJ^{+}20] Yuling Wang, Wei He, Minghu Jiang, Yunlong Huang, and Peijun Qiu. CHOpinionMiner: an unsupervised system for Chinese opinion target extraction. Concurrency and Computation: Practice and Experience, 32(7):e5582:1–e5582:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:CID

[WHZL21] Liping Wang, Wenzhun Huang, Shanwen Zhang, and Zhe Liu. Complex image denoising framework with CNN-wavelet under concurrency scenarios for informatics systems. Concurrency and Computation: Practice and Experience, 33(12): e5059:1–e5059:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:R

[WHZL23] Liping Wang, Wenzhun Huang, Shanwen Zhang, and Zhe Liu. Retraction. Concurrency and Computation: Practice and Experience, 33(12):e7750:1–e7750:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2022:DCA

[WJD22] Chenyang Wu, Xiaopeng Ji, and Yuewei Dai. Dynamic channel allocation strategy for quality of service enhancement in wireless ad hoc networks. Concurrency and Computation: Practice and Experience, 34(10):e6777:1–e6777:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:GGA

[WJJ23] Jing Wang, Haoxiang Jie, and Yue Jiang. Group-guided artificial bee colony algorithm with elastic adjustment strategy.

Concurrency and Computation: Practice and Experience, 35 (5):e7560:1–e7560:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:ACL

[WJL^{+}20] Xiancheng Wang, Zhangwei Jiang, Wei Li, Roozbeh Zarei, Guangyan Huang, Anwaar Ulhaq, Xiaoxia Yin, Bailing Zhang, Peng Shi, Mengjiao Guo, and Jing He. Active contours with local and global energy based-on fuzzy clustering and maximum a posterior probability for retinal vessel detection. Concurrency and Computation: Practice and Experience, 32(7):e5599:1–e5599:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:PWL

[WJLC21] Jianjun Wang, Zijie Jiang, Fan Li, and Weiming Chen. The prediction of water level based on support vector machine under construction condition of steel sheet pile cofferdam. Concurrency and Computation: Practice and Experience, 33(5): e6003:1–e6003:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:FSI

[WJS21] Zhibo Wang, Lin Jiang, and Bilial Suman. Foreword to the special issue of the intelligent systems for the Internet of Things (ISIT2018). Concurrency and Computation: Practice and Experience, 33(11):e6142:1–e6142:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

White:2020:OPP

[WK20] Sam White and Laxmikant V. Kale. Optimizing point-to-point communication between adaptive MPI endpoints in shared memory. Concurrency and Computation: Practice and Experience, 32(3):e4467:1–e4467:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2022:APB

[WKB^{+}22] Xingfu Wu, Michael Kruse, Prasanna Balaprakash, Hal Finkel, Paul Hovland, Valerie Taylor, and Mary Hall. Autotuning PolyBench benchmarks with LLVM Clang/Polly loop optimization pragmas using Bayesian optimization. Concurrency and Computation: Practice and Experience, 34(20): e6683:1–e6683:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Waqas:2022:BAD

[WKL^{+}22] Muhammad Waqas, Kamlesh Kumar, Asif Ali Laghari, Umair Saeed, Muhammad Malook Rind, Aftab Ahmed Shaikh, Fahad Hussain, Athaul Rai, and Abdul Qayoom Qazi. Botnet attack detection in Internet of Things devices over cloud environment via machine learning. Concurrency and Computation: Practice and Experience, 34(4):e6662:1–e6662:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:UOS

[WKY22] Honghao Wang, Majid Khayatnezhad, and Naser Youssefi. Using an optimized soil and water assessment tool by deep belief networks to evaluate the impact of land use and climate change on water resources. Concurrency and Computation: Practice and Experience, 34(10):e6807:1–e6807:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:HIN

[WL23] Yaodong Wang and Yamin Li. Hybrid interconnection networks for reducing hardware cost and improving path diversity based on fat-trees and hypercubes. Concurrency and Computation: Practice and Experience, 35(19):e7587:1– e7587:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:GRA

[WLCW20] Yupeng Wang, Tianlong Liu, Chang Choi, and Haoxiang Wang. Green resource allocation method for intelligent medical treatment-oriented service in a 5G mobile network. Concurrency and Computation: Practice and Experience, 32(1): e5057:1–e5057:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2022:BSC

[WLDW22] Rongzhi Wu, Binyuan Lv, Chaoming Dai, and Weigang Wang. Bayesian stochastic configuration networks for robust data modeling. Concurrency and Computation: Practice and Experience, 34(1):e6495:1–e6495:??, January 10, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:MAD

[WLG^{+}22] Wei Wang, Guangsong Li, Keke Gai, Yazhe Tang, Benchao Yang, and Xueming Si. Modelization and analysis of dynamic heterogeneous redundant system. Concurrency and Computation: Practice and Experience, 34(12):e6035:1–e6035:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:UCW

[WLJ20] Chao Wang, Wenjin Liu, and Giorgos Jimenez. Using chaos world cup optimization algorithm for medical images contrast enhancement. Concurrency and Computation: Practice and Experience, 32(5):e5482:1–e5482:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:SDD

[WLL21a] Liang Wang, Jiayan Liu, and Wenyuan Liu. Staged data delivery protocol: a blockchain-based two-stage protocol for non-repudiation data delivery. Concurrency and Computation: Practice and Experience, 33(13):e6240:1–e6240:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Weng:2021:WST

[WLL^{+}21b] Hui Weng, Lei Li, Huangwei Lei, Zhiming Luo, Candong Li, and Shaozi Li. A weakly supervised tooth-mark and crack detection method in tongue image. Concurrency and Computation: Practice and Experience, 33(16):e6262:1–e6262:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:USV

[WLLX21] Wei Wang, Xiangfeng Luo, Yang Li, and Shaorong Xie. Unmanned surface vessel obstacle avoidance with prior knowledge-based reward shaping. Concurrency and Computation: Practice and Experience, 33(9):e6110:1–e6110:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2020:IDA

[WLLZ20] Jinxia Wei, Chun Long, Jiawei Li, and Jing Zhao. An intrusion detection algorithm based on bag representation with ensemble support vector machine in cloud computing. Concurrency and Computation: Practice and Experience, 32(24): e5922:1–e5922:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2021:NCM

[WLX21] Gengkun Wu, Yongquan Liang, and Shaohua Xu. Numerical computational modeling of random rough sea surface based on JONSWAP spectrum and Donelan directional function. Concurrency and Computation: Practice and Experience, 33 (15):e5514:1–e5514:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2022:NRS

[WLY^{+}22] Yilin Wu, Jinghua Liu, Xiehua Yu, Yaojin Lin, and Shaozi Li. Neighborhood rough set based multi-label feature selection with label correlation. Concurrency and Computation: Practice and Experience, 34(22):e7162:1–e7162:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:OVT

[WLZ21a] Luyao Wang, Huchuan Lu, and Pingping Zhang. Online visual tracking via cross-similarity-based Siamese network. Concurrency and Computation: Practice and Experience, 34(1): e5617:1–e5617:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:ECS

[WLZ^{+}21b] Tian Wang, Jiakun Li, Mengyi Zhang, Aichun Zhu, Hichem Snoussi, and Chang Choi. An enhanced 3DCNN-ConvLSTM for spatiotemporal multimedia data analysis. Concurrency and Computation: Practice and Experience, 33(2):e5302:1– e5302:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:GPM

[WMC21] Shuangyan Wang, Gang Mei, and Salvatore Cuomo. A generic paradigm for mining human mobility patterns based on the GPS trajectory data using complex network analysis. Concurrency and Computation: Practice and Experience, 33(4): e5335:1–e5335:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:ART

[WMC^{+}23] Yuefeng Wang, Kuang Mao, Tong Chen, Yanglong Yin, Shuibing He, and Gang Chen. Accelerating real-time object detection in high-resolution video surveillance. Concurrency and Computation: Practice and Experience, 35(18):e6307:1– e6307:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:BBA

[WMQ^{+}22] Yingsen Wang, Yulan Ma, Yan Qiang, Juanjuan Zhao, Yi Li, and Keqin Li. BAC: a block alliance consensus mechanism for the mine consortium blockchain. Concurrency and Computation: Practice and Experience, 34(27):e7344:1–e7344:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:DUR

[WMS^{+}23] Yuqian Wang, JianWei Ma, Axyonov Sergey, Shaofei Zang, and Miao Zhang. DPA-UNet rectal cancer image segmentation based on visual attention. Concurrency and Computation: Practice and Experience, 35(21):e7670:1–e7670:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wada:2021:EIB

[WMY^{+}21] Takuma Wada, Naoki Matsumura, Ryota Yasudo, Koji Nakano, and Yasuaki Ito. Efficient implementations of Bloom filter using block RAMs and DSP slices on the FPGA. Concurrency and Computation: Practice and Experience, 33(12): e5475:1–e5475:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2021:RID

[WN21] Zhen Wei and Jinghuan Nie. Research on intelligent design mechanism of landscape lamp with regional cultural value based on interactive genetic algorithm. Concurrency and Computation: Practice and Experience, 33(16):e6273:1– e6273:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2023:RCF

[WN23] Zhen Wei and Jinghuan Nie. Research on color fusion model of historic and cultural blocks in Shanghai based on deep learning algorithm — take Tianzifang as an example. Concurrency and Computation: Practice and Experience, 35(4):e7530:1– e7530:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2024:ILD

[WNZ24] Zhen Wei, Jinghuan Nie, and Wei Zhang. Intelligent lighting design of cloud computing platform based on digital twin historic landscape blocks — take Yu Garden commercial street in Shanghai as an example. Concurrency and Computation: Practice and Experience, 36(7):e7975:1–e7975:??, March 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wickramasinghe:2022:HPI

[WPK^{+}22] Pulasthi Wickramasinghe, Niranda Perera, Supun Kamburugamuve, Kannan Govindarajan, Vibhatha Abeykoon, Chathura Widanage, Ahmet Uyar, Gurhan Gunduz, Selahattin Akkas, and Geoffrey Fox. High-performance iterative dataflow abstractions in Twister2:TSet. Concurrency and Computation: Practice and Experience, 34(12):e5998:1– e5998:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2020:PGN

[WPL20] Xinzhe Wu, Serge G. Petiton, and Yutong Lu. A parallel generator of non-Hermitian matrices computed from given spectra. Concurrency and Computation: Practice and Experience, 32(20):e5710:1–e5710:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:RCR

[WQS^{+}23] Zhi Wang, Yongfei Qin, Yan Shi, Ming Jiang, and Weigang Wang. Research on e-commerce recommendation system based on matrix factorization algorithm. Concurrency and Computation: Practice and Experience, 35(22):e7739:1– e7739:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:LFL

[WQY^{+}22] Jiagang Wang, Tu Qian, Anbang Yang, Hui Wang, and Jiangbo Qian. LSR-forest: an locality sensitive hashingbased approximate k-nearest neighbor query algorithm on high-dimensional uncertain data. Concurrency and Computation: Practice and Experience, 34(8):e5795:1–e5795:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:OFS

[WRE^{+}23] Chenxi Wang, Mengli Ren, Chen E, Lei Guo, Xiehua Yu, Yaojin Lin, and Shaozi Li. Online feature selection for hierarchical classification learning based on improved ReliefF. Concurrency and Computation: Practice and Experience, 35(27): e7844:1–e7844:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wright:2022:SIP

[Wri22] Steven A. Wright. Special issue on performance modeling, benchmarking, and simulation of high performance computing systems. Concurrency and Computation: Practice and Experience, 34(20):e7165:1–e7165:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wakrime:2020:CSC

[WRJ20] Abderrahim Ait Wakrime, Mouna Rekik, and Said Jabbour. Cloud service composition using minimal unsatisfiability and genetic algorithm. Concurrency and Computation: Practice and Experience, 32(15):e5282:1–e5282:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wyrzykowski:2021:TNG

[WS21] Roman Wyrzykowski and Boleslaw K. Szymanski. Taming next-generation HPC systems: Run-time system and algorithmic advancements. Concurrency and Computation: Practice and Experience, 33(11):e6153:1–e6153:??, June 10, 2021.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wyrzykowski:2024:AEC

[WS24] Roman Wyrzykowski and Boleslaw K. Szymanski. Advances into exascale computing. Concurrency and Computation: Practice and Experience, 36(11):e8040:1–e8040:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Weng:2021:ERT

[WSJ^{+}21] Yaoqing Weng, Ying Sun, Du Jiang, Bo Tao, Ying Liu, Juntong Yun, and Dalin Zhou. Enhancement of real-time grasp detection by cascaded deep convolutional neural networks. Concurrency and Computation: Practice and Experience, 33 (5):e5976:1–e5976:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2020:TCG

[WSL^{+}20] Haibo Wu, Shiliang Shi, Yi Lu, Yong Liu, and Weihong Huang. Top corner gas concentration prediction using tdistributed Stochastic Neighbor Embedding and Support Vector Regression algorithms. Concurrency and Computation: Practice and Experience, 32(14):e5705:1–e5705:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:SSA

[WSL21] Wenru Wang, Xiaohua Shi, and Xinghai Lu. A safety simulation analysis algorithm for Altarica language. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:CMB

[WSM^{+}20] Shuihua Wang, Junding Sun, Irfan Mehmood, Chichun Pan, Yi Chen, and Yu-Dong Zhang. Cerebral micro-bleeding identification based on a nine-layer convolutional neural network with stochastic pooling. Concurrency and Computation: Practice and Experience, 32(1):e5130:1–e5130:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wittwer:2024:AXR

[WSM^{+}24] Felix Wittwer, Nicholas K. Sauter, Derek Mendez, Billy K. Poon, Aaron S. Brewster, James M. Holton, Michael E. Wall, William E. Hart, Deborah J. Bard, and Johannes P. Blaschke. Accelerating X-ray tracing for exascale systems using Kokkos. Concurrency and Computation: Practice and Experience, 36 (5):e7944:1–e7944:??, February 29, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Song:2023:RFB

[wSYyCsD23] Ren wang Song, Lei Yang, Lin ying Chen, and Zeng shou Dong. Research on frequency band sparse optimization algorithm in health index of rolling bearing. Concurrency and Computation: Practice and Experience, 35(4):e7544:1– e7544:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:FDN

[WSZ^{+}23] Linna Wang, Yunfei Song, Yujia Zhu, Daoxun Xia, and Guoquan Han. A framework for deep neural network multiuser authorization based on channel pruning. Concurrency and Computation: Practice and Experience, 35(21):e7708:1–e7708:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2023:UEL

[WT23] Xingfu Wu and Valerie Taylor. Utilizing ensemble learning for performance and power modeling and improvement of parallel cancer deep learning CANDLE benchmarks. Concurrency and Computation: Practice and Experience, 35(15):e6516:1– e6516:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2023:PPM

[WTL23] Xingfu Wu, Valerie Taylor, and Zhiling Lan. Performance and power modeling and prediction using MuMMI and 10 machine learning methods. Concurrency and Computation: Practice and Experience, 35(15):e7254:1–e7254:??, July 10, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2022:ESI

[Wu22] Chase Wu. Editorial for special issue on “Workflows in support of large-scale science”. Concurrency and Computation: Practice and Experience, 34(14):e6924:1–e6924:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:NMT

[WWA22] Bingxian Wang, Peng Wu, and Maryam Arefzaeh. A new method for task scheduling in fog-based medical healthcare systems using a hybrid nature-inspired algorithm. Concurrency and Computation: Practice and Experience, 34(22): e7155:1–e7155:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2022:TFT

[WWF^{+}22] Jiashu Wu, Yang Wang, Xiaopeng Fan, Kejiang Ye, and Chengzhong Xu. Toward fast theta-join: a prefiltering and amalgamated partitioning approach. Concurrency and Computation: Practice and Experience, 34(17):e6996:1–e6996:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:CFA

[WWG^{+}20] Feiran Wang, Yiping Wen, Tianhang Guo, Jianxun Liu, and Buqing Cao. Collaborative filtering and association rule mining-based market basket recommendation on Spark. Concurrency and Computation: Practice and Experience, 32(7): e5565:1–e5565:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:FHG

[WWG21] Haizhen Wang, Xuerui Wu, and Fatemeh Gholinia. Forecasting hydropower generation by GFDL-CM3 climate model and hybrid hydrological-Elman neural network model based on Improved Sparrow Search Algorithm (ISSA). Concurrency and Computation: Practice and Experience, 33(24):e6476:1– e6476:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2020:IRM

[WWJ^{+}20] Shuaifang Wei, Wei Wu, Gwanggil Jeon, Awais Ahmad, and Xiaomin Yang. Improving resolution of medical images with deep dense convolutional neural network. Concurrency and Computation: Practice and Experience, 32(1):e5084:1– e5084:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wen:2020:CUP

[WWL^{+}20] Yiping Wen, Yuan Wang, Jianxun Liu, Buqing Cao, and Qi Fu. CPU usage prediction for cloud resource provisioning based on deep belief network and particle swarm optimization. Concurrency and Computation: Practice and Experience, 32(14):e5730:1–e5730:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:NPM

[WWM^{+}23] Hai Wang, Yan Wang, Yuhui Ma, Chenguang Qin, Kan Zhang, Ling Gao, and Jie Ren. Noise processing and multitask learning for far-field dialect classification. Concurrency and Computation: Practice and Experience, 35(13):e7274:1– e7274:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wen:2020:INR

[WWW^{+}20] Yiping Wen, Feiran Wang, Rui Wu, Jianxun Liu, and Buqing Cao. Improving the novelty of retail commodity recommendations using multiarmed bandit and gradient boosting decision tree. Concurrency and Computation: Practice and Experience, 32(14):e5703:1–e5703:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2023:HDS

[WWW^{+}23] Jiashu Wu, Yang Wang, Jinpeng Wang, Hekang Wang, and Taorui Lin. How does solid-state drives cluster perform for distributed file systems: an empirical study. Concurrency and Computation: Practice and Experience, 35(21):e7709:1– e7709:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2024:NSI

[WWW24] Ping Wei, Shou Fa Wang, and Yi Ming Wang. Numerical study on the influence of centroid-to-buoyancy center ratio on the motion stability of supercavitating vehicle. Concurrency and Computation: Practice and Experience, 36(4):e7930:1– e7930:??, February 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:CSR

[WWZ^{+}22] Yubiao Wang, Junhao Wen, Wei Zhou, Xibin Wang, Quanwang Wu, and Bamei Tao. A cloud service recommendation method based on extended multi-source information fusion. Concurrency and Computation: Practice and Experience, 34 (10):e6826:1–e6826:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2020:EEL

[WXCY20] Wenjia Wu, Weihong Xu, Zhouguo Chen, and Ming Yang. Energy-efficient link scheduling in time-variant dual-hop 60GHz wireless networks. Concurrency and Computation: Practice and Experience, 32(23):e5903:1–e5903:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:OAS

[WXH^{+}22] Wenru Wang, Xiaojie Xu, Jian Hu, Xinghai Lu, and Xiaohua Shi. Optimizing an AltaRica simulator. Concurrency and Computation: Practice and Experience, 34(8):e5337:1– e5337:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2021:CPS

[WXLD21] Xiaohui Wei, Zhewen Xu, Hongliang Li, and Zhaohui Ding. Coordinated process scheduling algorithms for coupled earth system models. Concurrency and Computation: Practice and Experience, 33(20):e6346:1–e6346:??, October 25, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:ISS

[WXS^{+}23] Shudi Wang, Manman Xu, Ying Sun, Guozhang Jiang, Yaoqing Weng, Xin Liu, Guojun Zhao, Hanwen Fan, Jun Li, Cejing Zou, Yuanmin Xie, Li Huang, and Baojia Chen. Improved single shot detection using DenseNet for tiny target detection. Concurrency and Computation: Practice and Experience, 35 (2):e7491:1–e7491:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:SSP

[WY20] Rui Wang and Shi Ying. SaaS software performance issue diagnosis using independent component analysis and restricted Boltzmann machine. Concurrency and Computation: Practice and Experience, 32(14):e5729:1–e5729:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:SSN

[WYE23] Zongxia Wang, Naigong Yu, and Firdaous Essaf. A somasynapses neuron model and FPGA implementation. Concurrency and Computation: Practice and Experience, 35(27): e7851:1–e7851:??, December 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:VCS

[WYL^{+}22] Yan Wang, Bingqing Yang, Jingxin Liu, Hui Zeng, and Changqing Xia. Virtual chain: a storage model supporting cross-blockchain transaction. Concurrency and Computation: Practice and Experience, 34(12):e5899:1–e5899:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2024:JJO

[WYLL24] Di Wu, Zhen Yang, Tong Li, and Junrui Liu. JOCP: a jointly optimized clustering protocol for industrial wireless sensor networks using double-layer selection evolutionary algorithm. Concurrency and Computation: Practice and Experience, 36 (4):e7927:1–e7927:??, February 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2020:ESI

[WYZAD20] Yulei Wu, Zheng Yan, Zhiwei Zhao, and Ahmed Al-Dubai. Editorial: Special issue on SDN-based wireless network virtualization. Concurrency and Computation: Practice and Experience, 32(16):e5444:1–e5444:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:RIT

[WZ20a] Yongchang Wang and Ligu Zhu. Research on improved text classification method based on combined weighted model. Concurrency and Computation: Practice and Experience, 32 (6):e5140:1–e5140:??, March 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2020:CIC

[WZ20b] Zhen Wei and Wei Zhang. Construction of an intelligent cellular automata model for historic landscape islands in Shanghai and fusion simulation. Concurrency and Computation: Practice and Experience, 32(19):e5801:1–e5801:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2021:RCF

[WZ21] Zhen Wei and Wei Zhang. Research on the constraint factors of cellular automaton-based Shanghai historic landscape island smart integration protection: Taking Shanghai Moller Villa as example. Concurrency and Computation: Practice and Experience, 33(5):e5977:1–e5977:??, March 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2024:EIA

[WZ24] Xudong Wang and Weiyi Zhong. Evolution and innovations in animation: a comprehensive review and future directions. Concurrency and Computation: Practice and Experience, 36 (2):e7904:1–e7904:??, January 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Woo:2021:SGD

[WZB21] Jungha Woo, Lan Zhao, and Gabriel J. Bowen. Streamlining geospatial data processing for isotopic landscape modeling. Concurrency and Computation: Practice and Experience, 33 (19):e6324:1–e6324:??, October 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:PSL

[WZC^{+}22a] Feng Wang, Lingling Zhang, Xingchen Chen, Ziming Wang, and Xin Xu. A personalized self-learning system based on knowledge graph and differential evolution algorithm. Concurrency and Computation: Practice and Experience, 34(8): e6190:1–e6190:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wen:2022:HEM

[WZC^{+}22b] Yingying Wen, Yiming Zhang, Guanjie Cheng, Shuiguang Deng, and Jianwei Yin. A holistic evaluation methodology for configuring production data centers. Concurrency and Computation: Practice and Experience, 34(25):e7257:1–e7257:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2021:DCD

[WZG^{+}21] Ling Wu, Qishan Zhang, Kun Guo, Erbao Chen, and Chaoyang Xu. Dynamic community detection method based on an improved evolutionary matrix. Concurrency and Computation: Practice and Experience, 33(8):e5314:1–e5314:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2021:NFT

[WZHL21] Desheng Wang, Weizhe Zhang, Hui He, and Chuanyi Liu. Node-fusion: Topology-aware virtual network embedding algorithm for repeatable virtual network mapping over substrate nodes. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:USG

[WZL^{+}22] Wei Wang, Han Zhang, Yang Li, Zhenyu Zhang, Xiangfeng Luo, and Shaorong Xie. USVs-Sim: a general simulation platform for unmanned surface vessels autonomous learning. Concurrency and Computation: Practice and Experience, 34 (3):e6567:1–e6567:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wan:2024:LSH

[WZL^{+}24] YunBo Wan, Zhong Zhao, Jie Liu, Laiping Zhang, Yong Zhang, and Jianqiang Chen. Large-scale homo- and heterogeneous parallel paradigm design based on CFD application PHengLEI. Concurrency and Computation: Practice and Experience, 36(5):e7933:1–e7933:??, February 29, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:AIT

[WZMJ23a] Yufeng Wang, Bo Zhang, Jianhua Ma, and Qun Jin. Artificial intelligence of things (AIoT) data acquisition based on graph neural networks: a systematical review. Concurrency and Computation: Practice and Experience, 35(23):e7827:1– e7827:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2023:LFR

[WZMJ23b] Yufeng Wang, Xu Zhang, Jianhua Ma, and Qun Jin. LDP-Fed+: a robust and privacy-preserving federated learning based classification framework enabled by local differential privacy. Concurrency and Computation: Practice and Experience, 35(19):e7429:1–e7429:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wu:2020:DGA

[WZSZ20] Wenxia Wu, Jinxiu Zhu, Xin Su, and Xuewu Zhang. DHGAN: Generative adversarial network with dark channel prior for single-image dehazing. Concurrency and Computation: Practice and Experience, 32(18):e5263:1–e5263:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2022:QAN

[WZX^{+}22] Wan Wei, Lian Zhanghua, Peng Xingyu, Yao Dongchi, and Liu Chang. Qualitative analysis of NET-BOWTIE risk model of dehydration and dehydrocarbon station based on improved compression algorithm. Concurrency and Computation: Practice and Experience, 34(9):e5808:1–e5808:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:OSF

[WZY^{+}22] Chenxi Wang, Xiaoqing Zhang, Liqin Ye, Yu Mao, Shaozi Li, and Yaojin Lin. Online streaming feature selection for multigranularity hierarchical classification learning. Concurrency and Computation: Practice and Experience, 34(17):e6994:1– e6994:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2020:MVN

[WZZ^{+}20] Cong Wang, Fanghui Zheng, Guangcong Zheng, Sancheng Peng, Zejie Tian, Yujia Guo, Guorui Li, and Ying Yuan. Modeling on virtual network embedding using reinforcement learning. Concurrency and Computation: Practice and Experience, 32(23):e6020:1–e6020:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wang:2022:SVD

[WZZ^{+}22] Lei Wang, Yunqiu Zhang, Xubin Zheng, Qi Yu, Shuhan Chen, and Junyao Ding. Singular value decomposition-based behavior-aware cloud service application programming interfaces recommendation for large-scale software cloud directory platforms. Concurrency and Computation: Practice and Experience, 34(21):e7121:1–e7121:??, September 25, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2022:NMB

[XA22] Yanfei Xu and Karlo Abnoosian. A new metaheuristic-based method for solving the virtual machines migration problem in the green cloud computing. Concurrency and Computation: Practice and Experience, 34(3):e6579:1–e6579:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2020:OAC

[XAC^{+}20] Jiajian Xiao, Philipp Andelfinger, Wentong Cai, Paul Richmond, Alois Knoll, and David Eckhoff. OpenABLext: an automatic code generation framework for agent-based simulations on CPU–GPU-FPGA heterogeneous platforms. Concurrency and Computation: Practice and Experience, 32(21): e5807:1–e5807:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xue:2020:UPD

[XCD^{+}20] Fei Xue, Yang Cao, Zhiming Ding, Hengliang Tang, Xi Yang, Lei Chen, and Juntao Li. Urban population density estimation based on spatio-temporal trajectories. Concurrency and Computation: Practice and Experience, 32(14):e5685:1– e5685:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2022:DRP

[XCG^{+}22] Xiaolong Xu, Hao Cao, Qingfan Geng, Xihua Liu, Fei Dai, and Chuanjian Wang. Dynamic resource provisioning for workflow scheduling under uncertainty in edge computing environment. Concurrency and Computation: Practice and Experience, 34(14):e5674:1–e5674:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2022:MDM

[XCJ22] Yaobin Xie, Rui Chang, and Liehui Jiang. A malware detection method using satisfiability modulo theory model checking for the programmable logic controller system. Concurrency and Computation: Practice and Experience, 34(16):e5724:1– e5724:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2020:IRC

[XCX^{+}20] Qilin Xiao, Yun Cheng, Minlei Xiao, Jun Zhang, Hongji Shi, Lihui Niu, Chenguang Ge, and Haitao Lang. Improved region convolutional neural network for ship detection in multiresolution synthetic aperture radar images. Concurrency and Computation: Practice and Experience, 32(22):e5820:1–e5820:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xia:2022:USL

[XCY22] Xiangli Xia, Wei Cheng, and Liu Yang. Urban spatial location service prediction algorithm based on fast adaptive genetic algorithm-least squares support vector machine under the background of Internet of Things. Concurrency and Computation: Practice and Experience, 34(13):e5946:1–e5946:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2021:IMC

[XCZ^{+}21] Xiaolong Xu, Qing Cai, Guoming Zhang, Jie Zhang, Wei Tian, Xiaorui Zhang, and Alex X. Liu. An incentive mechanism for crowdsourcing markets with social welfare maximization in cloud-edge computing. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2020:GDP

[XDH^{+}20] Zhipu Xie, Bowen Du, Shangfo Huang, Bo Huang, Leilei Sun, and Weifeng Lv. Guaranteeing differential privacy for sequence predictions in bike sharing systems. Concurrency and Computation: Practice and Experience, 32(19):e5770:1– e5770:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2021:MDI

[XDJY21] Hui Xu, Xiaofeng Ding, Hai Jin, and Qing Yu. A multidimensional index for privacy-preserving queries in cloud computing. Concurrency and Computation: Practice and Experience, 33(8):e5458:1–e5458:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2022:SPO

[XDL22] Gang Xu, Qianqian Dong, and Min Li. A simulated parameter optimization method-based manifold learning for a production process. Concurrency and Computation: Practice and Experience, 34(7):e5521:1–e5521:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2021:BCS

[XFH^{+}21] Xiaohua Xu, Baichuan Fan, Ping He, Yali Liang, Jie Ding, Yuan Lou, Zhijun Zhang, and Xincheng Chang. Bidirectional compressive sensing for classification of gene expression data. Concurrency and Computation: Practice and Experience, 33 (15):e5120:1–e5120:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiong:2024:SCG

[XFPY24] Tianhong Xiong, Shangqing Feng, Maolin Pan, and Yang Yu. Smart contract generation for inter-organizational process collaboration. Concurrency and Computation: Practice and Experience, 36(10):e7961:1–e7961:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xue:2023:GES

[XGCZ23] Zhaoyu Xue, Wanwan Guo, Zhihua Cui, and Wensheng Zhang. The global evaluation strategy for many-objective partial collaborative computation offloading problem. Concurrency and Computation: Practice and Experience, 35(2): e7474:1–e7474:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiang:2021:HNB

[XGX^{+}21] Yan Xiang, Jun-Jun Guo, Yan-Tuan Xian, Yu-Xin Huang, and Zheng-Tao Yu. Hybrid node-based tensor graph convolutional network for aspect-category sentiment classification of microblog comments. Concurrency and Computation: Practice and Experience, 33(21):e6431:1–e6431:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2020:CBD

[XGZ^{+}20] Xianlei Xu, Wenru Gao, Di Zhang, Taotao Li, and Xu Qiao. Cloud-based difference algorithm using big GPR data for roadbed damage detection. Concurrency and Computation: Practice and Experience, 32(23):e5545:1–e5545:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2022:RAM

[XHM22] Bing Xu, Darong Huang, and Bo Mi. Research on the application of mobile payment security system based on the Internet of Things. Concurrency and Computation: Practice and Experience, 34(9):e5779:1–e5779:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2020:FAN

[XHST20] Yibin Xu, Yangyu Huang, Jianhua Shao, and George Theodorakopoulos. A flexible n/2 adversary node resistant and halting recoverable blockchain sharding protocol. Concurrency and Computation: Practice and Experience, 32(19): e5773:1–e5773:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xia:2021:SSP

[XHZ^{+}21] Yao Xia, Zhiqiu Huang, Yonglong Zhang, Min Yuan, Shangguang Wang, and Yu Zhou. SPASC: Strategy-proof auction mechanism with cost and QoS incentive for service composition. Concurrency and Computation: Practice and Experience, 33(9):e6131:1–e6131:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xin-Hui:2022:SAE

[XHZHXBQX22] Zhao Xin-Hui, Wu Ze-Hui, Song Xiao-Bin, and Wang Qing-Xian. Secure analysis on entire software-defined network using coloring distribution model. Concurrency and Computation: Practice and Experience, 34(14):e5541:1–e5541:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiaoyan:2020:REM

[Xia20] He Xiaoyan. Research on the evolution mechanism of the electric vehicle market driven by big data. Concurrency and Computation: Practice and Experience, 32(6):e5148:1–e5148:??, March 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2022:RAI

[XJW^{+}22] Xin Xie, Xunyi Jiang, Weiru Wang, Bin Wang, Tiancheng Wan, Wenliang Tang, and Xianmin Wang. Research and application of intrusion detection method based on hierarchical features. Concurrency and Computation: Practice and Experience, 34(16):e5799:1–e5799:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2021:ACH

[XL21] Xiaodong Xu and Hongkai Li. Adaptive control for a hypersonic vehicle based on evolutionary algorithm and convex optimization. Concurrency and Computation: Practice and Experience, 33(12):e5555:1–e5555:??, June 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2023:R

[XL23] Xiaodong Xu and Hongkai Li. Retraction. Concurrency and Computation: Practice and Experience, 33(12):e7752:1– e7752:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2020:SCS

[XLL^{+}20] Xiuzhen Xie, Lei Li, Sheng Lian, Shaohao Chen, and Zhiming Luo. SERU: a cascaded SE-ResNeXT U-Net for kidney and tumor segmentation. Concurrency and Computation: Practice and Experience, 32(14):e5738:1–e5738:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2021:SMS

[XLL^{+}21] Xiaocong Xiao, Jianxun Liu, Deshun Liu, Yufei Tang, Juchuan Dai, and Fan Zhang. SSAE-MLP: Stacked sparse autoencoders-based multi-layer perceptron for main bearing temperature prediction of large-scale wind turbines. Concurrency and Computation: Practice and Experience, 33(17): e6315:1–e6315:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2023:UAD

[XLL^{+}23] Rongbin Xu, Zhiqiang Liu, Yuanmo Lin, Xianliang Feng, Hui Lin, and Ying Xie. Unified adaptive deep classification for industrial real-time situation awareness in edge environment. Concurrency and Computation: Practice and Experience, 35 (14):e6488:1–e6488:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2020:MAN

[XLWX20] Xiaoliang Xu, Chang Li, Yuxiang Wang, and Yixing Xia. Multiattribute approximate nearest neighbor search based on navigable small world graph. Concurrency and Computation: Practice and Experience, 32(24):e5970:1–e5970:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xia:2021:STO

[XLX^{+}21] Daoxun Xia, Haojie Liu, Lili Xu, Jiawen Li, and Linna Wang. Self-training with one-shot stepwise learning method for person re-identification. Concurrency and Computation: Practice and Experience, 33(17):e6296:1–e6296:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2020:IMB

[XLXZ20] Yi-Han Xu, Meng-Lian Liu, Jing-Wei Xie, and Jun Zhou. An IEEE 802.21 MIS-based mobility management for D2D communications over heterogeneous networks (HetNets). Concurrency and Computation: Practice and Experience, 32(5): e5552:1–e5552:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xue:2023:FII

[XLY^{+}23] Hanyu Xue, Bo Liu, Xin Yuan, Ming Ding, and Tianqing Zhu. Face image de-identification by feature space adversarial perturbation. Concurrency and Computation: Practice and Experience, 35(5):e7554:1–e7554:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2020:IEL

[XLZ20] Penghua Xie, Ming Lv, and Juanjuan Zhao. An improved energy-low clustering hierarchy protocol based on ensemble algorithm. Concurrency and Computation: Practice and Experience, 32(7):e5575:1–e5575:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2022:STT

[XLZL22] Xing Xu, Chengxing Liu, Yun Zhao, and Xiaoshu Lv. Shortterm traffic flow prediction based on whale optimization algorithm optimized BiLSTM Attention. Concurrency and Computation: Practice and Experience, 34(10):e6782:1–e6782:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiong:2022:PPR

[XLZZ22] Ping Xiong, Lin Liang, Yunli Zhu, and Tianqing Zhu. PriTxt: a privacy risk assessment method for text data based on semantic correlation learning. Concurrency and Computation: Practice and Experience, 34(5):e6680:1–e6680:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2023:LGA

[XPLX23] Lili Xu, Houfu Peng, Xing Lu, and Daoxun Xia. Learning to generalize aerial person re-identification using the metatransfer method. Concurrency and Computation: Practice and Experience, 35(12):e7687:1–e7687:??, May 30, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xue:2021:MGP

[XR21] Weicheng Xue and Christoper J. Roy. Multi-GPU performance optimization of a computational fluid dynamics code using OpenACC. Concurrency and Computation: Practice and Experience, 33(5):e6036:1–e6036:??, March 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2020:TCA

[XSGL20] Rongna Xie, Guozhen Shi, Yunchuan Guo, and Fenghua Li. A topic-centric access control model for the publish/subscribe paradigm. Concurrency and Computation: Practice and Experience, 32(9):e5614:1–e5614:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2020:NAF

[XSZ^{+}20] Jia Xu, Shouxin Song, Huaiyuan Zhai, Pengwei Yuan, and Mingli Chen. A new analytical framework for network vulnerability on subway system. Concurrency and Computation: Practice and Experience, 32(23):e5508:1–e5508:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2023:NSP

[XSZ^{+}23] Shangqun Xie, Hanxiao Suo, Rui Zhou, Juanhua Zhu, and Man Fan. Numerical study on over-pressure performance of premixed explosion in utility tunnel. Concurrency and Computation: Practice and Experience, 35(5):e7561:1–e7561:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2024:RBA

[XW24] Songyi Xiao and Wenjun Wang. Ranking-based architecture generation for surrogate-assisted neural architecture search. Concurrency and Computation: Practice and Experience, 36 (12):e8051:1–e8051:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xia:2022:TSA

[XWC^{+}22] Chengpeng Xia, Yalan Wu, Long Chen, Yawen Chen, and Jigang Wu. Three-stage auction scheme for computation offloading on mobile blockchain with edge computing. Concurrency and Computation: Practice and Experience, 34(25): e7253:1–e7253:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xing:2022:ESS

[XWD^{+}22] Longyue Xing, Zhaoshun Wang, Zhezhao Ding, Genshen Chu, Lingyu Dong, and Nan Xiao. An efficient sparse stiffness matrix vector multiplication using compressed sparse row storage format on AMD GPU. Concurrency and Computation: Practice and Experience, 34(23):e7186:1–e7186:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2023:VVF

[XWQ^{+}23] Binjing Xu, Qiong Wu, Hong Qin, Zhiyuan Liu, Lin Shi, and Shuang Li. VCSNet: V-fused color saliency net. Concurrency and Computation: Practice and Experience, 35(26):e7127:1– e7127:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2021:MAB

[XWW^{+}21] Minyang Xu, Wenjun Wang, Hui Wang, Songyi Xiao, and Zhikai Huang. Multipopulation artificial bee colony algorithm based on a modified probability selection model. Concurrency and Computation: Practice and Experience, 33(13):e6216:1– e6216:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2022:EOR

[XXD^{+}22] Fang Xu, Nan Xiao, Min Deng, Yong Xie, Zenggang Xiong, and Qiong Xu. Efficient opportunistic routing with social context awareness for distributed mobile social networks. Concurrency and Computation: Practice and Experience, 34(7): e5524:1–e5524:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiong:2023:SUB

[XXSL23] Runqun Xiong, Lan Xiong, Feng Shan, and Junzhou Luo. SBHA: an undetectable black hole attack on UANET in the sky. Concurrency and Computation: Practice and Experience, 35(13):e6700:1–e6700:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2021:UFE

[XY21] Li Xie and Sheng Yu. Unsupervised feature extraction with convolutional autoencoder with application to daily stock market prediction. Concurrency and Computation: Practice and Experience, 33(16):e6282:1–e6282:??, August 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2023:UAS

[XYFZ23] Jin Xu, Huiqun Yu, Guisheng Fan, and Jiayin Zhang. Uncertainty-aware scheduling of real-time workflows under deadline constraints on multi-cloud systems. Concurrency and Computation: Practice and Experience, 35(5):e7562:1– e7562:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xie:2021:AET

[XYLW21] Jiahao Xie, Xiai Yan, Yaping Lin, and Jianhao Wei. An accurate and efficient two-phase scheme for detecting Android cloned applications. Concurrency and Computation: Practice and Experience, 33(5):e6009:1–e6009:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xian:2024:FJJ

[XYY24] Gang Xian, Wenxiang Yang, and Jie Yu. FP-JSC: Job failure prediction on supercomputers through job application sequence correlation. Concurrency and Computation: Practice and Experience, 36(14):e8041:1–e8041:??, June 25, 2024. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2020:SIAb

[XZ20a] Zheng Xu and Qingyuan Zhou. Special issue on applications and techniques in cyber intelligence 2019. Concurrency and Computation: Practice and Experience, 32(23):e5682:1– e5682:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2020:SIAa

[XZ20b] Zheng Xu and Qingyuan Zhou. Special issue on applications and techniques in cyber intelligence Concurrency and Computation: Practice and Experience. Concurrency and Computation: Practice and Experience, 32(6):e5274:1–e5274:??, March 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2022:SPC

[XZ22] Bei Xu and Ziyuan Zhuang. Survey on psychotherapy chatbots. Concurrency and Computation: Practice and Experience, 34(7):e6170:1–e6170:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xia:2021:HDP

[XZD^{+}21] Yang Xia, Tianqing Zhu, Xiaofeng Ding, Hai Jin, and Deqing Zou. Heterogeneous differential privacy for vertically partitioned databases. Concurrency and Computation: Practice and Experience, 33(8):e5607:1–e5607:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2023:OPP

[XZG^{+}23] Limei Xiao, Yachao Zhang, Weizhe Gao, Dayou Xu, and Ce Li. An object perception and positioning method via deep perception learning object detection. Concurrency and Computation: Practice and Experience, 35(16):e6203:1–e6203:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiong:2020:MPM

[XZL^{+}20a] Jinbo Xiong, Yuanyuan Zhang, Li Lin, Jian Shen, Xuan Li, and Mingwei Lin. ms-PoSW: a multi-server aided proof of shared ownership scheme for secure deduplication in cloud. Concurrency and Computation: Practice and Experience, 32 (3):e4252:1–e4252:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2020:GEA

[XZL^{+}20b] Zifeng Xu, Fucai Zhou, Jin Li, Yuxi Li, and Qiang Wang. Graph encryption for all-path queries. Concurrency and Computation: Practice and Experience, 32(16):e5362:1–e5362:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2020:SIS

[xZlGCzJ20] Xin xin Zhang, Lei lei Gu, Hong Chen, and Guo zhu Jia. Study on the influence of surrounding urban SO_{2}, NO_{2}, and CO on haze formation in Beijing based on MF-DCCA and boosting algorithms. Concurrency and Computation: Practice and Experience, 32(24):e5921:1–e5921:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2020:EPE

[XZW^{+}20] Yang Xu, Quanrun Zeng, Guojun Wang, Cheng Zhang, Ju Ren, and Yaoxue Zhang. An efficient privacy-enhanced attribute-based access control mechanism. Concurrency and Computation: Practice and Experience, 32(5):e5556:1– e5556:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2021:MVS

[XZXV21] Lili Xiao, Huibiao Zhu, Shuangqing Xiang, and Phan Cong Vinh. Modeling and verifying SDN under multi-controller architectures using CSP. Concurrency and Computation: Practice and Experience, 33(2):e5334:1–e5334:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xian:2022:PPJ

[XZY^{+}22] Gang Xian, Xiaorong Zhang, Jie Yu, Guijuan Wang, Wenxiang Yang, Longfang Zhou, Yadong Wu, Xuejun Li, and Xin He. PreF: Predicting job failure on supercomputers with job path and user behavior. Concurrency and Computation: Practice and Experience, 34(23):e7202:1–e7202:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xu:2023:SSE

[XZYW23] Wanshan Xu, Jianbiao Zhang, Yilin Yuan, and Xiao Wang. Symmetric searchable encryption with supporting search pattern and access pattern protection in multi-cloud. Concurrency and Computation: Practice and Experience, 35(9): e7651:1–e7651:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yazici:2022:NVA

[YA22a] Ilkay Melek Yazici and Mehmet S. Aktas. A novel visualization approach for data provenance. Concurrency and Computation: Practice and Experience, 34(9):e6523:1–e6523:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yigit:2022:AIM

[YA22b] Gulsum Yigit and Mehmet Fatih Amasyali. Assessing the impact of minor modifications on the interior structure of GRU: GRU1 and GRU2. Concurrency and Computation: Practice and Experience, 34(20):e6775:1–e6775:??, September 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yalcin:2022:WPF

[Yal22] Sercan Yalçin. Weather parameters forecasting with time series using deep hybrid neural networks. Concurrency and Computation: Practice and Experience, 34(21):e7141:1– e7141:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yadav:2022:EEM

[YAR22] Subhash Kumar Yadav, Shabbir Ahmad, and Muhammad Riaz. On enhanced estimation and monitoring of population variance using robust auxiliary parameters based estimators. Concurrency and Computation: Practice and Experience, 34(24):e7221:1–e7221:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yalcin:2023:PBP

[YB23a] Emre Yalcin and Alper Bilge. Popularity bias in personality perspective: an analysis of how personality traits expose individuals to the unfair recommendation. Concurrency and Computation: Practice and Experience, 35(9):e7647:1– e7647:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yarava:2023:ETI

[YB23b] Anitha Yarava and C. Shoba Bindu. An efficient trust inference model in online social networks using fuzzy Petri nets. Concurrency and Computation: Practice and Experience, 35 (6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yeo:2023:CSG

[YBJ^{+}23] Sangho Yeo, Minho Bae, Minjoong Jeong, Oh-Kyoung Kwon, and Sangyoon Oh. Crossover-SGD: a gossip-based communication in distributed deep learning for alleviating large mini-batch problem and enhancing scalability. Concurrency and Computation: Practice and Experience, 35(15):e7508:1– e7508:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yildirim:2022:PAD

[YÇC22] Emre Yildirim, Ali Çalhan, and Murtaza Cicioglu. Performance analysis of disease diagnostic system using IoMT and real-time data analytics. Concurrency and Computation: Practice and Experience, 34(13):e6916:1–e6916:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2022:LMD

[YCL^{+}22a] Haichang Yao, Chen Chen, Shangdong Liu, Kui Li, Yimu Ji, Guangyan Huang, and Ruchuan Wang. Lane marking detection algorithm based on high-precision map and multisensor fusion. Concurrency and Computation: Practice and Experience, 34(8):e5797:1–e5797:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2022:PCL

[YCL^{+}22b] Haichang Yao, Shuai Chen, Shangdong Liu, Kui Li, Yimu Ji, Guangyong Hu, and Ruchuan Wang. Parallel compression for large collections of genomes. Concurrency and Computation: Practice and Experience, 34(2):e6339:1–e6339:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yi:2020:BOF

[YCY20] ZhengMing Yi, Fei Chen, and YiPing Yao. A barrier optimization framework for NUMA multi-core system. Concurrency and Computation: Practice and Experience, 32(5): e5527:1–e5527:??, March 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2020:ADD

[YF20] Guangzhou Yu and Weina Fu. Analysis of distributed database access path prediction based on recurrent neural network in Internet of Things. Concurrency and Computation: Practice and Experience, 32(1):e5116:1–e5116:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yokoyama:2023:IHS

[YFdP^{+}23] André M. Yokoyama, Mariza Ferro, Felipe B. de Paula, Vitor G. Vieira, and Bruno Schulze. Investigating hardware and software aspects in the energy consumption of machine learning: a green AI-centric analysis. Concurrency and Computation: Practice and Experience, 35(24):e7825:1–e7825:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yasmina:2022:WSS

[YFF22] Remaci Zeyneb Yasmina, Hadjila Fethallah, and Lahfa Fadoua. Web service selection and composition based on uncertain quality of service. Concurrency and Computation: Practice and Experience, 34(1):e6531:1–e6531:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2023:AMT

[YGS^{+}23] Qiang Yin, JiaXin Gu, Shaoyun Song, Yonglin Zhang, Gang Zhao, and ZhiQiang Hao. Acquisition method of target leveling height based on machine vision. Concurrency and Computation: Practice and Experience, 35(3):e7513:1–e7513:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2021:SRA

[YGZ^{+}21] Yuecheng Yu, Yu Gu, Huayu Zuo, Jinlei Wang, and Dongsheng Wang. Social recommendation algorithms with user feedback information. Concurrency and Computation: Practice and Experience, 33(22):e5934:1–e5934:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2021:CLL

[YHL^{+}21] Qiang Yin, Ao Hu, Qian Li, Xin Wei, Hongjun Yang, Beihai Wang, and Guoquan Zhang. Compound lower limb vibration training rehabilitation robot. Concurrency and Computation: Practice and Experience, 33(6):e6059:1–e6059:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2023:RLA

[YHL^{+}23] Xiaojing Yin, Jiwei Huang, Lei Liu, Wei He, and Lizhen Cui. An reinforcement learning approach for allocating software resources. Concurrency and Computation: Practice and Experience, 35(14):e6349:1–e6349:??, June 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2022:STS

[YHOY22] Shuiqiao Yang, Guangyan Huang, Bahadorreza Ofoghi, and John Yearwood. Short text similarity measurement using context-aware weighted biterms. Concurrency and Computation: Practice and Experience, 34(8):e5765:1–e5765:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yibin:2022:RCQ

[Yib22a] Xiang Yibin. Research on the construction of quantum induction universe network. Concurrency and Computation: Practice and Experience, 34(7):e5501:1–e5501:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yildirim:2022:AII

[YIB22b] Tulay Yildirim, Mirjana Ivanovic, and Ladjel Bellatreche. Advances on innovative issues in intelligent systems and applications. Concurrency and Computation: Practice and Experience, 34(14):e7042:1–e7042:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yildiz:2021:OBI

[Yil21a] Beytullah Yildiz. Optimizing bitmap index encoding for high performance queries. Concurrency and Computation: Practice and Experience, 33(18):e5943:1–e5943:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yilmaz:2021:IPA

[Yil21b] Volkan Yilmaz. Investigation of the performances of advanced image classification-based ground filtering approaches for digital terrain model generation. Concurrency and Computation: Practice and Experience, 33(13):e6219:1–e6219:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yilmaz:2021:NDS

[Yil21c] Volkan Yilmaz. A Non-Dominated Sorting Genetic Algorithm-II-based approach to optimize the spectral and spatial quality of component substitution-based pansharpened images. Concurrency and Computation: Practice and Experience, 33(5): e6030:1–e6030:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yildirim:2022:ACD

[Yil22a] Muhammed Yildirim. Automatic classification and diagnosis of heart valve diseases using heart sounds with MFCC and proposed deep model. Concurrency and Computation: Practice and Experience, 34(24):e7232:1–e7232:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yildiz:2022:RLU

[Yil22b] Beytullah Yildiz. Reinforcement learning using fully connected, attention, and transformer models in knapsack problem solving. Concurrency and Computation: Practice and Experience, 34(9):e6509:1–e6509:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yilmaz:2022:RII

[Yil22c] Volkan Yilmaz. The role of image interpolation in pansharpening. Concurrency and Computation: Practice and Experience, 34(11):e6892:1–e6892:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yilmaz:2023:NGT

[Yil23] Buse Yilmaz. A novel graph transformation strategy for optimizing SpTRSV on CPUs. Concurrency and Computation: Practice and Experience, 35(24):e7761:1–e7761:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yim:2023:QBE

[YJJ23] Yin-Goo Yim, Hyeon-Jun Jang, and Hyun-Wook Jin. QoS for best-effort batch jobs in container-based cloud. Concurrency and Computation: Practice and Experience, 35(15):e6422:1– e6422:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2020:ICR

[YJY^{+}20] Yongjia Yu, Vasu Jindal, I-Ling Yen, Farokh Bastani, Jie Xu, and Peter Garraghan. Integrating clustering and regression for workload estimation in the cloud. Concurrency and Computation: Practice and Experience, 32(23):e5931:1–e5931:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ye:2021:EGA

[YJY^{+}21] Ayong Ye, Junlin Jin, Zhijiang Yang, Ziwen Zhao, and LingYu Meng. Evolutionary game analysis on competition strategy choice of application providers. Concurrency and Computation: Practice and Experience, 33(8):e5446:1–e5446:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yenilmez:2023:NEG

[YK23] Ismail Yenilmez and Yeliz Mert Kantar. New exponentiated generalized censored regression models: Monte Carlo simulation and application. Concurrency and Computation: Practice and Experience, 35(1):e7436:1–e7436:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yadav:2023:GND

[YKL23] Kuldeep Singh Yadav, Anish Monsley Kirupakaran, and Rabul Hussain Laskar. GCR-Net: a deep learning-based bare hand detection and gesticulated character recognition system for human-computer interaction. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2020:HRA

[YKW20] Charlene Yang, Thorsten Kurth, and Samuel Williams. Hierarchical roofline analysis for GPUs: Accelerating performance optimization for the NERSC-9 Perlmutter system. Concurrency and Computation: Practice and Experience, 32 (20):e5547:1–e5547:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2020:ORS

[YL20a] Bo Yang and Chang Liu. Ontology reasoning of start-ups’ growth risk based on immune system. Concurrency and Computation: Practice and Experience, 32(23):e5518:1–e5518:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yong-lian:2020:MFD

[Yl20b] Luo Yong-lian. Multi-feature data mining for CT image recognition. Concurrency and Computation: Practice and Experience, 32(1):e4885:1–e4885:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2020:SCR

[YL20c] Zhao Yu and Xun Lin. Social car: The research of interaction design on the driver’s communication system. Concurrency and Computation: Practice and Experience, 32(6):e5139:1– e5139:??, March 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2020:SRC

[YLGY20] Li Yin, Cuiye Liu, Songtao Guo, and Yuanyuan Yang. Sparse random compressive sensing based data aggregation in wireless sensor networks. Concurrency and Computation: Practice and Experience, 32(3):e4455:1–e4455:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2022:VDC

[YLJ22] Chunxia Yin, Jian Liu, and Shunfu Jin. A virtualized data center energy-saving mechanism based on switching operating mode of physical servers and reserving virtual machines. Concurrency and Computation: Practice and Experience, 34(9): e5785:1–e5785:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yingying:2022:QBR

[YLJH22] Su Yingying, Han Lianjuan, Wang Jianan, and Wang Huimin. Quantum-behaved RS-PSO-LSSVM method for quality prediction in parts production processes. Concurrency and Computation: Practice and Experience, 34(7):e5522:1–e5522:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ye:2024:VRM

[YLKC24] Yunfeng Ye, Huifang Liu, Wenhui Kuang, and Wenxuan Chen. Virtual reality modeling application based on multi perspective and deep learning in the new media presentation and brand building of dongguan city memory. Concurrency and Computation: Practice and Experience, 36(11):e8015:1– e8015:??, May 15, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2020:RLO

[YLLL20] Lin Yang, Zhiming Luo, Wangqing Lin, and Shaozi Li. Research on line overload identification of power system based on improved neural network algorithm. Concurrency and Computation: Practice and Experience, 32(24):e5933:1–e5933:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2021:BBA

[YLM21] Guang Yang, Chunlei Li, and Kjell E. Marstein. A blockchainbased architecture for securing electronic health record systems. Concurrency and Computation: Practice and Experience, 33(14):e5479:1–e5479:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2023:AMD

[YLS^{+}23] Xuanxia Yao, Yang Li, Zhiguo Shi, Kaijun Liu, and Xiao-Jiang Du. Android malware detection based on sensitive features combination. Concurrency and Computation: Practice and Experience, 35(6):1, March 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2021:EAE

[YLT^{+}21] Chunrong Yao, Wantao Liu, Weiqing Tang, Jinrong Guo, Songlin Hu, Yijun Lu, and Wei Jiang. Evaluating and analyzing the energy efficiency of CNN inference on highperformance GPU. Concurrency and Computation: Practice and Experience, 33(6):e6064:1–e6064:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yuan:2021:EFD

[YLW^{+}21] Jiangjun Yuan, Weinan Liu, Jie Wang, Jiawen Shi, and Ling Miao. An efficient framework for data aggregation in smart agriculture. Concurrency and Computation: Practice and Experience, 33(10):e6160:1–e6160:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yi:2022:EAD

[YLW^{+}22] Shanwen Yi, Xiaole Li, Hua Wang, Yao Qin, and Jiaxin Yan. Energy-aware disaster backup among cloud datacenters using multiobjective reinforcement learning in software defined network. Concurrency and Computation: Practice and Experience, 34(3):e6588:1–e6588:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2020:HGR

[YLWL20] Bin Yu, Zhiming Luo, Huangbin Wu, and Shaozi Li. Hand gesture recognition based on attentive feature fusion. Concurrency and Computation: Practice and Experience, 32(22): e5910:1–e5910:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yongchareon:2020:UUF

[YLZ20] Sira Yongchareon, Chengfei Liu, and Xiaohui Zhao. Uni-FlexView: a unified framework for consistent construction of BPMN and BPEL process views. Concurrency and Computation: Practice and Experience, 32(11):e5646:1–e5646:??, June 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2021:EEE

[YLZ^{+}21] Hui Yan, Ya Li, Xiaomin Zhu, Dayu Zhang, Ji Wang, Huangke Chen, and Weidong Bao. EASE: Energy-efficient task scheduling for edge computing under uncertain runtime and unstable communication conditions. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ye:2023:NMO

[YLZC23] Tao Ye, Wenting Li, Jiangjiang Zhang, and Zhihua Cui. A novel multi-objective immune optimization algorithm for under sampling software defect prediction problem. Concurrency and Computation: Practice and Experience, 35(4): e7525:1–e7525:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2023:SFS

[YLZT23] Mengmeng Yang, Kwok-Yan Lam, Tianqing Zhu, and Chenghua Tang. SPoFC: a framework for stream data aggregation with local differential privacy. Concurrency and Computation: Practice and Experience, 35(5):e7572:1–e7572:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2023:OAE

[YLZW23] Zhongjun Yang, Zhi Liu, Xuejun Zong, and Guogang Wang. An optimized adaptive ensemble model with feature selection for network intrusion detection. Concurrency and Computation: Practice and Experience, 35(4):e7529:1–e7529:??, February 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yilmazer-Metin:2022:SES

[YM22] Ayse Yilmazer-Metin. sRSP: an efficient and scalable implementation of remote scope promotion. Concurrency and Computation: Practice and Experience, 34(9):e6483:1–e6483:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yogegowda:2022:EDC

[YMKH22] Prasad Adaguru Yogegowda, Jyoti Metan, Kurilinga Sannalingappa Ananda Kumar, and Shiva Prasad Kumbenahalli Hanumanthegowda. Early detection and classification of liver diseases in ultrasound images using hybrid machine learning techniques. Concurrency and Computation: Practice and Experience, 34(27):e7341:1–e7341:??, December 10, 2022. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2021:DSC

[YMWA21] Zhenkun Yang, Xianghua Ma, Kefan Wang, and Jing An. Dense short connection network for efficient image classification. Concurrency and Computation: Practice and Experience, 33(10):e6186:1–e6186:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2020:RRB

[YMZ^{+}20] Zhe Yang, Kun Ma, Xiaoli Zhang, Lizhen Cui, and Bo Yang. RSCVC: Row-based semantic cache with incremental versioning consistency. Concurrency and Computation: Practice and Experience, 32(17):e5672:1–e5672:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ye:2021:NMM

[YMZD21] Ayong Ye, Lingyu Meng, Jiaomei Zhang, and Yiqing Diao. A new γ-map mechanism for mobility traces privacy. Concurrency and Computation: Practice and Experience, 33(6): e5965:1–e5965:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yasudo:2023:DLD

[YNK^{+}23] Ryota Yasudo, Koji Nakano, Michihiro Koibuchi, Hiroki Matsutani, and Hideharu Amano. Designing low-diameter interconnection networks with multi-ported host-switch graphs. Concurrency and Computation: Practice and Experience, 35 (11):e6115:1–e6115:??, May 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yurtsever:2022:FST

[YÖT^{+}22] M. Mücahit Enes Yurtsever, Muhammet Özcan, Zübeyir Taruz, Süleyman Eken, and Ahmet Sayar. Figure search by text in large scale digital document collections. Concurrency and Computation: Practice and Experience, 34(1):e6529:1– e6529:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2022:CMB

[YOWY22] Yi Yin, Lin Ouyang, Zhixiang Wu, and Shuifang Yin. A classification method based on encoder-decoder structure with paper content. Concurrency and Computation: Practice and Experience, 34(9):e5737:1–e5737:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yoon:2021:ESA

[YPO21] Daegun Yoon, Gyudong Park, and Sangyoon Oh. Exploring a system architecture of content-based publish/subscribe system for efficient on-the-fly data dissemination. Concurrency and Computation: Practice and Experience, 33(18):e6090:1– e6090:??, September 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yadav:2021:FAD

[YPY21] Mahendra Pratap Yadav, Nisha Pal, and Dharmendra Kumar Yadav. A formal approach for Docker container deployment. Concurrency and Computation: Practice and Experience, 33 (20):e6364:1–e6364:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yoosuf:2021:LFC

[YR21] Mohamed Sirajudeen Yoosuf and Anitha R. Lightweight fogcentric auditing scheme to verify integrity of IoT healthcare data in the cloud environment. Concurrency and Computation: Practice and Experience, 33(24):e6450:1–e6450:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yoo:2023:ACD

[YRSO23] Sinwoo Yoo, Yeongbin Ryu, Sunghoon Shin, and Hyukjun Oh. Analysis and classification of drone sounds from digital media. Concurrency and Computation: Practice and Experience, 35(16):e6671:1–e6671:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Y:2023:ALA

[YRV^{+}23] Mallikarjuna Rao Y., Munagala Surya Bhupal Rao, P. Veeresh, Y. Narasimha Reddy, and M. M. Raghavendra. Adaptive leaf abnormality segmentation and improved plant disease classification by tuned long short-term memory with recurrent neural network. Concurrency and Computation: Practice and Experience, 35(9):e7639:1–e7639:??, April 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yadav:2022:ULF

[YS22] Ajit Kumar Singh Yadav and Marpe Sora. Unsupervised learning for financial statement fraud detection using manta ray foraging based convolutional neural network. Concurrency and Computation: Practice and Experience, 34(27):e7340:1– e7340:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yildiz:2023:HHM

[YS23] Gülçin Yildiz and Fatih Erdogan Sevilgen. Hyper-heuristic method for processor allocation in parallel tasks scheduling. Concurrency and Computation: Practice and Experience, 35(24):e7757:1–e7757:??, November 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yuan:2020:MSA

[YSC^{+}20] Jiangjun Yuan, Jiawen Shi, Juyan Che, Chonghuan Xu, and Jie Wang. Modeling and simulation analysis of public opinion polarization in a dynamic network environment. Concurrency and Computation: Practice and Experience, 32(19):e5771:1– e5771:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ye:2022:PCT

[YSH^{+}22] Xin Ye, Jianqi Shi, Yanhong Huang, Qin Li, Hansheng Wei, and Xinyu Chen. Parallel computational tree logic modelchecking on pushdown systems. Concurrency and Computation: Practice and Experience, 34(23):e7173:1–e7173:??, October 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2022:QAM

[YSK22] Ziming Yang, Yuxia Sun, and Qingxuan Kuang. Question answering model based on machine reading comprehension with knowledge enhancement and answer verification. Concurrency and Computation: Practice and Experience, 34(12): e5828:1–e5828:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2021:DQD

[YSL^{+}21a] Ying Yang, Xuanhua Shi, Wei Liu, Hai Jin, Yusheng Hua, and Yan Jiang. DDL-QoS: a dynamic I/O scheduling strategy of QoS for HPC applications. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yuan:2021:CNM

[YSL^{+}21b] Yun-Hao Yuan, Xiaobo Shen, Yun Li, Bin Li, Jianping Gou, Jipeng Qiang, Xinfeng Zhang, and Quan-Sen Sun. Composite nonlinear multiset canonical correlation analysis for multiview feature learning and recognition. Concurrency and Computation: Practice and Experience, 33(15):e5476:1–e5476:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2022:GRB

[YSLX22] Wenbin Yang, Suqin Sheng, Xiangfeng Luo, and Shaorong Xie. Geometric relation based point clouds classification and segmentation. Concurrency and Computation: Practice and Experience, 34(11):e6845:1–e6845:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yadav:2023:RAD

[YSP23] Uma Yadav, Ashish K. Sharma, and Dipti Patil. Review of automated depression detection: Social posts, audio and video, open challenges and future direction. Concurrency and Computation: Practice and Experience, 35(1):e7407:1– e7407:??, January 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

You:2021:KEE

[YSS^{+}21] Xindong You, Tian Sun, Dawei Sun, Xunyun Liu, Xueqiang Lv, and Rajkumar Buyya. K-ear: Extracting data access periodic characteristics for energy-aware data clustering and storing in cloud storage systems. Concurrency and Computation: Practice and Experience, 33(9):e6096:1–e6096:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yildiz:2021:IWE

[YT21] Beytullah Yildiz and Murat Tezgider. Improving word embedding quality with innovative automated approaches to hyperparameters. Concurrency and Computation: Practice and Experience, 33(18):e6091:1–e6091:??, September 25, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2020:TSM

[YTN^{+}20] Lei Yu, Fei Teng, Shangming Ning, Yunshu Li, Zhe Cui, and Shengdong Du. A two steps method of resources utilization predication for large Hadoop data center. Concurrency and Computation: Practice and Experience, 32(15):e5634:1– e5634:??, August 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2024:RTP

[YTW^{+}24] Baoyong Yan, Jialin Tian, Jun Wan, Yu Qiu, and Weiming Chen. Real-time prediction of horizontal drilling pressure based on convolutional transformer. Concurrency and Computation: Practice and Experience, 36(10):e8006:1–e8006:??, May 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2020:CCC

[Yu20] Linchen Yu. CCHybrid: CPU co-scheduling in virtualization environment. Concurrency and Computation: Practice and Experience, 32(3):e4213:1–e4213:??, February 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2021:MFA

[Yu21] Gan Yu. A modified firefly algorithm based on neighborhood search. Concurrency and Computation: Practice and Experience, 33(6):e6066:1–e6066:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yucelbas:2022:PDP

[Yüc22] Cüneyt Yücelbas. Pre-determination of power density and application time in laser applications using PSONN hybrid algorithm. Concurrency and Computation: Practice and Experience, 34(4):e6611:1–e6611:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yadav:2023:EFB

[YV23] Pratibha Yadav and Deo Prakash Vidyarthi. An efficient fuzzy-based task offloading in edge–fog–cloud architecture. Concurrency and Computation: Practice and Experience, 35(26):e7843:1–e7843:??, November 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2022:MCI

[YW22] Jie Yang and Hong Wang. Medical concept integrated residual short-long temporal convolutional networks for predicting clinical events. Concurrency and Computation: Practice and Experience, 34(21):e7055:1–e7055:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2023:PGM

[YWC^{+}23] Siyang Yu, Fan Wu, Baoding Chen, Ronghui Cao, Zhibang Yang, and Keqin Li. A parallel game model-based intrusion response system for cross-layer security in industrial internet of things. Concurrency and Computation: Practice and Experience, 35(28):e7826:1–e7826:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2021:DBE

[YWQ^{+}21] Shengying Yang, Jianfeng Wu, Huibin Qin, Qiangqiang Xie, Zhiwang Xu, and Yongzhu Hua. Distributed buildings energy storage charging load forecasting method considering parallel deep learning model. Concurrency and Computation: Practice and Experience, 33(12):e5580:1–e5580:??, June 25, 2021.

CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Xiao:2021:RSP

[yXlLyGX21] Shi yang Xiao, Cai lin Li, Bao yun Guo, and Han Xiao. A radix sorting parallel algorithm suitable for graphic processing unit computing. Concurrency and Computation: Practice and Experience, 33(6):e5818:1–e5818:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yao:2020:TMF

[YY20a] Zhen Yao and Zheng Yan. A trust management framework for software-defined network applications. Concurrency and Computation: Practice and Experience, 32(16):e4518:1– e4518:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yi:2020:SLN

[YY20b] ZhengMing Yi and YiPing Yao. A scalable lock on NUMA multicore. Concurrency and Computation: Practice and Experience, 32(24):e5964:1–e5964:??, December 25, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yucelbas:2022:ASD

[YY22] Sule Yücelbas and Cüneyt Yücelbas. Autism spectrum disorder detection using sequential minimal optimization-support vector machine hybrid classifier according to history of jaundice and family autism in children. Concurrency and Computation: Practice and Experience, 34(1):e6498:1–e6498:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2022:AEC

[YYLL22] Qin Yang, Shihao Yin, Qingpeng Li, and Yongping Li. Analysis of electricity consumption behaviors based on principal component analysis and density peak clustering. Concurrency and Computation: Practice and Experience, 34(21):e7126:1– e7126:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2024:TAG

[YYM^{+}24] Yan Yan, Pengbin Yan, Adnan Mahmood, Fei Xu, and Quan Z. Sheng. Towards achieving geo-indistinguishability for 3D GPS location: a 3D Laplace mechanism approach. Concurrency and Computation: Practice and Experience, 36(14): e8111:1–e8111:??, June 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2022:NSM

[YYPR22] Ruoqian Yang, Chao Yang, Xingfang Peng, and Amin Rezaeipanah. A novel similarity measure of link prediction in multi-layer social networks based on reliable paths. Concurrency and Computation: Practice and Experience, 34(10): e6829:1–e6829:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2023:EIE

[YYY^{+}23] Feng Yin, Rongjun Yang, Hongxin Yu, Wei Zhou, Yuanjun Zhao, and Shuai Zhang. Edge intelligence-enabled supply chain financial model based on Business-to-Business ebusiness platforms. Concurrency and Computation: Practice and Experience, 35(13):e6353:1–e6353:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2022:DTN

[YYZS22] Yong-Li Yang, Yu-Guang Yang, Yi-Hua Zhou, and Wei-Min Shi. Discrimination of two nonorthogonal states against amplitude-damping noise via feed-forward control. Concurrency and Computation: Practice and Experience, 34(10): e6810:1–e6810:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yu:2021:MOE

[YZ21] Weiwei Yu and Li Zhang. Many-objective evolutionary computation based on adaptive hypersphere dynamic angle vector dominance. Concurrency and Computation: Practice and Experience, 33(13):e6238:1–e6238:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yin:2021:ESH

[YZPL21] Lujia Yin, Yiming Zhang, Yuxing Peng, and Dongsheng Li. Elastic scheduler: Heterogeneous and dynamic deep learning in the cloud. Concurrency and Computation: Practice and Experience, 33(13):e6206:1–e6206:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yang:2022:EMW

[YZX^{+}22] Fan Yang, Chaoqun Zhu, Heng Xu, Yongfeng Qian, and Jun Song. An efficient malicious webpage static detection framework based on optimized Bayesian and hybrid machine learning. Concurrency and Computation: Practice and Experience, 34(10):e6792:1–e6792:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yuan:2022:EDP

[YZXL22] Yilin Yuan, Jianbiao Zhang, Wanshan Xu, and Zheng Li. Enable data privacy, dynamics, and batch in public auditing scheme for cloud storage system. Concurrency and Computation: Practice and Experience, 34(10):e6735:1–e6735:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yan:2021:TDO

[YZYT21] Zheping Yan, Jinzhong Zhang, Zewen Yang, and Jialing Tang. Two-dimensional optimal path planning for autonomous underwater vehicle using a whale optimization algorithm. Concurrency and Computation: Practice and Experience, 33(9): e6140:1–e6140:??, May 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2021:RPA

[yZyWD^{+}21] Zhi yun Zheng, Chen yu Wang, Yang Ding, Lun Li, and Dun Li. Research on partitioning algorithm based on RDF graph. Concurrency and Computation: Practice and Experience, 33 (8):e5612:1–e5612:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yuan:2023:PPM

[YZZA23] Liujie Yuan, Shaobo Zhang, Gengming Zhu, and Karim Alinani. Privacy-preserving mechanism for mixed data clustering with local differential privacy. Concurrency and Computation:

Practice and Experience, 35(19):e6503:1–e6503:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zerga:2022:DDT

[ZAB22] Hideyat Zerga, Asma Amraoui, and Badr Benmammar. Distributed, dynamic and trustworthy access control for telehealth systems. Concurrency and Computation: Practice and Experience, 34(28):e7352:1–e7352:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zamani:2020:SSB

[ZABT^{+}20] Ali Reza Zamani, Moustafa AbdelBaky, Daniel Balouek-Thomert, J. J. Villalobos, Ivan Rodero, and Manish Parashar. Submarine: a subscription-based data streaming framework for integrating large facilities and advanced cyberinfrastructure. Concurrency and Computation: Practice and Experience, 32(16):e5256:1–e5256:??, August 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2023:EIE

[ZALM23] Rongbo Zhu, Ashiq Anjum, Hongxiang Li, and Maode Ma. Edge intelligence-enabled cyber-physical systems. Concurrency and Computation: Practice and Experience, 35(13): e7500:1–e7500:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zaid:2023:BBI

[ZARR23] Muhammad Zaid, Muhammad Waheed Akram, Amna Rizvi, and Syed Khurram Jah Rizvi. Blockchain based integrity assurance framework for COVID-19 information management & decision making at National Command Operation Center, Pakistan. Concurrency and Computation: Practice and Experience, 35(8):e7632:1–e7632:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:DLM

[ZBC^{+}21] Qingchen Zhang, Changchuan Bai, Zhikui Chen, Peng Li, Hang Yu, Shuo Wang, and He Gao. Deep learning models for diagnosing spleen and stomach diseases in smart Chinese medicine with cloud computing. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zaman:2022:RRT

[ZBY22] Tolga Zaman, Hasan Bulut, and Subhash Kumar Yadav. Robust ratio-type estimators for finite population mean in simple random sampling: a simulation study. Concurrency and Computation: Practice and Experience, 34(25):e7273:1–e7273:??, November 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2023:HEP

[ZBY^{+}23] Jisheng Zhao, Colleen Bertoni, Jeffrey Young, Kevin Harms, Vivek Sarkar, and Brice Videau. HIPLZ: Enabling performance portability for exascale systems. Concurrency and Computation: Practice and Experience, 35(25):e7866:1– e7866:??, November 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhai:2020:DLB

[ZBZ^{+}20] Keke Zhai, Tania Banerjee, David Zwick, Jason Hackl, Rahul Koneru, and Sanjay Ranka. Dynamic load balancing for a mesh-based scientific application. Concurrency and Computation: Practice and Experience, 32(9):e5626:1–e5626:??, May 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zou:2022:ULP

[ZCD^{+}22] Yanliang Zou, Chen Chen, Tongliang Deng, Jian Zhang, Xiaomin Zhu, Si Chen, and Shu Yin. User-level parallel file system: Case studies and performance optimizations. Concurrency and Computation: Practice and Experience, 34(13): e6905:1–e6905:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2023:STL

[ZCH^{+}23] Jia Zhao, Pengyu Cheng, Jiazhen Hou, Tanghuai Fan, and Longzhe Han. Short-term load forecasting of multi-scale recurrent neural networks based on residual structure. Concurrency and Computation: Practice and Experience, 35(5): e7551:1–e7551:??, February 28, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2021:CAB

[ZCL21] Jing Zhou, Ke Chen, and Jinsheng Liu. A clustering algorithm based on the weighted entropy of conditional attributes for mixed data. Concurrency and Computation: Practice and Experience, 33(17):e6293:1–e6293:??, September 10, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:DRT

[ZCN22] Shiqiang Zhang, Dongzhi Cao, and Zhenhu Ning. A decentralized and reliable trust measurement for edge computing enabled Internet of Things. Concurrency and Computation: Practice and Experience, 34(24):e7238:1–e7238:??, November 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zou:2023:OHR

[ZCR23] Ting Zou, Zhihui Chen, and Saeid Razmjooy. Optimized hierarchical radial basis function neural networks by developing coronavirus herd immunity optimizer for solid oxide fuel cells. Concurrency and Computation: Practice and Experience, 35 (22):e7730:1–e7730:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zou:2023:SOA

[ZCW^{+}23] Le Zou, Yan-Ping Chen, Zhi-Ze Wu, Qian-Jing Huang, and Xiao-Feng Wang. A sweeping optimization algorithm for the global cosine fitting energy image segmentation model. Concurrency and Computation: Practice and Experience, 35(17): e6651:1–e6651:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2021:SNR

[ZCX^{+}21] Yun Zhao, Jiagui Chen, Xing Xu, Jingsheng Lei, and Wujie Zhou. SEV-Net: Residual network embedded with attention mechanism for plant disease severity detection. Concurrency and Computation: Practice and Experience, 33(10):e6161:1– e6161:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2022:DMP

[ZCZ^{+}22] Ying Zhao, Junjun Chen, Jiale Zhang, Di Wu, Michael Blumenstein, and Shui Yu. Detecting and mitigating poisoning attacks in federated learning using generative adversarial networks. Concurrency and Computation: Practice and Experience, 34(7):e5906:1–e5906:??, March 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2023:CMP

[ZCZC23] Rui Zhang, Yihao Cao, Weiquan Zhang, and Xingjuan Cai. Cross-modal person re-identification based on deep attention hash learning. Concurrency and Computation: Practice and Experience, 35(28):e7863:1–e7863:??, December 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:RCP

[ZCZW22] Chuyue Zhang, Manchun Cai, Xiaofan Zhao, and Dawei Wang. Research on case preprocessing based on deep learning. Concurrency and Computation: Practice and Experience, 34 (2):e6214:1–e6214:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zerrouki:2022:ELD

[ZDH^{+}22] Nabil Zerrouki, Abdelkader Dairi, Fouzi Harrou, Yacine Zerrouki, and Ying Sun. Efficient land desertification detection using a deep learning-driven generative adversarial network approach: a case study. Concurrency and Computation: Practice and Experience, 34(4):e6604:1–e6604:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2021:MCA

[ZDJ^{+}21] Jiahui Zhu, Qingyu Dou, Lihua Jian, Kai Liu, Farhan Hussain, and Xiaomin Yang. Multiscale channel attention network for infrared and visible image fusion. Concurrency and Computation: Practice and Experience, 33(22):e6155:1– e6155:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2022:PSA

[ZDL^{+}22] Xin Zhou, Yong Dou, Rongchun Li, Peng Zhang, and Yuntao Liu. A pipelining strategy for accelerating convolution neural networks on ARM CPUs. Concurrency and Computation: Practice and Experience, 34(2):e6102:1–e6102:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:CPT

[ZFF^{+}21] Liang Zhang, Ligang Fang, Zheng Fan, Wei Li, and Jing An. Conditional pre-trained attention based Chinese question generation. Concurrency and Computation: Practice and Experience, 33(20):e6374:1–e6374:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2023:NMA

[ZFH^{+}23] Huanlong Zhang, Yuxing Feng, Wanwei Huang, Jie Zhang, and Jianwei Zhang. A novel mutual aid Salp Swarm Algorithm for global optimization. Concurrency and Computation: Practice and Experience, 35(17):e6556:1–e6556:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2023:PDB

[ZFW23] Jun Zhou, Dan Feng, and Fang Wang. Probability distribution based resource management for multitenant cloud clusters. Concurrency and Computation: Practice and Experience, 35(21):e6360:1–e6360:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2020:REP

[ZFZ^{+}20] Changen Zhou, Haoyi Fan, Wen Zhao, Hongben Xu, Huangwei Lei, Zhaoyang Yang, Zuoyong Li, and Candong Li. Reconstruction enhanced probabilistic model for semisupervised tongue image segmentation. Concurrency and Computation: Practice and Experience, 32(22):e5844:1–e5844:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2023:MFA

[ZGG^{+}23] Yuerong Zhao, Hongbo Guo, Ling Gao, Hai Wang, Jie Zheng, Kan Zhang, and Yong Zheng. Multifeature fusion action recognition based on key frames. Concurrency and Computation: Practice and Experience, 35(21):e6137:1–e6137:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zou:2022:LTA

[ZGH^{+}22] Wenjie Zou, Longkun Guo, Peihuang Huang, Geng Lin, and Hengquan Mei. Linear time algorithm for computing minmax movement of sink-based mobile sensors for line barrier coverage. Concurrency and Computation: Practice and Experience, 34(2):e6175:1–e6175:??, January 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:SSM

[ZGLS21] Conghui Zhang, Shaopeng Guan, Yi Li, and Wenwen Sun. SMOSA: Spider monkey optimization-based scheduling algorithm for heterogeneous Hadoop. Concurrency and Computation: Practice and Experience, 33(20):e6368:1–e6368:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2023:OPS

[ZGWZ23] Zhenning Zhang, Longkun Guo, Linyang Wang, and Juan Zou. One-pass streaming algorithm for monotone lattice submodular maximization subject to a cardinality constraint. Concurrency and Computation: Practice and Experience, 35 (17):e6645:1–e6645:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2022:SCA

[Zha22] Jinfang Zhao. Study on the cost allocation method to maximize the benefit of e-commerce enterprises in the cloud computing environment. Concurrency and Computation: Practice and Experience, 34(13):e5889:1–e5889:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2024:ETS

[ZHDW24] Gang Zhao, Hui He, Bingbing Di, and Tong Wu. ERSChain: Towards secure and flexible educational resource sharing using consortium blockchain and revocable ciphertext-policy attribute-based encryption. Concurrency and Computation: Practice and Experience, 36(5):e7941:1–e7941:??, February 29, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2020:TDS

[ZHJ20] Maosheng Zhang, Ruimin Hu, and Lin Jiang. Threedimensional sound reproduction in vehicle based on data mining technique. Concurrency and Computation: Practice and Experience, 32(13):e5912:1–e5912:??, July 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:ICC

[ZHJW21] Maosheng Zhang, Ruimin Hu, Lin Jiang, and Xiaochen Wang. Intelligent cloud computing platform for three-dimensional sound reproduction. Concurrency and Computation: Practice and Experience, 33(11):e6021:1–e6021:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2023:NNF

[ZHT^{+}23] Huihui Zhu, Pengyun Hu, Xianpiao Tang, Daoxun Xia, and Haifeng Huang. NAGNet: a novel framework for real-time students’ sentiment analysis in the wisdom classroom. Concurrency and Computation: Practice and Experience, 35(22): e7727:1–e7727:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2021:HIM

[Zhu21] Zhuanghua Zhu. A hybrid indicator many-objective optimization algorithm for the selection and delivery of disaster relief materials problem. Concurrency and Computation: Practice and Experience, 33(6):e5948:1–e5948:??, March 25, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2022:RMS

[Zhu22a] Wenqiang Zhu. Research on multi-source mobile commerce service recommendation model of data fusion based on tree network. Concurrency and Computation: Practice and Experience, 34(13):e5862:1–e5862:??, June 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2022:MOE

[Zhu22b] Zhuanghua Zhu. Many-objective evolutionary algorithm assisted by a novel angle-based fitness strategy. Concurrency and Computation: Practice and Experience, 34(26):e7301:1– e7301:??, November 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2020:PRM

[ZHW^{+}20] Weijun Zhu, Yingjie Han, Huanmei Wu, Yang Liu, Xiaofei Nan, and Qinglei Zhou. Predicting the results of molecular specific hybridization using boosted tree algorithm. Concurrency and Computation: Practice and Experience, 32(1): e4982:1–e4982:??, January 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2022:SGP

[ZHWY22] Jianyong Zhu, Chunming Hu, Tianyu Wo, and Xiaoqiang Yu. ScaleReactor: a graceful performance isolation agent with interference detection and investigation for container-based scale-out workloads. Concurrency and Computation: Practice and Experience, 34(4):e6666:1–e6666:??, February 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:TME

[ZHX^{+}21] Shi Zhang, Jin Huang, Ruliang Xiao, Xin Du, Ping Gong, and Xinhong Lin. Toward more efficient locality-sensitive hashing via constructing novel hash function cluster. Concurrency and Computation: Practice and Experience, 33(20):e6355:1– e6355:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zeng:2023:ADH

[ZHX^{+}23] Zhixia Zeng, Ruohe Huang, Ruliang Xiao, Xinhong Lin, and Shi Zhang. Anomaly detection for high-dimensional dynamic data stream using stacked habituation autoencoder and union kernel density estimator. Concurrency and Computation: Practice and Experience, 35(22):e7718:1–e7718:??, October 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2023:DDL

[ZHXY23] Linghe Zhu, Haibo Hong, Mande Xie, and Jun Yu. DLPM: a dynamic location protection mechanism supporting continuous queries. Concurrency and Computation: Practice and Experience, 35(19):e7495:1–e7495:??, August 30, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zarzour:2021:CFR

[ZJ21] Hafed Zarzour and Yaser Jararweh. A conflict-free replicated data type for collaborative annotation systems. Concurrency and Computation: Practice and Experience, 33(2):e5670:1– e5670:??, January 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhongsheng:2020:TIT

[ZJSJ20] Wang Zhongsheng, Wang Jianguo, Yang Sen, and Gao Jiaqiong. Traffic identification and traffic analysis based on support vector machine. Concurrency and Computation: Practice and Experience, 32(2):e5292:1–e5292:??, January 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See retraction notice [ZJSJ21].

Zhongsheng:2021:RTI

[ZJSJ21] Wang Zhongsheng, Wang Jianguo, Yang Sen, and Gao Jiaqiong. Retracted: Traffic identification and traffic analysis based on support vector machine. Concurrency and Computation: Practice and Experience, 33(17):e6314:1–e6314:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). See [ZJSJ20].

Zhang:2024:STP

[ZK24] Man Zhang and Zheng Kou. Spatio-temporal pattern evolution and driving factors of A-level logistics enterprises in Yangtze River Delta. Concurrency and Computation: Practice and Experience, 36(7):e7980:1–e7980:??, March 25, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2023:PND

[ZKZ^{+}23] Mengyi Zhang, Zhaokai Kong, Wenjun Zhu, Fei Yan, and Chao Xie. Pulmonary nodule detection based on 3D feature pyramid network with incorporated squeeze-and-excitation-attention mechanism. Concurrency and Computation: Practice and Experience, 35(16):e6237:1–e6237:??, July 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhong:2022:SBC

[ZLC^{+}22] Han Zhong, Zheng Li, Peng Chen, Hao Lu, and Yijia Xu. The selection of burglary cases based on multidimensional features and PageRank. Concurrency and Computation: Practice and Experience, 34(10):e6723:1–e6723:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:GGE

[ZLCL21] Chaoli Zhang, Dazhen Lin, Donglin Cao, and Shaozi Li. Grammar guided embedding based Chinese long text sentiment classification. Concurrency and Computation: Practice and Experience, 33(21):e6439:1–e6439:??, November 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:WSC

[ZLCS21] Xiangping Zhang, Jianxun Liu, Buqing Cao, and Min Shi. Web service classification based on information gain theory and bidirectional long short-term memory with attention mechanism. Concurrency and Computation: Practice and Experience, 33(13):e6202:1–e6202:??, July 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:CBA

[ZLD22] Hao Lan Zhang, Jiming Liu, and Margaret G. Dowens. Complex brain activity analysis and recognition based on multiagent methods. Concurrency and Computation: Practice and Experience, 34(8):e5855:1–e5855:??, April 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:DCR

[ZLL^{+}22] Xiaowei Zhang, Luming Li, Hong Liu, Peng Yang, and Yuanyuan Gao. Disentangling classification and regression in Siamese-based network for visual tracking. Concurrency and Computation: Practice and Experience, 34(27):e7246:1– e7246:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:FDM

[ZLM22] Jun Zhang, Houda Labiod, and Weizhi Meng. For a few diversities more: a voting-attack-immune voting scheme for social question and answer websites. Concurrency and Computation: Practice and Experience, 34(16):e5791:1–e5791:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:IPV

[ZLO^{+}21] Jindan Zhang, Chuan Lin, Urszula Ogiela, Nadia Nedjah, Arun K. Sangaiah, and Xuan Wang. Improved publicly verifiable auditing protocol for cloud storage. Concurrency and Computation: Practice and Experience, 33(23):e6049:1– e6049:??, December 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:UIE

[ZLT21] Tingting Zhang, Yujie Li, and Shinya Takahashi. Underwater image enhancement using improved generative adversarial network. Concurrency and Computation: Practice and Experience, 33(22):e5841:1–e5841:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:VNS

[ZLTX21] Yujian Zhang, Chuanyou Li, Fei Tong, and Yuwei Xu. A variable neighborhood search algorithm for energy conscious task scheduling in heterogeneous computing systems. Concurrency and Computation: Practice and Experience, 33(24): e6456:1–e6456:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2022:FEF

[ZLV22] Yousheng Zhou, Kexin Liu, and Pandi Vijayakumar. FTPS: Efficient fault-tolerant dynamic phrase search over outsourced encrypted data with forward and backward privacy. Concurrency and Computation: Practice and Experience, 34(28): e7360:1–e7360:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2020:OUT

[ZLW^{+}20] Jiale Zhou, Hong Li, Chiyu Wang, Xinrong Li, Jiawen Shi, and Zhicheng Pang. Optimizing unbalanced text classification tasks by integrating critical data mining and restricted rewriting techniques. Concurrency and Computation: Practice and Experience, 32(24):e5952:1–e5952:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:CCN

[ZLW^{+}21] Ruoyu Zhang, Wenpeng Lu, Shoujin Wang, Xueping Peng, Rui Yu, and Yuan Gao. Chinese clinical named entity recognition based on stacked neural network. Concurrency and Computation: Practice and Experience, 33(22):e5775:1–e5775:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2023:TMT

[ZLW^{+}23] Xiao Zheng, Houkai Liu, Xiujun Wang, Xuangou Wu, and Feng Yu. Timo: In-memory temporal query processing for big temporal data. Concurrency and Computation: Practice and Experience, 35(13):e6758:1–e6758:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2024:PAB

[ZLX^{+}24] Rui Zhu, Hang Liu, Xiaolong Xu, Leilei Lin, Yeting Chen, and Wenxin Li. A-PGRD: Attention-based automatic business process model generation from RPA process description. Concurrency and Computation: Practice and Experience, 36(8): e7940:1–e7940:??, April 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2021:LSI

[ZLYS21] Wang Zhou, Jianping Li, Yujun Yang, and Fadia Shah. Leverage side information for top-n recommendation with latent Gaussian process. Concurrency and Computation: Practice and Experience, 33(12):e5534:1–e5534:??, June 25, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:AAG

[ZLZ^{+}22a] Yakun Zhang, Weijun Li, Liping Zhang, Xin Ning, Linjun Sun, and Yaxuan Lu. AGCNN: Adaptive Gabor convolutional neural networks with receptive fields for vein biometric recognition. Concurrency and Computation: Practice and Experience, 34(12):e5697:1–e5697:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2022:PEE

[ZLZ^{+}22b] Yuan Zhao, Bo Liu, Tianqing Zhu, Ming Ding, and Wanlei Zhou. Private-encoder: Enforcing privacy in latent space for human face images. Concurrency and Computation: Practice and Experience, 34(3):e6548:1–e6548:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2023:DDB

[ZML^{+}23] Yuqing Zhang, Zhaofeng Ma, Shoushan Luo, Shushuang Wang, Pengfei Duan, and Xiao Liu. DBSDS: a dualblockchain security data sharing model with supervision and privacy-protection. Concurrency and Computation: Practice and Experience, 35(21):e7706:1–e7706:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2023:APN

[ZMLW23] Yukang Zhang, Zhanyou Ma, Yilin Liu, and Shunzhi Wang. Analysis of P2P network system based on two-stage service and failure repairability. Concurrency and Computation: Practice and Experience, 35(23):e7765:1–e7765:??, October 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zaki:2022:HPP

[ZNDA22] John Zaki, Anand Nayyar, Surjeet Dalal, and Zainab H. Ali. House price prediction using hedonic pricing model and machine learning techniques. Concurrency and Computation: Practice and Experience, 34(27):e7342:1–e7342:??, December 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhong:2022:MCC

[ZNLL22] Han Zhong, Zhenhu Ning, Guijun Li, and Zheng Li. A method of core concept extraction based on semantic-weight ranking. Concurrency and Computation: Practice and Experience, 34 (1):e6504:1–e6504:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2021:ETC

[ZOS^{+}21] Huan Zhou, Xue Ouyang, Jinshu Su, Cees de Laat, and Zhiming Zhao. Enforcing trustworthy cloud SLA with witnesses: a game theory-based model using smart contracts. Concurrency and Computation: Practice and Experience, 33(14):e5511:1– e5511:??, July 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhong:2021:QGB

[ZPL21] Sheng-Hua Zhong, Jianfeng Peng, and Peiqi Liu. Question generation based on chat-response conversion. Concurrency and Computation: Practice and Experience, 34(1):e5584:1– e5584:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:PSP

[ZQL^{+}21] Changjian Zhang, Deyu Qi, Wenlin Li, Wenhao Huang, and Xinyang Wang. SimpleSync: a parallel delta synchronization method based on Flink. Concurrency and Computation: Practice and Experience, 33(20):e6327:1–e6327:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:LSN

[ZQW^{+}21] Haoxu Zhang, Chenchen Qiu, Chao Wang, Bin Wei, Zhibin Yu, Haiyong Zheng, and Juan Li. Learning spectral normalized adversarial systems with stacked structure for highquality 3D object generation. Concurrency and Computation: Practice and Experience, 33(15):e5430:1–e5430:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2023:LLA

[ZQX^{+}23] Weidong Zhu, Shiyou Qian, Jiawei Xu, Guangtao Xue, Jian Cao, Yanmin Zhu, and Wenjuan Li. Lap: a latency-aware parallelism framework for content-based publish/subscribe systems. Concurrency and Computation: Practice and Experience, 35(17):e6640:1–e6640:??, August 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zenith:2024:EIS

[ZR24] N. Meena Zenith and R. Radhika. Examining the impact of social media advertising using ANOVA analysis. Concurrency and Computation: Practice and Experience, 36(5):e7939:1– e7939:??, February 29, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:SSF

[ZRY^{+}22] Yuhang Zhang, Hongshuai Ren, Wensi Yang, Yang Wang, Kejiang Ye, and Cheng-Zhong Xu. The strong substructure and feature attention mechanism for image semantic segmentation. Concurrency and Computation: Practice and Experience, 34(12):e5920:1–e5920:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zobaed:2022:PPC

[ZS22] Sm Zobaed and Mohsen Amini Salehi. Privacy-preserving clustering of unstructured big data for cloud-based enterprise search solutions. Concurrency and Computation: Practice and Experience, 34(22):e7160:1–e7160:??, October 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:SAD

[ZSC^{+}21] Zhenxing Zhang, Shiyan Sun, Xunyu Chen, Tian Zhi, Qi Guo, and Yunji Chen. Space-address decoupled scratchpad memory management for neural network accelerators. Concurrency and Computation: Practice and Experience, 33(6):e6046:1– e6046:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:SDF

[ZSQ22] Yi Zhang, Mengdi Sun, and Xin Qi. Speedup of discrete Fourier transform by efficient modular arithmetic. Concurrency and Computation: Practice and Experience, 34(3): e6564:1–e6564:??, February 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zaman:2022:NEE

[ZSS22] Tolga Zaman, Murat Sagir, and Mehmet Sahin. A new exponential estimators for analysis of COVID-19 risk. Concurrency and Computation: Practice and Experience, 34(10): e6806:1–e6806:??, May 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zoghi:2023:PEI

[ZST^{+}23] Seyedeh Masoumeh Mirsadeghpour Zoghi, Masoud Sanei, Ghasem Tohidi, Shokoofeh Banihashemi, and Navideh Modarresi. Performance evaluation of investment with DEA models under pure jump processes. Concurrency and Computation: Practice and Experience, 35(3):e7502:1–e7502:??, February 1, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:AHS

[ZSX21] Yi Zhang, Mengdi Sun, and Yong Xu. Application of hybrid swarming algorithm on flexible job shop scheduling problems. Concurrency and Computation: Practice and Experience, 33 (20):e6348:1–e6348:??, October 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:LPP

[ZSZ^{+}22] Lefeng Zhang, Guanghua Song, Danyang Zhu, Wei Ren, and Ping Xiong. Location privacy preservation through kernel transformation. Concurrency and Computation: Practice and Experience, 34(16):e6014:1–e6014:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2020:DPC

[ZTF^{+}20] Jia Zhao, Jingjing Tang, Tanghuai Fan, Chenming Li, and Lizhong Xu. Density peaks clustering based on circular partition and grid similarity. Concurrency and Computation: Practice and Experience, 32(7):e5567:1–e5567:??, April 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2021:ONK

[ZTL^{+}21] Junwu Zhu, Ling Teng, Huimin Lu, Jieke Shi, and Bin Li. Ontology negotiation: Knowledge interchange between distributed ontologies through agent negotiation. Concurrency and Computation: Practice and Experience, 33(15):e5406:1– e5406:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zeutouo:2021:FSA

[ZTM21] Jerry Lacmou Zeutouo, Vianney Kengne Tchendji, and Jean Frédéric Myoupo. A fast sequential algorithm for the matrix chain ordering problem. Concurrency and Computation: Practice and Experience, 33(24):e6445:1–e6445:??, December 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2022:MTT

[ZTMC22] Haoyi Zhao, Bo Tao, Ruyi Ma, and Baojia Chen. Manipulator trajectory tracking based on adaptive sliding mode control. Concurrency and Computation: Practice and Experience, 34(21):e7051:1–e7051:??, September 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2023:MTT

[ZTMC23] Haoyi Zhao, Bo Tao, Ruyi Ma, and Baojia Chen. Manipulator trajectory tracking based on adaptive fuzzy sliding mode control. Concurrency and Computation: Practice and Experience, 35(8):e7620:1–e7620:??, April 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2023:VVF

[ZTP^{+}23] Zhou Zhou, Youliang Tian, Changgen Peng, Nan Yang, and Shigong Long. VFLF: a verifiable federated learning framework against malicious aggregators in Industrial Internet of Things. Concurrency and Computation: Practice and Experience, 35(20):e7193:1–e7193:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2022:LSI

[ZTS^{+}22] Zichen Zhao, Xiliang Tong, Ying Sun, Dongxu Bai, Xin Liu, Guojun Zhao, Hanwen Fan, Jun Li, Cejing Zou, and Baojia Chen. Large scale instance segmentation of outdoor environment based on improved YOLACT. Concurrency and Computation: Practice and Experience, 34(28):e7370:1–e7370:??, December 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2022:CEI

[ZTY^{+}22] Bin Zhou, Shishi Tan, Ruirong Yu, Shenggang Wan, and Qiang Cao. Cost-effectively improving solid state drive lifetime by hierarchical redundancy and heterogeneous memories. Concurrency and Computation: Practice and Experience, 34 (12):e6044:1–e6044:??, May 30, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zulfiqar:2023:SSP

[ZUTK23] Haidar Zulfiqar, Hafiz Mahfooz Ul Haque, Faiza Tariq, and Rashad Mahmood Khan. A survey on smart parking systems in urban cities. Concurrency and Computation: Practice and Experience, 35(15):e6511:1–e6511:??, July 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2022:ASQ

[ZWC^{+}22] Xiaoyu Zhu, Jie Wu, Wei Chang, Md Zakirul Alam Bhuiyan, Kim-Kwang Raymond Choo, Fang Qi, Qin Liu, and Guojun Wang. On authenticated skyline query processing over road networks. Concurrency and Computation: Practice and Experience, 34(14):e5747:1–e5747:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2023:ESC

[ZWC^{+}23] Shaopeng Zhu, Bangxuan Wei, Chaoxin Chen, Jian Gao, Wenbo Ning, and Zhi Huang. Emergency steering collision avoidance control based on distributed driving intelligent vehicles. Concurrency and Computation: Practice and Experience, 35(2):e7486:1–e7486:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2023:MOO

[ZWCC23] Zhixia Zhang, Jie Wen, Xingjuan Cai, and Zhihua Cui. A many-objective optimization algorithm with dual criteria and mixed distribution correction strategy. Concurrency and Computation: Practice and Experience, 35(21):e7704:1– e7704:??, September 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2020:ABM

[ZWCS20] Yizhou Zhang, Yibao Wang, Tinggui Chen, and Jiawen Shi. Agent-based modeling approach for group polarization behavior considering conformity and network relationship strength. Concurrency and Computation: Practice and Experience, 32 (14):e5707:1–e5707:??, July 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2021:SAC

[ZWJG21] Mengyue Zhao, Jiaji Wu, Gwanggil Jeon, and Xing Guo. Statistical analysis of cloud characteristics in Northwest China based on Fengyun satellite data. Concurrency and Computation: Practice and Experience, 33(4):e4972:1–e4972:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2020:MAB

[ZWL^{+}20] Junxing Zhu, Xiang Wang, Qiang Liu, Xiaoyong Li, Chengcheng Shao, and Bin Zhou. A multiview approach based on naming behavioral modeling for aligning Chinese user accounts across multiple networks. Concurrency and Computation: Practice and Experience, 32(22):e5819:1–e5819:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2023:MCD

[ZWL^{+}23] Binquan Zhang, Di Wu, Zhuoxuan Lan, Zhihua Cui, and Liping Xie. Malicious code detection based on many-objective transfer model. Concurrency and Computation: Practice and Experience, 35(22):e7728:1–e7728:??, October 10, 2023. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2020:NPA

[ZWO^{+}20] Jindan Zhang, Baocang Wang, Marek R. Ogiela, Xu An Wang, and Arun Kumar Sangaiah. New public auditing protocol based on homomorphic tags for secure cloud storage. Concurrency and Computation: Practice and Experience, 32(18):e5600:1–e5600:??, September 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zheng:2022:CSE

[ZWT22] Lijuan Zheng, Zihan Wang, and Senping Tian. Comparative study on electrocardiogram encryption using elliptic curves cryptography and data encryption standard for applications in Internet of Medical Things. Concurrency and Computation: Practice and Experience, 34(9):e5776:1–e5776:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:DHD

[ZWW^{+}21] Guangxue Zhang, Tian Wang, Guojun Wang, Anfeng Liu, and Weijia Jia. Detection of hidden data attacks combined fog computing and trust evaluation method in sensor-cloud system. Concurrency and Computation: Practice and Experience, 33(7):1, April 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2021:IIS

[ZWX21] Bin Zhou, Shenggang Wan, and Changsheng Xie. Isolation: Inexpensively separating cold data via garbage collection to improve the lifetime and performance of NAND flash SSDs. Concurrency and Computation: Practice and Experience, 33 (15):e5460:1–e5460:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2024:MOO

[ZWY^{+}24] Haochen Zhao, Jie Wen, Jinqian Yang, Xingjuan Cai, and Chunxia Liu. A many-objective optimization-based local tensor factorization model for skin cancer detection. Concurrency and Computation: Practice and Experience, 36(6):e7947:1– e7947:??, March 10, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2021:WEN

[ZWZ^{+}21] Yichao Zhou, Wanyin Wu, Jie Zou, Jianwang Qiao, and Jun Cheng. Weighted ensemble networks for multiview based tiny object quality assessment. Concurrency and Computation: Practice and Experience, 33(6):e5995:1–e5995:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zuo:2022:EHR

[ZWZ^{+}22] Chunxue Zuo, Fang Wang, Mai Zheng, Yuchong Hu, and Dan Feng. Ensuring high reliability and performance with low space overhead for deduplicated and delta-compressed storage systems. Concurrency and Computation: Practice and Experience, 34(5):e6706:1–e6706:??, February 28, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2020:MOI

[ZX20] Zhixia Zhang and Liping Xie. A many-objective integrated evolutionary algorithm for feature selection in anomaly detection. Concurrency and Computation: Practice and Experience, 32(22):e5861:1–e5861:??, November 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:CIR

[ZX21a] Shu Zhang and Youshen Xia. CT image reconstruction algorithms: a comprehensive survey. Concurrency and Computation: Practice and Experience, 33(8):e5506:1–e5506:??, April 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2021:DTI

[ZX21b] Jiangwei Zhao and Chuanpei Xu. Design of time-interleaved data acquisition system based on Network on Chip. Concurrency and Computation: Practice and Experience, 33(10): e6180:1–e6180:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2023:SFB

[ZX23a] Jianquan Zhang and Xiao Xiao. Soft fusion-based cooperative spectrum sensing using particle swarm optimization for cognitive radio networks in cyber-physical systems. Concurrency and Computation: Practice and Experience, 35(13):e6295:1– e6295:??, June 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2023:SCS

[ZX23b] Jingbo Zhang and Liping Xie. Survey on compressed sensing reconstruction method for 3D data. Concurrency and Computation: Practice and Experience, 35(2):e7479:1–e7479:??, January 25, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2021:DTB

[ZXL^{+}21a] Baoxin Zhao, Cheng-Zhong Xu, Siyuan Liu, Juanjuan Zhao, and Li Li. Dynamic traffic bottlenecks identification based on congestion diffusion model by influence maximization in metro-city scales. Concurrency and Computation: Practice and Experience, 33(6):e5790:1–e5790:??, March 25, 2021. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2021:RIU

[ZXL^{+}21b] Taoyun Zhou, Minlei Xiao, Yangyang Liu, Yun Cheng, and Yi Liu. Research on indoor UWB positioning based on expectation maximization in NLOS environment. Concurrency and Computation: Practice and Experience, 33(17):e6278:1– e6278:??, September 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2021:VSR

[ZXLD21] Hu Zhu, Guoxia Xu, Lu Liu, and Lizhen Deng. Video smoke removal based on low-rank tensor completion via spatialtemporal continuity constraint. Concurrency and Computation: Practice and Experience, 34(1):e6169:1–e6169:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:FRA

[ZYCS22] Jiwei Zhang, Xiaodan Yan, Zelei Cheng, and Xueqi Shen. A face recognition algorithm based on feature fusion. Concurrency and Computation: Practice and Experience, 34(14): e5748:1–e5748:??, June 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2023:EEP

[ZYH^{+}23] Junjie Zhang, Qiao Yin, Weicheng Hu, Yunfeng Li, Hu Li, Nan Ye, and Bingyao Cao. EPA: The effective pipeline architecture for CNN accelerator with high performance and computing efficiency based on FPGA. Concurrency and Computation: Practice and Experience, 35(18):e6198:1–e6198:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2024:CDA

[ZYPD24] Jiange Zhang, Qing Yi, Christina Peterson, and Damian Dechev. Compiler-driven approach for automating nonblocking synchronization in concurrent data abstractions. Concurrency and Computation: Practice and Experience, 36(5): e7935:1–e7935:??, February 29, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:ISR

[ZYX^{+}21] Qiao Zhang, Xiaomin Yang, Long Xiao, Feng Yang, Farhan Hussain, and Pyoung Won Kim. Image super-resolution with parallel convolution attention network. Concurrency and Computation: Practice and Experience, 33(22):e6109:1– e6109:??, November 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhao:2023:PMS

[ZYXX23] Juanjuan Zhao, Jiexia Ye, Minxian Xu, and Chengzhong Xu. Practical model with strong interpretability and predictability: an explanatory model for individuals’ destination prediction considering personal and crowd travel behavior. Concurrency and Computation: Practice and Experience, 35(18): e6151:1–e6151:??, August 15, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2024:CCE

[ZYY^{+}24] Tong Zhou, Yaning Yu, Haonan Yuan, Bing Liu, Hongyang Zhao, and Ruijin Wang. CCSFLF: Cloud-edge-terminal collaborative self-adaptive federated learning framework. Concurrency and Computation: Practice and Experience, 36(12): e8042:1–e8042:??, May 30, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2023:PPD

[ZYZC23] Yuanyuan Zhang, Zuobin Ying, Bowen Zhao, and Chun Lung Philip Chen. A privacy preserving data aggregation and query for metro passenger flow via mobile crowdsensing. Concurrency and Computation: Practice and Experience, 35(20): e6965:1–e6965:??, September 10, 2023. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2024:FMD

[ZZC24] Zhigang Zhang, Zhixia Zhang, and Zhihua Cui. Federated malware detection based on many-objective optimization in cross-architectural IoT. Concurrency and Computation: Practice and Experience, 36(3):e7919:1–e7919:??, February 1, 2024. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:MSA

[ZZK^{+}22] Qikun Zhang, Kunyuan Zhao, Xiaohui Kuang, Yongjiao Li, Yuanpan Zheng, Junling Yuan, and Ruifang Wang. Multidomain security authentication for the Internet of Things. Concurrency and Computation: Practice and Experience, 34(16): e5777:1–e5777:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2020:OEP

[ZZL^{+}20] Jinghui Zhang, Jun Zhan, Jiange Li, Jiahui Jin, and Lei Qian. Optimizing execution for pipelined-based distributed deep learning in a heterogeneously networked GPU cluster. Concurrency and Computation: Practice and Experience, 32(23): e5923:1–e5923:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:MCA

[ZZL^{+}22] Shunli Zhang, Shang Zhang, Yuanzhen Liu, Yuhe Zhang, and Haibo Zhang. Multi-core accelerated simulation of Xray projection based on Unigraphics NX model. Concurrency and Computation: Practice and Experience, 34(11):e6846:1– e6846:??, May 15, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhong:2021:WDD

[ZZLZ21] Lianlian Zhong, Yongquan Zhou, Qifang Luo, and Keyu Zhong. Wind driven dragonfly algorithm for global optimization. Concurrency and Computation: Practice and Experience, 33(6):e6054:1–e6054:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:CDD

[ZZLZ22a] Tao Zhang, Tianqing Zhu, Renping Liu, and Wanlei Zhou. Correlated data in differential privacy: Definition and analysis. Concurrency and Computation: Practice and Experience, 34(16):e6015:1–e6015:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhou:2022:OHS

[ZZLZ22b] Guo Zhou, Ruxin Zhao, Qifang Luo, and Yongquan Zhou. Optimal hydropower station dispatch using quantum social spider optimization algorithm. Concurrency and Computation: Practice and Experience, 34(9):e5782:1–e5782:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:IHD

[ZZQ^{+}22] Yanan Zhang, Juanjuan Zhao, Yan Qiang, Xiaotang Yang, Wei Wu, and Liye Jia. Improved heterogeneous data fusion and multi-scale feature selection method for lung cancer subtype classification. Concurrency and Computation: Practice and Experience, 34(1):e6535:1–e6535:??, January 10, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:DLO

[ZZS21a] Meng Zhang, Yi Zhang, and Yuming Sun. Dictionary lookup with one genome evolution operation. Concurrency and Computation: Practice and Experience, 33(6):e5840:1–e5840:??, March 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2021:MPS

[ZZS^{+}21b] Shufang Zhu, Shangbo Zhou, Jiaxing Shang, Limin Wang, and Baohua Qiang. A multiion particle swarm optimization algorithm based on repellent and attraction forces. Concurrency and Computation: Practice and Experience, 33(5): e5979:1–e5979:??, March 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:MLL

[ZZW^{+}22] Yanan Zhang, Juanjuan Zhao, Wei Wu, Yan Qiang, and Liye Jia. Multi-level learning based on 3D CT image integrated medical clinic information for accurate diagnosis of pulmonary nodules. Concurrency and Computation: Practice and Experience, 34(17):e6998:1–e6998:??, August 1, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2021:FDR

[ZZWZ21] Suofei Zhang, Wei Zhao, Xiaofu Wu, and Quan Zhou. Fast dynamic routing based on weighted kernel density estimation. Concurrency and Computation: Practice and Experience, 33 (15):e5281:1–e5281:??, August 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:MPS

[ZZY22a] Jianpei Zhang, Xiaoxian Zhang, and Jing Yang. Multiobjective particle swarm community discovery arithmetic based on representation learning. Concurrency and Computation: Practice and Experience, 34(9):e5788:1–e5788:??, April 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhang:2022:MBS

[ZZY^{+}22b] Lina Zhang, Xiangqin Zheng, Keping Yu, Wenjuan Li, Tao Wang, Xuan Dang, and Bo Yang. Modular-based secret image sharing in Internet of Things: a global progressive-enabled approach. Concurrency and Computation: Practice and Experience, 34(16):e6000:1–e6000:??, July 25, 2022. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Zhu:2020:SSS

[ZZZ^{+}20] Haiting Zhu, Yuan Zhang, Lu Zhang, Gaofeng He, Linfeng Liu, and Ning Liu. SA Sketch: a self-adaption sketch framework for high-speed network. Concurrency and Computation: Practice and Experience, 32(23):e5891:1–e5891:??, December 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).